A Complete Bibliography of Publications in Educational Studies in Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

22 October 2025 Version 1.73

Title word cross-reference

(−8)^{1/3} [TE97]. −2 = (−8)^{1/3} = (−8)^{2/6} = [(−8)^{2}]^{1/6} = 2 [GR97]. $109.99US [Cor22]. $149.99US [Cor22]. 2 [MP96]. 2 × 2 [Roo71]. $33.00 [Mah21a]. $35.00 [Mah21a]. 3^{ème} [Une74]. $49.99 [Ahm23]. $59.99 [Ahm23]. $70.00 [Mah21a]. dy/dx [Thu72]. f′ [RNRVM23, RNRVM23]. i [Sch11b]. R^{3} [SP16]. Z [Rev76]. n [Bac71, LBML24, Sch11b]. c̸ [dV18]. S_{4} [Hah75]. Σ [Gro77]. ε [Roh10]. x [Mos98]. Z [ZG01].

-D [MP96]. -Gons [Bac71]. -intercept [Mos98]. -ratio [VS79]. -strip [Roh10].

//doi.org/10.1007/978 [Ahm23, NX23]. //doi.org/10.1007/978-3-030-55100-1 [NX23]. //doi.org/10.1007/978-3-031-19078-0 [Ahm23]. //doi.org/10.1017/CBO9781139045360 [Sfa15]. //www.springer.com/gp/book/9783030550998 [NX23].

//www.springer.com/gp/book/9783030689551 [Cor22].

0 [Ahm23]. 09 [SM22b].

1 [DWMB05b, EH05a]. 102 [PM19]. 10th [RR05]. 10th-Grade [RR05]. 11-to-13-year-old [Rou68]. 11-Year-Olds [FB02]. 12 [Rut05, WN04]. 13-to-15-year-olds [Dem97]. 13-year-old [Kie01]. 130.00 [Shv22]. 14 [Blu02]. 16 [BT05]. 17th [HbL05]. 1861 [Ste13]. 19 [AC21, Bor21, BBL22, DTD21, GG22a, HMSBT21, KM21, KDM21, KHY21, SRP21, XXL21, dSBM21]. 1960s [FF21]. 1969 [Ano70, Fre69a]. 1970 [Ano71c]. 1971 [Ano72b]. 1972 [Ano74c]. 1973 [Ano74d]. 1974 [Ano74b]. 1977 [Fre78b]. 19th [Dam87]. 1er [Cas69c, Cas69b].

2 [DWMB05a, EH05b, Pea76]. 2001 [Rut02]. 2003 [Ano04c]. 2005 [Ano06b]. 2008 [GM10]. 2015 [Ano16b, Ano16c]. 23rd [vBP19]. 25 [Kae94]. 260 [Kae94].

3- [HVN^{+}24]. 3-space [MRCR11]. 346 [Uhl03]. 3D [PC10]. 3H [TTS21]. 3rd [FSY22].

4-year-old [HVN^{+}24].

5-year-olds [Wri94]. 50.3 [Uhl03]. 5ème [Une73]. 5th [Une73, OK05, RRP91, Wal73b].

6 [Bal90, Egs70]. 6- [VVV22]. 6-year-olds [LH18, Wri94]. 6ème [Une73]. 6th [Une73, Wal73b, ZS07].

7-Year-Old [Eva07]. 7-year-olds [VVV22]. 7th [BCFF^{+}98].

8- [SVVO23]. 8-year [WC08]. 8th [AF05, RRP91].

9 [DŽC25]. 9-year-old [SVVO23]. 90th [Pai18]. 978 [Ahm23]. 978-0-521-76166-6 [Sfa15]. 978-1-107-68569-7 [Sfa15]. 978-1-139-04536-0 [Sfa15]. 978-1-4758-5479-4 [Mah21a]. 978-1-4758-5480-0 [Mah21a]. 978-1-4758-5481-7 [Mah21a]. 978-3-030-55099-8 [NX23]. 978-3-030-55100-1 [NX23]. 978-3-030-55102-5 [NX23]. 978-3-030-68955-1 [Cor22]. 978-3-030-68956-8 [Cor22]. 978-3-031-19077-3 [Ahm23]. 978-3-031-19078-0 [Ahm23, Ahm23]. 978-3-031-19080-3 [Ahm23]. 978-90-04-45964-9 [Shv22]. 978-90-04-45965-6 [Shv22]. 978-90-04-45966-3 [Shv22].

être [TLM07]. a-didactical [MM25]. abacus [Eng75, Eng76]. Abbreviations [Nar00]. Abduction [Mey10, PR11]. abductive [KJ22]. Abilities [Eng97, AKL80, Bis80, SM72, SVVO23, VP13]. Ability [Cla07, MC01, Rut87, AG83, BCLH89, Che23, Dek87, EJ13, FQL^{+}17, HH78, Hod96, LC81, MCPGE24, PC10, PAR^{+}24, SVLO20, Tar09, VR90]. Aboriginal [Gra88]. above [HP03]. above-average [HP03]. Absolute [AI12, MH97, SBP11]. Abstract [Haz99, Arm69, Cob86, DL14, FC12, GE98, Hic24, SW14, Tic12, WR96]. Abstraction [Bre22, Haz99, HZ05, Kid08, MO06, ES96, GE98, MW00, NFFAP20, Sch16, WDvO19]. Abstractions [Nar00]. abused [TTL^{+}13]. Academic [CBSP12, RC04, AAB03, Cas04, Cob87, EPQH18, Eve99, Hah14, RC02, SPFGF14, Wak14]. accelerate [MB16]. acceptance [FTM81]. Accepting [ODV06]. access [SJS12]. accessibility [SHAC01]. accommodation [KRHD15]. accompagnant [Rob95]. accompanying [Rob95, Sin03]. According [LH97, SM22b, Tre87]. Accounting [Riv07, GB23]. Accounts [Llo06, LH22a]. accumulation [YS12]. Accuracy [McD89, RSL17, SGCA23]. Achievement [CRL06, Rob90, Ade98, Beg71, Cas04, Che88, Cra12, CLSM17, GRR23a, GMH23, Han89, IWF^{+}15, Kae93, Kae94, KAJHB10, KBK21, LC15, Lor82, Ma96, MK91, MMR21, NCMO15, QP22, SJ17, SC17, SM22b, Sou83, TA08, TF95, Tch11, WZC21, Zep81, ZKC18, ZC19]. achievements [Fre75, ZWC^{+}24]. achieving [Kar10, KB11, PFSS19]. Acknowledgements [Ano99a]. Acknowledgment [Fre78a]. ACODESA [HGM15, Zen18]. acquired [Dag96]. Acquiring [Wea90]. acquis [Dag96]. Acquisition [TDGV09, CLR16, Une74, Hee85, KT01, MB16, SL74, Ver79]. acquisitions [Bic82, Com80]. acronyms [ZR18]. across [DKL^{+}25, DGI23a, DŽC25, HMNWB20, JE14, LR20, PPSZ19, SVVO23]. act [MS99]. actie [Gre17]. Acting [Wat00, Kra93]. Action [SSLM^{+}05, BAZF04, BTvG^{+}17, Gre17, MM20, MCW25, OB11, RL00]. Actions [MMBD05, WC08, Dra14, EJ13, EE21, HMT13, MPdP17, SD95, SBDD24, WH24]. activating [Sáe09]. Active [Ant96, Ell13, JWFC^{+}25]. activité [Bal88, Her11, Kry88b, LS79]. Activities [MF97, APS22, AS11, Bon13, BB25, CEHM16, CSS^{+}14, DVF18, DWP21, Eng71, FP07, HS99, Kyn93, LBT22, MPS22, PK17, TTL^{+}17, VN19, ZC19, dPH13]. Activity [Ern06, Lab00, NBD^{+}04, Riv07, RR05, Sek21, APN01, Bal88, BWD10, Boa02, Bro91, Czo18, DT12, GABB25, GMHG09, Her11, Her02b, Hoy01, JP09, JS01, KM18, Kry88b, LaC14, LOMG^{+}20, LS79, RR11, Riv10, Roh10, RL04, Sha77, SE93, Wil09, Woo16, ZBBZ85, ZM25, vdS11, PJP23]. activity- [ZM25]. activity-based [Sha77]. actor [Loc11]. actor-oriented [Loc11]. acts [Kem81, WJW11, dKBvdS22]. actual [KT15, LMW91]. Actualising [FS16]. actualizations [ML20]. actually [LM11, SRB16]. Adam [RC04]. adapted [HA23]. Adapting [Büs24, HMNWB20]. Adaptive [Kru13b, Noe80b]. add [DVV20, Dem93]. Adding [Han12a]. Addition [LT86, Sel01, AL00, Bar88, CHM83, Csí16, Jer70, KBBK20, KBK^{+}24b, NK77, NGR82, PvdHPR12, PDT^{+}12, SPNH12, Tho94a, VDTV24, VBE21, VM13, Ves69]. additional [Bar03]. additionner [Dem93]. Additive [WG07, GCALB15, Tho93]. Address [Fre78b, Ala12, EK10, Fre69a]. addressing [Cd04, Gra20, HHT23, TR24]. adhere [VEBS24]. Adidactic [SS05a]. adjectives [FI17]. adjustment [Bro12]. Adler [Ano16c, Wal17]. administration [Høy07]. admissions [HYC^{+}97]. Adolescent [BCLH89, KPS83]. adolescents [FNM91, O’B72]. adoption [CEMP04, TW09]. Adult [TE14, Bur87, LBT22]. adults [Adl88, Kar02, VVV13, VM78]. advance [BA25b, Pim95]. Advanced [MM03, DOV24, FCN14, GPPT99, IS08, IS09, OLW20, PKP^{+}18, RT14, Rut89a]. advancements [Qua92]. advances [CL20, Men23]. advertising [ETC14]. Advice [Ano98a, Ano00a, Ano02a, Ano03a, Ano03b, Ano04a, Ano04b, Ano05, Rut98a, Rut98b, Rut99a, Rut99b, Rut99c]. Advise [Rut00a]. aesthetic [FF25]. aestheticizing [GPS24]. Affect [BTV19, CL20, DG06, Mal06, RW19, SBL22, ZBEH06, Ade90, GRR23b, HY22, Han19, HWJ20, Kle12, KSH25, Obe19, PB09, SG08, SAR19]. affected [VVV13]. Affecting [GD17, Del69a, FNM91, Gal86, LCD90, NHN03, Phe10, PSE87, SLS74]. Affective [GC00, GLH20, OPC^{+}98, SB20, WB12, WI01, DGI23b, FF25, For95, FM09, TL15, TF95]. affectivité [Nim77]. affectivity [Nim77]. affects [LS20, Mah21a, Mah21b]. Affine [Cas69b]. affines [Cas69b]. affordances [NCC22, PC23]. afforded [CEHM16]. Africa [Hyl24, Adl88, Ano16c, Ban11, Osi24, SA00, VAM21]. African [DBL22, HMG16, Hyl24, Nkh02, lRA16]. after [Fre69d, IE07, Kom17, Sfa13]. afterthoughts [Che90]. against [Atw24, Che23, WBM^{+}20]. age [Høy07, SGCA23, Tab22, Mal23a]. aged [Com77, Mas69, TJ94]. Agency [AT21, FF17, Nor15, WG07, BS15, CEHM16, LLL21, MYM12, NCC22, dFS12]. Agency-structure [AT21]. agenda [BSW16]. ages [Lev13, PSS^{+}19]. agree [EMRI22]. agricultural [dKV18]. Ahead [Kin74, Sil13]. Ahsbahs [Bak16, JV17]. Aid [Pas78, Ern85, Gan92, IIMR^{+}11, YC90]. Aide [Pas78, Bal80]. Aids [ACJO04]. Aimed [DEM04]. aims [Fle71]. Akihori [Han98]. Alan [GB19, Ano16b, Ano24a, Mur24]. Alangui [RC04]. aléatoire [LD88]. Alexander [Law16, Men20, Men23, Law24]. Alexandre [Sri18]. Alfred [SE93]. Algebra [Haz99, KB07, Lau74, MSdSGG16, SS05a, Uhl02, Uhl03, AZ20, AAM21, BS12, Bur02a, DRR02, EG23, FS16, FF21, Fre74b, FC12, Gal02, GKMJ24, GMH23, GE98, GB14, HL94, Hic24, JC22, KK12, Lev71, LH96, MMM25, MYZ23, NFFAP20, Nos86, PM16, Pir25, PBCPP25, RL00, Roo71, RS22, SL94, SM22b, SK69, SLR^{+}24, TAH08, TT91, Tic12, Van03a, VH16, WSR11, Yer00, Zah15, dAS22, vdHPKR13, Jon16]. Algebraic [BEW99, Dem97, DVF18, ES21, IB05, Lep90, RP07, THS06, TER98, WFB13, CMdR18, CB89, CBC25, CCRD22, CPPC18, CPPC23, CPPT^{+}13, EEST20, ES96, EE21, GT00, GAP24, Hah75, Hew12, HSZ17, LL99, MS97, MvdKM13, MS09, MP14, PK17, Rad00, RFP14, RDH09, SB97, SD94, Sla98, SGCA23, SJ04, Sut89, TW09, Vol69, WH24, ZL02, ZM25]. algébrique [CB89]. algébriques [SB97]. algorithm [Hee78, LL23, MGD^{+}23, NM98, Son16]. algorithmes [AL00]. Algorithmic [Sle82, FKIA15, Leh24, NJL25]. algorithmics [LL23]. algorithms [AL00, FVJL^{+}19, Ste83, Tho94a]. Alienation [RR17, Wil16b]. Aligning [KMN24]. alignment [GFJ13]. Alison [Art15a, Tab22]. alive [Whi87]. Allen [Ric20]. Allocution [Fre69a]. Almost [Sfa13]. along [dF04]. Aloud [ZS07, SZP10]. Alphonse [Hyl24]. also [Chr97]. alternance [Hah99]. Alternate [Sla97]. alternating [Hah99]. Alternative [Lén93, Rut05, Wat05, HSZ17, MK97]. always [AVM15, Gla76b, PG93, vLXD18]. am [AVM15, NX23, SM21]. ambiguity [BKGZ24]. ambitious [GCFS23]. améliorer [Kra02]. amenable [NK17]. America [Ste13]. American [RSD23, Her02a, HC14, RSD23]. Among [CBF07, And89, BCFF^{+}98, BJJ13, CH13, Hah99, JTP94, Loc11, LG16, MLL20, SMFL19, TAH08, VMN11]. amongst [KL14]. amounts [Joh15]. Amy [MDU^{+}25]. Ana [Tab22]. Analogical [ES96, Hic24]. analogy [LS11]. Analyse [CGS76, Rob95, BZ01, Cli17, DS02, Rev69, TB09]. Analysed [Rut05, Ped07]. Analyses [PRR08, CW96, Sim15]. analysi [Wit80]. Analysing [Dah04, MMC22, GB23, Reu23]. Analysis [BH05, Bus05, CGS76, DWMB05b, DWMB05a, Duv06, FV13, FG17, GBR05, II05, KČ12, MPZ16, NH94, RR05, Ryv06, SG98a, AV22, And79, Bec77, BZ01, Ber08, Bra11, BC12, CSP24, DS02, Dow96, Ern16b, FD24, FE09, Fin22, FT00, GT95, Gra91, GMHG09, GM84a, GM84b, GB13, GB17, HEWC10, Ior17, Kap70, KB20, Kir69, Kle02, Kle12, KL14, MCVM13, MM25, Miy17, MRC18, NK77, QP22, Rad00, RCH25, Rev69, RC94, SP16, SV25, Sek21, Ste69b, TK14, TB22, TB09, Tru15, TP15, UB23, VDWV04, VS13, Vol69, Vol76, Vou16, VGJ19, WS25, Wit80, XC19, YG99, ZWMR16, lRA16, Bec77, KM25, Rob95, VVM22]. analytic [Ng21a, Pir25, ZCPP13]. analytic-structural [Pir25]. Analytical [Tay91, Vin97, DF15, LH09]. analyze [EG23, MWVE09, RNRVM23]. Analyzes [PRR08]. Analyzing [BCLH89, BHLH85, Cle80, FCN14, HRR23a, HHB21, RNP13, Ste05, dV18, BTvG^{+}17, Büs24, CHK22, FBA10, GL10, Han12d, LE09, MYM12, Rob12, TMH21, VRK24, HRR23b]. anatomy [Sim15]. ancien [AG02]. ancient [Siu93]. andererseits [Ott76]. Anderson [MDU^{+}25, MND23]. Andreas [MR20]. Angelika [JV17, Bak16]. Angle [CB00, SC93, GB01, KZ16, MW00, Moo13, SC90]. angled [MRSP14]. Angles [Tsa07, Fyh08]. Angolan [Ger88b]. Anjum [MR17]. Ann [Sch24, SL95]. Anna [HJ12, Ric20, Wal17, Ano24b, Now92, Tur88, Win11]. Anne [Ahm23]. anniversary [Bag08a]. announce [FB17]. Announcement [Ano75a, Ano81a, Ano82a, Ano82b, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano85a, Ano85b, Ano86a, Ano86b, Ano86c, Ano87a, Ano87b, Ano88a, Ano88b, Ano90a, Ano90b, Ano91a, Ano91b, Ano92a, Ano92b, Ano93a, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano00b, Ano04c, Her04, HbL05, Hud94, Ano81c, Ano84d, Ano85c, Ano88f, Ano88g]. Announcements [Ano77a, Ano78a, Ano80a, Ano80b, Ano84b, Ano87c, Ano92c, Ano19]. anomaly [AAJW25]. ans [AV76, Bic82, Bui76, Com77, De 94, Dem93, Mas69, Rou68]. answer [Bag73, OVKR25]. Answering [Sfa15, FWCW22]. Answers [AR07, RDH10, Woo16]. Anthony [LX16]. Anthropology [Rad06, Con90]. Anthropomorphizing [Sri18]. anti [ISJ24]. anti-Asian [ISJ24]. anticipations [Cob85]. anticipatory [Tzu07]. anwendungsorientierte [Bec77]. Anxiety [New98, PSS^{+}19, Wil97, Bek10, GS20, LC15, TR24]. any [O’C01]. Anyway [Wil97]. août [Fre69a]. Aotearoa [PAR^{+}24]. apart [LK18]. APOS [BMP20, BGT17, DWMB05b, DWMB05a, KB20, KV23, MTBR16, Moo20]. APOS-Based [DWMB05b]. apparemment [Dum82]. Apparent [BG69]. apparently [Dum82]. Application [Bal71, LH06, Bec77, CEMP04, HJ93b, TW09]. application-oriented [Bec77]. Applications [Blu02, Als02, BN91, Bur80a, Eng68, Pol68, Pol69, Wit73a, Wit73b]. Applied [BN91, Les81, Lyn68, Eng69, Gri68]. apply [Wat87]. Applying [Nun96, SG98b, Wan25, VVV15]. appraisals [FI17]. Appraising [HEW10]. appreciation [Ano98b]. Apprehending [SS06]. apprehension [MCPGE24]. apprentice [Hah99]. apprentis [Hah99]. Apprentissage [Cha71, Dag96, Art97, CLR16, Dié70, Duv91, Mot71, Sch92, Ser76]. apprentissages [Tro04]. Approach [Adl97, CV09, Com77, Ler01, MD01, Mes04, PM97, BBF71, BA14, BS12, Bec24, BJ88, BC78, BMPT24, BC12, Bur80b, BBFG99, Cha22b, Col21, DVV20, DMHRV21, DS21, EN79, EG23, FQL^{+}17, FS16, FF17, Fle68, FDW19, GMH23, HYC^{+}97, Har73, HOEF89, HTM^{+}23, Hoy01, IB19, Jun91, KS69, Kon18, KBK^{+}24b, bL00, Ler85, Lib77, MW09, MWVE09, MAPHB21, Nar16, Neu99, Par16, RR80, Rev76, RNRVM23, Rog71, Rut89a, Sha77, SK69, VRP^{+}12, VPR21, Vil71, Vin75, Vin77, Wit80, YG99, Yer00, vSBvR00, AP75, CGS76]. Approaches [Mor00, TER98, Cra96, EK10, GG22b, Gra91, HPS^{+}19, HPRW25, ICR20, KBK^{+}17b, KFS01, KID25, KR11, MCKP19, MCVM13, Moo20, El 18, OCB25, RR17, TRDA20, KBK^{+}17a, JV17]. Approaching [Bja20a, Bja20b, FLM07, UMC24]. Approche [AP75, CGS76, Com77, BC78, bL00, Rev76]. appropriate [GBMB70]. Appropriating [Car10, Mos04]. Appropriation [BC82, SHG21, BC82]. approval [XXL21]. après [Fre69d]. Aptitude [BHX97]. APU [FB02]. Archimedes [Fre77b]. arcs [Moo13]. Area [DVJ98, ATDR19, BN96, CF00a, DPG89, FKC17, Kor03, KSL11, Leh23, PH96, SS17, Yim10]. area-comparison [CF00a]. Areas [II05, ZM91]. Argentina [SS93b]. argument [Bar07, MS20, Sho21, YE16]. Argumentation [Dou99, Web18a, Web18b, AS11, CSS^{+}14, IMRS07, Kru13a, MP14, MPZ16, NBZ12, Ped07, Reu23, WB11, ZC22]. argumentative [AWE22, PHS12, TRDA20]. arguments [Jan74, MPZ16, MS20, MYM12, ZWMR16]. Arithmetic [GL98, Gri87, GB17, SS05a, VVM22, BFG84, BMK21, Cob87, Fin22, Fro99, Gal02, GFWL11, GE98, Gra91, Gri13, GS18, HL94, JR72, Jer73a, Jer73b, JC22, Kir69, LCD90, LW89, Lev81, Lie90, Lin90, LH96, MP14, Nes76, NK77, PCP19, REL25, SSPS20, SLS74, Sfa15, SG25, Sla98, Tre87, Van03a, VGJ19, WMN22, WMN25, vLXD18, SB97]. Arithmetical [BEW99, Lan86a, FV13, FVJL^{+}19, LHS19, MS09, Sem84, TW09, VVV13, Ver79, VR90, Lab78]. arithmetico [HSZ17]. arithmetics [Vin75]. arithmétiques [Lab78, SB97]. array [BHHS09]. arrows [Bja20a, Bja20b]. art [CKS22, Fri21, KC22, LT20]. artefacts [MB09, MK14, Ola14, Rad14]. Arthur [Fit11, Gre17]. article [Pea76, Ril86]. Articles [Dre07]. Articulating [SK13]. artifact [SAD21]. Artifacts [Bon13, SBDD24]. artithmetical [Wol83]. ascesis [BVR24]. Asia [DKL^{+}25]. Asian [SMWM25, YK20, DKL^{+}25, ISJ24, Leu01, Leu05, LYBK23]. AsianCrit [ISJ24]. asking [IAP19]. Aspect [BC05, CGD21, Fri11, Lab78]. Aspects [El 86, Gal81, New98, Nis99, Ste91, Tru15, AG83, Bel79, BMK21, CBC25, GDK23, HR74, HP12, Lor82, Sch83, TPHS07, Van93a, Van69b, Vol69, WS18]. aspiration [Kle12]. aspirations [BWHM^{+}10, WS18]. assemblage [FF17]. assemblages [ML20]. Assessing [BFA22, Coo98, DKBD19, HI94, II05, SSPS20, SD13, SVLO20, VSS24, Bec23, DŽC25, GC23, Hay87, JTP94, JLTM97, JI15, SGB21, DKBD21]. Assessment [AKL80, Ban11, Eva07, Osi24, Wat05, WI01, BL22, BH95, BLP10, BJ22, BAZF04, CB19, CWYP92, EN79, Ern84a, EJ23, FKC17, FGA94, Fre13, ICR20, KSSO25, KID25, Lor82, MRFW^{+}12, MO94, PY22, Rut94, SS15, SGCA18, TJ94, TPHS07, Tzu07, Van94, Wil71, Bui76]. assessments [MM11]. Assessors [Wat00]. assimilation [Une73]. assist [HA23]. assisted [GS72, GS76]. Assisting [BSF96]. Associated [Mes04, Abt22, Kie81, SM22b, WM01]. associating [Joh15]. association [MWSS17, VWV21]. Associations [HP25]. associativity [SL74]. assumptions [Kle12]. Astonishment [Rot17]. astray [BSV95]. Asymmetry [SJ17]. asynchronous [CBC25]. ATD [BGT17]. Attached [Dim04]. attainers [Bac89]. attaining [DB86a, DB86b, SHG21]. Attainment [FB02, APS22, Bla85]. Attempt [SSLM^{+}05, Van94]. attending [JWFC^{+}25, SMFL19, WFC23]. Attention [NN07, Bar03, HKSS^{+}25, NJL25, PPPP14, VGJ19, WHE09]. Attitude [Han02, RMA98, TA08]. Attitudes [GH98, MRC18, PC98, Wat05, bMYT98, CPP09, Che88, DCHM19, GC90b, Gan92, GRR23a, HWR23, KM93, LC13, LC15, Mal93, MY84, MCL25, MN95, PMG^{+}94, QP22, Rel96, SC17, Wit71]. attitudinal [Min83]. attribute [PM15]. Attributed [Spe05]. attributes [LH18]. attributions [CL84, FPCL90]. Atweh [DKL^{+}25]. audiences [NX23]. audit [KHY21]. Audrey [KT01]. August [Fre69a, Fre78b]. Auslaufvorgänge [Bec77]. Australia [Ano16b, Art23a, Bla78, CBF07, MNRR95, Wat05]. Australian [Cla07, SV09]. Auswahl [Beh68]. authentic [Her02b, KMA02, WHC09]. authenticity [Pal08]. Author [Sch25, Uhl03, GDS22, Bro23]. Authority [AF05, LOMG^{+}20, SHG21]. Authors [Ano98a, Ano00a, Ano00f, Ano00g, Ano01b, Ano01c, Ano01d, Ano01e, Ano04e, Ano05, Rut98a, Rut98b, Rut99a, Rut99b, Rut99c, Rut00a, SS05b, Ano02a, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04f, Ano04g]. autistic [PBCPP25]. auto [Plu79]. auto-contrôle [Plu79]. automata [Hol81]. automatic [Pas78]. automatique [Pas78]. Autonomous [FP85, Woo16]. Autonomy [Sko04, Høy07]. autour [Cas69c]. aux [AL00, AV76, FPM69, FBM71]. auxiliary [FQL^{+}17]. Average [Wat07, HP03, Mak14, NPH99]. avoidable [Hew12]. avoidance [MZ18]. avoiding [HP25]. awakening [Mes89]. Award [Ano14a]. Awards [Ano06b, Art05, Art02]. Awareness [TER98, Tsa07, Ior18, LBT22, Neu99, Shr10]. away [SPNH12, Vog70]. awe [Car08]. axiom [Bul24]. Axiomatic [Rog71, Sto71]. axioms [GS72, Zas71]. axis [dF04].

B [Fit11, Ott76, Sel00, Sfa15, VDWV04, Web18a, Web18b, Zaz99]. Babies [Mat03]. Baccaglini [Ric20]. bachelor [PT09]. back [HFL^{+}23, Sil13]. Background [HM95, Kau71, HY22, Rad21, Vil10]. backgrounds [CW14a]. Backstage [Sie04]. Bagni [D’A10]. Baker [Web18a, Web18b]. Bakhtin [Bar16]. Bakhtinian [MA25]. Bakker [Gai11, Gre17]. Balacheff [dV18]. balance [Ell18, Orl71, Vla02]. Balanced [Wal76]. Balancierte [Wal76]. ball [Lén93]. Bands [FB02]. Bangladesh [Sha78]. Baptiste [Sta11]. Barber [Fit11]. Barbin [Fri21, Rut02]. baroque [Dho93]. barrier [Bar21]. barriers [Mai85, SKK^{+}23, Woo16, YST22]. Bartolini [Mor00, vBP19]. Barton [GM10, RC04]. Baruch [Rad11, Web18b, Web18a]. barycenter [Cas69a]. barycentre [Cas69a]. Base [NHN03, Sto22]. Based [DWMB05b, Leu05, VL04, Adl88, BL22, BMP20, BR17, BTvG^{+}17, CO10, Chr00, CWH19, CBSS96, DMHRV21, DŽC25, FQL^{+}17, FC12, GBE13, GFJ13, GAP24, HTM^{+}23, JN15, Jaw03, JC22, KID25, KH07, KSS25, KC22, KN21, LE09, LLW07, LL13, Lev10, MCW25, MP24, PY22, PFSS19, Ron15, RPP24, RH02, Sch13, Sha77, SLAY19, SS17, SLJ69, Sut89, TW09, WBM^{+}14, WFCW15, Whi09, WvdHPD15, Win03, Yer00, dPH13]. basic [CBSP12, FPS16, Hee85, Hof77, Isa87, SM72, Zav21]. Basis [BH05, Fid85, Pre92a, Sca71]. Basque [VMN11]. Basque- [VMN11]. Bastiaan [Fre77b]. Bastide [Gre17]. Bastide-van [Gre17]. battle [NRV24]. Bayesian [BEBL^{+}24]. Be [Hof06, Ano74b, Bec19, BMK21, BHX97, Cob86, CH02, CM92, Egs70, Fre68, HA23, Kla68, MP24, Miy17, MRCR11, O’C01, Osl12, Ped07, PC23, ST08, Ser68, Sfa05, SRB16, TLM07]. bear [MD01]. Beats [Tsa07]. beautiful [Yam09]. beauty [Pre07]. bec [Tro00]. become [GBC12, Mur24, Sch11b, WMPL08]. Becoming [AM25, McQ86, BW13, NRV24, Shv22, SRB16, dF12]. Bednarz [Eng99]. before [BCZ21, Fri21, Led92, ZR18]. beginners [TAH08]. Beginning [WM98, BS12, De 93, GB14, Güç13, LLL21, PK23b, Rot21, Wat80]. beginning-level [Güç13]. begint [Gre17]. Begle [Kei82]. Begriffe [JSV75]. Behavior [Car99, Add72, FP85, KC73, TY04]. behavioral [AKL80]. behaviour [Cla85, IM19, PT09, Tho85]. Behind [MW19, LNC23, dSBM21]. beholder [AE08]. Beijing [Leu95]. Being [ZSL02, JGF20]. Beirut [JS99]. Belgium [Bja20b, Bja20a]. Belief [Gat06, Bro16, CV09]. Beliefs [AS05, Bes07, Boa99, Gat06, Spe05, AS04, Bes12, CW14a, CPK08, CPP09, CP10, Cob85, FPCL90, For95, Goo07, HPS^{+}19, Kun12, LBH23, LSC20, LCK24, MF17, MS14, MCL25, PT09, Rot21, Sho21, Sia24, TB22, TLB^{+}24, Tie00, TG89, TTL^{+}13, VBE21, WZC21, WM09, WL90]. believe [Por90]. Believing [Har09b]. bending [Rot14]. benefits [LOW24a, LOW24b]. Berinderjeet [Hus25, LX16, YK20]. Bernoulli [Gos21]. best [CN16]. Beth [Bag08a]. Better [Bac89, Rut94, Kra02]. Between [AS05, ASP97, De 93, HMG16, MA25, SA00, Tho85, WW07b, AS04, AS17, Arm69, AWL15, Bal88, BML21, BWHM^{+}10, BW13, BLP10, BGT17, BP15, BB03, CCP23, Del69b, DVF18, DCHM19, DG89, Ell18, EVHRC^{+}25, FG17, FB17, For95, Fre71, GB23, GRR23a, GMH23, Gos21, HFL^{+}23, HMRRTT23, Her11, HL94, HMMG18, HGM15, IWF^{+}15, JM18, KBK21, KJvD20, KLBMK^{+}13, Kru13a, LLW07, Lev22, LH96, LL99, LLZ20, Mah21a, Mah21b, MCVM13, Mar13, MP14, MY84, MB19, Mon00, MWVE09, MK14, Ng16, Obe19, OAS25, Pai16, Par25, PCP19, Ped07, RFK07, RO09, RL12, Rob90, SBDM69, SZ07, Tch11, Tru15, VP13, VWV21, Wak14, XM17, YSL22, ZL02, ZWMR16, Zen18, Zep81, dA00]. between-group [Tru15]. Beyond [AMMS20, BCE^{+}19, BT05, BE23, Car08, Fin22, GC90a, Gat06, Jou25, KB12, MS99, Sim96, WD05, ZC06, REA09, Son16, Whi87]. Bharath [Pai18]. bias [Par25, VVV15]. Bielefeld [Ano74b]. bifocal [Kle12]. big [AGP15]. Bigger [HS13b]. Bijections [MDD04]. Bikner [Bak16, JV17]. Bikner-Ahsbahs [Bak16]. Bilingual [Men07, Zep82, Cla92, Dom11, DLK14, FKC17, Ng16, Zah15]. Bilingualism [Daw83, Wha96]. Bilinguals [Cla07, Jur77, TDEM13]. Bill [GM10, DKL^{+}25]. Billiard [KK13]. binary [FYAE12, MEH20]. biography [Kaa07]. biology [RT14, VN19]. biopolitics [ZK21]. biosocial [Pai19]. Birgit [Cha22b]. birth [Bag08a]. birthday [Pai18]. bisection [LL23]. Bishop [Ano16b, Ano24a]. bit [GLHM24]. Black [Bro11, WW07a, Nkh02, MJS21, McV24, Ort24, VCD24]. blaming [Car08]. blended [Pir25]. blending [ATDR19]. blends [YTD11]. blind [PW22]. blindness [MMM25]. block [GB19]. blocks [Loc92, WH24]. Blurred [DVC19]. board [DH20]. boards [Gil71]. Bock [Bja20a, Bja20b, OM24]. bodies [LT23, MPS12, Ste76]. Bodily [NBD^{+}04, RT09, Pre92a]. Body [Fer15, NFFAP20, HSNHHE15, KM18, Pre92a, SAD21, dFS13, dF16]. body-artifact [SAD21]. Bohlmann [Sri18]. bond [Bos93]. Book [AS87, ABW88, Ahm23, All99, Ano89a, Ano98c, Ano98d, Ano00c, ACS88, Art15a, Art19, Atw24, Bak16, Bec86, Bel85, Bja20a, De 24, BFG83, Bro11, Bro23, Bur03, Cha22b, Cob04, Cor22, Cot24, Cra85, Dav07, Daw94, DKL^{+}25, Dre93, Edd18, Edw93, Ell18, Ely10, Eng99, Fer15, FH90, Fit07, Fit11, Fit15b, Fit17, FH88, FKSS82, FJA86, Fri21, FSY22, Fur19, Fur22, Gai11, GV83, Goo20, Gra20, GB88, GM10, GB19, Gue19, Han12a, Han95, Han98, Han07, HSNHHE15, HJ12, Hes13, HFD82, Hus25, Hyl24, IPJ25, JS87, JC22, Jon16, JV17, Kai13, Kap09a, LG87, LT20, Law16, Law24, Law00, LW83, LX16, LP91, Lun82, Mah21a, MPS12, MDU^{+}25, Mal23a, Mal23b, MND23, MR17, Men20, Men23, Mor00, MR20, MH89, Mur24, Nol11, NX23, ND88]. Book [Nun95, OM24, Ott76, Pai18, PD91, PB84, Pim91, Pim93, Pre07, Rad11, Ric20, Ric10, RSD23, Rut02, Rut10, SF84, Sch24, SM83, Sel00, Sfa15, SJS12, Shv22, Sil25, SMWM25, SBS86, Sri18, Sta11, Sto22, Tab22, Tep20, Tsa02, TR24, Wal10, Wal12, Wal17, Wan25, Web18a, Whe96, Win11, WB84, Wit82, WQTV85, Wub92, Yam09, YK20, ZB19, dF12, vBP19, BP22, Fit09, Fit15a, Bja20b, Mah21b, Web18b]. books [Ste13]. Bophuthatswana [MK91]. Borwein [KMT16]. both [LLHL13, Roo19, RR23]. boundaries [PPSZ19, WPA^{+}23]. boundary [BA14, Bak16, HS22]. boundary-crossing [BA14]. Bourdieu [JGR14, Nol12, Wil12]. Bourdieuian [Nol16]. Bourdieusian [QP22]. bourgeois [Zev96]. box [Gar24]. Boxes [WW07a]. boxplots [ALR25]. boy [KLBMK^{+}13]. boys [CWT12, FPCL90, Han89, KAJHB10]. Braces [Fre69b]. brackets [GSH16a, GSH16b, PG20]. Brantina [Mur24, Hyl24]. Bratislava [Ano72b]. Brazil [LT20, D’A91, Men07, dSBM21]. Brazilian [D’A91, dKV18]. Breaking [Sie04]. Brian [Fit11, GB19, JMD08]. bridge [Zen18]. Bridges [Eva99]. Bridging [JP09, ZWMR16, BE23, CHK22, Ern16b, TP22]. Brief [VVV13, RD92, ZK21]. Bright [Led80, RT85]. Brill [Shv22]. bring [GT90]. Bringing [HHL22]. Britain [How78]. broaden [Boa02]. broadening [DR13]. Brodie [Han12a]. bronze [Høy07]. Brown [NX23, PR08, dF12]. Bucharest [Ano70]. Budapest [Ano71c, Gen70]. bugs [NM98]. build [DS17, WB11]. Building [Eva99, Flü05, GT00, Kap09b, LH08, Sel98, Bas17, Fra13, Loc92, Sha77, Sho21, WH15a, WH15b, vBP19]. bundle [HEWC10]. bundles [HEW10]. Bundoora [Art23a]. Burkina [TB09]. Burn [Ber06, DDLZ97]. business [Hah14]. Busprongs [Aus81]. Bussi [Mor00, vBP19].

C [All99, Art23a, Gai11, Rut02, KSSO25]. Cabri [FLM07, AG02, Ass05]. cadre [HK99]. Cai [Ahm23, Ell18, Sch24, SLS74]. calcul [CB89, Gla68]. calculating [Gla68]. Calculation [FB02, Mur04, Thr02, CW14b, Gla68, HB21, Hop85]. Calculator [DZ00, Rut05, WN04, BJJ13, Bro10a, Fie87, GT00, Hed85, Rut90]. calculators [FT03, MWSS17, QM94, Tro00]. calculatrices [Tro00]. calculer [Gla68]. Calculus [DGA05, GD99, Har09a, Kle01, Ber08, Dib19, Güç13, HRR23a, HRR23b, KV75, MWSS17, Mar72, MRRF94, PM00, Smi70, TK14, Tho94b, You68]. calculus-with-computer [Smi70]. Calder [Wal12]. California [Gar24]. Call [Ano88c, Ano90c, Ano92d, Ano93b, Ano94e, Ano96c, Ano98e, Ano15, Ano16a, Baz01, CL18, Fit09, Ano74a, FSB18, KP10]. called [Bau80]. Cambio [Cas98]. Cambridge [Sfa15, Osi24]. CAME [SA07]. camp [RT85]. Campbell [Tsa02]. Can [ACJO04, CM92, MCB05, MRCR11, O’C01, Ryv04, BF21, Cob86, Cd04, Egs70, For00, Gos21, HSM22, JI15, MP24, Osl12, Ped07, PK94, Pol69, Ram03, ST08, Ser68, SRB16, TLM07, Van76, WBM^{+}20, dF16]. Candia [PW03]. candidates [Isa87]. cannot [PC23]. Can’t [Dre99, Gar24]. Cantor [Jah01]. Cantoral [LQV04]. canyon [PF11]. Cao [IPJ25]. capabilities [Vek71]. capability [NS25]. capable [NK17]. capital [AM25]. capitalism [Pai19]. captivating [DSR^{+}23]. capture [JP09]. Capturing [ZL91]. card [AKL80, O’B75]. Cardinal [Jah01, BL22]. cardinality [SDVT25]. care [Nuc24]. career [BWHM^{+}10, Sch24]. careers [FLT14]. Caribbean [Ern84a]. Carl [Mal23b]. carpentry [DMHRV21]. carrier [Høy07]. Carson [AAB03]. Cartesian [Moo20]. cas [BZ01, Win03, DF15, JMA19]. cas-based [Win03]. Case [Ass05, BBEG05, CWL07, Dou99, Ern06, EH06, HZ05, Mos04, SS06, THS06, TE97, TJHD11, WI01, XM17, AS17, Ale81, Atw24, AK19, BZ01, BLP^{+}14, BBDD25, BBVD25, BGT17, BF96, Bro10b, ÇA20, CV09, Cha18, Cob86, DHD23, DK15, EPQH18, Eve90, FC12, GD17, GBE13, GP17, GS20, Gos21, GB17, Gue19, GBR22, GLH20, HS22, HRR23a, HRR23b, Har09a, HMMG18, HMSBR19, HC14, HHT23, ICR20, JMA19, Jon18, JS01, Kar02, Kar10, KT01, KM24a, KLLJ17, KK13, Kon21, Kon22, Lam15, Led87b, Lev13, Lev22, LLL21, MMM25, MCVM13, MZ18, MA09, Mur24, Ng21b, O’C01, PMG^{+}94, PFSS19, Rob12, RJWB20, Saj03, SB20, Sch12, Sek21, SL94, Sin03, SMW11, SK13, SLR^{+}24, TN15, TRDA20, TL15, Tsa01, TTL08, TT23, ZS15, ZER17, ZL08]. case [ZKC18, dKV18, rA20]. case-value [BBDD25, BBVD25]. cases [AAJW25, BKGZ24, Miy17, PGV22]. casserole [Tro00]. Castelnuovo [Ano14a, FM14]. catalog [Ves69, Ves71]. categorical [GB13]. categories [Fin22, NGR82]. categorization [Jan09a]. Categorizing [SKPS16, DF15]. Catherine [DKL^{+}25, SMWM25]. Cauchy [TK14]. Causal [CL84]. causes [Kon72]. cautionary [Led87a, She87]. cavalier [Bal88]. cavalière [Bal88]. CBAM [TW09, CEMP04]. CBO9781139045360 [Sfa15]. CD [Dim04]. celebrate [Bak19b]. celebrating [Pai18]. Celia [Sta11]. celui [Bal88]. Centenary [Bag08a]. Center [Kru13b, Vyš75]. centered [Cas69c, CWYP92]. central [Ste13]. centrality [Gra04]. centredness [Wal10]. centrée [Cas69c]. century [BP80, Col21, Dam87, Her02a, Ric10, VPR21]. CERME [HTM^{+}23]. Certains [Mar69]. certainty [Ern16a]. CF [Har73]. CF/PMM [Har73]. chain [ST08]. Challenge [GC90a, Jaw02, MG00a, Jou13, KR09, Sil25, VLY21]. challenged [MA09]. Challenges [De 24, Fau15, Sil25, vO13, HHT23, Kim20, kSL13, Ng21a, SSL25, dV18]. Challenging [BT05, Sri18]. Cham [Ahm23]. chance [Nil07, Ste91, WK04]. Change [Bla78, Cas98, Ehr78, El 78, Fey78, Fre78c, Fre78d, HV78, How78, IE07, Kap78, Ohu78, Pur78, Rev78, Ros78, Sen98, Sha78, Wal17, Wil78a, Wil78b, Bro10a, BBL22, CS94, CB17a, CB17b, GC90b, Joh15, JM18, KHLP08, MR17, MK97, MN95, Mwa02, Sch92, Sfa15, Tou90, VC19, VGJ19, WD14, Law24].

Changes [Gun78, SMS22, AAP21, CPP09, Cre03, Vou12, Beh68]. Changing [Bur87, Ger81, NK17, PC98, TW09, bMYT98, Gra20, NT01, ZSL02]. characterise [PK94]. Characterising [Bak14, Hem10]. Characteristics [Leu05, OK05, ZZ08, DSR^{+}23, Her23, JMRW22, Kem81, SDVT25, SG98b, WSU19]. Characterization [HPG05, MBF25, ZER17, vdS11]. characterizations [TP22]. Characterizing [Mes04, Öst06, WFCW15, NH15, VVV15]. Charles [Ahm23, How73a, How73b]. Charly [Atw24]. Charting [Sch24]. charts [dSBM21]. chat [Sch13, Wil12]. chat-based [Sch13]. Cheap [Mie23]. check [Gar24]. Checker [SK69]. checking [Plu79]. Chemistry [Wit05]. Cherokee [BR78]. Chevallard [Art23b]. chez [Aud83, Bic82, Gro77, Hah99, Sam85, TLM07]. chief [GDS22]. Child [Boo81, SO73, TY04, Bic82, Com80, SDVT25]. Child-methods [Boo81]. childhood [LBT22, VWV21]. Children [APN01, Bri06, Com77, Coo98, CH02, DEMW06, Ell86, EJDT06, Eng97, Eng06, FYAE12, Had77, HH78, Had78, HK84, Kyn93, LW99, Mur04, NT01, New98, Nte06, PDT^{+}12, Sel98, Sel01, Ste83, VDTV24, Yim10, Ade90, AG83, APS22, Ang89, AV76, BHHS09, Bar21, Bar88, BTV19, BJ88, Bla85, BBG^{+}17, BIBS^{+}19, BTvG^{+}17, Bur80b, CHM83, CN00, Cle80, Cob85, Cob87, CSK12, DC92, Dav71, DP92, DVV20, Dek87, DCHM19, Eke77, EN79, Eng91, FFL80, Fie79, FNM91, Fro99, Gra88, HVN^{+}24, Hay87, HA23, JTP94, JLTM97, Kra02, Lar86, LCD90, Lin88, Lin90, LLHL13, Mak14, Mas69, Mat83, MS14, NBE^{+}12, O’B77, Obe19, Por90, Rou68, SM72, Sia24, SR92, SVLO20, Tho94a, TTL^{+}13, TDGV09, TTTL10, TTL^{+}17, TP15, Urb93, VS79, Van69b, VMN11, VO17]. children [Vou12, VGJ19, VS23, VS25, YC95, YST22, vO13, Mon00, PSS^{+}19]. Chile [DVC19, MMR21]. chimera [DMHRV21]. China [FSY22, Law00, Bec80, CBF07, DHD23, FZ07, Jin07, Kei93, Leu87, LLZ20, LLL21, Siu93, Swe77, SY79, TLB^{+}24, VS13, XXL21, XH88, YSL22, ZKC18, ZM91]. Chinese [JC22, CDW14, CL23, DL14, FQL^{+}17, GLH20, JHC14, Leu21, Ma96, PXTW17, Qua92, Sun11, SC83, ZML87]. Chirinda [Hyl24, Mur24]. Choice [BST81, BFG84, Beh68, HEW10, KV75, Lev22, Ng21b, RR23, SJ17, ZZ08]. choices [Bak19b, FYAE12, JMRW22, MS19]. Choosing [BS15]. Christian [Fur22, MDU^{+}25]. Christine [JV17]. Christopher [Atw24]. circle [Cha98, MLK16b, TP15, KV23, MLK16a]. Circumventing [Tri95]. citizen [AC21]. city [SG08]. cK [dV18]. claims [VRK24]. Clairaut [ABOM20]. Clark [Art15a, Fur19, Tab22]. Clark-Wilson [Tab22]. Clarke [Jon16, LX16]. Clarkson [Atw24, MR17, Wan25]. clas [Fyh08]. Class [Ass05, RR05, SSLM^{+}05, SK07, Ano78b, AC95, Byu24a, Byu24b, Cli17, Deb06, FT03, Gra75, Jan09b, LW94, Mog17, QP22, Rou69, SZP10, Tur76, WFCW15, Zah15, Add82]. class-room [Tur76]. classe [CGS76, Deb06, For69, PRR08, Rou69, VD78, VD79]. Classes [Ham06, Add82, BCFC90, PB09, RL13, WAL18, XXL21]. classic [SGCA18]. classical [KV75, Kno77, LS11, Pro70, RT09]. classified [Bar88]. Classifying [SG25, AS17, FDW19, Vol77]. classique [AL00, Pro70]. Classroom [Adl97, AF05, BT05, Bau92, BR17, Bus05, Cha06, Elb03, Flü05, MCB05, MF97, McN00, NN07, ODV06, PBY03, SL06, See01, Sou83, TN15, WI01, ZS07, AI07, APRS09, AC95, AH76, AL78, Bau80, Bek10, BJ22, Bra14, BS94, BSF96, Bro96, Bro14b, ÇA20, CYW92, CW96, DT12, DE24, DK15, DDB^{+}10, ERS19, FP07, For95, FA01, FA03, Fur22, Ger88b, GBC12, Güç13, Güç16, Gue19, Han10, HEW10, HEWC10, HMMG18, Hud90, Jun91, Kem81, KJ22, KBK21, LZ08, LSH09, Leu95, kSL13, Lor82, MPdP17, MM25, MLL19, NK17, NR10, Nol16, Nor15, PRR08, Pla14, Rad11, RR11, RNM12, SG08, Sek21, Ste97, TL15, Tay96, TD12, TMH21, Tic12, TJHD11, TR24, WHE08, WHE09, Wal17, Was15, XC19, ZZ08, dA00, dFS12]. classroom [rA20, Fer15, HSNHHE15, MR20, SJS12, Rad11]. Classroom-based [BR17]. Classrooms [Bes07, GP01, Leu05, Llo06, Mor00, Nte06, ARKK23, BS25, CWT12, EPQH18, EE21, FQL^{+}17, FKIA15, FBA10, GABB25, GdA09, GMHG09, Han12d, HWJ20, HP08, HMT13, IPJ25, IWF^{+}15, Kap09a, Lam15, LOW24a, LOW24b, LX16, LC15, McC14, NT19, PPG95, Rot16, SA00, TZLT95, VCD24, Woo96, YLH24, vdKG22, MR17, Han12a, JC22]. Clements [Hus25, Ste13]. Clever [FSB18]. climbing [Fyh08]. Clinical [Coo98, Gal81, HS13b, KH07]. clip [KB11]. closer [FKC17]. Closing [FF21, Ban11]. club [Cal94]. cluster [Ng21a]. Clustering [HK25]. CMAS [PSS^{+}19]. CMAS-UK [PSS^{+}19]. Co [LLD14, CBC25, Gre17, HM13, Jaw03, Mog17]. co-construction [CBC25, HM13]. co-educational [Mog17]. co-learning [Jaw03]. co-ordinated [Gre17]. Co-thought [LLD14]. Cobb [SJS12, HJ12]. code [Bos87]. codes [And09, GBE13]. Coding [EVR16]. cognitifs [LD88]. Cognition [BR06, GS08, Tsa02, BC12, Car08, NF09, NEM99, Pro13, Rad09, SP19]. Cognitive [AAT99, Bag06, Dah04, DS05, Duv06, Küc81, Lec92, LH06, PPC09, SA07, Sid05, SW94, Wat07, AOJ22, CL23, CPPC23, CPPT^{+}13, CF00a, FG17, Fro99, FM09, Gal19, HL94, Ior17, Ior18, KBK^{+}17a, KBK^{+}17b, KC09, LD88, LH96, McD89, MCPGE24, PCP19, RL12, SKK^{+}23, Sfa15, Sou83, SRR22, TK14, TF95, Wat02, Yan12, ZSWC25, vLXD18]. cognitively [HMNWB20]. coherence [CDW14, RBWS18]. coherences [Mey10]. coherent [IB19]. coin [Sfa91]. Collaborating [JWFC^{+}25]. Collaboration [Atw24, HMRRTT23, Cor17, HHB21, KM18, TS18, Tab22]. Collaborative [DEM04, DEMW06, PDV07, PPSZ19, RL00, Bju04, CY23, Car10, Fra13, FDW19, GG96, GGR02, HKSS^{+}25, HGM15, IPJ25, LOMG^{+}20, MYM12, NCC22, OAS25, Ryv04, SB20]. collaboratively [CM20]. Collected [Fur22]. Collection [FB98, SG25]. collections [WP24]. Collective [BMR25, MT15, CSS^{+}14, ES21, Gel08, MRCR11, RWZ15, ZC22]. College [FMA20, Rut05, WW07b, WN04, Wil05, Bag73, CY23, Har09a, Hol81, MWSS17, O’B73, QM94, Rot14, Sha77, TJHD11, Wai70, WM09, dVSKvO18, SE93]. colleges [For69, For69]. colloquium [Fre69e]. Colombian [AV08]. coloniality [Osi24]. color [AZ20, Bat13]. colours [NFFAP20]. column [Her02a, Tre87, WHC09]. Combatting [RR80]. combinations [SRR22]. combinatoires [Dub84]. Combinatorial [FG97, GBR05, Kap70, LG16, BNPG97, Dub84, HH81]. combinatoric [Eng91]. combinatorics [LCW20]. Combined [RNRVM23]. Combining [PFSS19, Cli17, KD10]. Come [Mur04, Fer15, Kon20, TPHS07, Wal12]. comes [SS03]. Coming [Bur98, SS06, Whi87]. comme [Art97, TLM07]. Comment [AAB03, RC04, Ser68, Fit15a, Fit17, Kei82]. Commentary [Ber06, Col21, Fit14, Rod06, Sch12, See01, Tab13, Ver12, Adl19, AGP15, BA25b, FV22, Han19, Obe19, Pea76, PM19, Sfa09, Siu93]. Comments [Hof06, Mar72, Ril86, Uhl03, Bur02a, DRR02, Dra14, Gri78]. commercial [Kir69]. Commission [Ano95e, Ano04c, Ano04d, Art05, De 24, How84, Ric10]. commitment [RR11, Shv22]. Committee [Art05]. Commognition [Pre16]. Common [ZML87, Fit15b, Fur22]. commonalities [Bur02b]. commonly [Hew12]. commonsensical [KMN24]. communicate [BCLH89]. Communicating [SLP06, Sfa01, Win11, Woo12]. Communication [Big70, Kae94, Ong06, BH95, HP08, NR10, RNP13, SS15, Wil09]. Communicational [Par16, Ste05, TN16]. communications [FE09, HI94, Ng16]. communicative [Büs24, FT03, TDEM13]. communities [Bro14a, MJS21]. Community [Boa99, WM09, AT21, Elb03, FA01, GC90b, Gra04, Kim20, Mog17, Nar16, PFSS19]. community-based [PFSS19]. Commutative [SS06]. commutativity [SL74]. Comparative [And09, DAJ20, Miy17, War06, BFA22, Bra11, CM20, FZ07, Gal82, JI15, LCD90, MRRF94, PBCPP25, TLB^{+}24, YSL22]. compare [SVLO20]. compared [Fre75]. Comparing [BWD10, Hin15, WM98, ALR25, Fos18, KID25, SVVO23]. Comparison [BCFC90, ČK15, CBF07, Pug04, Tsa99, WAL18, AH00, BKK22, CR09, Cla92, CF00a, Cor17, DT08, FKC17, GCALB15, Han20, HPRW25, HC14, Jab23, KV75, PPDS^{+}25, SLC95, SHAC01]. comparisons [FLT14, Rob90]. compass [Cha98]. Compendium [Ell18, Ell18]. compensation [Dem93, WH24]. competence [Abr01, BB25, CGD21, CLR16, EN79, EPQH18, ERS19, FT03, Fos16, GMHG09, Hie88, KBK21, Obe19, TH13, Urb93, CLR16]. Competencies [Høj24, AC21, AV22, Bec23, CLK22, CKS22, CL23, FV22, GGD22, KBK^{+}17a, KBK^{+}17b, LLHL13, LK22, NH19, Sáe09, YSL22]. competency [DMHRV21, DBL22, GJ19, PK23b]. competency-based [DMHRV21]. competing [ES03]. competition [Kon20]. complement [DeB97, DeB97]. Complementarity [Ott03]. Complementary [Wit71, Kid08, Wit81]. complementizing [SP19]. completed [RW11a, RW11b]. completely [Sur77]. completeness [Ber08]. Complex [Lab00, DS17, Gol03, KRHD15, LMCG08, SRR22]. complexifying [SP19]. complexities [Mat03, TMH21]. Complexity [ODV06, BWD10, Gri83, HTM18, HM25, JP09, Kap09b, LE09, Pla18, Tho93, Was15]. complexity-discursive [HM25]. Components [Cop98, Kun12, Kry88b]. comportement [Add72, KC73]. Composantes [Cop98]. Composants [Kry88b]. composite [Leh23]. composition [MYZ23]. compound [CSP24, Par25]. Comprehending [Uhl02, Bur02a, DRR02, MGD^{+}23]. Comprehension [CF00b, Duv06, Öst06, DAJ20, Gel98, KCS22, MRFW^{+}12, PC23, Rob75, YL08, Yan12]. comprendre [Kra02]. comptage [TB09]. computation [CB89, MNRR95, PLW81]. Computational [Edw84, JS99, JM74, Zah15]. Computer [Ern86, GH98, GS76, VL04, AP75, BC92, CJ75, CBSS96, Dag96, Gan92, GBE13, GS72, HC95, HKG89, Jah83, Kor03, Kyn93, MG00b, Nos88, RH02, Sam85, Smi70, SW94, SLJ69, Sut89, TAH08, TT91, TPHS07, Thw69, Wal68, You68]. Computer-assisted [GS76, GS72]. Computer-Based [VL04, CBSS96, GBE13, RH02, SLJ69]. computer-provided [HC95]. Computerized [Kap01, HN96, Tro04]. computing [Art97]. conceiving [TLM07]. Concept [BM08, DWMB05b, DWMB05a, EJ23, FB98, Has97, HM95, LW99, PLW81, Prz05, TV81, BS80, DPG89, FYAE12, Fis77, FJC95, Fre69c, Fre93, Gel04, Güç13, HS22, KČ12, Kor03, MO81, ML20, MRCR11, Noe80a, Noe80b, Rad21, Rel96, Ryv04, Saj03, Sam85, Sem08, Sim17, Ste91, Str78, TGWN23, TP14, TT23, TOT24, Vin76, Vin77, WSR11, Wit73a, Wit73b]. conception [BDHN92, MPGDA12, Mos98, PP12, Sim17, TT07, Tzu07]. Conceptions [GKMJ24, GT95, ST02, Sch11a, AV08, AH00, Boa02, BHN10, BF96, GS18, HS13a, HK84, LK16, MK97, NH15, Par91, PB21, RMAGMR22, RT09, SGRS24, Sfa91, ST10, TK14, Tha10, Tho84, VS23, Woo12]. Concepts [Bag06, BH05, DDLZ97, Haz99, Kie81, RL12, BL22, Bur96, Cha98, Cla12, CBSS96, CBSP12, DC92, DB86a, DB86b, Dom25, lD95, Dri25, DDLZ94, FPS16, Fis93, GFWL11, GT95, GB14, Hah75, Han19, HC95, JSV75, KZ94, Kon18, LHZ95, Mas69, MW00, PM15, PH96, Sch16, Sch11a, Sem84, Shr10, SL74, SKPS16, Sin00, Tch11, TS93, UOF25, VA18, Ver79, WS08, WS25, YC95, Zen18]. Conceptual [Cas98, GMH23, PBR^{+}04, Vin97, BFA22, BP15, BE23, CG21, Edw09, HP25, HK99, JK17, KWF15, LMW91, Mar13, Nos86, PM00, RBWS18, RR80, Sáe09, SSPS20, SV13, Sle82, SKE25, VDTV24, VSS24, Vou12, ZW21]. conceptualisation [LK22]. Conceptualising [AGL17]. Conceptualization [Dou99, CL20]. conceptualizations [CN16]. Conceptualizing [BR12, Coo99, CKS22, RH19]. conceptually [Zah15]. conceptuel [HK99]. concernant [KC73]. concerning [CB17a, CB17b, KC73, Str78]. concerns [CP10, CEMP04, TW09]. concerns-based [TW09]. concevoir [TLM07]. conclusions [HM95]. conclusive [Vre69]. Concrete [Cob86, Arm69, Dié71, DL14, FM14, HRR23a, HRR23b, SW14]. concretization [Sem84]. condition [Her02b]. Conditional [AKL80, IS08, IS09, CV16, GT95, Had77, HH78, Had78, Mar86, YLH24]. conditionnelle [GT95]. conditions [Her11]. conducting [CW96, dPQ16]. conduit [Rot14]. Conference [Fre78b, Ano71b, FE09, Tab22]. Confidence [Fos16, FWCW22, Gra04]. configurations [Dub84]. Conflict [Wat07, Zev06, BW13, DC92, Gal19, Pre88, Wat02]. Conflicting [HMMG18]. conflicts [BG69, CN25, DC92]. confronted [Gel04]. confronts [Fit14]. congrès [Fre69a]. Congress [Ano91c, Fre69a, Fre69c]. congruence [Ste71, Zas71, ZL91]. congruent [CCRD22]. Conic [Ano84c]. conics [Pic71, CGGGG79a]. Coniques [CGGGG79a]. conjecture [Blo03, FG17]. conjectures [AI08, VE24]. Conjecturing [LS11]. Connecting [LL23, NHH97, Sut93, TP22, BP22, Gra20, HHL22, LLW07, MDPW96, Fur22]. connection [LW89, Moo20]. Connections [ASP97, Del69b, Pre06, DS17, EVR16, KL14, Loc11, MK14, RNRVM23, SG98b, Tch11]. connectives [Rob75, Zep82]. consciousness [BR12, SRP21]. consequences [D’A91, Gra91]. conservation [Kor03, KSL11]. consider [LW14, Nkh02]. consideration [BCE^{+}19]. considerations [DG03, LH18, NBZ12]. considered [RR17]. Considering [BLC20, HSM22]. Consolidating [THS06]. Consolidation [MO06, Kid08]. Constant [YG99]. Constantinos [Fur19, Gra20]. constitution [Str78]. Constraints [Ste05, RR23]. construct [APN01, Bec23, Büs24, DR13, McQ86, NS25, SM22a, UB23, ZW21]. constructed [SJ17, Sin03, Sin00, Zen18]. constructed-response [SJ17]. Constructing [GMHG09, Lam15, LSH09, Nos86, PGV22, Sch16, THS06, Abr09, AP01, ETC14, FQL^{+}17, Hun86, JS01, KP10, PK92, SAR19, Web01, WJW11]. Construction [Eng06, MDD04, NHH97, OB06, OPC^{+}98, WMN22, WMN25, AP25, AC95, Bar09, BZA01, BC96, CBC25, CHK22, GPPT99, HS99, HM13, Kom17, Kon93, KCS22, MB09, MRCR11, MYM12, WR96, XA20]. constructionist [MK14]. constructions [KD10, SS09, VSS24]. constructive [SS03, TJ94]. Constructivism [Ler89, Loc92, Ste98, Zev96, Har09b, Eng99]. Constructivist [MHDC04, ODV06, Ant96, Bau92, BJ88, PK92, Rot17, SS91, SS93a, ZM25]. constructs [RDH10, Sim17]. Consumer [Eng06, ETC14]. Contemporary [Bro23, Feh69, BW12, Ell18, NX23, Wil12]. Content [BE84, Dör03a, Boa93, BOL24, CLSM17, DVT^{+}18, GMNB18, GSW22, GB17, GS18, Jan09b, KS69, Kom16, Kun12, MM11, McD89, OAS25, VP13, dVSKvO18]. content-oriented [KS69]. content-specific [OAS25]. Contents [Ano01a, Ano02b, Ano04h, Ano06a, Ano00i, Ano00j, Ano01f, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g]. contests [Ano69]. Context [Cha06, Dou99, GD99, LK14a, Men07, SA07, Sty07, ZC06, AV22, AVM15, BFG84, Bis88c, Boa93, CV16, DVC19, FKC17, Kru13a, LSH09, LH18, LYBK23, LH96, LLL21, MP24, MMP15, MWVE09, MO94, Mwa02, Ngu24, NEM99, OSR71, PK17, SGRS24, SB97, SPFGF14, SVVO23, Swe95, TP14, TE14, Wed99, WvdHPD15, Yim10, ZWMR16, vO10]. context-based [MP24, WvdHPD15]. contexte [SB97]. Contexts [Baz90, DP99, Jur06, Lab00, Ano81b, Cob87, CF00a, DGI23a, DGI23b, HMNWB20, JC22, LR20, LL99, SSP04, SVVO23, dA00]. contextual [HB21, HS14, KB11, RH19, Sáe09, Van84]. Contextualization [RM21, LK14a, NR10]. contextualized [GC23]. contextualizing [SP19]. continue [FVJL^{+}19]. continues [Pre03]. Continuing [YSRU03]. continuité [Rev72]. continuity [HSZ17, LNC23, Rev72, TV81, TK14]. continuous [EEST20, Kid11]. contortions [Gar24]. contrôle [Plu79]. contracts [KM24b]. Contradiction [BW13, HHS00, SRB16]. Contradictions [SPS17, Pai11, PPSZ19]. contraposition [SSP04]. Contrasting [CH02, Jur06, EK10, Ng21a, PGV22]. contre [BZ01]. contre-exemples [BZ01]. contribuer [Add72]. contribute [Czo18, Jan74, dSBM21]. contributing [Add72, Buc87, Gli91, KAJHB10]. contribution [FF21, Kil01, Kra93, Pai16, Rot12, SDVT25]. Contributions [DP99, Gue19, Hof06, HS14, JMD08]. control [AS11]. convention [MZ12, ZR18]. conventional [CZ11, DC17]. conventions [LR20]. Convergence [AS04, AS05, Ber06, Prz05, Mar13]. conversation [Kem81, SZP10, TD12]. conversational [MFG22]. Convexity [Pat85, Wit76]. conviction [LCW20]. convincement [RL13]. convincing [Her93]. cooperative [LCK24, Ma96, TA08, Tar09]. coordinate [MWVE09]. Coordinating [CWL07, HK25, TRDA20]. Coordination [EJDT06, GL98, Ano68a, PK23a]. cope [Gal11, KZ16]. Coping [LH97, Ng21a]. core [Bja20a, Bja20b, Wit84]. coronavirus [RA21]. corporate [dKV18]. correct [CM92, OVKR25, RDH10]. correcting [Bar87]. Correction [Bja20b, BBDD25, Byu24a, DKBD21, HRR23b, HMG20, HMCT25a, HH24a, Kon22, LOW24a, Mah21b, RG20, Sch25, Web18b, WMN25, rA20, Kar16]. correlation [GT95, GT95]. Correspondance [Her11]. correspondances [Bal88]. Correspondence [Her11]. corresponding [Usi74]. Corrigendum [Ano98f]. Corte [Sel00]. coset [MGD^{+}23]. Costa [Mal23a, CN25]. Could [Hof06, NN07, Sfa05]. couldn’t [For00]. counsellor [JN15]. counted [JGF20]. counter [AZ20]. counter-story [AZ20]. counterexample [ZC08]. counterexamples [BZ01]. Counting [BT04, BL22, Jer70, Loc11, TB09, VMN11, Vou16]. counting-based [BL22]. countries [Ale81, DKL^{+}25, Han89, Hyl24, Kap09a, KDM21, VR90]. country [Atw24]. counts [MS20]. coup [Fre69d]. cours [BC78, BC82, Rob95]. Course [GD99, MHDC04, Wit05, BC78, BC82, FG17, Fre69e, Fre74a, HYC^{+}97, HWR23, Prz04, RSRVRVVR23, RPP24, Str93a, WS25, WR77, lRA16]. courses [FCN14, Har09a, HOS13, How75, Jan09b, JWFC^{+}25, MSdSGG16, Rob95, Wai70, WMM13, ZZ11]. Covariation [HGM15, BFA22]. Covariational [BEBL^{+}24, PGV22]. COVID [AC21, Bor21, BBL22, DTD21, GG22a, HMSBT21, KM21, KDM21, KHY21, SRP21, XXL21, dSBM21]. COVID-19 [AC21, Bor21, BBL22, DTD21, GG22a, HMSBT21, KM21, KDM21, KHY21, SRP21, XXL21, dSBM21]. cows [Vog70]. Create [Zas05, Ryv04]. Creating [Bro94, DZ00, LBML24, PK92, GD17, GP22, GGR02, NT01, TS93]. creation [Hic24, Wil21]. Creative [HR00, BLC20, BBAR23, Her11, Lit08, NJL25, PK92, Her11]. Creativity [Fre77c, LK22, OZ85, Pre03, VS13, BLC20, BHN10, Hay87, KLBMK^{+}13, Lev22, LBML24, MLL20, REL25, Shr10]. credit [CB19]. crises [Sko21]. crisis [AT21, CSW21, Cd04, DGI23b, Gos21, KMN24]. critères [Fre82]. Criteria [Fre82, Kno77]. critic [GP22]. Critical [Bra14, Che90, GFJ13, IE07, Kei82, RC04, SB12, AAB03, AAJW25, Bra11, BSW16, Bur90, CL23, CW14b, Cot24, GM84a, GM84b, HP12, HC82, Kap74, LK16, MW19, RK17, RPP24, Rot21, RC02, RNC21, Sko85, Sko94, SRP21, VDWV04, Wei17, lRA16, BMK21, Cot24, SK25]. Criticising [Kol16]. Criticism [Gra75, Gre17]. Criticisms [Pai11]. critiek [Gre17]. Critique [VS97, APS22, Lun12, Pai13, Wil16b]. Cross [EJDT06, KL14, RR05, CL84, HHB21, JC22, MCVM13, QP22, SV25, SGTK14, VRP^{+}12, VS25, YC95]. Cross-Analysis [RR05, KL14, MCVM13]. cross-cultural [HHB21, JC22, SGTK14, YC95]. cross-disciplinary [VRP^{+}12]. cross-grade-level [SV25]. Cross-Sectional [EJDT06]. cross-task [VS25]. Crossing [LH96, BA14, Bak16]. crossroads [LHZ95]. crux [Fre74a]. cryptography [Whi09]. CSMP [Ano71a, KS69]. cube [Gla70]. cubes [BHLH85]. cues [NT75]. Cultivating [YLH24]. Cultural [Ano81b, GL00, GB17, Ler01, Bis88c, Bro12, Bur90, Car10, CL84, Cra12, HHB21, JC22, MRC21, Nos88, Sfa15, SGTK14, XC19, YC95, dA00, Sfa15]. Culturally [Fit11, Cra12, Mat03, Mog17]. Culture [BHX97, Lou18, MHRSU99, SK98, SK07, Bis79b, Boa93, Che90, CM20, ETC14, Ger88a, GP22, Han12b, Leu21, MPS12, Mor21, NK17, Pre88, SJS12, vO01]. culture-conflict [Pre88]. Culture-Specific [BHX97]. cultures [Bau92, BJ22, CB17a, CB17b, Mit80, Nos01, TMH21]. cups [CO10]. curious [Mar86]. Current [XH88, CKS22, FF21, NK17]. curricula [BJJ13, Gri71, MRC18, NK17, Osi24, SMWM25, SS10, WAL18]. Curricular [BCFF^{+}98, Gue19, ZC06, CMW^{+}13, CHM16, NK17, Gue19]. Curriculum [Ano76a, Coo98, HBH07, Kap77, Kei82, NC06, Pop88, RC04, AAB03, CN25, CP10, Che89, CEMP04, Die15, DKBD19, DKBD21, GC90b, Leu87, MRC21, PMG^{+}94, PPSZ19, RK17, RC02, TTS21, Tay91, XH88, Zav21, dAS22, dKV18]. custom [Her02a]. cut [Bos93]. CXC [Ern84a, Isa87]. cycle [Bui76, Cas69c, Cas69b, Lev92, TSV82]. cycles [Sim94]. Cyclic [CB05]. cycloid [Bar75]. cyphering [Ste13]. Czechoslovakian [Vyš75].

D [Ott76, Dam87, MP96]. d’addition [AL00]. d’aire [DPG89]. d’Aix [Art23b]. d’Aix-Marseille [Art23b]. Dakar [VD78]. d’Alembert [Gos21]. d’analyse [GT95]. Danesi [Mal23a]. dans [Ano72a, AG02, BD72, Bos87, Bui76, Cal94, Cas69b, CGGGG74, Dag96, DS02, Dum82, DG89, FBM71, Hah99, Kon72, Kry68, Kry69b, Kry88b, bL00, Mas69, Mau85, Pro70, Rev72, Sar02, SB97, Ser76, Siw73, TSV82, Tro00, Van69b, VD78, VD79]. d’apprentissage [DPG89, Tro00]. Data [Eng12, KHRK15, Leu05, MC01, WM98, ALR25, AGP15, AP01, AWL15, BZA01, CPPC18, CVN21, FB02, GT95, GB13, GC23, HK25, Han10, HTM^{+}23, LH18, LT86, LMCG08, Min83, RT14, RNC21, SL19, SV25, VRP^{+}12]. Dataset [Rut05]. David [Jon16, LX16, Bye15, LT20]. Davidov [Ott76]. Davidov/D [Ott76]. Davydov [Col21, FF21, MRC21, Rad21, VPR21, VAM21]. Dawkins [MDU^{+}25]. day [Ano74b]. de-ritualization [HMCT25a, HMCT25b]. Deaf [Fro99, SLS74]. Dealing [Sch20]. Debate [LH06, HGM15, WMPL08]. debates [FF21]. debutants [Sam85]. decades [Law24, LK18]. Decentralising [UOF25]. d’échecs [Kon72]. decided [KP10]. Decimal [MM00, BST81, O’C01, SHAC01, Wea90]. Decimals [OK05, GT90]. Deciphering [CL23]. Decision [Bis76, CN25, STM06, FDW19, OCB25]. Decision-making [Bis76, FDW19, OCB25]. decisions [Ecc85, MZ18]. Deconstructing [Hin15]. deconstruction [MB19]. deconstructive [BA25a]. decor [Dum82, Dum82]. decreasing [AS17]. Decroly [Lib71]. decryption [Whi09]. Dede [Wan25]. dedicated [Sch24]. dédoublement [Duv83]. déductif [AM97, Duv91]. Deduction [Cha71, Sid05, Chr69, Mes89, Mot71, Cha71, Mes89, Mot71]. Deductive [AE08, Jon00, AM97, Duv91, Jan74, Kom16, MP14, MFJ17, Rot21, Sim96]. Deep [Sem08, WD05, Fre71, Mev83, SD13]. Deepening [ZC06]. deeper [PXTW17]. Deficit [Sol06]. deficits [Sch23, Sch25]. Define [TE97]. Defining [MF97, AS17, FDW19, OB11, Tay91, TGWN23, UMC24]. definite [Jon18]. Definition [OB06, GJ19, KV23, MRCR11, Roh10, SG14, TV81, Vin76, VEBS24, WSR11]. Definitional [BKGZ24, Vin77]. definitions [GDK23, OB11, Roh08, RR23, TN15, VE24, Yah25, ZL08]. deformable [MP96]. Deformation [O’B75]. defying [AMMS20]. degré [PRR08, Rev72]. degree [AM25, Ger10, XXL21]. Deleuzian [ML20]. d’élèves [De 94, DS02, GT95]. Deliberate [STM06, BC12]. Delphi [MM11]. demand [AOJ22, Küc81]. demanding [HMNWB20]. demands [CW14b, GG22a]. Démarches [DP77, ZP91]. democracy [Sko90]. democratic [Tay91]. Demonstrating [ATAB25]. demonstration [CLR16, Duv91, Pas78, CLR16, Duv91, Pas78, Rut02]. demonstrations [Bal80, Kon72, Bal80, Kon72]. Denkstufen [Van76]. denominators [Str82]. d’enseignants [PRR08]. d’enseignement [BC92, Rou68]. d’ensembles [DeB97]. d’entraînement [Cal94]. d’environnements [Art97]. Department [WD14]. Department-initiated [WD14]. departments [BWD10]. depicted [TP15]. depth [Bek10, DVJV02, Han12a, HMA13]. Derek [Jon16]. Derivative [ASP97, HRR23a, HRR23b, MZ12, Par15b, Par15a, Par16, Zen18]. Derivatives [Hir72, BMPT24]. Derive [Art97]. deriving [GRR23b]. Describe [HDM07]. describing [BWD10, Dra14, Hoy01]. description [vdS11]. Descriptions [Pug04, Bar09, JCM^{+}23, MK97]. descriptive [BR22, DB86a, Fri21]. design [Abr09, AGP15, Bel93b, CSK12, DBL22, Fre74a, Fur22, GGD22, Had78, Hah14, KSS25, LL23, MM11, MP24, PPSZ19, PHS12, Row08, Sin03, Str93a, TP22, Wit95, ZHT25]. design-based [KSS25]. designed [CM20, RIA17]. Designing [Bas17, LMCG08, SMW11, Wit05, HN96, Hoy01, JN15, Ric20, RTSK08, SV13, Van93a]. designs [MW09]. desolve [JMA19]. dessin [Bal88]. dessins [Bal88]. details [Ell18].

Determinants [Tie02, GLH20, Nes76]. determination [GP17, Kon72]. determine [CVN21, Kon72]. determining [SPNH12]. deterministic [MZ18]. d’étudiants [Cal94]. deux [Del68, TSV82]. devant [De 94, DeB97]. d’éveil [Mes89]. develop [TJ94, ZB19]. Developed [Ryv06]. developer [Art97]. Developing [BH05, BWHM^{+}10, BKH25, BC12, Fos18, GLZBZ15, GGD22, Gol03, IB19, LXT02, MMP15, Mor00, PBY03, PXTW17, SLP^{+}24, SS15, TD12, TH13, TGWN23, Vek71, Wat07, Wit01, Ale81, Bak14, CSK12, DHD23, GFJ13, Hie88, Hoy01, Jou13, Shr10, SS88, TN15, TDEM13, VR90, YTD11, Cor22]. Development [Ano04d, BDHN92, CBSS96, CB00, Dah04, DP99, Eng97, GRR23b, GBR22, II05, Jaw03, KSS25, Kru13b, LC13, MW00, OK05, Riv07, Rob75, SGRS24, Sfa15, SA07, Wal73a, WM01, Wat01, Yan19, Ano71a, Ano76a, AM25, Bar88, BL22, BR22, BB25, BS94, Bro14b, BTvG^{+}17, CPK08, Cha98, CSDK25, CWH19, Cra96, Ell13, ES96, Eve99, Fis69, Gel08, GE98, GGR02, HM95, HPS^{+}19, HM13, HMMG18, HMSBR19, HMNWB20, HHDG16, HMT13, JSV75, JP09, JE14, Kap77, KWF15, KV23, Kru13a, LDM92, LYBK23, Mal23b, Mas69, NT01, NGR82, Noe80a, Noe80b, PSS^{+}19, RL00, SM17, Sem08, Sev22, SS91, Sou83, SDVT25, SJ04, TS18, TLB^{+}24, TP14, UOF25, VRK24, VC19, Van69b, VVV15, VWV21, Vau71, VYZ21, WS25, Wil16b]. development [Win11, Wit69, WDvO19, Wri94, XH88, YC95, YK20, Zas71, ZHT25, dVSKvO18, DB86a]. developmental [FMA20, HYC^{+}97, Han19]. Developments [Kap80, Sfa13, BP15, CKS22, Fey89, Vol76]. développement [Fis69, LDM92, Mas69, Van69b]. deviance [Bul24]. Devices [Baz01]. devil [Fre71]. devils [Vog70]. Devising [AG14]. Devoir [Cop98]. devolved [SHG21]. devoted [BVR24, Beg71]. devraient [Kry88b]. d’exemples [BZ01]. d’expériences [AP75]. d’exploitation [Bra68]. d’habiletés [LDM92]. diagnostic [Bel93c, DB86a, DT08, FWCW22, SLP^{+}24]. Diagram [dFS12, FF25]. Diagrammatic [BH05, Kru13a]. diagrams [CH13, Fre69b, HK25, Ng16, PGE09, Van69a, Vre69]. dialect [Win11]. dialectic [AG02, Bar16]. dialectical [Rad21, SPS17]. dialectique [AG02]. dialog [HJ12]. dialogic [FDW19, KWF15]. dialogical [HMRRTT23]. Dialogue [BGT17, Zev06, Bar16, BB03, Con90, LK16, MYZ23, WHE08, ZG01, ZK15, Web18a, Web18b]. dice [Nil07]. Didactic [BBEG05, DGA05, Flü05, Gel04, HPG05, Jah01, Art97, BKH25, DG03, KM24b, Sch83, TH13, Van69b, Wal76, ZHT25]. Didactical [AL78, BG05, MD01, MW09, Usi85, ABM92, Fre93, MM25, PvdHPR12, Ste69b, Van94, Van03b, Win03]. Didactics [Fri07, Str08, BB03, Gof77, Gre17, GAP24, Tep20, WMM13]. didactiek [Gre17]. didactiques [Art97, TSV82, Van69b]. Didactising [YSRU03]. Didactizing [Pre03]. didaktische [Wal76]. didaktischen [Ott76]. Diego [GS20]. Difference [Sol06, VL04, Dör03a, ETC14, Mon00, PK23b, SPNH12]. Differences [BHX97, Mar13, Zev04, And89, Bur02b, Che89, Cle79, Eli85, FP85, GR71, HH88, Kae88, KM93, Led85, Lin88, LLHL13, Mon00, SC17, Say94, WL90]. Different [BCFF^{+}98, Ber19, FB02, GL00, II05, MCKP19, Nil07, SRR22, Van93a, BJJ13, BB03, CW14a, DGI23b, EG23, HPS^{+}19, Kid08, KAJHB10, KHRK15, Lam15, Lev13, MTBR16, MWVE09, SP16, Sfa91, ST20, SC93, Str82, TAH08, VN19, Wai70, Cas69a]. Différentes [Cas69a]. Differential [Ade98, Eva07, ZWC^{+}24, LOW24a, LOW24b, Min83, Zen18]. Differentially [Kni12]. Differentiation [Noe80a, BMP20, Ort83a, vdKG22]. difficult [GFWL11]. difficulté [DeB97, TLM07]. Difficultés [Add76, Gro77, SB97, VD78, VD79]. Difficulties [Ale81, Bag06, CCRD22, AG21, GL10, Gro77, Has81, Hew12, HM13, KR11, NBA23, PM15, SLP^{+}24, SB97, VD78, VD79, Add76]. difficulty [DeB97, Jan74, JR72, Nes76, TLM07, Web01]. Digging [Men20]. digit [Csí16, KBK24a, PDT^{+}12, Sel01]. Digital [HbL05, LK14a, MK14, Art15a, BdSCdA18, Cal12, GJ19, HFL^{+}23, Jou13, KML22, KL14, Nos01, SSA20, Tab22, Wal12, XXL21, Ric20]. dilemma [SPC^{+}16]. dimension [Ior18, Way94]. dimensional [Coo92, Mit80]. dimensions [Bau80, Boy16, Edd18, Sri18]. d’indépendance [Mau85]. d’individualisation [Gau69]. Ding [JC22]. d’initiation [AM97]. d’instruments [bL00]. d’intervention [Deb06]. Diophantus [Har87]. direct [Ler85, Mah21a, Mah21b]. direction [Bur80a, GCALB15]. directional [AGL17]. Directions [BRE20, CL20, Sfa13, SEC13]. Dirk [Bja20b, Bja20a, OM24]. disabilities [BJJ13]. disability [Lam15]. disabled [HM13, NRV24]. disadvantaged [SLS74]. Discernment [LBFM13, JM18, NHO22]. disciplinary [DR13, VRP^{+}12, VN19]. Disclosure [SSA20]. disconnections [Gra20]. discontinuity [LNC23]. discours [Bos87, Rob95]. Discourse [EPQH18, MHDC04, PW03, See01, AWE22, BF21, Bos87, Bra14, CH13, FF21, Güç13, Güç16, HRR23a, HRR23b, HEW10, HEWC10, HMSBR19, JB21, Jun91, KFS01, Kie01, KB11, KLwK24, MBF25, Par16, RC20, SG14, Sfa01, TJHD11, Wal17, XC19, Zev96, lRA16]. discourse-rich [HMSBR19]. Discourses [NBA23, Ryv06, Bar16, DVC19, KMN24, Kon21, Kon22, MAM24, Ngu24, Rob95, Ryv04, SECP^{+}23, SA00, VN19, Win11]. discoverer [Hol73]. discovering [Mey10]. discovery [FA76]. discrete [DH90, Kid11, Wit80]. Discursive [EMT06, Ler01, HRR23a, HRR23b, HM25, KB12, Ric04, Win03, XA20]. discursive-graphic [Ric04]. discursivo [Ric04]. discursivo-graphique [Ric04]. Discussing [Vil10, Yah25]. Discussion [Ano00d, Ano04d, BT05, Blu02, FvM97, ZS07, Ano68b, Bus96, CKS22, Ern84b, GABB25, Gut07, Hoy85, LW94, O’C01, Pai16, PS88, Vou16, Wal76]. Discussions [McN00, Byu24a, Byu24b, CBC25, Jan09b, MFG22, RPP24, SGB21, WMPL08, dPQ16]. Disentangling [GH98]. disjunctions [DC17]. Diskussion [Wal76]. disorder [Sri18]. disorders [REL25]. Displayed [MM00]. Dispositions [Nol12, SLM79, Zev11]. disproving [VV25].

Disruption [CM92]. Dissecting [KK13]. dissenting [dF04]. Distance [DTD21, SLW83, AAP21, RP21]. distances [RP21]. distinct [KRHD15, TLM07]. distracting [SLM79]. distractrices [SLM79]. Distribution [BH05, LLHL13, SGRS24]. distributions [NH15, SV25]. distributive [DL14, HL16, MM20, SM17]. district [Ngu24]. diverging [Gra91, GPPT99]. Diverse [PK23c, Cra12, DKL^{+}25, OM24]. diversity [Art23c, SJS12]. divide [Dem93, Ern16b, JP09]. diviser [Dem93]. Division [OK05, GT90, HC95, Jou25, KPH10, Lee17, Neu99, PM16, RL12, SW95, SS10, Sun11, TSV82, TG89, Ves71, Yim10]. Do [EMRI22, GG96, MB16, Mur04, Ala12, AK19, Bur98, CDW14, EK10, For00, HDM07, HH11, HBZE98, Kar02, KL20, Kon20, KR11, NCC22, Par15b, Par15a, Por90, Rad09, Rob86, WBM^{+}20, WJW11, YM16, ZB19, dF16]. Doctoral [AI08, GPW05, Her02b]. Document [Ano00d, Ano04d, BT05, FvM97, Blu02, GMNB18]. documentation [GMNB18, GT09, MR17]. Does [Ade90, Bro10a, FQL^{+}17, FG84, GSH16a, How80, IIMR^{+}11, KSH25, SAR19, VL04, AC21, BSV95, BA23, Czo18, DBL22, Eng76, FB17, GMH23, GBC12, Mon00, Mor06, Sca75, SWVO23, VLY21, YE16, GSH16b]. DOI [Sfa15]. doi.org [Ahm23, NX23, Sfa15]. Doing [Bur90, Cor17, LM11, OLW20, Ahm23]. Domain [HbL05, Mos05, SL19, BN96, Gal02, Kon18, Kun12, LLZ20, RK17, Mor00]. domain- [LLZ20]. domain-related [Kun12]. Domain-specific [SL19]. dominant [Abt22, Dom19, GDS22]. Donevska [Tab22]. Donevska-Todorova [Tab22]. données [GT95]. don’t [JB21, WHE08]. door [MW19, Swe77]. dosage [CW14b, HB21]. double [Fri11, LNC23, PW22, dKBvdS22]. double-blind [PW22]. double-edged [dKBvdS22]. douze [De 94]. down [Loc92]. Dr. [Bac21, Pea76]. draft [VE24]. drafts [VEBS24]. Dragana [Mal23a]. dragging [Ng16, PD23]. dragons [Vog70]. draw [SECP^{+}23]. Drawing [CPP07, ST10, Bal88, Bus96, Lén93, Mit80, MB19, RSL17, dKBvdS22]. drawings [Bal88, CSK12, Ger88b, Mar86, MB09, VO17]. Dreyfus [Fit15b, Rad11]. driven [LLL21]. dropout [GRR23b, GRR23a]. dual [Sfa91]. duality [SBDD24]. d’un [Aud83, For69, LDM92, ZP91]. d’une [BC92, Bra68, CLR16, Her11, LD88]. d’université [BZ01]. duoethnography [ZK15]. duplication [Duv83]. Durand [Mal23b]. During [DEM04, AWE22, BCZ21, CCW95, CY23, CPK08, DTD21, GB14, GAP24, HKSS^{+}25, HHL22, Jan09b, JMRW22, Kap01, KDM21, Kyn93, LOMG^{+}20, OAS25, PT09, Rad03, RPP24, SLP^{+}24, SB20, dVSKvO18, lDlzN21]. Dutch [vSBvR00]. Duval [Rut02]. Duvalian [Ior17]. Dyadic [THS06]. Dynamic [GPW05, HHS00, Jon00, Lab00, Mar00, ATAB25, BF21, BKH25, Fin22, FBA10, JDK^{+}24, Lab78, LBFM13, LL13, MG00b, MAHK08, NHO22, PPC09, PD23, Rot16, Sin03, TY04, VSS24, YS12, Zen18, HO23]. dynamics [AT21, OCA13, VGJ19]. dynamique [Lab78]. dyslexia [LH22a]. e-assessment [KID25]. E. [Bag08a]. each [Fra13, Sfa15]. Early [BBFG99, KPS83, Mos05, Neu99, VWV21, ZC19, AAM21, CPPC18, CS19, DOV24, FS16, FF21, Her02a, JC22, Kru13a, LBT22, NFFAP20, PBCPP25, Reu23, Sam85, SLR^{+}24, Tak18, Van03a, VAM21, VMN11, vdHPKR13]. ears [Sfa01]. earth [Ano84c]. East [Leu01, Leu05, LYBK23]. Eastern [Law24]. eBook [Ahm23, Cor22, Mah21a, NX23, Shv22]. école [AR82, AG02, BD72, Cas69b, CDGG76, FPM69, Gla76a, Mar69, Roy81, Fri21]. ecological [CSRlR24, GBC12]. Economic [II05, Kae90]. Economy [Ass05]. Ed [Ell18, Goo20, Gra20, Gue19, Law24, Mah21a, Mah21b, Men23, OM24, Pai18, Sfa13, Sil25, All99, FSY22, IPJ25]. edged [dKBvdS22]. edited [Bro11, JV17]. editor [Bak19b, Han98, Fit15a, Fit17, İşl23, Pre23, Tun04, Wil05]. Editorial [Ano85c, Ano88d, Ano98g, Ano98h, Ano99b, Ano00e, B.88, Bak19a, Bak20, BW21, BW08, Bis79a, Bis85, Bis88a, Bis88b, Bis89, Dör90, Dör92, Dör93a, Dör93b, Dör94, Dör95a, Dör95b, Dre06, Dre07, Dre08, Goo14, Goo15, Goo16, Goo17, Goo18a, Goo18b, JGM00, KFS01, PW22, PW24, Pre09, Pre10, Pre11, Pre12, Pre13a, Pre13b, Rut96, Rut00b, Rut00c, SLP06, Sie01, Sie02, Sie03, Str93b, WP22, WP23, WP24, WM25, Add87, Ano86d, Bel93a, BW12, Bur94, Cob94, Dre95, HJ93a, Lab92, Led95, Ler96, Mor14a, PM19, Rut89b, vG92]. editors [GDS22, MBV21]. editors-in-chief [GDS22]. Eds [Art15a, Art19, Bak16, Bja20a, De 24, Cha22b, Cor22, DKL^{+}25, Edd18, Fit11, Fit15b, Fri21, Fur19, GB19, HJ12, Hes13, Hus25, Hyl24, Jon16, LT20, Law16, LX16, MDU^{+}25, Mal23a, Mal23b, MND23, MR17, Men20, Rad11, Ric20, RSD23, Rut02, Sch24, SJS12, Sri18, Sta11, Sto22, Tab22, Wal17, Wan25, Web18a, YK20, vBP19, Eng99, SMWM25, Tsa02]. educated [BJJ13, VVV13]. Education [ABK^{+}05, ACJO04, AF05, Ano72a, Ano00h, Ano04d, Ano16b, Bro23, Bus05, CF03, CGGGG74, Ern99, Fre69a, Fre78b, Fre82, GSE07, Gau69, GL00, GD99, GPW05, HMRRTT23, Kai13, Kei82, Kru13b, Kry69a, LH97, LH06, Leu01, Men07, Mor06, Nis99, Ong06, Rad12, Ser69, SK98, SK07, Web18a, Web18b, ZBEH06, Zev06, AAP21, AT21, ARW22, AMMS20, Ano69, Ano76e, Ano81b, Ano88e, AS25, Art23c, Atw24, Aud83, AH79, BRE20, BGW^{+}14, BCZ21, Bal90, BBS07, BGSW94, Bec19, Beg69, BD72, Big70, BA25b, Bis80, Bis88c, BWHM^{+}10, Bla78, Blu02, De 24, Bor21, Boy16, BP80, Bro11, Bro08a, Bro10b, BC12, BW12, Bro12, BSW16, BG69, Bul24, Bur02b, Cas69c, CSW21, CPP09, Che90, CSRlR24, Cor22, D’A91, Dar16, Del68, DKL^{+}25, DS21, Dör03a]. education [Edd18, Ehr78, El 78, El 86, Ell13, Ely10, Eng69, Ern16b, ES03, Fen79, Fer15, Fey78, Fey89, Fis69, Fit11, Fit14, Fit15b, FB17, FC08, FB11, Fre78c, Fre78d, Fre79b, Fre81, Fur07, Fur19, Fur22, GCG^{+}92, Ger81, Ger88a, GK79, Gof93, Gol03, Goo20, Gra88, Gra20, Gri83, GB17, Gui17, GB01, Gun78, HV78, Han10, Har81, HEWJ^{+}15, HTM^{+}23, HJ12, HMSBT21, How73a, How73b, How78, How80, Hoy01, HHL22, Hyl24, Jah83, Jan09a, Jan81, ISJ24, Jin07, JMD08, JV17, Jon18, KR25, Kap80, Kap78, KS69, Kei93, KFS01, Kol16, Kol18, KM21, Kol21, KSS25, KDM21, KLBMK^{+}13, Kry68, LT20, Lar86, Law24, Ler89, LXT02, Ler14, Leu21, LYBK23, Lle12, Lle16, LMH70, LQV04, Lun12, Mac00, MCW25, MDU^{+}25, Mal23a, MCVM13]. education [MVCV24, MGI23, ME13, MFG22, Men71, Men20, MA25, Mor14a, Mor14b, MHH90, MHC24, Mur95, Mwa02, Nar16, Nol12, NX23, NEM99, Ohu78, Ort24, Ott07, PV12, Pai15, PA14, PvdHPR12, PW03, PS21, PM19, Pur78, Qua92, RR17, Rev71, Rev72, Rev78, Ric20, Roo19, Roo23, RL11, Ros78, Rot12, RW19, RJWB20, Rut94, Rut10, RNM12, SC17, Say94, Sca71, SSL25, See11, Sen96, Sev22, Sfa13, Sha78, Shv22, Sil13, Sil25, Sim94, Sim17, SEC13, Sko85, Sko90, Sko94, Sko16, Sol95, Sri18, Sta11, SW92, SB12, Str87, SMOR20, SS88, Swe77, SC83, Tab22, Tal08, Tao86, Tep20, Tou90, Tre93, Tur93, Van93a, Van03b, VVV15, VD78, VD79, VCS95, WBM^{+}20, WPA^{+}23, Wan25, Wil78a, Wil12, Wil16b, Wil78b, Wit84, Wit95, Wit01, WKB97]. education [XXL21, Yoo25, Zav21, ZK21, ZK24, ZM91, dSBM21, dF12, dFS13, dKV18, dV18, vO13, vSBvR00, Ano91c, Cot24, GDS22, SK25, BGW^{+}16, Fit11, All99, De 24, Eng99, Bak16, Cha22b, Cot24, Ell18, Hus25, Law16, Mal23b, MND23, Men23, RSD23, Sri18]. Educational [Ano02c, Fit15a, IE07, RH19, VR90, Wal17, vO01, BDS^{+}17, DVC19, DGI23a, HPRW25, Hus87, Kla71, MO81, Mog17, TE14, Pea76, Bak19b, Bec19, CL18, Kae94, MW19, MBV21]. educationally [HP12]. educative [YHL^{+}15]. educator [LYHC18, YHL^{+}15]. educator-researcher [LYHC18, YHL^{+}15]. Educators [HMRRTT23, Bag08b, GCFS23, MND23, TR24, ZZ11]. Edward [Sch24]. Effect [BNPG97, PM19, SGCA23, BCE^{+}19, CMW^{+}13, CHM16, CHM83, Gan92, HPS^{+}19, KB11, Kor03, Lin90, Mar86, Sim20, SHAC01, VVV13, ZSL02, ZM25]. effecting [BCLH89, BHLH85]. Effective [HBH07, LK16, NR10, RNP13, WL90, ZBBZ85]. effectiveness [Ade98, Wai70]. Effects [KBBK20, LC15, Mos05, PC98, Rel96, SKR15, SC93, BFG84, CV16, CM20, CF00a, CSK12, Hoz81, JMRW22, KMA02, LOW24a, LOW24b, Ler14, Ma96, NK17, PM00, QM94, TA08, Tar09, Tru15, YK24, ZWC^{+}24, RSL17]. efficacy [BLC20, CPK08, CP10, KSH25, LBH23, LLZ20, SLP^{+}12, SAR19, SMS22, TTL^{+}13, VSS20, XA20, ZC19]. efforts [GDS22, HMMG18]. Efraim [Ano98b, Gre01]. eigen [WS25]. eigen-esque-ish [WS25]. eigentheory [WS25]. eigenvectors [ST10]. eight [Han89]. Eighth [CO10, LLZ20, Min83, SV09]. eighth-grade [LLZ20, SV09]. einerseits [Ott76]. Einführung [Bec77]. Elaboration [LDM92, LDM92]. electrical [Rot14]. élémentaire [AR82, FPM69, Gla76a, Roy81, Sen96, TSV82]. Elementary [Cre06, FPM69, GB17, HDM07, HBH07, LH22b, MF17, NN07, Sel01, Sty07, Tie02, Uhl02, Uhl03, War06, Ala12, And79, BS25, BH94, BMK21, Bur02a, Cha18, CG21, Cor22, Csí16, DDH25, DHD23, DRR02, Eng71, Fre74b, FSY22, Gel98, Gla76a, GFWL11, GS76, GPPT99, HOS13, HC95, IWF^{+}15, Jan09b, JM74, JC22, LS70, Law00, LK18, MR20, Nim75, Nos86, Phe10, PS21, RL13, Row08, Roy81, Sen96, Ste97, TA08, TSV82, Tie00, TG89, Tou86, TP15, ZZ11]. Elements [Bar07, Šed69, TLM07, TLM07]. Elena [Mal23b]. Eleonora [Tab22]. élève [Aud83, BC78, BC82, Cop98]. Eleven [IT77, Dem93, PD23]. eleven-year-old [Dem93, PD23]. éléves [Sen96, Bal80, DeB97, Deb06, Dem93, Her11, Kon72, KC73, Rob86, Sam85, Siw73, VD78]. Elfriede [Ste94]. elicit [ST20]. eliciting [Leh24]. Elizabeth [Fer15, HSNHHE15]. Elke [Gre17]. Elkonin [Ott76, MRC21]. Ellerton [Ste13]. Ellipses [Rut02]. elliptic [KP10]. else [NHN03]. emancipation [Rad12]. embedded [RK17]. Embodied [Abr09, BR06, LT23, NEM99, PD23, SAD21, Wil09, NF09, Shv22, SBDD24]. embodiment [Bre22, TR11, dFS12, dLT08]. Emergence [SS05a, ATAB25, AGP15, Dam87, FGG13, HMA11, HO23, PA18, Pir25, Pro13, TTS21]. emergencies [ZK24]. Emergent [CB05, Zav21, vO10, DR13, HS99, LE09, PAW15, Rad00]. emerges [Kie01]. Emerging [ABK^{+}05, SP19, BF96, BJ22, Tab22]. Emeritus [Ano16b]. Emma [Ano14a, FM14]. Emotion [EMT06, Rod06, TL15]. Emotional [BR06, ODV06, WW02, KHLP08, LM11, MSdSGG16, PCP19, See11]. Emotions [Han06, ODV06, AWE22, Han02, RW19]. Empathy [DS05]. Emphasizing [Mos05]. Empirical [LPG05, Mes04, Cha93, CPPC18, Fre72, Fre79b, GP17, HY22, IPJ25, Jan10, KM93, Kom17, MRSP14, NS25, RMAGMR22, Roh10, WSU19]. Employing [FP07]. Empowering [TR24]. empty [RBK25, VBE21]. enable [Blo03]. enablers [YST22]. enabling [LDM92, NT19]. enacted [FCN14]. enactive [BC12]. enactivist [Pro13]. enactment [PA18]. enactments [TTL^{+}17]. encounter [RFK07]. encountered [SB97]. encounters [Nil07, dF16]. Encouraging [Boa93, TT91]. Encrypted [Whi09]. enculturation [PT09]. End [VS97, De 93, Une74, Her11, Lev92, Di 19, Sen96]. ended [NCC22]. ending [TT01]. Ends [Zaz98]. enfant [Bic82, Com80]. enfants [AV76, Com77, Kra02, Mas69, Rou68, Van69b]. enfeeblement [Ham68]. Engage [STM06, CBSP12, LMCG08, REL25, VE24]. Engagement [Cal94, HWJ20, WI01, Bra14, Cho22, CSS^{+}14, Dib19, DLK14, JCM^{+}23, TN15]. Engaging [Ell13]. engineer [BW13]. engineering [BBAR23, BP80, DBL22, Win03]. engineers [GK79, Hås68]. England [Coo98, HMA13]. English [Gre17, Sto22, AH00, Bar03, Daw83, Eva07, GDS22, HA23, LR81, Lyn68, VM13, WAL18, Wha96, Zep81, Zep82, ZML87, dAS22]. English-Medium [Eva07]. enhance [QM94]. enhanced [CWH19, WPL24]. Enhancing [MPdP17, Wal17]. enigma [ARHS14]. Enjoyment [Bar21, SLP^{+}12]. enlightenment [KD10]. enough [Als02, Dho93, MYZ23]. enquêtes [Del68]. Enquiry [FA76]. enrollment [Ecc85, Wai70]. enseignants [SB97]. Enseignement [Fis69, Pis68, Ano72a, Ano00h, AV76, Aud83, BD72, Cas69c, CGGGG74, Cas77, Del68, For69, Fre69a, Fre82, Gau69, HK99, Kry68, Kry69a, Kry69b, bL00, Mar69, Mas69, Pas78, Rev68, Rev72, Roy81, Sar02, Sen96, Ser69, Siw73, VD78, VD79]. enseigner [Ser68]. ensemble [TLM07].

Enseñanza [Cas98]. Entanglements [Fer15, HSNHHE15]. enterprise [AM14]. entitlement [PAR^{+}24]. entrance [SC83]. entre [Bal88, DG89, Her11]. entry [EH20, Kae90]. Entstehung [Dam87]. Entwicklung [JSV75]. Enumeration [MDD04, Par25]. Environment [GH98, Kap01, Mar00, BCFC90, BC96, Dag96, DG89, HFL^{+}23, HSZ17, KC22, Kyn93, LDM92, MG00b, MN95, NHO22, Nuc24, PD23, TAH08, Tro00, Wal12, Whi09, YS12, HO23]. Environments [Lab00, Art97, BdSCdA18, For95, HHS00, HN96, HS99, LH18, LBFM13, PK92, TY04, Tro04]. environnement [Dag96, DG89, LDM92, Tro00]. Environnements [Tro04]. envisioning [Ort24]. epidemic [Gos21, dSBM21]. episode [Cli17, DGFT13]. Epistemic [RL13, CBC25, RP21, RC94, SHAC01]. Epistemological [DGA05, Jah01, Ste97, Ste05, Ste06, Wil97, Hei75, NBZ12, PM19, Rad03, Rot21, Sie87, WM09, WMM13]. Epistemologies [Sol06, NPH99]. Epistemology [Han98, Kid16, Ott06, BP15, Bur95, MDU^{+}25, Sch11a, Ste91]. epos [Tou90]. épreuve [Mau85]. épreuves [Dum82]. Equal [SLW98, FMA20, LBML24, SLR^{+}24]. Equality [SLW98, Kie81]. Equation [GR97, PM97, NS25]. Equations [PM97, CO10, CVN21, HK25, Hew12, HC14, KB20, LH96, NS25, Rüe13, Ves73a, Ves73b, Vla02, dLT08]. equilibrium [ATAB25]. equiprobability [Par25]. equitable [Han12b]. Equity [Fen79, Bul24, FB11, SJS12, VCD24, XXL21, ZKC18, lDlzN21, Gra20]. Equivalence [Ham06, Asg09, PB21, SK13, SSP04, TP22]. equivalences [TGWN23]. Equivalent [Ham06, NS25]. era [Cor22, Cra96, GG22a, KK10, Art15a]. Erich [Fur22]. ERME [Tab22]. Erna [HJ12]. Ernest [Goo20]. Ernst [Ste98]. Errata [Ano75b, Ano75c]. Erratum [Ano76b, Ano95d, Ano96d, BGW^{+}16, CB17a, GSH16b, Kae94, KBK^{+}17a, MLK16a, Par15a, RW11a, SMSW13a, WH15a]. erreurs [TLM07]. Error [FZ05, GCALB15, GG22b, Kar16, SGCA18]. errors [ALR25, Bro14a, CVN21, Cla83, Cle80, DVJV02, HA23, LBT07, RBK25, Son13, TLM07, Wat80, ZML87]. ESM [BA25b, Dre06, GK20, HS02, LXT02]. espace [Bal88]. esque [WS25]. Essai [Gau69, Kon72]. essay [Han12c, LCW20, PW03, Yoo25]. Essays [Pai18, O’B76]. essential [HP25, Kry88b]. essentiel [Kry88b]. Establishing [Her02a, RNM12, SHG21]. Estelle [SG25]. estimates [PK23a]. Estimation [Lev81, RRP91, Edw84, Gli91, MGV^{+}25, MH97]. étapes [Dié70]. éteinte [Add93]. Ethical [Boy16, Abt22, Che23, OCB25, RR11, Shv22]. Ethiopian [MA09]. Ethnographic [JS99]. Ethnomathematic [VS97]. ethnomathematical [Vil10]. Ethnomathematics [RC04, AAB03, Bar96, Kni12, Pai11, PF11, RC02]. ethnopoetics [Osl12]. ethnoracial [Gar24]. étude [Bal80, BZ01, BC78, Tro00, LD88]. études [Add72]. étudiantes [TLM07]. étudiants [Gro77, TLM07]. Euclidean [KV23, Kyn93, Ste71, Tou90, Vau71]. EUR [NX23, Shv22, NX23]. Europe [Law24]. European [Law24]. Evaluate [LH22b, Kra02]. evaluating [AI08, CPP09, RTSK08]. evaluation [KM21, MHC24, Ott20, Plu79, ZBBZ85, Plu79, Bui76]. Évaluation-sondage [Bui76]. evaluations [SS09]. Évaluer [Kra02]. Évelyne [Fri21]. Even [Zaz98, MYZ23, TT23]. Evens [Zaz98]. event [NR10]. Events [De 24, DH97, Woo96, AMRM24, Par25, Yah25, dF16]. ever [Kap09b]. ever-increasing [Kap09b]. Every [FM09]. Everyday [Bra13, Lev81, LW99, GS08, Ste91, TB09, VS23]. Evidence [WBM^{+}14, Cha93, IS09, MRW14, MMR21, NCMO15, VVV13, VDD^{+}03, ZSWC25, ZHT25]. Evolution [BS12, CF03, KM24a, Ott11, AP75, GAP24, GBR22, Her02a, LQV04, LMW91, MP24, Di 19, Sch11a, TT07, AP75]. Evolutionary [SK98]. Evolving [Jon00, PPG95]. exactly [Thu72]. exam [LYBK23]. exam-oriented [LYBK23]. Examination [Zen18, Ern84a, Kap74, Kom17, PXTW17, Sty11, Osi24]. examinations [Bro10a, GM84a, GM84b, Isa87, MRRF94, SC83]. Examining [AAT99, BJJ13, ÇA20, CPPC18, CPPT^{+}13, Doe06, FE09, Güç13, HB21, LG16, MM00, PM00, RWZ15, XM17, dAS22, Bra14, CSS^{+}14, Dri25, GCFS23, Ngu24, rA20]. Example [DH97, GD99, RR05, Zaz99, Bak16, Bra68, Cha98, CV16, Dié71, FVJL^{+}19, FCN14, GM08, HK99, Jah83, JC22, MG12, PY22, Van03b, YE16]. example-based [JC22, PY22]. Examples [BK91, JV17, Siu93, Ste68, Ste69a, AI08, BZ01, DWP21, GKMJ24, IIMR^{+}11, KID25, MP84, Row08, SMSW13a, SMSW13b, Sun11, TOT24, VV25, Van93a, WS08, YM16, ZZ08]. excellence [Fra96]. exceptional [AM25]. exchange [Wil12]. exclusion [JGR14]. exemplary [CDW14, ZC08]. exemple [Bra68, HK99]. exemples [BZ01]. Exemplified [Ong06]. Exemplifying [ZL08]. exercises [Fos18, GS72, Lit03, Ste68, Ste69a, SLW83]. exhibit [HWJ20]. exist [How80]. existence [SPC^{+}16]. existential [AT21]. existing [VV25]. exit [EH20]. Expanding [VDD^{+}03]. expectations [Han02, JGF20, SLP^{+}12, SPFGF14, WZC21]. Experience [Cal94, AZ20, BVR24, Bro94, Bur87, Dör03b, Gli91, Hah14, MRC21, OB11, Pro70, RT09, SECP^{+}23, Vyš75, WB12, Cal94, Dou76, Pro70, Rou68]. experienced [Gro77, Hew12, YHL^{+}15]. Experiences [BCFF^{+}98, Bek10, DGI23a, Dör03b, FF25, FB11, JMRW22, MSdSGG16, MDPW96, PK23c, Shv22]. Experiencing [Asg09, KMT16]. Experiment [AP07, PBY03, Sel98, CSK12, Dou76, Hah14, Mat83, Moo13, Pro69, Rou68, Sha77, Smi70, ZM25]. Experimental [FZ05, Orl71, Fre77a, PRR08, RT14, Vyš75]. experimental-mathematics [PRR08]. Expérimentation [For69, For69]. expérimentés [PRR08]. experimenting [KMT16]. Experiments [Jak71, Wal73b, AP75, Bel93c, Fre77b, Jee71]. Expert [CLK22, BKK22, GLZBZ15, Hop85, KM21, Mar13, Rüe13, Sty11]. expert-novice [Rüe13]. expertise [BKK22, BR17, PXTW17, Hes13]. explain [EJ13, EPQH18, LLW07, TS99, VH16]. Explaining [Eli85, Kae93, KLBMK^{+}13, AS17, EPQH18, Her93, Kae94]. Explanation [Han00, ARHS14, EMRI22, FP85, LCW20]. Explanations [Jon00, Bel76, IAP19, Lev10, Lev13, RA21, SV09]. Explicating [Sim17, HP25]. Explicit [Ryv06, Güç16, MB16, TG89]. explicitly [IAP19]. exploitation [Bra68]. Exploration [Han00, CCP23, CB17a, CB17b, Hic24, LSS19, Loc11, SK25, Vog13]. Explorations [Llo06, HMG19, HMG20, Mak14]. Explorative [Reu23, BF21, CCP23, HMSBR19, Nil07]. Exploratory [RG19, Buc87, HWJ20, MvdKM13, Rut89a, SM72, VN19, WMR11, RG20]. Explore [Mos04, DS21]. Exploring [Boa02, Bye15, Cho22, Coo98, Cre03, Gat06, Gor98, KSL11, LLW07, Lev13, LBT22, Lev22, MCPGE24, OB06, PG20, Par25, REL25, RA21, RSD23, TTL^{+}13, WS18, YHL^{+}15, ZC06, Bre22, CY23, HPS^{+}19, Jou25, KMT16, Lam15, RK17, MDU^{+}25, Sto22]. Exponential [CS94]. exponentiation [Vin77]. exponents [MZ21]. expose [Kap09b]. Expository [SG98a, SS15]. expression [SG08]. Expressions [TER98, Hod96, PG20, Rüe13, SK13, TP22]. extended [RNRVM23]. extending [VV25]. Extension [LW99, Gal02, PPDS^{+}25]. external [Kae93, Kae94]. extinguished [Add93]. Extrema [TOT24]. extrinsic [WS18]. Eye [CSDK25, SKE25, SMOR20, ATAB25, AE08, BA25b, BBDD25, BBVD25, BOL24, HK25, MMM25, PPDS^{+}25, RIA17, SL19, SSL25, SV25, ZSWC25]. eye-mind [SL19]. eye-movement [HK25, RIA17]. Eye-tracking [SMOR20, BA25b, BBDD25, BBVD25, BOL24, PPDS^{+}25, SV25]. eyes [HH11].

F [Bur02a, DRR02, LH22b, Web18a, Web18b]. F-PosE [LH22b]. Fabrication [Sen96]. face [AAP21, Gos21, ZP91]. face-to-face [AAP21]. Facilitating [AAT99, DH97, PK23a, LYHC18, MPdP17]. facilitation [TS18]. facilitator [TLB^{+}24]. facilitators [SECP^{+}23]. Facility [WE95]. facing [Ale81]. fact [Fre69d]. factor [Ade90, Cas04, NT75]. Factors [Buc87, FNM91, GC90b, Gli91, HBZE98, LW14, Phe10, PSE87, AV08, FM09, Gal86, HB21, KČ12, LCD90, LBH23, MK91, Nkh02, PK23a, PCP19, RH19, SM22b, NHN03]. facts [Hof77, Tho90]. Faggiano [Tab22]. Failed [OM24]. Failure [HM25, Bur87, CL84, Dib19, Kon72, SBDM69]. fair [Gra20, SVVO23]. fair-sharing [SVVO23]. fall [Bja20a, Bja20b]. familiar [Kon18]. family [KLLJ17]. Fan [DKL^{+}25, ZB19]. fantasy [Edd18]. far [EJ13]. Farfán [LQV04]. Faso [TB09]. favorable [LDM92]. fear [Led80]. Features [ST20, Gal86]. feedback [HKG89, Kon21, Kon22, PC23, SC93, SKN21]. feel [KL20]. Feeling [WD10]. Felix [Ano04c, Ano06b, Art19, Ano16b, Art23a]. female [HS13a, Mal93, Sol12, SRB16]. feminine [SRB16]. feminist [Bur95]. Fermi [AG14, SF23]. Fiabilité [Fre82]. Fiber [Ham06]. Fictional [Llo06, ZK15]. fidelity [SHAC01]. Field [ABK^{+}05, BA25b, MCW25, Nol12, SL19, SB20, WPA^{+}23, Yoo25, DKL^{+}25]. fields [BB03, Eng69, HK25, MR17]. fieldwork [CPK08, RPP24]. fieldwork-based [RPP24]. Fifteenth [Ano04d]. Fifth [Eng97, Lev10, ČK18, GT90, MLL19, Yan19]. Fifth-Grade [Eng97, Lev10, ČK18, Yan19]. fighting [Høj24]. figural [BA25a, EJ13, Fis93, El 18, Wil21]. figurale [Ric04]. figurative [Ric04]. Figure [MM25, MCPGE24, MB19]. figured [GLHM24]. Figures [DVJ98, Par88]. Filling [LW99, MPS22, RBK25, SMW11]. fin [Une74, Her11, Lev92, Sen96]. Final [Ano68a]. Finding [Cla02, Sol12, Sta89, Ell18, GS72]. Findings [Jun91, Had78, HM95, Nar16, NH94, TTL^{+}17]. Fine [Tzu07]. finger [KBBK20]. Finite [Ste76, Tsa99, GR71, Šed69]. Fink [Mur24]. Finnish [ARHS14, KPH10, TL15]. First [GL98, AMRM24, Ano14a, BZ01, Bic82, BBG^{+}17, Bui76, Cas69c, Cas69b, Com80, Eng12, FKIA15, Ger84, Gla76b, GB14, JTP94, KBBK20, KBK^{+}24b, Pap71, PPDS^{+}25, RSRVRVVR23, Rev69, Ric10, Rob90, Tru15, VYZ21, Fre69a, Fre78b, MGD^{+}23]. first-grade [BBG^{+}17, Eng12, VYZ21]. first-year [BZ01, RSRVRVVR23]. Fischbein [Ano98b, Gre01]. fit [CN16]. Five [Ano78b, How84, VR90]. Flanders [DTD21]. flat [DPG89]. flesh [TR11]. flexibility [Lee17, SF23]. Flexible [Thr02]. flow [GD17, ML20, CGGGG74]. fluency [Fos18, TD12]. flux [CGGGG74]. Focal [NR10]. focus [BOL24, Dör03b, EH20, Gue19]. focused [Zah15]. Focusing [Van03a, Dra14, HVN^{+}24, LHS19, SWVO23, TMH21]. folding [HFL^{+}23]. follow [CWYP92, Wea90]. follow-up [CWYP92, Wea90]. following [Bag10, WAL18]. fonction [Bic82]. fonctionnement [Sam85]. Fonctions [Bos87]. fondamentaux [Mas69]. footsteps [HJ12]. force [Fre93, Vre69]. Forced [Sko04]. forces [MA25]. foreits [Lie90]. foreshadowing [Was22]. Foreword [Ano76c, Web18a, Web18b]. forge [KL14]. Forged [Dib19]. forging [MK14]. form [BFA22, BR12, BVR24, BE84, LC13, Nar16, Rob75, WS25, WHC09]. form-function [WS25]. Formal [Bar16, MD01, RC04, SZ07, Tal01, AAB03, Cha98, Dié71, GBTL23, Hew12, KT01, Led92, Moo94, RC02, WSR11, Siw73]. Formalization [Wit73b, Wit73a]. format [CV16]. formation [BS80, DS02, Del69a, Hah99, Has97, HK99, Høy07, Sem84]. Formatively [SGB21]. formed [Prz04]. formelle [Siw73]. former [Bak19b]. Forms [Ern99, Gra99, Nte06, Tir99, Blo03, HP08, KN21, RP07, Rut90, RDH09, vO01]. formula [SBA22]. formulae [Sch11b]. formulas [KJvD20]. formulated [Bal80]. formulées [Bal80]. forum [vdS11]. forward [Wan25]. foster [BF21, Güç16, Yah25]. Fostering [DBL22, Kom17, SA07, CKS22, FV22, GGD22, GABB25, WZ86]. Foucault [Kol16]. Foundation [Jon00, MDU^{+}25, Ste71, lRA16, vBP19]. Foundations [Cor22, Bos93, Cha18, Fle76]. founder [Bec19]. Four [ZM91, AKL80, Ano78b, LH09, O’B75, SWVO23]. four-card [AKL80]. Fourth [AAM21, GT90, Hyl24, Sem70, Hus25]. Fourth-grader [AAM21]. fraction [CRPPRG16, CR09, CGD21, CG21, CBSP12, Hun86, Lee17, O’C01, PB21, PH96, SL95, Sen96, Str78, Sun11, Wea90]. fractional [ES21, Rüe13, TP15]. Fractions [Bri06, CPP07, EJDT06, XM17, Ala12, ČK18, GP17, HL16, Has81, KT01, KZ94, MLL19, PP12, PGM13, PM16, SS10, Str82, Yim10]. Fragile [SC90].

Frame [McN00, Sch23, Sch25]. frames [AM14, BC12, Cha22a, HMMG18, KL14]. Framework [BC05, CB05, Ong06, PBR^{+}04, Ryv06, Ant96, BE23, Büs24, CSS^{+}14, DF15, DB86a, Dra14, EH20, Ell13, GB23, GGD22, Hay87, HWJ20, HK99, JN15, Jaw03, JLTM97, Kol16, Kol18, Kom16, Lit08, MDPW96, MYM12, Nos86, Reu23, RL12, Ron15, SLP^{+}24, SV13, Sut89, TTS21, Tha10, VRK24, VV25, VA18, WqWLR18, LH22b]. frameworks [Bec24, Cli17, Kid08, TMH21, Van84]. France [Art23b, Han20, Rev78]. Frank [Ric20, EH06]. Frederick [Jon16]. free [HMT13, Mar72]. Freitas [Fer15, HSNHHE15]. French [Add72, Add75, Add76, Add82, Add93, Ano72a, Ano00h, AP75, AM97, AL00, AV76, AR82, Art97, AG02, Aud83, Bal80, Bal87, Bal88, BC92, BZ01, BD72, BC78, BC82, Bic82, Bos87, Bra68, Bui76, Cal94, CB89, Cas69a, Cas69c, Cas69b, CGGGG74, CDGG76, Cas77, CGGGG79a, CGGGG79b, Cha71, CGS76, Com77, Com80, Cop98, CLR16, Dag96, De 94, DeB97, DS02, Deb06, Del68, Del69a, Dem93, Dié70, Dou76, DPG89, Dub84, Dum82, DG89, DP77, Duv83, Duv91, Une73, Une74, Fis69, FPM69, FBM71, For69, Fre69a, Fre69d, Fre82, Gau69, Gla68, Gla76b, Gla76a, GT95, Gro77, Hah99, Her11, Hol69, HK99, Kon72, Kra02, Kry68, Kry69a, KC73, Kry69b, Kry88b, Kry88a, Lab78, bL00, LD88, LDM92, Lev92, LS79, Mar69]. French [Mas69, Mau85, Mes89, Miy17, Mot71, Nim77, PRR08, Pas78, Pis68, Plu79, Pro70, Rev68, Rev69, Rev72, Rev76, Ric04, Rob95, RR05, Rob86, Rom74, Rou68, Rou69, Roy81, Sam85, Sar02, SB97, Sch92, Sen96, Ser68, Ser69, Ser76, Siw73, Str08, SLM79, TSV82, TLM07, TB09, Tro00, Tro04, Van69b, VD78, VD79, ZP91]. frequencie [Fre72]. frequencies [RBK25]. frequency [PK23a]. freshman [Gof77]. Freudenthal [Add93, Ano04c, Ano06b, Ano16c, Art23b, Gre17, Van93b, Add93, Ano90d, Bec19, Bos93, Kei93, Pea76, Str93b, Tre93, Van93a]. Fried [Fit15b]. frontiers [Ely10]. fruitful [Tab22]. fulfill [BF21]. Fulvia [De 24, Men20]. Fun [Moy01]. Function [ASP97, Sla97, Bic82, BDHN92, Car18, Güç16, Lun12, MZ18, MPG12, MAPHB21, PSH^{+}18, Par15b, Par15a, Prz04, Ron15, Saj03, San11, TN15, TK14, WS25, Yer00, ZER17]. functional [AAM21, AWE22, Ber04, BBG^{+}17, DDH25, DHD23, SAD21, Vol86, Wil16a, YG99]. Functionality [Gor98]. functioning [Sam85]. Functions [AWL15, BBEG05, Ham06, Mes04, Mos04, Blo03, BC96, CS94, Cor17, Eve90, FLM07, FBA10, GK71, Har09a, Kar02, Kol18, MYZ23, SD95, Sho21, TMP10, UMC24, WAL18, WT14, Whi09, Bos87]. Fundamental [DDLZ97, II05, Bur96, DDLZ94, FSY22, Gri13, Hah75, Hei75, Law00, Mas69, MW09, Tho94b, Wit80]. Furinghetti [De 24, Men20]. Further [Bur06, PSS^{+}19, Del69b, Gen70, IS09, MCW25, MRW14, Tha10]. fuss [Sfa09]. Future [BCZ21, Bor21, CL20, Ell18, Fle71, GC90b, GCG^{+}92, How75, KDM21, LS20, LK14b, SB97, Str03, SS17]. future-oriented [LK14b]. futures [MCW25]. futuristic [Hyl24]. futurologist [Tal08]. futurs [SB97]. fuzziness [MZ18].

G [Art19, Mor00, Sfa15, vBP19]. Gabriel [Sch24, Gue19]. Gabriele [Hes13]. Gagné’s [DŽC25]. gains [NCMO15, SL94]. Game [Har09b, SE93, Gal73, KML22, LS79, Nil07, vdHPKR13]. Games [Bas69, SL06, Kni12, SLAY19, SK69]. gaming [Ern86]. Gap [SZ07, ABvP02, FF21, HL94, LLW07, LH96, LW99, MPS22, MZ21, MMR21, ZWMR16]. gap-filling [MPS22]. gaps [BDS^{+}17]. Garrison [Eng99]. Gaspard [Fri21]. gaucher [Tro00]. gears [BBFG99]. Geert [Bja20a, Bja20b]. Gemert [Gre17]. Gender [BHX97, Che89, For95, HH88, LLHL13, Rel96, Say94, TZLT95, Tie00, Tie02, VL04, ZKC18, Ade98, AC95, Jun91, KM93, MMR21, SC17, TF95]. Gender-related [Tie00]. gendering [FLT14]. general [BD72, Bre22, FC08, LBT07, MPG12, MP84, BD72]. generalisation [FS16, FF17, MW00]. Generalising [WC08, Sta89]. Generalization [ZL02, BA25a, EJ13, JE14, MPdP17, El 18, PBCPP25, Riv10]. generalizations [ČK15, TAH08, ZM25]. Generalizing [WH24, BIBS^{+}19, lD95, Wil16a]. generate [Nar16]. Generated [Ber06, Bur05, Ott20, WS08]. Generating [KJ22, SMSW13b, SLC95, HKG89, IIMR^{+}11, KID25, Shr10, SMSW13a]. Generation [DH97]. generations [QP22]. generative [RL04]. Generators [BC05]. Generic [MP84, DWP21, YE16]. genesis [SAD21, vDDDB21]. Genetic [Sch83, Zas71, FP07, MDU^{+}25]. genre [Bul24, PW03]. geoboard [Har74, Har75]. GeoGebra [Zen18]. Geology [Wit05]. géomètre [AG02]. Geometric [Lev71, MMBD05, Nar00, BMPT24, Car10, CSDK25, FQL^{+}17, HKG89, KR11, KN21, McD89, VO17, WR96, Yah25, Yan11]. Geometrical [Eng71, Kyn93, TMP10, CH13, GP90, Ger88a, GG22b, GB01, HY22, KSL11, LL13, MCPGE24, Mit80, MB19, PPPP14]. géometrie [Aud83, CDGG76, CLR16, Hol69, HK99, Mes89]. geometries [KV23, Ste76]. Geometrieunterricht [Wal76]. Geometry [Ano71b, Ano74b, Bar07, BH11, Fre71, HHS00, HO23, Jon00, Lab00, LR81, Ano71a, AG02, Aud83, BS80, Bju04, Bos93, Bra11, CDGG76, Cha93, Chi98, CLR16, Cox71, DH20, Egs70, Ern86, Fie79, Fle71, Fle76, Fri21, GL10, Gal11, Gal19, Gla70, Goo07, HKSS^{+}25, Her02a, Hol69, HK99, Hoz81, KP10, Kel71, Kle71, Kom17, KR11, LG13, LBFM13, LL13, MG00b, MM25, Men71, Mes89, Miy17, NHO22, Par88, PPC09, PC10, PD23, Pro69, Rev71, Rob12, Rog71, Roo71, Rot14, SPC^{+}16, SL19, Sch71, ST08, Sem70, SC90, Sin03, SLAY19, Ste71, Sto71, TY04, Tou90, Tri95, Vau71, Vil71, Vol76, Wal73a, Wal76, WH15b, YL08, Yan12, WH15a]. Georg [Tab22]. German [Bec77, Beh68, Bra13, Dam87, JSV75, Sel01, Van76, Wal76]. Germany [Ano74b, ARKK23, DTD21, Han20, KCS22, TLB^{+}24, YSL22, LT20]. Gert [Law16]. Gestalt [Sha98]. Gestion [DG89]. gesture [YTD11, dFS12]. Gestures [APRS09, Edw09, ART15b, CH13, CSRlR24, FS16, HH11, Jou25, LLD14, MB09, Ng16, Rad09, Sfa09, SLR^{+}24, WFCW15, YTD11]. Gesturing [ZW11]. Getting [SWG05]. Ghislaine [Cha22b]. Ghosts [Har87, Vog70]. Giacardi [De 24]. gifted [ÇA20, Lar86, SOC86, Tao86, WZ86, rA20]. giftedness [Pre86]. Gilah [Art23a]. Giorgio [D’A10]. Giorgiutti [Rut02]. girl [FSB18, PSE87]. Girls [JGF20, CWT12, Fen79, FPCL90, Han89, KAJHB10, Led80]. Give [NN07]. given [CMdR18]. giving [MM20, Van77]. Glasersfeld [Ste98]. Glenda [LX16]. glimpses [SMWM25]. Global [GB19, Hud90, Als02, BZA01, ETC14, Kun12, MVCV24, MA25, SV25, TE14]. go [BSV95, BA23]. goal [Jan10, PA18]. Goals [Han06, MHDC04, Kei82, LE09]. Godfrey [How73a, How73b]. Godoy [LT20]. goes [Ano16b, Ano16c, Art23a, Art23b]. Going [Gat06]. gone [Her02b]. Gons [Bac71]. Good [Bat13, Rad11, Sfa05, BW13, MS20]. goodness [Kno77]. Gosia [Wan25]. government [Baz88]. Governmental [Lar86]. gp [Cor22, NX23]. Grade [Alb00, AF05, BCFF^{+}98, CRL06, Eng97, GL98, Nte06, PBY03, RR05, Sel98, TP22, ZS07, BS25, BBG^{+}17, ČK18, CO10, Che89, CWYP92, CGS76, CBSP12, CSK12, Une74, Eng12, FV13, FG17, For95, For69, Han89, Has97, JHC14, JTP94, JNF25, JE14, KBK^{+}24b, LH71, LHS19, Lev10, LC15, LLZ20, MLL19, MS20, Min83, Ott20, RRP91, Rog71, SV25, Sem70, Sin00, SV09, TD12, VD78, VD79, VYZ21, VBE21, Yan19, DŽC25]. Grader [OK05, SLW98, AAM21, OMB20]. graders [GT90, PPDS^{+}25]. Grades [Rut05, WN04, Fre74b, Jur77, Pet91, TF95, Tch11, Var72, Wal73b, vBP19]. gradient [MAPHB21]. Graduate [Car99, DS05, RA21]. graduates [Gal82]. graduating [KLLJ17]. Grail [Lle12]. grain [Tzu07]. grammar [TRDA20]. grand [Jou13]. Graph [Nim75, FBA10, JDK^{+}24, Kar16, RT14, WWFC16, YS12]. Graphen [Wal76]. graphes [Bal80, CGGGG74]. graphic [Blo03, GT00, MK97, Ott20, Ric04, Rut90]. Graphical [ASP97, Ola14, Riv07, Jon18, SBL22, TP22, ZWMR16, Cal94]. graphics [Bro10a, FT03]. Graphing [DZ00, KJvD20, QM94, RDH09, SBL22]. graphique [Ric04]. graphiques [Cal94]. Graphs [CGGGG74, KHY21, Abe78, Bal80, Moo20, NBHK07, RNRVM23, RL04, Šed69, Wal76, WT14]. grasp [TO13]. grasping [Bre22]. Gravemeijer [HJ12]. gravitation [CGGGG79a]. great [Law16, How78]. greater [MBF25, Tab13]. Greek [Sid05, SK07, Tou90]. Greer [Fit11, Sel00]. Griffiths [JMD08]. ground [Fur22, Fit15b]. Grounded [YTD11]. Grounding [Baz01, NEM99, WD10]. Group [Dah04, DDLZ97, Fre78b, JS99, MEH20, Nar00, Bal90, BJ88, Bur96, CY23, Dek87, DDLZ94, GGR02, HKSS^{+}25, Hah75, Hoy85, IT77, LHZ95, MGP75, MGD^{+}23, MLL20, O’C01, RNP13, Sha77, SM21, TS18, Tru15, WMPL08]. grouping [OAS25, PAR^{+}24, Wit73a, Wit73b]. groupings [Wit75]. Groups [Cre06, GPW05, LBML24, ZWC^{+}24]. groupwork [EH20]. Growing [KH80, MT15, Rot16, Wil21, Sto22]. Growing-making [Rot16]. Growth [ART15b, BIBS^{+}19, PK94, Ron15, BKK22, LYHC18, WC08, Win11]. Guerrier [Mal23b]. guessing [Kom16]. Guest [Add87, Ano86d, Ano99b, Bel93a, Bur94, Cob94, Dre95, HJ93a, Han98, JGM00, KFS01, Lab92, Led95, Ler96, Mor14a, Rut89b, SLP06, vG92, BW12]. Gueudet [Cha22b]. Guidance [GPW05, HFL^{+}23, ZBBZ85]. Guide [AAT99, Eng06, Ahm23, HRW25]. Guided [IM19, CCP23, SHG21]. Guiding [DC17, Kol16]. Guinea [Sfa15, Kae93, Sou83, Atw24, Kae94]. Guinean [Cla83].

H [Art19, Mor00]. H. [Add93]. H.-G [Art19]. Habermas [Büs24, UB23]. habitus [Zev11]. Hackenberg [MDU^{+}25]. Halai [MR17]. hand [ATAB25, Hed85, KJvD20]. hand-held [Hed85]. handbook [Hus25, Law16]. handed [Tro00]. handicapped [Bar88]. handle [Nil07]. Handling [BG05, OVKR25]. Hands [TTS21]. Hans [Ano04c, Ano06b, Gre17, Tab22, Ano90d, Ano16c, Art23b, Bec19, Kei93, Van93b]. Hans-Georg [Tab22]. happens [JMA19]. Hardback [Sfa15]. Hardcover [Ahm23, Cor22, Mah21a, NX23, Shv22]. harmful [Abt22]. harmonious [Fur19]. Harsh [dF12]. hater [AVM15]. Hating [Lun12]. Hauke [Sri18]. Having [MT15, Moy01]. Hawai’i [WL90]. Head [TTS21, Lan86a]. Health [ZK21]. Heart [TTS21]. Heather [Bro11, Mur24]. Hedges [Row95]. Heidegger [Bag10]. held [Ano74b, Hed85, Mur93, Mur95]. Help [HPG05, NN07, Gos21, GSH16a, GSH16b, HSM22, JI15, PG93, Ram03, vdS11]. helpful [vLXD18]. Helping [Mur24]. Hemispheric [Fid85]. her [FC12, FM09]. here [GLHM24, SRB16]. Hermeneutics [Bro91]. Herscovics [Ber94]. Hersh [Pai18]. Hershkowitz [Rad11]. hétérogénéités [Sar02]. Heterogeneities [Sar02]. heterogeneous [LOW24a, LOW24b]. Heuristic [Hug74, Bal71, CM92, Fle68]. Heuristics [FD24]. HF [Gof93]. hiatus [FB17]. Hidden [Bau80, ERS19, Hus87, ISJ24, SBA22]. Hiebert [Ahm23]. Hierarchies [Har81]. High [BBEG05, Cha93, Cla07, Dah04, Jur06, Lev71, NHO22, AI12, BZA01, Bra14, DH20, FJC95, For69, Gar24, Goo07, Gut92, HS13a, HWJ20, HMMG18, Hil71, ICR20, JGF20, JNF25, Kar02, Kel71, Kim20, Kle71, Kra93, Ma96, MLL20, MRC18, Par91, Pre92b, Pro70, RA21, SGRS24, SM22b, VS13, ZER17, dKV18]. high-level [JGF20]. high-school [FJC95, MRC18]. higher [BWHM^{+}10, Jin07, MHC24, SC83]. highly [Bon14, GDS22, KS69]. Hindrance [Vla02]. hint [PG93]. histograms [BBDD25, BBVD25]. Historical [BG06, DWMB05b, DWMB05a, Hyl24, RP07, AI07, BP15, Cle79, Ell18, GB01, MAHK08, PGM13, Swe95, TT07]. Historically [Ber06, Bur05]. History [Cla12, Fri07, Gui17, KB07, Kle01, Sid05, Web18a, Web18b, AOJ22, CPP09, FP07, FvM97, Fur07, Fur19, GS20, Gos21, GB01, Jan09a, Jan10, KB12, Law16, LC15, LLL21, Men20, Men23, Ott07, Sch11a, Ste13, Str03, WqWLR18, ZK21, dV18]. Hochmuth [Mal23b]. Hoff [Fur19]. Hohensee [Ahm23]. holes [Mie23]. holistic [RW19]. Hollebrands [MND23]. Holton [Jon16]. Holy [Lle12]. home [DMHRV21, HHL22, ZC19]. homework [Lan13, LM11, Ma96]. Homo [Add93]. homogeneous [NT01]. homomorphism [RS22]. Hong [Che88, Che89, Leu95, LNC23, YSL22]. Hoong [YK20]. hope [NBA23]. horizon [Sch16]. horses [Gri87]. hospital [WSU19]. Houdebine [Rut02]. Howe [Cor22]. hows [Jan09a]. Hoyles [Sta11, Gai11]. HSLS [SM22b]. https [Ahm23, Cor22, NX23, Sfa15]. Hua [vBP19]. Huang [Mur24]. Hui [Siu93]. Human [Win11, Cra96, dF25]. Humanities [Sie87]. Humanizing [Pai18]. humans [Bor21, Bye15]. humans-with-media [Bor21]. humans-with-non-living-things [Bor21]. hundred [Ano00h, De 24]. Hungarian [AH00, CSK12]. Hungary [Ano88e, HV78]. Husserlian [ZS15]. Hwang [Ahm23, MPS12]. hybridity [SRB16]. Hybrids [CCP23]. hypothese [Boa02]. hypothesis [SL19].

I.G.P.M.E [Fre78b]. Ibero [RSD23, RSD23]. Ichishkíin [RJWB20]. ICME [Bal90]. ICME-6 [Bal90]. ICMI [Ano00d, Ano04c, Ano04d, Ano14b, Art05, GL00, HbL05, vBP19, Ano69, Ano98i, Ano06b, Ano14a, Art02, Art05, Art23c, BT05, Blu02, De 24, FvM97]. IDeA [Kru13b, GT95]. ideal [Kla71]. Ideas [Ehr78, El 78, Fey78, Fre78d, Gun78, Mon01, Ohu78, Pur78, Ram03, Ros78, Wil78a, Wil78b, Bla78, Egs70, Fra13, Fre78c, HV78, Hei75, How78, Kap78, MGI23, Mas69, Rev78, Sha78, VAM21, Wit80, Sfa15]. idées [Mas69]. Identification [WZ86]. identifications [dF12]. identify [EJ13, MWVE09]. identifying [GRR23b, LBH23, Reu23]. identities [BWHM^{+}10, BW13, Bro11, GB23, HTM18, HRW25, KPS17, RL11]. Identity [Dar16, Leu01, AVM15, AM25, BWHM^{+}10, ETC14, HP08, HM13, HMA13, LK14b, LK18, PM19, RBWS18, SJS12, Tak18, dF12]. ideological [Bur90, Lun12]. ideology [Lou18, Pai13, Roo19]. idiosyncratic [Tho94a]. IEA [Baz90, Fre75, Pea76, Rob90]. If [Par15a, Abt22, Par15b]. ignored [PvdHPR12]. II [Baz90, DB86b, HH78, Noe80b]. III [Had78]. I’ll [Hic24]. illuminate [HTM18, RDH10]. illuminating [Cla12]. Illusion [VDD^{+}03]. illusions [Rad12]. Illustration [SL06]. im [Dam87, Van76, GLHM24]. Image [Fis77, MT15, BM08, KD14, MRCR11, Sem08, TV81]. Imagery [ASP97, LC81]. Images [Mur93, Mur95, Nar00, Prz04, Tho94b, ETC14, Jon18, Mar13, PB00, Roh08, TT23, TOT24, WSR11, WT14]. Imagination [NBD^{+}04, KC09, NF09, Pre92a]. imaginative [Pre92b]. imagineer [Tal08]. imagining [dF25]. imitative [Lit08]. immigrant [GdA09, Tak18, VCD24]. IMO [Ano77c, Ano77b, Ano78d]. Impact [Bus05, FKIA15, Mal93, Pal08, Cla85, FV13, FB11, IM19, ICR20, Lam15, MHC24, SS93a, VDWV04, WSU19]. implement [MCKP19]. implementation [GGD22, Gue19, PFSS19, Zaz99]. Implementing [MG00a, CP10, HN08, kSL13]. Implication [DS02, GT95, Dum82, DG03, HK02, Vre69]. Implications [ES03, GC90b, KB07, O’B77, Sim15, Abr01, Har09b, Kla71, PF11, Sin03, SV13, TK14, Usi74, dA00]. Implicit [TPHS07, BNPG97, BMP20, Güç16]. impliquées [BD72]. importance [Bat13, BDS^{+}17, GBMB70, Han12d, IMRS07, JWFC^{+}25, KWF15, MS99]. important [EPQH18, Fey89, ZB19]. imported [HEWJ^{+}15, VM13]. impose [Rüe13]. imposing [RR23]. Improper [DVJV02]. Improve [ACJO04, FQL^{+}17, FG84, Kra02, RH19]. Improvement [Van94, Wal17, Beg69]. Improvements [KBK^{+}24b]. Improving [KML22, WD05, Jak71, KCS22, PK23c, Gai11, SJS12]. in-action [OB11]. in-depth [DVJV02, HMA13]. In-service [Gut92, BKK22, Gra04]. in/for [Kle02]. in/of [dFS13]. inaccuracy [Pal20]. Inappropriately [VVV15]. incarnate [BR12]. incident [Tur76]. incidents [RPP24]. Include [LW99, AMRM24, NBE^{+}12]. Inclusion [Roo19, GABB25, Roo23]. inclusive [Sol95, LT20]. incompatible [KZ16]. incompréhension [Add75]. inconsistent [RDH10, dKBvdS22]. inconvenient [Kle12]. incorporation [Mat03]. incorrect [OVKR25]. Increasing [Kom16, AS17, Kap09b]. Independence [Tay91, Mau85]. Independent [DS05]. India [BS25, Cho22, Kap78]. Indian [PF11]. indicating [EE21]. indicative [Kno77]. indicators [Vol86]. indices [Gro77, Sch11b]. Indies [Wil78b]. Indigenous [ME13, MFG22, Men07, AT21, Lan81, RJWB20, HH24a, HH24b]. indirect [Ler85, PvdHPR12]. Individual [Kru13b, MLL20, GP17, RWZ15, DP77]. individualistic [LYBK23]. Individualization [Gau69]. individualized [KS69]. individuelles [DP77]. Indonesia [Sia24]. Indonesian [SS88]. induce [CPP09, Sfa15]. Induction [Chr69, AS25, AH76, AL78, Ern84b]. Inductive [Rot21, Kyn93, MP14, Sim96, SK69, SLM79]. inductrices [SLM79]. industrial [Hyl24, Kla71]. industry [Pol76, Zev11]. ineffective [Gal19]. Inequalities [Riv07, AI12, QP22, SBP11]. inequality [Moo20]. Inference [WM98, And09, IS08, IS09, Jur77, Ric04, ST02, dVSKvO18, Ric04]. Inferences [Dri25, MMP15]. Inferential [Wat01, Wat02, GBTL23, Mak14]. inferentialism [Dri25]. inferred [FYAE12]. Infinite [Tsa99, MPGDA12, RW11a, RW11b, Tal80, Tsa01]. Infinitely [Kle01]. infinitesimal [KK10, TK14]. Infinities [Jah01, Tal01]. infinitude [ZK15]. Infinity [DWMB05b, DWMB05a, Fis01, Mon01, TT01, FTH79, KT15, KČ12, LMW91, Tal80]. inflection [TO13]. Influence [CBF07, Deb06, Gat06, GCALB15, Mau85, Cha18, Dum82, DBL22, HBZE98, Jin07, Kae90, LBH23, Mor21, Nos88, Rut90, Wat02, Wha96, Deb06, Mau85, Sch24]. Influences [Car99, GC00, Hyl24, Lin88]. influencing [Cas04, Jaw03, KČ12]. Informal [ABvP02, Wat00, Way94, Bar16, GBTL23, KC22, Mai85, Mak14, MMP15, MO94, O’B76, UOF25, Van03a, dVSKvO18]. informatics [BF96]. Information [Art02, SOC86, AC21, BCLH89, Cra96, HY22, Lin90, Mal23a]. informatique [Dag96, Sam85]. informatiques [Art97, bL00]. informatisés [Tro04]. informed [AAB03, CWH19, RC02, RC04]. infrastructure [Kap09b]. infrastructures [HP08]. Inhelder [Kap74]. inherited [YK24]. initial [Hod96, PH96]. initiated [APS22, WD14]. initiating [Mey10]. Initiation [EH20, FPM69, Gla68, Gla69, AM97, FBM71, Hol69, vO01, FPM69]. Initiation-entry-focus-exit [EH20]. initiative [KZH^{+}19]. inner [SG08]. inner-city [SG08]. Innocence [VS97]. Innovation [BEW99, Ern84a, vO13].

Innovations [FV22]. innovative [HPS^{+}19, MCKP19]. innumeracy [Tre91]. Inquiry [Adl97, DŽC25, Tou71, Zaz98, CP10, Elb03, GFJ13, Han12a, HMRRTT23, SLAY19]. Inquiry-based [DŽC25, GFJ13]. inserted [BF96]. inservice [MCL25, SS93a]. Inside-Out [Alb00]. Insider [LH22a]. Insight [Mac78, Gri71]. Insights [GT90, JC22, KZH^{+}19, Wal17, Woo96, YM16, CPP09, CPPC18, DKL^{+}25, Gue23, HJ12, MS14, WWFC16]. Inspired [Ber06, Bur05]. inspiring [KMT16]. instance [Mos98]. instances [SL74]. instantané [Sch92]. instantaneous [Sch92]. Institute [Van93a]. Institutional [Pop88, RG01, AV08, Har09a]. Institutions [Cas04, SC83]. Instruction [Ano04c, Art05, De 24, Ano78b, Bat13, Beg71, BN91, BR89, CHM83, Chr69, CR09, CBSP12, Del69b, FKIA15, Gel98, GBE13, GS72, GS76, HKSS^{+}25, Han12b, Han12d, Hes13, HMNWB20, HC82, Kim20, Kra93, KMA02, Lev81, Mat03, MLL19, NK17, RH19, SDVT25, lDlzN21, vdKG22, Ano95e, Ano04d, How84, Cor22, Mah21a, Tsa02, Mah21b, Ric10]. Instructional [AAT99, Han10, MHDC04, PK23b, Ade98, Bra11, CDW14, CBSS96, HPS^{+}19, HMSBR19, Kae90, KBK21, OCB25, SKK^{+}23, SKN21, Tho84]. Instructions [Ano00f, Ano00g, Ano01b, Ano01c, Ano01d, Ano01e, Ano04e, Dim04, SS05b, WT22, Ano02d, Ano02e, Ano02f, Ano03c, Ano03d, Ano04f, Ano04g]. instructor [FC12, HOS13, TJHD11, WFCW15]. instructors [MG12]. Instrument [BP80, AV22, DB86a]. instrumental [DDB^{+}10, Her23, MB19, vDDDB21]. Instrumentalism [MO81]. instrumentation [PD23, SAD21]. instrumenté [CLR16]. Instrumented [Riv07, CLR16, GBR22, Kid08]. instruments [FLM07]. integer [Lie90]. integers [Gal02]. integral [Jon18]. Integrated [FGA94, Kry69a, Tre87, HP08]. Integrating [Eve99, NBZ12, CP10, LLL21, Sia24, WqWLR18, Wit84, bL00]. Integration [Ass05, AG02, BA14, Boa93, Edw09, Fre69c, GBE13, GBTL23, LE09, NK17, Ort83b, Sch12, YST22, Zen18, AG02, Cas69c, Fre69d, bL00, Cas69c, Fre69d]. intégré [Kry69a]. intellectual [Fis69, Ham68, Høy07, Kap09b]. intellectually [SOC86]. intellectuel [Fis69]. intelligence [MY84]. intelligent [HOEF89]. intensive [MAM24, TAH08]. intention [PB09]. intentionalities [BRE20]. intentionality [Sko16]. interacting [Sch11a, VSS20]. Interaction [CYW92, DEM98, Jun91, MHRSU99, Mos04, Rad11, SS05a, Ste06, Bar03, Elb03, KP10, KV23, KH07, Led87b, Lor82, RR11, Ste97, TL15, TZLT95, Rad11]. Interactional [Byu24b, Kle12, Byu24a]. Interactions [AS04, AS05, AS17, BdSCdA18, Flü05, Sol06, XM17, Bat13, CY23, CF00a, GB23, HM13, OAS25, RNM12, WSR11, ZHT25, dA00]. Interactive [BS94, OPC^{+}98, BF21, BH94, LSH09, PPG95]. intercept [Mos98]. Intercorporeality [RR11]. interdependence [Tay91]. interdisciplinary [Con90, KLBMK^{+}13]. interest [CSP24, KCS22, LS20, SLP^{+}12, ZC19]. interface [MPS12, Wak14]. interfering [NT75]. Intergenerational [Zev04]. interior [dF04]. intermediate [Hed85, LL23]. internal [Kae93, Kae94].

International [Ano70, Ano71c, Ano72b, Ano74c, Ano74d, Ano76h, Ano76g, Ano77c, Ano77b, Ano78d, Ano91c, Ano04c, Ano04d, De 24, Cor22, Fre78b, Han98, OM24, Osi24, Ric10, RD92, Tep20, BCZ21, Bal90, BGSW94, CM20, DT08, FLT14, Fre69a, Hes13, Hus25, Men20, RSD23, TE14, Ano74b, Ano95e, Art05, How84]. internationally [Fre75]. Internet [BBAR23, EH05a, EH05b]. Interplay [MMBD05, Ste05, CH13, HFL^{+}23, Kid11, MYM12, Ng16, SDVT25]. Interpret [NC06, AC21, Son13]. Interpretation [Coo98, Rut05, LT86, RT14, SL19, SSL25, Sko16]. Interpretations [Jon00, SLW98, Cal12, Deb06, ES03, GC23, Deb06]. interpreted [BGT17]. Interpreting [RL04, SL06, Wal12, BCE^{+}19, BBDD25, BBVD25, SMFL19, WM01]. Interrelated [HMT13]. Interrogating [VCD24]. interrogation [Wal10]. intersection [Pai16, Wei17]. intersections [dF04]. Intertextuality [RFP14]. Intertwining [CBC25]. intervening [Bis76]. intervention [BGW^{+}14, BGW^{+}16, BJ88, Deb06, FQL^{+}17, Feh69, Ott20, SS91, dVSKvO18]. intervention-based [FQL^{+}17]. Interventions [DEM04, Lun77, NK17, SS17]. Interviews [Coo98, Dah04, HBZE98, HS13b, KH07]. into/influencing [Jaw03]. intrinsic [Kyn93, WS18]. introduce [WS08]. Introducing [AP01, Har75, Prz05, WBM^{+}14, ZK15]. Introduction [Dör03b, FWM85, LPG05, LH08, Mar00, RFK07, REA09, SSL25, VBT12, ZBEH06, AL00, BBS07, BTV19, Bec77, BW08, Bro10a, GB14, HWR23, HMG19, HMG20, Kru13b, Law16, Pap71, Pic71, RD92, VPR21, Gla68]. introductory [CJ75, Sha77]. Intuiting [Gri13]. Intuition [LW99, Tsa07, BR78, FTH79, Mac78, MF10, Sem08, Tal80, ZS15]. Intuitions [FBM71, FG97, Fis99, FG84, Kid11, VVV13]. Intuitive [LH09, MD01, SZ07, TS99, TTL08, Zaz99, Bar19, FTM81, GS18, Jou25, Lie90, Tru15, VDWV04, Wal73b, Wit81]. invariant [NHO22]. invariants [LBFM13]. inventaire [BD72]. invented [Son16]. inventory [BD72, LC13]. Inverse [PSH^{+}18, KB20, MZ18, RL12, SPNH12]. Inversion [Gre12, NBE^{+}12, Sch12]. Inversive [Cox71]. Investigating [BA25a, Bon14, CWL07, DDH25, Gra04, Gue08, KC09, Nte06, PB00, Tha10, Wat80, LNC23, Mey10, WSR11, LOW24a, LOW24b, MPS22, ZB19]. Investigation [DS06, PW03, Ang89, Cob87, FT03, Gal81, Hod96, LS70, LG13, Mat81, MRSP14, Ng21a, Rot17, SM72, Ste97, VP13, WSU19, Wil71, YC95, dPH13, PW03]. investigations [GBR22, HI94, Hir81, Kon21, Kon22, vSBvR00]. invisible [RM15]. invitation [WHE08]. invitations [PKP^{+}18]. invite [MCW25]. inviting [Byu24a, Byu24b]. involved [BD72, Ior17, Joh15]. involvement [KLLJ17, Tak18, Tao86, DS02]. Involving [DVJ98, AG14, AAM21, Dum82, OB11, SW95, WH24]. Iran [Goo07, Ros78]. IRE [TL15]. Irish [RO09]. Irrational [MZ21, SZ07, FJC95]. irresistibility [DVJV02]. ISBN [Ahm23, Cor22, NX23, Sfa15, Shv22]. ish [WS25]. Isoda [RSD23]. Isomorphism [LS70, MGD^{+}23, LHZ95, RS22]. Israel [MA09]. Israeli [Als24, RA21, WAL18]. Issue [BA25b, FV22, Hof06, NBD^{+}04, Obe19, AP01, Bak19b, Bro08b, CL18, Col21, HMG19, PR08, TP14, HMG20]. Issues [DWMB05b, DWMB05a, Spe05, AL78, BBS07, BN91, DGI23b, GB14, HS14, Pop88, PM19, dV18]. Italian [Baz88]. Items [Eva07, DH90, Küc81, LBML24, Ola14, SJ17]. Iteration [Wei91]. iterative [RW11a, RW11b]. itself [FB17]. IUFM [Art23b]. I’ve [NRV24]. IX [Rog71].

J [Ano16b, Ano24a, Cor22, Eng99, Gue19, MDU^{+}25, MR20, Rut02, Sch24, Tal08, Web18a, Web18b]. J. [Sfa13]. Jackson [GB19]. Jahrhundert [Dam87]. James [Ahm23, Cor22, Tal08]. Jana [Tab22]. Japan [Ala12, MNRR95, Min83, WL90]. Japanese [RSD23, AAJW25, BMR25, Hin15, Miy17, RSD23, Sek21, SLC95]. Jean [Bac21, Sta11, Wit73a, Wit73b]. Jean-Baptiste [Sta11]. Jenlink [Mah21a, Mah21b]. Jennifer [Sto22]. jeu [LS79]. Jewish [Leu21]. Jill [Wal17, Ano16c]. Jim [LH08]. Jinfa [Ahm23, Sch24, Ell18]. Jiu [Siu93]. Johan [Hus25, Gof77]. Johannesburg [Ano16c]. Johnny [Dre99, Kon18, LH97]. joke [RC20]. Jonathan [KMT16]. Jordan [IE07]. José [Dek87]. jouer [Kry88b]. joueur [ZP91]. Journal [BR89, Han98, SWG05, Uhl03, AMMS20]. journals [GDS22]. journey [Ahm23, HJ12, KMT16, Sta11]. judgement [BFA22, DAJ20, JI15, Rut94]. judgements [FNM91]. judging [BOL24]. judgment [Jan74]. judgments [LD88]. Jugements [LD88]. July [Ano70, Ano72b, Ano74c, Ano74d]. June [Ano71c]. Junior [BZA01, Che89]. Jurdak [Edd18]. Just [CH02, WHE08, Gar24, Hic24]. Justice [All99, MFG22, YK24]. Justification [KD10, Cha93, MPdP17, Rot21]. Justifications [MS19, AAM21]. justifying [TMH21]. juxtaposition [HOS13].

Kaiser [Hes13]. Kakoma [Hyl24]. Kanamori [Han98]. kann [Van76]. Kaput [HN08, LH08, Tal08]. Kara [GB19]. Karen [MND23, TR24]. Karin [Han12a]. Karp [Law16, Law24, Men20, Men23]. Kathleen [Fur19]. Katie [For00]. Kaur [Hus25, LX16, YK20]. Kay [Atw24]. keen [VM78]. Keeping [Gra20]. Keiko [GB19]. Keitel [All99]. Ken [Hus25, Ste13]. Kenney [Sch24]. Kent [Gai11]. Kenya [NCMO15]. Keszthely [Ano71c]. Kevin [MND23]. Key [Coo98, Ram03, Kaa07, Pla18, SJS12]. kid [GB19]. kind [dF16]. Kindergarten [Bra13, CEHM16, KAJHB10, LHS19, LBML24, Di 19, SM72, TTL^{+}13, TTTL10, TTL^{+}17]. kindergartener [MB16]. kinematic [GK71]. kinetic [ST10]. King [SE93]. Kingdom [Mor21]. Kjeldsen [Fur19]. Klaus [Fri21]. Klein [Ano04c, Ano06b, Ano16b, Art23a, Art19]. Knigge [LT20]. Knipping [JV17]. Knocking [Loc92]. knots [MP96]. Know [Bur98, DS06, MM00, Bro23, Pol76, Swe95, Wed99]. Knowing [Coo99, Gra99, Par88, Boa02, Bro94, MS99, Rad14, Shv22, Sim96, FSY22, Law00]. knowing-to [MS99]. Knowledge [Adl97, Boa99, BC05, DP99, Ern99, Flü05, Fri07, GL98, GPPT99, Mos05, RK17, SZ07, THS06, Tir99, ABOM20, BA14, Bak14, BN96, Blo03, Bro10b, BOL24, CGD21, CLSM17, Dag96, DR13, DVF18, DVT^{+}18, DK15, EVHRC^{+}25, Eve90, ET95, Eve99, Fau15, FMA20, GC00, GSW22, GB13, GS18, Hah14, HM95, HPS^{+}19, HHB21, HH24a, HH24b, Hun86, KMT14, KRHD15, KML22, KD10, KLLJ17, KM21, KJ22, LLW07, MvdKM13, Mah21a, Mah21b, MM11, MS99, NBHK07, Ott94, OCA13, PA14, PM00, Rad03, Rad11, RSL17, Rot21, Sáe09, SSPS20, Sch20, Sch12, SC90, Ste91, Tch11, TP15, VDTV24, VP13, VSS24, Web01, ZZ11, ZB19, dVSKvO18, MMC22]. knowledgeable [AGL17]. knows [For00]. Koeno [HJ12]. Kollosche [LT20]. Kong [Che88, Che89, Leu95, LNC23, YSL22]. Korea [KHY21, SS10]. Korean [HC14, Wha96]. Krygowska [Now92, Tur88]. kujitegemea [Phy71]. Kuwait [Ala12]. Kuwaiti [AR07, HA23]. Kwon [YK20]. Kyungmee [Jon16].

L [Ott76, Sel00, Web18a, Web18b]. laboratory [BGW^{+}14, BGW^{+}16, RT14]. Laborde [Rut02]. Lacan [Bro08a]. Lagrange [Sta11]. Lakatos [Nun96]. Lam [SMWM25]. lancement [CGGGG79b]. landscape [Jou13, WFB13]. Langage [LS79, Dum82]. Langford [Ril86]. Language [Ade90, AH79, Cla92, HP25, Pla18, SA00, SK07, War06, All85, AMRM24, Art23c, Bar03, Bro94, BH11, CJ75, Daw83, DOV24, DS21, Dum82, EPQH18, GDS22, GM10, Gut07, Jon82, Kni12, LOW24a, LOW24b, Mah21a, Mah21b, Mor21, Ng16, OSR71, OKL87, Rad03, RO09, RG19, RG20, Rob12, Sfa13, SGCA23, Sou83, VM13, Zep81, dAS22, LS79, Sfa13, Mor00]. language-responsive [LOW24a, LOW24b]. Languages [Cla07, Mos07, Pla14, RJWB20, SA00]. Lanka [Gun78]. Large [Kle01, MGP75, AG14, BCE^{+}19, CVN21, FKC17, Fre72, GAP24, LLZ20, PDT^{+}12, Wai70, WSU19]. Large-group [MGP75]. large-scale [CVN21, FKC17, GAP24, LLZ20]. Larochelle [Eng99]. lasting [NK17]. late [Bla78, Ehr78, El 78, Fey78, Fre78c, Fre78d, Gun78, HV78, How78, Kap78, Ohu78, Pur78, Rev78, Ros78, Sha78, Wil78a, Wil78b]. Latent [KM25]. later [SWVO23, ZC19]. Latino [DLK14, TDEM13]. Latino/a [DLK14, TDEM13]. launch [CGGGG79b]. Laura [Bro11]. Lave [Car08, GS08]. law [Fre72, MM20, SM17, WSU19]. layer [SSA20]. layer-model [SSA20]. layering [Cal12]. lead [Bro10a]. leader [Eve99]. Leading [Car99, BWHM^{+}10, BW13, SGB21]. Learn [CRL06, Gra99, MCB05, McN00, Bas17, Bye15, CCP23, EPQH18, HM25, LK14a, NT19, Sfa01, Way94, WFC23, WvdHPD15, WJW11, YM16]. learnable [Nos01]. learned [MP24, NRV24, Osl12, dF12]. Learner [Ott20, Sho21, AS04, Bro14a, EG23, GB23, Han19, HMG16, Hin15, Ng16, RBWS18, RL11, WS08, dAS22, AS05]. Learner-generated [Ott20]. learners [Bur87, Sto22]. Learning [Adl19, AGP15, Ano76d, Ano98i, AAJW25, AI07, BH05, BZBM15, Boa99, Bro14a, BEW99, CBF07, Cla07, Cre03, Dah04, DH97, DEM04, DEMW06, Duv06, Elb03, EE21, Fra13, GH98, GSE07, Haz99, HA23, Jaw03, Jon82, KFS01, LaC14, Lab00, LJW13, Leh23, Ler01, LHZ95, LSC20, LW94, MMBD05, Mor21, Mos04, Mos07, NBD^{+}04, Nic98, NC06, OK05, PBY03, PBR^{+}04, PDV07, Pla14, SLP06, Sim94, Ste05, Sto71, Str03, STM06, Tir99, Van77, VM13, Vin97, Wat87, WD05, WMPL08, Zas05, vSBvR00, ATAB25, All85, AMRM24, And09, Ant96, Arc03, Art97, AWE22, Bar21, Bel79, BCFC90, Bju04, BdSCdA18, BC08, BC12, Bro16, BE23, CMW^{+}13, CHM16, Car10, Cha18, CW14a, Che88, Chr69, Cla02, CYW92, CBSP12, CLR16, DDH25, Dié70, DPG89, DŽC25, DDLZ94, Duv91]. learning [Ell13, EPQH18, Eva99, EJ23, FFL80, FvM97, FP85, FF17, FG17, Fis77, FBA10, For95, FT00, GP90, Gal11, Gal19, GRR23b, GFWL11, GMNB18, Gra04, Gre12, GAP24, Güç16, GBR22, Had78, HFL^{+}23, Han12b, Han12d, HS13a, Har73, HC95, Hea81, HPS^{+}19, Hee78, HBZE98, HS13b, HS99, Hew12, HM13, HMG19, HMSBR19, HMG20, HMCT25a, HMCT25b, HGM15, Hof77, HP03, Hoy82, Hoz81, HHB21, HHT23, HHL22, ICR20, Ior17, Ior18, Jee71, JWFC^{+}25, Jou25, Kae88, Kae90, KK12, KT01, KM24a, KB12, KC09, KZH^{+}19, KR09, Kru13b, KBBK20, KBK^{+}24b, KR11, KL14, LL23, LH71, LOW24a, LOW24b, LCK24, Loc92, MND23, MG00b, MDPW96, MJS21, MS14, MR17, ML20, Mot71, Neu99, NCMO15, NBA23, Nos88, Nuc24, OLW20, Ott11, Pim90, Pla18, PSE87]. learning [RSS11, Rad14, RL00, Rel96, Reu23, RO09, RMAGMR22, RFP14, REL25, RIA17, RH02, SGRS24, SBDM69, Sca75, Sch92, Sch12, Sch24, Sch11a, Sek21, Ser76, Sfa01, SD13, SAD21, SBDD24, Sia24, SW94, Str84, Str85, SMS22, TN16, Tab22, Tak18, TA08, Tar09, Tic12, TMP10, Tro00, Tro04, TE14, Tzu07, VAM21, Voi94, Whe92, WM09, Whi09, WS18, Wil71, WPL24, Woo96, YST22, dAS22, Cha71, Dag96, HbL05, Wal17, Shv22, Wal17, Men23, Tsa02]. learnt [BMK21]. Lebanon [SC17]. lecture [FC12, Wai70, WFCW15]. lecture-based [FC12]. lecturers [Kei93]. lectures [OLW20, PKP^{+}18, WFC23, WWFC16]. lecturing [MGP75]. Leder [Art23a]. Leen [Dör03b, Pre03, PvdHP03, YSRU03]. left [Tro00]. left-handed [Tro00]. Legacy [PvdHP03, Art19, Bag08b, Fri21, Gre01, Str93b]. Legitimacy [Baz01]. Leikin [Sil25]. Length [DVJ98, Jer73a, Jer73b]. lens [De 24, BBL22, CL23, Cor17, CHK22, GBC12, Kle12, McC14, MTBR16, Nol12, NH15, Pre16, REL25, SLAY19, Sty11, Vou16, dAS22, vDDDB21]. lenses [DGFT13, KHRK15]. Leone [Wil78a]. Leong [YK20]. less [Jon82, MPS22, SLM79]. lesson [CM20, MW09, RSD23, ZHT25]. Lessons [Ort24, BW20, BMR25, Chr00, CM20, DSR^{+}23, EK10, Hin15, LE09, dF12]. Letter [İşl23, Pre23, Tun04, Wil05, KM25]. Letter-as-Object [KM25]. Letting [MD01, LJW13]. létude [HK99]. Leung [Jon16, Ric20]. leur [VD78]. Level [Ano98i, DEM98, DEM04, Haz99, Nar00, Ade98, AOJ22, BN77, lD95, GS76, Güç13, Har09a, Hed85, JGF20, JE14, KO21, Kra93, El 18, PRR08, Par91, RWZ15, Rüe13, SV25, Sem08, Sha77, Ste68, Vil71, Vin77, XXL21]. Levels [Miy00, PP12, KZH^{+}19, MCPGE24, MY84, OKL87, Van76, Was15]. leveraging [SBL22, Wal17]. Lewis [Cor22]. Lexical [HEWC10, HEW10]. Li [Cor22, Hes13]. Lianghuo [DKL^{+}25, ZB19]. liberal [Zev96]. liberate [LT23]. liberatory [McV24, Ort24]. liées [Add76, TLM07]. liés [Art97]. Life [LW99, AWE22, Cot24, Dom25, FB11, NCC22, TB09, VS23]. lifelong [TE14]. light [Add93, FF21, GM08, MCL25, Sch16, TPHS07, Wil21]. like [MF17, Shr10]. Lima [CRL06]. Limit [MD01, FPS16, Güç13, Kid08, KT15, Prz04, Roh08, Roh10, TK14]. limitations [Gue19, Tri95]. Limits [BBEG05, Ber06, Bur05, Har09a, Pai18, Rad03, Sie87, TV81]. line [CN16, Ern85, KP10, MGV^{+}25, TT07, TP15, VBE21]. linéaire [Bic82]. Linear [DVJ98, Lau74, MSdSGG16, PM97, Uhl02, Uhl03, AZ20, AWL15, Bic82, BMPT24, Bur02a, CO10, DVJV02, DRR02, GKMJ24, GB14, Hew12, Kar16, Kar02, Pir25, SD94, Sta89, Vla02, WSR11, ZER17, dLT08]. Linearity [VDD^{+}03]. Lines [Car93, CP10, DVC19, FQL^{+}17]. Linguistic [JM74, LCD90, LR20, Row95, Jan74, Jur77, Lep90]. link [IWF^{+}15, JM18, Mah21a, Mah21b, MB19]. Linking [Hah14, OKL87, Ron15, Van03a]. links [BN91, KBK21]. lion [MF17]. Liping [FSY22, Law00]. lire [Rut02]. listen [AI07, SM21]. Listening [Nic98, SM21, HRW25]. listing [LG16]. literacies [Wei17, Gai11]. Literacy [Men07, Büs24, Gel04, HMSBT21, Kil01, VRP^{+}12, Wei17]. literate [Kae88, Kae90]. literature [CKS22, Fre13, GB01, HTM^{+}23, MGI23, MHC24, RBWS18, Sia24, SMOR20, TP85, YST22]. Littlefield [Mah21a]. Liu [Siu93]. Live [NN97, LSS19]. lived [AZ20]. Livia [De 24]. living [Bor21]. load [vLXD18]. local [Als02, GBE13, GB19, KSS25, MA25, SV25]. locally [BMPT24]. lockdown [DTD21]. Locutions [SLM79]. logbook [AAP21]. Logic [Tsa07, Var72, AKL80, CJ75, DC17, Dum82, Fre69e, Gil71, GS76, Kap76, Šed69, SS95, Siw73, Ser69]. Logical [Ano72a, OSR71, O’B72, O’B73, OK05, DG03, HK02, Mey10, PHS12, Rob75, Roh10, Vil71, Zep82, ZML87]. logiciel [Art97]. Logics [SO73]. Logique [Ser69, Siw73, Ano72a, Dum82]. Logo [Nos86, Sut93, DG89, LDM92, OKL87, DG89, LDM92]. Logomathematics [MN95]. London [Leu95]. long [AMRM24, Fur19, NK17, Str84, Str85]. long-lasting [NK17]. long-term [AMRM24]. Longitudinal [War06, Wat01, Bla85, CW96, MMR21, Ng21b, Pet91, SPNH12, SMS22, TF95, Van03b, VS25, Yer00, ZWC^{+}24]. look [FZ07, FKC17, KL82, MW19]. Looking [MJS21, Sfa01, ART15b, Sil13]. lors [BC92]. lost [Sur77]. Loto [Plu79, Plu79]. Loto-questionnaires [Plu79, Plu79]. loving [Lun12]. low [And09, Bac89, DB86a, DB86b, JNF25, Kar10, KB11, PFSS19, SHG21, SPFGF14]. low-achieving [Kar10, KB11, PFSS19]. low-attaining [SHG21]. Lower [Miy00, ARKK23, Fre74b, HC95, KBK21, Miy17, Var72]. Lowrie [Hus25]. Luc [Cha22b]. Luis [Shv22]. lumière [Add93]. Luneta [Hyl24]. lycée [Her11, Pro70]. lycées [For69]. Lyn [Sto22]. Lyon [Fre69a, Fre69a].

M [Art19, Eng99, Fit15b, Fur19, KMT16, LQV04, Mah21a, Mah21b, Mor00, Ste13, Dre06]. M. [Hus25]. Ma [FSY22, Law00]. maîtres [DS02, Del69a, HK99, Kry69a]. Macdonald [Gri78]. machine [Hol81]. machines [Gla68]. Macro [Flü05, JP09, Was15, dA00]. macro- [JP09]. macro-levels [Was15]. Macro-Situation [Flü05]. Madden [Cor22]. made [BHLH85, Cla83, Del68, Ell86, ISJ24, LH18, MS19]. made-up [Ell86]. magic [dLT08]. magnitude [MGV^{+}25]. Magnitudes [Ste69b]. Mainland [YSL22]. mainstream [Edd18]. maintaining [Als24]. Major [Bag06, Fre81]. Make [RSL17, VL04, Mon00]. Makes [Ste06, PK23b, WS25, ZC08]. Making [Bar96, Güç16, KOMH22, MT15, Moo94, Moo13, MM00, Ryv06, SW14, Wak14, WG07, ZG01, Bis76, CN25, EG23, FDW19, FC12, KO23, KBBK20, OCB25, PK17, Pal08, RSS11, Rot16, Sch16, See11, VC19, Sel00]. Malawi [Mwa02]. Malawian [Kaz07]. malentendus [Add75]. man [Add93, Van76]. manage [HSM22]. Management [Ass05, dF12]. managing [CN25, DG89]. mangoes [TB09]. mangues [TB09]. Manipulation [IB05, DF15, Dié71, LDM92, MP96, WFB13]. Manipulatives [Moy01, ZM25]. manufactured [KMN24]. Manufacturing [Sen96]. many [LHZ95]. map [FF21, LLD14]. Mapping [GB13, Ler14, DKL^{+}25, HM95, Ryv04]. mappings [Bal88, Ste71]. maps [EJ23]. Marcel [Mal23a]. march [Hol73]. marge [Del68, Mot71]. margin [Edd18, Mot71]. margins [Del68]. Maria [vBP19]. Marianne [Gre17]. Markers [BR06]. marks [BC08]. Marschall [Wan25]. Marseille [Art23b]. Marta [Fri21]. Martinovic [Mal23a]. Masami [RSD23]. Mastery [MHDC04, Dri25]. Material [Fer15, HSNHHE15, dF16, Gel04, PFSS19]. material- [PFSS19]. Materialien [Dam87]. materialist [FF17, Rad21, dFS13]. Materials [Eva07, NC06, WD05, Ano74a, CN25, Dam87, FM14, LOW24a, LOW24b, Rot16, Sin03, SHAC01, dAS22]. Math [BHX97, HBH07, Nic02, BSV95, BAZF04, CY23, Cha22a, Chr00, FMA20, Lev81, Mac21, MHH90, SRP21, Whi87, WJW11]. mathematic [HR00]. Mathematical [Abr01, Adl97, Alb00, Ano76h, Ano76g, Ano77c, Ano77b, Ano91c, Ano95e, Ano00h, Ano04c, Ano04d, AS25, Art05, AH76, AL78, BS25, De 24, Bur80a, Bus96, CMW^{+}13, CHM16, Cal94, Car99, CSP24, Cha97, DP99, DP92, DG06, De 94, DEM98, DEM04, Die15, Doe06, Eng06, Ern84b, Ern06, ETC14, FB98, Fis99, FB16, Fre69a, Fre78b, GABB25, Gra88, Gri71, Han20, HO23, HMSBT21, How84, Jon00, Kaa07, KAJHB10, KO21, Leh24, Lit00, MM03, MTBR16, MHRSU99, Mos04, Mos05, NN97, NF09, NH19, NHH97, Öst06, Ott94, Ott06, Ott07, OB06, PBY03, PT09, Phy71, PDV07, PS88, Rad03, Ric10, Ryv06, Ser69, Sfa15, SS06, Sko85, Sko90, Ste05, Ste06, SY79, Swe95, Tir99, TP10, Usi85, Vog13, War06, WD05, ZC06, Abe78]. mathematical [Add72, Add76, ARKK23, AG83, AC21, AP25, AV22, Ano69, Ano78d, AP75, ART15b, AH79, BF21, Bak14, Bal80, Bar19, Bar21, Bar09, Bar16, BR12, BVR24, Bel76, Ber04, BLC20, BBAR23, BP22, BWHM^{+}10, BIBS^{+}19, BN91, Boa02, BLP10, Bos87, BJJ13, BB25, Bra13, BP80, Bro91, Bro10a, BOL24, Buc87, BE23, BTvG^{+}17, BE85, ČK15, CL20, ÇA20, Cal12, CV09, CCW95, CBC25, CEHM16, Car18, CKS22, Cha18, Cha98, Cha93, Che23, Chr69, Cla85, CWS93, Cle79, Cle80, CW14b, CSS^{+}14, CGD21, CB17a, CB17b, Cor22, Cra12, Cre03, CSK12, CHK22, DC92, DT12, Daw83, DC17, DOV24, Del68, Del69b, DT08, Dom25, DS17, DKBD19, DKBD21, Dri25, DBL22, Duv83, Edw09, EN79, Ell13, Eng69, ES21, ERS19, EVHRC^{+}25, FD24]. mathematical [FF25, FTM81, FGG13, Fos16, Fos18, Fra13, FV22, Gai11, GG22a, Gal82, Gal86, GB23, GGD22, Gel04, GJ19, GDK23, Gli91, GK79, GFWL11, GG96, GK20, GSW22, GM10, Gre12, GB17, Gue19, GBR22, GLH20, Hah99, HI94, HH88, HTM18, HRW25, Ham68, Han12a, Han12b, HR74, HWR23, HM95, Hay87, HH11, HMA11, HBZE98, HS13b, Her11, Her02b, HM13, HMG19, HMG20, Hie88, How73a, How73b, HN92, Hoy01, HHB21, IMRS07, IS08, IS09, Jab23, JK17, Kap78, KPS17, KM18, Kem81, Kie01, Kil01, KID25, KL20, KLLJ17, KK13, KB11, KJ22, Kon93, Kon18, KOMH22, KMA02, KDM21, KCS22, KSH25, KLBMK^{+}13, Kru13a, KC73, Kry88b, KM24b, KL14, LaC14, LM11, LC81, Les81, kSL13, Lev22, LBML24, LR20, Lit03, LLZ20, LT23, LS79, LMW91]. mathematical [LK14b, MPS12, Mar13, MT15, MB09, Mas69, MPdP17, Mey10, MS20, Mie23, Min83, MJ15, MLL20, MWVE09, Mur93, Ng16, Ng21a, NBA23, Nos88, Nun96, O’C01, Obe19, OAS25, Orl71, OB11, PA18, Pal20, PGE09, Pet91, PK94, Pre86, Prz04, RT85, Rad09, RSS11, RCH25, RWZ15, RG01, RSL17, Reu23, Rev71, RO09, Rob95, RNRVM23, Rom74, REL25, RT09, RR23, Sáe09, SLP^{+}24, SG14, SM72, Sch16, SP19, Sch23, Sch25, Sch11a, Sch11b, SS95, Sfa91, Sfa01, ST20, SE93, Shr10, SL74, Sil25, SKPS16, Sim17, Sim15, SM21, SOC86, SW94, Ste91, Sto22, Sty11, SSA20, TN15, TD12, TP02, TTL^{+}13, Tro04, TS93, Tur80, Tzu07, UOF25, UB23, VRK24, VS13, VP13, VWV21, VN19, Voi94, VEBS24, WB12, WB11, Was15]. mathematical [WT22, WR77, WDvO19, Yah25, Yer00, YTD11, ZG01, ZS15, ZER17, ZZ11, ZBBZ85, Zev11, ZSWC25, ZKC18, dPQ16, dF25, rA20, vO01, vO10, Fis69, Kry69b, KN21, Sil25, Bro11]. Mathematically [PW03, SLP06, Zev04, AAJW25, Bye15, GSH16a, GSH16b, Kis86, LaC14, Lev10, Ryv04, SGTK14, WZ86]. mathematician [Bes12, Dho93, Hem10, JWFC^{+}25, Kla71, Kon21, Kon22, MJ15, Pas78, Pol76, ST10, SRB16, WFC23]. Mathematicians [Bur98, HMRRTT23, Add72, Als24, DW17, Ell86, EMRI22, Gri68, JWFC^{+}25, LW14, LH22a, MRW14, PB00, Shr10, Wat80, WMR11, WFC23, WJW11]. mathématicien [Pas78]. mathématiciens [Add72]. mathematicizing [Kry68]. Mathematics [ABK^{+}05, ACJO04, All99, AF05, Ano72a, Ano74b, Ano76e, Ano88e, Ano98i, Ano02c, Ano04d, Art23c, AAT99, Ass05, Ayd90, Bag10, BT05, Baz01, Bec80, Bes07, Bis88c, Boa99, BW12, Bro23, Bur98, Bus05, CF03, CBF07, Car99, CWL07, CSW21, Cha06, Che88, Cla07, Coo98, CH02, Cop98, Cre06, CRL06, DS06, Dör03a, DS05, Duv06, EH05a, EH05b, Ern99, EMT06, Eva07, Fre82, Fri07, GH98, GSE07, Gat06, Goo07, GP01, Gra99, GL00, GD99, Hås68, Han89, Han98, Han06, HDM07, Har73, HZ05, HMRRTT23, HbL05, IE07, Jaw02, Jaw03, Kai13, KP10, Kei82, KC22, LH22b, Led92, Ler01, LH97, LH06, Leu01, Leu05, LCK24, Lib71, LYHC18, Llo06, LQV04, MCW25, MG00a, MDPW96, McN00, Miy00, Mor00, Mor06, MHDC04, Mos07, Moy01, Nar00, NBD^{+}04]. Mathematics [New98, NN07, Nic98, NC06, Nis99, Nte06, OM24, Ong06, ODV06, Osl12, OPC^{+}98, OCB25, Pai19, PBY03, Pea76, PBR^{+}04, PSS^{+}19, PC98, Pre06, RR05, RD92, Rod06, Rog85, RC04, SLP06, SC17, See01, Sel98, Sen98, SSLM^{+}05, SK98, SA07, SG98a, Sid05, Sko04, Sko88, Sko21, Sol06, Sou83, Spe05, SWG05, SK07, SK25, Str08, Sty07, STM06, SC83, TF95, Tie02, Tur93, VL04, Vin97, WW02, Wat00, Wat05, WWFC16, WI01, WW07a, WW07b, Wit05, Wit95, bMYT98, ZBEH06, Zas05, ZC06, Zev06, ZM91, ZS07, dKV18, ATAB25, Abt22, AAB03, Add75, Ade98, AZ20, Adl88, Adl19, AOJ22, AV08, Ala12, AAP21, Ale81, AT21, Als24, AMRM24, ARW22, AMMS20, AH00, And09, ARHS14, APS22, Ano68a, Ano68c, Ano81b, Ano14b]. mathematics [Arc03, AM25, Arm69, ABOM20, Art15a, APRS09, AC95, Bac89, BRE20, Bag73, Bag08b, Bak16, BCZ21, Bal90, Bal71, Bar21, Bar87, Bar03, BBS07, Bas17, BKK22, BGSW94, BTV19, BMR25, Bat13, Bau80, Baz90, Bec24, Bec19, Beg69, Bek10, Bel79, BCFC90, Ber19, BKGZ24, BD72, BWD10, Bes12, Big70, BA25b, Bis79b, Bis80, BH94, Bja20a, Bja20b, Bla78, Bla85, BN91, Blu02, Boa93, Boa02, De 24, BHN10, Bon14, Boo81, BR89, Bor21, BF96, Boy16, BR17, BKH25, BS15, BP15, Bra11, Bra14, BS94, Bro94, BSF96, Bro96, Bro08a, Bro08b, Bro10b, BC12, Bro12, Bro16, BSW16, Bul24, Bur80a, Bur80b, Bur87, Bur95, Bur02b, BB03, BE84, Bye15, Byu24a, Byu24b, CDW14, ÇA20, Cas04, Cas69c, Cas77, CW14a, CPK08, CPP09, CP10, Cha22b]. mathematics [CWT12, CL23, Che89, Che90, Cho22, Chr69, CEMP04, CWH19, Cla12, Cla85, CL84, Cla92, CM20, CWYP92, CYW92, CSRlR24, Con90, CF00b, Cd04, Cor17, Cor22, Cot24, CLSM17, D’A91, Dam87, Dar16, DVC19, Dav71, DR13, DIW19, Deb06, Del69a, DKL^{+}25, DGI23a, DGI23b, DMHRV21, Die15, DSR^{+}23, DHD23, DS21, Dom11, Dom19, lD95, DCHM19, Dow96, DŽC25, Dra14, DE24, DK15, DDB^{+}10, DTD21, DP77, Ecc85, Edd18, Ehr78, El 78, El 86, Elb03, Eli85, Ell86, Ell18, Ely10, Eng68, Eng69, EPQH18, Ern84a, Ern16a, Ern16b, Eva99, EJ23, ES03, EK10, FZ07, FP07, FvM97, Feh69, Fen79, FP85, FPCL90, FE09, Fer15, Fey78, Fey89, Fid85, Fis77, Fit11, Fit14, Fit15b, Fit15a, FB17, FKIA15, FC08, FBA10, FB11, For95, FLT14, FA01]. mathematics [For69, FGA94, FI17, Fra96, Fre68, Fre78c, Fre78d, Fre81, FA03, FSY22, FCN14, Fur07, FM14, Fur19, Fur22, GP90, Gai11, Gal82, GC90b, GCG^{+}92, Gar24, GRR23b, GRR23a, Gel98, Gel08, Gen70, GBMB70, Ger10, Ger81, Ger84, Ger88a, Ger88b, GCFS23, GP22, GLHM24, Gof77, Gof93, Gol03, GC00, GS20, GFJ13, Goo20, GdA09, Gos21, Gra20, Gra04, Gra75, GPPT99, GM10, Gre17, GS08, GMHG09, Gri83, GM84a, GM84b, Gro77, GAP24, GT09, Gue23, Gui17, Gun78, GPS24, Gut07, Gut92, HYC^{+}97, HRW25, HV78, Ham68, Han02, Han19, Han10, Han12d, Har09b, HS13a, Har81, HOS13, Hea81, HWJ20, HPS^{+}19, HP08, HOEF89, HEW10, HEWC10, HEWJ^{+}15, HSM22, HTM^{+}23, HJ12, Hes13, HM13, HMMG18]. mathematics [HMSBT21, HMCT25a, HM25, HMCT25b, HC82, HB69, Hir81, HPRW25, Hol81, HMA13, HP03, How78, How80, Hoy82, Hoy85, Hoy01, Høy07, Hud90, Hug74, HHL22, HH24a, HH24b, HMT13, Hus25, IM19, IPJ25, IWF^{+}15, Ior17, Ior18, Jan09a, Jan09b, JCM^{+}23, Jan81, JGF20, JP09, JK17, JM74, ISJ24, Jin07, JWFC^{+}25, Jon82, JMD08, JV17, Jon18, JGR14, Jou13, KSSO25, Kae88, Kae90, Kae93, Kae94, KM93, KBK^{+}17a, KBK^{+}17b, KR25, KMT14, Kap09a, Kap70, Kar02, Kar10, KS69, KWF15, KT01, Kei93, KL82, KFS01, Kie01, Kim20, KMN24, KB12, Kla68, Kle02, Kle12, KO21, KH07, Kol16, Kol18, KM21, Kol21, Kom16, KSS25, KBK21, Kon20, KMT16, KC09, Kra93, KR09, KLBMK^{+}13, Kru13b, Kry88b, Küc81, KHY21, LK14a, LL23, Lam15]. mathematics [LT20, LS70, Lan81, Lan13, LM11, LOMG^{+}20, LZ08, LSH09, Law16, Law24, Law00, LBH23, Led80, Led85, LJW13, LLW07, LOW24a, LOW24b, Ler89, LXT02, Ler14, Leu87, Leu95, Leu21, Lev81, LX16, LSC20, LNC23, Lib77, LS20, LC13, LC15, LYBK23, Lin88, Lle12, Lle16, LW94, LMH70, Lor82, Lou18, LLL21, Lun12, Lun77, Lyn68, Ma96, Mac78, MBF25, MGP75, Mac00, Mah21a, Mah21b, MMC22, Mai85, MDU^{+}25, Mal23b, Mal93, MK91, MCVM13, MP24, MND23, Mar69, MF17, MVCV24, MSdSGG16, MGI23, MS99, Mat03, MJS21, McC14, MLL19, MS14, McQ86, McV24, ME13, MR17, MFG22, MRFW^{+}12, MMR21, Men20, Men23, MG12, MA25, MY84, Mog17, MCL25, MAM24, MAHK08, Mor14a, Mor14b, MO94, MR20, MRC18, MA09, MHC24, Mur93, Mur95, Mwa02, NT19]. mathematics [NHO22, NBZ12, Nar16, Ng21b, Ng21a, Ngu24, NCMO15, NRV24, NR10, Nol12, Nol16, NX23, Nor15, Nos01, Nuc24, NEM99, Ohu78, OVKR25, OLW20, OKL87, Ort24, Osi24, PV12, Pai13, Pai15, Pai18, PKP^{+}18, PRR08, PK23b, PA14, PXTW17, Phe10, PPG95, PB09, Pim90, PW03, PK23c, PK92, Pis68, Pla14, Pla18, Plu79, Pol68, Pol69, PSE87, PAR^{+}24, PMG^{+}94, Pop88, PPSZ19, PJP23, PNH98, Pre88, Pre92b, Pre07, PR08, Pro13, PM19, Pur78, Qua92, QP22, RR17, RBWS18, RW11a, RW11b, RSRVRVVR23, RL00, Rel96, RK17, RR80, Reu23, Rev68, Rev78, Ric20, RL13, Rob95, RG19, RG20, Rob90, RMAGMR22, Roo19, Roo23, RL11, Ros78, Rot12, RM15, Rot16, Rot17, RW19, Rot21, Rou68, Rou69, Row95, Row08, RC02, RNC21, RM21, RJWB20, RMA98, RSD23, Rut87, Rut89a]. mathematics [Rut94, RH02, RDH09, Rut10, RC20, RNM12, RNP13, RH19, SMFL19, SJ17, SS15, Sar02, SBL22, Say94, SBDM69, Sca75, SGC22, SSL25, Sch12, Sch24, Sch11a, SM22a, SECP^{+}23, ST08, Šed76, See11, Sek21, Ser68, Ser76, SA00, Sev22, Sfa13, SJS12, Sha78, SGB21, SD13, SAD21, Shv22, SBDD24, Sia24, Sil13, Sil25, SS91, SS93a, Sim94, Sim17, SS03, SEC13, SMWM25, Siw73, Sko94, Sko16, Sol95, Sol12, Sri18, SV09, Sta11, SW92, Ste13, Ste97, SRP21, SG98b, SB12, SPS17, SKN21, SPFGF14, Str87, Str93a, SMS22, SMOR20, Sun11, SLW83, SS88, SW14, Swe77, TN16, TS18, TL15, Tak18, TV81, Tal08, TZLT95, TA08, Tar09, Tav68, Tay91, Tay96, Tch11, Tep20, TR11, Tho84, TPHS07, TB22, TLB^{+}24, Thw69, Tie00]. mathematics [Tre93, Tre83, TJHD11, TR24, UOF25, VRK24, VC19, VCD24, Van93a, Van03b, Van69b, Van76, VD78, VD79, Var76, VA18, VM78, Vin76, VCS95, Voi94, Vol76, WHE08, WHE09, Wag12, WBM^{+}20, WPA^{+}23, Wai70, Wak14, WBM^{+}14, Wal12, Wal17, WqWLR18, Wan25, Was22, Wed99, WW11, WFCW15, WHC09, Wha96, Whe92, WM09, WL90, WvdHPD15, WJW11, WS18, Wil21, Wil71, Wil78a, Wil12, Wil16b, Wil78b, WPL24, Wit69, Wit81, Wit84, Wit01, Woo96, Woo12, WKB97, XA20, XXL21, XH88, XC19, YHL^{+}15, YK20, YSL22, YST22, Yan11, YLH24, YK24, Yoo25, Zav21, ZZ11, Zep81, Zev11, ZWC^{+}24, ZB19, ZC19, ZHT25, ZK21, ZK24, ZZ08, dA00, dF04, dF12, dFS12, dFS13, dF16, dV18, lRA16, lDlzN21, rA20, vO13, vdS11, vdKG22, Atw24]. Mathematics [Fer15, Fur19, GDS22, HSNHHE15, Hyl24, Mal23a, MF17, NRV24, Nim77, Sta11, Tab22, TB09, Kae94, Gue19, Bak19b, Bec19, CL18, Edd18, Fit11, Fit15b, MW19, MBV21]. Mathematics-for-Teaching [DS06, Osl12]. mathematics-related [Han19]. mathematics-teaching [Rou68]. Mathematiklehrerberufs [Dam87]. Mathematikunterricht [Van76]. Mathématique [Nim77, Add72, BD72, Bos87, Cal94, De 94, Del68, DP77, Fis69, Fre69a, Fre82, Her11, Kry69b, Kry88b, LS79, Rou68, Rou69, Ser68, Ser69, Ser76, Siw73, VD78, Ano00h]. Mathématiques [Cop98, TB09, Add75, Add76, Ano72a, AP75, Bal80, Cas69c, Cas77, Deb06, Del69a, Duv83, For69, Gro77, Hah99, KC73, Mar69, Mas69, PRR08, Pis68, Plu79, Rev68, Rob95, Rom74, Sar02, Tro04, Van69b, VD79]. mathématiques/réalité [Hah99]. mathématisation [Kry68]. mathematization [Ste68, Ste69a, Tre83]. Mathematizing [Pre03, Wit76, HMG16, Win11]. mathophobia [RR80]. Maths [MRRF94, SE93, All85, Als02, AVM15, BG69, Chr00, JN15]. matrices [Fle68, Roo71]. matrix [KB20]. Matter [Bes07, WW02, BN96, BS94, Eve90, ET95, Fer15, Rad09, RJWB20, Fit11]. matters [Dom11, LS20]. mature [Kae93, Kae94, SBP11]. maturity [Hee78]. Maximizing [McN00]. may [BA25b, BHX97, PB09, dSBM21]. McCallum [Art19]. McNamara [dF12]. me [AAP21, Fri11, SM21]. Mean [BG06, PLW81]. Meaning [Bec24, Bre98, Chr97, DZ00, KR25, Rad06, RP07, Abr09, BP22, Cla02, EG23, FA76, HP25, Han12d, KP10, MM20, MB09, MWVE09, Pre92a, RFK07, RSS11, See11, Sko16, SSA20, VMN11, Voi94, Wea90, ZG01, Sam85]. meaning-making [EG23, RSS11]. meaning-related [HP25]. Meanings [CMdR18, Kaz07, NHH97, APN01, ART15b, AP01, Fra96, MGD^{+}23, NPH99, PSH^{+}18, Rad00, Rad09, RG01, Tou71, VS23]. means [Abr09, BH95, Hah75, HS13b, Kid08, Šed69, Ste71, Van77]. measurable [Ste76]. Measure [CB00, BLC20, CMW^{+}13, CHM16, FTM81, Leh23, Moo13, WBM^{+}14, VYZ21]. measured [Min83]. measurement [BN96, CBSS96, IB19, LK22, PM15, PPC09, VLY21]. measures [GRR23b, HC82]. Measuring [AV22, ALO^{+}91, KPS17, CKS22, Fau15, FV22, LBH23, Moo13, Tal80]. mechanics [Fre93, WS25]. mechanism [Noe80b]. Medal [Ano16b, Ano16c, Art23a, Art23b]. Medals [Ano04c, Ano06b]. media [ARW22, Bor21, Cal12, GG22a, HMSBT21, SK69, dF16]. Median [BG06]. mediated [GGR02, HO23]. Mediating [VA18, GRR23a]. Mediation [Adl97, Mar00, ZS07, AM25, CEHM16, Cha98, FLM07, MG12, SZP10, SLR^{+}24]. mediators [BF21, GdA09, RNP13, ZC19]. medication [CW14b]. Medium [Eva07, BCE^{+}19, RO09]. meet [Nar16]. Meeting [CW14b, GC90a, Sil25, Tre91]. meets [Con90, MMM25, Sfa01, Wil12]. Megan [Sto22]. Meixia [JC22]. Melbourne [XC19]. Mellin [HN95]. meme [BBAR23]. memes [BBAR23]. mémoire [Kry88a]. memoriam [Ano90d, Ber94]. Memory [BE85, Flü05, Kry88a, Men20]. Mendick [Bro11]. Menghini [Art19, Fri21]. Mental [ASP97, Bar88, FB02, MNRR95, Mur04, Pro13, Thr02, Chi98, Csí16, Hop85, MP96, PG20, Str78]. Mental-addition [Bar88]. mentally [Bar88]. mere [BE23, MS99]. Mesa [Hus25]. Mesopotamian [Høy07]. met-before [ZR18]. Meta [DG06, lD95, GDK23, KB12]. Meta-Affect [DG06]. meta-discursive [KB12]. Meta-mathematical [GDK23]. metacognition [GGR02, HO23]. Metacognitive [Kap01, KMA02, SG98b, Yan12, GG96]. metalevel [Güç16]. metaphor [Die15]. Metaphors [Baz01, BEW99, Cha97, FBA10, OLW20, WW07b, Pre92b, SGC22]. method [CW96, DVV20, GT95, GC23, HGM15, KB20, KB12, TMH21, WM09, Zen18, dKBvdS22]. méthode [GT95]. méthodes [Rou69]. Methodological [Ryv06, APS22, HMRRTT23]. Methodology [Han98, JV17, MM11, MCVM13, SMOR20, WWFC16]. Methods [FB02, Sel01, Spe05, WD05, Wat05, Ade98, Bal71, Boo81, Han19, HM95, Kal83, RFP14, Rou69, SLP^{+}12, TA08, TTTL10, ZZ11, JV17]. metonymies [Pre92b]. metonymy [NH15]. metric [Pic71]. metrics [AMMS20]. Mexican [MA25, RRP91]. Mexico [AC21, Gut92]. Miao [GLH20]. Michael [Bro12, MPS12, Web18a, Web18b]. Michel [LT20]. Micro [XXL21, BKH25, JP09, SMS22, Was15, dA00, dF12]. micro- [Was15]. micro-divide [JP09]. micro-longitudinal [SMS22]. micro-management [dF12]. micro-teaching [BKH25]. microethnographical [Jun91]. Microgenetic [VGJ19]. microworld [Kor03]. microworlds [HN92, SW94]. Middle [AR07, Mes04, VL04, Ell13, FKIA15, GAP24, HHDG16, KK12, MvdKM13, MRC18, PFSS19, RCH25, Sty11, Tch11, WS18, WPL24, Yan11]. middle-school [Ell13]. midpoint [Bar19]. mieux [Kra02]. might [HA23]. Milieu [SS05a, Blo03]. military [Als24]. millennium [GCG^{+}92]. Milo [HS22]. mimic [Hic24]. mind [ABvP02, Pre92a, SL19, MPS12]. mindfulness [YLH24]. Minding [Kon21, Kon22]. minds [TW09]. Minicomputer [Pap69]. mining [DS21]. minoritized [ARW22, Byu24a, Byu24b]. minority [MA09]. Minoru [LX16]. minus [ZG01]. Minute [BAZF04]. minutes [Ano78b, IT77]. Miriam [LT20]. mis [TT23, TOT24, TT23]. mis-in [TOT24, TT23]. mis-out [TOT24, TT23]. misalignment [HMMG18]. misconception [KM25]. Misconceptions [Sha77, AI12, Mev83]. mises [Bal88]. misinterpretation [Sim20]. MISP [Rog85]. missing [BS25, CVN21, IWF^{+}15]. Misunderstandings [Add75]. misunderstood [Add75]. mixed [DVV20, Dek87, SD94]. mixed-method [DVV20]. mixtures [SVVO23]. mobility [Bla85]. mobilizes [PA18]. modal [Wil09]. Modalités [De 94]. Modalities [De 94]. Model [AAT99, CPP07, Dah04, DVJ98, DEM98, Ecc85, FB98, Gat06, GPW05, SSLM^{+}05, dKBvdS22, BNPG97, BR22, CO10, Car93, CHK22, DT08, Hah75, Han12b, JP09, JK17, Jer70, JDK^{+}24, JTP94, KP10, KCS22, Lie90, Lun77, MRFW^{+}12, MPZ16, Mev83, NBZ12, NHN03, RL04, RH02, Sev22, Sha77, SGCA23, SSA20, TW09, VH16, Vla02, WMM13, YL08, dV18]. modèles [AP75, LD88, Rob86]. Modeling [BJ22, Eng06, ZC06, AV22, Czo18, GP22, Han20, KO21, LH18, LMCG08, RSRVRVVR23, Sev22, ST20, Sto22, VN19, YSL22, vDDDB21]. Modelling [GC90a, HPRW25, IMRS07, MHRSU99, BN91, Blu02, Bur80a, CLK22, CKS22, DBL22, Eng03, Eng12, Fre13, FB16, FV22, GGD22, GSW22, Han10, HGM15, HSZ17, Jab23, KCS22, KSH25, Leh24, LK22, RSL17, SLP^{+}12, SKR15, SKK^{+}23, TJ94]. Models [Fis01, GC90a, WW07b, AP75, Chi98, Cra12, Eli85, Eng03, GP90, GK71, Jou25, Kid11, LD88, Lec92, Lén93, Rob86, SPNH12, Sev22, ST20, SHAC01, Van03b, Ves69, Ves71, Ves73a, Ves73b, Ves76]. models-and-modeling [Sev22]. moderate [GMH23]. modern [Bja20a, Bja20b, D’A91, Eng69, Ham68, Men71, OM24]. Modes [Fre13, MM00, SV09, BP22, CBC25, Cho22]. modify [LG13]. moins [SLM79]. moment [CB19, FP07, MS99]. Monash [Ano16b]. Money [Bre98]. Monge [Fri21]. monitoring [EH20]. monolingual [Cla92]. Monster [Yam09]. Montrouge [Dou76]. more-than-human [dF25]. Morgan [PW03]. morphisms [BBF71]. Morris [Ahm23]. Moschkovich [Sfa13]. Moscow [Ano74d]. most [LOW24a, LOW24b, ZB19]. motion [NFFAP20, WFB13]. Motivated [HDM07]. Motivation [Han06, ZC06, Ger10, Jan09b, KB12, LC15, VSS20]. motivational [LHS19, Phe10, WS18]. Motivations [Lau74, Rou69, Cob85]. motor [CSDK25]. Movement [OM24, Bja20a, Bja20b, D’A91, HK25, HMT13, PA18, Rot16, RIA17]. movements [CSDK25, SKE25, ZSWC25]. moves [FC12, Was15]. Moving [Bur95, Son16, Wan25]. Mozambique [Ger81, Ger84]. much [Als02, KSSO25]. Muke [Atw24]. Mukhopadhyay [Fit11]. Multi [Nar00, AGL17, HO23, KM18, Wil09, YS12]. multi-directional [AGL17]. Multi-Level [Nar00]. multi-modal [Wil09]. multi-party [KM18]. multi-representation [YS12]. multi-touch [HO23]. multidigit [Cli17, JTP94, Tha10]. Multidimensional [CB05, NS25, TB22]. multifaceted [WE95]. multilevel [KM25]. Multilingual [Adl97, BBS07, GP01, EE21, Han12d, LT23, MR17, Mor07, Nor15, SA00]. Multimedia [RIA17, BH94]. multimodality [FKC17]. Multiple [DHD23, FKIA15, Gut07, Mos05, Obe19, Sfa13, TTTL10, BC96, Cli17, DK15, FA01, FT00, Hin15, KV75, LCW20, SJ17, SKR15, SAR19, SLC95, VPR21, WSU19, Wil16a]. multiple-choice [SJ17]. multiple-perspective [FT00]. multiplicatifs [Lev92]. multiplication [Ang89, BHHS09, Cli17, GT90, GPS24, HP25, Hee78, Hee85, IB19, Jou25, LT86, O’B70, RL12, RSD23, SW95, SS10, TG89, Ves71]. multiplicative [CRPPRG16, CS94, DS17, GCALB15, HS22, Lev92, ML20, PS21, ZW21]. Multiplicatively [EJDT06]. multiplier [Dem93]. multiply [DVV20, Dem93, SM22a]. multistep [CGD21]. multivariate [MWVE09]. Murad [Edd18]. music [CBSP12, Pre07]. Muslim [BW13]. Mutual [UOF25]. My [Dör03b]. mysteries [Hop85]. mythbreaking [Tay96]. Mythmaking [Tay96]. mythologizing [WHE09].

N [Eng99, Fit15b, Sfa13]. Na’ilah [SJS12]. Naiman [Rut05]. naive [Vin75, Vin76]. naïves [Del68]. Nam [YK20]. names [HA23, MB16]. NanoRoboMath [KML22]. Nardi [Mal23b]. Narrating [Sol12]. Narrative [DSR^{+}23, BS15, Kru13a, LK18, Nar16]. Narratives [BEW99, AVM15, KO23, WWFC16]. Nasir [SJS12]. Nathalie [Art15a, Fer15, HSNHHE15]. National [Coo98, CG21, Tep20, WAL18]. nationwide [KZH^{+}19, MRRF94]. native [Dag96]. Natural [Rob12, Tal01, Wit75, Gal02, MGV^{+}25, Siw73, VVV15]. natural-number [Gal02]. Nature [CB05, IB05, Nis99, NX23, Sen98, Boa02, Bor86, Dag96, DVJV02, FA03, GFWL11, KK12, LBT07, Miy17, Sfa91, VS23, Wat87]. naturel [Siw73]. navigate [LLD14]. navigating [Ahm23, Par25, Was15]. NCTM [Ell18]. near [EJ13]. necessary [Her02b]. necessity [PHS12]. Need [DS06, BF21, Con90, Gai11, HHS00, LMCG08, Web01]. Needs [Rod06, AC21, BRE20, DVC19, Dav71, dAS22]. Negative [LW99, SL74, Bek10, BLP^{+}14, HC95, MRCR11]. neglected [TTL^{+}13]. Negotiating [VN19, Cha22a, LW94]. Negotiation [Chr97, Voi94, KP10]. Nelson [Fit11]. neo [Kae88, Kae90, RR17, Wil16b]. neo-literate [Kae88, Kae90]. neo-Vygotskian [RR17, Wil16b]. neocolonial [Atw24]. neoconservative [KMN24]. neoliberal [KMN24, dKV18]. nerd [FSB18]. Nerida [Ste13]. n’est [Gla76b]. nesting [TTS21]. Netherlands [DKBD19, DKBD21, DTD21, Fre78d, Tep20]. network [WWFC16]. networked [BDS^{+}17]. Networking [AM14, Bak16, TMH21, Kid16, TRDA20, Bak16]. networks [BDS^{+}17]. Netz [Sid05]. Neubrand [Art19]. neurology [Tho85]. never [TT01]. never-ending [TT01]. Newly [Ryv06]. news [ARW22, GG22a, KHY21]. Newspaper [Adl88]. Newspaper-based [Adl88]. newspapers [KHY21]. Next [HN08, Cor22, Hoy01]. Nexus [GH98]. Nicolas [Ber94]. Nigel [Wal12]. Nigeria [Ale81, Ohu78]. Nigerian [Mal93]. Nina [Sri18]. nineteenth [BP80]. no [Pea76]. noble [Car08]. nombre [BC78, BC82, Com77, Com80]. nominal [GB13]. Nominations [Ano14a]. Nomograms [Fle76]. Non [Gor98, Add72, AL00, Bor21, CL23, Dom19, GDS22, KZ94, PGE09, PBCPP25, Rob95, SKE25, Ste76]. non-autistic [PBCPP25]. non-cognitive [CL23]. non-conceptual [SKE25]. non-dominant [Dom19, GDS22]. non-mathematicians [Add72]. non-mathématiciens [Add72]. non-measurable [Ste76]. non-routine [PGE09]. non-standard [KZ94]. non-strictly [Rob95]. Non-Visual [Gor98]. nonexamples [TTL08]. nonmathematical [TTL^{+}13]. Nonprofessional [Kar10]. nonviolent [Gui17]. Nonvisual [AS05]. Norma [JV17]. Normal [Wil97]. Normalising [APN01]. normality [Zav21]. normative [DH20]. Norms [RR23, ÇA20, DW17, HS99, LYBK23, LW94, SG14, VEBS24, rA20]. North [Ste13]. Norton [MDU^{+}25]. Noss [Gai11]. notation [BBG^{+}17, Gro77, Hew12, MS97, ZL02]. Notations [Bri06, Gro77, VS25]. note [Ano96e, IM19, Led87a, Šed76, She87]. note-taking [IM19]. notebooks [FA03, YM16]. notes [IM19, Mai85, O’B70]. Nothing [PM97]. Notice [Ano80c, Ano89b, Ano92e, CY23, Sch23, Sch25]. Noticing [MT15, AV22, BKK22, CLK22, CY23, Dom19, DK15, HY22, Lou18, RPP24, Yah25]. Notion [Sty07, Bic82, Cas69a, Com80, DG03, GT95, MPG12, Mau85, Pap71, Pla18, Rev72, Tal80, TT07, ZER17]. notions [BD72]. nouveau [AG02]. novice [BKK22, EG23, Nol16, Rüe13, SECP^{+}23, ZHT25]. Number [Ern06, GD17, LW99, Mos05, Tsa02, BL22, BFG84, BC82, Car93, CPPT^{+}13, CVN21, CS19, Com80, DB86a, DB86b, Ern85, Fre72, Gal02, HC95, HS13b, HA23, Hof77, JTP94, JNF25, KML22, KBK24a, MGV^{+}25, MB16, Sfa15, SS93b, Son16, Tal80, TT07, Tho90, TP15, VVV15, VAM21, VBE21, WD10, Yan19, Zaz99, Com77]. Numbers [JB21, LW99, Sel01, SZ07, Van84, Zaz98, AG14, Ano14b, BST81, Ber08, BLP^{+}14, BP22, BC08, Cha18, DVT^{+}18, Eke77, FJC95, KBBK20, MGV^{+}25, O’B70, Ott03, Ste69b, Tha10, TT23, Usi74, Ves69, Ves71, Wat87, WSU19, Wit75, ZCPP13, vBP19]. Numeracy [GB19, II05, IB05, Men07, Zev04, Cra18, Edw84, GCG^{+}92, HMG16, MB16, TE14, ZC19, GB19]. numerals [VS23]. Numeration [GL98, AR82, AV76, BJ82, BJ88, SHAC01, AV76]. numeric [SK13, WR96]. Numerical [Ber06, Flü05, IB05, MGV^{+}25, Une74, Gla68, HVN^{+}24, HC95, LBT22, LL99, LLHL13, El 18, Urb93, Wri94, JSV75]. numérique [Gla68]. numériques [Une74]. numerosity [LHS19]. nursery [WDvO19]. nursing [CW14b, NPH99]. nurturing [JTP94, JLTM97].

Obituary [D’A10]. Object [LPG05, AP25, Bal88, SBDD24, TLM07, KM25]. Objectification [San11, MRSP14, Shv22, ZS15]. objectified [JMA19]. Objective [And79, And89]. objectives [And09]. objectivity [Dri25]. Objects [Ior17, Bal88, Duv83, FGG13, GFWL11, HRR23a, HRR23b, Jab23, MTBR16, SD95, Sfa91, TN15, WMPL08, Whi09]. objet [Bal88, TLM07]. objets [Duv83]. obligation [HEW10]. obligations [JWFC^{+}25]. observable [And09]. observateur [Add82]. Observation [Pas78, Add72, Add82, Pas78]. observational [Str78, WBM^{+}14]. observations [Del68]. Observer [Add82]. Observing [BTvG^{+}17, SK07]. obstacle [Duv83]. Obstacles [DP99, Abr01, PK23a, Sie87, YC90]. occasioning [Lev22]. occasionnées [Gro77]. Occasions [TD02]. occupation [Als24]. October [Van93b]. Odd [Zaz98, MYZ23]. Odds [Zaz98]. Offer [Mor06, IAP19, NCC22]. official [AC21]. Oh [YK20]. Ohtani [LX16]. Oksapmin [Sfa15]. Old [De 94, Eva07, AG02, Dem93, HVN^{+}24, Kie01, Mat83, PD23, Rou68, Sch16, SD95, SVVO23, Urb93, WC08]. Olds [FB02, Dem97, DB86a, DB86b, LH18, SS93b, VVV22, Wri94]. Ole [Cot24, LT20]. Olfos [RSD23]. Oliver [MND23]. Olsen [HN95]. Olwen [dF12]. Olympiad [Ano70, Ano71c, Ano72b, Ano74c, Ano74d, Ano77c, Ano77b, Ano78d, Ano76h, Ano76g]. olympiades [Rom74]. olympiads [Ano69, Bec80, Ger84, Rom74, SY79]. omnipresence [KM25]. Once [ZG01]. One [Ano00h, De 24, BR78, BT04, DGFT13, Pla14, Dho93, DG03, FPS16, FC12, GMH23, Lev13, Pal20, RL00, SW95, Van76, VMN11, Vla02, Wai70, Wea90, Ano00h, PM97]. one-step [SW95]. oneself [ZW11]. Ongoing [DS06, HHDG16, Pai16]. Online [PY22, Sfa15, FWCW22, HHT23, MND23, Nuc24, OAS25, RSRVRVVR23, RL11, vdS11, vdHPKR13, MND23]. ont [Rob86]. Onto [GBR05, EG23, MWVE09, RNRVM23]. Onto-Semiotic [GBR05, EG23, MWVE09, RNRVM23]. ontological [RP21, SD95]. ontologies [dFS13]. onze [Dem93]. Open [SKK^{+}23, DVV20, KL20, MW09, NCC22, REL25, Swe77, vdS11]. open-door [Swe77]. open-ended [NCC22]. Opening [Dim04, KL20, Wag12, Ban11, RL11]. Operating [LH96]. operation [BST81, BFG84, MEH20, Sla98]. operational [SK69, Tho94b]. Operations [GL98, CG21, FVJL^{+}19, GSH16a, GSH16b, Par25, VVV13, Ves69, Ves73a, Ves73b, ZR18]. operative [BS80]. operator [Car93]. Opportunities [CRL06, AMRM24, Bas17, CCP23, KK12, LW94, NT19, Pla14, WMPL08]. opportunities-to-learn [NT19]. Opportunity [WvdHPD15, Zas05, Dag96, GBC12]. Opportunity-to-learn [WvdHPD15]. opposing [Sch20]. oppositions [SPS17]. optimization [MF10, WR77]. Options [Abr01]. Oral [Vou16, ICR20]. orchestrating [LCK24]. orchestration [CEHM16]. orchestrations [DDB^{+}10]. Order [ZR18, Asg09, BLP^{+}14, GSH16a, GSH16b, HVN^{+}24, TT07]. ordering [YK24]. Ordinal [Jah01, BC82, VS25]. ordinales [BC82]. Ordinary [HPG05]. ordinated [Gre17]. ordinateur [AP75, BC92]. organisation [BC92]. organising [GMNB18]. organization [BC92, Gen70]. organizing [Asg09, Sty11]. Orientations [BR06, Gel08, KHLP08, LHS19, Zav21]. oriented [Bec77, Cho22, CHK22, KS69, LYBK23, Loc11, LK14b, PPSZ19, WW11, WR77]. origin [Her23, MA09, VMN11]. original [ABOM20]. Ornella [Art15a]. OSA [MTBR16]. other [AGL17, BN91, Fra13, GMH23, GS08, Sfa15, TE14, WFC23]. Our [RC04, Yoo25]. out-of-school [Cha22a]. outcome [LG16]. Outcomes [Che88, Baz90, DŽC25, HS14, KRHD15, TTTL10]. Outflow [Bec77]. Outside-In [Alb00]. overboard [Mat03]. Overcoming [YC90, FB17]. overlooked [CGD21]. overshadow [GS18]. Overview [Han00, Lan81, Led85, Men20]. Owens [Atw24]. Own [DEMW06, AS04, AS05, FM09, TGWN23].

P [Ber06, DKL^{+}25, Gai11, SMWM25, Uhl03]. pages [Ahm23, Cor22, Mah21a, NX23, Sfa15, Shv22]. Pais [Sri18]. palpability [Rad09]. Pandemic [BW20, AC21, BCZ21, CSW21, GG22a, HMSBT21, HHT23, HHL22, MJS21, XXL21, lDlzN21]. Panel [Ano68b]. paper [BP80, FS16, Gri78, HSZ17, KID25, TPHS07, WP24]. paper-based [KID25]. Paperback [Ahm23, Mah21a, Sfa15, Shv22]. Papers [Ano98e, Ano90c, Ano92d, Ano93b, Ano94e, Ano96c, Ano15, Ano16a, CL18, Fur22, SG25]. Papua [Atw24, Kae94, Sfa15, Cla83, Kae93, Sou83]. parabola [Tro00]. parabole [Tro00]. paradigm [PS21, VLY21]. paradigmatic [NH15]. paradise [Ort24]. Paradox [Add82]. paradoxe [Add82]. Paradoxes [DWMB05b, DWMB05a, MHH90]. Paraide [Atw24]. parallelism [TT07]. parameters [Šed76]. Parent [TY04, APS22, Tak18]. parent-initiated [APS22]. Parental [CBF07, Cra12, Tao86, KLLJ17, WZC21]. paring [GCFS23]. Paris [Rut02]. Park [Jon16]. part [ČK18, CRPPRG16, DKBD19, DKBD21, EPQH18, GP17, SR92, SM21, Sko88, Wol83, DB86a, DB86b, DWMB05b, DWMB05a, EH05a, EH05b, HH78, Had78, Noe80a, Noe80b]. part-of-part [GP17]. part-whole [ČK18, CRPPRG16]. partial [BMPT24, MGD^{+}23]. Partially [RDH10]. participants [Ano68a]. participate [Jan09b, Nol11, PKP^{+}18]. Participation [Boa99, BF21, EH20, ERS19, GABB25, GMHG09, HP08, HMSBR19, IWF^{+}15, MBF25, VN19, ZC22, Her02b, Bro11]. Participatory [Adl97, Tzu07]. particular [FC08, MP84, TV81]. partitioning [CN00, DH90, MGD^{+}23]. Partitions [Ham06]. partnered [Kie01]. partnership [Fur19]. partnerships [Jaw03]. parts [MBF25, Tab13]. party [KM18, Pim95]. Pascal [OCA13]. passage [Dié71]. passion [Rot17]. past [RFK07, SS17]. path [GLZBZ15]. pathology [ZK21]. Pathways [FB11, DHD23, GP17, Obe19]. patriarchal [Mog17]. Patricia [Sch24, Atw24]. Patrick [Mah21b, Mah21a]. Pattern [Pre07, DVF18, EJ13, FS16, FF17, JE14, El 18, PPDS^{+}25, Riv10, SWVO23, WC08]. Patterning [JNF25, VWV21]. Patterns [Bar03, FZ05, MHDC04, BA25a, Bra14, Dib19, GG22b, HWJ20, Jur77, Kle12, KBBK20, Sta89, TTL^{+}17, TJHD11, WC08, Wil21, ZL02]. Paul [Goo20, HJ12, SJS12, MDU^{+}25]. Pausing [Nol11]. pavages [De 94]. PDF [Shv22]. Pedagogic [Chr00, BBF71, Kle02]. Pedagogical [Kun12, AP25, Bra13, BOL24, CLSM17, DVT^{+}18, Ern84b, FC12, GSW22, GS18, HR74, HK84, Hem10, KBK21, LW14, LCD90, LR20, MM11, MPZ16, NBZ12, Pal20, PGM13]. pedagogies [Lam15, YK24]. pedagogised [TE14]. Pedagogy [Zev06, HN92, LT23, Siu93, Sol95, Swe95, ZB19]. peer [GB23, RP21, TJ94]. Peers [SS05a, PBCPP25, TTL^{+}13]. Peircean [Ott06, Ott11]. pencil [HSZ17, TPHS07]. pensée [Ano72a, De 94, Van69b]. Penteado [LT20]. People [De 24, Mon01, SRP21, Dör03a, Fit11, Fit14, HMG16, Ram03, Rot16, Bec80, Swe77, SY79]. people-pleasing [HMG16]. Pepin [Cha22b]. perceive [Phe10]. Perceived [CBF07, Tsa01, WZC21, YST22]. Perceiving [KRHD15]. percentage [Hah99, Van94, Van03b]. percents [SZP10]. Perception [MHDC04, BH11, CW14a, CSDK25, GL10, LL13, VD78, VD79, XXL21]. Perceptions [IE07, PB09, Tie02, BCFC90, For95, Har09a, MEH20, SK25]. perceptual [LH18, ML20]. perezhivanie [Ng21b]. Performance [Eva07, Jur06, Ade90, BCLH89, BHLH85, Buc87, Cle79, Csí16, DCHM19, Gal86, HH88, ICR20, KSH25, LC81, LSC20, LLL21, MWSS17, MNRR95, Pet91, PPC09, QM94, RSL17, RRP91, RO09, RBK25, SKR15, SLS74, SF23, VGJ19]. performance-driven [LLL21]. performances [AKL80]. performed [TTL^{+}13]. performing [Tru15]. Perhaps [STM06]. perimeter [ATDR19]. period [Dho93]. Periodical [BC05]. Periodicity [Sha98]. permanence [Sem84]. permitting [RR23]. permutations [KRHD15]. perseverance [Bar21, BE23, TP02]. persistence [BE23, Her02b]. persistent [McC14]. person [WSU19]. Personal [WG07, CZ11, Fra96, GB23, MK91, McQ86, OCA13, RG01, Rüe13]. personalities [SS88]. Personality [Hea81]. Perspective [Boa99, DG06, Ern06, Gal11, Hof06, LH97, ODV06, PBR^{+}04, SK98, Ste06, Bal88, Bau92, Bus96, CPPC23, CL84, Cle79, DG03, Ell86, Eng03, FT00, GL10, HS99, HM25, Hin15, HHB21, Ior17, KR25, KV23, Kni12, MA25, MAPHB21, MK14, Nol16, NX23, PT09, PJP23, RR11, RCH25, RG19, Rot16, RNC21, SD95, SPNH12, Sev22, Shv22, SBDD24, SS91, SS93a, SS88, Wil09, ZS15, vLXD18, RG20, Hes13, YK20]. Perspectives [AR07, CWL07, Ern99, Jur06, LPG05, Mos05, SLP06, AGP15, Als02, BWD10, Fri11, GDS22, GB19, HOS13, Hem10, Hud90, KZ16, LH09, LYHC18, MVCV24, MBV21, MTBR16, MAHK08, PGM13, PK23c, SP19, Sch20, Sfa13, SMWM25, SK13, TN16, VPR21, WHC09, Wil16b, ZZ11, lDlzN21]. pertinence [Fre82]. perturbed [CM92]. Peru [CRL06, CLSM17, Zav21]. PH.D [Her02b]. phase [BR22]. phases [HKSS^{+}25, VGJ19]. PhD [Ger10]. phenomena [Art97, TTS21, ZS15]. phenomènes [Art97]. phenomenographic [Neu99, VS23]. phenomenological [RT09]. Phenomenology [BG06, Usi85, Bro96, Fre93]. Philip [Atw24, MR17, Wan25]. philosopher [AMMS20]. Philosophical [Ern99, Kni12, Vil10]. Philosophy [Han98, Fre77a, Goo20, Ott07, Pai18]. Photovoice [HS13a]. phrases [HP25]. physical [SHAC01]. physiciens [Pis68]. physicist [Tav68]. physicists [Pis68]. physics [Ano68a, Van69b]. physique [Van69b]. PIAAC [TE14]. Piaget [Kap74, MDU^{+}25, Wit73a, Wit73b]. Piagetian [HC82, O’B77]. pictographs [WM01]. Pictorial [vLXD18]. pieces [Sch20]. pièges [Rev68]. pioneering [JMD08]. pipe [LaC14]. pipes [LaC14]. PISA [Sáe09, ARHS14, Bec24, CL23, KMT14]. pitfalls [HH81, Rev68, SL94, Ric20]. Place [NN97, GLHM24, HA23, KMN24, MB16, Sca75, TP14, Vol76, YC95]. place-value [YC95]. placement [HYC^{+}97]. places [Cha22a]. Plan [Kin74, AG14]. Plane [DVJ98, DPG89, Par88, Rob12, WE95]. planes [SP16]. planetary [dF25]. planned [Was22]. planning [Ahm23, vdKG22]. platform [LL13]. Plato [OCA13]. platonic [Vin75]. play [Bra13, Kry88b, SW94, Vog13, vO10]. player [ZP91]. pleasing [HMG16]. plot [Mie23]. plots [BBDD25, BBVD25]. Plotting [dF04]. pluralism [Pai18]. plus [Ano78b, HS13b, SLM79]. PME [NBD^{+}04]. PMM [Har73]. poems [HTM18]. Pohl [Sri18]. Point [Ott06, ART15b, HRR23a, HRR23b, Hoy85, Ott11]. pointers [CR09, Row95]. points [BCE^{+}19, BKH25, LK18, Ron15, TO13, TOT24]. Poland [Ehr78, Tur93]. polemic [Gos21]. policies [Lar86]. policy [HSM22, JB21, Ler14, PPSZ19, RH19, Swe77, TE14]. policymaking [Ngu24]. Political [PV12, Ern16b, Kol16, Kol18, Pai16, dF04]. political-technical [Ern16b]. Politicizing [Gut07]. Politics [PF11, GPS24]. polygons [BML21]. Polynomial [Riv07]. polysemous [Kon18]. polytechnic [Fri21]. polytechnique [Fri21]. poor [AK19]. popularization [Jin07]. population [AP75]. populations [AP75]. por [Cas98]. portant [Dum82]. portraits [Han12a]. PosE [LH22b, Cre03, MP24]. Posing [DS17, Eng97, LH22b, XM17, BR22, BLC20, Bon13, CMW^{+}13, CHM16, CL20, ÇA20, CMdR18, Ell13, GLH20, HWJ20, Kar16, KL20, KK13, Kon20, KSH25, kSL13, LSC20, LLZ20, RCH25, SB20, Sil13, SEC13, SV13, TH13, VS13, VP13, VSS20, ZSWC25, dPH13, rA20]. positieve [Gre17]. position [Rev71, VS25]. positionality [FB11]. positioned [Kni12]. Positioning [EMT06, HEWJ^{+}15, DE24, GLHM24, Nor15, WHE09]. positive [Bag10, Bur80a, Gre17]. possibilité [Dag96]. possibilities [Dav71, Wil71]. possible [CR09, FP85, FTM81, GRR23b, Ger88b]. post [Kae88, KK10, MLL20, Rot12, Rot17]. post-constructivist [Rot17]. post-high [MLL20]. post-secondary [Kae88]. post-structuralist [Rot12]. post-triumvirate [KK10]. postcolonialism [Atw24]. posters [BH95]. postmodern [SB12]. poststructuralist [Kle02]. Potential [Bus05, McN00, Ric20, Bar21, Bar87, FFL80, KT15, Pal20, PvdHPR12, PC23, SSL25, SD13, Wal17, WPL24]. potpourri [Sch71]. pour [Add72, Art97, HK99, Kra02, Kry88b, Pis68, Plu79, Rev76, Ser68, Tro04]. pourcentage [Hah99]. pouvant [TLM07]. poverty [Bat13]. Powell [Fit11]. Power [Tak18, AV08, Han12b, Law16, MJS21, PM16, Sko12, VDWV04, YHL^{+}15]. Powerful [Kra93, Mur24]. pp [Kae94]. pp. [Sri18]. practicable [NK17]. Practical [Sfa05, Eve99, lRA16]. practically [Lev10]. Practice [AAT99, BBEG05, BC05, GS08, HPG05, Jaw03, Kap77, Llo06, MHRSU99, Woo12, AMRM24, AI07, Bak16, Bes12, BR17, BB03, Car08, Eng03, Ern86, ES03, GFJ13, GdA09, Gra04, GBC12, GB19, Har09b, HPS^{+}19, HMMG18, HPRW25, JGR14, Kra02, Lan13, LLW07, MDPW96, NJL25, NPH99, PA14, PNH98, Rad03, RDH09, SW92, Sut93, Tab22, Tch11, Tho84, TB09, TW09, VA18, Wak14, XC19, dPQ16, dF25, Web18a, Web18b, Wan25]. Practices [Bur98, Cha06, EMT06, Mes04, Mos04, Wat00, Bas17, Bon14, Bra13, Byu24a, Byu24b, CY23, Cre03, Ell18, FGG13, FA03, Har09a, HH11, HMSBR19, IWF^{+}15, KMT14, McC14, MJ15, Mor14b, Nol16, OCB25, PRR08, SA00, Sou83, Sty11, TLB^{+}24, TP10, VEBS24, WBM^{+}14, Was22]. practitioner [RH02]. pratique [Kra02, TB09]. pratiques [PRR08]. praxeological [AAJW25, HHB21]. praxeologies [BGT17, GBE13, ZHT25]. praxis [SB12]. Pre [AG21, BHN10, DC92, DH90, GCFS23, GSW22, JMRW22, KZ94, LK16, Lee17, LK14b, OVKR25, PSH^{+}18, SF23, YSL22, vdKG22, AV22, BKK22, Bek10, Bis79b, BKH25, DP92, DVT^{+}18, Ell13, KPH10, KN21, MB16, MvdKM13, MLK16a, MLK16b, Phe10, SG14, Sin03, Wil21, dVSKvO18]. pre-constructed [Sin03]. pre-kindergartener [MB16]. Pre-paring [GCFS23]. pre-school [DP92]. Pre-schoolers [DH90]. Pre-service [AG21, BHN10, DC92, GSW22, JMRW22, KZ94, LK16, Lee17, LK14b, OVKR25, PSH^{+}18, SF23, YSL22, vdKG22, AV22, BKK22, Bek10, BKH25, DVT^{+}18, Ell13, KPH10, KN21, MvdKM13, MLK16a, MLK16b, Phe10, SG14, Wil21, dVSKvO18]. pre-technological [Bis79b]. preapprentice [LaC14]. precalculus [QM94]. precision [RC20]. precollege [MWSS17]. preconceptions [Yan11]. predict [SWVO23, TS99]. Predicting [JR72, WM01]. Prediction [BC05]. predictions [KK12]. predictive [OMB20, VDWV04]. predictors [LHS19]. Prediger [Bak16]. Predominance [DVJ98]. Preface [BE03, Fre76b]. Preference [Gra91, DVV20, VVV22]. preferences [Lev10]. Preformal [BK91, PM16]. Preliminary [Ano81c, Ano84d, Mai85, DOV24]. premier [Bui76, Fre69a, Gla76b]. première [BZ01]. premières [Bic82, Com80]. premiers [Rev69]. Préparation [Kry69a, Ayd90, Mah21a, Mah21b]. préparatoire [BC78, BC82]. Preparatory [PC98, BC78, BC82]. prepare [Isa87]. preparedness [VH16]. Preparing [WKB97, SGC22, Kry69a]. Preschool [VS23, KBBK20, LLHL13, SWVO23]. preschooler [Tar09]. presence [FT00, GD17]. present [Ano74b, DH20, Fle71, Hyl24, RFK07, SS10, SS17]. present-day [Ano74b]. presentation [AWL15, Add76]. presentations [BJ22, ET95]. presented [SD94]. presenting [Add76, KB11, LW14]. Preservice [Gra99, HDM07, Llo06, MHH90, NC06, Son16, TG89, CLK22, CPK08, CPP09, Cre03, Gli91, Kle12, MCL25, Son13, Tha10, GLHM24]. preserving [SHG21]. Presmeg [JV17]. Press [Sfa15]. Pressure [Tou90]. pretty [ZG01]. Preußen [Dam87]. preuve [Bal87]. previous [Gli91]. primacy [Dar82, Sch83]. primaire [AG02, DS02, Lev92]. primaires [FBM71]. Primary [Ass05, BML21, Dou99, Eng06, XM17, vdHPKR13, Ano14b, AG02, BHHS09, Bar03, BMR25, Baz88, BJ82, BJ88, Big70, BHN10, Bus96, BBFG99, Cho22, DS02, DOV24, DVF18, GG22b, HRR23a, HRR23b, HOEF89, JMRW22, KM24a, LBH23, Lev92, MM20, Di 19, NCMO15, PG20, SLP^{+}24, SG14, SA00, Tre91, TR24, VVV15, Var76, VA18, Wal12, WH24, XXL21, YST22, dVSKvO18, vBP19, FBM71]. primary-school [DS02]. primes [ZK15]. principle [BS80, Fre77b, NBE^{+}12, Sem84, SDVT25]. Principled [WD05]. Principles [Bel93b]. prior [Wil16a]. Private [Hus87, Cas04, ERS19, FA03, YM16, ZWC^{+}24, Cop98]. Privées [Cop98]. Probabilidad [Cas98]. probabilistes [LD88]. probabilistic [CM92, Eng75, Eng76, FG84, FNM91, Gre01, RBK25, VDD^{+}03, LD88]. probabilité [GT95]. probabilités [FPM69, FBM71, Gla76a]. probabilities [CZ11, FBM71, Par25, RBK25, SVLO20, Gla76a]. Probability [Cas98, FPM69, Kaz07, BN77, CV16, EK10, FFL80, FYAE12, FG84, Fre74a, GT95, HK84, JLTM97, KR73, Mor21, PK23a, Par25, Sha77, SVVO23, Var72]. probably [SWVO23]. Probing [CWS93]. Problem [BG05, BR78, Cal94, CB05, Cha97, CWT12, DG06, Eng97, Jer73a, Jer73b, JS01, Jur06, Kap01, Kon20, LT86, MDD04, Mos05, NHN03, Pug04, SE93, Sil13, SEC13, SLJ69, XM17, Yer00, bMYT98, dPH13, AKL80, Aud83, BMR25, BR22, BFG84, BLC20, Bju04, BN91, Bon13, Buc87, Bur80b, ČK15, CMW^{+}13, CHM16, CL20, ÇA20, CV09, CCW95, CMdR18, Cha22a, Chi98, CV16, CWYP92, CW14b, CF00a, CSK12, Dav71, DP92, Ell13, Ern16a, Eva99, FV13, FZ07, FD24, FKC17, FC08, FF21, FM09, GG96, GGR02, GLH20, GS18, HKSS^{+}25, HH81, HWJ20, Hin15, Hus87, IPJ25, Isa87, JSV75, JM74, JC22, Jon82, JI15, Kar16, Kie01, KK13, KSH25, Lec92, Lee17, Les81, kSL13, LSC20, Lib77, LLZ20, NT75, NCC22, Noe80b, Nun96, O’B75, OK13, Ott94, Pal08, PCP19]. problem [PT09, RR17, RCH25, SLP^{+}24, SB20, SB97, Sch13, SLC95, SV13, SGCA23, SJ04, Sty11, Tar09, TH13, UB23, VS13, VP13, VVV22, Ves76, VVM22, VSS20, Vou12, WSU19, WR77, Whi09, Wit71, ZP91, ZSWC25, ZM91, rA20, vdHPKR13]. problem-centered [CWYP92]. problem-oriented [WR77]. Problem-Posing [Eng97, Kon20, Sil13, SEC13, Bon13, CMdR18, RCH25, SV13, VP13, VSS20]. Problem-Solving [CB05, LT86, SLJ69, BMR25, Buc87, ČK15, CW14b, FZ07, NCC22, Noe80b, PCP19, PT09, Sch13, SV13, Tar09, UB23, VSS20, Whi09]. problem-structure [Noe80b]. problematic [Gal11, Gal19]. Problematik [JSV75]. Problématique [Ser76]. problématisée [Her11]. Problematising [Lle16]. problematized [Her11]. Problematizing [Wei17, Osi24]. problème [Aud83, ZP91]. Problèmes [Ano72a, Cal94, De 94, Del69a, KC73, Lev92, Mar69, SB97, Sen96, TSV82]. Problems [Cha06, CH02, De 94, DVJ98, Duv06, GBR05, Gor98, GD99, LH22b, Pim90, TP02, Wat00, Wil71, YG99, Abe78, Ade90, AG83, AG14, BST81, BFG84, BSV95, Bor86, CHM83, CV16, CGD21, Cre03, DVV20, Del69a, Dom11, DS17, Ell86, FQL^{+}17, Fey89, Fle71, Fre81, Fro99, GBMB70, Gli91, HC95, Hyl24, JR72, Jer73a, Jer73b, JHC14, Kal83, KPS83, Kle71, KL20, KO21, Kon20, KC73, Lan86a, LCD90, LG13, Lep90, Lev92, Lin90, Loc11, MF10, Mar69, Mar86, Di 19, MJ15, Nes76, NK77, NH94, Ott20, PGE09, PvdHPR12, Pol68, RO09, SSPS20, SLP^{+}12, SKR15, SAR19, SKK^{+}23, SLS74, SF23, Sen96, SD94, SG25, SW95, SGCA23, SOC86, Sta89, SRR22, Sun11, TSV82, Van94, VM13, Wha96, Wol83, dKBvdS22, lRA16]. problems [vLXD18, Ano72a, Par88, Ser76, Sel00]. Procedural [Vou12, dLT08, GMH23, LDM92, PM00, Sáe09]. procédurale [LDM92]. procedure [PvdHPR12]. Procedures [ABvP02, SS05a, Dem97, FZ07, Fos16, JMA19, MS19, RL12]. proceedings [Str08]. procept [Kid08]. Process [Bal87, DEM98, Kap01, PDV07, Riv07, Sha98, BR22, Bel79, BBAR23, Bju04, Boa93, BSV95, BDHN92, Bre22, Czo18, Dam87, DPG89, FBA10, GP90, Gal81, HMCT25a, HMCT25b, HGM15, HSZ17, Jak71, KT15, Kle02, Kry68, Lun77, MW19, RP21, Rob12, SMSW13b, SBDD24, Str84, Str85, UB23, Wit01, ZS15, Aud83, SMSW13a, MR17]. Processes [Alb00, BG05, Dah04, GBR05, OB06, OPC^{+}98, Pug04, THS06, Vin97, Bec77, CSDK25, EG23, FDW19, Han20, Ior17, KWF15, KAJHB10, KC22, MH97, MM25, Mey10, OB11, PR11, PW22, RW11b, RCH25, Reu23, SJ17, Sch16, Sch13, Sfa91, ST20, Tro00, Whi09, ZSWC25, vDDDB21, RW11a]. processing [CCW95, SOC86, SLW83, Wal12]. Processus [Aud83, Bal87, Kry68, DPG89, Tro00]. produced [MG00b]. producing [DVC19, NK17]. product [MP24]. Production [BZ01, IIMR^{+}11, RFP14, BZ01]. Productions [WA04, DS02, Deb06, PM16, WB12]. productive [Kle02, LG16, Ryv04, SM21]. products [GG22a]. Produire [Rut02]. profess [LW14]. Professed [Spe05]. professeurs [VD79]. profession [Dam87]. Professional [Ano04d, KBK^{+}17b, TS18, TLB^{+}24, ABOM20, BR17, Bro14a, CWH19, DK15, FB16, Gel08, Hah14, HPS^{+}19, HMMG18, HMSBR19, HMNWB20, HHDG16, Høy07, KL82, KM24a, KR09, PK23b, VC19, YSL22, dPQ16, KBK^{+}17a, YK20]. professionalisation [Dam87]. Professionalisierung [Dam87]. Professor [Ano16b]. proficiencies [Zep81]. proficiency [Cor22, EPQH18, Isa87, JNF25, RO09]. profile [DT08, Way94]. profiles [MBF25, Phe10]. Profound [DR13]. Program [Lev71, PC98, CN25, CPP09, Eve99, GS72, HMNWB20, Loc92, SS93a, Vil10]. programmation [DG89]. programme [GR71, HHDG16, HN08, JN15, PT09]. programmes [Une73]. programming [BB25, DG89, GBR22, Kyn93, Nos86, SD94]. programs [Baz88, BF96, Une73, Mah21a, Mah21b]. Progress [Han98, WD05, Lle16, RWZ15]. progressing [Dra14]. progression [BBG^{+}17, DDH25, El 18]. progressive [MW00, Tre87]. Project [VCS95, CWYP92, HS14, Lan81, IB05, Rog85, SA07]. projectiles [CGGGG79b]. Projection [CDGG76]. projective [CDGG76]. projects [DMHRV21, HI94, HB69]. prologue [TN16]. promise [McC14, VCD24]. promises [Cra18]. promising [vSBvR00]. Promote [MHDC04, Fra13, SD13, ZC22]. promoted [HMT13]. Promoting [BB25, HHDG16, GK20, HHS00, RPP24, SKPS16]. prompting [SKR15]. prompts [DS17, MFG22, SKK^{+}23]. Proof [Ber06, Bur05, DH20, HJ93b, Han00, HP03, KD14, Mar00, MMBD05, Miy00, Siu93, Sol06, Sty07, SS09, Uhl02, Uhl03, WA04, AS11, AK19, Bel76, Bro14b, Bur02a, Cha93, DAJ20, DW17, Dho93, DRR02, FG17, FC12, GD17, HWR23, Hem10, Her02a, IIMR^{+}11, Kom17, KJ22, KLwK24, LG13, LR20, LCW20, MP14, MRFW^{+}12, MGD^{+}23, Miy17, MFJ17, Moo94, NK17, Ott94, Ped07, PvdHPR12, PC23, Ram03, RG01, Rob12, SPC^{+}16, Sho21, Sim15, SS17, Van77, WFCW15, WHC09, YL08, Yan12, Bal87, Han98]. proof-based [WFCW15]. Proof-Generated [Bur05]. Proofs [LZ08, MG00b, DH20, Kom16, Kon21, Kon22, LW14, Ler85, MPS22, MRW14, RP21, RIA17, VE24, Web01, WMR11, WT22, ZWMR16]. propedeutic [CLR16]. propédeutique [CLR16]. Proper [NH15]. properties [BC82, CLR16, NHO22, UMC24, Usi74, VVV15]. Property [MT15, Ber08, DL14, WH24]. Proportion [AP07, Sin00, Son13, SVVO23, VS79, VLY21]. Proportional [BGW^{+}14, BCFF^{+}98, DLK14, TP85, HHDG16, KPS83, MW09, Noe80a, Noe80b, SR92, SVVO23, VWV21, BGW^{+}16]. proportionality [VRP^{+}12, Hah99]. Proportionnalité [Hah99]. Proportions [Tou86]. propos [Sch92]. proposals [Ano88c]. proposed [Zav21]. Proposing [JDK^{+}24, VH16]. Propositions [Ano68c]. propriétés [BC82, CLR16]. Prospective [Ano98a, Ano00a, Ano05, ABOM20, CRPPRG16, Gel98, Jan09b, LBH23, PPPP14, PC98, Rut98a, Rut98b, Rut99a, Rut99b, Rut99c, Rut00a, Tsa99, Tsa07, XM17, Yan11, Ano02a, Ano03a, Ano03b, Ano04a, Ano04b, BOL24, CN25, Cla12, FJC95, Gal82, KBK^{+}17a, KBK^{+}17b, LNC23, PGM13, RPP24, SMFL19, Sch23, Sch25, SGB21, Shr10, Nic02]. protagonist [BF21]. protest [Che23]. Prototype [Jon18]. Prototypes [Pre92b]. Prove [Dre99, Blo03, FG17, HHS00, vSBvR00]. provide [SKN21]. provided [HC95, WvdHPD15]. Providing [Jon00, Sut89]. Proving [Her93, Lab00, AI08, BK91, DWP21, FM09, Gal81, HWR23, Her02a, NK17, PR11, Rob12, SMSW13a, SMSW13b, VV25, MR20]. proximal [GGR02]. Prozeß [Dam87]. Prussia [Dam87]. Prytz [Hus25]. Pseudo [Vin97]. Pseudo-Analytical [Vin97]. Pseudo-Conceptual [Vin97]. psychoanalysis [KLBMK^{+}13]. psychological [HR74, HK84, Lor82]. Psychology [Fre78b, Ler01, LH06, Bal90, GBC12, RW19, VM78, Wit73a, Wit73b]. Public [Cop98, FLT14, ARW22, ERS19, FA03, JB21]. publication [MW19]. Publications [Ano76f]. Publiques [Cop98]. publish [Bec19, GDS22]. published [GK20]. Publisher [Ano88f, Ano88g, Ano89b, Ano92e, Ano96e]. Publishers [Mah21a]. Publishing [Cor22, Ahm23]. punctuation [BC08]. Pupil [Flü05, Fre75, Di 19, Pet91, Bar87, Bel76, BCFC90, BC78, BC82, CVN21, DS02, Hoy82, OVKR25, PB00, SS88, Vek71]. Pupils [Dah04, De 94, Eva07, Aud83, Bal80, BNPG97, ČK18, Her11, KC73, LJW13, RT85, Sen96, Siw73, SOC86, VD78]. purchases [DMHRV21]. Pure [Gri68]. purely [Lec92]. Purifying [Win11]. purpose [Row08]. Purposes [BR06]. pursuit [BSW16, Lle16]. Putting [MLK16b, MLK16a]. puzzle [BA23]. pyramid [MB09]. Pythagorean [MRSP14].

Q [WM09]. quadratic [HC14]. quadratics [Wil21]. quadrilaterals [FDW19, Has97, Mon00, PY22]. qualification [IMRS07]. Qualitative [ZBBZ85, Fin22, JV17, VS23, VS25]. Quality [ACJO04, HBH07, LS20, NCMO15, AMMS20, EMRI22, Kle02, KBK21, PK23b, SC93, VSS24]. quantified [SM22a]. Quantitative [Tho93, Eke77, JM18, NBE^{+}12, QP22]. Quantities [EJDT06, CHK22, Joh15]. quantities-oriented [CHK22]. quantity [PPDS^{+}25, VS25, WD10]. quantum [WS25]. Quartet [MMC22].

Quasi [SO73]. Quasi-Child [SO73]. que [SLM79]. qu’elle [Ser68]. Quelques [Add72, AV76, Del68, KC73, TLM07, Van69b, VD78, Del69a, VD79]. Quels [Rob86, Tro04]. question [Bag73, Mau85, OK13, Sfa15, Wed99]. question-and-answer [Bag73]. Questioning [Nic98, PKP^{+}18, ZC22]. questionnaires [Plu79]. Questions [SS05a, Add76, FWCW22, RSRVRVVR23, SEC13]. qui [Kry88b]. quite [PAW15]. quotidienne [TB09]. quotient [MGD^{+}23].

R [Ber06, Gai11, Rut02, Tsa02, DT08]. rôle [Kry88b]. Race [MVCV24]. Rachel [Hun86]. racial [Als24]. Racialization [Osi24, MVCV24]. Racialized [Bul24, LT23]. racializing [YK24]. racism [ISJ24, MVCV24, PAR^{+}24, WBM^{+}20, ZK24]. Radford [Shv22]. Radical [Ste98, TR11]. Raimundo [RSD23]. Raising [DEM98, DEM04, Mac00]. raisonnement [AM97, Duv91, Ric04, Siw73]. Raisonnements [SB97]. random [LD88, Lec92]. randomness [SWVO23]. rankings [AMMS20]. rapprochement [Nar16]. Rasch [II05]. Rate [YG99, HP12, Joh15, KPS83, Sch92, Tho94b]. rates [CS94, VR90]. Ratio [AP07, Car93, Noe80a, Noe80b, Sin00, Son13, Str84, Str85, VS79]. Rational [Mos05, Car93, DVT^{+}18, KML22, MGV^{+}25, PG20, SK13, Ste69b, VVV15, ZC22]. Rationalising [Lun77]. Rationality [Bro16, LH06, Büs24, Pre92b, UB23]. Re [Rot12, WHE09, KMN24, LK14a, Lle12, Nar16]. re-contextualization [LK14a]. Re-mythologizing [WHE09]. re-rendering [KMN24]. re-storying [Nar16]. Re/writing [Rot12]. Reaction [DDLZ97, LQV04, VVV13]. reaction-time [VVV13]. reactions [SB97]. read [ABOM20, MRW14, WMR11]. reader [WW11]. reader-oriented [WW11]. readiness [HC82]. Reading [Dim04, Öst06, Ber19, BP22, Bro08b, KCS22, MAM24, PR08, RNC21, YL08, Yan12]. ready [CH02]. Real [Jur06, AWE22, Bal88, Ber08, Boa93, Gra20, HR00, Jab23, KO21, KCS22, Mac21, NCC22, Pai15, SAR19, Shr10, SG98b, Usi74, ZCPP13, ZWMR16]. real-life [AWE22, NCC22]. real-world [KO21, KCS22, SAR19]. realisation [Bla78, Ehr78, El 78, Fey78, Fre78c, Fre78d, Gun78, HV78, How78, Kap78, Ohu78, Pur78, Rev78, Ros78, Sha78, Wil78a, Wil78b]. Realistic [CH02, GD99, Tre93, Van03b, WKB97]. réalité [Hah99]. realities [Fit14]. reality [Bar87, Bau80, Bos93, Hah99, dSBM21, Edd18]. Realization [MDD04, HRR23a, HRR23b]. Really [Rad11, Fit11, PAW15, Por90, Van77]. Reals [Rob86]. reanalysis [FB02]. Reason [MC01, AGP15, BZBM15, BLP^{+}14, ČK18, Pre92a]. Reasonable [AR07]. Reasonableness [AR07]. reasoners [SR92]. Reasoning [AS05, BH05, BCFF^{+}98, BG05, DGA05, FG97, Fis99, Jon00, Lit00, Wat01, AG21, AKL80, AS04, AP25, AM97, AE08, BGW^{+}14, BGW^{+}16, BS25, BA25a, BHHS09, Bar21, BNPG97, BLP10, BDS^{+}17, BEBL^{+}24, Car18, CN16, CH13, CPPT^{+}13, CM92, Daw83, DC17, DVJV02, DLK14, Duv91, EEST20, ES96, ES21, GLZBZ15, GBTL23, HL16, Had77, HH78, Had78, Han12a, HHDG16, JK17, JM18, JDK^{+}24, JE14, KPH10, KPS83, Kol21, KJ22, Lit03, Lit08, Mak14, Mar86, MPdP17, MMP15, MW09, El 18, NP86, Noe80a, Noe80b, NBE^{+}12, OVKR25, OMB20, PAW15, PC10, PHS12, RW11a, RW11b, Ric04, SLP^{+}24, Sim96, Siw73, SV09, SWVO23, SVVO23, Tho93, TS99, TP85, Tru15, UOF25, VRK24, VDD^{+}03, VWV21, WM01, Wat02, ZW21]. Reasons [NN07]. Rechenfertigkeit [JSV75]. Recherche [Gro77, Aud83, Fre82, Kon72]. Rechnung [Van76]. recipes [WT22]. reciprocal [Dom19]. recognised [GDS22]. Recognising [Bur02b]. recognition [CSDK25]. recognize [DH20]. Recommendations [Wat00, Ano68a, Ano71b, HSM22, LLW07]. reconceived [LS11]. reconceptualization [RT09, Sev22]. Reconciling [Eng03, Bur02b]. Reconfiguring [KM18]. Reconstruction [PDV07]. reconstructions [GP17]. recorder [Fie87]. recreated [CCP23]. rectangle [TP15, Yim10]. rectangles [Son13]. rectangular [BHLH85]. recursion [Kil85]. Recursive [Sen98]. redescription [Vou16]. Redirecting [Dra14]. Reducing [Haz99, HZ05]. Reel [SG25]. réels [Bal88, Rob86]. reference [BKH25, TV81]. referenced [MO94]. References [Kem81]. referent [Lee17]. Referents [MM00]. refined [SL19]. Refinements [RW11b, RW11a]. refining [Tha10]. Reflected [Han06]. Reflecting [GG22b, Kra93, LS20, Abt22]. Reflection [AAT99, GSE07, Kil85, TZLT95, Bag10, BC12, Elb03, GG22b, Güç16, HGM15, Str93a, Whe92, Hah99]. Reflections [ABK^{+}05, Ano02c, Bur06, Eva99, Sfa91, SK07, AV76, Bju04, BK91, FA03, HS14, KDM21, LK16, Roo71, Tep20]. Reflective [Sen98, ZCPP13, GP90, GE98, RPP24, Wit81]. Réflexion [Hah99]. reflexions [AV76]. Reform [IE07, MG00a, Sen98, Ban11, Bra11, BP80, BSF96, CP10, Cho22, D’A91, DVC19, Del69b, DKBD19, DKBD21, Goo07, How73a, How73b, Lun12, MRC18, PPSZ19, SS10, Swe77, SMWM25]. reform-oriented [Cho22, PPSZ19]. reformatting [RNC21]. reformed [VCD24]. reforming [SW92]. reframing [RNC21]. refutation [KJ22, PC23]. refutations [Kom16, LZ08]. Regarding [DWMB05b, DWMB05a, CEMP04, Kor03, MCL25, VO17]. regime [KMT14]. regimes [YK24]. register [TD12]. registers [MAPHB21, ST20, SPFGF14]. regularities [SWVO23]. Regulate [DEMW06]. Regulated [PBY03, KR09]. Regulation [Mal06, BH11]. Regulative [HMCT25b, HMCT25a]. Reification [Sla97, GE98, SL94]. Reifying [SBDD24]. reimagined [MJS21]. reimagining [RNC21]. Reinhard [Mal23b]. reinterpreting [BA23]. Reinventing [PM16]. reinvention [DC17, Fyh08, SHG21]. Rejoinder [Fre76a]. relate [KR11, Nol11]. Related [Bag06, EJDT06, And89, Art97, BA14, Bek10, ÇA20, Dar82, Eli85, FP85, FLT14, GRR23b, HP25, Han19, KO21, Kun12, Led85, SG14, Sie87, Tie00, TTL^{+}13, TLM07, VDTV24, WZC21, WMN22, WMN25, rA20]. Relating [AWE22, AOJ22, Dav71, FI17, JDK^{+}24, Mau85, RNRVM23, Sch11a]. relation [Arm69, BLP10, Cor17, Dör03a, FC08, GRR23b, GRR23a, Gol03, Hah99, HMSBT21, HMCT25a, HMCT25b, Kie01, KJvD20, PCP19, PC10, RL12, Roy81, SPNH12, TT07, Tou71, ZBBZ85]. Relational [DLK14, Bat13, CSP24, DVV20, Her23, PS21, VVV22, VYZ21]. relationality [Che23]. Relations [AF05, KR73, Lab78, Bas69, BP15, DG89, EVHRC^{+}25, GG22b, KBBK20, KBK24a, LOMG^{+}20, NBE^{+}12, TY04]. Relationship [MP14, Tch11, AAM21, Bes12, BWHM^{+}10, CRPPRG16, For95, FA03, GMH23, Hah99, Kru13a, Lev22, LL99, LLZ20, MY84, Ped07, RO09, ST02, SBDM69, Tho84]. Relationships [DCHM19, MK91, MM00, SMFL19, Zep81, dA00, AWL15, BML21, BBG^{+}17, BIBS^{+}19, Bro11, DG89, GB23, MZ12, OAS25, PGV22, SR92, SSA20, VP13, Wil16a]. Relative [MH97, Wai70, JR72, Mau85, PK23a]. relearn [SM22a]. Releasing [ZW21]. Relevance [Fre82, Eng69, Tal80]. relevant [LT86, Mat03, Men71, Mog17]. reliability [DOV24, Kno77, WBM^{+}14, Fre82]. reliable [HH78]. reliance [ČK18]. reliant [Wit69]. relief [GS20]. religions [CW14a]. religious [CW14a]. remarks [KC73, TS93]. remarques [KC73]. ReMath [AM14, HS14]. remember [Kar02]. remote [lDlzN21]. Renaissance [Rad03]. renarrating [RNC21]. Renato [LT20]. rencontrées [SB97]. rendering [JSV75, KMN24]. renewal [Bro10b]. Renuka [Edd18]. Repeating [TTL^{+}17, Dib19, PPDS^{+}25]. repetition [Han12c]. Reply [Lan86b, Pre23]. réponse [DP77]. Report [VVV13, Ano69, Bal90, Fre79b]. reporting [MO94]. represent [NT01, PK94]. Representation [FZ07, Mos05, Par91, BHHS09, Fin22, FKIA15, Jab23, Jon18, LCW20, MM20, MAPHB21, Par88, Pim90, Sty11, YS12]. Representational [DG06, BC96, HP08, MAHK08, Vou16]. Representations [Bag06, Cal94, PBCPP25, SR92, Arc03, BZA01, BP22, CO10, Cas69a, DC92, DT12, DL14, DK15, EG23, FV13, GdA09, Ior17, KZ94, MP96, MK97, MK14, Ott20, Rad14, Ron15, SP16, SBL22, TP22, TMP10, TP15, Wei91, Wil16a, vLXD18, Cal94, Cas69a]. representativeness [CM92]. represented [AWL15, CCRD22, ST20]. Representing [WM01, BIBS^{+}19, VS25]. reproducibility [ABM92]. reproduction [MM25]. Republic [Bec80, Swe77, SY79]. require [SRR22, VLY21]. requiring [DS17]. Research [Ahm23, BB03, Bus05, FZ05, Fre82, GPW05, Jaw03, Kon93, Kru13b, LPG05, LMH70, MJ15, Mor00, Mor06, Nis99, Sch24, See01, Sfa05, SS17, Tsa02, AMMS20, APS22, AS25, Aud83, Bak16, BCZ21, Beg69, BA25b, Boa02, BGT17, BAZF04, Bro08a, BDS^{+}17, CL20, CWH19, Dar82, Dar16, DKL^{+}25, Ell18, Eng03, ES03, FA76, Fre74b, Fre79b, Gol03, GK20, Had78, HM95, HS14, HTM^{+}23, HMRRTT23, HJ12, HPRW25, Hoy01, HN08, JN15, JV17, KR25, KFS01, Kil01, Kle71, KSS25, LT20, LH22a, LXT02, Lit08, Lun77, MCW25, MGI23, MR17, MA25, Nar16, NH94, NBA23, O’B77, PV12, Pai19, PH96, Pop88, PPSZ19, Pre16, PM19, RBWS18, RL00, Ron15, RTSK08, Sca71, Sfa13, Sil13, Sim17, SMW11, SEC13, SB12, Sto22, SMOR20]. research [TP22, Tur93, WBM^{+}20, WPA^{+}23, WP24, Wal17, WMM13, Woo96, Yoo25, Gro77, Mal23b, NX23, Sri18, MDU^{+}25, Sfa13, Bro23]. Research-based [SS17, JN15, Ron15]. research-informed [CWH19].

Researcher [Bar09, Ahm23, FB11, LYHC18, YHL^{+}15]. Researchers [Wal17, BGSW94, Fur22]. Researching [ABK^{+}05, PV12, RL11, Men20]. reservoirs [Bec77]. residue [TP02]. resistance [Bra14, Tou90, VC19]. resisting [Lam15, Wag12]. Resnick [Web18a, Web18b]. Résolution [TSV82, DC92, Cal94, Lev92, SB97, ZP91, Lev92]. resource [Ano74a, Cha22b, GMNB18, KM24a, PXTW17, Pla18, PY22, Tab22, ZHT25]. resources [APRS09, Chr00, Cra12, Ell18, Gue19, PPG95, RIA17, RH02, WJW11, Yer00]. respect [Van94]. respond [Son13]. Responding [Nic98, SMFL19]. Response [Led87a, MM00, Uhl03, Wil05, Bro12, Bro23, CHM16, Car08, Che23, Ola14, PR08, RC04, SJ17, SC93, SS03, XXL21]. response-ability [Che23]. Responses [CH02, Nar00, WW02, AWL15, DP77, GKMJ24, Ng21a, RCH25, SLC95, Son16]. responsibility [Han10, Han12d, ME13]. responsive [Fit11, LOW24a, LOW24b]. Responsiveness [OPC^{+}98]. Restrictions [BBEG05]. restructuring [Noe80b]. résultats [CGS76]. Results [BGSW94, BJ88, KM93, SA07, Sut93, CGS76, CG21, DOV24, GAP24, Han89, SM22b, Sin03, Str78]. retail [Zev11]. retention [McD89]. Rethinking [Bra11, HbL05, ISJ24, Rut94, Sta11, VLY21]. retrieval [WP24]. retrospective [MRC18]. Reuben [Pai18]. réunion [DeB97]. revealed [GT95]. Revealing [HP12, Sie04]. révélateur [Art97]. revelation [RM15]. revelations [HEWJ^{+}15]. révélées [GT95]. reversal [GCALB15, SGCA18]. reverse [BBAR23]. reversibility [SKPS16]. reversible [SJ17]. Reviel [Sid05]. Review [ABW88, Ahm23, All99, Ano98c, Ano98d, Ano00c, Atw24, Bak16, Bec86, Bja20a, Bja20b, De 24, Bur03, Cha22b, Cob04, Cor22, Cra85, Dre07, Ell18, Eng99, Fer15, Fit11, Fit15b, Fri21, FSY22, Fur19, Fur22, Gai11, GV83, Goo20, Gra20, GM10, Gre17, GB19, Gue19, Han12a, Han98, HSNHHE15, HJ12, Hes13, Hus25, Hyl24, IPJ25, JC22, Jon16, JV17, Kai13, Law16, Law24, Law00, LX16, Mah21a, Mah21b, MPS12, MDU^{+}25, Mal23a, Mal23b, MND23, MR17, Men20, Men23, Mor00, MR20, MH89, NX23, OM24, Ott76, Pai18, Rad11, Ric20, RC04, RSD23, Rut02, Sel00, Sfa15, Shv22, Sil25, SMWM25, Sta11, Ste13, Sto22, Str08, Tab22, Tep20, Tsa02, TR24, Wal12, Wal17, Wan25, Win11, Yam09, YK20, ZB19, dF12, vBP19, AAB03, Ala12]. review [Ano89a, AS25, ACS88, Art15a, Art19, Bel85, Bis80, Bro11, Bro23, CKS22, Cot24, Dar82, Dav07, Daw94, DKL^{+}25, Dre93, Edd18, Edw93, Ely10, Fit07, Fit15a, Fit17, FH88, Fre13, GK20, Han95, Han07, HK84, HTM^{+}23, HC82, HFD82, JS87, KSSO25, Kap09a, LT20, LP91, Lun82, MGI23, MHC24, Mur24, Nol11, ND88, Nun95, PD91, Pim91, Pim93, PW03, PH96, PW22, Pre07, RBWS18, RP21, Ric10, RC02, Rut10, Sch24, SJS12, SM22b, SBS86, Sri18, SMOR20, TTS21, TP85, Wal10, Web18a, Web18b, Whe96, Wit82, Wub92, Sfa13, Sri18]. reviewers [Fit09]. Reviews [LW83, AS87, BFG83, FH90, FKSS82, FJA86, GB88, LG87, MBV21, PB84, SF84, SM83, WB84, WQTV85]. revising [VE24, VEBS24]. revisited [BM08, FYAE12, NBZ12, NH19, TT07]. Revisiting [Car08, Kei93, Ryv06, TP14, GS08, RC20]. revolution [Hyl24]. Rewriting [Ste13]. rhetoric [HR00, Kaa07]. Rhetorical [Ern99, RP21]. Rhythmic [PA18]. Rica [CN25]. Ricardo [LQV04]. Rich [Lab00, DDB^{+}10, Fos18, HS99, HMSBR19]. Richard [Bag08b]. right [Cha22a, DG03, MRSP14, Woo16]. right-angled [MRSP14]. rigidity [Hoz81]. Riley [Lan86b]. Rina [Rad11]. Ring [SS06]. Rings [Fre79a]. rise [Bja20a, Bja20b, Zah15]. risk [ALO^{+}91, NBA23]. Ritual [NT19, HMSBR19, LSS19, McC14, MLL19]. Ritual-enabling [NT19]. Ritualisation [CS19]. ritualization [Adl19, HMCT25a, HMCT25b]. ritualized [VN19]. Rituals [HMG19, Woo16, Adl19, CCP23, HMG20]. road [HH81, Pat85]. Robin [MND23]. robotting [Sko12]. robust [CR09, Mur24]. robustness [RTSK08]. Robutti [Art15a]. Rods [Bja20b, Bja20a]. Roger [Cor22]. Rogers [GB19]. Role [AS05, Ano74b, Bur02a, Flü05, Lep90, ODV06, Sol06, Uhl02, Uhl03, Wat07, AS04, Als24, Arc03, AM25, ABM92, BF21, Bar21, Beg69, Bro14b, DT12, DWP21, DRR02, FvM97, FM14, GRR23a, GCFS23, HHS00, HB21, HH11, HN96, HM13, HKG89, Hol81, HMA13, Jou25, Kid08, Kla71, KOMH22, KCS22, Kry88b, LH18, MF10, MS09, MFG22, MGD^{+}23, NH94, Nol16, Nuc24, PR11, Rad14, RP21, RTSK08, RNP13, Sáe09, Šed76, Sla98, SHG21, Ste13, Thw69, TP10, VV25, VVV22, WH15a, WH15b, Whe92, ZS15]. roles [BdSCdA18, BC08, GE98, HS99, Høy07, KT15, LCK24, Sun11, Wit81]. Romania [Rom74]. Romany [SK07]. room [Tur76]. Roos [Dek87]. roots [Str84, Str85]. Rorty [Bag08b]. Rosa [LQV04]. Rotation [Gor98, SC93]. rote [Han12c]. Roth [MPS12, Bro12]. Roumanie [Rom74]. Route [Sla97]. routine [FT03, PGE09]. Routines [Gel08, LSS19, CCP23, EH20, HMCT25a, HMCT25b]. Rowlands [AAB03]. Rowman [Mah21a]. Roza [Sil25]. ruin [ZP91]. ruine [ZP91]. Ruiz [Mur24]. Rule [DGA05, Bar19, SSP04, WC08, Zaz99]. Rules [Rot14, Güç16, GSH16a, GSH16b, KB12, KLwK24, TS99, VDWV04, Zaz99]. run [Zah15]. Rupture [HSZ17, FG17]. rural [KMN24]. Russian [Kai13].

S [Jon16, Tsa02, DIW19, KR25, Dre06]. sa [Kra02]. Sacha [Gre17]. safety [CW14b]. safety-critical [CW14b]. Saharan [Osi24]. Same [Tsa01, Tsa07, BJ88, Sfa91, VDWV04]. Sameness [RS22]. Sample [CRL06, CZ11, CG21, LK16, RRP91, ST02]. samples [AGP15, BZBM15, GLZBZ15, MMP15]. sampling [AG21, GLZBZ15, MMP15, NH15, PAW15, SGRS24]. sand [Ger88b]. Sandra [Mur24]. Sara [SWG05]. satisfaction [GRR23a]. savages [Car08]. Saxe [Sfa15]. say [LJW13]. scaffolding [GMNB18, LCK24, VE24]. Scale [PSS^{+}19, CVN21, CWH19, FKC17, GAP24, HSM22, LLZ20, MCKP19, RH19, ZSL02]. Scaling [EEST20, HPS^{+}19, KZH^{+}19, NK17, PFSS19, RTSK08]. Scaling-continuous [EEST20]. Schantz [Web18a, Web18b]. scheduled [Ano74b]. schema [KV23, Wol83]. Schemas [Chi98, ČK15]. Schemata [FG97, Fis99]. Schematic [FV13, Ng21a, SJ04]. schematisation [Tre87]. schematization [GP17]. scheme [Sim15, Sty11]. Schemes [Riv07, GBR22, HN96, HP03, Hun86, Kid08, MPZ16, NH94, PD23, RL13, SL95, Ste83]. Schmittau [Bac21]. Schoenfeld [Mur24]. Scholastic [BHX97]. School [AR07, Ass05, Baz01, Dah04, DVJ98, Dou99, Eng06, EMT06, FPM69, FWCW22, GS08, HZ05, Jur06, Lev71, Lit00, Mes04, Miy00, Ngu24, Pre88, Prz05, SK07, SPFGF14, Sty07, STM06, Tie02, War06, Wat01, AI12, AAJW25, Arm69, AG02, Bal71, Bar03, BNPG97, Baz90, BJ82, BJ88, BZA01, BML21, BD72, Bes12, Bla85, Boa93, BBDD25, BBVD25, Bra14, Bur80a, Bus96, BBFG99, Cas69b, CDGG76, Cha18, Cha22a, Cha93, Cla12, CG21, Cra12, DP92, DS02, DVJV02, DOV24, DVF18, DGI23a, DGI23b, DH20, Dow96, Ell13, Eng71, FJC95, FKIA15, Fre74b, GC90b, Gar24, Gla76a, Goo07, GAP24, Gui17, Gut92, Hah14, Han12a, HS13a, Hay87, HWJ20, Her02a, Her11, HMMG18, HMSBT21, Hil71, HHDG16, Hol81, HC14, Hoz81, HHL22, JK17]. school [JGR14, JS01, Kap76, Kap70, Kar02, KK12, Kel71, KL82, KM24a, Kim20, KMN24, Kle71, KO21, Kom16, Kom17, Küc81, LS70, LM11, LLW07, Leu87, Lun12, Ma96, MGV^{+}25, MM20, MvdKM13, Mal93, MK91, Mar69, MG00b, Di 19, MGI23, MDPW96, MS20, MCPGE24, MY84, MLL20, MCL25, MAM24, MRC18, NHO22, Nim75, PG20, Par91, Pop88, PFSS19, Pre92b, Pro70, RCH25, RL13, RA21, Roy81, RDH09, RC20, Saj03, SGRS24, SM22a, ST08, Ste13, Ste68, Sty11, SW14, Thw69, Tie00, Tou86, Tre91, TP15, VRK24, VS13, Var76, Whi87, Wil16a, WS18, WPL24, XXL21, XH88, YST22, Yan11, ZER17, dVSKvO18, vdHPKR13, Lib71, SMWM25]. school-level [KO21, XXL21]. school-type [Mal93]. schoolbooks [Bar07]. schooled [TDGV09]. schooler [SM22b]. schoolers [DH90]. Schooling [Nol16, Che88, Cla85, JMRW22, Led92]. Schools [BBEG05, HBH07, Baz88, Cd04, Dav71, Fid85, For69, Gen70, Ham68, Kar10, LR81, Lyn68, Mal93, Miy17, NCMO15, Sek21, Str08, Tay91, dKV18]. Schoolteacher [PM15]. Schorcht [Fur19]. Schubring [Art19, Law16]. Schwarz [Rad11, Web18a, Web18b]. Science [Kry69a, RD92, Ano81b, BP80, HTM^{+}23, LMCG08, Wit95, Han98, Fur22]. sciences [Bag10, Feh69, Kry69a, Mur93]. scientific [HGM15, JSV75, MPS12]. scoping [MHC24]. score [WBM^{+}14]. Scores [Lie90]. scream [LM11]. scripts [KLwK24]. sea [Fre71]. séances [PRR08]. Search [Leu01, Str84, Str85, Vol86, Zev06, ARHS14, Kon72, Lle12, Sim96]. searching [Fit15b]. Sebastian [Fur19]. Second [All85, CWYP92, Daw83, HS22, Has97, Jon82, PRR08, Rev72, RG19, RG20, Rob90]. second-grade [CWYP92]. secondaire [Aud83, BD72, Cas69c, Cas69b, CGGGG74, VD78, VD79]. secondaires [FBM71]. Secondary [Bes07, BBVD25, CGGGG74, Cor17, DVJ98, JCM^{+}23, Miy00, PBR^{+}04, Prz05, Wat05, ARKK23, Ano69, ABOM20, Aud83, BKK22, BNPG97, BD72, BN77, Boo81, CLK22, Car18, Cas69c, Cas69b, CSP24, Che89, Cla85, CB17a, CB17b, DVJV02, DVF18, DSR^{+}23, FBM71, FT03, Fra96, Gal82, Gen70, GSW22, Gue08, Gue23, Gui17, Han12a, Her11, HC14, KPH10, Kae88, Kar10, KL82, KO21, Kom17, KBK21, KO23, Küc81, LR81, Leu87, LNC23, MGV^{+}25, MMC22, Mal93, MK91, MG00b, MCPGE24, Miy17, MLK16a, MLK16b, Nuc24, Orl71, PB09, PK23c, Rev72, Rüe13, RDH09, SLP^{+}24, Saj03, SMFL19, ST08, Sek21, Ste68, SG98b, Tou90, VVV15, VD78, VD79, Vil71, VR90, WH15a, WH15b, Wil16a, Wil21, XH88, vSBvR00, BBDD25]. secondary-education [Aud83]. secondary-school [MGV^{+}25, RDH09]. Secondary-to-tertiary [Cor17]. secret [Dom25]. Sectional [EJDT06]. sections [Ano84c]. seduction [BH11, Wag12]. see [GL10, HH11, LaC14, PAW15]. Seeing [Han12b, JC22, KBK24a, LNC23, MP84, Wil21, Par88]. Seeking [Ely10, MS14]. seemingly [VGJ19]. seems [ZG01]. seen [De 24, KHRK15]. Seizing [Zas05]. Select [LH22b]. selected [Jur77, TF95]. selecting [LH18, MJ15]. Selection [Kis86, JDK^{+}24]. Self [Fri07, HC95, Mal06, Ng21a, PBY03, BLC20, BS15, BAZF04, HGM15, KMT14, KR09, KSH25, LLZ20, Plu79, Rel96, SLP^{+}12, SAR19, Sol12, SMS22, TJ94, TTL^{+}13, VSS20, Wit69, XA20, ZC19]. self-assessment [BAZF04, TJ94]. self-checking [Plu79]. self-concept [Rel96]. self-efficacy [BLC20, KSH25, LLZ20, SLP^{+}12, SAR19, SMS22, TTL^{+}13, VSS20, XA20, ZC19]. Self-Knowledge [Fri07]. self-reflection [HGM15]. Self-Regulated [PBY03, KR09]. Self-Regulation [Mal06]. self-reliant [Wit69]. Self-schematic [Ng21a]. sellers [Hah99]. selling [TB09]. Semantic [BEW99, Nar00, SW95, WA04, FD24, Min83, NK77, NGR82, SLAY19, WB11]. semantically [CCRD22]. seminal [OM24]. semio [Ior17, Ior18]. semio-cognitive [Ior17, Ior18]. Semiotic [AS11, BH05, Ern06, GBR05, Hof06, Ong06, Ott06, SLP06, Sch13, SLR^{+}24, Win03, APRS09, CN25, EG23, FLM07, Fri11, FDW19, MWVE09, MAPHB21, MK14, Rad00, RNRVM23, San11, See11, ST20, SZP10, GFWL11]. semiotic/dialogic [FDW19]. Semiotics [Mor06, Pre06, Ott11, PR08, Sch11a, TR11]. Senegal [VD79, VD79]. senior [MMC22]. Sense [MM00, PK17, Sel00, WG07, Bar96, CPPT^{+}13, FC12, JTP94, JNF25, KO23, KOMH22, KJvD20, LL99, Moo13, Pal08, PPDS^{+}25, RFP14, Sch16, SM17, Sim96, Sla98, WD10, Wak14, Yan19]. sense-making [KO23]. sensemaking [BE23, MZ21, RP21]. Sensitivity [Jaw02]. sensory [CSDK25, GPS24]. sensory-motor [CSDK25]. Sensuous [RP07, Rad09]. SEO [MA09]. Seoul [XC19]. separate [Joh15]. separation [Van74]. September [Ano74b, Van93b]. Sequence [MD01, Prz05, BC92, Roh08, BC92]. Sequences [AS05, Ber06, Bur05, AS04, AS17, Sch11b, TL15, Vou16, Wei91]. sequential [AWL15]. serendipity [WP24]. Series [AS05, AS04, Car10, DL14, GMNB18, Mar13, MPGDA12, vBP19]. servants [SECP^{+}23]. serves [ME13]. service [AG21, AV22, BKK22, Bek10, BHN10, BKH25, DC92, DVT^{+}18, Ell13, Gra04, GSW22, Gut92, JMRW22, KPH10, KZ94, KN21, LK16, Lee17, LK14b, MvdKM13, MLK16a, MLK16b, OVKR25, PSH^{+}18, Phe10, SG14, SF23, Wil21, YSL22, dVSKvO18, vdKG22]. ses [TLM07]. Sesotho [Zep82]. sessions [GB14, PRR08]. s’est [Add93]. Set [Bag06, Kap76, Mon72, Ber08, CZ11, FB98, Hod96, TLM07]. Set-theory [Kap76, Mon72]. Sets [Ham06, Tsa99, WM98, ALR25, Bja20a, Bja20b, DeB97, Fre69e, Ott03, Tsa01]. setting [Jab23, KC22, MM25, Nil07, WK04]. settings [Cra12, Fra13, KM18, WR96]. seven [Mat83, SBA22]. Seventh [Alb00, PBY03, Ano91c]. Seventh-Grade [Alb00, PBY03]. Seventy [How84]. Sex [And89, Cle79, Kae88, Led85, CWT12, Eli85, FP85, Mal93]. Sex-related [And89, Led85, Eli85, FP85]. Sfard [HJ12, Wal17, Win11]. Shame [PAR^{+}24, JMRW22]. Shanghai [HHT23, WqWLR18, XC19]. Shape [VAM21, CSDK25, KZ16, KO23, Vol77]. Shape-shifting [VAM21]. shapes [BA23, GG22b, Leh23, Tri95, VO17, WE95]. Shaping [Sid05, ETC14, GC90b, HS14]. shared [TDEM13]. sharing [NCC22, SVVO23, Yoo25]. Sharon [Fit11, Lev13]. sharpens [BOL24]. Shedding [GM08, MCL25]. shift [AMRM24, LOMG^{+}20]. Shifting [AVM15, Sch23, HMNWB20, VAM21, Sch25]. Shifts [NP86, NJL25, PT09, SPS17, VGJ19]. Short [HbL05, Kae94, Wil05, LC13]. shortcut [TDGV09]. shortest [Pat85]. Should [FVJL^{+}19, Gra99, Kry88b]. Shu [Siu93]. Sides [Tsa07, Sfa91]. Sierpińska [Ano24b, Mor00]. Sierpiński [ATDR19]. Sierra [Wil78a]. Sign [Ste06, Ber04, FMA20, Gro77, SLR^{+}24]. signe [Gro77]. signed [NM98]. Significance [Pre06, HEWC10, SGRS24, ZZ11]. Signification [Sam85, Wal73a]. Signifying [Bro08b, PR08, RSS11, YS12]. Signs [Fri11, Rad00, Ior17]. Silence [Sie04]. Silver [Sch24]. simcalc [SG08]. Similar [DVJ98, Ham68, KV75, KAJHB10, Son13]. similarité [GT95]. Similarities [WL90]. similarity [GT95, LH18, Vol77]. simple [CSP24, Dub84, Fro99, Gra91, Jan74, Jer70, vLXD18]. simples [Dub84]. simplicity [Bar07]. Simplification [Gri83]. Simplifying [TER98]. simulating [ZP91]. simulation [AP75, ZP91]. simulations [HSM22]. simultaneity [EVR16]. since [Bla78, Bor21, Ehr78, El 78, Fey78, Fre78c, Fre78d, Gun78, HV78, How78, Kap78, Ohu78, Pur78, Rev78, Ros78, Sha78, Wil78a, Wil78b]. Sinclair [Art15a, Fer15, HSNHHE15]. Singapore [FZ07]. Singaporean [JHC14, VVM22]. Single [Mos05, PW22, CWT12, PDT^{+}12]. Single-blind [PW22]. single-digit [PDT^{+}12]. single-sex [CWT12]. sinusoidal [Car18]. Siqi [Mur24]. site [BS15]. Situated [Jur06, Tic12, KBK^{+}17a, KBK^{+}17b]. situatedness [NEM99]. Situating [NBHK07]. Situation [Bus05, Cop98, Flü05, NHN03, Bra68, Gos21, Kun12, LH18, LD88, Mes89, MW09]. situation-specific [Kun12]. situational [Fra96]. Situations [AM97, BG05, HPG05, LPG05, Ste05, ABM92, AWE22, Bal87, Bel76, Bra13, BEBL^{+}24, CYW92, DeB97, Gal11, Gal19, GRR23b, Jan81, JDK^{+}24, Lec92, NT01, Pre88, Sch13, SKN21, TSV82, Tre83, Vog13]. Situative [PBR^{+}04]. Six [Car99, Dié70, Jur77, Mat83, Sin00, Urb93]. six-years-old [Urb93]. sixième [CGS76, For69]. Sixth [CRL06, CGS76, For69, JHC14, LH71]. sixth-grade [JHC14]. size [BFG84, Sim20]. sizes [BCE^{+}19, PM19]. skepticism [Bro14b]. sketches [Sin03]. skill [Sle82, WE95]. skills [BMK21, Bro10a, DT08, Ern86, Ham68, HVN^{+}24, Hee85, HHDG16, JSV75, LDM92, LHS19, SMFL19, SGB21, SRR22, TJ94]. Skovsmose [Cot24, LT20]. sky [Ano84c]. Slope [WG07, CN16, MRCR11, Zah15, ZSL02]. sloppy [ZSL02]. Slow [ZK24]. Small [Kle01, BCE^{+}19, BHLH85, CY23, GGR02, Sha77, TS18]. small-group [Sha77, TS18]. smallpox [Gos21]. smoothie [ATDR19]. so-called [Bau80]. Social [BC05, CF00a, Gat06, GdA09, Har09b, Kol18, Mor06, Mor14a, MHDC04, QP22, YK24, AV08, AC95, Bau92, CBC25, GB23, GB19, Hoy01, JGR14, Lan13, LW94, MFG22, MK14, Nol12, PV14, Rad03, See11, SJS12, Ste91, TB09, ZHT25, All99]. social/institutional [AV08]. sociale [TB09]. Socialist [How80]. socialization [AMRM24]. Socially [GGR02, HO23]. Societal [Lin88]. Society [Rog85, Kae88, Kae90, Tay91, TE14]. Socio [CSRlR24, II05, KHLP08, ODV06, ÇA20, Kae90, Kol16, Kol18, MK91, TJ94, dA00, rA20]. socio-constructive [TJ94]. Socio-Constructivist [ODV06]. socio-cultural [dA00]. Socio-ecological [CSRlR24]. Socio-Economic [II05, Kae90]. Socio-emotional [KHLP08]. socio-mathematical [ÇA20, rA20]. socio-personal [MK91]. socio-political [Kol16, Kol18]. Sociocultural [KFS01, Ler01, DGI23b, JP09]. sociolinguistic [RG19, RG20]. Sociological [PA14, Bec24, Dow96]. sociology [Wil12]. Sociomathematical [SG14, VE24, HS99, LYBK23]. Sociopolitical [Edd18, Sri18]. soft [Ham68]. Softcover [NX23]. Software [Jon00, Mar00, Art97, RDH09]. soit [Ser68]. solid [Rog71]. solids [BHLH85, JS01]. Solomon [Bro11]. Solution [Tho90, BSV95, CV16, Lin90, PGE09, TTTL10]. Solutions [DVJ98, Tsa07, AI12, CHM83, CV16, Hin15, Hyl24, SKR15, SAR19, SLC95]. solve [AG83, AG14, ČK15, JMA19, VM13]. solver [FM09]. Solving [BSV95, BG05, Cal94, CB05, Cha97, DG06, GBR05, Gor98, HKG89, Jur06, Kap01, MDD04, Mos05, NHN03, Pug04, Riv07, XM17, bMYT98, Abe78, BMR25, Bju04, BN91, BA23, BR78, Buc87, Bur80b, ČK15, CV09, CCW95, CWT12, Chi98, CW14b, CF00a, CSK12, DP92, FV13, FZ07, FQL^{+}17, FD24, FF21, GG96, GGR02, GS18, HKSS^{+}25, HH81, HC95, Hew12, Hin15, IPJ25, Isa87, JM74, JHC14, JC22, JI15, JS01, Kal83, KB20, Kie01, KMA02, KBK24a, Lee17, LT86, Les81, Lib77, MF10, NT75, NCC22, Noe80b, Nun96, OAS25, Pal08, PCP19, PvdHPR12, PT09, SLP^{+}24, SSPS20, SB20, SB97, Sch13, SF23, SE93, SV13, SGCA23, SOC86, SJ04, Sty11, SLJ69, Tar09, UB23, VVV22, Ves76, VVM22, Vla02, VSS20, Vou12, Wha96, Whi09, Wit71]. solving [Yer00, dKBvdS22, vdHPKR13, TSV82, ZP91]. Somatic [BR06]. Some [AG83, Bag06, Bel93c, Ber06, Bur02a, Bur05, DRR02, DWMB05b, DWMB05a, Egs70, FA03, GB14, Gri78, HR74, HS99, Jee71, Leu05, Min83, O’B70, PS21, Sin03, Str78, Str84, Str85, STM06, TS93, Usi74, Vol69, El 86, Lor82, MK91, Mon00, Pol68, VD79, Add72, AV76, Del68, Del69a, Hof06, KC73, Mar69, PF11, TLM07, Van69b, VD78]. sondage [Bui76]. Sophia [Tur88]. sophisticated [DS17]. sought [Ano14a]. sound [BR12]. Source [Lab00, Fre69d, ZB19]. sources [AI07, ABOM20, Bon13, ET95, SGCA18]. soustraction [AL00]. soustraire [Dem93]. South [Ano16c, DBL22, Adl88, Ban11, HMG16, Nkh02, SA00, VAM21, lRA16]. Souza [LT20]. Soviet [FF21, Fre74b]. Space [MZ18, Bal88, CZ11, FCN14, KC09, LH71, MRCR11, Pap71, Par88, Par91, Pir25, RL11, ST10, TDEM13]. spaces [Ban11, ERS19, EVHRC^{+}25, FD24, Gla69, GM08, Lec92, NT01, YTD11, KN21]. span [GKMJ24]. Spanish [BBEG05, VVM22, VMN11]. Spanish-speaking [VMN11]. Spatial [Bis80, LC81, BCLH89, BH11, BDS^{+}17, PC10, SMW11]. spatializing [WPL24]. speak [AAJW25, JB21]. speaker [Pla14]. Speaking [XC19, NX23, VMN11]. Special [BA25b, FV22, Hof06, NBD^{+}04, Obe19, PvdHPR12, Rod06, AP01, BRE20, Bak19b, Bro08b, CL18, Col21, DVC19, HMG19, HMG20, PR08]. specialised [EVHRC^{+}25]. Specific [BHX97, DGA05, Kun12, LLZ20, OAS25, SL19, SLP^{+}12]. Spécificité [BC92]. Specificity [BC92]. speck [AOJ22]. speech [Kem81, MB09]. speed [JHC14]. splitting [CB19]. Spontaneous [DH90, HVN^{+}24, LHS19, Tre83, HP25, HWJ20, SWVO23, UOF25, VO17]. sportifs [CGS76]. sporting [CGS76]. spout [Tro00]. spout-casserole [Tro00]. spread [Fri21, RA21]. spreadsheet [LE09, CB89]. spreadsheet-based [LE09]. spring [ZW21]. Springer [Ahm23, Cor22, Gre17, NX23, Cot24]. square [BA23, ZL08]. Squared [BP80]. Sri [Gun78]. Sriraman [Pai18]. stabilité [PRR08]. Stability [MK97, Fre72, McD89, PRR08]. stable [VGJ19]. Stacy [Mal23a]. STAD [TA08]. Staehler [Sri18]. Stage [Coo98]. Stages [KB07, Dié70, Ger10, Noe80a, Noe80b]. stagesin [Tzu07]. stakes [ICR20]. stance [HEWC10, TL15]. stances [McV24]. stand [AAB03, RC02, RC04]. Standard [GC90a, FVJL^{+}19, KZ94, Pre06]. standards [Mac00, MO94, WBM^{+}14]. standards-based [WBM^{+}14]. standing [Car08]. started [LM11]. starts [Gre17]. State [BN91, Nis99, CKS22, Høy07, LT20]. state-of-the-art [CKS22, LT20]. statement [Bis79a, Bis85, FTM81]. statements [CCRD22, SM22a, SS95]. states [RL13, FSY22, MNRR95, YC95, Law00]. static [Lab78, MAHK08]. statique [Lab78]. Statistical [BH05, WM98, WK04, vDDDB21, BA14, Büs24, GG22a, GLZBZ15, GAP24, Kap80, Mev83, ST02, Wei17, dVSKvO18]. statistics [CGS76, Hah14, MRRF94, MHC24, Tru15, VM78, dSBM21]. statistiques [CGS76]. status [Kae90, Mwa02, XH88]. Steinbring [Mor00]. STEM [Fri21, Fri21, KPS17]. step [Hoy01, MM25, NH94, Ric04, SW95]. Stephen [Ahm23]. Steps [Ber06, Bur05, HN08, Rev69]. Stereotypes [Tie02]. stereotyping [Rut87]. Stieg [HN95]. still [JGF20, VVV13]. Stimulating [DVT^{+}18]. stochastic [Hei75]. Stoffe [Beh68]. stone [AMMS20]. Stories [BVR24, Llo06, FSB18, KK13, Mie23, Osl12]. story [AZ20, AOJ22, BF21, Die15, FF25, Fur19, HS02, KB11, LXT02, Lin90, MZ12, RL00, Yoo25]. storying [Nar16]. Storylines [ARW22]. straight [KP10]. Strategic [CGD21, RSL17, Web01]. Strategies [ABvP02, CB00, Csí16, Fie79, Gor98, JS99, Sel01, Rev76, BBDD25, BBVD25, CN00, CWT12, CVN21, CR09, Deb06, DS17, Eng91, Fro99, Gal86, GG96, GS18, HK25, HO23, HKG89, HP03, JB21, JHC14, JNF25, KPH10, Kor03, KSL11, kSL13, Noe80b, Nol16, OMB20, PPG95, PFSS19, Pro13, SS15, Sch16, SD94, Tho90, TDGV09, TR24, Van03a, Vol86, Yer00, Yim10, ZC22, vdHPKR13, Deb06, Rev76]. Strategy [Mur04, Ano76a, RRP91, Son16, VDTV24, Vin75, Yan12]. Streefland [Pre03, PvdHP03, YSRU03]. Street [GB19, JS99]. strengths [Kim20, MCW25, Sch23, Sch25]. strengths-based [MCW25]. strewn [HH81]. strictement [Rob95]. strictly [Rob95]. string [Fre79a]. strip [Roh10]. Strongly [KM24b]. Structural [Jan74, JGR14, Jur77, SLS74, Jer73a, Jer73b, KBK^{+}24b, Lep90, Pir25, RC94]. structuralist [Rot12]. Structure [Duv91, LL99, PPDS^{+}25, Sha98, SS06, AT21, BFG84, BIBS^{+}19, CWS93, GCALB15, GB13, Hic24, LS70, LH71, Mar13, Mev83, MCPGE24, MFJ17, Mor14b, Noe80b, Roh10, Rüe13, SM17, VS23, Duv91]. Structured [ABvP02, HKG89, Hin15, Jee71]. Structures [Ham06, Usi85, WG07, Yan12, YG99, CSP24, Dié70, Une74, HP25, HMA11, KMT14, McD89, MP96, Nol16, PA18, PS21, Sch20, SW95, Tho93]. Structuring [TD02, DT12]. struggle [HMG16, TT01]. struggles [Kon18]. Stuck [MZ12]. Student [Alb00, BH05, BN96, Bju04, BMPT24, BG05, Cal94, CY23, CN16, CPP07, CR09, CB00, Cop98, DVJ98, Doe06, FDW19, GP17, HKSS^{+}25, HL16, Har09a, Hod96, HK02, IWF^{+}15, IB05, Joh15, Jon00, Kaz07, KV23, KID25, LL13, Lit03, Loc11, MS97, MMBD05, MPGDA12, MPG12, MC01, McN00, MFJ17, Mos98, Mos05, Ort83a, Ort83b, PLW81, Pug04, RP21, Riv07, Roh08, Roo23, RBK25, Ryv06, SG08, SV25, SM17, SG98a, Sol06, SK25, UMC24, VL04, Wal68, Wat07, Web01, WT14, Wil16a, Zaz98, ZC06, ALR25, AI08, AI12, AMRM24, AVM15, AWL15, AWE22, BA14, BA25a, Bar03, Bec23, Beg71, BCLH89, BHLH85, BZA01, BML21, BWHM^{+}10, BBDD25, BBVD25, BC96, BMP20, BP15, Bra14, BAZF04, Büs24, CO10, CMW^{+}13, CHM16, CLK22, CV09, CBC25]. student [Car10, Car18, Cas04, Cha93, CH13, CL23, Cla85, CSS^{+}14, Csí16, DVC19, DC17, DVJV02, Deb06, DGI23b, DT08, Dra14, Dri25, Ecc85, EJ13, EEST20, FQL^{+}17, FD24, FKC17, FE09, FWCW22, FA03, FMA20, Gan92, GLZBZ15, GKMJ24, Ger10, GBTL23, GMNB18, GK20, GMHG09, GC23, GAP24, Güç16, GS18, HK25, HTM18, HRW25, Han20, HP25, HOS13, HWR23, HBZE98, HS13b, HM13, Hic24, HHDG16, HMA13, HMT13, IM19, ICR20, JCM^{+}23, JHC14, JM18, JDK^{+}24, JE14, KPH10, KP10, KR25, KK12, KPS17, KB11, Kon21, Kon22, KOMH22, KJvD20, Kor03, KSL11, KC22, KR11, LOMG^{+}20, Led87b, LK16, Lev10, Lev13, LC15, LK22, MBF25, Mal93, MWSS17, MF17, Mar13, MSdSGG16, MPdP17, MMP15, MEH20, Mev83, MK97, MS20, MCPGE24, MW09]. student [MAPHB21, El 18, NCC22, NRV24, NH15, NJL25, Ola14, OAS25, PP12, Par91, PvdHPR12, PT09, PAW15, PPDS^{+}25, PD23, PMG^{+}94, PR08, QM94, Rad00, RW11a, RW11b, RCH25, RSL17, RRP91, Roh10, RDH10, RC94, SLP^{+}24, Saj03, SMFL19, SGRS24, SJ17, SBL22, Sch23, Sch25, SLP^{+}12, SKR15, Son13, Son16, SMS22, SSP04, Sty11, SLW83, TA08, Tch11, TT07, TS99, Tru15, TO13, Tzu07, VS13, VP13, VYZ21, VN19, VR90, Wal10, WZC21, WAL18, WSR11, Way94, WM09, WS18, WH24, XXL21, YM16, ZWC^{+}24, ZC22, vdS11, vdHPKR13, GT95, IPJ25, LX16]. student-centredness [Wal10]. student-teachers [HMA13]. Students [BCFF^{+}98, BH95, CBF07, Car99, Cla07, CRL06, GL98, GBR05, GPW05, II05, Jur06, MM00, STM06, WB11, Wat01, Wit05, ARW22, AK19, Bat13, BVR24, Bek10, Blo03, BLP10, Bon14, BJJ13, BSF96, Bro08b, Byu24a, Byu24b, CY23, Cla83, Cla92, CBSP12, CSK12, CLSM17, DW17, DeB97, Dem93, DDH25, Dib19, DH20, Dom11, DLK14, Dom19, DS17, DBL22, Ell13, Eng12, ET95, FV13, FJC95, Fra13, Fra96, GdA09, GG22b, Gro77, GLH20, Hah14, HRW25, HC95, Her02b, Hew12, HP03, IAP19, JGF20, JNF25, Kae93, Kae94, Kar10, Kis86, Kon72, LH71, LBT07, LMCG08, LLZ20, LG16, LLD14, Ma96, MGV^{+}25, MK91, MG00b, ME13, MFG22, Min83, MY84, MLL20, MRC18, MA09, Mur24, Ng21a, Nil07, Nkh02, Nuc24, O’B73, PG20, PBCPP25, PFSS19, RO09, Rob86]. students [REL25, Rüe13, Sam85, SM22a, SLS74, SGTK14, Sie87, SBP11, SLC95, SS93a, Sin00, SHG21, SG98b, TJ94, TLM07, Tur80, VCD24, VE24, WZ86, Wal12, Yan19, YM16, dPH13, FPS16]. Studien [Dam87]. Studies [Ano02c, Bak19b, Bec19, CL18, Kae94, LPG05, MW19, MBV21, Pea76, Add72, And09, CV09, DB86b, Fit15a, Kol16, Kol18, Mur24, Obe19, Prz04, Rob90, Sch24, Sfa15, Dam87]. studio [KC22]. Study [Ano98i, Ano04d, Ass05, BT05, Blu02, CPP07, Cre06, Dou99, DS05, DGA05, EJDT06, EH06, GL00, HbL05, JS99, LD88, Mos04, RSRVRVVR23, RD92, Sem70, Sid05, SS06, Spe05, THS06, War06, Wat05, WI01, WMM13, Ale81, Ano14b, Art23c, AK19, BGW^{+}14, BGW^{+}16, Bal80, Bal71, Baz90, Bel76, BZ01, BC78, Bla85, BBDD25, BBVD25, BAZF04, BOL24, Buc87, Car10, Cha18, CM20, Cob86, Cor17, CHK22, DVJV02, DMHRV21, DH20, DHD23, FQL^{+}17, Fra96, FC12, GD17, GG22a, Gal82, GCG^{+}92, GK71, GAP24, GBR22, HWJ20, HSM22, HMSBR19, HK99, ICR20, IS08, IS09, Jan10, KM93, Kar10, KT01, KM24a, Kon21, Kon22, KN21, LL23, Lam15, LCD90, Led87b, LLZ20, LLL21, MGV^{+}25, MvdKM13, MCVM13, MRW14, MMR21, ML20, MRRF94, Ng21b]. study [NS25, O’C01, Ott20, Par16, PA14, Pet91, PPDS^{+}25, PBCPP25, PMG^{+}94, Pop88, RL13, RMAGMR22, Roh10, RT14, RIA17, Rüe13, RJWB20, RSD23, RDH09, Saj03, SBL22, SB20, Sek21, SGTK14, SS91, Sin03, SMS22, TF95, TLB^{+}24, Tro00, TJHD11, VVV13, VYZ21, VMN11, VO17, VS23, VS25, WK04, WMR11, WM09, Wri94, YSL22, ZS15, ZWC^{+}24, dA00, vBP19, Ano00d, Ano14b, FvM97, Leu05, IPJ25]. Studying [HP08, Ler01, RMA98, Adl19, Gof77, Rot21, VA18, MRCR11]. style [Chr00, Kae90]. Styles [Kol21, Ber19, CPPT^{+}13, Hem10, Isa87, PPC09]. Stylianides [Gue19, MR20, Sch24]. Suad [SJS12]. Suan [Siu93]. Sub [Osi24]. Sub-Saharan [Osi24]. subconstructs [ČK18]. subitized [WMN22, WMN25]. Subject [Eve90, ET95, Ng21b, BN96, BS94, JMRW22, Ngu24, Pai16, Rot12, TH13]. subject-choices [JMRW22]. Subject-matter [ET95]. subjectivation [VC19]. Subjectivity [Bro12, Bec24, Bro08a, BH11, PR08]. subjects [Beh68, BN91, BF96]. subsequent [SBDM69]. Subspace [WSR11]. substance [Nar16]. substantive [SKN21]. substitution [TGWN23]. Subtest [BHX97]. subtract [Dem93]. Subtracting [Str82]. Subtraction [FZ05, Sel01, Tho90, AL00, CHM83, FVJL^{+}19, KBBK20, KBK^{+}24b, KBK24a, Mat81, Mat83, NK77, NGR82, PvdHPR12, PDT^{+}12, SPNH12, Son16, VDTV24, VBE21, VM13, Ves69, WH24]. subtraction-compensation [WH24]. subtraction-related [VDTV24]. subtractions [PDT^{+}12]. Success [Car99, MA09, Sel01, Bur87, CL84, Gli91, KK13, Led80, Nkh02, SJS12]. Successful [Car99, Dah04, Bon14, Dib19, Nkh02]. successfully [LLD14]. successive [Noe80b]. such [Tsa01]. Sudan [El 78, PSE87]. suggestions [LBT22]. Suh [Sto22]. suitable [EVHRC^{+}25]. sum [MBF25, Tab13]. summaries [DAJ20]. summative [CB19]. summer [Str08]. Summing [Rut10]. Sun [vBP19]. SungWon [MPS12]. Sunshadows [Dou99]. superfluous [GSH16a, GSH16b, Lin90]. Supervised [Cop98]. supervision [FE09]. Supervisors [GPW05]. supplementary [AAJW25]. Support [HBH07, Kap01, MO94, BA14, CWH19, CSS^{+}14, GRR23b, MHC24, RIA17, RH02, Vla02, VEBS24, WC08]. supported [GABB25, RDH09, SS15, vdHPKR13]. Supporting [GDS22, LH22b, LLD14, MND23, MC01, SDVT25, EEST20, GCFS23, HFL^{+}23, PFSS19]. supports [Woo16]. supposer [YC90]. suppression [WHE08]. surface [DPG89, HP25]. Surveillé [Cop98]. surveilling [Sko12]. survey [BCZ21, BGSW94, Bui76, Fey89, FB02, GB01, MCW25, MRW14]. surveys [Del68, TE14]. survive [NRV24]. Susanne [Bak16]. sustainable [MCW25]. Swapna [Fit11]. Sweden [AAJW25]. Swedish [Han10]. Sweep [Vog70]. Switch [Gil71]. Switzerland [SS93b]. sword [dKBvdS22]. Sydney [Wat05]. syllabuses [Hol81]. syllabusism [Høj24]. Symbol [SLW98, BP80, Kie81, KJvD20]. Symbolic [NHH97, Sko12, Bos87, DF15, MM20, Rob75, Rut90, SSP04, TAH08, TP22, Tro00, VYZ21, ZWMR16]. symbolically [CCRD22]. symbolique [Bos87]. symboliques [Tro00]. Symbolising [Pai15]. symbolism [CMdR18, CB17a, CB17b, HR74, KT15]. symbols [Hie88, MM20, Wea90]. symmetric [Hah75]. symmetry [Coo92, Ros69]. symposium [Ano74b]. Synergies [MCVM13]. Syntactic [WA04, SK13]. Syntax [RP07]. Synthese [Han98]. System [Van74, Bro10a, BBL22, CB19, Dub84, Gen70, HOEF89, KRHD15, PGV22, Sfa15, SAD21, SS93b]. system-wide [Bro10a]. Systematic [Eng68, AS25, CKS22, KSSO25, MGI23, RBWS18, SM22b, SMOR20, WSU19]. systématique [Dub84]. systeme [Ott76]. systemic [PAR^{+}24, Wit01, ZCPP13]. Systems [Gat06, CV09, Dié71, GMHG09, GT09, KB20, LCW20, MWVE09, PXTW17, Sek21, SK69].

T [Fit15b]. table [Lan86a]. Tableur [CB89]. Tacit [Fis01, Kid11, PA18]. Tackling [TR24]. TAI [TA08]. tails [LLHL13]. Taiwan [KCS22, YC95]. Taiwanese [YST22]. take [Sca75]. taken [Kon18]. Taking [SPNH12, ALO^{+}91, IM19]. tale [Gri87, HRW25]. talented [ÇA20, Ell86, Kis86, Lar86, SGTK14, rA20]. Tales [KDM21, GM10]. Talk [Cre06, DE24, Edw09, LK16, RG19, RG20, Row95]. Tall [Bye15, KH80]. Tamàs [Kry88a]. tangents [Hir72]. tangram [BA23]. Tanzania [Phy71]. Tao [Cle84]. tape [Fie87]. Task [Chr97, KK13, MM00, VBE21, AGP15, Bro08a, DOV24, FB16, GGD22, KH07, KSH25, KBK24a, LL13, LLZ20, OAS25, SLP^{+}12, SMW11, SS88, TGWN23, TTTL10, VGJ19, VS25, WSU19]. task-based [KH07, LL13]. task-specific [LLZ20, SLP^{+}12]. Tasks [Doe06, II05, Lit00, MDD04, ARKK23, AOJ22, AAM21, BLP10, BOL24, Cle80, EJ13, FI17, Fos18, GG22b, Gue19, HC82, HKG89, JE14, KID25, KL20, KSL11, Kra93, KMA02, LLW07, LMCG08, Lev22, LG16, LLD14, MP24, NCC22, NJL25, PPPP14, PPDS^{+}25, Ric20, REL25, SMW11, SV13, TTL^{+}13, VVV15, Vol77, VGJ19, WvdHPD15, WH24, YLH24]. Taught [Mur04, Egs70, Miy17, MRC18]. taux [Sch92]. Taxicab [KV23]. Taxonomy [EH05a, Ola14, Bec23, DŽC25]. Taylor [Mar13]. Teach [Moy01, Nic98, NC06, PBR^{+}04, Ano76d, Cd04, FVJL^{+}19, Fre68, HM25, LK14a, LBH23, LSC20, LC15, Par15b, Par15a, Pol69, RSD23, RDH09, Ser68, Sim94, Swe95, Van76]. teach/learn [HM25]. Teacher [ABK^{+}05, AR07, ABM92, AAT99, BBEG05, BKK22, Bes07, Bes12, CWL07, CP10, CEMP04, CSS^{+}14, Coo99, CG21, Cre06, CLSM17, DEM04, DK15, EJ13, FPCL90, FI17, Fre77a, Fur07, GSE07, Gat06, GS18, HDM07, HHT23, HMT13, IE07, Ior18, Kim20, LE09, Law00, Led87b, Llo06, LLL21, MCB05, MMBD05, PKP^{+}18, PC98, PMG^{+}94, RR05, Sen98, SSLM^{+}05, Sia24, Spe05, Tie02, TER98, Tsa07, Wal10, WG07, WHC09, ZC22, ZS07, dPQ16, dF12, AG21, AV08, AAP21, AV22, AH00, And09, Art15a, Ayd90, BN96, Bek10, BCFC90, Bju04, BHN10, BF96, BR17, BKH25, BC12, BBL22, BOL24, BE23, CLK22, CEHM16, CRPPRG16, CSP24, CPK08, CPP09, Cor22, Cre03, DC92, Dam87, DR13, DVT^{+}18, DWP21, DDB^{+}10, EH20, Ell13, ET95, Eve99]. teacher [Fau15, FSY22, Gel98, Gel08, GCFS23, GLHM24, Gof77, GMNB18, GS20, Goo07, Gra04, GSW22, GB17, Gui17, Gut92, HFL^{+}23, Har09b, HPS^{+}19, HM13, HMMG18, HMSBR19, HK99, Høy07, HHB21, IWF^{+}15, IAP19, JN15, Jan09b, JCM^{+}23, JGF20, JMRW22, KHLP08, KPH10, KM24a, KZ94, Kle12, KOMH22, KR09, KR11, LBH23, LK16, Lee17, LLW07, Ler14, Lev22, LSC20, LCK24, LNC23, LS20, LYHC18, Lou18, LK14b, LK18, MvdKM13, Mah21a, Mah21b, Mal93, MM11, MPdP17, McQ86, MPZ16, MCL25, MLK16a, MLK16b, MN95, NHO22, NBZ12, Nol12, Nuc24, OVKR25, OK13, Osl12, OCB25, PSH^{+}18, PPPP14, PA14, PXTW17, PB09, PR08, RL00, RPP24, Rut10, RNM12, SMFL19, SG14, Sca71, SF23, SGB21, SGTK14, SZP10, Shr10, SS91, SS93a, Sim94, Son16]. teacher [Tch11, Tha10, Tho84, TB22, TH13, TG89, VC19, Var76, Was15, Was22, WL90, Wil21, XA20, Yah25, YHL^{+}15, YSL22, YST22, Yan11, YM16, ZZ11, ZB19, ZZ08, dVSKvO18, lDlzN21, vdKG22]. Teacher-centredness [Wal10]. teacher-education [Ell13]. Teacher-student [HMT13, LK16]. Teacherly [Che23]. Teachers [Ano04d, DS06, Gra99, HS13b, Kry69a, LH22b, LG13, kSL13, MP24, Men07, Moy01, Nic02, NC06, Pre06, Sko04, Tsa99, Wat00, Wat05, XM17, Ale81, ABOM20, Ban11, BKK22, Bec23, BH94, Bon14, BSF96, Bro08b, CDW14, CN25, CW14a, Cla12, CSS^{+}14, CG21, Cor17, DS02, Del69b, Del69a, DH20, Dra14, EK10, FE09, FJC95, Fra96, Fre69e, Gal11, Gal82, Gen70, Gli91, Gri68, GT09, HOS13, HBZE98, HSM22, HHDG16, HMA13, How75, Jan09b, KBK^{+}17a, KBK^{+}17b, KL82, KL20, KH07, KZ16, KN21, McV24, MS20, MHH90, Mur93, Mur95, Nol16, O’B77, OAS25, PRR08, PGM13, PK23b, Phe10, PK23c, Rel96, Rüe13, Sch23, Sch25, SB97, Son13, SKN21, TS18, Tie00, TDEM13, VD79, WKB97, YK20, ZB19, ZHT25, dF04, dF12, dVSKvO18, TR24]. teachers- [YK20]. teachers-student [OAS25]. Teaching [AP07, ACJO04, Als02, Ano74b, Ano98i, Ass05, Bag73, BP22, Blo03, BBL22, Cas98, CWL07, Cha97, Chr00, Cla02, Cli17, DS06, Doe06, EH05a, EH05b, EVHRC^{+}25, GP01, HPG05, HBH07, HbL05, Jaw02, Jaw03, KB07, Kle02, LPG05, Ler01, Mac21, MMBD05, MR17, NBE^{+}12, PBY03, PGM13, PM97, SLP^{+}12, Šed69, Sel98, SGTK14, SBP11, Ste05, TB22, TER98, Tir99, Tsa99, Tsa02, Uhl02, Uhl03, Ves76, VM78, VEBS24, WD05, Wit84, Abt22, Adl19, AV08, All85, AMRM24, AH00, ARHS14, Ano68a, Ano68c, AV76, BS12, Bat13, Beh68, Bel93b, Bel93c, BC92, BS80, BN77, BKH25, BR78, BE23, Bur80a, Bur02a, Bur80b, Cas77, CJ75, Cha18, CW14a, CPK08, CWH19, CGD21, DVF18, DB86b, DMHRV21, DS21, lD95, DRR02, DTD21, DBL22]. teaching [EVR16, Elb03, Eng68, Ern85, Eve90, EK10, FvM97, Fie79, FG84, Fle71, FBA10, For69, Fre74b, Fre93, FSY22, FC12, FCN14, FM14, GP90, Gan92, GBMB70, GMH23, GCFS23, GMNB18, GFJ13, GSW22, Gri87, GB13, GSH16a, GSH16b, Hah75, Han10, HOEF89, HSM22, HMG19, HMG20, Hof77, HK99, HHB21, HHT23, Hug74, Ior17, Isa87, Jak71, JP09, Jon16, Jou13, Kla68, KO23, KZH^{+}19, Kry69b, bL00, Law00, LCK24, Lib77, LLL21, Lun77, MCKP19, MMC22, Mar69, Mas69, Mat03, MLL19, MG12, Mog17, Moo13, MRSP14, Mur24, NCMO15, Orl71, Osl12, Pas78, PXTW17, Pim90, Pol68, PY22, PPSZ19, PJP23, Pro69, RSS11, RL00, Rel96, RK17, Rev68, RMAGMR22, Rog71, Roo19, Rou68, Row08, Roy81, RH02, RNM12, Sar02, SGC22]. teaching [Sch24, Sch71, SECP^{+}23, Šed76, Sek21, Shv22, Sia24, Siw73, SHAC01, Ste91, Ste97, SK69, Sto71, SPS17, Str03, SLR^{+}24, Sut93, TN16, TK14, Tch11, TSV82, Tho84, TLB^{+}24, Tri95, Usi74, VRP^{+}12, Van76, Vil71, Vin75, Vol69, Vol76, Wai70, WBM^{+}14, Wal73b, Wal17, Wal76, WqWLR18, WH15a, WH15b, WL90, Wit69, Wit81, YST22, ZZ11, ZM25, Han12a, JC22, Pis68, PJP23, RSD23]. Teaching/learning [Cla02]. Teasing [Han12c]. tech [KO21]. tech-related [KO21]. Technical [Ano02d, Ano02e, Ano02f, Ano03c, Ano03d, Ano04f, Ano04g, SS05b, Ern16b, RNP13]. technique [Wai70]. Techniques [MDD04, bL00, Lev81, PSH^{+}18, UMC24, Wai70]. techno [Gai11]. techno-mathematical [Gai11]. Technological [Baz01, MM03, Bis79b, KM24a, TP10]. technologically [GABB25]. technologically-supported [GABB25]. Technologies [EH05a, HbL05, Jou13, KL14, LK14a, Ric20, Wal12]. Technologizing [Zev04]. Technology [Fey89, GH98, Jah83, SK98, SLAY19, Wil05, BdSCdA18, BR17, BKH25, Bur90, CEHM16, CWH19, DDB^{+}10, FT00, HTM^{+}23, HSZ17, bL00, LE09, MCVM13, PB09, RDH09, SGRS24, Sko88, Sta11, Str87, TB22, TLB^{+}24]. Technology-based [SLAY19]. technology-enhanced [CWH19]. technology-rich [DDB^{+}10]. technology-supported [RDH09]. teletype [SLJ69]. television [GR71, Wai70]. Tell [Bur98, BF21, AAP21]. Telling [FF25, GM10]. templates [Riv10]. temporarily [HA23]. ten [Jur77, KBBK20]. Tending [Bar19]. tension [ZL02]. tensions [Abt22, HSM22, MLL19, Zev11]. tenth [TD12]. tenth-grade [TD12]. Terence [Cle84]. term [AMRM24, Str84, Str85]. Terminology [FA76]. terms [Ger10, JSV75, RNP13, Vau71, WM09]. terrain [Sta11]. tertiary [Ade98, CB17a, Cor17, CB17b, Gue08, Gue23, KO23, LNC23, LYBK23, MG12, PK23c]. tesselation [Har74]. Test [BHX97, HH78, KV75, Mau85, Ola14, Wal73a, Yan19, Kon72]. tested [Bro10a]. Testing [Hee78, And79, DT08, HYC^{+}97, JDK^{+}24, VH16, Gau69]. Tests [Coo98, WW02, And89, Bec23, CF00b, Dum82, SGRS24]. Text [NHN03, Ber19, DS21, Kry69b, Mor14b, SGCA23]. Textbook [MG12, DL14, GKMJ24, Lit03, Par16, Sun11, Tur80]. Textbooks [Mes04, NN07, NC06, Ala12, Cho22, DF15, DHD23, FZ07, HRR23a, HRR23b, HC14, SD13, SV09, VVM22, WW11, WvdHPD15]. texte [Kry69b]. textes [Rut02]. Texts [Öst06, AK19, Bra11, Dow96, Kon93, Wag12]. Thailand [Pur78]. Thales [EVHRC^{+}25]. Their [Cla07, HDM07, Jon00, LW99, AS04, AS05, AWE22, Bes12, Bju04, BWHM^{+}10, Cha93, DEMW06, FV13, FYAE12, Fra13, FA03, Gue19, HH11, HM13, HMCT25a, HMCT25b, IM19, Jab23, JMRW22, JM18, Kar02, Kon21, Kon22, KO23, Kra02, KBBK20, LBT07, LJW13, LK16, Lev22, LCK24, Lin88, MY84, Ng21a, Nkh02, NBE^{+}12, PXTW17, Phe10, PC10, PBCPP25, Roh08, ST02, SECP^{+}23, SGTK14, Sol06, Sun11, TTL^{+}13, Tsa07, TDEM13, VDTV24, VP13, VD78, VYZ21, WZC21, WT14, Wei91, WM09, WS18, YM16, YK24, Zev11, ZB19]. Them [HDM07, IAP19, LBH23, Pol76]. Thema [Wal76]. theme [Chr00]. theme-based [Chr00]. themes [BCZ21, Jou13]. themselves [JB21]. Theorem [MRSP14, EVHRC^{+}25, Gri13, LL23, Tho94b, MGD^{+}23]. théorèmes [Pas78]. Theorems [GB17, Dav93, Pas78]. theoretic [AS11, RL04, SJ04]. Theoretical [CPP07, Eli85, Jaw03, LPG05, AM14, BWD10, BBFG99, Cha18, CSDK25, HY22, HPRW25, KL14, MDU^{+}25, MCVM13, MM25, MTBR16, PK23a, Reu23, Sch20, Sim17, Str93a, Vou16]. Theories [Ely10, Bak16, BGT17, GBE13, Kid16, KSS25, LXT02, RSD23]. theorising [SM21]. Theorizing [Dom19, MLL19, PJP23, dFS12]. Theory [Bag06, BSW16, Bro23, DDLZ97, GD17, HPG05, Nar00, Spe05, Tsa02, Web18a, Web18b, Bas17, Boa02, BMP20, BW12, BTvG^{+}17, Bur96, BB03, DE24, DDLZ94, Eng03, ES03, Fis93, HEWJ^{+}15, Hie88, Hol81, JP09, Kap76, Kap77, Kei82, KN21, Lan13, LLW07, LYHC18, Mai85, McQ86, MEH20, Mon72, Moo20, MAPHB21, Mor14a, Nim75, Nol12, NX23, Nun96, PV14, Pro13, RNRVM23, Rot12, Shv22, SB12, Str84, Str85, Sut93, TR11, VA18, Ves76, WB11, WW11, WH15a, WH15b, Wil09, Zaz99, PJP23, Tab22]. theory-based [LLW07]. theory-building [Bas17]. Therapy [Wil97]. There [MMM25, Sfa01, BS25, ST08]. these [MDPW96]. thesis [Dar82, Sch83]. things [Bor21]. think [Bye15, DE24, ZG01]. thinkers [Mur24]. Thinking [Ano72a, De 94, Doe06, EJDT06, IE07, Lan86a, MM03, PBY03, RP07, SZP10, ZS07, Bar19, Bar09, BBG^{+}17, BBFG99, CLK22, CBC25, CSP24, CPPC18, CPPC23, DVF18, DDH25, DHD23, DS17, DKBD19, DKBD21, ES21, EE21, GP90, GLZBZ15, Ger88a, GK20, GPPT99, Gre01, Gre12, HY22, HS13b, HSZ17, JLTM97, KAJHB10, KOMH22, Kru13a, KC22, Kyn93, Leh24, LH09, MvdKM13, MH97, OSR71, O’B72, O’B73, Pir25, PC10, Rad00, RSS11, Rot21, SGC22, Sch23, Sch25, Sfa01, SKE25, SJ04, Sut89, TT91, Tav68, TW09, Van69b, VYZ21, Vol86, WC08, WH24, Wit69, Wit81, ZCPP13, ZL02, vO10, Win11]. Thinking-Aloud [ZS07]. thinking-based [KC22]. thinking-language [OSR71]. Third [Gre17, OMB20, SLW98, BR78, CBSP12, CSK12, Une74, El 86, GCG^{+}92, VD78, VD79, VBE21, RD92, Ano14b]. third-grade [CBSP12, CSK12, VD79, VBE21]. Third-grader [OMB20]. Thirteenth [GL00]. Thomas [Hus25]. Those [Cd04, Bat13, HH11]. thou [Fie87]. Thought [Alb00, Vin97, Car93, For00, GE98, LLD14, MS09, Sla98, Van76]. Thoughts [Fre93, PF11, Str84, Str85]. thread [Sur77]. threads [HH24a, HH24b]. Three [BT04, Coo92, FPS16, Hem10, II05, Mit80, Nes76, O’B76, Sel98, Sel01, Van93a, BJ88, BKGZ24, Bye15, Csí16, Jou13, Kid08, KDM21, KBK24a, PK23b, PXTW17, SSA20, VMN11, Yan19, ZG01, DeB97]. Three-digit [Sel01, Csí16, KBK24a]. Three-dimensional [Coo92, Mit80]. three-tier [Yan19]. tier [Yan19]. tiles [CO10]. Tiling [De 94]. Time [Ass05, Beg71, WW02, CSW21, HHT23, HHL22, JMRW22, LK18, Mac21, SRP21, VVV13]. Timed [WW02]. Times [Zev04, AT21, Kle02]. TIMSS [RD92, ARHS14, DT08, Leu05, Mac00]. TIMSS-R [DT08]. Tin [SMWM25]. Tinne [Fur19]. today [Ano88e, Ayd90, BW20, ZM91, Goo20]. toddlers [BP22]. Todorova [Tab22]. Together [Joh15, HH24a, HH24b, SZP10, lDlzN21]. Toh [SMWM25]. toilet [FS16]. Tomaso [D’A10]. Tommy [Rad11]. tomorrow [BW20]. Tonga [Mor21]. tonight [KLLJ17]. Tony [NX23, PR08, dF12]. Too [Als02]. tool [Car10, DDB^{+}10, HMRRTT23, Høy07, KL14, LC15, Rob12, RPP24, SSA20]. toolkit [CWH19]. Tools [PNH98, Cha98, DC17, FLM07, GFWL11, HS99, Kor03, bL00, PGE09, PA14, RH02, TP10]. topic [SMS22, Sun11, Tab22, Wal76]. topics [Güç16]. topological [Dar82, Pap71, Sch83]. Topology [Hil71, Kel71, Kap74, Wal73b]. topos [RFK07]. Toruń [Ano74c]. totally [TE14]. touch [HO23]. touchant [Del69a]. Touching [NPH99]. toujours [Gla76b]. Toulmin [MPZ16, NBZ12, Sim15]. tous [Kry88b]. towers [WH24]. TPACK [KSSO25]. Tracing [CPK08]. tracking [BA25b, BBDD25, BBVD25, BOL24, PPDS^{+}25, SL19, SSL25, SV25, SMOR20]. trades [LaC14]. traditional [Bra11, FKIA15, Ger88b, MRC18, Son16]. traditionally [Bon14, TDGV09]. Traditions [GL00]. tragen [Van76]. trainee [PM15]. Training [BH94, BGSW94, CN25, DS02, Del69b, Del69a, Fre77a, Gal11, Gen70, Gof77, Gui17, Gut92, HK99, KL82, KR09, LaC14]. traits [VH16]. trajectories [AM25, MZ21]. trajectory [Van03b]. transcoding [HA23]. transcript [Ano78b, IT77]. transcription [DH20]. transcripts [CW96]. Transfer [CV16, Boa93, Eva99, Loc11, TP14, Van74]. transferring [Dag96]. transfert [Dag96]. Transformation [Har74, Rad11, Ern86, FQL^{+}17, ML20, NS25, TZLT95, TGWN23, ZER17]. transformations [BC96, Cas69b, Kel71]. Transforming [NS25, Wal17, Shv22]. transiting [MWVE09]. Transition [DGI23a, DL14, DS05, KM25, SS05a, Tsa99, Ber08, CB17a, CB17b, DVF18, DGI23b, Gal02, GRR23b, Gue08, Gue23, KO23, LNC23, LYBK23, MGI23, MAM24, Moo94, PK23c, SM22a, ST08, VR90]. transitional [Ger10, Mos98]. Transitioning [KLwK24]. Transitions [GE98, TAH08, Sla98]. transitive [Roy81]. Translated [Gre17]. Translation [WW07b, CCRD22, Gre17, Rut90]. translations [HP25, Vau71, Yan11]. transposition [MRC21]. trash [Ham68]. trauma [LM11]. Travail [Cop98]. Travels [Dho93]. treasured [Kid11]. Treasures [Yoo25]. treat [SBA22]. tree [HRR23a, HRR23b]. trend [APN01]. Trends [JE14, El 18, BN91, Han19, MGI23]. Trgalová [Tab22]. Triad [PJP23]. Triadic [KH07, Ong06]. triangle [MRSP14]. triangles [Has97, TTL08]. tribe [Win11]. Tribute [Ano78c, Bac21]. tricks [Mie23]. trigger [PG20]. triggers [Kon20]. trigonometric [GK71]. trigonometry [FG17]. triumvirate [KK10]. Trobe [Art23a]. Trois [DeB97]. troisième [VD78, VD79]. Troubling [HMA13]. Trouche [Cha22b]. Truth [Bro10b, YK24]. try [Hic24]. Tunis [Wil05]. Tunisian [KM24a]. Turkey [Ayd90]. Turkish [DT08, TA08]. Turn [CB00, KZ16, CBSS96]. turned [O’C01, Sch11b]. turtle [SC90, ZL91]. Tutor [Mos04, KO23]. Tutorial [Jaw02, GB14]. tutoring [HOEF89, Hus87, Kar10, ZWC^{+}24]. Twelfth [Ano00d]. Twelve [De 94, Kap09a]. Twelve-Year-Old [De 94]. twentieth [Her02a]. Twenty [HTM^{+}23, Han89, Ano14b]. Twenty-third [Ano14b]. Two [BT04, Cob85, Kal83, Mos07, PB21, Bag06, CV09, CP10, Dek87, Del68, DGFT13, FQL^{+}17, HRW25, Han12c, Her02a, HPRW25, LK18, MPG12, MTBR16, Moo20, MAPHB21, MRRF94, NH94, PGV22, PK23b, Pla14, SECP^{+}23, SPNH12, Sin00, SHAC01, TRDA20, TSV82, TMP10, TTL^{+}17, VMN11, Wai70, WM98, WK04, WT14, WHC09, ZHT25]. two-column [Her02a, WHC09]. two-step [NH94]. two-variable [Moo20, TMP10, WT14]. two-year [WK04]. type [Dum82, Mal93, RSL17, SDVT25]. Types [ARKK23, Cla83, Gli91, PC10, BLP10, Bro10a, ČK15, SC93, TSV82]. typical [GG22b, RBK25]. Tzanakis [Fur19]. u [Dam87]. U.S. [CDW14, SLC95]. U.S.A. [Fey78]. Uhlig [Bur02a, DRR02]. UK [PSS^{+}19, Sfa15, Cd04, FB02]. Unboxing [GP22]. Uncertainty [Zas05, HHS00, Row95, SB12]. Uncontrollable [ASP97]. unconventional [AL00]. uncovers [HK25]. Undergraduate [EH05a, EH05b, HWR23, HB69, Hir81, MSdSGG16, Nar00, RMAGMR22, SSP04, AK19, GBTL23, HRW25, KO23, LZ08, LG16, MRFW^{+}12, RW11a, RW11b, Vin77]. undergraduates [And89, EMRI22]. underlies [SPC^{+}16]. underlying [Sch16, VGJ19]. underserved [Bon14]. Understand [SS06, HBZE98, HS13b, JGR14, Kap09b, Kra02, Ram03, Van77, Wal12]. Understanding [ALR25, BDS^{+}17, CWL07, DB86a, DB86b, GK20, Gre01, Kil01, MD01, MN95, Mor14b, OPC^{+}98, Sha98, SJS12, SS93b, Sin00, TY04, Tir99, Wat07, WW11, Zaz98, ZSWC25, Ang89, BFA22, BHHS09, BJ82, BML21, BIBS^{+}19, BMP20, BMPT24, BE85, CRPPRG16, Cla12, CR09, CG21, Cra12, DC92, DR13, Dri25, EH20, Eke77, FPS16, Fra13, FSY22, GFWL11, GT00, HP25, HS13a, HS14, Her23, HMA13, Jou13, KPH10, KV23, KM25, KČ12, Law00, LCD90, Lin88, Lle12, MS97, MGV^{+}25, MT15, MPG12, McC14, MFJ17, MAPHB21, Ort83a, Ort83b, PXTW17, PS88, PK94, Pla18, PLW81, Roh08, Roh10, RT14, Saj03, SMFL19, Sle82, SSP04, Tho94b, Tzu07, Vol77, WAL18, WJW11, Zah15, Zep82, Mur24]. Understandings [CPP07, ZC06, Gol03, HBZE98, HK02, VYZ21, YTD11]. unexpected [RCH25, SBDM69]. unfamiliar [RL04]. Unfettering [MFG22]. Unfolding [CPPC23]. unidirectional [JM18]. unified [ZER17]. unifying [lD95]. union [DeB97]. Unit [Bus05, CBSS96, CS94, Ern16b, MLK16a, MLK16b]. United [FSY22, Law00, MNRR95, YC95]. units [Lee17, WR96, WMN22, WMN25, Wit84]. unity [FG17]. universal [Add93, CGGGG79a]. Universalis [Add93]. universelle [CGGGG79a]. Universitaires [TLM07]. universities [Cd04, El 86, Ham68]. University [Ano98i, Ano16b, Ano16c, Art23a, GBR05, How75, Jaw02, Kae93, Kae94, MAPHB21, Sfa15, TO13, bMYT98, BVR24, BZ01, DGI23a, DGI23b, EJ23, GRR23b, GRR23a, GLHM24, GS76, GB14, GM84a, GM84b, HYC^{+}97, IM19, JMRW22, Jin07, JWFC^{+}25, Lit03, Mal23b, MK91, MGI23, MAM24, Mur93, Mur95, Nar16, NRV24, PK23c, PJP23, Prz04, RSRVRVVR23, RPP24, SLW83, TLM07, VCS95, dPH13, lRA16]. unknown [LH96, Vla02]. unknowns [HL16]. Unleashing [WPL24]. Unpacking [Lle12, MAM24, SBA22, SS95, Was15, Was22, HRW25, GBC12]. unravel [dSBM21]. Unraveling [DVV20]. Unravelling [Hop85]. unsuccessful [FM09]. Unterricht [Beh68]. upon [SD95]. Upper [Car18, Eng71, FT03, KPH10, VR90]. upper-secondary [FT03]. uptake [Kle12]. USA [Ala12, SS10, VS13]. Usage [Rut05, WN04, Abe78, Dem93, Gla68]. usages [Tro04]. Use [Cla07, Dah04, Jan81, LH22b, Mor00, Mos04, Moy01, Mur04, NC06, Wat05, Wil12, AI08, Art97, Ber04, BLP10, BJJ13, BKH25, CN16, CB17b, DF15, DIW19, DVJV02, DS17, Dra14, Eng68, FKC17, Fro99, Gal19, Gal86, Gla68, GK71, Gue19, GB01, GSH16a, GSH16b, Gut07, HP25, HS13b, JM18, JWFC^{+}25, Kle71, LLW07, Lev10, LYHC18, MWSS17, McQ86, Mos98, Orl71, Pai13, PG20, PvdHPR12, PDT^{+}12, Pic71, PB09, PNH98, RRP91, RNRVM23, Row08, Rut90, RH02, SL19, SRP21, TPHS07, TDGV09, Van03b, VDTV24, VBE21, VO17, VS23, Wal68, Wil16a, Yan12, YE16, ZC22, dV18, Dem93, CB17a]. use-value [Pai13]. Used [FB02, ČK15, Dem97, Hah75, KPH10, PM16]. useful [Als02, Bec19, Fre68, Kla68, Ser68]. usefulness [Her23]. uselessness [Van69a]. uses [Ger88b, Tro04]. Using [BLP^{+}14, CPP09, CVN21, Doe06, Dom11, DE24, EG23, GP90, HY22, HTM18, Isa87, Jon00, Jou13, LW99, MM11, Mog17, MWVE09, Mos07, PGE09, RPP24, RDH09, TJ94, Ves73a, Ves73b, WS08, XM17, Bal80, BC92, BH95, BH94, Boa02, BDS^{+}17, Bur80b, Cas69a, CM92, CSK12, DVV20, DT08, FT03, Hah14, HRW25, Han10, JB21, Jan09a, Jan10, JP09, JGR14, KML22, KB20, KM24a, KB11, KBK^{+}24b, KL14, LC15, Loc11, MMC22, Mar86, MLL19, MPZ16, MHH90, MRC18, Nar16, PPG95, QM94, REL25, SMSW13a, SMSW13b, SV25, Sta89, TN15, TT91, TA08]. utile [Ser68]. utilisant [Cas69a]. utilisation [Art97, BC92]. utility [LS20]. utilizing [Moo20]. Utrecht [Fre78b]. Uworabayeho [Hyl24].

V [GB19, Ott76]. vague [GG22a]. validating [And09, Czo18]. validation [Bal87]. validité [Fre82]. Validity [Fre82, DOV24, Kno77, Sho21, WBM^{+}14]. value [AI12, BBDD25, BBVD25, CVN21, GB01, HA23, Kno77, KSH25, LL23, LS20, MB16, Pai13, SLP^{+}12, SBP11, TP14, Wil12, YC95]. Values [DW17, SS88, Han02, Lev22, Mac00, MA25, Wan25, Wan25]. valuing [Wan25]. Vanpaemel [Bja20a, Bja20b]. Varga [Kry88a]. variability [AG21, GLZBZ15, PAW15, RT14, Tru15]. variable [Bis76, BBG^{+}17, Jer73a, Jer73b, MS09, Mar72, Moo20, Sam85, TMP10, WT14]. variable-free [Mar72]. variables [DG89, GT00, HGM15, Jan74, JM74, Jur77, Kae90, Kei82, KM25, Lep90, MPG12, MAPHB21, SLS74, SGCA18, TF95, Wil16a, Win03]. variance [Tru15]. Variation [Sun11, EEST20, JM18, NPH99, Sch92, SLAY19, VS23, WK04]. variational [JM18]. variations [VS25]. various [KZH^{+}19, RP21, Zep81]. Varna [Fre69c]. vector [Gla69, HK25, LH71]. Vectors [Ros69, SP16, Tay91]. vehicle [Leh24]. vendeurs [Hah99]. Vendors [JS99]. Venn [Fre69b, Van69a, Vre69]. vente [TB09]. venture [KLBMK^{+}13]. ventures [Loc92]. Verbal [NT75, Pug04, BST81, BFG84, Jan74, JR72, Jer73a, Jer73b, MK97, Nes76, Rob75, SSP04, ZWMR16]. verbal-symbolic [ZWMR16]. verbally [CCRD22]. verification [Orl71]. versatile [GT00, TT91]. Verschaffel [Sel00]. verseur [Tro00]. Versions [Eva07, Han12c]. versus [BW13, FKIA15, KBK^{+}17a, KBK^{+}17b, Lan86a, Sko85]. vertical [WH24]. very [SRB16]. via [FCN14, KLwK24, LS11, Roh10, UOF25, YTD11]. Vice [Ber06, Bur05, Bur06]. victim [Car08]. Victoria [Art23a, Gri87]. Video [Leu05, GCFS23, KB11, PPG95]. Videodisc [Str87]. Videopapers [Dim04]. videorecordings [CW96]. vie [TB09]. Vietnamese [Cla07]. View [AF05, Jah01, Ott06, CDW14, FM14, Hei75, Hoy82, Kap74, LBT07, Lar86, Di 19, Mey10, Ott11, PS21, Ram03, SJ04, TJ94, Yer00]. viewed [CB17a, CB17b, Ves76]. Viewing [Dim04, Nol12]. Views [VL04, BZA01, Cha93, FLT14, Mon00, PMG^{+}94, RT09, RS22, RC94, SBA22, SV25, SGTK14, WM09]. Vilma [Hus25]. Vincken [Gre17]. violence [Che23]. Violet [BLP^{+}14]. viral [CSW21]. Virtual [SK98, Bak19b, BdSCdA18, FS16, RL11, ZK15]. visible [ISJ24, KBBK20, RM15]. Vision [CF03, LH08, LQV04, Mac00]. Visit [Nic02]. Vistro [DKL^{+}25, SMWM25]. Vistro-Yu [DKL^{+}25]. Visual [CCW95, Dav93, Gor98, NHH97, RP07, Riv10, SD94, AS04, Arc03, DT12, GL10, HKSS^{+}25, LC81, MH97, MB09, PHS12, RNP13, SMW11, Tri95, WC08, YC90]. Visualisation [Pre86]. Visualising [Bis79b]. Visualization [NN97, KT15]. visualizations [RBK25, RNC21]. visualize [Hod96]. Visualizing [BHLH85]. visually [SD94]. visuospatial [Fin22, KC22]. vitality [Gal82]. Vithal [Edd18]. Viviane [Mal23b]. Vocabulary [Kaz07]. vocational [Bak14, BGW^{+}14, BGW^{+}16, Fit14]. voice [LX16, Sol12, dF04]. Voices [HDM07, FA01, Roo23]. vol [Pea76, Uhl03]. Volkert [Fri21]. Volume [Ano00i, Ano00j, Ano01a, Ano01f, Ano02b, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g, Ano04h, Ano06a]. volumes [Bak19b]. voluntary [VBE21]. voting [Ste76]. vs [KZ16, LH09, Par88, SV25, Sle82]. Vygotskian [Shv22, Cra96, RR17, Wil16b]. Vygotsky [Bar16].

W [Art19, Bag08a, Cor22]. Wait [Cha22a]. Wandlungen [Beh68]. want [Pol76]. wants [Bro23, Pol76]. Warrant [NBZ12]. warrants [WB11, WMPL08]. Warsaw [Ano74c]. was [Abt22, FT03]. Wasserbehältern [Bec77]. water [Bec77]. way [FM09, GG96, PG20, PAW15, Pir25, Roo19, TS99]. Ways [Coo99, Fre79b, Cla02, Cor17, MCKP19, NT01, Nil07, Sch23, Sch25, SBA22]. weapon [SRP21]. Weaving [HH24b, HH24a]. web [CWH19]. web-based [CWH19]. Wei [Mur24]. Weigand [Art19, Tab22]. weird [ZG01]. Welcome [GG22a]. Welsh [Eva07]. Welsh-Medium [Eva07]. Wenzelburger [Ste94]. were [Bek10]. West [Ano74b, Wil78b]. Western [Sfa15]. Where [AAB03, BSV95, Her02b, Nar16, NN97, Nic02, RC02, RC04, SS03, Fit11, Fit14, FT03, Kon20, BA23]. Which [DG03, DBL22, Nil07, RM21, SKN21]. while [Bek10, BC08]. Whither [PV14]. Who [LOW24b, Mor07, Cd04, HH11, VM78, Bro23, LOW24a]. Whole [PM97, vBP19, Ano14b, Byu24a, Byu24b, ČK18, CRPPRG16, Cha18, Jan09b, KM18, MBF25, O’B70, SZP10, SR92, Son16, Tab13, Tha10, Ves69, Ves71, Wol83]. whole-body [KM18]. whole-class [Byu24a, Byu24b, Jan09b, SZP10]. whom [NX23]. whose [RM21]. whys [Jan09a]. Wicks [TR24]. Wickstrom [Sto22]. wide [Bro10a]. Wie [Van76]. will [Han12c]. William [Gre17]. Wilson [Art15a, Tab22, Rut05]. window [BDS^{+}17, CR09]. windows [PXTW17]. Winsløw [Mal23b]. Wiskobas [Tre93]. wiskunde [Gre17]. wissenschaftlicher [JSV75]. Within [BEW99, MG00a, AGL17, AV22, Bat13, BN96, Cho22, Gra04, HO23, LPG05, MA09, PJP23, RNM12, THS06, Tab22, Tru15, VS25, WFB13]. within- [Tru15]. within-task [VS25]. without [BJJ13, IAP19, MPS22]. Wittmann [Fur22]. Witwatersrand [Ano16c]. Wolff [MPS12, Bro12]. woman [BW13]. Women [Fen79, AZ20, ERS19, Mai85, McV24]. Word [Cha06, DVJ98, Sel00, YG99, Ade90, CHM83, CGD21, CSK12, FF21, GS18, Lan86a, LCD90, Lep90, NK77, Ott20, Pal08, RO09, SSPS20, SLS74, SG25, SW95, SGCA23, SRR22, VVV22, VM13, VVM22, Wha96, dKBvdS22]. word-problem [FF21, VVV22]. wording [SGCA23]. words [FS16, MPS22, REA09]. Work [Ass05, Cop98, BA14, Bar87, Cas04, CS19, Dra14, Eng76, FD24, Fit14, FB17, Gai11, GMNB18, HJ12, HMA13, IT77, KR11, KN21, KM24b, LK14b, LK18, PXTW17, Pre03, RNP13, SM21, Sty11, TL15, VCS95, YSRU03, YM16]. work-related [BA14]. workers [Zev11]. Working [Gof93, KN21, MB09, McN00, Shr10, Zev04, Boa02, EVHRC^{+}25, LOMG^{+}20, Lun77]. Workplace [MM03, Nic02, WW07a, WW07b, CW14b, FB17, JS01, NBHK07, Rot14, TP10, TP14, Woo12]. World [Jur06, LH97, Boa93, El 86, GLHM24, KO21, KCS22, Lar86, RM21, SAR19, Sta11, SG98b, Gra20, Jon16, LX16]. worlds [Bye15, dF25]. Worldviews [CW14a]. would [AAB03, RC02, RC04]. write [AK19]. Writing [MBV21, Nte06, SG98a, AK19, BR89, CWS93, KLwK24, LR20, PM00, Rot12, SS15, Way94, PW03]. writing-to-learn [Way94]. Written [FZ05, Pug04, BC08, Cle80, FVJL^{+}19, HRR23a, HRR23b, Hie88, RIA17, SS15, SS93b, Tho94a, VS23]. www.springer.com [Cor22, NX23].

Xenofontos [Gra20]. Xenomathematics [dF25]. XIIIth [Ano72b]. XIIth [Ano71c]. Xinyu [Mur24]. XIth [Ano70]. XIVth [Ano74c]. XIXth [Ano77b]. Xu [vBP19]. xvi [Shv22]. xvii [Ahm23, Cor22]. XVIIIth [Ano77c]. XVIIth [Ano76g]. XVIth [Ano76h]. XVth [Ano74d]. xxi [NX23, Col21, VPR21]. XXth [Ano78d].

Yackel [HJ12]. Yasukawa [GB19]. Year [De 94, Eva07, FB02, SS93b, BJ88, BZ01, Dem93, Dem97, DB86a, DB86b, GAP24, HVN^{+}24, Kie01, LH18, Mat83, PD23, RSRVRVVR23, Rou68, SVVO23, Tru15, VVV22, WC08, WK04, Wea90, Wri94]. Year-olds [SS93b]. Years [IE07, VL04, Ano00h, AV76, Bic82, De 24, Bui76, DOV24, HRR23a, HRR23b, HTM^{+}23, How84, Kru13a, PK23b, PSS^{+}19, Sfa13, SWVO23, Tak18, Urb93, Wal12]. Yeping [Hes13]. Yew [YK20]. Yiming [IPJ25]. Yiping [Cor22]. Young [Bri06, CN00, Eng91, Han19, Hew12, JS99, Mak14, Mon01, Tho94a, VS79, VS25, Zev11, APS22, Ang89, BTV19, Bur80b, Cob87, CSK12, DCHM19, FFL80, JLTM97, KLBMK^{+}13, MS14, Obe19, Sto22, SVLO20, TTL^{+}13, Tur80, VO17, Vou12, VGJ19, vO13]. younger [Fie79]. Yu [DKL^{+}25, SMWM25]. Yüksel [Wan25]. Yves [Art23b]. Yvette [Bro11].

Zahlbegriff [JSV75]. Zambia [Say94]. Zankov [Ott76]. Zazkis [Tsa02].

Zealand [IB05, PAR^{+}24]. zeitgeist [DS21]. zero [OB11]. zero-definitions [OB11]. Zhang [Siu93]. Zilina [Ano72b]. Zofia [Now92]. zones [GGR02, HMT13]. Zooming [KK10]. ZPD [AGL17]. zum [Dam87, Wal76]. Zuñiga [Mur24]. zur [JSV75].

References

Adam:2003:CRC

[AAB03] S. Adam, W. Alangui, and B. Barton. A comment on: Rowlands & Carson, “Where would formal, academic mathematics stand in a curriculum informed by ethnomathematics? A critical review”. Educational Studies in Mathematics, 52(3):327– 335, April 2003. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1024308220169.  See

[RC02].

Aoki:2025:LSM

[AAJW25] Mayu Aoki, Yukiko Asami-Johansson, and Carl Winsløw. Learning to speak mathematically at the Japanese supplementary school in Sweden: critical cases of praxeological anomaly. Educational Studies in Mathematics, 119(1):23–39, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10365-x.

Ayala-Altamirano:2021:FGJ

[AAM21] Cristina Ayala-Altamirano and Marta Molina. Fourth-graders’ justifications in early algebra tasks involving a functional relationship. Educational Studies in Mathematics, 107(2):359– 382, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10036-1.

Albano:2021:TMA

[AAP21] Giovanna Albano, Samuele Antonini, and Anna Pierri. “Tell me about”: a logbook of teachers’ changes from face-to-face to distance mathematics education. Educational Studies in Mathematics, 108(1–2):15–34, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10108-2.

Artzt:1999:CME

[AAT99] Alice F. Artzt and Eleanor Armour-Thomas. A cognitive model for examining teachers’ instructional practice in mathematics: A guide for facilitating teacher reflection. Educational Studies in Mathematics, 40(3):211–235, November 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003871918392.

Abele:1978:UGS

[Abe78] A. Abele. The usage of graphs in solving mathematical problems. Educational Studies in Mathematics, 9(1):7–15, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00352189.

Adler:2005:REF

[ABK^{+}05] Jill Adler, Deborah Ball, Konrad Krainer, Fou-Lai Lin, and Jarmila Novotna. Reflections on an emerging field: Researching mathematics teacher education. Educational Studies in Mathematics, 60(3):359–381, November 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5072-6.

Arsac:1992:TRR

[ABM92] Gilbert Arsac, Nicolas Balacheff, and Michel Mante. Teacher’s role and reproducibility of didactical situations. Educational Studies in Mathematics, 23(1):5–29, February 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302312.

Arnal-Bailera:2020:PSM

[ABOM20] Alberto Arnal-Bailera and Antonio M. Oller-Marcén. Prospective secondary mathematics teachers read Clairaut: professional knowledge and original sources. Educational Studies in Mathematics, 105(2):237–259, October 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09988-7.

Abrantes:2001:MCA

[Abr01] Paulo Abrantes. Mathematical competence for all: Options, implications and obstacles. Educational Studies in Mathematics, 47(2):125–143, July 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1014589220323.

Abrahamson:2009:EDC

[Abr09] Dor Abrahamson. Embodied design: constructing means for constructing meaning. Educational Studies in Mathematics, 70 (1):27–47, January 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9137-1.

Abtahi:2022:WIW

[Abt22] Yasmine Abtahi. What if I was harmful? Reflecting on the ethical tensions associated with teaching the dominant mathematics. Educational Studies in Mathematics, 110(1):149– 165, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10117-1.

Anghileri:2002:ISS

[ABvP02] Julia Anghileri, Meindert Beishuizen, and Kees van Putten. From informal strategies to structured procedures: mind the gap! Educational Studies in Mathematics, 49(2):149–170, February 2002. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1016273328213.

Adda:1988:BR

[ABW88] Josette Adda, N. Balacheff, and Erich Wittmann. Book review. Educational Studies in Mathematics, 19(1):105–114, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00428388.

Atweh:1995:CGS

[AC95] Bill Atweh and Tom Cooper. The construction of gender, social class and mathematics in the classroom. Educational Studies in Mathematics, 28(3):293–310, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274178.

Aguilar:2021:WMC

[AC21] Mario Sánchez Aguilar and Apolo Castaneda. What mathematical competencies does a citizen needs to interpret Mexico’s official information about the COVID-19 pandemic? Educational Studies in Mathematics, 108(1–2):227–248, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10082-9.

Ahmed:2004:HCT

[ACJO04] Afzal Ahmed, Alison Clark-Jeavons, and Adrian Oldknow. How can teaching aids improve the quality of mathematics education. Educational Studies in Mathematics, 56(2–3): 313–328, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040412.
39121.e0.

Artigue:1988:BR

[ACS88] M. Artigue, M. F. Costeroy, and Detlef D. Spalt. Book review. Educational Studies in Mathematics, 19(2):277–286, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00751237.

Adda:1972:QEP

[Add72] Josette Adda. Quelques études pour contribuer à l’observation du comportement mathématique des non-mathématiciens. (French) [Some studies contributing to the observation of mathematical behavior of non-mathematicians]. Educational Studies in Mathematics, 4(3):368–392, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302585.

Adda:1975:LMM

[Add75] Josette Adda. L’incompréhension en mathématiques et les malentendus. (French) [Misunderstandings in mathematics and the misunderstood]. Educational Studies in Mathematics, 6(3): 311–326, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01793614.

Adda:1976:DLP

[Add76] Josette Adda. Difficultés liées à la présentation des questions mathématiques. (French) [Difficulties in presenting mathematical questions]. Educational Studies in Mathematics, 7(1–2): 3–22, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00144355.

Adda:1982:LCP

[Add82] Josette Adda. L’observation de classes et le paradoxe de l’observateur. (French) [Class observation and observer paradox]. Educational Studies in Mathematics, 13(1):21–32, February 1982. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305496.

Adda:1987:GE

[Add87] Josette Adda. Guest editorial. Educational Studies in Mathematics, 18(3):223–227, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386195.

Adda:1993:LSE

[Add93] Josette Adda. Une lumière s’est éteinte. H. Freudenthal — Homo Universalis. (French) [A light is extinguished: H. Freudenthal — universal man]. Educational Studies in Mathematics, 25(1–2):9–19, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274099.

Adetula:1990:LFD

[Ade90] Lawal O. Adetula. Language factor: Does it affect children’s performance on word problems? Educational Studies in Mathematics, 21(4):351–365, August 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304263.

Adedayo:1998:DEG

[Ade98] O. A. Adedayo. Differential effectiveness by gender of instructional methods on achievement in mathematics at tertiary level. Educational Studies in Mathematics, 37(1):83–91, October 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003295525989.

Adler:1988:NBM

[Adl88] Jill Adler. Newspaper-based mathematics for adults in South Africa. Educational Studies in Mathematics, 19(1):59–78, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00428385.

Adler:1997:PIA

[Adl97] Jill Adler. A participatory–inquiry approach and the mediation of mathematical knowledge in a multilingual classroom. Educational Studies in Mathematics, 33(3):235–258, September 1997. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1002976114883.

Adler:2019:LAM

[Adl19] Jill Adler. Learning about mathematics teaching and learning from studying rituals and ritualization? A commentary. Educational Studies in Mathematics, 101(2):291–299, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9876-6.

Ayalon:2008:DRE

[AE08] Michal Ayalon and Ruhama Even. Deductive reasoning: in the eye of the beholder. Educational Studies in Mathematics, 69(3):235–247, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9136-2.

Amit:2005:AAR

[AF05] Miriam Amit and Michael N. Fried. Authority and authority relations in mathematics education: A view from an 8th grade classroom. Educational Studies in Mathematics, 58(2): 145–168, February 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-3618-2.

Afekenstam:1983:SAC

[AG83] Adolf Afekenstam and Karl Greger. Some aspects of children’s ability to solve mathematical problems. Educational Studies in Mathematics, 14(4):369–384, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368235.

Assude:2002:DAN

[AG02] Teresa Assude and Jean-Michel Gelis. La dialectique ancien– nouveau dans l’intégration de Cabri — géomètre à l’école primaire. (French) [The old–new dialectic in the integration of Cabri — geometry in primary school]. Educational Studies in Mathematics, 50(3):259–287, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021293215485.

Albarracin:2014:DPS

[AG14] Lluís Albarracín and Núria Gorgorió. Devising a plan to solve Fermi problems involving large numbers. Educational Studies in Mathematics, 86(1):79–96, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9528-9.

Abu-Ghalyoun:2021:PST

[AG21] Omar Abu-Ghalyoun. Pre-service teachers’ difficulties in reasoning about sampling variability. Educational Studies in Mathematics, 108(3):553–577, November 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10067-8.

Abtahi:2017:CMK

[AGL17] Yasmine Abtahi, Mellony Graven, and Stephen Lerman. Conceptualising the more knowledgeable other within a multidirectional ZPD. Educational Studies in Mathematics, 96(3):

275–287, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9768-1.

Ainley:2015:LRS

[AGP15] Janet Ainley, Robert Gould, and Dave Pratt. Learning to reason from samples: commentary from the perspectives of task design and the emergence of “big data”. Educational Studies in Mathematics, 88(3):405–412, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9592-4.

Avital:1976:MIC

[AH76] Shmuel Avital and Rodney T. Hansen. Mathematical induction in the classroom. Educational Studies in Mathematics, 7(4):399–411, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00452222.

Austin:1979:LME

[AH79] J. L. Austin and A. G. Howson. Language and mathematical education. Educational Studies in Mathematics, 10(2): 161–197, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00230986.

Andrews:2000:CHE

[AH00] Paul Andrews and Gillian Hatch. A comparison of Hungarian and English teachers’ conceptions of mathematics and its teaching. Educational Studies in Mathematics, 43(1):31– 64, July 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017575231667.

Ahmed:2023:BRP

[Ahm23] Salah A. M. Ahmed. Book review: From planning to publishing: navigating the researcher’s journey. James Hiebert, Jinfa Cai, Stephen Hwang, Anne K Morris, Charles Hohensee (2023) Doing Research: A New Researcher’s Guide, Cham: Springer. https://doi.org/10.1007/978-3-031-19078-0, xvii + 136 pages. Hardcover: ISBN: 978-3-031-19077-3, $59.99 US. Paperback:

ISBN: 978-3-031-19080-3, $49.99 US. eBook: ISBN: 978-3-031-19078-0. Educational Studies in Mathematics, 114(1):173–179, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10248-7.

Arcavi:2007:LLH

[AI07] Abraham Arcavi and Masami Isoda. Learning to listen: from historical sources to classroom practice. Educational Studies in Mathematics, 66(2):111–129, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9075-8.

Alcock:2008:DSU

[AI08] Lara Alcock and Matthew Inglis. Doctoral students’ use of examples in evaluating and proving conjectures. Educational Studies in Mathematics, 69(2):111–129, October 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9149-x.

Almog:2012:AVI

[AI12] Nava Almog and Bat-Sheva Ilany. Absolute value inequalities: high school students’ solutions and misconceptions. Educational Studies in Mathematics, 81(3):347–364, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9404-z.

Azrou:2019:WDS

[AK19] Nadia Azrou and Abdelkader Khelladi. Why do students write poor proof texts? A case study on undergraduates’ proof writing. Educational Studies in Mathematics, 102(2):257–274, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-019-09911-9.

Adi:1980:CLA

[AKL80] Helen Adi, Robert Karplus, and Anton E. Lawson. Conditional logic abilities on the four-card problem: Assessment of behavioral and reasoning performances. Educational Studies in Mathematics, 11(4):479–496, November 1980.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00231218.

Avital:1978:MIC

[AL78] Shmuel Avital and Shlomo Libeskind. Mathematical induction in the classroom: Didactical and mathematical issues. Educational Studies in Mathematics, 9(4):429–438, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410588.

Arsenault:2000:INC

[AL00] Cathy Arsenault and Gisèle Lemoyne. Une introduction non classique aux algorithmes d’addition et de soustraction. (French) [An unconventional introduction to addition and subtraction algorithms]. Educational Studies in Mathematics, 42 (3):269–296, May 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017554903071.

Alajmi:2012:HDE

[Ala12] Amal Hussain Alajmi. How do elementary textbooks address fractions? A review of mathematics textbooks in the USA, Japan, and Kuwait. Educational Studies in Mathematics, 79(2):239–261, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9342-1.

Albert:2000:OIS

[Alb00] Lillie R. Albert. Outside-in–inside-out: Seventh-grade students’ mathematical thought processes. Educational Studies in Mathematics, 41(2):109–141, February 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003860225392.

Ale:1981:DFM

[Ale81] Sam O. Ale. Difficulties facing mathematics teachers in developing countries — a case study of Nigeria. Educational Studies in Mathematics, 12(4):479–489, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00308144.

Alladina:1985:SLT

[All85] Safder Alladina. Second language teaching through maths — learning maths through a second language. Educational Studies in Mathematics, 16(2):215–219, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02400939.

Allen:1999:BRK

[All99] Barbara Allen. Book review: Keitel, C. (ed.), Social Justice and Mathematics Education. Educational Studies in Mathematics, 40(1):91–92, September 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003705013341.

Atkins:1991:MRT

[ALO^{+}91] Warren J. Atkins, Gilah C. Leder, Peter J. O’Halloran, Graham H. Pollard, and Peter Taylor. Measuring risk taking. Educational Studies in Mathematics, 22(3):297–308, June 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368343.

Abt:2025:USE

[ALR25] Martin Abt, Katharina Loibl, and Frank Reinhold. Understanding student errors in comparing data sets with boxplots. Educational Studies in Mathematics, 120(1):169–193, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10387-z.

Alsina:2002:TME

[Als02] Claudi Alsina. Too much is not enough teaching maths through useful applications with local and global perspectives. Educational Studies in Mathematics, 50(2):239–250, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021114525054.

Alshwaikh:2024:RRM

[Als24] Jehad Alshwaikh. The racial role of mathematicians and mathematics in maintaining the Israeli military occupation. Educational Studies in Mathematics, 116(3):391–402, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10255-8.

Arsac:1997:SIA

[AM97] Gilbert Arsac and Michel Mante. Situations d’initiation au raisonnement déductif. (French) [situations of initiation to deductive reasoning]. Educational Studies in Mathematics, 33 (1):21–43, June 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1002967421661.

Artigue:2014:NTF

[AM14] Michèle Artigue and Maria Alessandra Mariotti. Networking theoretical frames: the ReMath enterprise. Educational Studies in Mathematics, 85(3):329–355, March 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9522-2.

Archer:2025:BER

[AM25] Louise Archer and Heather Mendick. Becoming exceptional: the role of capital in the development and mediation of mathematics identity and degree trajectories. Educational Studies in Mathematics, 118(1):119–137, January 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10360-2.

Andrade-Molina:2020:BQM

[AMMS20] Melissa Andrade-Molina, Alex Montecino, and Mario Sánchez Aguilar. Beyond quality metrics: defying journal rankings as the philosopher’s stone of mathematics education research. Educational Studies in Mathematics, 103(3):359–374, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09932-9.

Aminifar:2024:LTS

[AMRM24] Elahe Aminifar, Mohsen Malaki, Ulrika Ryan, and Hamid Mesgarani. A long-term shift to include students’ first language in the mathematics teaching practice: socialization events and learning opportunities. Educational Studies in Mathematics, 116(1):113–135, May 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10307-7.

Anderson:1979:OTE

[And79] Johnston Anderson. Objective testing in elementary analysis. Educational Studies in Mathematics, 10(2):227–243, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00230988.

Anderson:1989:SRD

[And89] Johnston Anderson. Sex-related differences on objective tests among undergraduates. Educational Studies in Mathematics, 20(2):165–177, May 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00579461.

Andrews:2009:CSM

[And09] Paul Andrews. Comparative studies of mathematics teachers’ observable learning objectives: validating low inference codes. Educational Studies in Mathematics, 71(2):97–122, June 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9165-x.

Anghileri:1989:IYC

[Ang89] Julie Anghileri. An investigation of young children’s understanding of multiplication. Educational Studies in Mathematics, 20(4):367–385, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00315607.

Anonymous:1968:FRT

[Ano68a] Anonymous. Final recommendations of the participants about the coordination of the teaching of mathematics and physics.

Educational Studies in Mathematics, 1(1–2):245–246, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426248;   http://link.springer.
com/article/10.1007/BF00426248.

Anonymous:1968:PD

[Ano68b] Anonymous. Panel discussion. Educational Studies in Mathematics, 1(1–2):61–79, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00426232.

Anonymous:1968:PTM

[Ano68c] Anonymous. Propositions on the teaching of mathematics. Educational Studies in Mathematics, 1(1–2):244, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/BF00426247;   http://link.springer.
com/content/pdf/10.1007/BF00426247.pdf.

Anonymous:1969:IRM

[Ano69] Anonymous. ICMI report on mathematical contests in secondary education (olympiads) I. Educational Studies in Mathematics, 2(1):80–114, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368987.

Anonymous:1970:XIO

[Ano70] Anonymous. XIth International Olympiad Bucharest, 5–20 July 1969. Educational Studies in Mathematics, 3(1):128– 129, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00381599.

Anonymous:1971:CDG

[Ano71a] Anonymous. The CSMP development of geometry. Educational Studies in Mathematics, 3(3–4):281–285, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302297.

Anonymous:1971:GCR

[Ano71b] Anonymous. Geometry conference recommendations. Educational Studies in Mathematics, 3(3–4):286–287, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302298.

Anonymous:1971:XIO

[Ano71c] Anonymous. XIIth International Olympiad Budapest– Keszthely, 10–24 June 1970. Educational Studies in Mathematics, 3(2):131–132, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305441.

Anonymous:1972:PPL

[Ano72a] Anonymous. Problèmes de la pensée logique dans l’enseignement des mathématiques. (French) [Problems of logical thinking in mathematics education]. Educational Studies in Mathematics, 4(4):491–500, December 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00567003.

Anonymous:1972:XIO

[Ano72b] Anonymous. XIIIth International Olympiad Bratislava–Zilina, 7–20 July 1971. Educational Studies in Mathematics, 4(3): 398–399, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302587.

Anonymous:1974:CRM

[Ano74a] Anonymous. A call for resource materials. Educational Studies in Mathematics, 5(3):390, March 1974. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424559.

Anonymous:1974:ISR

[Ano74b] Anonymous. International symposium on ‘The Role of Geometry in present-day Mathematics Teaching’ scheduled to be held in Bielefeld (West Germany) from September 16 to 20, 1974. Educational Studies in Mathematics, 5(3):389, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424558.

Anonymous:1974:XIOa

[Ano74c] Anonymous. XIVth International Olympiad Warsaw–Toruń, 7–18 July 1972. Educational Studies in Mathematics, 5(3): 385–386, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01424556.

Anonymous:1974:XIOb

[Ano74d] Anonymous. XVth International Olympiad Moscow, 4–17 July 1973. Educational Studies in Mathematics, 5(3):387–388, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01424557.

Anonymous:1975:A

[Ano75a] Anonymous. Announcement. Educational Studies in Mathematics, 6(2):253–257, July 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302546.

Anonymous:1975:Ea

[Ano75b] Anonymous. Errata. Educational Studies in Mathematics, 6 (1):125, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00590031.pdf.

Anonymous:1975:Eb

[Ano75c] Anonymous. Errata. Educational Studies in Mathematics, 6(3):387–388, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF01793619.pdf.

Anonymous:1976:CDS

[Ano76a] Anonymous. Curriculum development — a strategy. Educational Studies in Mathematics, 7(3):351–363, August 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00553921.

Anonymous:1976:E

[Ano76b] Anonymous. Erratum. Educational Studies in Mathematics, 7 (4):535, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00452230.pdf.

Anonymous:1976:F

[Ano76c] Anonymous. Foreword. Educational Studies in Mathematics, 7(3):191–192, August 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00553918.

Anonymous:1976:LT

[Ano76d] Anonymous. Learning to teach. Educational Studies in Mathematics, 7(3):332–350, August 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00553920.

Anonymous:1976:ME

[Ano76e] Anonymous. Mathematics education. Educational Studies in Mathematics, 7(3):193–258, August 1976. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00553919.

Anonymous:1976:P

[Ano76f] Anonymous. Publications. Educational Studies in Mathematics, 7(3):364–367, August 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00553922.

Anonymous:1976:XIMb

[Ano76g] Anonymous. XVIIth International Mathematical Olympiad. Educational Studies in Mathematics, 6(4):501–502, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00411095.

Anonymous:1976:XIMa

[Ano76h] Anonymous. XVIth International Mathematical Olympiad. Educational Studies in Mathematics, 6(4):499–500, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00411094.

Anonymous:1977:A

[Ano77a] Anonymous. Announcements. Educational Studies in Mathematics, 8(2):239–240, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00241028.

Anonymous:1977:XIMb

[Ano77b] Anonymous. XIXth International Mathematical Olympiad (IMO). Educational Studies in Mathematics, 8(4):511–512, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00310951.

Anonymous:1977:XIMa

[Ano77c] Anonymous. XVIIIth International Mathematical Olympiad (IMO). Educational Studies in Mathematics, 8(4):509–510, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00310950.

Anonymous:1978:A

[Ano78a] Anonymous. Announcements. Educational Studies in Mathematics, 9(1):141–142, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF00352197.

Anonymous:1978:FPF

[Ano78b] Anonymous. Five plus four minutes class instruction — a transcript. Educational Studies in Mathematics, 9(1):85–95, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00352194.

Anonymous:1978:T

[Ano78c] Anonymous. Tribute. Educational Studies in Mathematics, 9 (4):504, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00410592.

Anonymous:1978:XIM

[Ano78d] Anonymous. XXth International mathematical Olympiad (IMO). Educational Studies in Mathematics, 9(4):505–506, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00410593.

Anonymous:1980:Aa

[Ano80a] Anonymous. Announcements. Educational Studies in Mathematics, 11(1):133–135, February 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00369163.

Anonymous:1980:Ab

[Ano80b] Anonymous. Announcements. Educational Studies in Mathematics, 11(2):255–256, May 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00304358.

Anonymous:1980:N

[Ano80c] Anonymous. Notice. Educational Studies in Mathematics, 11 (4):497, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00231219.

Anonymous:1981:A

[Ano81a] Anonymous. Announcement. Educational Studies in Mathematics, 12(4):513, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00308146.

Anonymous:1981:CCS

[Ano81b] Anonymous. Cultural contexts of science and mathematics education. Educational Studies in Mathematics, 12(2):266, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305626.

Anonymous:1981:PA

[Ano81c] Anonymous. Preliminary announcement. Educational Studies in Mathematics, 12(2):265, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305625.

Anonymous:1982:Aa

[Ano82a] Anonymous. Announcement. Educational Studies in Mathematics, 13(1):117–118, February 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305501.

Anonymous:1982:Ab

[Ano82b] Anonymous. Announcement. Educational Studies in Mathematics, 13(4):449, November 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00366623.

Anonymous:1983:Aa

[Ano83a] Anonymous. Announcement. Educational Studies in Mathematics, 14(1):107, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00704706.

Anonymous:1983:Ab

[Ano83b] Anonymous. Announcement. Educational Studies in Mathematics, 14(3):323, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00410546.

Anonymous:1983:Ac

[Ano83c] Anonymous. Announcement. Educational Studies in Mathematics, 14(3):324, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00410547.

Anonymous:1983:Ad

[Ano83d] Anonymous. Announcement. Educational Studies in Mathematics, 14(4):434, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368239.

Anonymous:1984:Ab

[Ano84a] Anonymous. Announcement. Educational Studies in Mathematics, 15(2):211, May 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305897.

Anonymous:1984:Aa

[Ano84b] Anonymous. Announcements. Educational Studies in Mathematics, 15(1):103–104, February 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00380442.

Anonymous:1984:CSS

[Ano84c] Anonymous. Conic sections in the sky and on the earth. Educational Studies in Mathematics, 15(2):191–210, May 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305896.

Anonymous:1984:PA

[Ano84d] Anonymous. Preliminary announcement. Educational Studies in Mathematics, 15(4):447, November 1984. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311118.

Anonymous:1985:Aa

[Ano85a] Anonymous. Announcement. Educational Studies in Mathematics, 16(2):229, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02400942.

Anonymous:1985:Ab

[Ano85b] Anonymous. Announcement. Educational Studies in Mathematics, 16(4):434, November 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00417198.

Anonymous:1985:EA

[Ano85c] Anonymous. Editorial announcement. Educational Studies in Mathematics, 16(1):iv, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00354879.

Anonymous:1986:Aa

[Ano86a] Anonymous. Announcement. Educational Studies in Mathematics, 17(2):209, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311524.

Anonymous:1986:Ab

[Ano86b] Anonymous. Announcement. Educational Studies in Mathematics, 17(2):210, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311525.

Anonymous:1986:Ac

[Ano86c] Anonymous. Announcement. Educational Studies in Mathematics, 17(4):453, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311330.

Anonymous:1986:GE

[Ano86d] Anonymous. Guest editorial. Educational Studies in Mathematics, 17(3):211–212, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF00305069.

Anonymous:1987:Ab

[Ano87a] Anonymous. Announcement. Educational Studies in Mathematics, 18(4):454, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF00240992.pdf.

Anonymous:1987:Ac

[Ano87b] Anonymous. Announcement. Educational Studies in Mathematics, 18(4):455, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF00240993.pdf.

Anonymous:1987:Aa

[Ano87c] Anonymous. Announcements. Educational Studies in Mathematics, 18(1):107–108, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367918.

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcement. Educational Studies in Mathematics, 19(3):418, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312457.

Anonymous:1988:Ab

[Ano88b] Anonymous. Announcement. Educational Studies in Mathematics, 19(4):512, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00578698.

Anonymous:1988:CP

[Ano88c] Anonymous. Call for proposals. Educational Studies in Mathematics, 19(4):511, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00578697.

Anonymous:1988:E

[Ano88d] Anonymous. Editorial. Educational Studies in Mathematics, 19(1):1, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00428381.

Anonymous:1988:MEH

[Ano88e] Anonymous. Mathematics education in Hungary today. Educational Studies in Mathematics, 19(3):291–298, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312449.

Anonymous:1988:PAa

[Ano88f]  Anonymous.  Publisher’s announcement.  Educational

Studies in Mathematics, 19(2):115, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00751226.

Anonymous:1988:PAb

[Ano88g]  Anonymous.  Publisher’s announcement.  Educational

Studies in Mathematics, 19(4):iii, November 1988. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578689.

Anonymous:1989:BR

[Ano89a] Anonymous. Book review. Educational Studies in Mathematics, 20(4):469–474, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00315611.

Anonymous:1989:PN

[Ano89b] Anonymous. Publisher’s notice. Educational Studies in Mathematics, 20(4):481, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF00315612.pdf.

Anonymous:1990:Aa

[Ano90a] Anonymous. Announcement. Educational Studies in Mathematics, 21(1):101–107, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311019.

Anonymous:1990:Ab

[Ano90b] Anonymous. Announcement. Educational Studies in Mathematics, 21(4):391–392, August 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF00304267.

Anonymous:1990:CP

[Ano90c] Anonymous. Call for papers. Educational Studies in Mathematics, 21(4):393, August 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00304268.

Anonymous:1990:MHF

[Ano90d] Anonymous. In memoriam: Hans Freudenthal. Educational Studies in Mathematics, 21(6):599–601, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00315947.

Anonymous:1991:Aa

[Ano91a] Anonymous. Announcement. Educational Studies in Mathematics, 22(1):104–105, February 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302719.

Anonymous:1991:Ab

[Ano91b] Anonymous. Announcement. Educational Studies in Mathematics, 22(2):209, April 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00555724.

Anonymous:1991:SIC

[Ano91c] Anonymous. Seventh International Congress on Mathematical Education. Educational Studies in Mathematics, 22(4):407, August 1991. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00369298.

Anonymous:1992:Ab

[Ano92a] Anonymous. Announcement. Educational Studies in Mathematics, 23(5):553–555, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF00571473.

Anonymous:1992:Ac

[Ano92b] Anonymous. Announcement. Educational Studies in Mathematics, 23(6):625–629, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00540065.

Anonymous:1992:Aa

[Ano92c] Anonymous. Announcements. Educational Studies in Mathematics, 23(4):431–441, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF00302444.

Anonymous:1992:CP

[Ano92d] Anonymous. Call for papers. Educational Studies in Mathematics, 23(2):211, April 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00588057.

Anonymous:1992:PN

[Ano92e] Anonymous. Publisher’s notice. Educational Studies in Mathematics, 23(6):iii, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00540059.

Anonymous:1993:A

[Ano93a] Anonymous. Announcement. Educational Studies in Mathematics, 24(2):i, February 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273695.

Anonymous:1993:CP

[Ano93b] Anonymous. Call for papers. Educational Studies in Mathematics, 25(3):279, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273866.

Anonymous:1994:Aa

[Ano94a] Anonymous. Announcement. Educational Studies in Mathematics, 26(1):104, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273304.

Anonymous:1994:Ab

[Ano94b] Anonymous. Announcement. Educational Studies in Mathematics, 27(1):115–116, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01284533.

Anonymous:1994:Ac

[Ano94c] Anonymous. Announcement. Educational Studies in Mathematics, 27(3):316, October 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273734.

Anonymous:1994:Ad

[Ano94d] Anonymous. Announcement. Educational Studies in Mathematics, 27(4):451, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273383.

Anonymous:1994:CP

[Ano94e] Anonymous. Call for papers. Educational Studies in Mathematics, 27(4):452, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273384.

Anonymous:1995:Aa

[Ano95a] Anonymous. Announcement. Educational Studies in Mathematics, 28(1):99, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273860.

Anonymous:1995:Ab

[Ano95b] Anonymous. Announcement. Educational Studies in Mathematics, 28(4):433, June 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274082.

Anonymous:1995:Ac

[Ano95c] Anonymous. Announcement. Educational Studies in Mathematics, 29(1):89–92, July 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273902.

Anonymous:1995:E

[Ano95d] Anonymous. Erratum. Educational Studies in Mathematics, 29(3):295, October 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF01274096.pdf.

Anonymous:1995:ICM

[Ano95e] Anonymous. The International Commission on Mathematical Instruction. Educational Studies in Mathematics, 28(1): 91–98, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273859.

Anonymous:1996:Aa

[Ano96a] Anonymous. Announcement. Educational Studies in Mathematics, 30(1):3–4, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00163750.

Anonymous:1996:Ab

[Ano96b] Anonymous. Announcement. Educational Studies in Mathematics, 30(2):211, March 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302631.

Anonymous:1996:CP

[Ano96c] Anonymous. Call for papers. Educational Studies in Mathematics, 30(4):415, June 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00570832.

Anonymous:1996:E

[Ano96d] Anonymous. Erratum. Educational Studies in Mathematics, 30(1):107, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00163755.pdf.

Anonymous:1996:PN

[Ano96e] Anonymous. Publisher’s note. Educational Studies in Mathematics, 31(3):iii, October 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00376321.

Anonymous:1998:APA

[Ano98a] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 37(1):1–6, October 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017170803924.

Anonymous:1998:AEF

[Ano98b] Anonymous. An appreciation of Efraim Fischbein, 1920– 1998. Educational Studies in Mathematics, 37(1):9–10, October 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1026468219099.

Anonymous:1998:BRa

[Ano98c] Anonymous. Book review. Educational Studies in Mathematics, 36(3):291–298, September 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017113623143.

Anonymous:1998:BRb

[Ano98d] Anonymous. Book review. Educational Studies in Mathematics, 37(3):295–299, December 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017250918191.

Anonymous:1998:CP

[Ano98e]  Anonymous.  A call for papers.  Educational Stud-

ies in Mathematics, 37(1):93–94, October 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026440025960.

Anonymous:1998:C

[Ano98f] Anonymous. Corrigendum. Educational Studies in Mathematics, 36(3):299, September 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017102606537.

Anonymous:1998:Ea

[Ano98g] Anonymous. Editorial. Educational Studies in Mathematics, 36(1):VII, June 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017152809884.

Anonymous:1998:Eb

[Ano98h] Anonymous. Editorial. Educational Studies in Mathematics, 37(1):7–8, October 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1026408620008.

Anonymous:1998:IST

[Ano98i] Anonymous. ICMI study on the teaching and learning of mathematics at university level. Educational Studies in Mathematics, 36(1):91–103, June 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017182500022.

Anonymous:1999:A

[Ano99a] Anonymous. Acknowledgements. Educational Studies in Mathematics, 40(3):v, November 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1023/A%3A1017214705912.

Anonymous:1999:GE

[Ano99b] Anonymous. Guest editorial. Educational Studies in Mathematics, 39(1–3):5–8, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017241105828.

Anonymous:2000:APA

[Ano00a] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 42(3):iii–iv, May 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017502031211.

Anonymous:2000:A

[Ano00b] Anonymous. Announcement. Educational Studies in Mathematics, 42(3):297, May 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017529212379.

Anonymous:2000:BR

[Ano00c] Anonymous. Book review. Educational Studies in Mathematics, 41(3):309–310, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017285915625.

Anonymous:2000:DDT

[Ano00d] Anonymous. Discussion document for the Twelfth ICMI Study. Educational Studies in Mathematics, 42(2):215–224, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017295119128.

Anonymous:2000:E

[Ano00e] Anonymous. Editorial. Educational Studies in Mathematics, 43(3):223–224, November 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017458824644.

Anonymous:2000:IAa

[Ano00f] Anonymous. Instructions for authors. Educational Studies in Mathematics, 43(1):117–124, July 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017506116190.

Anonymous:2000:IAb

[Ano00g] Anonymous. Instructions for authors. Educational Studies in Mathematics, 44(1–3):163–170, December 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017226704804.

Anonymous:2000:OHY

[Ano00h] Anonymous. One hundred years of L’enseignement mathématique. (French) [One hundred years of mathematical education]. Educational Studies in Mathematics, 41(3):311–312, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017230502289.

Anonymous:2000:VCa

[Ano00i] Anonymous. Volume contents. Educational Studies in Mathematics, 42(3):299–300, May 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017555030303.

Anonymous:2000:VCb

[Ano00j] Anonymous. Volume contents. Educational Studies in Mathematics, 43(3):337–338, November 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017217002987.

Anonymous:2001:CV

[Ano01a] Anonymous. Contents of volume 48. Educational Studies in Mathematics, 48(2–3):331, November 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021955027665.

Anonymous:2001:IAa

[Ano01b] Anonymous. Instructions for authors. Educational Studies in Mathematics, 45(1–3):147–154, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017304417670.

Anonymous:2001:IAb

[Ano01c] Anonymous. Instructions for authors. Educational Studies in Mathematics, 46(1–3):299–306, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017374306283.

Anonymous:2001:IAc

[Ano01d] Anonymous. Instructions for authors. Educational Studies in Mathematics, 47(1):117–124, May 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017516811735.

Anonymous:2001:IAd

[Ano01e] Anonymous. Instructions for authors. Educational Studies in Mathematics, 48(1):121–128, October 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017388117187.

Anonymous:2001:VC

[Ano01f] Anonymous. Volume contents. Educational Studies in Mathematics, 47(3):373–374, September 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017347015369.

Anonymous:2002:APA

[Ano02a] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 51(1–2):1–2, July 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022465016593.

Anonymous:2002:CV

[Ano02b] Anonymous. Contents of volume 49. Educational Studies in Mathematics, 49(3):403–404, March 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021843019802.

Anonymous:2002:RES

[Ano02c] Anonymous. Reflections on educational studies in mathematics. Educational Studies in Mathematics, 50(3):251–257, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021259630296.

Anonymous:2002:TIAa

[Ano02d] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 49(1):141–148, January 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016083411624.

Anonymous:2002:TIAb

[Ano02e] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 50(1):115–122, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020547224345.

Anonymous:2002:TIAc

[Ano02f] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 51(1–2):175–182, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022440200663.

Anonymous:2002:VCa

[Ano02g] Anonymous. Volume contents. Educational Studies in Mathematics, 50(3):347–348, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1021276914775.

Anonymous:2002:VCb

[Ano02h] Anonymous. Volume contents. Educational Studies in Mathematics, 51(3):291–292, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023699711566.

Anonymous:2003:APAa

[Ano03a] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 52(1):1–2, January 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023649921439.

Anonymous:2003:APAb

[Ano03b] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 53(1):1–2, May 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024635232622.

Anonymous:2003:TIAa

[Ano03c] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 52(1):93–100, January 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023637007810.

Anonymous:2003:TIAb

[Ano03d] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 53(1):83–90, May 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024625621865.

Anonymous:2003:VCa

[Ano03e] Anonymous. Volume contents. Educational Studies in Mathematics, 52(3):337–338, April 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1024248819622.

Anonymous:2003:VCb

[Ano03f] Anonymous. Volume contents. Educational Studies in Mathematics, 53(3):275–276, September 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026016230525.

Anonymous:2003:VCc

[Ano03g] Anonymous. Volume contents. Educational Studies in Mathematics, 54(2–3):315–316, December 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006238.61235.48.

Anonymous:2004:APAa

[Ano04a] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 55(1–3):1–2, March 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017806.47464.98.

Anonymous:2004:APAb

[Ano04b] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 56(1):1–2, May 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000028414.83972.77.

Anonymous:2004:AIC

[Ano04c] Anonymous. Announcement: The International Commission on Mathematical Instruction: The ICMI Felix Klein and Hans Freudenthal Medals for 2003. Educational Studies in Mathematics, 56(2–3):373–377, July 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000040626.90147.e9.

Anonymous:2004:DDF

[Ano04d] Anonymous. Discussion document: The Fifteenth ICMI Study: The Professional Education and Development of Teachers of Mathematics: The International Commission on Mathematical Instruction. Educational Studies in Mathematics, 56(2–3):359–372, July 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000040631.89460.9f.

Anonymous:2004:IA

[Ano04e] Anonymous. Instructions for authors. Educational Studies in Mathematics, 57(1):149–155, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000047183.33437.e2.

Anonymous:2004:TIAa

[Ano04f] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 55(1–3):275–280, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017722.41489.1c.

Anonymous:2004:TIAb

[Ano04g] Anonymous. Technical instructions for authors. Educational Studies in Mathematics, 56(1):123–128, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000028339.20486.16.

Anonymous:2004:VC

[Ano04h] Anonymous. Volume contents. Educational Studies in Mathematics, 56(2–3):379–380, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3AEDUC.0000040617.
44560.9d.

Anonymous:2005:APA

[Ano05] Anonymous. Advice to prospective authors. Educational Studies in Mathematics, 60(1):1–2, September 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-3697-0.

Anonymous:2006:CV

[Ano06a] Anonymous. Contents of volume 61. Educational Studies in Mathematics, 61(3):409–410, March 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-8200-z.

Anonymous:2006:IAF

[Ano06b] Anonymous. ICMI awards the 2005 Felix Klein and Hans Freudenthal Medals. Educational Studies in Mathematics, 62 (2):231, June 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-8408-y.

Anonymous:2014:NSF

[Ano14a] Anonymous. Nominations sought for the first ICMI Emma Castelnuovo Award. Educational Studies in Mathematics, 87 (2):249, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-014-9571-1.pdf.

Anonymous:2014:TTI

[Ano14b] Anonymous. The Twenty-third ICMI Study: primary mathematics study on whole numbers. Educational Studies in Mathematics, 86(3):307–309, July 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9562-2.

Anonymous:2015:CP

[Ano15] Anonymous. Call for papers. Educational Studies in Mathematics, 89(2):149–150, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-015-9609-z.
pdf.

Anonymous:2016:CP

[Ano16a] Anonymous. Call for papers. Educational Studies in Mathematics, 91(3):295–297, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9686-7.

Anonymous:2016:FKM

[Ano16b] Anonymous. The Felix Klein Medal for 2015 goes to Alan J. Bishop, Emeritus Professor of Education, Monash University, Australia. Educational Studies in Mathematics, 91(1):5–6, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9662-7.

Anonymous:2016:HFM

[Ano16c] Anonymous. The Hans Freudenthal Medal for 2015 goes to Jill Adler, University of the Witwatersrand, Johannesburg, South Africa. Educational Studies in Mathematics, 91(1):3– 4, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9661-8.

Anonymous:2019:A

[Ano19] Anonymous. Announcements. Educational Studies in Mathematics, 100(1):3–6, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9867-7.

Anonymous:2024:AJB

[Ano24a] Anonymous. Alan J. Bishop 1937–2023. Educational Studies in Mathematics, 115(1):9–11, January 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10302-y.

Anonymous:2024:ASF

[Ano24b] Anonymous. Anna Sierpińska, 1947–2023. Educational Studies in Mathematics, 115(1):5–7, January 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10301-z.

Anthony:1996:ALC

[Ant96] Glenda Anthony. Active learning in a constructivist framework. Educational Studies in Mathematics, 31(4):349–369, December 1996. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00369153.

Agterberg:2022:SSR

[AOJ22] D. A. Agterberg, R. J. Oostdam, and F. J. J. M. Janssen. From speck to story: relating history of mathematics to the cognitive demand level of tasks. Educational Studies in Mathematics, 110(1):49–64, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10093-6.

Anxolabehere:1975:AMM

[AP75] D. Anxolabéhère and G. Périquet. Approche des modèles mathématiques de l’évolution des populations par la simulation d’expériences sur ordinateur. (French) [Approach of mathematical models of population evolution by simulation with computer experiments]. Educational Studies in Mathematics, 6(3):327–338, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01793615.

Aunley:2001:ISI

[AP01] Janet Aunley and Dave Pratt. Introducing a special issue on constructing meanings from data. Educational Studies in Mathematics, 45(1–3):1–13, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1013882709661.

Adjiage:2007:ETR

[AP07] Robert Adjiage and François Pluvinage. An experiment in teaching ratio and proportion. Educational Studies in Mathematics, 65(2):149–175, June 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9049-x.

Alderton:2025:CMR

[AP25] Julie Alderton and Nick Pratt. The construction of mathematical reasoning as a pedagogical object. Educational Studies in Mathematics, 120(2):??, October 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10420-1.

Ainley:2001:NCA

[APN01] Janet Ainley, Dave Pratt, and Elena Nardi. Normalising: Children’s activity to construct meanings for trend. Educational Studies in Mathematics, 45(1–3):131–146, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1013822512833.

Arzarello:2009:GSR

[APRS09] Ferdinando Arzarello, Domingo Paola, Ornella Robutti, and Cristina Sabena. Gestures as semiotic resources in the mathematics classroom. Educational Studies in Mathematics, 70 (2):97–109, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9163-z.

Andrews:2022:MCR

[APS22] Paul Andrews, Jöran Petersson, and Judy Sayers. A methodological critique of research on parent-initiated mathematics activities and young children’s attainment. Educational Studies in Mathematics, 109(1):23–40, January 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10080-x.

Artigue:1982:NEE

[AR82] Michèle Artigue and Jacqueline Robinet. Numération a l’école élémentaire. (French). Educational Studies in Mathematics, 13 (2):155–175, May 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00460709.

Alajmi:2007:RRA

[AR07] Amal Alajmi and Robert Reys. Reasonable and reasonableness of answers: Kuwaiti middle school teachers’ perspectives. Educational Studies in Mathematics, 65(1):77–94, May 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9042-4.

Arcavi:2003:RVR

[Arc03] A. Arcavi. The role of visual representations in the learning of mathematics. Educational Studies in Mathematics, 52(3):

215–241, April 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1024312321077.

Andrews:2014:PTF

[ARHS14] Paul Andrews, Andreas Ryve, Kirsti Hemmi, and Judy Sayers. PISA, TIMSS and Finnish mathematics teaching: an enigma in search of an explanation. Educational Studies in Mathematics, 87(1):7–26, September 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9545-3.

Adleff:2023:TMT

[ARKK23] Ann-Kristin Adleff, Natalie Ross, Johannes König, and Gabriele Kaiser. Types of mathematical tasks in lower secondary classrooms in Germany. Educational Studies in Mathematics, 114(3):371–392, November 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10254-9.

Armitage:1969:RBA

[Arm69] J. V. Armitage. The relation between abstract and ‘concrete’ mathematics at school. Educational Studies in Mathematics, 2(2–3):180–188, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00303456.

Artigue:1997:LDC

[Art97] Michèle Artigue. Le logiciel ‘Derive’ comme révélateur de phenomènes didactiques liés à l’utilisation d’environnements informatiques pour l’apprentissage. (French) [The Derive software as a developer of didactic phenomena related to the use of computing environments for learning]. Educational Studies in Mathematics, 33(2):133–169, July 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002996128978.

Artigue:2002:IAI

[Art02] Michèle Artigue. Information about the ICMI Awards. Educational Studies in Mathematics, 51(1–2):173–174, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022449423963.

Artigue:2005:IIC

[Art05] Michèle Artigue. ICMI: The International Commission on Mathematical Instruction, ICMI Awards Committee. Educational Studies in Mathematics, 58(3):421, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-6732-2.

Artigue:2015:BRM

[Art15a] Michèle Artigue. Book review: The mathematics teacher in the digital era, Alison Clark–Wilson, Ornella Robutti and Nathalie Sinclair (Eds.) (2014). Educational Studies in Mathematics, 90(3):357–362, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9632-0.

Arzarello:2015:GPG

[ART15b] Ferdinando Arzarello, Ornella Robutti, and Mike Thomas. Growth point and gestures: looking inside mathematical meanings. Educational Studies in Mathematics, 90(1):19–37, September 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9611-5.

Artigue:2019:BRL

[Art19] Michèle Artigue. Book review: The legacy of Felix Klein. H.-G. Weigand, W. McCallum, M. Menghini, M. Neubrand and G. Schubring (Eds.) (2019). Educational Studies in Mathematics, 102(1):147–152, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-019-09902-w.

Artigue:2023:FKM

[Art23a] Michèle Artigue. The Felix Klein Medal for 2009 goes to Gilah C. Leder, La Trobe University, Bundoora, Victoria, Australia. Educational Studies in Mathematics, 114(1):S1–S2, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-010-9245-6.

Artigue:2023:HFM

[Art23b] Michèle Artigue. The Hans Freudenthal Medal for 2009 goes to Yves Chevallard, IUFM d’Aix-Marseille, France. Educational Studies in Mathematics, 114(1):S7–S8, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-010-9244-7.

Artigue:2023:ISM

[Art23c] Michèle Artigue. ICMI study 21 mathematics education and language diversity. Educational Studies in Mathematics, 114 (1):S3–S6, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-010-9246-5.

Andersson:2022:SPN

[ARW22] Annica Andersson, Ulrika Ryan, and David Wagner. Storylines in public news media about mathematics education and minoritized students. Educational Studies in Mathematics, 111 (2):323–343, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10161-5.

Adda:1987:BR

[AS87] Josette Adda and Anna Sierpińska. Book reviews. Educational Studies in Mathematics, 18(2):209–221, May 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00314727.

Alcock:2004:CSS

[AS04] Lara Alcock and Adrian Simpson. Convergence of sequences and series: Interactions between visual reasoning and the learner’s beliefs about their own role. Educational Studies in Mathematics, 57(1):1–32, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000047051.07646.92.

Alcock:2005:CSS

[AS05] Lara Alcock and Adrian Simpson. Convergence of sequences and series 2: Interactions between nonvisual reasoning and the Learner’s beliefs about their own role. Educational Studies in Mathematics, 58(1):77–100, January 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-2813-5.

Arzarello:2011:STC

[AS11] Ferdinando Arzarello and Cristina Sabena. Semiotic and theoretic control in argumentation and proof activities. Educational Studies in Mathematics, 77(2–3):189–206, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9280-3.

Alcock:2017:IBD

[AS17] Lara Alcock and Adrian Simpson. Interactions between defining, explaining and classifying: the case of increasing and decreasing sequences. Educational Studies in Mathematics, 94 (1):5–19, January 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9709-4.

Arnesen:2025:MIE

[AS25] Kristin Krogh Arnesen and Øystein Ingmar Skartsæterhagen. Mathematical induction in education research: a systematic review. Educational Studies in Mathematics, 119(1):79–100, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10373-x.

Asghari:2009:EEO

[Asg09] Amir H. Asghari. Experiencing equivalence but organizing order. Educational Studies in Mathematics, 71(3):219–234, July 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9173-x.

Aspinwall:1997:UMI

[ASP97] Leslie Aspinwall, Kenneth L. Shaw, and Norma C. Presmeg. Uncontrollable mental imagery: Graphical connections between a function and its derivative. Educational Studies in Mathematics, 33(3):301–317, September 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002976729261.

Assude:2005:TMW

[Ass05] Teresa Assude. Time management in the work economy of a class, a case study: Integration of Cabri in primary school mathematics teaching. Educational Studies in Mathematics, 59 (1–3):183–203, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-5888-0.

Allen:2021:ASD

[AT21] Piata Allen and Tony Trinick. Agency-structure dynamics in an indigenous mathematics education community in times of an existential crisis in education. Educational Studies in Mathematics, 108(1–2):351–368, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10098-1.

Abdu:2025:DML

[ATAB25] Rotem Abdu, Sofia Tancredi, Dor Abrahamson, and Ramesh Balasubramaniam. Demonstrating mathematics learning as the emergence of eye–hand dynamic equilibrium. Educational Studies in Mathematics, 118(3):505–528, March 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10279-0.

Apkarian:2019:SSB

[ATDR19] Naneh Apkarian, Michal Tabach, Tommy Dreyfus, and Chris Rasmussen. The Sierpiński smoothie: blending area and perimeter. Educational Studies in Mathematics, 101(1):19– 34, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-019-09889-4.

Atweh:2024:BRC

[Atw24] Bill Atweh. Book review: Collaboration against postcolonialism. Patricia Paraide, Kay Owens, Charly Muke, Philip Clarkson, & Christopher Owens (2022) Mathematics education in a neocolonial country: The case of Papua New Guinea. Educational Studies in Mathematics, 115(3):517–524, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10283-4.

Audibert:1983:PRP

[Aud83] Gerard Audibert. Processus de recherche d’un problème de géometrie chez l’élève de l’enseignement secondaire. (French) [Process of research of a geometry problem in secondaryeducation pupils]. Educational Studies in Mathematics, 14 (2):155–181, May 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303684.

Austin:1981:B

[Aus81] Keith Austin. Busprongs. Educational Studies in Mathematics, 12(3):369–371, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311066.

Artigue:1976:QRE

[AV76] Michèle Artigue and Jacqueline Viennot. Quelques reflexions sur l’enseignement de la numération aux enfants de 7, 8 ou 9 ans. (French) [Some reflections on the teaching of numeration to children of 7, 8 or 9 years]. Educational Studies in Mathematics, 6(4):415–428, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00411090.

Agudelo-Valderrama:2008:PCM

[AV08] Cecilia Agudelo-Valderrama. The power of Colombian mathematics teachers’ conceptions of social/institutional factors of teaching. Educational Studies in Mathematics, 68(1):37– 54, May 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-007-9107-z.

Alwast:2022:MPS

[AV22] Alina Alwast and Katrin Vorhölter. Measuring pre-service teachers’ noticing competencies within a mathematical modeling context — an analysis of an instrument. Educational Studies in Mathematics, 109(2):263–285, February 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10102-8.

Andersson:2015:AMH

[AVM15] Annica Andersson, Paola Valero, and Tamsin Meaney. “I am [not always] a maths hater”: Shifting students’ identity narratives in context. Educational Studies in Mathematics, 90(2): 143–161, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9617-z.

Ayalon:2022:RSE

[AWE22] Michal Ayalon, Karina J. Wilkie, and Katrin Hajjar Eid. Relating students’ emotions during argumentative discourse to their learning of real-life functional situations. Educational Studies in Mathematics, 110(1):23–48, May 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10121-5.

Ayalon:2015:FRL

[AWL15] Michal Ayalon, Anne Watson, and Steve Lerman. Functions represented as linear sequential data: relationships between presentation and student responses. Educational Studies in Mathematics, 90(3):321–339, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9628-9.

Aydin:1990:MTP

[Ayd90] Yusuf Aydin. Mathematics teacher preparation in Turkey today. Educational Studies in Mathematics, 21(5):471–478, October 1990. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00398865.

Adiredja:2020:LEL

[AZ20] Aditya P. Adiredja and Michelle Zandieh. The lived experience of linear algebra: a counter-story about women of color in mathematics. Educational Studies in Mathematics, 104(2): 239–260, June 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-020-09954-3.

B:1988:E

[B.88] A. J. B. Editorial. Educational Studies in Mathematics, 19 (2):117–118, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00751227.

Bakker:2014:BCA

[BA14] Arthur Bakker and Sanne F. Akkerman. A boundary-crossing approach to support students’ integration of statistical and work-related knowledge. Educational Studies in Mathematics, 86(2):223–237, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-013-9517-z.
pdf.

Bofferding:2023:WDS

[BA23] Laura Bofferding and Mahtob Aqazade. “Where does the square go?”: reinterpreting shapes when solving a tangram puzzle. Educational Studies in Mathematics, 112(1):25–47, January 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10166-0.

Barahmand:2025:ISR

[BA25a] Ali Barahmand and Nargessadat Attari. Investigating students’ reasoning in generalization of deconstructive figural patterns. Educational Studies in Mathematics, 118(1):53–70, January 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10355-z.

Bikner-Ahsbahs:2025:HET

[BA25b] Angelika Bikner-Ahsbahs. How eye tracking may advance the field of mathematics education: a commentary on the ESM special issue on eye-tracking research in mathematics education. Educational Studies in Mathematics, 118(3):555–574, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10386-0.

Bachmann:1971:G

[Bac71] Friedrich Bachmann. n-gons. Educational Studies in Mathematics, 3(3–4):288–309, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302299.

Backhouse:1989:BML

[Bac89] John K. Backhouse. Better mathematics for low attainers. Educational Studies in Mathematics, 20(1):105–110, February 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00356044.

Bachman:2021:TDJ

[Bac21] Rachel Bachman. A tribute to Dr. Jean Schmittau (1937– 2015). Educational Studies in Mathematics, 106(3):479–483, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-10025-w.

Bagnato:1973:TCM

[Bag73] Robert A. Bagnato. Teaching college mathematics by question-and-answer. Educational Studies in Mathematics, 5 (2):185–192, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01421025.

Bagni:2006:SCD

[Bag06] Giorgio T. Bagni. Some cognitive difficulties related to the representations of two major concepts of set theory. Educational Studies in Mathematics, 62(3):259–280, July 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-8545-3.

Bagni:2008:CBA

[Bag08a] Giorgio T. Bagni. Centenary birth anniversary of E. W. Beth (1908–1964). Educational Studies in Mathematics, 69(1):73– 76, September 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9135-3.

Bagni:2008:RRH

[Bag08b] Giorgio T. Bagni. Richard Rorty (1931–2007) and his legacy for mathematics educators. Educational Studies in Mathematics, 67(1):1–2, January 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-007-
9096-y.

Bagni:2010:MPS

[Bag10] Giorgio T. Bagni. Mathematics and positive sciences: a reflection following Heidegger. Educational Studies in Mathematics, 73(1):75–85, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-009-
9214-0.

Bakker:2014:CDV

[Bak14] A. Bakker. Characterising and developing vocational mathematical knowledge. Educational Studies in Mathematics, 86 (2):151–156, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-014-9560-4.pdf.

Bakker:2016:BRN

[Bak16] Arthur Bakker. Book review: Networking theories as an example of boundary crossing. Angelika Bikner-Ahsbahs and Susanne Prediger (Eds.) (2014). Networking of theories as a research practice in mathematics education. Educational Studies in Mathematics, 93(2):265–273, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9715-6.

Bakker:2019:E

[Bak19a] Arthur Bakker. Editorial. Educational Studies in Mathematics, 100(1):1–2, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9870-
z;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9870-z.pdf.

Bakker:2019:VSI

[Bak19b] Arthur Bakker. A virtual special issue to celebrate 100 volumes of Educational Studies in Mathematics: former editors’ choices. Educational Studies in Mathematics, 101(1):1–5, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-9882-3;   http://link.springer.com/
content/pdf/10.1007/s10649-019-9882-3.pdf.

Bakker:2020:E

[Bak20] Arthur Bakker. Editorial. Educational Studies in Mathematics, 103(1):1–5, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09928-
0;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09928-0.pdf.

Balk:1971:AHM

[Bal71] G. D. Balk. Application of heuristic methods to the study of mathematics at school. Educational Studies in Mathematics, 3 (2):133–146, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305442.

Balacheff:1980:EAG

[Bal80] N. Balacheff. Une étude, à l’aide de graphes, de démonstrations mathématiques formulées par des élèves. (French) [A study, using graphs, of mathematical demonstrations formulated by pupils]. Educational Studies in Mathematics, 11(1):91–111, February 1980. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00369161.

Balacheff:1987:PPS

[Bal87] Nicolas Balacheff. Processus de preuve et situations de validation. (French) [Proof process and validation situations]. Educational Studies in Mathematics, 18(2):147–176, May 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00314724.

Baldy:1988:EDD

[Bal88] Rene Baldy. De l’espace du dessin à celui de l’objet. Une activité de mises en correspondances entre des dessins en perspective cavalière et des objects réels. (French) [From the space of the drawing to that of the object. An activity of mappings between cavalier perspective drawings and real objects]. Educational Studies in Mathematics, 19(1):43–57, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00428384.

Balacheff:1990:IRI

[Bal90] N. Balacheff. ICME-6 report of the international group psychology of mathematics education. Educational Studies in Mathematics, 21(2):193–197, April 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304901.

Bansilal:2011:ARS

[Ban11] Sarah Bansilal. Assessment reform in South Africa: opening up or closing spaces for teachers? Educational Studies in Mathematics, 78(1):91–107, September 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9311-8.

Barra:1975:C

[Bar75] Mario Barra. The cycloid. Educational Studies in Mathematics, 6(1):93–98, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00590027.

Baruk:1987:RPC

[Bar87] Stella Baruk. The reality and potential of correcting pupils’ work in mathematics. Educational Studies in Mathematics, 18(3):273–303, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00386199.

Baroody:1988:MAD

[Bar88] Arthur J. Baroody. Mental-addition development of children classified as mentally handicapped. Educational Studies in Mathematics, 19(3):369–388, August 1988. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312453.

Barton:1996:MSE

[Bar96] Bill Barton. Making sense of ethnomathematics: Ethnomathematics is making sense. Educational Studies in Mathematics, 31(1–2):201–233, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00143932.

Barwell:2003:PAI

[Bar03] R. Barwell. Patterns of attention in the interaction of a primary school mathematics student with English as an additional language. Educational Studies in Mathematics, 53(1): 35–59, May 2003. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1024659518797.

Barbin:2007:ASE

[Bar07] Evelyne Barbin. On the argument of simplicity in elements and schoolbooks of geometry. Educational Studies in Mathematics, 66(2):225–242, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9074-9.

Barwell:2009:RDC

[Bar09] Richard Barwell. Researchers’ descriptions and the construction of mathematical thinking. Educational Studies in Mathematics, 72(2):255–269, November 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9202-4.

Barwell:2016:FIM

[Bar16] Richard Barwell. Formal and informal mathematical discourses: Bakhtin and Vygotsky, dialogue and dialectic. Educational Studies in Mathematics, 92(3):331–345, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9641-z.

Barahmand:2019:TMI

[Bar19] Ali Barahmand. “Tending to the midpoint”: an intuitive rule in mathematical thinking. Educational Studies in Mathematics, 102(2):221–238, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-019-09909-3.

Barnes:2021:ELM

[Bar21] Alison Barnes. Enjoyment in learning mathematics: its role as a potential barrier to children’s perseverance in mathematical reasoning. Educational Studies in Mathematics, 106(1):45–63, January 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09992-x.

Bastier:1969:GR

[Bas69] J. Bastier. Games and relations. Educational Studies in Mathematics, 1(4):415–421, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02396315.

Bass:2017:DOL

[Bas17] Hyman Bass. Designing opportunities to learn mathematics theory-building practices. Educational Studies in Mathematics, 95(3):229–244, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Battey:2013:GMT

[Bat13] Dan Battey. “Good” mathematics teaching for students of color and those in poverty: the importance of relational interactions within instruction. Educational Studies in Mathematics, 82(1):125–144, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9412-z.

Bauersfeld:1980:HDC

[Bau80] Heinrich Bauersfeld. Hidden dimensions in the so-called reality of a mathematics classroom. Educational Studies in Mathematics, 11(1):23–41, February 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00369158.

Bauersfeld:1992:CCS

[Bau92] Heinrich Bauersfeld. Classroom cultures from a social constructivist’s perspective. Educational Studies in Mathematics, 23(5):467–481, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00571468.

Bazzini:1988:NGP

[Baz88] Luciana Bazzini. The new government programs for Italian primary schools. Educational Studies in Mathematics, 19(3): 389–393, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00312454.

Bazzini:1990:ISM

[Baz90] Luciana Bazzini. The IEA study of mathematics II: Contexts and outcomes of school mathematics. Educational Studies in Mathematics, 21(4):383–390, August 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304266.

Bazzini:2001:GMT

[Baz01] Luciana Bazzini. From grounding metaphors to technological devices: A call for legitimacy in school mathematics. Educational Studies in Mathematics, 47(3):259–271, September 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015143318759.

Brookhart:2004:MMA

[BAZF04] Susan M. Brookhart, Marissa Andolina, Megan Zuza, and Rosalie Furman. Minute math: An action research study of student self-assessment. Educational Studies in Mathematics, 57(2):213–227, September 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000049293.55249.d4.

Bussi:2003:RPT

[BB03] Maria G. Bartolini Bussi and Luciana Bazzini. Research, practice and theory in didactics of mathematics: Towards dialogue between different fields. Educational Studies in Mathematics, 54(2–3):203–223, December 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006169.86495.49.

Brandsaeter:2025:PMC

[BB25] Andreas Brandsæter and Runar Lie Berge. Promoting mathematical competence development through programming activities. Educational Studies in Mathematics, 119(2):225–247, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10380-y.

Bini:2023:HMI

[BBAR23] Giulia Bini, Angelika Bikner-Ahsbahs, and Ornella Robutti. “How to meme it”: reverse engineering the creative process of mathematical Internet memes. Educational Studies in Mathematics, 112(1):141–174, January 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10173-1.

Boels:2025:CSS

[BBDD25] Lonneke Boels, Arthur Bakker, Wim Van Dooren, and Paul Drijvers. Correction to: Secondary school students’ strategies when interpreting histograms and case-value plots: an eye-tracking study. Educational Studies in Mathematics, 119 (1):177–178, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10379-5. See

[BBVD25].

Barbe:2005:DRT

[BBEG05] Joaquim Barbé, Marianna Bosch, Lorena Espinoza, and Josep Gascón. Didactic restrictions on the teacher’s practice: The case of limits of functions in Spanish high schools. Educational Studies in Mathematics, 59(1–3):235–268, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5889-z.

Baker:1971:PAM

[BBF71] J. E. Baker, M. Bruckheimer, and H. G. Flegg. A pedagogic approach to morphisms. Educational Studies in Mathematics, 4(2):252–263, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00303390.

Bussi:1999:EAT

[BBFG99] Maria G. Bartolini Bussi, Mara Boni, Franca Ferri, and Rossella Garuti. Early approach to theoretical thinking: gears in primary school. Educational Studies in Mathematics, 39 (1–3):67–87, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003707727896.

Blanton:2017:PFG

[BBG^{+}17] Maria Blanton, Bárbara M. Brizuela, Angela Murphy Gardiner, Katie Sawrey, and Ashley Newman-Owens. A progression in first-grade children’s thinking about variable and variable notation in functional relationships. Educational Studies in Mathematics, 95(2):181–202, June 2017. CODEN EDS-MAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Brunetto:2022:TSC

[BBL22] Domenico Brunetto, Giulia Bernardi, and Peter Liljedahl. Teaching as a system: COVID-19 as a lens into teacher change. Educational Studies in Mathematics, 110(1):65–81, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10107-3.

Barwell:2007:MIM

[BBS07] Richard Barwell, Bill Barton, and Mamokgethi Setati. Multilingual issues in mathematics education: introduction. Educational Studies in Mathematics, 64(2):113–119, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9065-x.

Boels:2025:SSS

[BBVD25] Lonneke Boels, Arthur Bakker, Wim Van Dooren, and Paul Drijvers. Secondary school students’ strategies when interpreting histograms and case-value plots: an eye-tracking study. Educational Studies in Mathematics, 118(3):479–503, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10351-3. See correction [BBDD25].

Bessot:1978:EAD

[BC78] Annie Bessot and Claude Comiti. Une étude sur l’approche du nombre par l’élève du cours préparatoire. (French) [A study of the pupil’s approach in the preparatory course]. Educational Studies in Mathematics, 9(1):17–39, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00352190.

Bessot:1982:APO

[BC82] A. Bessot and C. Comiti. Appropriation des propriétés ordinales du nombre par l’élève du cours préparatoire. (French) [Appropriation of the ordinal properties of number by the pupil in the preparatory course]. Educational Studies in Mathematics, 13(1):59–88, February 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305498.

Bellemain:1992:SOS

[BC92] F. Bellemain and B. Capponi. Spécificité de l’organisation d’une séquence d’enseignement lors de l’utilisation de l’ordinateur. (French) [Specificity of the organization of a teaching sequence when using the computer]. Educational Studies in Mathematics, 23(1):59–97, February 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302314.

Borba:1996:SCT

[BC96] Marcelo C. Borba and Jere Confrey. A student’s construction of transformations of functions in a multiple representational environment. Educational Studies in Mathematics, 31(3):319– 337, October 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00376325.

Buendia:2005:PPA

[BC05] Gabriela Buendía and Francisco Cordero. Prediction and the periodical aspect as generators of knowledge in a social practice framework. Educational Studies in Mathematics, 58(3): 299–333, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-2295-5.

Brizuela:2008:RPM

[BC08] Bárbara M. Brizuela and Gabrielle A. Cayton. The roles of punctuation marks while learning about written numbers. Educational Studies in Mathematics, 68(3):209–225, July 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9109-x.

Brown:2012:DDA

[BC12] Laurinda Brown and Alf Coles. Developing “deliberate analysis” for learning mathematics and for mathematics teacher education: how the enactive approach to cognition frames reflection. Educational Studies in Mathematics, 80(1–2):217– 231, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9389-7.

Bakker:2019:BSM

[BCE^{+}19] Arthur Bakker, Jinfa Cai, Lyn English, Gabriele Kaiser, Vilma Mesa, and Wim Van Dooren. Beyond small, medium, or large: points of consideration when interpreting effect sizes. Educational Studies in Mathematics, 102(1):1–8, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-019-09908-4;   http://link.springer.
com/content/pdf/10.1007/s10649-019-09908-4.pdf. See

commentary [PM19].

Ben-Chaim:1990:CTP

[BCFC90] David Ben-Chaim, Barbara Fresko, and Miriam Carmeli. Comparison of teacher and pupil perceptions of the learning environment in mathematics classes. Educational Studies in Mathematics, 21(5):415–429, October 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00398861.

Ben-Chaim:1998:PRA

[BCFF^{+}98] David Ben-Chaim, James T. Fey, William M. Fitzgerald, Catherine Benedetto, and Jane Miller. Proportional reasoning among 7th grade students with different curricular experiences. Educational Studies in Mathematics, 36(3):247–273, September 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003235712092.

Ben-Chaim:1989:AAC

[BCLH89] David Ben-Chaim, Glenda Lappan, and Richard T. Houang. Adolescents’ ability to communicate spatial information: Analyzing and effecting students’ performance. Educational Studies in Mathematics, 20(2):121–146, May 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00579459.

Bakker:2021:FTM

[BCZ21] Arthur Bakker, Jinfa Cai, and Linda Zenger. Future themes of mathematics education research: an international survey before and during the pandemic. Educational Studies in Mathematics, 107(1):1–24, May 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10049-w.

Bernet:1972:INI

[BD72] Théo Bernet and André Delessert. Un inventaire des notions impliquées dans un enseignement général de mathématique à l’école secondaire. (French) [An inventory of the notions involved in a general education of mathematics in secondary school]. Educational Studies in Mathematics, 4(4):450–465, December 1972. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00567000.

Breidenbach:1992:DPC

[BDHN92] Daniel Breidenbach, Ed Dubinsky, Julie Hawks, and Devilyna Nichols. Development of the process conception of function. Educational Studies in Mathematics, 23(3):247–285, June 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02309532.

Bruce:2017:UGR

[BDS^{+}17] Catherine D. Bruce, Brent Davis, Nathalie Sinclair, Lynn McGarvey, David Hallowell, Michelle Drefs, Krista Francis, Zachary Hawes, Joan Moss, Joanne Mulligan, Yukari Okamoto, Walter Whiteley, and Geoff Woolcott. Understanding gaps in research networks: using “spatial reasoning” as a window into the importance of networked educational research. Educational Studies in Mathematics, 95(2):143–161, June 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Borba:2018:IVL

[BdSCdA18] Marcelo C. Borba, Aparecida Santana de Souza Chiari, and Helber Rangel Formiga Leite de Almeida. Interactions in virtual learning environments: new roles for digital technology. Educational Studies in Mathematics, 98(3):269–286, July 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9812-9.

Byers:1984:CFM

[BE84] Victor Byers and Stanley Erlwanger. Content and form in mathematics. Educational Studies in Mathematics, 15(3): 259–275, August 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00312077.

Byers:1985:MMU

[BE85] Victor Byers and Stanley Erlwanger. Memory in mathematical understanding. Educational Studies in Mathematics, 16(3): 259–281, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00776733.

Ball:2003:P

[BE03] Deborah Loewenberg Ball and Ruhama Even. Preface. Educational Studies in Mathematics, 54(2–3):139–146, December 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006236.67566.bd.

Buenrostro:2023:BMP

[BE23] Patricia Buenrostro and Nadav Ehrenfeld. Beyond mere persistence: a conceptual framework for bridging perseverance and mathematical sensemaking in teaching and teacher learning. Educational Studies in Mathematics, 114(2):199–221, October 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10240-1.

Buchter:2024:CRB

[BEBL^{+}24] Theresa Büchter, Andreas Eichler, Katharina Böcherer-Linder, Markus Vogel, Karin Binder, Stefan Krauss, and Nicole Steib. Covariational reasoning in Bayesian situations. Educational Studies in Mathematics, 115(3):481–505, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10274-5.

Beck:1977:AWA

[Bec77] Uwe Beck. Auslaufvorgänge aus Wasserbehältern — eine anwendungsorientierte Einführung in die Analysis. (German) [outflow processes from water reservoirs — an applicationoriented introduction to analysis]. Educational Studies in Mathematics, 8(4):439–460, December 1977. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310947.

Becker:1980:MOP

[Bec80] Jerry P. Becker. 1978 mathematics olympiads in the People’s Republic of China. Educational Studies in Mathematics, 11(4): 463–478, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00231217.

Becker:1986:BR

[Bec86] Jerry P. Becker. Book review. Educational Studies in Mathematics, 17(1):89–96, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302381.

Beckers:2019:WPM

[Bec19] Danny Beckers. Why to publish on mathematics education so as to be useful? Educational Studies in Mathematics and its founder Hans Freudenthal. Educational Studies in Mathematics, 101(1):7–17, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-019-9881-
4;   http://link.springer.com/content/pdf/10.1007/
s10649-019-9881-4.pdf.

Becevic:2023:WTC

[Bec23] Semir Becevic. When teachers construct tests for assessing students’ competencies: a taxonomy. Educational Studies in Mathematics, 114(2):315–336, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10233-0.

Beccuti:2024:MSP

[Bec24] Francesco Beccuti. Meaning and subjectivity in the PISA mathematics frameworks: a sociological approach. Educational Studies in Mathematics, 116(1):49–65, May 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10296-z.

Begle:1969:RRI

[Beg69] E. G. Begle. The role of research in the improvement of mathematics education. Educational Studies in Mathematics, 2(2–3): 232–244, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303460.

Begle:1971:TDI

[Beg71] E. G. Begle. Time devoted to instruction and student achievement. Educational Studies in Mathematics, 4(2):220–224, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00303386.

Behnke:1968:WAS

[Beh68] H. Behnke. Wandlungen in der Auswahl der Stoffe für den Unterricht. (German) [Changes in the choice of subjects for teaching]. Educational Studies in Mathematics, 1(1–2):161– 165, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00426241.

Bekdemir:2010:PST

[Bek10] Mehmet Bekdemir. The pre-service teachers’ mathematics anxiety related to depth of negative experiences in mathematics classroom while they were students. Educational Studies in Mathematics, 75(3):311–328, December 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9260-7.

Bell:1976:SPP

[Bel76] A. W. Bell. A study of pupils’ proof — explanations in mathematical situations. Educational Studies in Mathematics, 7 (1–2):23–40, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00144356.

Bell:1979:LPA

[Bel79] A. W. Bell. The learning of process aspects of mathematics. Educational Studies in Mathematics, 10(3):361–387, August 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00314662.

Bell:1985:BR

[Bel85] A. W. Bell. Book review. Educational Studies in Mathematics, 16(1):103–110, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00354887.

Bell:1993:GE

[Bel93a] Alan Bell. Guest editorial. Educational Studies in Mathematics, 24(1):1–4, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273292.

Bell:1993:PDT

[Bel93b] Alan Bell. Principles for the design of teaching. Educational Studies in Mathematics, 24(1):5–34, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273293.

Bell:1993:SED

[Bel93c] Alan Bell. Some experiments in diagnostic teaching. Educational Studies in Mathematics, 24(1):115–137, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273297.

Bergeron:1994:MNH

[Ber94] Jacques C. Bergeron. In memoriam Nicolas Herscovics 1935–1994. Educational Studies in Mathematics, 27(2):v–vi, September 1994. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01278917.

Berger:2004:FUM

[Ber04] Margot Berger. The functional use of a mathematical sign. Educational Studies in Mathematics, 55(1–3):81–102, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017672.49486.f2.

Berge:2006:CNS

[Ber06] Analía Bergé. Convergence of numerical sequences — a commentary on The Vice: Some Historically Inspired and Proof Generated Steps to Limits of Sequences” by R. P. Burn. Educational Studies in Mathematics, 61(3):395–402, March 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-8754-9. See [Bur05, Bur06].

Berge:2008:CPS

[Ber08] Analía Bergé. The completeness property of the set of real numbers in the transition from calculus to analysis. Educational Studies in Mathematics, 67(3):217–235, March 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9101-5.

Berger:2019:DRS

[Ber19] Margot Berger. Different reading styles for mathematics text. Educational Studies in Mathematics, 100(2):139–159, February 2019. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9871-y.

Beswick:2007:TBM

[Bes07] Kim Beswick. Teachers’ beliefs that matter in secondary mathematics classrooms. Educational Studies in Mathematics, 65 (1):95–120, May 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9035-3.

Beswick:2012:TBA

[Bes12] Kim Beswick. Teachers’ beliefs about school mathematics and mathematicians’ mathematics and their relationship to practice. Educational Studies in Mathematics, 79(1):127–147, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9333-2.

Brown:1999:SIA

[BEW99] Tony Brown, Frank Eade, and Dave Wilson. Semantic innovation: Arithmetical and algebraic metaphors within narratives of learning. Educational Studies in Mathematics, 40(1): 53–70, September 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003738611811.

Bottino:1996:ETC

[BF96] Rosa Maria Bottino and Fulvia Furinghetti. The emerging of teachers’ conceptions of new subjects inserted in mathematics programs: the case of informatics. Educational Studies in Mathematics, 30(2):109–134, March 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302626.

Baccaglini-Frank:2021:TSY

[BF21] Anna Baccaglini-Frank. To tell a story, you need a protagonist: how dynamic interactive mediators can fulfill this role and foster explorative participation to mathematical discourse. Educational Studies in Mathematics, 106(2):291–312, February 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-10009-w.

Bagossi:2022:ACF

[BFA22] Sara Bagossi, Federica Ferretti, and Ferdinando Arzarello. Assessing covariation as a form of conceptual understanding through comparative judgement. Educational Studies in Mathematics, 111(3):469–492, November 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10178-w.

Borovcnik:1983:BR

[BFG83] M. Borovcnik, R. Fischer, and G. Glaeser. Book reviews. Educational Studies in Mathematics, 14(1):101–105, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00704705.

Bell:1984:COV

[BFG84] Alan Bell, Efraim Fischbein, and Brian Greer. Choice of operation in verbal arithmetic problems: the effects of number size, problem structure and context. Educational Studies in Mathematics, 15(2):129–147, May 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305893.

Bruckheimer:1969:ACM

[BG69] M. Bruckheimer and N. Gowar. Apparent conflicts in maths education. Educational Studies in Mathematics, 2(1):115– 122, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00368988.

Brousseau:2005:DHS

[BG05] Guy Brousseau and Patrick Gibel. Didactical handling of students’ reasoning processes in problem solving situations. Educational Studies in Mathematics, 59(1–3):13–58, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-2532-y.

Bakker:2006:HPM

[BG06] Arthur Bakker and Koeno P. E. Gravemeijer. An historical phenomenology of mean and median. Educational Studies in Mathematics, 62(2):149–168, June 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-7099-8.

Batanero:1994:TRM

[BGSW94] M. C. Batanero, J. D. Godino, H. G. Steiner, and E. Wenzelburger. The training of researchers in mathematics education — results from an international survey. Educational Studies in Mathematics, 26(1):95–102, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273302.

Bosch:2017:DBT

[BGT17] Marianna Bosch, Josep Gascón, and María Trigueros. Dialogue between theories interpreted as research praxeologies: the case of APOS and the ATD. Educational Studies in Mathematics, 95(1):39–52, May 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Bakker:2014:PRL

[BGW^{+}14] Arthur Bakker, Djonie Groenveld, Monica Wijers, Sanne F. Akkerman, and Koeno P. E. Gravemeijer. Proportional reasoning in the laboratory: an intervention study in vocational education. Educational Studies in Mathematics, 86(2):211– 221, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-012-9393-y.pdf. See erra-

tum [BGW^{+}16].

Bakker:2016:EPR

[BGW^{+}16] Arthur Bakker, Djonie Groenveld, Monica Wijers, Sanne F. Akkerman, and Koeno P. E. Gravemeijer. Erratum to: Proportional reasoning in the laboratory: an intervention study in vocational education. Educational Studies in Mathematics, 93 (3):399, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-016-9723-6.pdf. See

[BGW^{+}14].

Bitter:1994:TEM

[BH94] Gary G. Bitter and Mary M. Hatfield. Training elementary mathematics teachers using interactive multimedia. Educational Studies in Mathematics, 26(4):405–409, June 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01279523.

Berry:1995:SUP

[BH95] John Berry and Ken Houston. Students using posters as a means of communication and assessment. Educational Studies in Mathematics, 29(1):21–27, July 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273898.

Bakker:2005:DRB

[BH05] Arthur Bakker and Michael H. G. Hoffmann. Diagrammatic reasoning as the basis for developing concepts: A semiotic analysis of students’ learning about statistical distribution. Educational Studies in Mathematics, 60(3):333–358, November 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-5536-8.

Brown:2011:GSS

[BH11] Tony Brown and David Heywood. Geometry, subjectivity and the seduction of language: the regulation of spatial perception. Educational Studies in Mathematics, 77(2–3):351–367, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9281-2.

Barmby:2009:ARP

[BHHS09] Patrick Barmby, Tony Harries, Steve Higgins, and Jennifer Suggate. The array representation and primary children’s understanding and reasoning in multiplication. Educational Studies in Mathematics, 70(3):217–241, April 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9145-1.

Ben-Haim:1985:VRS

[BHLH85] D. Ben-Haim, G. Lappan, and R. T. Houang. Visualizing rectangular solids made of small cubes: Analyzing and effecting students’ performance. Educational Studies in Mathematics, 16(4):389–409, November 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00417194.

Bolden:2010:PSP

[BHN10] David S. Bolden, Tony V. Harries, and Douglas P. Newton. Pre-service primary teachers’ conceptions of creativity in mathematics. Educational Studies in Mathematics, 73(2): 143–157, March 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9207-z.

Byrnes:1997:GDM

[BHX97] James P. Byrnes, Li Hong, and Shaoying Xing. Gender differences on the math subtest of the scholastic aptitude test may be culture-specific. Educational Studies in Mathematics, 34 (1):49–66, October 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003011300372.

Blanton:2019:GCU

[BIBS^{+}19] Maria Blanton, Isil Isler-Baykal, Rena Stroud, Ana Stephens, Eric Knuth, and Angela Murphy Gardiner. Growth in children’s understanding of generalizing and representing mathematical structure and relationships. Educational Studies in Mathematics, 102(2):193–219, October 2019. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-019-09894-7.

Bicco:1982:PAN

[Bic82] Graciela Bicco. Les premières acquisitions de la notion de fonction linéaire chez l’enfant de 7 à 11 ans. (French) [The first acquisitions of the notion of linear function in the child of 7 to 11 years]. Educational Studies in Mathematics, 13(3): 289–327, August 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311246.

Biggs:1970:CPE

[Big70] E. E. Biggs. Communication on primary education in mathematics. Educational Studies in Mathematics, 2(4):469–475, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00417628.

Bishop:1976:DMI

[Bis76] Alan J. Bishop. Decision-making, the intervening variable. Educational Studies in Mathematics, 7(1–2):41–47, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144357.

Bishop:1979:ES

[Bis79a] Alan Bishop. Editorial statement. Educational Studies in Mathematics, 10(1):1, February 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311172.

Bishop:1979:VMP

[Bis79b] Alan J. Bishop. Visualising and mathematics in a pretechnological culture. Educational Studies in Mathematics, 10 (2):135–146, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00230984.

Bishop:1980:SAM

[Bis80] Alan J. Bishop. Spatial abilities and mathematics education — a review. Educational Studies in Mathematics, 11(3):257–269, August 1980. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00697739.

Bishop:1985:ES

[Bis85] Alan Bishop. Editorial statement. Educational Studies in Mathematics, 16(4):335, November 1985. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417190.

Bishop:1988:Ea

[Bis88a] Alan J. Bishop. Editorial. Educational Studies in Mathematics, 19(3):287, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312447.

Bishop:1988:Eb

[Bis88b] Alan J. Bishop. Editorial. Educational Studies in Mathematics, 19(4):419, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00578690.

Bishop:1988:MEC

[Bis88c] Alan J. Bishop. Mathematics education in its cultural context. Educational Studies in Mathematics, 19(2):179–191, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00751231.

Bishop:1989:E

[Bis89] Alan J. Bishop. Editorial. Educational Studies in Mathematics, 20(3):345, August 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00310877.

Bednarz:1982:UNP

[BJ82] Nadine Bednarz and Bernadette Janvier. The understanding of numeration in primary school. Educational Studies in Mathematics, 13(1):33–57, February 1982. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305497.

Bednarz:1988:CAN

[BJ88] Nadine Bednarz and Bernadette Janvier. A constructivist approach to numeration in primary school: Results of a three year intervention with the same group of children. Educational Studies in Mathematics, 19(3):299–331, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312450.

Brady:2022:MPT

[BJ22] Corey Brady and Hyunyi Jung. Modeling presentations: toward an assessment of emerging classroom cultures of modeling. Educational Studies in Mathematics, 109(2):237–261, February 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10056-x.

Bjarnadottir:2020:BRA

[Bja20a] Kristín Bjarnadóttir. Book review: Approaching the core of the modern mathematics movement. Dirk De Bock and Geert Vanpaemel (Eds.) (2019) Rods, sets and arrows — the rise and fall of modern mathematics in Belgium. Educational Studies in Mathematics, 104(3):431–438, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09961-4.

Bjarnadottir:2020:CBR

[Bja20b] Kristín Bjarnadóttir. Correction to: Book Review: Approaching the core of the modern mathematics movement. Dirk De Bock and Geert Vanpaemel (2019) Rods, sets and arrows — the rise and fall of modern mathematics in Belgium. Educational Studies in Mathematics, 105(2):285, October 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09985-w.

Bouck:2013:ECU

[BJJ13] Emily C. Bouck, Gauri S. Joshi, and Linley Johnson. Examining calculator use among students with and without disabilities educated with different mathematical curricula. Educational Studies in Mathematics, 83(3):369–385, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9461-3.

Bjuland:2004:STR

[Bju04] Raymond Bjuland. Student teachers’ reflections on their learning process through collaborative problem solving in geometry. Educational Studies in Mathematics, 55(1–3):199–225, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017690.90763.c1.

Blum:1991:PPE

[BK91] Werner Blum and Arnold Kirsch. Preformal proving: Examples and reflections. Educational Studies in Mathematics, 22 (2):183–203, April 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00555722.

Bergman:2024:DAM

[BKGZ24] Anna Marie Bergman, Andrew Kercher, Keith Gallagher, and Rina Zazkis. Definitional ambiguity in mathematics: three cases. Educational Studies in Mathematics, 115(1):93–110, January 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10241-0.

Bozkurt:2025:DPS

[BKH25] Gülay Bozkurt, Melike Yiğit Koyunkaya, and Mariam Haspekian. Developing pre-service mathematics teachers’ didactic reference points for the use of dynamic technology through micro-teaching. Educational Studies in Mathematics, 120(3):429–454, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-025-10429-
6.

Bastian:2022:TNG

[BKK22] Anton Bastian, Gabriele Kaiser, and Johannes König. Teacher noticing and its growth toward expertise: an expert–novice comparison with pre-service and in-service secondary mathematics teachers. Educational Studies in Mathematics, 110 (2):205–232, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10128-y.

Lagrange:2000:LII

[bL00] Jean baptiste Lagrange. L’intégration d’instruments informatiques dans l’enseignement: Une approche par les techniques. (French) [Integrating IT tools into teaching: An approach through technology]. Educational Studies in Mathematics, 43 (1):1–30, July 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1012086721534.

Baroody:2022:DAC

[BL22] Arthur J. Baroody and Menglung Lai. The development and assessment of counting-based cardinal number concepts. Educational Studies in Mathematics, 111(2):185–205, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10153-5.

Blakers:1978:CME

[Bla78] A. L. Blakers. Change in mathematics education since the late 1950’s — ideas and realisation Australia. Educational Studies in Mathematics, 9(2):147–158, May 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221155.

Blane:1985:LSC

[Bla85] Dudley Blane. A longitudinal study of children’s school mobility and attainment in mathematics. Educational Studies in Mathematics, 16(2):127–142, May 1985. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02400934.

Bicer:2020:CMC

[BLC20] Ali Bicer, Yujin Lee, and Robert M. Capraro. Considering mathematical creative self-efficacy with problem posing as a measure of mathematical creativity. Educational Studies in Mathematics, 105(3):457–485, November 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09995-8.

Bloch:2003:TFG

[Blo03] I. Bloch. Teaching functions in a graphic milieu: What forms of knowledge enable students to conjecture and prove? Educational Studies in Mathematics, 52(1):3–28, January 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023696731950.

Boesen:2010:RBT

[BLP10] Jesper Boesen, Johan Lithner, and Torulf Palm. The relation between types of assessment tasks and the mathematical reasoning students use. Educational Studies in Mathematics, 75(1):89–105, September 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9242-9.

Bishop:2014:UOR

[BLP^{+}14] Jessica Pierson Bishop, Lisa L. Lamb, Randolph A. Philipp, Ian Whitacre, and Bonnie P. Schappelle. Using order to reason about negative numbers: the case of Violet. Educational Studies in Mathematics, 86(1):39–59, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9519-x.

Blum:2002:ISA

[Blu02] Werner Blum. ICMI Study 14: Applications and modelling in mathematics education — discussion document. Educational Studies in Mathematics, 51(1–2):149–171, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022435827400.

Bingolbali:2008:CIR

[BM08] Erhan Bingolbali and John Monaghan. Concept image revisited. Educational Studies in Mathematics, 68(1):19–35, May 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9112-2.

Bjorklund:2021:WLC

[BMK21] Camilla Björklund, Ference Marton, and Angelika Kullberg. What is to be learnt? Critical aspects of elementary arithmetic skills. Educational Studies in Mathematics, 107(2):261– 284, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10045-0.

Bernabeu:2021:PSS

[BML21] Melania Bernabeu, Mar Moreno, and Salvador Llinares. Primary school students’ understanding of polygons and the relationships between polygons. Educational Studies in Mathematics, 106(2):251–270, February 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-020-10012-

Borji:2020:SUI

[BMP20] Vahid Borji and Rafael Martínez-Planell. On students’ understanding of implicit differentiation based on APOS theory. Educational Studies in Mathematics, 105(2):163–179, October 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09991-y.

Borji:2024:SGU

[BMPT24] Vahid Borji, Rafael Martínez-Planell, and María Trigueros. Students’ geometric understanding of partial derivatives and the locally linear approach. Educational Studies in Mathematics, 115(1):69–91, January 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-023-10242-
z.

Batteau:2025:CPS

[BMR25] Valérie Batteau, Takeshi Miyakawa, and Minbom Ryu. Collective problem-solving in Japanese primary mathematics lessons. Educational Studies in Mathematics, 119(3):421–443, July 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10400-5.

Yusof:1998:CAU

[bMYT98] Yudariah bt. Mohammad Yusof and David Tall. Changing attitudes to university mathematics through problem solving. Educational Studies in Mathematics, 37(1):67–82, October 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003456104875.

Bognar:1977:TPS

[BN77] K. Bognár and T. Nemetz. On the teaching of probability at secondary level. Educational Studies in Mathematics, 8(4): 399–404, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00310944.

Blum:1991:AMP

[BN91] Werner Blum and Mogens Niss. Applied mathematical problem solving, modelling, applications, and links to other subjects — state, trends and issues in mathematics instruction. Educational Studies in Mathematics, 22(1):37–68, February 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302716.

Baturo:1996:STS

[BN96] Annette Baturo and Rod Nason. Student teachers’ subject matter knowledge within the domain of area measurement. Educational Studies in Mathematics, 31(3):235–268, October 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00376322.

Batanero:1997:EIC

[BNPG97] Carmen Batanero, Virginia Navarro-Pelayo, and Juan D. Godino. Effect of the implicit combinatorial model on combinatorial reasoning in secondary school pupils. Educational Studies in Mathematics, 32(2):181–199, February 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002954428327.

Boaler:1993:ETS

[Boa93] Jo Boaler. Encouraging the transfer of ‘school’ mathematics to the ‘real world’ through the integration of process and content, context and culture. Educational Studies in Mathematics, 25(4):341–373, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273906.

Boaler:1999:PKB

[Boa99] Jo Boaler. Participation, knowledge and beliefs: A community perspective on mathematics learning. Educational Studies in Mathematics, 40(3):259–281, November 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003880012282.

Boaler:2002:ENM

[Boa02] Jo Boaler. Exploring the nature of mathematical activity: using theory, research and ‘working hypotheses’ to broaden conceptions of mathematics knowing. Educational Studies in Mathematics, 51(1–2):3–21, July 2002.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022468022549.

Brunner:2024:HPC

[BOL24] Kirsten Brunner, Andreas Obersteiner, and Timo Leuders. How pedagogical content knowledge sharpens prospective teachers’ focus when judging mathematical tasks: an eyetracking study. Educational Studies in Mathematics, 115(2): 177–196, February 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10281-6.

Bonotto:2013:ASP

[Bon13] Cinzia Bonotto. Artifacts as sources for problem-posing activities. Educational Studies in Mathematics, 83(1):37–55, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9441-7.

Bonner:2014:IPH

[Bon14] Emily P. Bonner. Investigating practices of highly successful mathematics teachers of traditionally underserved students. Educational Studies in Mathematics, 86(3):377–399, July 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9533-7.

Booth:1981:CMS

[Boo81] Lesley R. Booth. Child-methods in secondary mathematics. Educational Studies in Mathematics, 12(1):29–41, February 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386044.

Borasi:1986:NP

[Bor86] Raffaělla Borasi. On the nature of problems. Educational Studies in Mathematics, 17(2):125–141, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311517.

Borba:2021:FME

[Bor21] Marcelo C. Borba. The future of mathematics education since COVID-19: humans-with-media or humans-with-non-living-things. Educational Studies in Mathematics, 108(1–2): 385–400, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10043-2.

Boschet:1987:FDC

[Bos87] Françoise Boschet. Fonctions du code symbolique dans le discours mathématique. (French) [Functions of symbolic code in mathematical discourse]. Educational Studies in Mathematics, 18(1):19–34, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00367912.

Bos:1993:BRC

[Bos93] Henk J. M. Bos. ‘The bond with reality is cut’ — Freudenthal on the foundations of geometry around 1900. Educational Studies in Mathematics, 25(1–2):51–58, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274101.

Boylan:2016:EDM

[Boy16] Mark Boylan. Ethical dimensions of mathematics education. Educational Studies in Mathematics, 92(3):395–409, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9678-z.

Brock:1980:SPN

[BP80] William H. Brock and Michael H. Price. Squared paper in the nineteenth century: Instrument of science and engineering, and symbol of reform in mathematical education. Educational Studies in Mathematics, 11(4):365–381, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00231211.

Braating:2015:RBH

[BP15] Kajsa Bråting and Johanna Pejlare. On the relations between historical epistemology and students’ conceptual developments in mathematics. Educational Studies in Mathematics, 89(2): 251–265, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9600-8.

Bjorklund:2022:TTM

[BP22] Camilla Björklund and Hanna Palmér. Teaching toddlers the meaning of numbers — connecting modes of mathematical representations in book reading. Educational Studies in Mathematics, 110(3):525–544, July 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10147-3.

Brown:1978:OTC

[BR78] Stephen I. Brown and Gerald R. Rising. One third Cherokee: Problem solving, teaching, and intuition. Educational Studies in Mathematics, 9(4):391–409, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410585.

Borasi:1989:JWM

[BR89] Raffaella Borasi and Barbara J. Rose. Journal writing and mathematics instruction. Educational Studies in Mathematics, 20(4):347–365, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00315606.

Brown:2006:ECS

[BR06] Laurinda Brown and David A. Reid. Embodied cognition: Somatic markers, purposes and emotional orientations. Educational Studies in Mathematics, 63(2):179–192, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9027-3.

Bautista:2012:CSF

[BR12] Alfredo Bautista and Wolff-Michael Roth. Conceptualizing sound as a form of incarnate mathematical consciousness. Educational Studies in Mathematics, 79(1):41–59, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9337-y.

Bozkurt:2017:CBP

[BR17] Gulay Bozkurt and Kenneth Ruthven. Classroom-based professional expertise: a mathematics teacher’s practice with technology. Educational Studies in Mathematics, 94(3):309– 328, March 2017. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-016-9732-5.pdf.

Baumanns:2022:PPP

[BR22] Lukas Baumanns and Benjamin Rott. The process of problem posing: development of a descriptive phase model of problem posing. Educational Studies in Mathematics, 110(2):251– 269, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10136-y.

Brailly:1968:EEL

[Bra68] A. Brailly. Un exemple d’exploitation d’une ‘situation’. (French) [An example of exploitation of a ‘situation’]. Educational Studies in Mathematics, 1(1–2):237–243, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426246;   http://link.springer.
com/article/10.1007/BF00426246.

Brantlinger:2011:RCM

[Bra11] Andrew Brantlinger. Rethinking critical mathematics: a comparative analysis of critical, reform, and traditional geometry instructional texts. Educational Studies in Mathematics, 78(3):395–411, December 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9331-4.

Brandt:2013:EPP

[Bra13] Birgit Brandt. Everyday pedagogical practices in mathematical play situations in German “Kindergarten”. Educational Studies in Mathematics, 84(2):227–248, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9490-6.

Brantlinger:2014:CMD

[Bra14] Andrew Brantlinger. Critical mathematics discourse in a high school classroom: examining patterns of student engagement and resistance. Educational Studies in Mathematics, 85(2): 201–220, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9506-2.

Brenner:1998:MM

[Bre98] Mary E. Brenner. Meaning and money. Educational Studies in Mathematics, 36(2):123–155, July 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003176619818.

Bagger:2020:DIS

[BRE20] Anette Bagger, Helena Roos, and Margareta Engvall. Directions of intentionalities in special needs education in mathematics. Educational Studies in Mathematics, 104(1):41–63, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09945-4;   http://link.
springer.com/content/pdf/10.1007/s10649-020-09945-
4.pdf.

Breive:2022:AEE

[Bre22] Svanhild Breive. Abstraction and embodiment: exploring the process of grasping a general. Educational Studies in Mathematics, 110(2):313–329, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10137-
x.

Brizuela:2006:YCN

[Bri06] Bárbara M. Brizuela. Young children’s notations for fractions. Educational Studies in Mathematics, 62(3):281–305, July 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-9003-3.

Brown:1991:HMA

[Bro91] Tony Brown. Hermeneutics and mathematical activity. Educational Studies in Mathematics, 22(5):475–480, October 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367909.

Brown:1994:CKM

[Bro94] Tony Brown. Creating and knowing mathematics through language and experience. Educational Studies in Mathematics, 27 (1):79–100, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01284529.

Brown:1996:PMC

[Bro96] Tony Brown. The phenomenology of the mathematics classroom. Educational Studies in Mathematics, 31(1–2):115–150, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00143929.

Brown:2008:LST

[Bro08a] Tony Brown. Lacan, subjectivity and the task of mathematics education research. Educational Studies in Mathematics, 68 (3):227–245, July 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9111-3.

Brown:2008:SST

[Bro08b] Tony Brown. Signifying “students”, “teachers” and “mathematics”: a reading of a special issue. Educational Studies in Mathematics, 69(3):249–263, November 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9130-8.

Brown:2010:DIG

[Bro10a] Roger G. Brown. Does the introduction of the graphics calculator into system-wide examinations lead to change in the types of mathematical skills tested? Educational Studies in Mathematics, 73(2):181–203, March 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9220-2.

Brown:2010:TRK

[Bro10b] Tony Brown. Truth and the renewal of knowledge: the case of mathematics education. Educational Studies in Mathematics, 75(3):329–343, December 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9259-0.

Brodie:2011:BRM

[Bro11] Karin Brodie. Book review: Mathematical relationships in education: identities and participation, edited by Laura Black, Heather Mendick, & Yvette Solomon (2009). Educational Studies in Mathematics, 76(2):237–241, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9270-5.

Brown:2012:SCA

[Bro12] Tony Brown. Subjectivity and cultural adjustment in mathematics education: a response to Wolff–Michael Roth. Educational Studies in Mathematics, 80(3):475–490, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9400-3.

Brodie:2014:LAL

[Bro14a] Karin Brodie. Learning about learner errors in professional learning communities. Educational Studies in Mathematics, 85(2):221–239, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9507-1.

Brown:2014:SRD

[Bro14b] Stacy A. Brown. On skepticism and its role in the development of proof in the classroom. Educational Studies in Mathematics, 86(3):311–335, July 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9544-4.

Brown:2016:RBL

[Bro16] Tony Brown. Rationality and belief in learning mathematics. Educational Studies in Mathematics, 92(1):75–90, May 2016.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-015-9670-7.pdf.

Brown:2023:BRR

[Bro23] Tony Brown. Book review response. Who wants to know? Author’s response to review of A Contemporary Theory of Mathematics Education Research. Educational Studies in Mathematics, 113(2):365–367, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-023-10221-

4. See [NX23].

Bender:1980:POC

[BS80] Peter Bender and Alfred Schreiber. The principle of operative concept formation in geometry teaching. Educational Studies in Mathematics, 11(1):59–90, February 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00369160.

Bromme:1994:IDS

[BS94] Rainer Bromme and Heinz Steinbring. Interactive development of subject matter in the mathematics classroom. Educational Studies in Mathematics, 27(3):217–248, October 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273730.

Banerjee:2012:ETA

[BS12] Rakhi Banerjee and K. Subramaniam. Evolution of a teaching approach for beginning algebra. Educational Studies in Mathematics, 80(3):351–367, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9353-y.

Braathe:2015:CMN

[BS15] Hans Jørgen Braathe and Yvette Solomon. Choosing mathematics: the narrative of the self as a site of agency. Educational Studies in Mathematics, 89(2):151–166, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9585-8.

Banerjee:2025:MRE

[BS25] Rakhi Banerjee and Kaye Stacey. Mathematical reasoning in elementary grade classrooms in India: what is there and what is missing? Educational Studies in Mathematics, 118(1):5–27, January 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10362-0.

Brown:1996:ATS

[BSF96] Catherine A. Brown, Mary Kay Stein, and Ellice Ann Forman. Assisting teachers and students to reform the mathematics classroom. Educational Studies in Mathematics, 31(1–2):63– 93, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00143927.

Bell:1981:COV

[BST81] Alan Bell, Malcolm Swan, and Glenda Taylor. Choice of operation in verbal problems with decimal numbers. Educational Studies in Mathematics, 12(4):399–420, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00308139.

Boekaerts:1995:SMP

[BSV95] Monique Boekaerts, Gerard Seegers, and Harriet Vermeer. Solving math problems: Where and why does the solution process go astray? Educational Studies in Mathematics, 28 (3):241–262, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274175.

Brown:2016:TME

[BSW16] Tony Brown, Yvette Solomon, and Julian Williams. Theory in and for mathematics education: in pursuit of a critical agenda. Educational Studies in Mathematics, 92(3):287–297, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-016-9706-7.pdf.

Bruce:2004:OTT

[BT04] R. A. Bruce and J. Threlfall. One, two, three and counting. Educational Studies in Mathematics, 55(1–3):3–26, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017676.79430.dc.

Barbeau:2005:ISC

[BT05] Edward J. Barbeau and Peter J. Taylor. ICMI Study 16: Challenging mathematics in and beyond the classroom: Discussion document. Educational Studies in Mathematics, 60(1):125– 139, September 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-8481-7.

Batchelor:2019:AMY

[BTV19] Sophie Batchelor, Joke Torbeyns, and Lieven Verschaffel. Affect and mathematics in young children: an introduction. Educational Studies in Mathematics, 100(3):201–209, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9864-x.

Bunck:2017:OAC

[BTvG^{+}17] M. J. A. Bunck, E. Terlien, M. van Groenestijn, S. W. M. Toll, and J. E. H. Van Luit. Observing and analyzing children’s mathematical development, based on action theory. Educational Studies in Mathematics, 96(3):289–304, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9763-6.

Buchanan:1987:FCM

[Buc87] Nina Kay Buchanan. Factors contributing to mathematical problem-solving performance: An exploratory study. Educational Studies in Mathematics, 18(4):399–415, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00240987.

Buisson:1976:ESP

[Bui76] Pierre Buisson. Évaluation-sondage dans le premier cycle (12– 16 ans). (French) [Assessment survey in the first cycle (12–16 years)]. Educational Studies in Mathematics, 6(4):429–445, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00411091.

Bullock:2024:RDA

[Bul24] Erika C. Bullock. Racialized deviance as an axiom in the mathematics education equity genre. Educational Studies in Mathematics, 116(3):333–350, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-023-10260-
x.

Burghes:1980:MMP

[Bur80a] David N. Burghes. Mathematical modelling: A positive direction for the teaching of applications of mathematics at school. Educational Studies in Mathematics, 11(1):113–131, February 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00369162.

Burton:1980:TMY

[Bur80b] Leone Burton. The teaching of mathematics to young children using a problem solving approach. Educational Studies in Mathematics, 11(1):43–58, February 1980. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00369159.

Burton:1987:FSC

[Bur87] Leone Burton. From failure to success: Changing the experience of adult learners of mathematics. Educational Studies in Mathematics, 18(3):305–316, August 1987. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386200.

Burt:1990:DCC

[Bur90] Gordon Burt. Doing critical cultural and ideological technology. Educational Studies in Mathematics, 21(3):289–298, June 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305094.

Burton:1994:GE

[Bur94] Leone Burton. Guest editorial. Educational Studies in Mathematics, 27(4):317–319, December 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273375.

Burton:1995:MTF

[Bur95] Leone Burton. Moving towards a feminist epistemology of mathematics. Educational Studies in Mathematics, 28(3):275– 291, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01274177.

Burn:1996:WFC

[Bur96] Bob Burn. What are the fundamental concepts of group theory? Educational Studies in Mathematics, 31(4):371– 377, December 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00369154. See reac-

tion [DDLZ97].

Burton:1998:PMW

[Bur98] Leone Burton. The practices of mathematicians: What do they tell us about coming to know mathematics? Educational Studies in Mathematics, 37(2):121–143, November 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003697329618.

Burn:2002:SCR

[Bur02a] R. P. Burn. Some comments on the “Role of proof in comprehending and teaching elementary linear algebra’ by F. Uhlig. Educational Studies in Mathematics, 51(3):183–184, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1023652816270.  See

[Uhl02, Uhl03].

Burton:2002:RCR

[Bur02b] Leone Burton. Recognising commonalities and reconciling differences in mathematics education. Educational Studies in Mathematics, 50(2):157–175, March 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021158009837.

Burn:2003:BR

[Bur03] B. Burn. Book review. Educational Studies in Mathematics, 52 (2):211–214, March 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1024013009990.

Burn:2005:VSH

[Bur05] Bob Burn. The vice: Some historically inspired and proof-generated steps to limits of sequences. Educational Studies in Mathematics, 60(3):269–295, November 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-7923-6. See commentary

[Ber06] and reply [Bur06].

Burn:2006:FRV

[Bur06] R. P. Burn. Further reflections on “The Vice”. Educational Studies in Mathematics, 61(3):403–407, March 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-1924-y.

Bussi:1996:MDP

[Bus96] Maria G. Bartolini Bussi. Mathematical discussion and perspective drawing in primary school. Educational Studies in Mathematics, 31(1–2):11–41, September 1996. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00143925.

Bussi:2005:WCS

[Bus05] Maria G. Bartolini Bussi. When classroom situation is the unit of analysis: The potential impact on research in mathematics education. Educational Studies in Mathematics, 59(1– 3):299–311, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-5478-1.

Buscher:2024:AHC

[Büs24] Christian Büscher. Adapting Habermas’ construct of communicative rationality into a framework for analyzing students’ statistical literacy. Educational Studies in Mathematics, 117 (1):121–141, September 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10325-5.

Beccuti:2024:SDU

[BVR24] Francesco Beccuti, Paola Valero, and Ornella Robutti. Stories of devoted university students: the mathematical experience as a form of ascesis. Educational Studies in Mathematics, 115 (1):51–67, January 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10259-4.

Bills:2008:EI

[BW08] Liz Bills and Anne Watson. Editorial introduction. Educational Studies in Mathematics, 69(2):77–79, October 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9147-z.

Brown:2012:MEC

[BW12] Tony Brown and Margaret Walshaw. Mathematics education and contemporary theory: guest editorial. Educational Studies in Mathematics, 80(1–2):1–8, May 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-012-9397-7.pdf.

Black:2013:CCB

[BW13] Laura Black and Julian Williams. Contradiction and conflict between ‘leading identities’: becoming an engineer versus becoming a ‘good Muslim’ woman. Educational Studies in Mathematics, 84(1):1–14, September 2013. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-013-9481-7.pdf.

Bakker:2020:PLT

[BW20] Arthur Bakker and David Wagner. Pandemic: lessons for today and tomorrow? Educational Studies in Mathematics, 104(1):1–4, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09946-
3;   http://link.springer.com/content/pdf/10.1007/
s10649-020-09946-3.pdf.

Bakker:2021:E

[BW21] Arthur Bakker and David Wagner. Editorial. Educational Studies in Mathematics, 106(1):1–3, January 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10015-y.

Beswick:2010:CTP

[BWD10] Kim Beswick, Anne Watson, and Els De Geest. Comparing theoretical perspectives in describing mathematics departments: complexity and activity. Educational Studies in Mathematics, 75(2):153–170, November 2010. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9248-3.

Black:2010:DLI

[BWHM^{+}10] Laura Black, Julian Williams, Paul Hernandez-Martinez, Pauline Davis, Maria Pampaka, and Geoff Wake. Developing a ‘leading identity’: the relationship between students’ mathematical identities and their career and higher education aspirations. Educational Studies in Mathematics, 73(1): 55–72, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9217-x.

Byers:2015:HHL

[Bye15] William Byers. How humans learn to think mathematically: Exploring the three worlds of mathematics, by David tall. (2013). Educational Studies in Mathematics, 88(1):159–162, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9579-6.

Byun:2024:CIP

[Byu24a] Sunghwan Byun. Correction to: Interactional practices of inviting minoritized students to whole-class mathematics discussions. Educational Studies in Mathematics, 115(3):351, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10305-9. See [Byu24b].

Byun:2024:IPI

[Byu24b] Sunghwan Byun. Interactional practices of inviting minoritized students to whole-class mathematics discussions. Educational Studies in Mathematics, 115(3):321–350, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10292-3. See correction [Byu24a].

Benbachir:2001:PEC

[BZ01] Amina Benbachir and Moncef Zaki. Production d’exemples et de contre-exemples en analyse: étude de cas en première d’université. (French) [Production of examples and counterexamples in analysis: first-year university case study]. Educational Studies in Mathematics, 47(3):273–295, September 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015198520576.

Ben-Zvi:2001:JHS

[BZA01] Dani Ben-Zvi and Abraham Arcavi. Junior high school students’ construction of global views of data and data representations. Educational Studies in Mathematics, 45(1–3):35–65, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1013809201228.

Ben-Zvi:2015:LRS

[BZBM15] Dani Ben-Zvi, Arthur Bakker, and Katie Makar. Learning to reason from samples. Educational Studies in Mathematics, 88 (3):291–303, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-015-9593-3.pdf.

Cakir:2020:ESM

[ÇA20] Aslı Çakır and Hatice Akkoç. Examining socio-mathematical norms related to problem posing: a case of a gifted and talented mathematics classroom. Educational Studies in Mathematics, 105(1):19–34, September 2020. CODEN EDSMAN.

ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09965-
0.

Callejo:1994:RGR

[Cal94] María Luz Callejo. Les représentations graphiques dans la résolution de problèmes: Une expérience d’entraînement d’étudiants dans un club mathématique. (French) [Graphical representations in problem solving: An experience of student engagement in a mathematical club]. Educational Studies in Mathematics, 27(1):1–33, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01284526.

Calder:2012:LMI

[Cal12] Nigel Calder. The layering of mathematical interpretations through digital media. Educational Studies in Mathematics, 80 (1–2):269–285, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9365-7.

Carraher:1993:LTR

[Car93] David William Carraher. Lines of thought: A ratio and operator model of rational number. Educational Studies in Mathematics, 25(4):281–305, December 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273903.

Carlson:1999:MBS

[Car99] Marilyn P. Carlson. The mathematical behavior of six successful mathematics graduate students: Influences leading to mathematical success. Educational Studies in Mathematics, 40(3):237–258, November 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003819513961.

Carraher:2008:BBV

[Car08] David W. Carraher. Beyond ‘blaming the victim’ and ‘standing in awe of noble savages’: a response to “Revisiting Lave’s ‘cognition in practice”’. Educational Studies in Mathematics, 69(1):23–32, September 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9126-4.

Carlsen:2010:AGS

[Car10] Martin Carlsen. Appropriating geometric series as a cultural tool: a study of student collaborative learning. Educational Studies in Mathematics, 74(2):95–116, June 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9230-0.

Carlsen:2018:USS

[Car18] Martin Carlsen. Upper secondary students’ mathematical reasoning on a sinusoidal function. Educational Studies in Mathematics, 99(3):277–291, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9844-
1.

Castelnuovo:1969:DRU

[Cas69a] Emma Castelnuovo. Différentes représentations utilisant la notion de barycentre. (French) [Different representations using the notion of barycenter]. Educational Studies in Mathematics, 2(2–3):307–332, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00303466.

Castelnuovo:1969:TAC

[Cas69b] Emma Castelnuovo. Les transformations affines dans le 1er cycle de l’école secondaire. (French) [Affine transformations in the first cycle of secondary school]. Educational Studies in Mathematics, 1(3):274–288, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00558313.

Castelnuovo:1969:LCA

[Cas69c]  Emma Castelnuovo.  L’intégration centrée autour des

mathématiques du 1er cycle de l’enseignement secondaire. (French) [Integration centered on the mathematics of the first cycle of secondary education]. Educational Studies in Mathematics, 1(3):340–342, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00558319.

Castelnuovo:1977:LMF

[Cas77] Emma Castelnuovo. L’enseignement des mathématiques. (French) [The teaching of mathematics]. Educational Studies in Mathematics, 8(1):41–50, April 1977. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302504.

Castro:1998:TPC

[Cas98] César Sáenz Castro. Teaching probability for conceptual change: La enseñanza de la probabilidad por cambio conceptual. Educational Studies in Mathematics, 35(3):233–254, March 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003182219483.

Castela:2004:IIM

[Cas04] Corine Castela. Institutions influencing mathematics students’ private work: A factor of academic achievement. Educational Studies in Mathematics, 57(1):33–63, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000047050.70008.59.

Capponi:1989:TCA

[CB89] B. Capponi and N. Balacheff. Tableur et calcul algébrique. (French) [Spreadsheet and algebraic computation]. Educational Studies in Mathematics, 20(2):179–210, May 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00579462.

Clements:2000:SDS

[CB00] Douglas H. Clements and Barbara A. Burns. Students’ development of strategies for turn and angle measure. Educational Studies in Mathematics, 41(1):31–45, January 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003938415559.

Carlson:2005:CNP

[CB05] Marilyn P. Carlson and Irene Bloom. The cyclic nature of problem solving: An emergent multidimensional problemsolving framework. Educational Studies in Mathematics, 58 (1):45–75, January 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-0808-x.

Corriveau:2017:EST

[CB17a] Claudia Corriveau and Nadine Bednarz. Erratum to: The secondary–tertiary transition viewed as a change in mathematical cultures: an exploration concerning symbolism and its use. Educational Studies in Mathematics, 95(3):351, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-017-9766-3.pdf.

Corriveau:2017:STT

[CB17b] Claudia Corriveau and Nadine Bednarz. The secondary– tertiary transition viewed as a change in mathematical cultures: an exploration concerning symbolism and its use. Educational Studies in Mathematics, 95(1):1–19, May 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). See erratum [CB17a].

Cabral:2019:CSS

[CB19] Tânia C. B. Cabral and Roberto Ribeiro Baldino. The credit system and the summative assessment splitting moment. Educational Studies in Mathematics, 102(2):275–288, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09907-5.

Caputo:2025:ISS

[CBC25] Sara Gagliani Caputo, Laura Branchetti, and Annalisa Cusi. Intertwining students’ social modes of co-construction and epistemic aspects of algebraic thinking in asynchronous mathematical discussions. Educational Studies in Mathematics, 120 (1):109–135, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-025-10395-z.

Cao:2007:PPI

[CBF07] Zhongjun Cao, Alan Bishop, and Helen Forgasz. Perceived parental influence on mathematics learning: A comparison among students in China and Australia. Educational Studies in Mathematics, 64(1):85–106, January 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-006-9033-5.pdf.

Courey:2012:AMM

[CBSP12] Susan Joan Courey, Endre Balogh, Jody Rebecca Siker, and Jae Paik. Academic music: music instruction to engage third-grade students in learning basic fraction concepts. Educational Studies in Mathematics, 81(2):251–278, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9395-9.

Clements:1996:DTT

[CBSS96] Douglas H. Clements, Michael T. Battista, Julie Sarama, and Sudha Swaminathan. Development of turn and turn measurement concepts in a computer-based instructional unit. Educational Studies in Mathematics, 30(4):313–337, June 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00570828.

Christiansen:2023:HBR

[CCP23] Iben Maj Christiansen, Claudia Corriveau, and Kerstin Pettersson. Hybrids between rituals and explorative routines: opportunities to learn through guided and recreated exploration. Educational Studies in Mathematics, 112(1):49–72, January 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10167-z.

Castro:2022:DSC

[CCRD22] Encarnación Castro, María C. Cañadas, and Susana Rodríguez-Domingo. Difficulties in semantically congruent translation of verbally and symbolically represented algebraic statements. Educational Studies in Mathematics, 109(3):593–609, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10088-3.

Campbell:1995:VPD

[CCW95] K. Jennifer Campbell, Kevin F. Collis, and Jane M. Watson. Visual processing during mathematical problem solving. Educational Studies in Mathematics, 28(2):177–194, March 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01295792.

Cooper:2004:WCT

[Cd04] Paul Cooper and Ray d’Inverno. Those who can, teach: addressing the crisis in mathematics in UK schools and universities. Educational Studies in Mathematics, 56(2–3): 343–357, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040392.
42967.fb.

Castelnuovo:1976:GPE

[CDGG76] Emma Castelnuovo, Daniela, and Claudio Gori-Giorgi. La géométrie projective à l’école. (French) [Projection geometry at school]. Educational Studies in Mathematics, 7(4):443–463, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00452225.

Cai:2014:HDE

[CDW14] Jinfa Cai, Meixia Ding, and Tao Wang. How do exemplary Chinese and U.S. mathematics teachers view instructional coherence? Educational Studies in Mathematics, 85(2):265–280, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9513-3.

Carlsen:2016:KTO

[CEHM16] Martin Carlsen, Ingvald Erfjord, Per Sigurd Hundeland, and John Monaghan. Kindergarten teachers’ orchestration of mathematical activities afforded by technology: agency and mediation. Educational Studies in Mathematics, 93(1):1–17, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-016-9692-9.

Christou:2004:TCR

[CEMP04] Constantinos Christou, Maria Eliophotou-Menon, and George Philippou. Teachers’ concerns regarding the adoption of a new mathematics curriculum: An application of CBAM. Educational Studies in Mathematics, 57(2):157–176, September 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000049271.01649.dd.

Cobo:2000:SIC

[CF00a] Pedro Cobo and Josep M. Fortuny. Social interactions and cognitive effects in contexts of area-comparison problem solving. Educational Studies in Mathematics, 42(2):115–140, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004187711956.

Conradie:2000:CTM

[CF00b] Jurie Conradie and John Frith. Comprehension tests in mathematics. Educational Studies in Mathematics, 42(3):225–235, May 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017502919000.

Cantoral:2003:MEV

[CF03] Ricardo Cantoral and Rosa M. Farfáan. Mathematics education: A vision of its evolution. Educational Studies in Mathematics, 53(3):255–270, September 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026008829822. See reaction [LQV04].

Copur-Gencturk:2021:TCU

[CG21] Yasemin Copur-Gencturk. Teachers’ conceptual understanding of fraction operations: results from a national sample of elementary school teachers. Educational Studies in Mathematics, 107(3):525–545, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10033-

Copur-Gencturk:2021:SCM

[CGD21] Yasemin Copur-Gencturk and Tenzin Doleck. Strategic competence for multistep fraction word problems: an overlooked aspect of mathematical knowledge for teaching. Educational Studies in Mathematics, 107(1):49–70, May 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10028-1.

Castelnuovo:1974:GFE

[CGGGG74] Emma Castelnuovo, Daniela Gori-Giorgi, and Claudio Gori-Giorgi. Les graphes de flux dans l’enseignement secondaire. (French) [Flow graphs in secondary education]. Educational Studies in Mathematics, 5(4):467–492, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01420659.

Castelnuovo:1979:CGU

[CGGGG79a] E. Castelnuovo, D. Gori-Giorgi, and C. Gori-Giorgi. Coniques et gravitation universelle. (French) [Conics and universal gravitation]. Educational Studies in Mathematics, 10(3):323–359, August 1979. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00314661.

Castelnuovo:1979:LPF

[CGGGG79b] Emma Castelnuovo, Daniela Gori-Giorgi, and Claudio Gori-Giorgi. Le lancement des projectiles. (French) [The launch of projectiles]. Educational Studies in Mathematics, 10(2): 147–159, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00230985.

Comiti:1976:ASC

[CGS76] Claude Comiti, Marie-Rose Guillermard, and Alain Solano. Approche des statistiques en classe de sixième: Analyse de résultats sportifs. (French) [Approach to statistics in sixth grade: Analysis of sporting results]. Educational Studies in Mathematics, 7(4):497–521, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00452227.

Cooper:2002:CRC

[CH02] Barry Cooper and Tony Harries. Children’s responses to contrasting ‘realistic’ mathematics problems: Just how realistic are children ready to be? Educational Studies in Mathematics, 49(1):1–23, January 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016013332659.

Chen:2013:IAG

[CH13] Chia-Ling Chen and Patricio Herbst. The interplay among gestures, discourse, and diagrams in students’ geometrical reasoning. Educational Studies in Mathematics, 83(2):285–307, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9454-2.

Chaye:1971:ADF

[Cha71] Jacques Chayé. Apprentissage de la déduction. (French) [Learning deduction]. Educational Studies in Mathematics, 4(2):225–231, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00303387.

Chazan:1993:HSG

[Cha93] Daniel Chazan. High school geometry students’ justification for their views of empirical evidence and mathematical proof. Educational Studies in Mathematics, 24(4):359–387, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273371.

Chapman:1997:MTM

[Cha97] Olive Chapman. Metaphors in the teaching of mathematical problem solving. Educational Studies in Mathematics, 32(3): 201–228, March 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1002991718392.

Chassapis:1998:MTD

[Cha98] Dimitris Chassapis. The mediation of tools in the development of formal mathematical concepts: The compass and the circle as an example. Educational Studies in Mathematics, 37(3): 275–293, December 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003696507155.

Chapman:2006:CPC

[Cha06] Olive Chapman. Classroom practices for context of mathematics word problems. Educational Studies in Mathematics, 62 (2):211–230, June 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-7834-1.

Chambris:2018:ITM

[Cha18] Christine Chambris. The influence of theoretical mathematical foundations on teaching and learning: a case study of whole numbers in elementary school. Educational Studies in Mathematics, 97(2):185–207, February 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9790-3.

Chapman:2022:WMP

[Cha22a] K. Chapman. “Wait — it’s a math problem, right?”: negotiating school frames in out-of-school places. Educational Studies in Mathematics, 109(3):661–676, March 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10099-0.

Chazan:2022:BRL

[Cha22b] Daniel I. Chazan. Book review: Luc Trouche, Ghislaine Gueudet, & Birgit Pepin (Eds.) (2019) The ‘resource’ approach to mathematics education. Educational Studies in Mathematics, 111(3):567–572, November 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10146-4.

Cheung:1988:OSM

[Che88] K. C. Cheung. Outcomes of schooling: Mathematics achievement and attitudes towards mathematics learning in Hong Kong. Educational Studies in Mathematics, 19(2):209–219, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00751233.

Cheung:1989:GDJ

[Che89] K. C. Cheung. Gender differences in the junior secondary (grade 7) mathematics curriculum in Hong Kong. Educational Studies in Mathematics, 20(1):97–103, February 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00356043.

Chevallard:1990:MEC

[Che90] Yves Chevallard. On mathematics education and culture: Critical afterthoughts. Educational Studies in Mathematics, 21 (1):3–27, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311013.

Chen:2023:TRA

[Che23] Grace A. Chen. Teacherly response-ability: ethical relationality as protest against mathematical violence. Educational Studies in Mathematics, 114(2):277–296, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10230-3.

Chinnappan:1998:SMM

[Chi98] Mohan Chinnappan. Schemas and mental models in geometry problem solving. Educational Studies in Mathematics, 36(3):201–217, September 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1003134323371.

Czocher:2022:BSA

[CHK22] Jennifer A. Czocher, Hamilton L. Hardison, and Sindura S. Kularajan. A bridging study analyzing mathematical model construction through a quantities-oriented lens. Educational Studies in Mathematics, 111(2):299–321, October 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10163-3.

Carpenter:1983:EIC

[CHM83] T. P. Carpenter, J. Hiebert, and J. M. Moser. The effect of instruction on children’s solutions of addition and subtraction word problems. Educational Studies in Mathematics, 14(1): 55–72, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00704702.

Cai:2016:MPP

[CHM16] Jinfa Cai, Stephen Hwang, and John C. Moyer. Mathematical problem posing as a measure of curricular effect on students’ learning: a response. Educational Studies in Mathematics, 91 (1):9–10, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9642-y.

Chowdhuri:2022:EME

[Cho22] Meghna Nag Chowdhuri. Exploring modes of engagement within reform-oriented primary mathematics textbooks in India. Educational Studies in Mathematics, 110(2):331–351, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10135-z.

Christiansen:1969:IDL

[Chr69] Bent Christiansen. Induction and deduction in the learning of mathematics and in mathematical instruction. Educational Studies in Mathematics, 2(2–3):139–159, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303454.

Christiansen:1997:WNM

[Chr97] Iben Maj Christiansen. When negotiation of meaning is also negotiation of task. Educational Studies in Mathematics, 34 (1):1–25, October 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1002944413332.

Chronaki:2000:TMT

[Chr00] A. Chronaki. Teaching maths through theme-based resources: Pedagogic style, ‘theme’ and ‘maths’ in lessons.

Educational Studies in Mathematics, 42(2):141–163, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004124328308.

Cater:1975:CLT

[CJ75] R. M. Cater and H. W. Johnstone, Jr. A computer language for teaching introductory logic. Educational Studies in Mathematics, 6(1):87–91, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00590026.

Cadez:2015:CTG

[ČK15] Tatjana Hodnik Čadež and Vida Manfreda Kolar. Comparison of types of generalizations and problem-solving schemas used to solve a mathematical problem. Educational Studies in Mathematics, 89(2):283–306, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9598-y.

Cadez:2018:HFG

[ČK18] Tatjana Hodnik Čadež and Vida Manfreda Kolar. How fifthgrade pupils reason about fractions: a reliance on part-whole subconstructs. Educational Studies in Mathematics, 99(3): 335–357, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9838-z.

Cevikbas:2022:SLR

[CKS22] Mustafa Cevikbas, Gabriele Kaiser, and Stanislaw Schukajlow. A systematic literature review of the current discussion on mathematical modelling competencies: state-of-the-art developments in conceptualizing, measuring, and fostering. Educational Studies in Mathematics, 109(2):205–236, February 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10104-6.

Clarkson:1984:CAS

[CL84] P. Clarkson and G. C. Leder. Causal attributions for success and failure in mathematics: A cross cultural perspective. Educational Studies in Mathematics, 15(4):413–422, November 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311115.

Cai:2018:CPE

[CL18] Jinfa Cai and Roza Leikin. Call for papers: Educational Studies in Mathematics special issue. Educational Studies in Mathematics, 99(3):243–244, November 2018. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-018-9858-8;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9858-8.pdf.

Cai:2020:AMP

[CL20] Jinfa Cai and Roza Leikin. Affect in mathematical problem posing: conceptualization, advances, and future directions for research. Educational Studies in Mathematics, 105(3):287–301, November 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-10008-x.

Chen:2023:DCN

[CL23] Zhihui Chen and Frank Okai Larbi. Deciphering the cognitive and non-cognitive competencies of Chinese students’ mathematics through a critical lens of PISA. Educational Studies in Mathematics, 113(3):519–528, July 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10210-7.

Clarkson:1983:TEM

[Cla83] Philip Clarkson. Types of errors made by Papua New Guinean students. Educational Studies in Mathematics, 14(4):355–367, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00368234.

Clarke:1985:ISS

[Cla85] David Clarke. The impact of secondary schooling and secondary mathematics on student mathematical behaviour. Educational Studies in Mathematics, 16(3):231–257, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00776732.

Clarkson:1992:LMC

[Cla92] Philip C. Clarkson. Language and mathematics: A comparison of bilingual and monolingual students of mathematics. Educational Studies in Mathematics, 23(4):417–429, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302443.

Clarkson:2002:TLF

[Cla02] P. S. Clarkson. Teaching/learning: Finding new ways to meaning. Educational Studies in Mathematics, 51(3):287–290, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1023608506239.

Clarkson:2007:AVS

[Cla07] Philip C. Clarkson. Australian Vietnamese students learning mathematics: High ability bilinguals and their use of their languages. Educational Studies in Mathematics, 64(2):191– 215, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-4696-5.

Clark:2012:HMI

[Cla12] Kathleen Michelle Clark. History of mathematics: illuminating understanding of school mathematics concepts for prospective mathematics teachers. Educational Studies in Mathematics, 81(1):67–84, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9361-y.

Clements:1979:SDM

[Cle79] M. A. (Ken) Clements. Sex differences in mathematical performance: An historical perspective. Educational Studies in Mathematics, 10(3):305–322, August 1979. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00314660.

Clements:1980:ACE

[Cle80] M. A. (Ken) Clements. Analyzing children’s errors on written mathematical tasks. Educational Studies in Mathematics, 11 (1):1–21, February 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00369157.

Clements:1984:TT

[Cle84] M. A. (Ken) Clements. Terence Tao. Educational Studies in Mathematics, 15(3):213–238, August 1984. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312075.

Clivaz:2017:TMM

[Cli17] Stéphane Clivaz. Teaching multidigit multiplication: combining multiple frameworks to analyse a class episode. Educational Studies in Mathematics, 96(3):305–325, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9770-7.

Cai:2022:EPS

[CLK22] Jinfa Cai, Raymond LaRochelle, and Gabriele Kaiser. Expert and preservice secondary teachers’ competencies for noticing student thinking about modelling. Educational Studies in Mathematics, 109(2):431–453, February 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10071-y.

Coutat:2016:LIP

[CLR16] Sylvia Coutat, Colette Laborde, and Philippe R. Richard. L’apprentissage instrumenté de propriétés en géométrie: propédeutique à l’acquisition d’une compétence de démonstration. (French) [The instrumented learning of properties in geometry: propedeutic to the acquisition of a competence in demonstration]. Educational Studies in Mathematics, 93(2):195–221, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-016-9684-9.

Cueto:2017:TPC

[CLSM17] Santiago Cueto, Juan León, M. Alejandra Sorto, and Alejandra Miranda. Teachers’ pedagogical content knowledge and mathematics achievement of students in Peru. Educational Studies in Mathematics, 94(3):329–345, March 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9735-2.

Cox:1992:DRH

[CM92] Caryl Cox and John T. Mouw. Disruption of the representativeness heuristic: Can we be perturbed into using correct probabilistic reasoning? Educational Studies in Mathematics, 23(2):163–178, April 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00588054.

Clivaz:2020:ECM

[CM20] Stéphane Clivaz and Takeshi Miyakawa. The effects of culture on mathematics lessons: an international comparative study of a collaboratively designed lesson. Educational Studies in Mathematics, 105(1):53–70, September 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-020-09980-1;   http://link.springer.
com/content/pdf/10.1007/s10649-020-09980-1.pdf.

Canadas:2018:MGA

[CMdR18] María C. Cañadas, Marta Molina, and Aurora del Río. Meanings given to algebraic symbolism in problem-posing. Educational Studies in Mathematics, 98(1):19–37, May 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9797-9.

Cai:2013:MPP

[CMW^{+}13] Jinfa Cai, John C. Moyer, Ning Wang, Stephen Hwang, Bikai Nie, and Tammy Garber. Mathematical problem posing as a measure of curricular effect on students’ learning. Educational Studies in Mathematics, 83(1):57–69, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9429-3.

Charles:2000:YCP

[CN00] Kathy Charles and Rod Nason. Young children’s partitioning strategies. Educational Studies in Mathematics, 43(2):191– 221, September 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017513716026.

Casey:2016:SUS

[CN16] Stephanie A. Casey and Courtney Nagle. Students’ use of slope conceptualizations when reasoning about the line of best fit. Educational Studies in Mathematics, 92(2):163–177, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9679-y.

Cespedes:2025:DMM

[CN25] María José Castillo Céspedes and María Burgos Navarro. Decision making in managing semiotic conflicts in curriculum materials: a training program with prospective teachers in Costa Rica. Educational Studies in Mathematics, 118(1):71– 99, January 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10363-z.

Caglayan:2010:EGS

[CO10] Günhan Caglayan and John Olive. Eighth grade students’ representations of linear equations based on a cups and tiles model. Educational Studies in Mathematics, 74(2):143–162, June 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9231-z.

Cobb:1985:TCA

[Cob85] Paul Cobb. Two children’s anticipations, beliefs, and motivations. Educational Studies in Mathematics, 16(2):111–126, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02400933.

Cobb:1986:CCA

[Cob86] Paul Cobb. Concrete can be abstract: A case study. Educational Studies in Mathematics, 17(1):37–48, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302377.

Cobb:1987:IYC

[Cob87] Paul Cobb. An investigation of young children’s academic arithmetic contexts. Educational Studies in Mathematics, 18 (2):109–124, May 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00314722.

Cobb:1994:GE

[Cob94] Paul Cobb. Guest editorial. Educational Studies in Mathematics, 26(2–3):105–109, March 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273659.

Coben:2004:BR

[Cob04] Diana Coben. Book review. Educational Studies in Mathematics, 57(2):295–300, September 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000049327.22222.48.

Coles:2021:CSI

[Col21] Alf Coles. Commentary on a special issue: Davydov’s approach in the XXI century. Educational Studies in Mathematics, 106(3):471–478, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-020-10018-
9.

Comiti:1977:ADN

[Com77] Claude Comiti. Approche du nombre par des enfants de 6 à 7 ans. (French) [Number approach by children aged 6– 7]. Educational Studies in Mathematics, 8(2):181–198, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241024.

Comiti:1980:PAN

[Com80] Claude Comiti. Les premières acquisitions de la notion de nombre par l’enfant. (French) [The first acquisitions of the notion of number by the child]. Educational Studies in Mathematics, 11(3):301–318, August 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00697742.

Connors:1990:WMM

[Con90] Jeanne Connors. When mathematics meets anthropology: The need for interdisciplinary dialogue. Educational Studies in Mathematics, 21(5):461–469, October 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00398864.

Cooper:1992:TDS

[Coo92] Martin Cooper. Three-dimensional symmetry. Educational Studies in Mathematics, 23(2):179–202, April 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00588055.

Cooper:1998:ANC

[Coo98] Barry Cooper. Assessing national curriculum mathematics in England: Exploring children’s interpretation of key stage 2 tests in clinical interviews. Educational Studies in Mathematics, 35(1):19–49, January 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002945216595.

Cooney:1999:CTW

[Coo99] Thomas J. Cooney. Conceptualizing teachers’ ways of knowing. Educational Studies in Mathematics, 38(1–3):163–187, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003504816467.

Coppe:1998:CPP

[Cop98] Sylvie Coppé. Composantes privées et publiques du travail de l’Élève en situation de devoir surveillé de mathématiques. (French) [Private and public components of student work in supervised mathematics]. Educational Studies in Mathematics, 35(2):129–151, February 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1003086628291.

Corriveau:2017:STC

[Cor17] Claudia Corriveau. Secondary-to-tertiary comparison through the lens of ways of doing mathematics in relation to functions: a study in collaboration with teachers. Educational Studies in Mathematics, 94(2):139–160, February 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9719-2.

Corven:2022:BRF

[Cor22] Julien Corven. Book review: Foundations for the next era of elementary mathematics teacher education. Yiping Li, Roger E. Howe, W. James Lewis, and James J. Madden (Eds.) (2021) Developing mathematical proficiency for elementary instruction. Springer International Publishing https://www.springer.com/gp/book/9783030689551 xvii + 298 pages. Hardcover: ISBN 978-3-030-68955-1 $149.99US. eBook: ISBN 978-3-030-68956-8 $109.99US. Educational Studies in Mathematics, 110(1):193–203, May 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10103-7.

Cotton:2024:BRL

[Cot24] Anthony Cotton. Book review: A life in critical mathematics education. Ole Skovsmose (2023) Critical Mathematics Education. Springer. Educational Studies in Mathematics, 116 (1):137–143, May 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10268-3.

Coxeter:1971:IG

[Cox71] H. S. M. Coxeter. Inversive geometry. Educational Studies in Mathematics, 3(3–4):310–321, June 1971. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302300.

Charalambous:2010:TCE

[CP10] Charalambos Y. Charalambous and George N. Philippou. Teachers’ concerns and efficacy beliefs about implementing a mathematics curriculum reform: integrating two lines of inquiry. Educational Studies in Mathematics, 75(1):1–21, September 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9238-5.

Charalambous:2008:TDP

[CPK08] Charalambos Y. Charalambous, George N. Philippou, and Leonidas Kyriakides. Tracing the development of preservice teachers’ efficacy beliefs in teaching mathematics during fieldwork. Educational Studies in Mathematics, 67(2):125–142, February 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-007-9084-2.

Charalambous:2007:DTM

[CPP07] Charalambos Y. Charalambous and Demetra Pitta-Pantazi. Drawing on a theoretical model to study students’ understandings of fractions. Educational Studies in Mathematics, 64(3): 293–316, March 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9036-2.

Charalambous:2009:UHM

[CPP09] Charalambos Y. Charalambous, Areti Panaoura, and George Philippou. Using the history of mathematics to induce changes in preservice teachers’ beliefs and attitudes: insights from evaluating a teacher education program. Educational Studies in Mathematics, 71(2):161–180, June 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9170-0.

Chimoni:2018:EEA

[CPPC18] Maria Chimoni, Demetra Pitta-Pantazi, and Constantinos Christou. Examining early algebraic thinking: insights from empirical data. Educational Studies in Mathematics, 98(1):57– 76, May 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9803-x.

Chimoni:2023:UAT

[CPPC23] Maria Chimoni, Demetra Pitta-Pantazi, and Constantinos Christou. Unfolding algebraic thinking from a cognitive perspective. Educational Studies in Mathematics, 114(1):89–108, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10218-z.

Chrysostomou:2013:ENS

[CPPT^{+}13] Marilena Chrysostomou, Demetra Pitta-Pantazi, Chara Tsingi, Eleni Cleanthous, and Constantinos Christou. Examining number sense and algebraic reasoning through cognitive styles. Educational Studies in Mathematics, 83(2):205– 223, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9448-0.

Clarke:2009:SFC

[CR09] Doug M. Clarke and Anne Roche. Students’ fraction comparison strategies as a window into robust understanding and possible pointers for instruction. Educational Studies in Mathematics, 72(1):127–138, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9198-9.

Crawford:1985:BR

[Cra85] Kathryn Crawford. Book review. Educational Studies in Mathematics, 16(4):431–433, November 1985. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417197.

Crawford:1996:VAH

[Cra96] Kathryn Crawford. Vygotskian approaches in human development in the information era. Educational Studies in Mathematics, 31(1–2):43–62, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00143926.

Crafter:2012:PCM

[Cra12] Sarah Crafter. Parental cultural models and resources for understanding mathematical achievement in culturally diverse school settings. Educational Studies in Mathematics, 81(1): 31–46, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9359-5.

Craig:2018:PN

[Cra18] Jeffrey Craig. The promises of numeracy. Educational Studies in Mathematics, 99(1):57–71, September 2018. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-018-9824-5.

Crespo:2003:LPM

[Cre03] S. Crespo. Learning to pose mathematical problems: Exploring changes in preservice teachers’ practices. Educational Studies in Mathematics, 52(3):243–270, April 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024364304664.

Crespo:2006:ETT

[Cre06] Sandra Crespo. Elementary teacher talk in mathematics study groups. Educational Studies in Mathematics, 63(1):29–56, September 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-9006-0.

Cueto:2006:OLA

[CRL06] Santiago Cueto, Cecilia Ramirez, and Juan Leon. Opportunities to learn and achievement in mathematics in a sample of sixth grade students in Lima, Peru. Educational Studies in Mathematics, 62(1):25–55, May 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-7922-2.

Castro-Rodriguez:2016:PTU

[CRPPRG16] Elena Castro-Rodríguez, Demetra Pitta-Pantazi, Luis Rico, and Pedro Gómez. Prospective teachers’ understanding of the multiplicative part-whole relationship of fraction. Educational Studies in Mathematics, 92(1):129–146, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9673-4.

Confrey:1994:EFR

[CS94] Jere Confrey and Erick Smith. Exponential functions, rates of change, and the multiplicative unit. Educational Studies in Mathematics, 26(2–3):135–164, March 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273661.

Coles:2019:REN

[CS19] Alf Coles and Nathalie Sinclair. Ritualisation in early number work. Educational Studies in Mathematics, 101 (2):177–194, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9815-
6;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9815-6.pdf.

Chumachenko:2025:EMD

[CSDK25] Dmitry Chumachenko, Anna Shvarts, Anna Dreneva, and Anatoly Krichevets. Eye movements in the development of geometric shape recognition: from sensory-motor processes to theoretical perception. Educational Studies in Mathematics, 118(3):529–554, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10298-5.

Csikos:2016:SPE

[Csí16] Csaba Csíkos. Strategies and performance in elementary students’ three-digit mental addition. Educational Studies in Mathematics, 91(1):123–139, January 2016. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9658-3.

Csikos:2012:EUD

[CSK12] Csaba Csíkos, Judit Szitányi, and Rita Kelemen. The effects of using drawings in developing young children’s mathematical word problem solving: A design experiment with thirdgrade Hungarian students. Educational Studies in Mathematics, 81(1):47–65, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9360-z.

Cavalcante:2024:MSS

[CSP24] Alexandre Cavalcante, Annie Savard, and Elena Polotskaia. Mathematical structures of simple and compound interest: An analysis of secondary teachers’ relational thinking. Educational Studies in Mathematics, 116(2):215–235, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10308-6.

Coles:2024:SEG

[CSRlR24] Alf Coles, Armando Solares-Rojas, and Kate le Roux. Socioecological gestures of mathematics education. Educational Studies in Mathematics, 116(2):165–183, June 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10318-4.

Conner:2014:TSC

[CSS^{+}14] AnnaMarie Conner, Laura M. Singletary, Ryan C. Smith, Patty Anne Wagner, and Richard T. Francisco. Teacher support for collective argumentation: A framework for examining how teachers support students’ engagement in mathematical activities. Educational Studies in Mathematics, 86 (3):401–429, July 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9532-8.

Chan:2021:MET

[CSW21] Man Ching Esther Chan, Cristina Sabena, and David Wagner. Mathematics education in a time of crisis — a viral pandemic. Educational Studies in Mathematics, 108(1–2):1–13, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10113-5.

Callejo:2009:AMP

[CV09] María Luz Callejo and Antoni Vila. Approach to mathematical problem solving and students’ belief systems: two case studies. Educational Studies in Mathematics, 72(1):111–126, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-009-9195-z.

Chow:2016:TSC

[CV16] Alan F. Chow and James P. Van Haneghan. Transfer of solutions to conditional probability problems: effects of example problem format, solution format, and problem context. Educational Studies in Mathematics, 93(1):67–85, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9691-x.

Chval:2021:ULS

[CVN21] Martin Chvál, Naďa Vondrová, and Jarmila Novotná. Using large-scale data to determine pupils’ strategies and errors in missing value number equations. Educational Studies in Mathematics, 106(1):5–24, January 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10000-5.

Cobb:1996:MCL

[CW96] Paul Cobb and Joy W. Whitenack. A method for conducting longitudinal analyses of classroom videorecordings and transcripts. Educational Studies in Mathematics, 30(3):213–228, April 1996. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00304566.

Chan:2014:WRB

[CW14a] Yip-Cheung Chan and Ngai-Ying Wong. Worldviews, religions, and beliefs about teaching and learning: perception of mathematics teachers with different religious backgrounds. Educational Studies in Mathematics, 87(3):251–277, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9555-1.

Coben:2014:MMD

[CW14b] Diana Coben and Keith Weeks. Meeting the mathematical demands of the safety-critical workplace: medication dosage calculation problem-solving for nursing. Educational Studies in Mathematics, 86(2):253–270, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9537-3.

Clark-Wilson:2019:RIW

[CWH19] Alison Clark-Wilson and Celia Hoyles. A research-informed web-based professional development toolkit to support technology-enhanced mathematics teaching at scale. Educational Studies in Mathematics, 102(3):343–359, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9836-1;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9836-1.pdf.

Cavey:2007:ITM

[CWL07] Laurie O. Cavey, Joy W. Whitenack, and Louann Lovin. Investigating teachers’ mathematics teaching understanding: A case for coordinating perspectives. Educational Studies in Mathematics, 64(1):19–43, January 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9031-7.

Clarke:1993:PSM

[CWS93] David J. Clarke, Andrew Waywood, and Max Stephens. Probing the structure of mathematical writing. Educational Studies in Mathematics, 25(3):235–250, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273863.

Che:2012:PSS

[CWT12] Megan Che, Elaine Wiegert, and Karen Threlkeld. Problem solving strategies of girls and boys in single-sex mathematics classrooms. Educational Studies in Mathematics, 79(2): 311–326, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9346-x.

Cobb:1992:FAS

[CWYP92] Paul Cobb, Terry Wood, Erna Yackel, and Marcela Perlwitz. A follow-up assessment of a second-grade problem-centered mathematics project. Educational Studies in Mathematics, 23(5):483–504, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00571469.

Campbell:2023:SNC

[CY23] Tye G. Campbell and Sheunghyun Yeo. Student noticing of collaborative practices: exploring how college students notice during small group interactions in math. Educational Studies in Mathematics, 113(3):405–423, July 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10206-3.

Cobb:1992:ILM

[CYW92] Paul Cobb, Erna Yackel, and Terry Wood. Interaction and learning in mathematics classroom situations. Educational Studies in Mathematics, 23(1):99–122, February 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302315.

Chernoff:2011:PCP

[CZ11] Egan J. Chernoff and Rina Zazkis. From personal to conventional probabilities: from sample set to sample space. Educational Studies in Mathematics, 77(1):15–33, May 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9288-8.

Czocher:2018:HDV

[Czo18] Jennifer A. Czocher. How does validating activity contribute to the modeling process? Educational Studies in Mathematics, 99(2):137–159, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9833-4.

DAmbrosio:1991:MMR

[D’A91] Beatriz Silva D’Ambrosio. The modern mathematics reform movement in Brazil and its consequences for Brazilian mathematics education. Educational Studies in Mathematics, 22(1): 69–85, February 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302717.

deAbreu:2000:RBM

[dA00] Guida de Abreu. Relationships between macro and micro socio-cultural contexts: implications for the study of interactions in the mathematics classroom. Educational Studies in Mathematics, 41(1):1–29, January 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003875728720.

DAmore:2010:OGT

[D’A10] Bruno D’Amore. Obituary: Giorgio Tomaso Bagni (1958– 2009). Educational Studies in Mathematics, 73(1):73–74, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-009-9218-9.pdf.

Dagher:1996:AEI

[Dag96] Antoine Dagher. Apprentissage dans un environnement informatique: possibilité nature transfert des acquis. (French) [Learning in a computer environment: native opportunity for transferring acquired knowledge]. Educational Studies in Mathematics, 30(4):367–398, June 1996. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00570830.

Dahl:2004:ACL

[Dah04] Bettina Dahl. Analysing cognitive learning processes through group interviews of successful high school pupils: Development and use of a model. Educational Studies in Mathematics, 56(2– 3):129–155, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040393.
88227.05.

Davies:2020:CJP

[DAJ20] Ben Davies, Lara Alcock, and Ian Jones. Comparative judgement, proof summaries and proof comprehension. Educational Studies in Mathematics, 105(2):181–197, October 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-09984-x.

Damerow:1987:EMI

[Dam87] Peter Damerow. Die Entstehung des Mathematiklehrerberufs im 19. Jahrhundert. Studien u. Materialien zum Prozeß d. Professionalisierung in Preußen (1810–1870). (German) [The emergence of the mathematics teacher’s profession in the 19th century. Studies and materials for the process of professionalisation in Prussia (1810–1870)]. Educational Studies in Mathematics, 18(4):451–453, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00240991.

Darke:1982:RRR

[Dar82] Ian Darke. A review of research related to the topological primacy thesis. Educational Studies in Mathematics, 13(2): 119–142, May 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00460707.

Darragh:2016:IRM

[Dar16] Lisa Darragh. Identity research in mathematics education. Educational Studies in Mathematics, 93(1):19–33, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9696-5.

deAraujo:2022:EEL

[dAS22] Zandra de Araujo and Erin Smith. Examining English language learners’ learning needs through the lens of algebra curriculum materials. Educational Studies in Mathematics, 109 (1):65–87, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10081-w.

Davis:1971:PRM

[Dav71] Robert B. Davis. The problem of relating mathematics to the possibilities and needs of schools and children. Educational Studies in Mathematics, 3(3–4):322–336, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302301.

Davis:1993:VT

[Dav93] Philip J. Davis. Visual theorems. Educational Studies in Mathematics, 24(4):333–344, December 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273369.

Davis:2007:BR

[Dav07] Brent Davis. Book review. Educational Studies in Mathematics, 64(3):373–379, March 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9061-1.

Dawe:1983:BMR

[Daw83] Lloyd Dawe. Bilingualism and mathematical reasoning in English as a second language. Educational Studies in Mathematics, 14(4):325–353, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368233.

Dawson:1994:BR

[Daw94] A. J. Dawson. Book review. Educational Studies in Mathematics, 27(3):307–315, October 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273733.

Denvir:1986:UNCa

[DB86a] B. Denvir and M. Brown. Understanding of number concepts in low attaining 7–9 year olds: Part I. Development of descriptive framework and diagnostic instrument. Educational Studies in Mathematics, 17(1):15–36, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302376.

Denvir:1986:UNCb

[DB86b] B. Denvir and M. Brown. Understanding of number concepts in low attaining 7–9 year olds: Part II. The teaching studies. Educational Studies in Mathematics, 17(2):143–164, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311518.

Durandt:2022:FMM

[DBL22] Rina Durandt, Werner Blum, and Alfred Lindl. Fostering mathematical modelling competency of south African engineering students: which influence does the teaching design have? Educational Studies in Mathematics, 109(2):361–381, February 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10068-7.

DAmbrosio:1992:PST

[DC92] Beatriz S. D’Ambrosio and Tânia Maria Mendonça Campos. Pre-service teachers’ representations of children’s understanding of mathematical concepts: conflicts and conflict resolution. Educational Studies in Mathematics, 23(3):213–230, June 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02309530.

Dawkins:2017:GRC

[DC17] Paul Christian Dawkins and John Paul Cook. Guiding reinvention of conventional tools of mathematical logic: students’ reasoning about mathematical disjunctions. Educational Studies in Mathematics, 94(3):241–256, March 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9722-7.

Dowker:2019:RBA

[DCHM19] Ann Dowker, Olivia Cheriton, Rachel Horton, and Winifred Mark. Relationships between attitudes and performance in young children’s mathematics. Educational Studies in Mathematics, 100(3):211–230, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-019-9880-
5;   http://link.springer.com/content/pdf/10.1007/
s10649-019-9880-5.pdf.

Drijvers:2010:TTI

[DDB^{+}10] Paul Drijvers, Michiel Doorman, Peter Boon, Helen Reed, and Koeno Gravemeijer. The teacher and the tool: instrumental orchestrations in the technology-rich mathematics classroom. Educational Studies in Mathematics, 75(2):213–234, November 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-010-9254-5.pdf.

Deng:2025:ILP

[DDH25] Xixi Deng, Rui Ding, and Rongjin Huang. Investigating a learning progression of functional thinking for elementary students. Educational Studies in Mathematics, 120(1):81–107, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10376-8.

Dubinsky:1994:LFC

[DDLZ94] Ed Dubinsky, Jennie Dautermann, Uri Leron, and Rina Zazkis. On learning fundamental concepts of group theory. Educational Studies in Mathematics, 27(3):267–305, October 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273732.

Dubinsky:1997:RBW

[DDLZ97] Ed Dubinsky, Jennie Dautermann, Uri Leron, and Rina Zazkis. A reaction to Burn’s “What Are The Fundamental Concepts of Group Theory?”. Educational Studies in Mathematics, 34(3):249–253, December 1997. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002940611963. See [Bur96].

DeLange:1993:BEB

[De 93] J. De Lange. Between end and beginning. Educational Studies in Mathematics, 25(1–2):137–160, March 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274106.

DeBlock-Docq:1994:MPM

[De 94] Christine De Block-Docq. Modalités de la pensée mathématique d’élèves de douze ans devant des problèmes de pavages. (French) [Modalities of mathematical thinking of twelveyear-old pupils in tiling problems]. Educational Studies in Mathematics, 27(2):165–189, September 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01278920.

Bock:2024:BRO

[De 24] Dirk De Bock. Book review: One hundred years of mathematics education seen through the lens of ICMI. Fulvia Furinghetti & Livia Giacardi (Eds.) (2022). The International Commission on Mathematical Instruction, 1908–2008: People, Events, and Challenges in Mathematics Education.

Educational Studies in Mathematics, 115(3):507–515, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10267-4.

Drageset:2024:UPT

[DE24] Ove Gunnar Drageset and Fiona Ell. Using positioning theory to think about mathematics classroom talk. Educational Studies in Mathematics, 115(3):353–385, March 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10295-0.

DeBlois:1997:TED

[DeB97] Lucie DeBlois. Trois élèves en difficulté devant des situations de réunion et de complément d’ensembles. (French) [Three students in difficulty in situations of union and complement of sets]. Educational Studies in Mathematics, 34(1):67–96, October 1997. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003061919494.

Deblois:2006:IIP

[Deb06] Lucie Deblois. Influence des interprétations des productions des élèves sur les stratégies d’intervention en classe de mathématiques. (French) [Influence of interpretations of students’ productions on strategies of intervention in a mathematics class]. Educational Studies in Mathematics, 62(3):307– 329, July 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-7833-2.

Dekker:1987:RJT

[Dek87] Rijkje Dekker. Roos and José, two children in a mixed ability group. Educational Studies in Mathematics, 18(3):317–324, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00386201.

Delessert:1968:QON

[Del68] A. Delessert. Quelques observations naïves en marge de deux enquêtes sur l’enseignement mathématique. (French) [Some observations made in the margins of two surveys on mathematical education]. Educational Studies in Mathematics, 1 (1–2):80–88, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00426233.

Delessert:1969:QPT

[Del69a] A. Delessert. De quelques problèmes touchant à la formation des maîtres de mathématiques. (French) [Some problems affecting the training of mathematics teachers]. Educational Studies in Mathematics, 2(2–3):245–256, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303461.

Delessert:1969:CBR

[Del69b] André Delessert. Connections between the reform of mathematical instruction and the further training of teachers. Educational Studies in Mathematics, 1(4):365–377, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02396309.

Demby:1993:LCA

[Dem93] Agnieszka Demby. L’usage de la compensation additionner– soustraire et multiplier–diviser par les élèves de onze ans. (French) [Use of compensation to add–subtract and multiply– divide by eleven-year-old students]. Educational Studies in Mathematics, 24(3):239–249, September 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01275424.

Demby:1997:APU

[Dem97] Agnieszka Demby. Algebraic procedures used by 13-to-15-year-olds. Educational Studies in Mathematics, 33(1):45–70, June 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002963922234.

Dekker:1998:PMI

[DEM98] Rijkje Dekker and Marianne Elshout-Mohr. A process model for interaction and mathematical level raising. Educational Studies in Mathematics, 35(3):303–314, March 1998.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003187204737.

Dekker:2004:TIA

[DEM04] Rijkje Dekker and Marianne Elshout-Mohr. Teacher interventions aimed at mathematical level raising during collaborative learning. Educational Studies in Mathematics, 56 (1):39–65, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000028402.
10122.ff.

Dekker:2006:HCR

[DEMW06] Rijkje Dekker, Marianne Elshout-Mohr, and Terry Wood. How children regulate their own collaborative learning. Educational Studies in Mathematics, 62(1):57–79, May 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-1688-4.

deFreitas:2004:PIA

[dF04] Elizabeth de Freitas. Plotting intersections along the political axis: The interior voice of dissenting mathematics teachers. Educational Studies in Mathematics, 55(1–3):259–274, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017694.18374.d6.

deFreitas:2012:BRT

[dF12] Elizabeth de Freitas. Book review: Teacher identity: Harsh lessons learned from the micro-management of teachers and teacher education: A review of Tony Brown and Olwen McNamara’s Becoming a mathematics teacher: Identity and identifications (2011). Educational Studies in Mathematics, 81 (1):133–137, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9374-6.

Davis:2015:AFC

[DF15] Jon D. Davis and Nicole L. Fonger. An analytical framework for categorizing the use of CAS symbolic manipulation in textbooks. Educational Studies in Mathematics, 88(2):239–258, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9581-z.

deFreitas:2016:MEM

[dF16] Elizabeth de Freitas. Material encounters and media events: what kind of mathematics can a body do? Educational Studies in Mathematics, 91(2):185–202, February 2016. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-015-9657-4.pdf.

deFreitas:2025:XPP

[dF25] Elizabeth de Freitas. Xenomathematics as a planetary practice: imagining more-than-human mathematical worlds. Educational Studies in Mathematics, 120(3):551–566, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10424-x.

deFreitas:2012:DGA

[dFS12] Elizabeth de Freitas and Nathalie Sinclair. Diagram, gesture, agency: theorizing embodiment in the mathematics classroom. Educational Studies in Mathematics, 80(1–2):133–152, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9364-8.

deFreitas:2013:NMO

[dFS13] Elizabeth de Freitas and Nathalie Sinclair. New materialist ontologies in mathematics education: the body in/of mathematics. Educational Studies in Mathematics, 83(3):453–470, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9465-z.

Dupuis:1989:GRE

[DG89] Claire Dupuis and Dominique Guin. Gestion des relations entre variables dans un environnement de programmation logo. (French) [Managing relationships between variables in a Logo programming environment]. Educational Studies in Mathematics, 20(3):293–316, August 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00310875.

Durand-Guerrier:2003:WNI

[DG03] V. Durand-Guerrier. Which notion of implication is the right one? From logical considerations to a didactic perspective. Educational Studies in Mathematics, 53(1):5–34, May 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024661004375.

DeBellis:2006:AMA

[DG06] Valerie A. DeBellis and Gerald A. Goldin. Affect and metaaffect in mathematical problem solving: a representational perspective. Educational Studies in Mathematics, 63(2):131– 147, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9026-4.

Durand-Guerrier:2005:EDS

[DGA05] Viviane Durand-Guerrier and Gilbert Arsac. An epistemological and didactic study of a specific calculus reasoning rule. Educational Studies in Mathematics, 60(2):149–172, October 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5614-y.

Drijvers:2013:OET

[DGFT13] Paul Drijvers, Juan D. Godino, Vicenç Font, and Luc Trouche. One episode, two lenses. Educational Studies in Mathematics, 82(1):23–49, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-012-9416-8.
pdf.

DiMartino:2023:TSUa

[DGI23a] Pietro Di Martino, Francesca Gregorio, and Paola Iannone. Transition from school into university mathematics: experiences across educational contexts. Educational Studies in Mathematics, 113(1):1–5, May 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10217-0.

DiMartino:2023:TSUb

[DGI23b] Pietro Di Martino, Francesca Gregorio, and Paola Iannone. The transition from school to university mathematics in different contexts: affective and sociocultural issues in students’ crisis. Educational Studies in Mathematics, 113(1):79–106, May 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10179-9.

Davis:1990:SPP

[DH90] Gary Ernest Davis and Robert Paul Hunting. Spontaneous partitioning: Pre-schoolers and discrete items. Educational Studies in Mathematics, 21(4):367–374, August 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304264.

Dahlberg:1997:FLE

[DH97] Randall P. Dahlberg and David L. Housman. Facilitating learning events through example generation. Educational Studies in Mathematics, 33(3):283–299, September 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002999415887.

Dimmel:2020:PTH

[DH20] Justin K. Dimmel and Patricio G. Herbst. Proof transcription in high school geometry: a study of what teachers recognize as normative when students present proofs at the board. Educational Studies in Mathematics, 105(1):71–89, September 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09975-y.

Ding:2023:MPD

[DHD23] Rui Ding, Rongjin Huang, and Xixi Deng. Multiple pathways for developing functional thinking in elementary mathematics textbooks: a case study in China. Educational Studies in Mathematics, 114(2):223–248, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10237-w.

Dhombres:1993:OPE

[Dho93] Jean Dhombres. Is one proof enough? Travels with a mathematician of the baroque period. Educational Studies in Mathematics, 24(4):401–419, December 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273373.

Martino:2019:PVP

[Di 19] Pietro Di Martino. Pupils’ view of problems: the evolution from kindergarten to the end of primary school. Educational Studies in Mathematics, 100(3):291–307, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9850-3.

Dibbs:2019:FFE

[Dib19] Rebecca Dibbs. Forged in failure: engagement patterns for successful students repeating calculus. Educational Studies in Mathematics, 101(1):35–50, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-019-9877-
0.

Dienes:1970:SEA

[Dié70] Z. P. Diénès. Les six étapes de l’apprentissage des structures. (French) [The six stages of learning structures]. Educational Studies in Mathematics, 3(1):12–42, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381592.

Dienes:1971:EPC

[Dié71] Zoltan P. Diénès. An example of the passage from the concrete to the manipulation of formal systems. Educational Studies in Mathematics, 3(3–4):337–352, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302302.

Dietiker:2015:MSM

[Die15] Leslie Dietiker. Mathematical story: a metaphor for mathematics curriculum. Educational Studies in Mathematics, 90(3):

285–302, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9627-x.

Dimattia:2004:IOR

[Dim04] Cara Dimattia. Instructions for opening, reading and viewing the videopapers on the attached CD. Educational Studies in Mathematics, 57(3):323–324, November 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-004-5934-3.

Dawkins:2019:UM

[DIW19] Paul Christian Dawkins, Matthew Inglis, and Nicholas Wasserman. The use(s) of is in mathematics. Educational Studies in Mathematics, 100(2):117–137, February 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9868-6.

Dreher:2015:TPK

[DK15] Anika Dreher and Sebastian Kuntze. Teachers’ professional knowledge and noticing: The case of multiple representations in the mathematics classroom. Educational Studies in Mathematics, 88(1):89–114, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9577-8.

Drijvers:2019:AMT

[DKBD19] Paul Drijvers, Hanneke Kodde-Buitenhuis, and Michiel Doorman. Assessing mathematical thinking as part of curriculum reform in The Netherlands. Educational Studies in Mathematics, 102(3):435–456, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09905-
7;   http://link.springer.com/content/pdf/10.1007/

s10649-019-09905-7.pdf. See correction [DKBD21].

Drijvers:2021:CAM

[DKBD21] Paul Drijvers, Hanneke Kodde-Buitenhuis, and Michiel Doorman. Correction to: Assessing mathematical thinking as part of curriculum reform in the Netherlands. Educational Studies in Mathematics, 107(2):425, June 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10061-0. See [DKBD19].

deKoning:2022:MMD

[dKBvdS22] Björn B. de Koning, Anton J. H. Boonen, and Menno van der Schoot. Model method drawing acts as a doubleedged sword for solving inconsistent word problems. Educational Studies in Mathematics, 111(1):29–45, September 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10150-8.

Dewantara:2025:BRM

[DKL^{+}25] Andi Harpeni Dewantara, Kartianom, Yustina Dwisofiani Lawung, Magy Gaspersz, and Sitti Busyrah Muchsin. Book review: mapping mathematics education research in Asia: insights across diverse countries. Bill Atweh, Lianghuo Fan, Catherine P. Vistro-Yu (Eds.) (2023) Asian research in mathematics education: mapping the field. Educational Studies in Mathematics, 119(2):411–417, June 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10364-y.

deToledoeToledo:2018:MEN

[dKV18] Neila de Toledo e Toledo, Gelsa Knijnik, and Paola Valero. Mathematics education in the neoliberal and corporate curriculum: the case of Brazilian agricultural high schools. Educational Studies in Mathematics, 99(1):73–87, September 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9825-4.

Ding:2014:TCA

[DL14] Meixia Ding and Xiaobao Li. Transition from concrete to abstract representations: the distributive property in a Chinese textbook series. Educational Studies in Mathematics, 87(1):103–121, September 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9558-y.

Dominguez:2014:REP

[DLK14] Higinio Dominguez, Carlos A. López Leiva, and Lena Licón Khisty. Relational engagement: Proportional reasoning with bilingual Latino/a students. Educational Studies in Mathematics, 85(1):143–160, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9501-7.

deLima:2008:PEM

[dLT08] Rosana Nogueira de Lima and David Tall. Procedural embodiment and magic in linear equations. Educational Studies in Mathematics, 67(1):3–18, January 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9086-0.

Diego-Mantecon:2021:CCB

[DMHRV21] José Manuel Diego-Mantecón, Elena Haro, and Avenilde Romo-Vázquez. The chimera of the competency-based approach to teaching mathematics: a study of carpentry purchases for home projects. Educational Studies in Mathematics, 107(2):339–357, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10032-5.

Doerr:2006:ETT

[Doe06] Helen M. Doerr. Examining the tasks of teaching when using students’ mathematical thinking. Educational Studies in Mathematics, 62(1):3–24, May 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-4437-9.

Dominguez:2011:UWM

[Dom11] Higinio Domínguez. Using what matters to students in bilingual mathematics problems. Educational Studies in Mathematics, 76(3):305–328, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9284-z.

Dominguez:2019:TRN

[Dom19] Higinio Dominguez. Theorizing reciprocal noticing with non-dominant students in mathematics. Educational Studies in Mathematics, 102(1):75–89, September 2019. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-019-09896-5.

Dominguez:2025:SLM

[Dom25] Higinio Dominguez. The secret life of mathematical concepts. Educational Studies in Mathematics, 120(3):567–586, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10413-0.

Dorfler:1990:E

[Dör90] W. Dörfler. Editorial. Educational Studies in Mathematics, 21 (1):1–2, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311012.

Dorfler:1992:E

[Dör92] W. Dörfler. Editorial. Educational Studies in Mathematics, 23(4):315, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302438.

Dorfler:1993:Ea

[Dör93a] W. Dörfler. Editorial. Educational Studies in Mathematics, 24(1):v, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273291.

Dorfler:1993:Eb

[Dör93b] W. Dörfler. Editorial. Educational Studies in Mathematics, 25 (1–2):v–vi, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274097.

Dorfler:1994:E

[Dör94] W. Dörfler. Editorial. Educational Studies in Mathematics, 27 (2):iii, September 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01278916.

Dorfler:1995:Ea

[Dör95a] W. Dörfler. Editorial. Educational Studies in Mathematics, 28(1):v, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273853.

Dorfler:1995:Eb

[Dör95b] W. Dörfler. Editorial. Educational Studies in Mathematics, 29 (4):iii, December 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273908.

Dorfler:2003:MME

[Dör03a] Willi Dörfler. Mathematics and mathematics education: Content and people, relation and difference. Educational Studies in Mathematics, 54(2–3):147–170, December 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006118.25919.07.

Dorfler:2003:IFE

[Dör03b] Willibald Dörfler. Introduction: A focus on experiences — my experience with Leen. Educational Studies in Mathematics, 54 (1):5–7, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000005234.
84301.6b.

Douady:1976:EMF

[Dou76] Régine Douady. Une expérience à Montrouge. (French) [An experiment at Montrouge]. Educational Studies in Mathematics, 7(1–2):49–58, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00144358.

Douek:1999:ACC

[Dou99] Nadia Douek. Argumentation and conceptualization in context: a case study on sunshadows in primary school. Educational Studies in Mathematics, 39(1–3):89–110, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003800814251.

DeKeersmaeker:2024:TAM

[DOV24] Karen De Keersmaeker, Patrick Onghena, and Wim Van Dooren. Towards an advanced mathematical language task for the early primary school years: preliminary results on reliability and validity. Educational Studies in Mathematics, 117 (3):511–531, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10342-4.

Dowling:1996:SAS

[Dow96] Paul Dowling. A sociological analysis of school mathematics texts. Educational Studies in Mathematics, 31(4):389–415, December 1996. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00369156.

Duval:1977:DIR

[DP77] R. Duval and F. Pluvinage. Démarches individuelles de réponse en mathématique. (French) [Individual responses in mathematics]. Educational Studies in Mathematics, 8(1):51– 116, April 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302505.

Davis:1992:MPS

[DP92] Gary Davis and Kristine Pepper. Mathematical problem solving by pre-school children. Educational Studies in Mathematics, 23(4):397–415, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302442.

Dapueto:1999:COC

[DP99] Carlo Dapueto and Laura Parenti. Contributions and obstacles of contexts in the development of mathematical knowledge. Educational Studies in Mathematics, 39(1–3):1–21, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003702003886.

Douady:1989:PAD

[DPG89] Regine Douady and Marie-Jeanne Perrin-Glorian. Un processus d’apprentissage du concept d’aire de surface plane. (French) [A learning process for the concept of the area of a flat surface]. Educational Studies in Mathematics, 20(4):387– 424, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00315608.

daPonte:2013:PPB

[dPH13] João Pedro da Ponte and Ana Henriques. Problem posing based on investigation activities by university students. Educational Studies in Mathematics, 83(1):145–156, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9443-5.

daPonte:2016:TPP

[dPQ16] João Pedro da Ponte and Marisa Quaresma. Teachers’ professional practice conducting mathematical discussions. Educational Studies in Mathematics, 93(1):51–66, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9681-z.

Davis:2013:PUE

[DR13] Brent Davis and Moshe Renert. Profound understanding of emergent mathematics: broadening the construct of teachers’ disciplinary knowledge. Educational Studies in Mathematics, 82(2):245–265, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9424-8.

Drageset:2014:RPF

[Dra14] Ove Gunnar Drageset. Redirecting, progressing, and focusing actions — a framework for describing how teachers use students’ comments to work with mathematics. Educational Studies in Mathematics, 85(2):281–304, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9515-1.

Dreyfus:1993:BR

[Dre93] Tommy Dreyfus. Book review. Educational Studies in Mathematics, 24(2):231–234, February 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273694.

Dreyfus:1995:GE

[Dre95] Tommy Dreyfus. Guest editorial. Educational Studies in Mathematics, 29(2):93–95, September 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274208.

Dreyfus:1999:WJC

[Dre99] Tommy Dreyfus. Why Johnny can’t prove. Educational Studies in Mathematics, 38(1–3):85–109, March 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003660018579.

Dreyfus:2006:MEE

[Dre06] Tommy Dreyfus. E, S, and M as in ESM– an editorial. Educational Studies in Mathematics, 62(1):1–2, May 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9044-2.

Dreyfus:2007:ERA

[Dre07] Tommy Dreyfus. Editorial: Review articles. Educational Studies in Mathematics, 66(1):1, September 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9095-z.

Dreyfus:2008:E

[Dre08] Tommy Dreyfus. Editorial. Educational Studies in Mathematics, 69(3):195–196, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9156-y.

Drimalla:2025:IMO

[Dri25] James Drimalla. Inferences, mastery, and objectivity: examining student understanding of mathematical concepts with inferentialism. Educational Studies in Mathematics, 119(2):393– 410, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10377-7.

Dorier:2002:SCR

[DRR02] Jean-Luc Dorier, Aline Robert, and Marc Rogalski. Some comments on ‘The role of proof in comprehending and teaching elementary linear algebra’ by F. Uhlig. Educational Studies in Mathematics, 51(3):185–192, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023624601930. See [Uhl02, Uhl03].

DeBlois:2002:IAP

[DS02] Lucie DeBlois and Hassan Squalli. Implication de l’analyse de productions d’élèves dans la formation des maîtres du primaire. (French) [Involvement of the analysis of pupil productions in the training of primary-school teachers]. Educational Studies in Mathematics, 50(2):212–237, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021116014594.

Duffin:2005:CET

[DS05] Janet Duffin and Adrian Simpson. Cognitive empathy and the transition to independent graduate study in mathematics. Educational Studies in Mathematics, 58(1):121–135, January 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-2384-5.

Davis:2006:MTO

[DS06] Brent Davis and Elaine Simmt. Mathematics-for-teaching: an ongoing investigation of the mathematics that teachers (need to) know. Educational Studies in Mathematics, 61(3):293–319, March 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-2372-4.

Downton:2017:PCP

[DS17] Ann Downton and Peter Sullivan. Posing complex problems requiring multiplicative thinking prompts students to use sophisticated strategies and build mathematical connections. Educational Studies in Mathematics, 95(3):303–328, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Dobie:2021:LMT

[DS21] Tracy E. Dobie and Bruce Sherin. The language of mathematics teaching: a text mining approach to explore the zeitgeist of US mathematics education. Educational Studies in Mathematics, 107(1):159–188, May 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-020-10019-
8.

daSilva:2021:CEB

[dSBM21] Alexandre Sousa da Silva, Maria Tereza Serrano Barbosa, and Marcos Nascimento Magalhães. The COVID-19 epidemic in Brazil: how statistics education may contribute to unravel the reality behind the charts. Educational Studies in Mathematics, 108(1–2):269–289, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10112-6.

Dietiker:2023:NCC

[DSR^{+}23] Leslie Dietiker, Rashmi Singh, Meghan Riling, Hector I. Nieves, and Erin Barno. Narrative characteristics of captivating secondary mathematics lessons. Educational Studies in Mathematics, 112(3):481–504, March 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10184-y.

Dogan:2008:ICU

[DT08] Enis Dogan and Kikumi Tatsuoka. An international comparison using a diagnostic testing model: Turkish students’ profile of mathematical skills on TIMSS-R. Educational Studies in Mathematics, 68(3):263–272, July 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9099-8.

David:2012:RVR

[DT12] Maria Manuela David and Vanessa Sena Tomaz. The role of visual representations for structuring classroom mathematical activity. Educational Studies in Mathematics, 80(3):413– 431, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9358-6.

Drijvers:2021:DMT

[DTD21] Paul Drijvers, Daniel Thurm, and Michiel Doorman. Distance mathematics teaching in Flanders, Germany, and The Netherlands during COVID-19 lockdown. Educational Studies in Mathematics, 108(1–2):35–64, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10094-5.

Dubois:1984:SCC

[Dub84] Jean-Guy Dubois. Une systématique des configurations combinatoires simples. (French) [A system of simple combinatorial configurations]. Educational Studies in Mathematics, 15(1): 37–57, February 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00380438.

Dumont:1982:LDD

[Dum82] Bernard Dumont. L’influence du ‘décor’ et du langage dans des épreuves de type ‘logique’ portant apparemment sur l’implication. (French) [The influence of ‘decor’ and language in ‘logic’ type tests apparently involving implication]. Educational Studies in Mathematics, 13(4):409–429, November 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00366620.

Duval:1983:LDD

[Duv83] Raymond Duval. L’obstacle du dédoublement des objets mathématiques. (French) [The obstacle of the duplication of mathematical objects]. Educational Studies in Mathematics, 14(4):385–414, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368236.

Duval:1991:SDR

[Duv91] Raymond Duval. Structure du raisonnement déductif et apprentissage de la démonstration. (French) [Structure of deductive reasoning and learning of demonstration]. Educational Studies in Mathematics, 22(3):233–261, June 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368340.

Duval:2006:CAP

[Duv06] Raymond Duval. A cognitive analysis of problems of comprehension in a learning of mathematics. Educational Studies in Mathematics, 61(1–2):103–131, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-0400-z.

deVittori:2018:AUH

[dV18] Thomas de Vittori. Analyzing the use of history in mathematics education: issues and challenges around Balacheff’s cK c̸ model. Educational Studies in Mathematics, 99(2):125–136, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9831-6.

Darragh:2019:BLP

[DVC19] Lisa Darragh and Luz Valoyes-Chávez. Blurred lines: producing the mathematics student through discourses of special educational needs in the context of reform mathematics in Chile. Educational Studies in Mathematics, 101(3):425– 439, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9875-7.

Demonty:2018:ATP

[DVF18] Isabelle Demonty, Joëlle Vlassis, and Annick Fagnant. Algebraic thinking, pattern activities and knowledge for teaching at the transition between primary and secondary school. Educational Studies in Mathematics, 99(1):1–19, September 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9820-9.

DeBock:1998:PLM

[DVJ98] Dirk De Bock, Lieven Verschaffel, and Dirk Janssens. The predominance of the linear model in secondary school students’ solutions of word problems involving length and area of similar plane figures. Educational Studies in Mathematics, 35(1): 65–83, January 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003151011999.

DeBock:2002:IUL

[DVJV02] Dirk De Bock, Wim Van Dooren, Dirk Janssens, and Lieven Verschaffel. Improper use of linear reasoning: An in-depth study of the nature and the irresistibility of secondary school students’ errors. Educational Studies in Mathematics, 50(3): 311–334, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1021205413749.

deVetten:2018:DIS

[dVSKvO18] Arjen de Vetten, Judith Schoonenboom, Ronald Keijzer, and Bert van Oers. The development of informal statistical inference content knowledge of pre-service primary school teachers during a teacher college intervention. Educational Studies in Mathematics, 99(2):217–234, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9823-6;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9823-6.pdf.

Depaepe:2018:SPS

[DVT^{+}18] Fien Depaepe, Patrick Van Roy, Joke Torbeyns, Thilo Kleickmann, Wim Van Dooren, and Lieven Verschaffel. Stimulating pre-service teachers’ content and pedagogical content knowledge on rational numbers. Educational Studies in Mathematics, 99(2):197–216, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9822-7.

Degrande:2020:AMO

[DVV20] Tine Degrande, Lieven Verschaffel, and Wim Van Dooren. To add or to multiply in open problems? Unraveling children’s relational preference using a mixed-method approach. Educational Studies in Mathematics, 104(3):405–430, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09966-z.

Dawkins:2017:VNP

[DW17] Paul Christian Dawkins and Keith Weber. Values and norms of proof for mathematicians and students. Educational Studies in Mathematics, 95(2):123–142, June 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Dubinsky:2005:SHIb

[DWMB05a] Ed Dubinsky, Kirk Weller, Michael A. McDonald, and Anne Brown. Some historical issues and paradoxes regarding the concept of infinity: An APOS analysis: Part 2. Educational Studies in Mathematics, 60(2):253–266, October 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-0473-0.

Dubinsky:2005:SHIa

[DWMB05b] Ed Dubinsky, Kirk Weller, Michael A. Mcdonald, and Anne Brown. Some historical issues and paradoxes regarding the concept of infinity: An APOS-based analysis: Part 1. Educational Studies in Mathematics, 58(3):335–359, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-2531-z.

Dogan:2021:RGE

[DWP21] Muhammed Fatih Dogan and Caro Williams-Pierce. The role of generic examples in teachers’ proving activities. Educational Studies in Mathematics, 106(1):133–150, January 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-10002-3.

Doerr:2000:CMG

[DZ00] Helen M. Doerr and Roxana Zangor. Creating meaning for and with the graphing calculator. Educational Studies in Mathematics, 41(2):143–163, February 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003905929557.

Doz:2025:IBL

[DŽC25] Daniel Doz, Amalija Žakelj, and Mara Cotič. Inquiry-based learning in Grade 9 mathematics: assessing outcomes across Gagné’s taxonomy. Educational Studies in Mathematics, 120 (2):269–298, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-025-10417-w.

Eccles:1985:MSM

[Ecc85] Jacquelynne Eccles. Model of students’ mathematics enrollment decisions. Educational Studies in Mathematics, 16(3): 311–314, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00776739.

Eddy:2018:BRM

[Edd18] Colleen M. Eddy. Book review: Mathematics for all: From fantasy to reality. Murad Jurdak and Renuka Vithal (Eds.) (2018) Sociopolitical dimensions of mathematics education: from the margin to the mainstream. Educational Studies in Mathematics, 99(3):359–362, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9837-0.

Edwards:1984:CEN

[Edw84] A. Edwards. Computational estimation for numeracy. Educational Studies in Mathematics, 15(1):59–73, February 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00380439.

Edwards:1993:BR

[Edw93] Laurie D. Edwards. Book review. Educational Studies in Mathematics, 24(3):319–327, September 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01275430.

Edwards:2009:GCI

[Edw09] Laurie D. Edwards. Gestures and conceptual integration in mathematical talk. Educational Studies in Mathematics, 70 (2):127–141, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9124-6.

Eriksson:2021:LAI

[EE21] Helena Eriksson and Inger Eriksson. Learning actions indicating algebraic thinking in multilingual classrooms. Educational Studies in Mathematics, 106(3):363–378, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10007-y.

Ellis:2020:SCV

[EEST20] Amy Ellis, Robert Ely, Brandon Singleton, and Halil Tasova. Scaling-continuous variation: supporting students’ algebraic reasoning. Educational Studies in Mathematics, 104(1):87– 103, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09951-6.

Erbilgin:2023:USA

[EG23] Evrim Erbilgin and Serigne M. Gningue. Using the ontosemiotic approach to analyze novice algebra learners’ meaningmaking processes with different representations. Educational Studies in Mathematics, 114(2):337–357, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10247-8.

Egsgard:1970:SIG

[Egs70] John C. Egsgard. Some ideas in geometry that can be taught from K–6. Educational Studies in Mathematics, 2(4):478–495, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00417630.

Engelbrecht:2005:TUMa

[EH05a] Johann Engelbrecht and Ansie Harding. Teaching undergraduate mathematics on the Internet. Part 1: Technologies and taxonomy. Educational Studies in Mathematics, 58(2):235– 252, February 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-6456-3.

Engelbrecht:2005:TUMb

[EH05b] Johann Engelbrecht and Ansie Harding. Teaching undergraduate mathematics on the Internet. Part 2. Educational Studies in Mathematics, 58(2):253–276, February 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-6457-2.

Eynde:2006:CSF

[EH06] Peter Op’t Eynde and Markku S. Hannula. The case study of Frank. Educational Studies in Mathematics, 63(2):123–129, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9030-8.

Ehrenfeld:2020:IEF

[EH20] Nadav Ehrenfeld and Ilana S. Horn. Initiation-entry-focus-exit and participation: a framework for understanding teacher groupwork monitoring routines. Educational Studies in Mathematics, 103(3):251–272, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09939-
2.

Ehrenfeucht:1978:CME

[Ehr78] Aniela Ehrenfeucht. Change in mathematics education since the late 1950’s — ideas and realisation Poland. Educational Studies in Mathematics, 9(3):283–295, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241031.

ElMouhayar:2013:TAI

[EJ13] Rabih Raif El Mouhayar and Murad Eid Jurdak. Teachers’ ability to identify and explain students’ actions in near and far figural pattern generalization tasks. Educational Studies in Mathematics, 82(3):379–396, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9434-6.

Evans:2023:CMA

[EJ23] Tanya Evans and Inae Jeong. Concept maps as assessment for learning in university mathematics. Educational Studies in Mathematics, 113(3):475–498, July 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10209-0.

Empson:2006:FCM

[EJDT06] Susan B. Empson, Debra Junk, Higinio Dominguez, and Erin Turner. Fractions as the coordination of multiplicatively related quantities: A cross-sectional study of children’s thinking. Educational Studies in Mathematics, 63(1):1–28, September 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9000-6.

Even:2010:WMD

[EK10] Ruhama Even and Tova Kvatinsky. What mathematics do teachers with contrasting teaching approaches address in probability lessons? Educational Studies in Mathematics, 74(3): 207–222, July 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-010-9234-9.

Ekenstam:1977:CQU

[Eke77] Adolf Af Ekenstam. On children’s quantitative understanding of numbers. Educational Studies in Mathematics, 8(3): 317–332, October 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00385928.

ElSawi:1978:CME

[El 78] M. El Sawi. Change in mathematics education since the late 1950’s — ideas and realisation Sudan. Educational Studies in Mathematics, 9(3):317–330, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241034.

ElTom:1986:AME

[El 86] M. E. A. El Tom. Aspects of mathematics education in some third world universities. Educational Studies in Mathematics, 17(2):165–191, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311519.

Mouhayar:2018:TPS

[El 18] Rabih El Mouhayar. Trends of progression of student level of reasoning and generalization in numerical and figural reasoning approaches in pattern generalization. Educational Studies in Mathematics, 99(1):89–107, September 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9821-8.

Elbers:2003:CIR

[Elb03] Ed Elbers. Classroom interaction as reflection: Learning and teaching mathematics in a community of inquiry. Educational Studies in Mathematics, 54(1):77–99, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000005211.95182.90.

Elizabeth:1985:ESR

[Eli85]  Fennema Elizabeth.  Explaining sex-related differences

in mathematics: Theoretical models. Educational Studies in Mathematics, 16(3):303, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00776735.

Ellerton:1986:CMM

[Ell86] Nerida F. Ellerton. Children’s made-up mathematics problems — a new perspective on talented mathematicians. Educational Studies in Mathematics, 17(3):261–271, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305073.

Ellerton:2013:EPS

[Ell13] Nerida F. Ellerton. Engaging pre-service middle-school teacher-education students in mathematical problem posing: development of an active learning framework. Educational Studies in Mathematics, 83(1):87–101, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9449-z.

Ellerton:2018:BRN

[Ell18] Nerida F. Ellerton. Book review: NCTM’s Compendium: finding a balance between historical details, contemporary practices, and future resources. Jinfa Cai (Ed.) (2017) Compendium for research in mathematics education. Educational Studies in Mathematics, 99(1):109–123, September 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9827-2.

Ely:2010:BRT

[Ely10] Robert Ely. Book review: Theories of mathematics education: Seeking new frontiers. Educational Studies in Mathematics, 75(2):235–240, November 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9250-9.

Evans:2022:DMU

[EMRI22] Tanya Evans, Juan Pablo Mejía-Ramos, and Matthew Inglis. Do mathematicians and undergraduates agree about explanation quality? Educational Studies in Mathematics, 111(3): 445–467, November 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10164-2.

Evans:2006:DPE

[EMT06] Jeff Evans, Candia Morgan, and Anna Tsatsaroni. Discursive positioning and emotion in school mathematics practices. Educational Studies in Mathematics, 63(2):209–226, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9029-1.

Ekenstam:1979:NAA

[EN79] Adolf A. F. Ekenstam and Margita Nilsson. A new approach to the assessment of children’s mathematical competence. Educational Studies in Mathematics, 10(1):41–66, February 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311174.

Engel:1968:SUA

[Eng68] Arthur Engel. Systematic use of applications in mathematics teaching. Educational Studies in Mathematics, 1(1–2): 202–221, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00426244;   http:/
/link.springer.com/article/10.1007/BF00426244.

Engel:1969:RMF

[Eng69] Arthur Engel. The relevance of modern fields of applied mathematics for mathematical education. Educational Studies in Mathematics, 2(2–3):257–269, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303462.

Engel:1971:GAU

[Eng71] Arthur Engel. Geometrical activities for the upper elementary school. Educational Studies in Mathematics, 3(3–4):353– 394, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302303.

Engel:1975:PA

[Eng75] Arthur Engel. The probabilistic abacus. Educational Studies in Mathematics, 6(1):1–22, March 1975. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00590021.

Engel:1976:WDP

[Eng76] Arthur Engel. Why does the probabilistic abacus work? Educational Studies in Mathematics, 7(1–2):59–69, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144359.

English:1991:YCC

[Eng91] Lyn D. English. Young children’s combinatoric strategies. Educational Studies in Mathematics, 22(5):451–474, October 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367908.

English:1997:DFG

[Eng97] Lyn D. English. The development of fifth-grade children’s problem-posing abilities. Educational Studies in Mathematics, 34(3):183–217, December 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002963618035.

English:1999:BRL

[Eng99] Lyn English. Book review: Larochelle, M., Bednarz, N. and Garrison, J. (eds.), Constructivism and Education. Educational Studies in Mathematics, 40(1):93–99, September 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003750912432.

English:2003:RTR

[Eng03] Lyn D. English. Reconciling theory, research, and practice: A models and modelling perspective. Educational Studies in Mathematics, 54(2–3):225–248, December 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006167.14146.7b.

English:2006:MMP

[Eng06] Lyn D. English. Mathematical modeling in the primary school: Children’s construction of a consumer guide. Educational Studies in Mathematics, 63(3):303–323, November 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9013-1.

English:2012:DMF

[Eng12] Lyn D. English. Data modelling with first-grade students. Educational Studies in Mathematics, 81(1):15–30, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9377-3.

Erath:2018:DCI

[EPQH18] Kirstin Erath, Susanne Prediger, Uta Quasthoff, and Vivien Heller. Discourse competence as important part of academic language proficiency in mathematics classrooms: the case of explaining to learn and learning to explain. Educational Studies in Mathematics, 99(2):161–179, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9830-7;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9830-7.pdf.

Ernest:1984:CME

[Ern84a] Paul Ernest. The CXC mathematics examination: A Caribbean innovation in assessment. Educational Studies in Mathematics, 15(4):397–412, November 1984. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311114.

Ernest:1984:MIP

[Ern84b] Paul Ernest. Mathematical induction: A pedagogical discussion. Educational Studies in Mathematics, 15(2):173–189, May 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305895.

Ernest:1985:NLT

[Ern85] Paul Ernest. The number line as a teaching aid. Educational Studies in Mathematics, 16(4):411–424, November 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417195.

Ernest:1986:CGP

[Ern86] Paul Ernest. Computer gaming for the practice of transformation geometry skills. Educational Studies in Mathematics, 17 (2):205–207, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311523.

Ernest:1999:FKM

[Ern99] Paul Ernest. Forms of knowledge in mathematics and mathematics education: Philosophical and rhetorical perspectives. Educational Studies in Mathematics, 38(1–3):67–83, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003577024357.

Ernest:2006:SPM

[Ern06] Paul Ernest. A semiotic perspective of mathematical activity: The case of number. Educational Studies in Mathematics, 61(1–2):67–101, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
6423-7.

Ernest:2016:PCM

[Ern16a] Paul Ernest. The problem of certainty in mathematics. Educational Studies in Mathematics, 92(3):379–393, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9651-x.

Ernest:2016:UAM

[Ern16b] Paul Ernest. The unit of analysis in mathematics education: bridging the political-technical divide? Educational Studies in Mathematics, 92(1):37–58, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9689-4.

Ernest:2019:HCW

[ERS19] Jessica Brooke Ernest, Daniel L. Reinholz, and Niral Shah. Hidden competence: women’s mathematical participation in public and private classroom spaces. Educational Studies in Mathematics, 102(2):153–172, October 2019. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-019-09910-w.

English:1996:ARD

[ES96] Lyn D. English and Patrick V. Sharry. Analogical reasoning and the development of algebraic abstraction. Educational Studies in Mathematics, 30(2):135–157, March 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302627.

Even:2003:ICI

[ES03] Ruhama Even and Baruch B. Schwarz. Implications of competing interpretations of practice for research and theory in mathematics education. Educational Studies in Mathematics, 54(2–3):283–313, December 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006166.50134.7a.

Eriksson:2021:AFT

[ES21] Helena Eriksson and Lovisa Sumpter. Algebraic and fractional thinking in collective mathematical reasoning. Educational Studies in Mathematics, 108(3):473–491, November 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10044-1.

Even:1995:SMK

[ET95] Ruhama Even and Dina Tirosh. Subject-matter knowledge and knowledge about students as sources of teacher presentations of the subject-matter. Educational Studies in Mathematics, 29(1):1–20, July 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273897.

Evans:2014:MIA

[ETC14] Jeff Evans, Anna Tsatsaroni, and Barbara Czarnecka. Mathematical images in advertising: constructing difference and shaping identity, in global consumer culture. Educational Studies in Mathematics, 85(1):3–27, January 2014. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9496-0.

Evans:1999:BBR

[Eva99] Jeff Evans. Building bridges: Reflections on the problem of transfer of learning in mathematics. Educational Studies in Mathematics, 39(1–3):23–44, June 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003755611058.

Evans:2007:DPI

[Eva07] Siôn Wyn Evans. Differential performance of items in mathematics assessment materials for 7-year-old pupils in Englishmedium and Welsh-medium versions. Educational Studies in Mathematics, 64(2):145–168, February 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-1634-5.

Even:1990:SMK

[Eve90] Ruhama Even. Subject matter knowledge for teaching and the case of functions. Educational Studies in Mathematics, 21(6): 521–544, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00315943.

Even:1999:IAP

[Eve99] Ruhama Even. Integrating academic and practical knowledge in a teacher leaders’ development program. Educational Studies in Mathematics, 38(1–3):235–252, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003665225190.

Espinoza-Vasquez:2025:TTT

[EVHRC^{+}25] Gonzalo Espinoza-Vásquez, Carolina Henríquez-Rivas, Nuria Climent, Rodrigo Ponce, and Paula Verdugo-Hernández. Teaching Thales’s theorem: relations between suitable mathematical working spaces and specialised knowledge. Educational Studies in Mathematics, 118(2):271–293, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10367-9.

Ekdahl:2016:CTS

[EVR16] Anna-Lena Ekdahl, Hamsa Venkat, and Ulla Runesson. Coding teaching for simultaneity and connections. Educational Studies in Mathematics, 93(3):293–313, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9700-0.

Finegold:1976:EDR

[FA76] Menahem Finegold and Shmuel Avital. Enquiry, discovery and research: Terminology and meaning. Educational Studies in Mathematics, 7(4):389–397, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00452221.

Forman:2001:MVM

[FA01] Ellice Forman and Ellen Ansell. The multiple voices of a mathematics classroom community. Educational Studies in Mathematics, 46(1–3):115–142, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014097600732.

Fried:2003:SRM

[FA03] Michael N. Fried and Miriam Amit. Some reflections on mathematics classroom notebooks and their relationship to the public and private nature of student practices. Educational Studies in Mathematics, 53(2):91–112, July 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1025572900956.

Fauskanger:2015:CMT

[Fau15] Janne Fauskanger. Challenges in measuring teachers’ knowledge. Educational Studies in Mathematics, 90(1):57–73, September 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9612-4.

Fischbein:1998:MCS

[FB98] Efraim Fischbein and Madlen Baltsan. The mathematical concept of set and the ‘collection’ model. Educational Studies in Mathematics, 37(1):1–22, October 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003421206945.

Foxman:2002:MCM

[FB02] Derek Foxman and Meindert Beishuizen. Mental calculation methods used by 11-year-olds in different attainment bands: a reanalysis of data from the 1987 APU survey in the UK. Educational Studies in Mathematics, 51(1–2):41–69, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022416021640.

Foote:2011:PEM

[FB11] Mary Q. Foote and Tonya Gau Bartell. Pathways to equity in mathematics education: how life experiences impact researcher positionality. Educational Studies in Mathematics, 78(1):45–68, September 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9309-2.

Frejd:2016:MMP

[FB16] Peter Frejd and Christer Bergsten. Mathematical modelling as a professional task. Educational Studies in Mathematics, 91 (1):11–35, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9654-7.

FitzSimons:2017:WMD

[FB17] Gail E. FitzSimons and Lisa Björklund Boistrup. In the workplace mathematics does not announce itself: towards overcoming the hiatus between mathematics education and work. Educational Studies in Mathematics, 95(3):329–349, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Font:2010:MMC

[FBA10] Vicenç Font, Janete Bolite, and Jorge Acevedo. Metaphors in mathematics classrooms: analyzing the dynamic process of teaching and learning of graph functions. Educational Studies in Mathematics, 75(2):131–152, November 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9247-4.

Fischbein:1971:IPI

[FBM71] E. Fischbein, Ileana Bărbat, and I. Mînzat. Intuitions primaires et intuitions secondaires dans l’initiation aux probabilités. (French) [Primary and secondary intuitions in initiation in probabilities]. Educational Studies in Mathematics, 4(2):264–280, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00303391.

Font:2008:PPR

[FC08] Vicenç Font and Ángel Contreras. The problem of the particular and its relation to the general in mathematics education. Educational Studies in Mathematics, 69(1):33–52, September 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9123-7.

Fukawa-Connelly:2012:CSO

[FC12] Timothy Patrick Fukawa-Connelly. A case study of one instructor’s lecture-based teaching of proof in abstract algebra: making sense of her pedagogical moves. Educational Studies in Mathematics, 81(3):325–345, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9407-9.

Fukawa-Connelly:2014:ATA

[FCN14] Timothy Patrick Fukawa-Connelly and Charlene Newton. Analyzing the teaching of advanced mathematics courses via the enacted example space. Educational Studies in Mathematics, 87(3):323–349, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9554-2.

Favier:2024:HSS

[FD24] Stéphane Favier and Jean-Luc Dorier. Heuristics and semantic spaces for the analysis of students’ work in mathematical problem solving. Educational Studies in Mathematics, 115(3):407– 431, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10297-y.

Fujita:2019:SCD

[FDW19] Taro Fujita, Jonathan Doney, and Rupert Wegerif. Students’ collaborative decision-making processes in defining and classifying quadrilaterals: a semiotic/dialogic approach. Educational Studies in Mathematics, 101(3):341–356, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-019-09892-9;   http://link.springer.
com/content/pdf/10.1007/s10649-019-09892-9.pdf.

Fernandez:2009:ESM

[FE09] Maria Lorelei Fernandez and Evrim Erbilgin. Examining the supervision of mathematics student teachers through analysis of conference communications. Educational Studies in Mathematics, 72(1):93–110, September 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9185-1.

Fehr:1969:CMI

[Feh69] Howard F. Fehr. Contemporary mathematics and its intervention in the sciences. Educational Studies in Mathematics, 1(3): 347–351, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00558321.

Fennema:1979:WGM

[Fen79] Elizabeth Fennema. Women and girls in mathematics — equity in mathematics education. Educational Studies in Mathematics, 10(4):389–401, November 1979. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417086.

Ferrara:2015:BRW

[Fer15] Francesca Ferrara. Book review: When mathematics and mathematics education come to matter: A review of Elizabeth de Freitas and Nathalie Sinclair’s Mathematics and the Body: Material Entanglements in the Classroom (2014). Educational Studies in Mathematics, 90(2):215–220, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9619-x.

Fey:1978:CME

[Fey78] James T. Fey. Change in mathematics education since the late 1950’s — ideas and realisation U.S.A. Educational Studies in Mathematics, 9(3):339–353, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241036.

Fey:1989:TME

[Fey89] James T. Fey. Technology and mathematics education: A survey of recent developments and important problems. Educational Studies in Mathematics, 20(3):237–272, August 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310873.

Ferrara:2017:AAP

[FF17] Francesca Ferrara and Giulia Ferrari. Agency and assemblage in pattern generalisation: a materialist approach to learning. Educational Studies in Mathematics, 94(1):21–36, January 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9708-5.

Freiman:2021:CGM

[FF21] Viktor Freiman and Olga O. Fellus. Closing the gap on the map: Davydov’s contribution to current early algebra discourse in light of the 1960s Soviet debates over word-problem solving. Educational Studies in Mathematics, 106(3):343–361, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09989-6.

Ferrara:2025:TSD

[FF25] Francesca Ferrara and Giulia Ferrari. Telling the story of a diagram: affective and aesthetic mathematical experiences. Educational Studies in Mathematics, 119(2):269–285, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10383-9.

Falk:1980:PLP

[FFL80] Ruma Falk, Raphael Falk, and Iris Levin. A potential for learning probability in young children. Educational Studies in Mathematics, 11(2):181–204, May 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304355.

Fischbein:1984:DTP

[FG84] E. Fischbein and A. Gazit. Does the teaching of probability improve probabilistic intuitions? Educational Studies in Mathematics, 15(1):1–24, February 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00380436.

Fischbein:1997:SIC

[FG97] Efraim Fischbein and Aline Grossman. Schemata and intuitions in combinatorial reasoning. Educational Studies in Mathematics, 34(1):27–47, October 1997. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002914202652.

Fiallo:2017:ACU

[FG17] Jorge Fiallo and Angel Gutiérrez. Analysis of the cognitive unity or rupture between conjecture and proof when learning to prove on a grade 10 trigonometry course. Educational Studies in Mathematics, 96(2):145–167, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9755-6.

Fortuny:1994:IAM

[FGA94] Josep M. Fortuny, Joaquim Giménez, and Claudi Alsina. Integrated assessment on mathematics 12–16. Educational Studies in Mathematics, 27(4):401–412, December 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273380.

Font:2013:EOM

[FGG13] Vicenç Font, Juan D. Godino, and Jesús Gallardo. The emergence of objects from mathematical practices. Educational Studies in Mathematics, 82(1):97–124, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9411-0.

Fletcher:1988:BR

[FH88] T. J. Fletcher and K. M. Hart. Book review. Educational Studies in Mathematics, 19(2):269–275, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00751236.

Fischbein:1990:BR

[FH90] E. Fischbein and Gila Hanna. Book reviews. Educational Studies in Mathematics, 21(5):479–489, October 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00398866.

Foster:2017:TAA

[FI17] Colin Foster and Matthew Inglis. Teachers’ appraisals of adjectives relating to mathematics tasks. Educational Studies in Mathematics, 95(3):283–301, July 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-017-9750-y.pdf.

Fidelman:1985:HBS

[Fid85] Uri Fidelman. Hemispheric basis for schools in mathematics. Educational Studies in Mathematics, 16(1):59–74, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00354883.

Fielker:1979:STG

[Fie79] David S. Fielker. Strategies for teaching geometry to younger children. Educational Studies in Mathematics, 10(1):85–133, February 1979. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311177.

Fielker:1987:CTR

[Fie87] David S. Fielker. A calculator, a tape recorder, and thou. Educational Studies in Mathematics, 18(4):417–437, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00240988.

Finesilver:2022:BCD

[Fin22] Carla Finesilver. Beyond categories: dynamic qualitative analysis of visuospatial representation in arithmetic. Educational Studies in Mathematics, 110(2):271–290, June 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10123-3.

Fischbein:1969:EMD

[Fis69] E. Fischbein. Enseignement mathématique et développement intellectuel. (French) [Mathematical education and intellectual development]. Educational Studies in Mathematics, 2(2–3): 290–306, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303465.

Fischbein:1977:ICL

[Fis77] E. Fischbein. Image and concept in learning mathematics. Educational Studies in Mathematics, 8(2):153–165, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241022.

Fischbein:1993:TFC

[Fis93] Efraim Fischbein. The theory of figural concepts. Educational Studies in Mathematics, 24(2):139–162, February 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273689.

Fischbein:1999:ISM

[Fis99] Efraim Fischbein. Intuitions and schemata in mathematical reasoning. Educational Studies in Mathematics, 38(1–3):11– 50, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003488222875.

Fischbein:2001:TMI

[Fis01] Efraim Fischbein. Tacit models and infinity. Educational Studies in Mathematics, 48(2–3):309–329, November 2001.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016088708705.

FitzSimons:2007:BR

[Fit07] Gail E. FitzSimons. Book review. Educational Studies in Mathematics, 65(3):385–392, July 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9073-x.

FitzSimons:2009:CBR

[Fit09] Gail FitzSimons. Call for book reviewers. Educational Studies in Mathematics, 71(2):95, June 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10649-009-
9194-0.pdf.

FitzSimons:2011:BRM

[Fit11] Gail FitzSimons. Book review: Mathematics education where people really matter. Brian Greer, Swapna Mukhopadhyay, Arthur B. Powell & Sharon Nelson–Barber (Eds.) (2009). Culturally responsive mathematics education. Educational Studies in Mathematics, 76(2):243–246, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9285-y.

FitzSimons:2014:CVM

[Fit14] Gail E. FitzSimons. Commentary on vocational mathematics education: where mathematics education confronts the realities of people’s work. Educational Studies in Mathematics, 86 (2):291–305, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9556-0.

FitzSimons:2015:ESM

[Fit15a] Gail FitzSimons. Educational studies in mathematics book review: Editor’s comment. Educational Studies in Mathematics, 90(2):213, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-015-9622-2.pdf.

FitzSimons:2015:BRM

[Fit15b] Gail E. FitzSimons. Book review: Mathematics & mathematics education: searching for common ground. Fried, M. N., & Dreyfus, T. (Eds.) (2014). Educational Studies in Mathematics, 90(1):95–103, September 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9614-2.

FitzSimons:2017:BRE

[Fit17] Gail FitzSimons. Book review Editor’s comment. Educational Studies in Mathematics, 96(3):391, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-017-9783-2.pdf;   http://link.springer.
com/article/10.1007/s10649-017-9783-2.

Freudenthal:1986:BR

[FJA86] Hans Freudenthal, Konrad Jacobs, and Paulo Abrantes. Book reviews. Educational Studies in Mathematics, 17(3):323–333, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305077.

Fischbein:1995:CIN

[FJC95] Efraim Fischbein, Ruth Jehiam, and Dorit Cohen. The concept of irrational numbers in high-school students and prospective teachers. Educational Studies in Mathematics, 29(1):29– 44, July 1995. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01273899.

Fernandes:2017:CLB

[FKC17] Anthony Fernandes, Leslie H. Kahn, and Marta Civil. A closer look at bilingual students’ use of multimodality in the context of an area comparison problem from a large-scale assessment. Educational Studies in Mathematics, 95(3):263–282, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Flores:2015:MRI

[FKIA15] Raymond Flores, Esther Koontz, Fethi A. Inan, and Mara Alagic. Multiple representation instruction first versus traditional algorithmic instruction first: Impact in middle school mathematics classrooms. Educational Studies in Mathematics, 89(2):267–281, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9597-z.

Freudenthal:1982:BR

[FKSS82] Hans Freudenthal, N. Knoche, Ian Stewart, and Frank J. Swetz. Book reviews. Educational Studies in Mathematics, 13(4):443–448, November 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00366622.

Fletcher:1968:HAM

[Fle68] T. J. Fletcher. A heuristic approach to matrices. Educational Studies in Mathematics, 1(1–2):166–180, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426242.

Fletcher:1971:TGP

[Fle71] T. J. Fletcher. The teaching of geometry present problems and future aims. Educational Studies in Mathematics, 3(3– 4):395–412, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302304.

Fletcher:1976:NFG

[Fle76] T. J. Fletcher. Nomograms and the foundations of geometry. Educational Studies in Mathematics, 7(1–2):71–82, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144360.

Falcade:2007:AFC

[FLM07] Rossana Falcade, Colette Laborde, and Maria Alessandra Mariotti. Approaching functions: Cabri tools as instruments of semiotic mediation. Educational Studies in Mathematics, 66(3):317–333, November 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9072-y.

Forgasz:2014:PVG

[FLT14] Helen Forgasz, Gilah Leder, and Hazel Tan. Public views on the gendering of mathematics and related careers: international comparisons. Educational Studies in Mathematics, 87(3):369–388, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9550-6.

Fluckiger:2005:MSN

[Flü05] Annick Flückiger. Macro-situation and numerical knowledge building: The role of pupils’ didactic memory in classroom interactions. Educational Studies in Mathematics, 59(1–3):59– 84, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-5885-3.

Furinghetti:2009:EUP

[FM09] Fulvia Furinghetti and Francesca Morselli. Every unsuccessful problem solver is unsuccessful in his or her own way: affective and cognitive factors in proving. Educational Studies in Mathematics, 70(1):71–90, January 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9134-4.

Furinghetti:2014:RCM

[FM14] Fulvia Furinghetti and Marta Menghini. The role of concrete materials in Emma Castelnuovo’s view of mathematics teaching. Educational Studies in Mathematics, 87(1):1–6, September 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9569-8.

Fyfe:2020:CDM

[FMA20] Emily R. Fyfe, Percival G. Matthews, and Eric Amsel. College developmental math students’ knowledge of the equal sign. Educational Studies in Mathematics, 104(1):65–85, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09947-2.

Fischbein:1991:FAP

[FNM91] Efraim Fischbein, Maria Sainati Nello, and Maria Sciolis Marino. Factors affecting probabilistic judgements in children and adolescents. Educational Studies in Mathematics, 22(6): 523–549, December 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00312714.

Fort:1969:EEM

[For69] Jacques Fort. Expérimentation d’un enseignement des mathématiques en classe de sixième des lycées et collèges. (French) [Experimentation in teaching of mathematics in sixth grade of high schools and colleges]. Educational Studies in Mathematics, 1(3):262–273, January 1969. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00558312.

Forgasz:1995:GRB

[For95] Helen J. Forgasz. Gender and the relationship between affective beliefs and perceptions of grade 7 mathematics classroom learning environments. Educational Studies in Mathematics, 28(3):219–239, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274174.

Forster:2000:KTS

[For00] Patricia A. Forster. Katie thought she couldn’t do it but now she knows she can. Educational Studies in Mathematics, 43(3): 225–242, November 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1011903213430.

Foster:2016:CCM

[Fos16] Colin Foster. Confidence and competence with mathematical procedures. Educational Studies in Mathematics, 91(2): 271–288, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-015-9660-9.pdf.

Foster:2018:DMF

[Fos18] Colin Foster. Developing mathematical fluency: comparing exercises and rich tasks. Educational Studies in Mathematics, 97(2):121–141, February 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-017-9788-
x;   https://link.springer.com/content/pdf/10.1007/
s10649-017-9788-x.pdf.

Fennema:1985:ALB

[FP85] Elizabeth Fennema and Penelope L. Peterson. Autonomous learning behavior: A possible explanation of sex-related differences in mathematics. Educational Studies in Mathematics, 16(3):309–311, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00776738.

Farmaki:2007:EGM

[FP07] Vassiliki Farmaki and Theodorus Paschos. Employing genetic ‘moments’ in the history of mathematics in classroom activities. Educational Studies in Mathematics, 66(1):83–106, September 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9056-y.

Fennema:1990:TAB

[FPCL90] Elizabeth Fennema, Penelope L. Peterson, Thomas P. Carpenter, and Cheryl A. Lubinski. Teachers’ attributions and beliefs about girls, boys, and mathematics. Educational Studies in Mathematics, 21(1):55–69, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311015.

Fischbein:1969:IPE

[FPM69] E. Fischbein, Ileana Pampu, and I. Mînzat. Initiation aux probabilités à l’école élémentaire. (French) [Initiation in probability in elementary school]. Educational Studies in Mathematics, 2(1):16–31, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368983.

Fernandez-Plaza:2016:TCO

[FPS16] José Antonio Fernández-Plaza and Adrian Simpson. Three concepts or one? Students’ understanding of basic limit concepts. Educational Studies in Mathematics, 93(3):315–332, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-016-9707-6.

Fan:2017:DTA

[FQL^{+}17] Lianghuo Fan, Chunxia Qi, Xiaomei Liu, Yi Wang, and Mengwei Lin. Does a transformation approach improve students’ ability in constructing auxiliary lines for solving geometric problems? an intervention-based study with two Chinese classrooms. Educational Studies in Mathematics, 96(2):229–248, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-017-9772-5.

Franks:1996:SSM

[Fra96] Douglas R. Franks. A situational study of the meanings of personal excellence for secondary mathematics teachers and students. Educational Studies in Mathematics, 31(3):295–317, October 1996. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00376324.

Francisco:2013:LCS

[Fra13] John M. Francisco. Learning in collaborative settings: students building on each other’s ideas to promote their mathematical understanding. Educational Studies in Mathematics, 82(3):417–438, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9437-3.

Freudenthal:1968:WTM

[Fre68] Hans Freudenthal. Why to teach mathematics so as to be useful. Educational Studies in Mathematics, 1(1–2):3–8, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426224.

Freudenthal:1969:ADP

[Fre69a] Hans Freudenthal. Allocution du premier congrès international de l’enseignement mathématique, Lyon, 24–31 août 1969. (French) [Address to the First International Congress of Mathematical Education, Lyon, 24–31 August 1969]. Educational Studies in Mathematics, 2(2–3):135–138, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303453.

Freudenthal:1969:BVD

[Fre69b] Hans Freudenthal. Braces and Venn diagrams. Educational Studies in Mathematics, 1(4):408–414, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02396314.

Freudenthal:1969:CIV

[Fre69c] Hans Freudenthal. The concept of integration at the Varna Congress. Educational Studies in Mathematics, 1(3):338–339, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00558318.

Freudenthal:1969:LAC

[Fre69d] Hans Freudenthal. L’intégration après coup ou à la source. (French) [Integration after the fact or at the source]. Educational Studies in Mathematics, 1(3):327–337, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00558317.

Freudenthal:1969:TCC

[Fre69e] Hans Freudenthal. A teachers course colloquium on sets and logic. Educational Studies in Mathematics, 2(1):32–58, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368984.

Freudenthal:1971:GBD

[Fre71] Hans Freudenthal. Geometry between the devil and the deep sea. Educational Studies in Mathematics, 3(3–4):413–435, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302305.

Freudenthal:1972:ELL

[Fre72] H. Freudenthal. The ‘empirical law of large numbers’ or ‘the stability of frequencies’. Educational Studies in Mathematics, 4(4):484–490, December 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00567002.

Freudenthal:1974:CCD

[Fre74a] H. Freudenthal. The crux of course design in probability. Educational Studies in Mathematics, 5(3):261–277, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424549.

Freudenthal:1974:SRT

[Fre74b] H. Freudenthal. Soviet research on teaching algebra at the lower grades of the elementary school. Educational Studies in Mathematics, 5(4):391–412, December 1974. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01420653.

Freudenthal:1975:PAI

[Fre75] Hans Freudenthal. Pupils’ achievements internationally compared — the IEA. Educational Studies in Mathematics, 6 (2):127–186, July 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302542. See com-

ments [Pea76].

Freudenthal:1976:R

[Fre76a] H. Freudenthal. Rejoinder. Educational Studies in Mathematics, 7(4):529–533, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00452229.

Freudenthal:1976:P

[Fre76b] Hans Freudenthal. Preface. Educational Studies in Mathematics, 7(3):189–190, August 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00553917.

Freudenthal:1977:TTE

[Fre77a] H. Freudenthal. Teacher training — an experimental philosophy. Educational Studies in Mathematics, 8(4):369–376, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00310941.

Freudenthal:1977:BEA

[Fre77b] Hans Freudenthal. Bastiaan’s experiments on Archimedes’ principle. Educational Studies in Mathematics, 8(1):3–16, April 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302500.

Freudenthal:1977:C

[Fre77c] Hans Freudenthal. Creativity. Educational Studies in Mathematics, 8(1):1, April 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302499.

Freudenthal:1978:A

[Fre78a] Hans Freudenthal. Acknowledgment. Educational Studies in Mathematics, 9(4):503–504, November 1978. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410591.

Freudenthal:1978:AFC

[Fre78b] Hans Freudenthal. Address to the First Conference of I.G.P.M.E. (International Group for the Psychology of Mathematical Education), at Utrecht 29 August 1977. Educational Studies in Mathematics, 9(1):1–5, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00352188.

Freudenthal:1978:CMEa

[Fre78c] Hans Freudenthal. Change in mathematics education since the late 1950’s — ideas and realisation. Educational Studies in Mathematics, 9(2):143–145, May 1978. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221154.

Freudenthal:1978:CMEb

[Fre78d] Hans Freudenthal. Change in mathematics education since the late 1950’s — ideas and realisation The Netherlands.

Educational Studies in Mathematics, 9(3):261–270, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241029.

Freudenthal:1979:RS

[Fre79a] Hans Freudenthal. Rings and string. Educational Studies in Mathematics, 10(1):67–70, February 1979. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311175.

Freudenthal:1979:WRE

[Fre79b] Hans Freudenthal. Ways to report on empirical research in education. Educational Studies in Mathematics, 10(3):275–303, August 1979. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00314659.

Freudenthal:1981:MPM

[Fre81] Hans Freudenthal. Major problems of mathematics education. Educational Studies in Mathematics, 12(2):133–150, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305618.

Freudenthal:1982:FVP

[Fre82] Hans Freudenthal. Fiabilité, validité et pertinence — critères de la recherche sur l’enseignement de la mathématique. (French) [Reliability, validity, and relevance — criteria for research in mathematics education]. Educational Studies in Mathematics, 13(4):395–408, November 1982. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00366619.

Freudenthal:1993:TTM

[Fre93] Hans Freudenthal. Thoughts on teaching mechanics didactical phenomenology of the concept of force. Educational Studies in Mathematics, 25(1–2):71–87, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274103.

Frejd:2013:MMA

[Fre13] Peter Frejd. Modes of modelling assessment — a literature review. Educational Studies in Mathematics, 84(3):413–438, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9491-5.

Fried:2007:DHM

[Fri07] Michael N. Fried. Didactics and history of mathematics: Knowledge and self-knowledge. Educational Studies in Mathematics, 66(2):203–223, October 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9025-5.

Fried:2011:SYS

[Fri11] Michael N. Fried. Signs for you and signs for me: the double aspect of semiotic perspectives. Educational Studies in Mathematics, 77(2–3):389–397, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9319-0.

Fried:2021:BRE

[Fri21] Michael N. Fried. Book review: École polytechnique and descriptive geometry: STEM before STEM. Évelyne Barbin, Marta Menghini, & Klaus Volkert (Eds.). (2019) Descriptive geometry, the spread of a polytechnic art: the legacy of Gaspard Monge. Educational Studies in Mathematics, 106(2): 313–321, February 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-10013-0.

Frostad:1999:DCU

[Fro99] Per Frostad. Deaf children’s use of cognitive strategies in simple arithmetic problems. Educational Studies in Mathematics, 40(2):129–153, October 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003609532442.

Ferrara:2016:EAA

[FS16] Francesca Ferrara and Nathalie Sinclair. An early algebra approach to pattern generalisation: Actualising the virtual through words, gestures and toilet paper. Educational Studies in Mathematics, 92(1):1–19, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9674-3.

Foyn:2018:CGS

[FSB18] Trine Foyn, Yvette Solomon, and Hans Jørgen Braathe. Clever girls’ stories: the girl they call a nerd. Educational Studies in Mathematics, 98(1):77–93, May 2018. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-017-9801-4;   http://link.springer.com/
content/pdf/10.1007/s10649-017-9801-4.pdf.

Fuadah:2022:BRR

[FSY22] Ulfah Samrotul Fuadah, Yovika Sukma, and Fikri Yanda. Book review: A review of the book by Liping Ma (2020) Knowing and teaching elementary mathematics: teachers’ understanding of fundamental mathematics in China and the United States (3rd ed.). Educational Studies in Mathematics, 110 (2):379–391, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10116-2.

Forster:2000:MPA

[FT00] Patricia A. Forster and Peter C. Taylor. A multipleperspective analysis of learning in the presence of technology. Educational Studies in Mathematics, 42(1):35–59, April 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004053819533.

Forster:2003:ICC

[FT03] Patricia A. Forster and Peter C. Taylor. An investigation of communicative competence in an upper-secondary class where using graphics calculators was routine. Educational Studies in Mathematics, 52(1):57–77, January 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023676010819.

Fischbein:1979:II

[FTH79] E. Fischbein, D. Tirosh, and P. Hess. The intuition of infinity. Educational Studies in Mathematics, 10(1):3–40, February 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311173.

Fischbein:1981:IPM

[FTM81] E. Fischbein, D. Tirosh, and U. Melamed. Is it possible to measure the intuitive acceptance of a mathematical statement? Educational Studies in Mathematics, 12(4):491–512, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00308145.

Furinghetti:2007:TET

[Fur07] Fulvia Furinghetti. Teacher education through the history of mathematics. Educational Studies in Mathematics, 66(2): 131–143, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9070-0.

Furinghetti:2019:BRL

[Fur19] Fulvia Furinghetti. Book review: The long story of the history in mathematics education. Kathleen M. Clark, Tinne Hoff Kjeldsen, Sebastian Schorcht, & Constantinos Tzanakis (Eds.) (2018). Mathematics, education and history. Towards a harmonious partnership. Educational Studies in Mathematics, 100(1):109–116, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9852-1.

Furinghetti:2022:BRC

[Fur22] Fulvia Furinghetti. Book review: A common ground of mathematics and mathematics education for the classroom and for researchers. Erich Christian Wittmann (2021) Connecting mathematics and mathematics education. Collected papers on mathematics education as a design science. Educational Studies in Mathematics, 111(2):367–374, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10143-7.

Fagnant:2013:SRA

[FV13] Annick Fagnant and Joëlle Vlassis. Schematic representations in arithmetical problem solving: Analysis of their impact on grade 4 students. Educational Studies in Mathematics, 84(1):149–168, September 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9476-4.

Frejd:2022:CSI

[FV22] Peter Frejd and Pauline Vos. A commentary on the special issue “Innovations in measuring and fostering mathematical modelling competencies”. Educational Studies in Mathematics, 109(2):455–468, February 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10122-
4.

Fischer:2019:SWC

[FVJL^{+}19] Jean-Paul Fischer, Bruno Vilette, Sophie Joffredo-Lebrun, Mireille Morellato, Céline Le Normand, Calliste Scheibling-Seve, and Jean-François Richard. Should we continue to teach standard written algorithms for the arithmetical operations? The example of subtraction. Educational Studies in Mathematics, 101(1):105–121, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-019-09884-9.

Fauvel:1997:RHM

[FvM97] John Fauvel and Jan van Maanen. The role of the history of mathematics in the teaching and learning of mathematics: Discussion document for an ICMI Study (1997–2000). Educational Studies in Mathematics, 34(3):255–259, December 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003038421040.

Foster:2022:SSC

[FWCW22] Colin Foster, Simon Woodhead, and Alison Clark-Wilson. School students’ confidence when answering diagnostic questions online. Educational Studies in Mathematics, 109(3):491– 521, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10084-7.

Fennema:1985:I

[FWM85] Elizabeth Fennema, Herbert Walberg, and Cora Marrett. Introduction. Educational Studies in Mathematics, 16(3):303– 304, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00776736.

Falk:2012:CCP

[FYAE12] Ruma Falk, Patricia Yudilevich-Assouline, and Adily Elstein. Children’s concept of probability as inferred from their binary choices — revisited. Educational Studies in Mathematics, 81(2):207–233, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9402-1.

Fyhn:2008:CCR

[Fyh08] Anne Birgitte Fyhn. A climbing class’ reinvention of angles. Educational Studies in Mathematics, 67(1):19–35, January 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9087-z.

Fiori:2005:ERE

[FZ05] Carla Fiori and Luciana Zuccheri. An experimental research on error patterns in written subtraction. Educational Studies in Mathematics, 60(3):323–331, November 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-7530-6.

Fan:2007:RPS

[FZ07] Lianghuo Fan and Yan Zhu. Representation of problem-solving procedures: A comparative look at China, Singapore, and US mathematics textbooks. Educational Studies in Mathematics, 66(1):61–75, September 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9069-6.

Giberti:2025:MDC

[GABB25] Chiara Giberti, Ferdinando Arzarello, Silvia Beltramino, and Giorgio Bolondi. Mathematical discussion in classrooms as a technologically-supported activity fostering participation and inclusion. Educational Studies in Mathematics, 118(2): 201–228, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10356-y.

Gainsburg:2011:BRH

[Gai11] Julie Gainsburg. Book review: Hoyles, C., Noss, R., Kent, P., & Bakker. A. (2010). Improving mathematics at work: The need for techno-mathematical literacies. Educational Studies in Mathematics, 76(1):117–122, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9266-1.

Galtier:1973:AG

[Gal73] Jean Galtier. Around a game. Educational Studies in Mathematics, 5(1):65–70, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01425412.

Galbraith:1981:APC

[Gal81] P. L. Galbraith. Aspects of proving: A clinical investigation of process. Educational Studies in Mathematics, 12(1): 1–28, February 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00386043.

Galbraith:1982:MVS

[Gal82] P. L. Galbraith. The mathematical vitality of secondary mathematics graduates and prospective teachers: A comparative study. Educational Studies in Mathematics, 13(1):89–112, February 1982. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305499.

Galbraith:1986:UMS

[Gal86] P. L. Galbraith. The use of mathematical strategies: factors and features affecting performance. Educational Studies in Mathematics, 17(4):413–441, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311328.

Gallardo:2002:ENN

[Gal02] Aurora Gallardo. The extension of the natural-number domain to the integers in the transition from arithmetic to algebra. Educational Studies in Mathematics, 49(2):171–192, February 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016210906658.

Gal:2011:APT

[Gal11] Hagar Gal. From another perspective — training teachers to cope with problematic learning situations in geometry. Educational Studies in Mathematics, 78(2):183–203, November 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9321-6.

Gal:2019:WUC

[Gal19] Hagar Gal. When the use of cognitive conflict is ineffective — problematic learning situations in geometry. Educational Studies in Mathematics, 102(2):239–256, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09904-8.

Ganguli:1992:ESA

[Gan92] Aparna B. Ganguli. The effect on students’ attitudes of the computer as a teaching aid. Educational Studies in Mathematics, 23(6):611–618, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00540063.

Grugeon-Allys:2024:ALM

[GAP24] Brigitte Grugeon-Allys and Julia Pilet. The algebraic learning of middle school students’ evolution during a school year: a statistical large-scale study based on results in mathematics didactics. Educational Studies in Mathematics, 117(3):533–564, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10347-z.

Gargroetzi:2024:YCJ

[Gar24] Emma C. Gargroetzi. “You can’t just check the box”: the mathematics of ethnoracial contortions at a California high school. Educational Studies in Mathematics, 116(3):403–421, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10258-5.

Gates:2006:GBB

[Gat06] Peter Gates. Going beyond belief systems: Exploring a model for the social influence on mathematics teacher beliefs. Educational Studies in Mathematics, 63(3):347–369, November 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9007-z.

Gauthier:1969:EIE

[Gau69] R. Gauthier. Essai d’individualisation de l’enseignement. (French) [Testing the individualization of education]. Educational Studies in Mathematics, 2(2–3):189–200, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303457.

Greer:1988:BR

[GB88] Brian Greer and George Booker. Book reviews. Educational Studies in Mathematics, 19(4):499–510, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578696.

Gulikers:2001:HAS

[GB01] Iris Gulikers and Klaske Blom. ‘A historical angle’, a survey of recent literature on the use and value of history in geometrical education. Educational Studies in Mathematics, 47 (2):223–258, July 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1014539212782.

Groth:2013:MSK

[GB13] Randall E. Groth and Jennifer A. Bergner. Mapping the structure of knowledge for teaching nominal categorical data analysis. Educational Studies in Mathematics, 83(2):247–265, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9452-4.

Grenier-Boley:2014:SIA

[GB14] Nicolas Grenier-Boley. Some issues about the introduction of first concepts in linear algebra during tutorial sessions at the beginning of university. Educational Studies in Mathematics, 87(3):439–461, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9564-0.

Guala:2017:CAM

[GB17] Elda Guala and Paolo Boero. Cultural analysis of mathematical content in teacher education: the case of elementary arithmetic theorems. Educational Studies in Mathematics, 96(2): 207–227, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9767-2.

Grotluschen:2019:BRN

[GB19] Anke Grotlüschen and Klaus Buddeberg. Book review: Numeracy as social practice. global and local perspectives. Keiko Yasukawa, Alan Rogers, Kara Jackson, & Brian V. Street (Eds.) (2018) Numeracy: the new kid on the block? Educational Studies in Mathematics, 101(3):441–448, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09891-w.

Gardee:2023:FAR

[GB23] Aarifah Gardee and Karin Brodie. A framework for analysing the relationships between peer interactions and learners’ mathematical identities: accounting for personal and social identities. Educational Studies in Mathematics, 113(3):443–473, July 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10201-0.

Gresalfi:2012:WDO

[GBC12] Melissa Sommerfeld Gresalfi, Jacqueline Barnes, and Dionne Cross. When does an opportunity become an opportunity? Unpacking classroom practice through the lens of ecological psychology. Educational Studies in Mathematics, 80(1–2):249– 267, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9367-5.

Gellert:2013:TLI

[GBE13] Uwe Gellert, Joaquim Barbé, and Lorena Espinoza. Towards a local integration of theories: codes and praxeologies in the case of computer-based instruction. Educational Studies in Mathematics, 82(2):303–321, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9427-5.

Georgescu-Buzau:1970:IAP

[GBMB70] E. Georgescu-Buzāu, N. Matei, and Gr. Bānescu. The importance of appropriate problems in the teaching of mathematics. Educational Studies in Mathematics, 2(4):430–437, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417625.

Godino:2005:SAC

[GBR05] Juan D. Godino, Carmen Batanero, and Rafael Roa. An onto-semiotic analysis of combinatorial problems and the solving processes by university students. Educational Studies in Mathematics, 60(1):3–36, September 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5893-3.

Gueudet:2022:DEI

[GBR22] Ghislaine Gueudet, Chantal Buteau, and Marisol Santacruz Rodriguez. Development and evolution of instrumented schemes: a case study of learning programming for mathematical investigations. Educational Studies in Mathematics, 110(2):353–377, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10133-1.

Gomez-Blancarte:2023:IUS

[GBTL23] Ana Luisa Gómez-Blancarte and María Guadalupe Tobías-Lara. The integration of undergraduate students’ informal and formal inferential reasoning. Educational Studies in Mathematics, 113(2):251–269, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-022-10205-
w.

Galbraith:1990:BSM

[GC90a] P. L. Galbraith and N. J. Clatworthy. Beyond standard models — meeting the challenge of modelling. Educational Studies in Mathematics, 21(2):137–163, April 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304899.

Galbraith:1990:FSC

[GC90b] Peter L. Galbraith and D. Chant. Factors shaping community attitudes to school mathematics: Implications for future curriculum change. Educational Studies in Mathematics, 21(4): 299–318, August 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00304261.

Gomez-Chacon:2000:AIK

[GC00] Inés María Gómez-Chacón. Affective influences in the knowledge of mathematics. Educational Studies in Mathematics, 43(2):149–168, September 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017518812079.

Groth:2023:MAS

[GC23] Randall E. Groth and Yoojin Choi. A method for assessing students’ interpretations of contextualized data. Educational Studies in Mathematics, 114(1):17–34, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10234-z.

Gonzalez-Calero:2015:IAM

[GCALB15] José Antonio González-Calero, David Arnau, and Belén Laserna-Belenguer. Influence of additive and multiplicative structure and direction of comparison on the reversal error. Educational Studies in Mathematics, 89(1):133–147, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9596-0.

Gibbons:2023:PPA

[GCFS23] Lynsey Gibbons, Suzanne Chapin, Ziv Feldman, and Rachel Starks. Pre-paring for ambitious mathematics teaching: examining the role of video in supporting mathematics teacher educators. Educational Studies in Mathematics, 113(2):331– 351, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10229-w.

Galbraith:1992:TNT

[GCG^{+}92] Peter L. Galbraith, Marjorie C. Carss, Richard D. Grice, Lovie Endean, and Merle C. Warry. Towards numeracy for the third millennium: A study of the future of mathematics and mathematics education. Educational Studies in Mathematics, 23(6): 569–593, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00540061.

Gravemeijer:1999:CPR

[GD99] Koeno Gravemeijer and Michiel Doorman. Context problems in realistic mathematics education: A calculus course as an example. Educational Studies in Mathematics, 39(1–3):111– 129, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003749919816.

Gabel:2017:AFP

[GD17] Mika Gabel and Tommy Dreyfus. Affecting the flow of a proof by creating presence — a case study in number theory. Educational Studies in Mathematics, 96(2):187–205, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-016-9746-z.

Gorgorio:2009:SRM

[GdA09] Núria Gorgorió and Guida de Abreu. Social representations as mediators of practice in mathematics classrooms with immigrant students. Educational Studies in Mathematics, 72(1): 61–76, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9183-3.

Gilboa:2023:MMA

[GDK23] Nava Gilboa, Tommy Dreyfus, and Ivy Kidron. Metamathematical aspects of definitions. Educational Studies in Mathematics, 114(3):461–481, November 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10252-x.

Geiger:2022:SEN

[GDS22] Vince Geiger, Catherine Delzoppo, and Rudolf Straesser. Supporting English non-dominant language authors’ efforts to publish: perspectives from the editors-in-chief of highly recognised journals in Mathematics Education. Educational Studies in Mathematics, 111(3):543–565, November 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10174-0.

Goodson-Espy:1998:RRR

[GE98] Tracy Goodson-Espy. The roles of reification and reflective abstraction in the development of abstract thought: Transitions from arithmetic to algebra. Educational Studies in Mathematics, 36(3):219–245, September 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003473509628.

Gellert:1998:PET

[Gel98] Uwe Gellert. Prospective elementary teachers’ comprehension of mathematics instruction. Educational Studies in Mathematics, 37(1):23–43, October 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026492400537.

Gellert:2004:DMC

[Gel04] Uwe Gellert. Didactic material confronted with the concept of mathematical literacy. Educational Studies in Mathematics, 55(1–3):163–179, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017693.32454.01.

Gellert:2008:RCO

[Gel08] Uwe Gellert. Routines and collective orientations in mathematics teachers’ professional development. Educational Studies in Mathematics, 67(2):93–110, February 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9089-x.

Genzwein:1970:SOF

[Gen70] Ferenc Genzwein. The system and the organization of further training for the mathematics teachers of the secondary schools in Budapest. Educational Studies in Mathematics, 2(4):419– 429, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00417624.

Gerdes:1981:CME

[Ger81] Paulus Gerdes. Changing mathematics education in Mozambique. Educational Studies in Mathematics, 12(4):455–477, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00308143.

Gerdes:1984:FMO

[Ger84] Paulus Gerdes. The first mathematics olympiads in Mozambique. Educational Studies in Mathematics, 15(2):149–172, May 1984. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305894.

Gerdes:1988:CGT

[Ger88a] Paulus Gerdes. On culture, geometrical thinking and mathematics education. Educational Studies in Mathematics, 19 (2):137–162, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00751229.

Gerdes:1988:PUT

[Ger88b] Paulus Gerdes. On possible uses of traditional Angolan sand drawings in the mathematics classroom. Educational Studies in Mathematics, 19(1):3–22, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00428382.

Geraniou:2010:TSP

[Ger10] Eirini Geraniou. The transitional stages in the PhD degree in mathematics in terms of students’ motivation. Educational Studies in Mathematics, 73(3):281–296, April 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9205-1.

Goodchild:2013:CAI

[GFJ13] Simon Goodchild, Anne Berit Fuglestad, and Barbara Jaworski. Critical alignment in inquiry-based practice in developing mathematics teaching. Educational Studies in Mathematics, 84(3):393–412, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9489-z.

Godino:2011:WLE

[GFWL11] Juan D. Godino, Vicenç Font, Miguel R. Wilhelmi, and Orlando Lurduy. Why is the learning of elementary arithmetic concepts difficult? Semiotic tools for understanding the nature of mathematical objects. Educational Studies in Mathematics, 77(2–3):247–265, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9278-x.

Goos:1996:DIW

[GG96] Merrilyn Goos and Peter Galbraith. Do it this way! Metacognitive strategies in collaborative mathematical problem solving. Educational Studies in Mathematics, 30(3):229–260, April 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304567.

Gal:2022:WEV

[GG22a] Iddo Gal and Vince Geiger. Welcome to the era of vague news: a study of the demands of statistical and mathematical products in the COVID-19 pandemic media. Educational Studies in Mathematics, 111(1):5–28, September 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10151-7.

Gotz:2022:RGS

[GG22b] Daniela Götz and Hedwig Gasteiger. Reflecting geometrical shapes: approaches of primary students to reflection tasks and relations to typical error patterns. Educational Studies in Mathematics, 111(1):47–71, September 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10145-5.

Geiger:2022:DTD

[GGD22] Vince Geiger, Peter Galbraith, and Catherine Delzoppo. Developing a task design and implementation framework for fostering mathematical modelling competencies. Educational Studies in Mathematics, 109(2):313–336, February 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10039-y.

Goos:2002:SMM

[GGR02] Merrilyn Goos, Peter Galbraith, and Peter Renshaw. Socially mediated metacognition: creating collaborative zones of proximal development in small group problem solving. Educational Studies in Mathematics, 49(2):193–223, February 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016209010120.

Galbraith:1998:DNA

[GH98] Peter Galbraith and Chris Haines. Disentangling the nexus: Attitudes to mathematics and technology in a computer learning environment. Educational Studies in Mathematics, 36(3):275–290, September 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003198120666.

Giles:1971:SBL

[Gil71] Geoffrey Giles. Switch boards and logic. Educational Studies in Mathematics, 3(2):184–191, April 1971. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305447.

Geraniou:2019:TDM

[GJ19] Eirini Geraniou and Uffe Thomas Jankvist. Towards a definition of “mathematical digital competency”. Educational Studies in Mathematics, 102(1):29–45, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-019-09893-8;   http://link.springer.
com/content/pdf/10.1007/s10649-019-09893-8.pdf.

Grabovskij:1971:UKM

[GK71] M. A. Grabovskij and P. M. Kotel’nikov. The use of kinematic models in the study of trigonometric functions. Educational Studies in Mathematics, 3(2):147–160, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305443.

Gnedenko:1979:MEE

[GK79] B. V. Gnedenko and Z. Khalil. The mathematical education of engineers. Educational Studies in Mathematics, 10(1): 71–83, February 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311176.

Goos:2020:UPS

[GK20] Merrilyn Goos and Sila Kaya. Understanding and promoting students’ mathematical thinking: a review of research published in ESM. Educational Studies in Mathematics, 103(1): 7–25, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-019-09921-7.

Gerami:2024:CSL

[GKMJ24] Saba Gerami, Eric Khiu, Vilma Mesa, and Thomas Judson. Conceptions of span in linear algebra: from textbook examples to student responses. Educational Studies in Mathematics, 116(1):67–89, May 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10306-8.

Giroux:1998:CKN

[GL98] Jacinthe Giroux and Gisèle Lemoyne. Coordination of knowledge of numeration and arithmetic operations on first grade students. Educational Studies in Mathematics, 35(3):283–301, March 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1002939706074.

Graf:2000:TIS

[GL00] Klaus-D. Graf and Frederick K. S. Leung. The Thirteenth ICMI Study on Mathematics Education in Different Cultural Traditions. Educational Studies in Mathematics, 43(1):95– 116, July 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017929002002.

Gal:2010:SSA

[GL10] Hagar Gal and Liora Linchevski. To see or not to see: analyzing difficulties in geometry from the perspective of visual perception. Educational Studies in Mathematics, 74(2):163– 183, June 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9232-y.

Glaymann:1968:IAC

[Gla68] Maurice Glaymann. Initiation au calcul numérique et usage des machines à calculer. (French) [Introduction to numerical calculation and use of calculating machines]. Educational Studies in Mathematics, 1(1–2):118–125, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426239.

Glaymann:1969:IVS

[Gla69] Maurice Glaymann. Initiation to vector spaces. Educational Studies in Mathematics, 2(1):69–79, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368986.

Glaymann:1970:GC

[Gla70] Maurice Glaymann. A geometry on the cube. Educational Studies in Mathematics, 3(1):84–94, September 1970.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381596.

Glaymann:1976:PEE

[Gla76a] Maurice Glaymann. Les probabilités à l’école élémentaire. (French) [Probabilities in elementary school]. Educational Studies in Mathematics, 6(4):389–393, March 1976. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00411087.

Glaymann:1976:OPE

[Gla76b] Maurice Glaymann. Ou le premier n’est pas toujours premier . . .. (French) [Or the first is not always first …]. Educational Studies in Mathematics, 7(1–2):83–88, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144361.

Guo:2020:ADM

[GLH20] Meng Guo, Frederick K. S. Leung, and Xiang Hu. Affective determinants of mathematical problem posing: the case of Chinese Miao students. Educational Studies in Mathematics, 105 (3):367–387, November 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-09972-1.

Gildehaus:2024:IBP

[GLHM24] Lara Gildehaus, Michael Liebendörfer, and Einat Heyd-Metzuyanim. “I’m a bit out of place here.” — Preservice teachers’ positioning in the figured world of university mathematics. Educational Studies in Mathematics, 116(2):237–255, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10315-7.

Gliner:1991:FCS

[Gli91] Gail S. Gliner. Factors contributing to success in mathematical estimation in preservice teachers: Types of problems and previous mathematical experience. Educational Studies in Mathematics, 22(6):595–606, December 1991. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312717.

Garfield:2015:DSR

[GLZBZ15] Joan Garfield, Laura Le, Andrew Zieffler, and Dani Ben-Zvi. Developing students’ reasoning about samples and sampling variability as a path to expert statistical thinking. Educational Studies in Mathematics, 88(3):327–342, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9541-7.

Griffiths:1984:CAUa

[GM84a] H. B. Griffiths and R. R. McLone. A critical analysis of university examinations in mathematics. Educational Studies in Mathematics, 15(3):291–311, August 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312079.

Griffiths:1984:CAUb

[GM84b] H. B. Griffiths and R. R. McLone. A critical analysis of university examinations in mathematics. Educational Studies in Mathematics, 15(4):423–442, November 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311116.

Goldenberg:2008:SLE

[GM08] Paul Goldenberg and John Mason. Shedding light on and with example spaces. Educational Studies in Mathematics, 69(2): 183–194, October 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9143-3.

Greer:2010:BRL

[GM10] George Brian Greer and Swapna Mukhopadhyay. Book review: The language of mathematics: Telling mathematical tales. Bill Barton. 2008. Educational Studies in Mathematics, 73(2):211– 215, March 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-009-9219-8.

Gerasimova:2023:CPT

[GMH23] Daria Gerasimova, Angela D. Miller, and Margret A. Hjalmarson. Conceptual and procedural teaching: does one teaching approach moderate the relationship between the other teaching approach and algebra achievement? Educational Studies in Mathematics, 114(2):181–198, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10219-y.

Gresalfi:2009:CCA

[GMHG09] Melissa Gresalfi, Taylor Martin, Victoria Hand, and James Greeno. Constructing competence: an analysis of student participation in the activity systems of mathematics classrooms. Educational Studies in Mathematics, 70(1):49–70, January 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9141-5.

Gonzalez-Martin:2018:RDS

[GMNB18] Alejandro S. González-Martín, Elena Nardi, and Irene Biza. From resource to document: scaffolding content and organising student learning in teachers’ documentation work on the teaching of series. Educational Studies in Mathematics, 98 (3):231–252, July 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9813-8.

Goffree:1977:JTT

[Gof77] F. Goffree. Johan — a teacher training freshman studying mathematics and didactics. Educational Studies in Mathematics, 8(2):117–152, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00241021.

Goffree:1993:HWM

[Gof93] Fred Goffree. HF: Working on mathematics education. Educational Studies in Mathematics, 25(1–2):21–49, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274100.

Goldin:2003:DCU

[Gol03] Gerald A. Goldin. Developing complex understandings: On the relation of mathematics education research to mathematics. Educational Studies in Mathematics, 54(2–3):171– 202, December 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000006180.
20493.3c.

Gooya:2007:MTB

[Goo07] Zahra Gooya. Mathematics teachers’ beliefs about a new reform in high school geometry in Iran. Educational Studies in Mathematics, 65(3):331–347, July 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9055-z.

Goos:2014:E

[Goo14] Merrilyn Goos. Editorial. Educational Studies in Mathematics, 85(1):1–2, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-013-9524-0.pdf.

Goos:2015:E

[Goo15] Merrilyn Goos. Editorial. Educational Studies in Mathematics, 88(1):1–2, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-014-9584-9.pdf.

Goos:2016:E

[Goo16] Merrilyn Goos. Editorial. Educational Studies in Mathematics, 91(1):1–2, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-015-9669-0.pdf.

Goos:2017:E

[Goo17] Merrilyn Goos. Editorial. Educational Studies in Mathematics, 94(1):1–3, January 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-016-9739-y.pdf.

Goos:2018:Ea

[Goo18a] Merrilyn Goos. Editorial. Educational Studies in Mathematics, 97(1):1–3, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10649-017-
9800-5.pdf;   http://link.springer.com/article/10.
1007/s10649-017-9800-5.

Goos:2018:Eb

[Goo18b] Merrilyn Goos. Editorial. Educational Studies in Mathematics, 99(3):241–242, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9857-
9;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9857-9.pdf.

Goodchild:2020:BRP

[Goo20] Simon Goodchild. Book review: The philosophy of mathematics education today. Paul Ernest (Ed.) (2018). Educational Studies in Mathematics, 103(1):109–119, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09919-1.

Gorgorio:1998:EFV

[Gor98] Núria Gorgorió. Exploring the functionality of visual and non-visual strategies in solving rotation problems. Educational Studies in Mathematics, 35(3):207–231, March 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003132603649.

Gosztonyi:2021:HHM

[Gos21] Katalin Gosztonyi. How history of mathematics can help to face a crisis situation: the case of the polemic between Bernoulli and d’Alembert about the smallpox epidemic. Educational Studies in Mathematics, 108(1–2):105–122, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10077-6.

Gagatsis:1990:UGM

[GP90] A. Gagatsis and T. Patronis. Using geometrical models in a process of reflective thinking in learning and teaching mathematics. Educational Studies in Mathematics, 21(1):29–54, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311014.

Gorgorio:2001:TMM

[GP01] Núria Gorgorió and Núria Planas. Teaching mathematics in multilingual classrooms. Educational Studies in Mathematics, 47(1):7–33, May 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017980828943.

Glade:2017:SIS

[GP17] Matthias Glade and Susanne Prediger. Students’ individual schematization pathways — empirical reconstructions for the case of part-of-part determination for fractions. Educational Studies in Mathematics, 94(2):185–203, February 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9716-5.

Gibbs:2022:UMC

[GP22] Antonnette M. Gibbs and Joo Young Park. Unboxing mathematics: creating a culture of modeling as critic. Educational Studies in Mathematics, 110(1):167–192, May 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10119-z.

Gray:1999:KCD

[GPPT99] Eddie Gray, Marcia Pinto, Demetra Pitta, and David Tall. Knowledge construction and diverging thinking in elementary & advanced mathematics. Educational Studies in Mathematics, 38(1–3):111–133, March 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003640204118.

Gunes:2024:SPM

[GPS24] Canan Güneş, Kelly Paton, and Nathalie Sinclair. The sensory politics of mathematics: aestheticizing multiplication. Educational Studies in Mathematics, 117(2):239–261, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10326-4.

Grevholm:2005:DME

[GPW05] Barbro Grevholm, Lars-Erik Persson, and Peter Wall. A dynamic model for education of doctoral students and guidance of supervisors in research groups. Educational Studies in Mathematics, 60(2):173–197, October 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-4497-2.

Graham:1971:TPF

[GR71] Alexander Graham and Graham A. Read. A television programme on finite differences. Educational Studies in Mathematics, 3(2):206–219, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305449.

Goel:1997:E

[GR97] Sudhir K. Goel and Michael S. Robillard. The equation: −2 = (−8)^{1/3} = (−8)^{2/6} = [(−8)^{2}]^{1/6} = 2. Educational Studies in Mathematics, 33(3):319–320, September 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002932719513.

Gray:1975:CMC

[Gra75] J. D. Gray. Criticism in the mathematics class. Educational Studies in Mathematics, 6(1):77–86, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00590025.

Graham:1988:MEA

[Gra88] Beth Graham. Mathematical education and Aboriginal children. Educational Studies in Mathematics, 19(2):119–135, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00751228.

Gray:1991:ADA

[Gra91] Edward M. Gray. An analysis of diverging approaches to simple arithmetic: Preference and its consequences. Educational Studies in Mathematics, 22(6):551–574, December 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312715.

Graeber:1999:FKM

[Gra99] Anna Graeber. Forms of knowing mathematics: What preservice teachers should learn. Educational Studies in Mathematics, 38(1–3):189–208, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003624216201.

Graven:2004:IMT

[Gra04] Mellony Graven. Investigating mathematics teacher learning within an in-service community of practice: The centrality of confidence. Educational Studies in Mathematics, 57(2):177– 211, September 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000049277.
40453.4b.

Grant:2020:BRK

[Gra20] Helena A. Grant. Book review: Keeping mathematics education real and fair: connecting the disconnections. Constantinos Xenofontos (Ed.) (2019) Equity in mathematics education: addressing a changing world. Educational Studies in Mathematics, 104(3):439–444, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09958-
z;   http://link.springer.com/content/pdf/10.1007/
s10649-020-09958-z.pdf.

Greer:2001:UPT

[Gre01] Brian Greer. Understanding probabilistic thinking: The legacy of Efraim Fischbein. Educational Studies in Mathematics, 45 (1–3):15–33, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1013801623755.

Greer:2012:IMT

[Gre12] Brian Greer. Inversion in mathematical thinking and learning. Educational Studies in Mathematics, 79(3):429–438, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9317-2.

Greer:2017:RTA

[Gre17] Brian Greer. Review of “All positive action starts with criticism: Hans Freudenthal and the didactics of mathematics” (2015). Translated from “Elke positieve actie begint met critiek: Hans Freudenthal en de didactiek van de wiskunde” by Sacha la Bastide-van Gemert (Springer, 2006). English translation by Marianne Vincken and William Third, co-ordinated by Arthur Bakker. Educational Studies in Mathematics, 95 (1):113–122, May 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Griffiths:1968:PMT

[Gri68] H. B. Griffiths. Pure mathematicians as teachers of applied mathematicians. Educational Studies in Mathematics, 1(1– 2):18–23, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00426227.

Griffiths:1971:MIM

[Gri71] H. Brian Griffiths. Mathematical insight and mathematical curricula. Educational Studies in Mathematics, 4(2):153–165, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00303381.

Griffiths:1978:SCM

[Gri78] Brian Griffiths. Some comments on Macdonald’s paper. Educational Studies in Mathematics, 9(4):421–427, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410587.

Griffiths:1983:SCM

[Gri83] H. B. Griffiths. Simplification and complexity in mathematics education. Educational Studies in Mathematics, 14(3): 297–317, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00410544.

Griffiths:1987:THA

[Gri87] Rachel Griffiths. A tale of 804.573 horses: Arithmetic teaching in Victoria 1860–1914. Educational Studies in Mathematics, 18 (2):191–207, May 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00314726.

Griffiths:2013:IFT

[Gri13] Martin Griffiths. Intuiting the fundamental theorem of arithmetic. Educational Studies in Mathematics, 82(1):75–96, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9410-1.

Grossel:1977:RDO

[Gro77] J. Grossel. Recherche sur les difficultés occasionnées chez les étudiants par les notations avec indices et par le signe ‘Σ’ en mathématiques. (French) [Research on the difficulties experienced by students in notation with indices and by the sign ‘Σ’ in mathematics]. Educational Studies in Mathematics, 8(3): 333–357, October 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00385929.

Geisler:2023:RBA

[GRR23a] Sebastian Geisler, Stefanie Rach, and Katrin Rolka. The relation between attitudes towards mathematics and dropout from university mathematics — the mediating role of satisfaction and achievement. Educational Studies in Mathematics, 112 (2):359–381, February 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10198-6.

Geisler:2023:DAT

[GRR23b] Sebastian Geisler, Katrin Rolka, and Stefanie Rach. Development of affect at the transition to university mathematics and its relation to dropout — identifying related learning situations and deriving possible support measures. Educational Studies in Mathematics, 113(1):35–56, May 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10200-1.

Goldberg:1972:CAI

[GS72] Adele Goldberg and Patrick Suppes. A computer-assisted instruction program for exercises on finding axioms. Educational Studies in Mathematics, 4(4):429–449, December 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00566999.

Goldberg:1976:CAI

[GS76] Adele Goldberg and Patrick Suppes. Computer-assisted instruction in elementary logic at the university level. Educational Studies in Mathematics, 6(4):447–474, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00411092.

Greiffenhagen:2008:SME

[GS08] Christian Greiffenhagen and Wes Sharrock. School mathematics and its everyday other? Revisiting Lave’s ‘cognition in practice’. Educational Studies in Mathematics, 69(1):1– 21, September 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9115-7.

Gvozdic:2018:WIC

[GS18] Katarina Gvozdic and Emmanuel Sander. When intuitive conceptions overshadow pedagogical content knowledge: Teachers’ conceptions of students’ arithmetic word problem solving strategies. Educational Studies in Mathematics, 98(2): 157–175, June 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9806-7.

Gonzalez:2020:HTR

[GS20] María S. García González and Gustavo Martínez Sierra. The history of a teacher’s relief of his mathematics anxiety: the case of Diego. Educational Studies in Mathematics, 103(3): 273–291, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-020-09941-8.

Garcia:2007:LTR

[GSE07] Mercedes García, Victoria Sánchez, and Isabel Escudero. Learning through reflection in mathematics teacher education. Educational Studies in Mathematics, 64(1):1–17, January 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9021-9.

Gunnarsson:2016:DIH

[GSH16a] Robert Gunnarsson, Wang Wei Sönnerhed, and Bernt Hernell. Does it help to use mathematically superfluous brackets when teaching the rules for the order of operations? Educational Studies in Mathematics, 92(1):91–105, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9667-2. See er-

ratum [GSH16b].

Gunnarsson:2016:EDI

[GSH16b] Robert Gunnarsson, Wang Wei Sönnerhed, and Bernt Hernell. Erratum to: Does it help to use mathematically superfluous brackets when teaching the rules for the order of operations? Educational Studies in Mathematics, 92(1):107, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-016-9685-8.pdf. See [GSH16a].

Greefrath:2022:PSS

[GSW22] Gilbert Greefrath, Hans-Stefan Siller, and Raphael Wess. Preservice secondary teachers’ pedagogical content knowledge for the teaching of mathematical modelling. Educational Studies in Mathematics, 109(2):383–407, February 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10038-z.

Graeber:1990:IFF

[GT90] Anna O. Graeber and Dina Tirosh. Insights fourth and fifth graders bring to multiplication and division with decimals. Educational Studies in Mathematics, 21(6):565–588, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00315945.

Gras:1995:CEN

[GT95] Regis Gras and Andre Totohasina. Conceptions d’élèves sur la notion de probabilité conditionnelle révélées par une méthode d’analyse des données: Implication — similarité — corrélation. (French) [Student concepts on the idea of conditional probability revealed by a method of data analysis: Implication — similarity — correlation]. Educational Studies in Mathematics, 28 (4):337–363, June 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274078.

Graham:2000:BVU

[GT00] Alan T. Graham and Michael O. J. Thomas. Building a versatile understanding of algebraic variables with a graphic calculator. Educational Studies in Mathematics, 41(3):265–282, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1004094013054.

Gueudet:2009:TND

[GT09] Ghislaine Gueudet and Luc Trouche. Towards new documentation systems for mathematics teachers? Educational Studies in Mathematics, 71(3):199–218, July 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9159-8.

Gucler:2013:EDL

[Güç13] Beste Güçler. Examining the discourse on the limit concept in a beginning-level calculus classroom. Educational Studies in Mathematics, 82(3):439–453, March 2013. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9438-2.

Gucler:2016:MIM

[Güç16] Beste Güçler. Making implicit metalevel rules of the discourse on function explicit topics of reflection in the classroom to foster student learning. Educational Studies in Mathematics, 91 (3):375–393, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9636-9.

Gueudet:2008:IST

[Gue08] Ghislaine Gueudet. Investigating the secondary–tertiary transition. Educational Studies in Mathematics, 67(3):237–254, March 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-007-9100-6.

Gueudet:2019:BRC

[Gue19] Ghislaine Gueudet. Book review: Curricular resources and their classroom implementation: Contributions (and limitations) of a focus on mathematical tasks. Gabriel J. Stylianides (Ed.) (2016) Curricular resources and classroom use. The case of mathematics. Educational Studies in Mathematics, 100(1): 101–108, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9842-3.

Gueudet:2023:NIA

[Gue23] Ghislaine Gueudet. New insights about the secondary–tertiary transition in mathematics. Educational Studies in Mathematics, 113(1):165–179, May 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-023-10223-
2.

Guillemette:2017:HMS

[Gui17] David Guillemette. History of mathematics in secondary school teachers’ training: towards a nonviolent mathematics education. Educational Studies in Mathematics, 96(3):349– 365, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9774-3.

Gunawardena:1978:CME

[Gun78] A. J. Gunawardena. Changes in mathematics education since the late 1950’s — ideas and realisation Sri Lanka. Educational Studies in Mathematics, 9(3):303–316, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241033.

Guttenberger:1992:SHS

[Gut92] Elfriede Wenzelburger Guttenberger. In-service high school mathematics teacher training in Mexico. Educational Studies in Mathematics, 23(3):299–306, June 1992. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02309534.

Gutstein:2007:MLU

[Gut07] Eric Gutstein. Multiple language use and mathematics: Politicizing the discussion. Educational Studies in Mathematics, 64(2):243–246, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9063-z.

Glaeser:1983:BR

[GV83] G. Glaeser and Joop Van Dormolen. Book review. Educational Studies in Mathematics, 14(2):213–215, May 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303686.

Hewitt:2023:LEK

[HA23] Dave Hewitt and Amal Hussain Alajmi. Learning from English and Kuwaiti children’s transcoding errors: how might number names be temporarily adapted to assist learning of place value? Educational Studies in Mathematics, 114(1):149–172, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10231-2.

Hadar:1977:CCR

[Had77] Nitsa Hadar. Children’s conditional reasoning. Educational Studies in Mathematics, 8(4):413–438, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310946.

Hadar:1978:CCRb

[Had78] N. Hadar. Children’s conditional reasoning. Part III: A design for research on children’s learning of conditional reasoning and research findings. Educational Studies in Mathematics, 9(1):

115–140, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00352196.

Hahn:1975:MSG

[Hah75] H. P. Hahn. A model of the symmetric group S_{4} used as a means of teaching fundamental algebraic concepts. Educational Studies in Mathematics, 6(1):99–111, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00590028.

Hahn:1999:PPC

[Hah99] Corinne Hahn. Proportionnalité et pourcentage chez des apprentis vendeurs. Réflexion sur la relation mathématiques/ réalité dans une formation ‘en alternance’. (French) [Proportionality and percentage among apprentice sellers. Reflection on the mathematical reality relationship in an ‘alternating’ formation]. Educational Studies in Mathematics, 39(1–3):229– 249, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003777202978.

Hahn:2014:LAK

[Hah14] Corinne Hahn. Linking academic knowledge and professional experience in using statistics: a design experiment for business school students. Educational Studies in Mathematics, 86(2): 239–251, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9363-9.

Hammersley:1968:EMS

[Ham68] J. M. Hammersley. On the enfeeblement of mathematical skills by modern mathematics and by similar soft intellectual trash in schools and universities. Educational Studies in Mathematics, 1(1–2):17, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00426226.pdf.

Hamdan:2006:ESS

[Ham06] May Hamdan. Equivalent structures on sets: Equivalence classes, partitions and fiber structures of functions. Educational Studies in Mathematics, 62(2):127–147, June 2006.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-5798-9.

Hanna:1989:MAG

[Han89] Gila Hanna. Mathematics achievement of girls and boys in grade eight: results from twenty countries. Educational Studies in Mathematics, 20(2):225–232, May 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00579464.

Hanna:1995:BR

[Han95] G. Hanna. Book review. Educational Studies in Mathematics, 28(1):87–90, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273858.

Hanna:1998:BRA

[Han98] Gila Hanna. Book review: Akihori Kanamori (Guest editor), Proof and Progress in Mathematics Synthese: An International Journal for Epistemology, Methodology and Philosophy of Science. Educational Studies in Mathematics, 36 (2):197–200, July 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003291502227.

Hanna:2000:PEE

[Han00] Gila Hanna. Proof, explanation and exploration: An overview. Educational Studies in Mathematics, 44(1–3):5–23, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012737223465.

Hannula:2002:ATM

[Han02] Markku S. Hannula. Attitude towards mathematics: emotions, expectations and values. Educational Studies in Mathematics, 49(1):25–46, January 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1016048823497.

Hannula:2006:MMG

[Han06] Markku S. Hannula. Motivation in mathematics: Goals reflected in emotions. Educational Studies in Mathematics, 63 (2):165–178, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
9019-8.

Hanna:2007:BR

[Han07] Gila Hanna. Book review. Educational Studies in Mathematics, 64(1):107–111, January 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9066-9.

Hansson:2010:IRM

[Han10] Åse Hansson. Instructional responsibility in mathematics education: modelling classroom teaching using Swedish data. Educational Studies in Mathematics, 75(2):171–189, November 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-010-9249-2.pdf.

Hand:2012:BRA

[Han12a] Victoria Hand. Book review: Adding depth to portraits of mathematical inquiry. Karin Brodie (2010). Teaching mathematical reasoning in secondary school classrooms. Educational Studies in Mathematics, 79(1):149–155, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9344-z.

Hand:2012:SCP

[Han12b] Victoria Hand. Seeing culture and power in mathematical learning: toward a model of equitable instruction. Educational Studies in Mathematics, 80(1–2):233–247, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9387-9.

Handa:2012:TRR

[Han12c] Yuichi Handa. Teasing out repetition from rote: an essay on two versions of will. Educational Studies in Mathematics, 79(2):263–272, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9343-0.

Hansson:2012:MMI

[Han12d] Åse Hansson. The meaning of mathematics instruction in multilingual classrooms: analyzing the importance of responsibility for learning. Educational Studies in Mathematics, 81(1):103–125, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-012-9385-y.
pdf.

Hannula:2019:YLM

[Han19] Markku S. Hannula. Young learners’ mathematics-related affect: A commentary on concepts, methods, and developmental trends. Educational Studies in Mathematics, 100(3):309–316, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9865-
9;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9865-9.pdf.

Hankeln:2020:MMG

[Han20] Corinna Hankeln. Mathematical modeling in Germany and France: a comparison of students’ modeling processes. Educational Studies in Mathematics, 103(2):209–229, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-019-09931-5;   http://link.springer.
com/content/pdf/10.1007/s10649-019-09931-5.pdf.

Hartnett:1973:CPA

[Har73] William E. Hartnett. The CF/PMM approach to learning mathematics. Educational Studies in Mathematics, 5(1):1– 22, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01425408.

Harkin:1974:TTG

[Har74] J. B. Harkin. Transformation and tesselation on the geoboard. Educational Studies in Mathematics, 5(4):461–465, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01420658.

Harkin:1975:IG

[Har75] J. B. Harkin. Introducing the geoboard. Educational Studies in Mathematics, 6(1):113–118, March 1975. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00590029.

Hart:1981:HME

[Har81] K. Hart. Hierarchies in mathematics education. Educational Studies in Mathematics, 12(2):205–218, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305622.

Harper:1987:GD

[Har87] Eon Harper. Ghosts of Diophantus. Educational Studies in Mathematics, 18(1):75–90, February 1987. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367915.

Hardy:2009:SPI

[Har09a] Nadia Hardy. Students’ perceptions of institutional practices: the case of limits of functions in college level calculus courses. Educational Studies in Mathematics, 72(3):341–358, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9199-8.

Harkness:2009:SCB

[Har09b] Shelly Sheats Harkness. Social constructivism and the believing game: a mathematics teacher’s practice and its implications. Educational Studies in Mathematics, 70(3):243– 258, April 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9151-3.

Haastad:1968:ME

[Hås68] Matts Håstad. Mathematics and engineers. Educational Studies in Mathematics, 1(1–2):93–97, May 1968. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426235.

Hasemann:1981:DF

[Has81] Klaus Hasemann. On difficulties with fractions. Educational Studies in Mathematics, 12(1):71–87, February 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386047.

Hasegawa:1997:CFT

[Has97] Junichi Hasegawa. Concept formation of triangles and quadrilaterals in the second grade. Educational Studies in Mathematics, 32(2):157–179, February 1997. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002923229779.

Haylock:1987:FAM

[Hay87] Derek W. Haylock. A framework for assessing mathematical creativity in school children. Educational Studies in Mathematics, 18(1):59–74, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00367914.

Hazzan:1999:RAL

[Haz99] Orit Hazzan. Reducing abstraction level when learning abstract algebra concepts. Educational Studies in Mathematics, 40(1):71–90, September 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003780613628.

Hirst:1969:UPM

[HB69] Keith Hirst and Norman Biggs. Undergraduate projects in mathematics. Educational Studies in Mathematics, 1(3):252– 261, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00558311.

Harries:2021:ERC

[HB21] Catherine Sara Harries and Julia Hilary Botha. Examining the role of contextual factors in dosage calculation. Educational Studies in Mathematics, 107(3):487–502, July 2021.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10054-z.

Hook:2007:QMC

[HBH07] William Hook, Wayne Bishop, and John Hook. A quality math curriculum in support of effective teaching for elementary schools. Educational Studies in Mathematics, 65(2):125– 148, June 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9050-4.

Hoyles:2005:ISD

[HbL05] Celia Hoyles and Jean baptiste Lagrange. 17th ICMI Study Digital Technologies and Mathematics Teaching and Learning: Rethinking the domain — short announcement. Educational Studies in Mathematics, 60(2):267–268, October 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-4491-8.

Heid:1998:FIT

[HBZE98] M. Kathleen Heid, Glendon W. Blume, Rose Mary Zbiek, and Barbara S. Edwards. Factors that influence teachers learning to do interviews to understand students’ mathematical understandings. Educational Studies in Mathematics, 37(3): 223–249, December 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003657820047.

Hiebert:1982:PTR

[HC82] James Hiebert and Thomas P. Carpenter. Piagetian tasks as readiness measures in mathematics instruction: A critical review. Educational Studies in Mathematics, 13(3):329–345, August 1982. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311247.

Hativa:1995:SLN

[HC95] Nira Hativa and Dorit Cohen. Self learning of negative number concepts by lower division elementary students through solving computer-provided numerical problems. Educational Studies in Mathematics, 28(4):401–431, June 1995.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274081.

Hong:2014:CKA

[HC14] Dae S. Hong and Kyong Mi Choi. A comparison of Korean and American secondary school textbooks: the case of quadratic equations. Educational Studies in Mathematics, 85(2):241– 263, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9512-4.

Harkness:2007:PET

[HDM07] Shelly Sheats Harkness, Beatriz D’Ambrosio, and Anastasia S. Morrone. Preservice elementary teachers’ voices describe how their teacher motivated them to do mathematics. Educational Studies in Mathematics, 65(2):235–254, June 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9045-1.

Head:1981:PLM

[Hea81] John Head. Personality and the learning of mathematics. Educational Studies in Mathematics, 12(3):339–350, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311064.

Hedren:1985:HHC

[Hed85] Rolf Hedren. The hand-held calculator at the intermediate level. Educational Studies in Mathematics, 16(2):163–179, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02400936.

Heege:1978:TML

[Hee78] Hans Ter Heege. Testing the maturity for learning the algorithm of multiplication. Educational Studies in Mathematics, 9(1):75–83, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00352193.

Heege:1985:ABM

[Hee85] Hans Ter Heege. The acquisition of basic multiplication skills. Educational Studies in Mathematics, 16(4):375–388, November 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00417193.

Heitele:1975:EVF

[Hei75] Dietger Heitele. An epistemological view on fundamental stochastic ideas. Educational Studies in Mathematics, 6(2): 187–205, July 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302543.

Hemmi:2010:TSC

[Hem10] Kirsti Hemmi. Three styles characterising mathematicians’ pedagogical perspectives on proof. Educational Studies in Mathematics, 75(3):271–291, December 2010. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9256-3.

Hersh:1993:PCE

[Her93] Reuben Hersh. Proving is convincing and explaining. Educational Studies in Mathematics, 24(4):389–399, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273372.

Herbst:2002:ECP

[Her02a] Patricio G. Herbst. Establishing a custom of proving in American school geometry: evolution of the two-column proof in the early twentieth century. Educational Studies in Mathematics, 49(3):283–312, March 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020264906740.

Herzig:2002:WAS

[Her02b] A. H. Herzig. Where have all the students gone? Participation of doctoral students in authentic mathematical activity as a necessary condition for persistence toward the PH.D.

Educational Studies in Mathematics, 50(2):177–212, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021126424414.

Herbst:2004:A

[Her04] P. Herbst. Announcement. Educational Studies in Mathematics, 57(1):147–148, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3AEDUC.0000047186.
31826.89.

Hersant:2011:CEE

[Her11] Magali Hersant. Correspondance entre élèves: conditions d’une activité mathématique ^{≪} créative ^{≫} et problématisée à la fin du lycée. (French) [Correspondence between pupils: conditions of a ‘creative’ and problematized mathematical activity at the end of the secondary school]. Educational Studies in Mathematics, 78(3):343–370, December 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9327-0.

Herheim:2023:OCU

[Her23] Rune Herheim. On the origin, characteristics, and usefulness of instrumental and relational understanding. Educational Studies in Mathematics, 113(3):389–404, July 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10225-0.

Hess:2013:BRE

[Hes13] Kristin King Hess. Book review: Expertise in mathematics instruction: an international perspective. Yeping Li and Gabriele Kaiser (Eds.) (2011). Educational Studies in Mathematics, 83 (3):487–492, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9467-5.

Herbel-Eisenmann:2010:ALB

[HEW10] Beth Herbel-Eisenmann and David Wagner. Appraising lexical bundles in mathematics classroom discourse: obligation and choice. Educational Studies in Mathematics, 75(1):43– 63, September 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-010-9240-y.

Hewitt:2012:YSL

[Hew12] Dave Hewitt. Young students learning formal algebraic notation and solving linear equations: are commonly experienced difficulties avoidable? Educational Studies in Mathematics, 81(2):139–159, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9394-x.

Herbel-Eisenmann:2010:LBA

[HEWC10] Beth Herbel-Eisenmann, David Wagner, and Viviana Cortes. Lexical bundle analysis in mathematics classroom discourse: the significance of stance. Educational Studies in Mathematics, 75(1):23–42, September 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9253-6.

Herbel-Eisenmann:2015:PME

[HEWJ^{+}15] Beth A. Herbel-Eisenmann, David Wagner, Kate R. Johnson, Heejoo Suh, and Hanna Figueras. Positioning in mathematics education: revelations on an imported theory. Educational Studies in Mathematics, 89(2):185–204, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9588-5.

Hilton:1982:BR

[HFD82] Peter Hilton, Hans Freudenthal, and W. Dörfler. Book review. Educational Studies in Mathematics, 13(2):223–232, May 1982. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00460712.

Hahkioniemi:2023:IBG

[HFL^{+}23] Markus Hähkiöniemi, John Francisco, Antti Lehtinen, Pasi Nieminen, and Salla Pehkonen. The interplay between the guidance from the digital learning environment and the teacher in supporting folding back. Educational Studies in Mathematics, 112(3):461–479, March 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10193-x.

Hitt:2015:CBV

[HGM15] Fernando Hitt and Alejandro S. González-Martín. Covariation between variables in a modelling process: The ACODESA (collaborative learning, scientific debate and self-reflection) method. Educational Studies in Mathematics, 88(2):201–219, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9578-7.

Hadar:1978:CCRa

[HH78] N. Hadar and L. Henkin. Children’s conditional reasoning. Part II: Towards a reliable test of conditional reasoning ability. Educational Studies in Mathematics, 9(1):97–114, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00352195.

Hadar:1981:RSC

[HH81] N. Hadar and R. Hadass. The road to solving a combinatorial problem is strewn with pitfalls. Educational Studies in Mathematics, 12(4):435–443, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00308141.

Hall:1988:GDM

[HH88] Cathy W. Hall and Cynthia Hoff. Gender differences in mathematical performance. Educational Studies in Mathematics, 19(3):395–401, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312455.

Healy:2011:RGM

[HH11] Lulu Healy and Solange Hassan Ahmad Ali Fernandes. The role of gestures in the mathematical practices of those who do not see with their eyes. Educational Studies in Mathematics, 77 (2–3):157–174, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-010-9290-1.

Hunter:2024:CWT

[HH24a] Jodie Hunter and Roberta (Bobbie) Hunter. Correction to: Weaving together the threads of Indigenous knowledge and mathematics. Educational Studies in Mathematics, 117(2):339, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10272-7. See [HH24b].

Hunter:2024:WTT

[HH24b] Jodie Hunter and Roberta (Bobbie) Hunter. Weaving together the threads of Indigenous knowledge and mathematics. Educational Studies in Mathematics, 116(3):501–518, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10256-7. See correction [HH24a].

Huang:2021:ATL

[HHB21] Xingfeng Huang, Rongjin Huang, and Marianna Bosch. Analyzing a teacher’s learning through cross-cultural collaboration: a praxeological perspective of mathematical knowledge for teaching. Educational Studies in Mathematics, 107(3):427– 446, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10057-w.

Hilton:2016:PMS

[HHDG16] Annette Hilton, Geoff Hilton, Shelley Dole, and Merrilyn Goos. Promoting middle school students’ proportional reasoning skills through an ongoing professional development programme for teachers. Educational Studies in Mathematics, 92 (2):193–219, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9694-7.

Hunter:2022:BHS

[HHL22] Jodie Hunter, Roberta Hunter, and Generosa Leach. Bringing the home into school: learning and connecting through mathematics education during the time of a pandemic. Educational Studies in Mathematics, 111(2):207–224, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10157-1.

Hadas:2000:RCU

[HHS00] Nurit Hadas, Rina Hershkowitz, and Baruch B. Schwarz. The role of contradiction and uncertainty in promoting the need to prove in dynamic geometry environments. Educational Studies in Mathematics, 44(1–3):127–150, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012781005718.

Huang:2023:TLA

[HHT23] Xingfeng Huang, Rongjin Huang, and Luc Trouche. Teachers’ learning from addressing the challenges of online teaching in a time of pandemic: a case in Shanghai. Educational Studies in Mathematics, 112(1):103–121, January 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10172-2.

Haines:1994:AMC

[HI94] Chris Haines and John Izard. Assessing mathematical communications about projects and investigations. Educational Studies in Mathematics, 27(4):373–386, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273378.

Hicks:2024:IJT

[Hic24] Michael D. Hicks. “I’ll just try to mimic that”: an exploration of students’ analogical structure creation in abstract algebra. Educational Studies in Mathematics, 117(2):303–327, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10345-1.

Hiebert:1988:TDC

[Hie88] James Hiebert. A theory of developing competence with written mathematical symbols. Educational Studies in Mathematics, 19(3):333–355, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312451.

Hilton:1971:THS

[Hil71] Peter Hilton. Topology in the high school. Educational Studies in Mathematics, 3(3–4):436–453, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302306.

Hino:2015:CMS

[Hin15] Keiko Hino. Comparing multiple solutions in the structured problem solving: Deconstructing Japanese lessons from learner’s perspective. Educational Studies in Mathematics, 90(2):121–141, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9626-y.

Hirst:1972:DT

[Hir72] K. E. Hirst. Derivatives and tangents. Educational Studies in Mathematics, 4(3):299–305, April 1972. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302579.

Hirst:1981:UIM

[Hir81] K. E. Hirst. Undergraduate investigations in mathematics. Educational Studies in Mathematics, 12(3):373–387, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311067.

Hanna:1993:GE

[HJ93a] Gila Hanna and H. Niels Jahnke. Guest editorial. Educational Studies in Mathematics, 24(4):329–331, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273368.

Hanna:1993:PA

[HJ93b] Gila Hanna and Hans Niels Jahnke. Proof and application. Educational Studies in Mathematics, 24(4):421–438, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273374.

Hershkowitz:2012:BRD

[HJ12] Rina Hershkowitz and Barbara Jaworski. Book review: a dialog in the footsteps of the book “A journey in mathematics education research — insights from the work of Paul Cobb; Erna Yackel, Koeno Gravemeijer and Anna Sfard (Eds.); (2011). Educational Studies in Mathematics, 81(3):407–420, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9406-x.

Hawkins:1984:CCP

[HK84] Anne S. Hawkins and Ramesh Kapadia. Children’s conceptions of probability — a psychological and pedagogical review. Educational Studies in Mathematics, 15(4):349–377, November 1984. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311112.

Houdement:1999:ECC

[HK99] Catherine Houdement and Alain Kuzniak. Un exemple de cadre conceptuel pour létude de l’enseignement de la géométrie en formation des maîtres. (French) [An example of a conceptual framework for the study of teaching of geometry in teacher training]. Educational Studies in Mathematics, 40(3): 283–312, November 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003851228212.

Hoyles:2002:SUL

[HK02] Celia Hoyles and Dietmar Küchemann. Students’ understandings of logical implication. Educational Studies in Mathematics, 51(3):193–223, November 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023629608614.

Hahn:2025:CEM

[HK25] Larissa Hahn and Pascal Klein. Clustering eye-movement data uncovers students’ strategies for coordinating equations and diagrams of vector fields. Educational Studies in Mathematics, 118(3):359–385, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10243-y.

Hillel:1989:SSG

[HKG89] Joel Hillel, Carolyn Kieran, and Jean-Luc Gurtner. Solving structured geometric tasks on the computer: The role of feedback in generating strategies. Educational Studies in Mathematics, 20(1):1–39, February 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00356039.

Haataja:2025:SVA

[HKSS^{+}25] Eeva S. H. Haataja, Anniina Koskinen-Salmia, Visajaani Salonen, Miika Toivanen, and Markku S. Hannula. Student visual attention during group instruction phases in collaborative geometry problem solving. Educational Studies in Mathematics, 118(3):387–407, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10337-1.

Herscovics:1994:CGB

[HL94] Nicolas Herscovics and Liora Linchevski. A cognitive gap between arithmetic and algebra. Educational Studies in Mathematics, 27(1):59–78, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01284528.

Hackenberg:2016:SDR

[HL16] Amy J. Hackenberg and Mi Yeon Lee. Students’ distributive reasoning with fractions and unknowns. Educational Studies in Mathematics, 93(2):245–263, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9704-9.

Hasemann:1995:CMR

[HM95] Klaus Hasemann and Helen Mansfield. Concept mapping in research on mathematical knowledge development: Background, methods, findings and conclusions. Educational Studies in Mathematics, 29(1):45–72, July 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273900.

Heyd-Metzuyanim:2013:CCL

[HM13] Einat Heyd-Metzuyanim. The co-construction of learning difficulties in mathematics — teacher–student interactions and their role in the development of a disabled mathematical identity. Educational Studies in Mathematics, 83(3):341–368, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9457-z.

Heyd-Metzuyanim:2025:FTL

[HM25] Einat Heyd-Metzuyanim. Failure to teach/learn mathematics: a complexity-discursive perspective. Educational Studies in Mathematics, 119(3):515–533, July 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10404-1.

Hegedus:2011:EMS

[HMA11] Stephen John Hegedus and Luis Moreno-Armella. The emergence of mathematical structures. Educational Studies in Mathematics, 77(2–3):369–388, July 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9297-7.

Hossain:2013:TUM

[HMA13] Sarmin Hossain, Heather Mendick, and Jill Adler. Troubling “understanding mathematics in-depth”: Its role in the identity work of student-teachers in England. Educational Studies in Mathematics, 84(1):35–48, September 2013. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-013-9474-6.pdf.

Heyd-Metzuyanim:2025:CRR

[HMCT25a] Einat Heyd-Metzuyanim, Barak Cohen, and Michal Tabach. Correction to: Regulative routines and their relation to the process of de-ritualization in mathematics learning. Educational Studies in Mathematics, 120(3):549–550, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10448-3. See [HMCT25b].

Heyd-Metzuyanim:2025:RRT

[HMCT25b] Einat Heyd-Metzuyanim, Barak Cohen, and Michal Tabach. Regulative routines and their relation to the process of deritualization in mathematics learning. Educational Studies in Mathematics, 120(3):525–548, November 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.

1007/s10649-025-10401-4. See correction [HMCT25a].

Heyd-Metzuyanim:2016:BPP

[HMG16] Einat Heyd-Metzuyanim and Mellony Graven. Between people-pleasing and mathematizing: South African learners’ struggle for numeracy. Educational Studies in Mathematics, 91 (3):349–373, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9637-8.

Heyd-Metzuyanim:2019:REM

[HMG19] Einat Heyd-Metzuyanim and Mellony Graven. Rituals and explorations in mathematical teaching and learning: introduction to the special issue. Educational Studies in Mathematics, 101(2):141–151, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-019-09890-x. See

[HMG20].

Heyd-Metzuyanim:2020:CRE

[HMG20] Einat Heyd-Metzuyanim and Mellony Graven. Correction to: Rituals and explorations in mathematical teaching and learning: introduction to the special issue. Educational Studies in Mathematics, 103(1):137, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09916-
4;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09916-4.pdf. See [HMG19].

Heyd-Metzuyanim:2018:CFC

[HMMG18] Einat Heyd-Metzuyanim, Charles Munter, and James Greeno. Conflicting frames: a case of misalignment between professional development efforts and a teacher’s practice in a high school mathematics classroom. Educational Studies in Mathematics, 97(1):21–37, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-017-9777-0.

Heyd-Metzuyanim:2020:APD

[HMNWB20] Einat Heyd-Metzuyanim, Talli Nachlieli, Merav Weingarden, and Rinat Baor. Adapting a professional development program for cognitively demanding instruction across shifting contexts. Educational Studies in Mathematics, 104(3):385– 403, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09967-y.

Hernandez-Martinez:2023:CBM

[HMRRTT23] Paul Hernandez-Martinez, Svitlana Rogovchenko, Yuriy Rogovchenko, and Stephanie Treffert-Thomas. Collaboration between mathematicians and mathematics educators: dialogical inquiry as a methodological tool in mathematics education research. Educational Studies in Mathematics, 114(1):129–148, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10245-w.

Heyd-Metzuyanim:2019:REP

[HMSBR19] Einat Heyd-Metzuyanim, Margaret Smith, Victoria Bill, and Lauren B. Resnick. From ritual to explorative participation in discourse-rich instructional practices: a case study of teacher learning through professional development. Educational Studies in Mathematics, 101(2):273–289, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9849-9.

Heyd-Metzuyanim:2021:MML

[HMSBT21] Einat Heyd-Metzuyanim, Aviv J. Sharon, and Ayelet Baram-Tsabari. Mathematical media literacy in the COVID-19 pandemic and its relation to school mathematics education. Educational Studies in Mathematics, 108(1–2):201–225, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10075-8.

Hussain:2013:TSD

[HMT13] Mohammed Abdul Hussain, John Monaghan, and John Threlfall. Teacher-student development in mathematics classrooms: Interrelated zones of free movement and promoted actions. Educational Studies in Mathematics, 82(2):285–302, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9426-6.

Hoyles:1992:PMM

[HN92] Celia Hoyles and Richard Noss. A pedagogy for mathematical microworlds. Educational Studies in Mathematics, 23(1): 31–57, February 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302313.

Hoyles:1995:SMO

[HN95] Celia Hoyles and Marilyn Nickson. Stieg Mellin–Olsen. Educational Studies in Mathematics, 28(4):335–336, June 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274077.

Hershkovitz:1996:RSD

[HN96] Sara Hershkovitz and Pearla Nesher. The role of schemes in designing computerized environments. Educational Studies in Mathematics, 30(4):339–366, June 1996. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00570829.

Hoyles:2008:NSI

[HN08] Celia Hoyles and Richard Noss. Next steps in implementing Kaput’s research programme. Educational Studies in Mathematics, 68(2):85–97, June 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-007-
9102-4.

Hegedus:2023:MSE

[HO23] Stephen J. Hegedus and Yenny Otálora. Mathematical strategies and emergence of socially mediated metacognition within a multi-touch Dynamic Geometry Environment. Educational Studies in Mathematics, 112(2):289–307, February 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10170-4.

Hodgson:1996:SAV

[Hod96] Ted Hodgson. Students’ ability to visualize set expressions: An initial investigation. Educational Studies in Mathematics, 30 (2):159–178, March 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302628.

Hennessy:1989:ITS

[HOEF89] Sara Hennessy, Tim O’Shea, Rick Evertsz, and Ann Floyd. An intelligent tutoring system approach to teaching primary mathematics. Educational Studies in Mathematics, 20(3): 273–292, August 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00310874.

Hoffman:1977:TLB

[Hof77] Nathan Hoffman. The teaching and learning of basic number facts. Educational Studies in Mathematics, 8(4):391–398, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00310943.

Hoffmann:2006:WSP

[Hof06] Michael H. G. Hoffmann. What is a “semiotic perspective”, and what could it be? Some comments on the contributions to this special issue. Educational Studies in Mathematics, 61(1–2):279–291, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
1456-5.

Hojgaard:2024:CFS

[Høj24] Tomas Højgaard. Competencies and fighting syllabusism. Educational Studies in Mathematics, 115(3):459–479, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10285-2.

Hollinger:1969:IGF

[Hol69] A. Hollinger. Sur l’initiation à la géométrie. (French) [On initiation to geometry]. Educational Studies in Mathematics, 1 (4):386–393, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02396311.

Holden:1973:MD

[Hol73] Lyman Holden. The march of the discoverer. Educational Studies in Mathematics, 5(2):193–205, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01421026.

Holcombe:1981:RAM

[Hol81] M. Holcombe. The role of automata and machine theory in school and college mathematics syllabuses. Educational Studies in Mathematics, 12(2):219–234, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305623.

Hope:1985:UME

[Hop85] Jack A. Hope. Unravelling the mysteries of expert mental calculation. Educational Studies in Mathematics, 16(4):355– 374, November 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00417192.

Hart:2013:JIS

[HOS13] Lynn C. Hart, Susan Oesterle, and Susan L. Swars. The juxtaposition of instructor and student perspectives on mathematics courses for elementary teachers. Educational Studies in Mathematics, 83(3):429–451, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9464-0.

Howson:1973:CGRa

[How73a] A. G. Howson. Charles Godfrey (1873–1924) and the reform of mathematical education. Educational Studies in Mathematics, 5(1):157–180, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00684695.

Howson:1973:CGRb

[How73b] A. G. Howson. Charles Godfrey (1873–1924) and the reform of mathematical education. Educational Studies in Mathematics, 5(2):157–180, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01421023.

Howson:1975:UCF

[How75] A. G. Howson. University courses for future teachers. Educational Studies in Mathematics, 6(3):273–292, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01793612.

Howson:1978:CME

[How78] A. G. Howson. Change in mathematics education since the late 1950’s — ideas and realisation Great Britain. Educational Studies in Mathematics, 9(2):183–223, May 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221158.

Howson:1980:SME

[How80] A. G. Howson. Socialist mathematics education: Does it exist? Educational Studies in Mathematics, 11(3):285–299, August 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00697741.

Howson:1984:SFY

[How84] A. G. Howson. Seventy five years of the International Commission on Mathematical Instruction. Educational Studies in Mathematics, 15(1):75–93, February 1984. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00380440.

Hoyles:1982:PVM

[Hoy82] Celia Hoyles. The pupil’s view of mathematics learning. Educational Studies in Mathematics, 13(4):349–372, November 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00366617.

Hoyles:1985:WPG

[Hoy85] Celia Hoyles. What is the point of group discussion in mathematics? Educational Studies in Mathematics, 16(2):205– 214, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02400938.

Hoyles:2001:DDM

[Hoy01] Celia Hoyles. From describing to designing mathematical activity: the next step in developing a social approach to research in mathematics education? Educational Studies in Mathematics, 46(1–3):273–286, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014019612105.

Hoyrup:2007:RMB

[Høy07] Jens Høyrup. The roles of Mesopotamian bronze age mathematics tool for state formation and administration — carrier of teachers’ professional intellectual autonomy. Educational Studies in Mathematics, 66(2):257–271, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9090-4.

Hoz:1981:ERS

[Hoz81] Ron Hoz. The effects of rigidity on school geometry learning. Educational Studies in Mathematics, 12(2):171–190, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305620.

Housman:2003:PSL

[HP03] David Housman and Mary Porter. Proof schemes and learning strategies of above-average mathematics students. Educational Studies in Mathematics, 53(2):139–158, July 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1025541416693.

Hegedus:2008:SNF

[HP08] Stephen J. Hegedus and William R. Penuel. Studying new forms of participation and identity in mathematics classrooms with integrated communication and representational infrastructures. Educational Studies in Mathematics, 68(2):171– 183, June 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9120-x.

Herbert:2012:REC

[HP12] Sandra Herbert and Robyn Pierce. Revealing educationally critical aspects of rate. Educational Studies in Mathematics, 81(1):85–101, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9368-4.

Hankeln:2025:LEA

[HP25] Corinna Hankeln and Susanne Prediger. Language is essential for avoiding surface translations: Associations of students’ spontaneous use of meaning-related phrases for explicating structures with conceptual understanding of multiplication. Educational Studies in Mathematics, 120(1):57–79, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10414-z.

Hersant:2005:COT

[HPG05] Magali Hersant and Marie-Jeanne Perrin-Glorian. Characterization of an ordinary teaching practice with the help of the theory of didactic situations. Educational Studies in Mathematics, 59(1–3):113–151, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
2183-z.

Hochmuth:2025:MME

[HPRW25] Reinhard Hochmuth, Jana Peters, Frode Rønning, and Carl Winsløw. Modelling mathematics for educational research and practice: a comparison of two theoretical approaches. Educational Studies in Mathematics, 118(2):153–168, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10368-8.

Heck:2019:SIL

[HPS^{+}19] Daniel J. Heck, Courtney L. Plumley, Despina A. Stylianou, Adrienne A. Smith, and Gwendolyn Moffett. Scaling up innovative learning in mathematics: exploring the effect of different professional development approaches on teacher knowledge, beliefs, and instructional practice. Educational Studies in Mathematics, 102(3):319–342, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09895-6.

Harkin:1974:SPP

[HR74] Joseph B. Harkin and Gerald R. Rising. Some psychological and pedagogical aspects of mathematical symbolism. Educational Studies in Mathematics, 5(3):255–260, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424548.

Huckstep:2000:CMR

[HR00] Peter Huckstep and Tim Rowland. ‘creative mathematics’ — real or rhetoric? Educational Studies in Mathematics, 42 (1):81–100, April 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004085404106.

Haghjoo:2023:AWD

[HRR23a] Saeid Haghjoo, Farzad Radmehr, and Ebrahim Reyhani. Analyzing the written discourse in calculus textbooks over 42 years: the case of primary objects, concrete discursive objects, and a realization tree of the derivative at a point. Educational Studies in Mathematics, 112(1):73–102, January 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10168-y. See correction [HRR23b].

Haghjoo:2023:CAW

[HRR23b] Saeid Haghjoo, Farzad Radmehr, and Ebrahim Reyhani. Correction to: Analyzing the written discourse in calculus textbooks over 42 years: the case of primary objects, concrete discursive objects, and a realization tree of the derivative at a point. Educational Studies in Mathematics, 113(2):369, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10226-z. See corrigendum [HRR23a].

Hall:2025:TTS

[HRW25] Jennifer Hall, Travis Robinson, and Jane Wilkinson. A tale of two students: unpacking undergraduate mathematics students’ mathematical identities using the Listening Guide. Educational Studies in Mathematics, 120(3):635–656, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10425-w.

Hershkowitz:1999:EPR

[HS99] Rina Hershkowitz and Baruch Schwarz. The emergent perspective in rich learning environments: Some roles of tools and activities in the construction of sociomathematical norms. Educational Studies in Mathematics, 39(1–3):149–166, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003769126987.

Hanna:2002:SE

[HS02] Gila Hanna and Nathan Sidoli. The story of ESM. Educational Studies in Mathematics, 50(2):123–156, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021162617070.

Harkness:2013:PUH

[HS13a] Shelly Sheats Harkness and James Stallworth. Photovoice: understanding high school females’ conceptions of mathematics and learning mathematics. Educational Studies in Mathematics, 84(3):329–347, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9485-3.

Heng:2013:BNM

[HS13b] Mary Anne Heng and Akhila Sudarshan. “Bigger number means you plus!” — teachers learning to use clinical interviews to understand students’ mathematical thinking. Educational Studies in Mathematics, 83(3):471–485, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9469-3.

Healy:2014:TUS

[HS14] Lulu Healy and Ana Isabel Sacristán. Towards an understanding of the shaping of research outcomes by contextual issues: reflections on the contributions of the ReMath project. Educational Studies in Mathematics, 85(3):423–435, March 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9538-2.

Hackenberg:2022:BSM

[HS22] Amy J. Hackenberg and Serife Sevinc. A boundary of the second multiplicative concept: the case of Milo. Educational Studies in Mathematics, 109(1):177–193, January 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10083-8.

Herbst:2022:HCT

[HSM22] Patricio Herbst, Mollee Shultz, and Amanda Milewski. How can teaching simulations help us study at scale the tensions mathematics teachers have to manage when considering policy recommendations? Educational Studies in Mathematics, 110(1):1–21, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10118-0.

Harper:2015:BRR

[HSNHHE15] Frances K. Harper, Heejoo Suh, Julie Nurnberger-Haag, and Beth Herbel-Eisenmann. Book review: A review of the book by Elizabeth de Freitas and Nathalie Sinclair (2014). Mathematics and the body: Material entanglements in the classroom. Educational Studies in Mathematics, 90(2):221–230, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9621-3.

Hitt:2017:RCA

[HSZ17] Fernando Hitt, Mireille Saboya, and Carlos Cortés Zavala. Rupture or continuity: The arithmetico–algebraic thinking as an alternative in a modelling process in a paper and pencil and technology environment. Educational Studies in Mathematics, 94(1):97–116, January 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9717-4.

Hall:2018:UPI

[HTM18] Jennifer Hall, Jo Towers, and Lyndon C. Martin. Using I poems to illuminate the complexity of students’ mathematical identities. Educational Studies in Mathematics, 99(2): 181–196, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9839-y.

Herfort:2023:TYR

[HTM^{+}23] Jonas Dreyøe Herfort, Andreas Lindenskov Tamborg, Florian Meier, Benjamin Brink Allsopp, and Morten Misfeldt. Twenty years of research on technology in mathematics education at CERME: a literature review based on a data science approach. Educational Studies in Mathematics, 112(2):309–336, February 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10202-z.

Hudson:1990:GPM

[Hud90] Brian Hudson. Global perspectives in the mathematics classroom. Educational Studies in Mathematics, 21(2):129–136, April 1990. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00304898.

Hudson:1994:A

[Hud94] Brian Hudson. Announcement. Educational Studies in Mathematics, 27(1):113–114, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01284532.

Hughes:1974:HTM

[Hug74] Barnabas B. Hughes. Heuristic teaching in mathematics. Educational Studies in Mathematics, 5(3):291–299, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424551.

Hunting:1986:RSC

[Hun86] Robert P. Hunting. Rachel’s schemes for constructing fraction knowledge. Educational Studies in Mathematics, 17(1): 49–66, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302378.

Hussein:1987:PTH

[Hus87] Mansour G. A. Hussein. Private tutoring a hidden educational problem. Educational Studies in Mathematics, 18(1): 91–96, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00367916.

Hussein:2025:BRM

[Hus25] Hisham Barakat Bishr Hussein. Book review: M. A. (Ken) Clements, Berinderjeet Kaur, Thomas Lowrie, Vilma Mesa, and Johan Prytz (Eds.) (2024) Fourth international handbook of mathematics education. Educational Studies in Mathematics, 119(3):557–564, July 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-025-10388-
y.

Halmos:1978:CME

[HV78] Mária Halmos and Tamás Varga. Change in mathematics education since the late 1950’s — ideas and realisation Hungary. Educational Studies in Mathematics, 9(2):225–244, May 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221159.

Harju:2024:SFN

[HVN^{+}24] Heidi Harju, Jo Van Hoof, Cristina E. Nanu, Jake McMullen, and Minna Hannula-Sormunen. Spontaneous focusing on numerical order and numerical skills of 3- to 4-year-old children. Educational Studies in Mathematics, 117(1):43–65, September 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10327-3.

Headrick:2020:EAP

[HWJ20] Lorna Headrick, Adi Wiezel, and Amanda Jansen. Engagement and affect patterns in high school mathematics classrooms that exhibit spontaneous problem posing: an exploratory framework and study. Educational Studies in Mathematics, 105(3):435–456, November 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09996-7.

Hasa:2023:USA

[HWR23] Jokke Hásá, Lín Westlin, and Johanna Rámó. Undergraduate students’ attitudes towards mathematical proving in an introduction to proof course. Educational Studies in Mathematics, 114(3):393–415, November 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10239-8.

Haj-Yahya:2022:UTE

[HY22] Aehsan Haj-Yahya. Using theoretical and empirical background information to affect noticing of geometrical thinking. Educational Studies in Mathematics, 111(3):493–513, November 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10176-y.

Haeck:1997:DAM

[HYC^{+}97] Wim Haeck, Nan Yeld, Jurie Conradie, Neill Robertson, and Adrienne Shall. A developmental approach to mathematics testing for university admissions and course placement. Educational Studies in Mathematics, 33(1):71–91, June 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002974021690.

Hylemon:2024:BRH

[Hyl24] Jennifer T. H. Hylemon. Book review: Historical influences, present problems, and futuristic solutions in African countries. Brantina Chirinda, Kakoma Luneta, and Alphonse Uworabayeho (Eds.). (2022). Mathematics education in Africa: the fourth industrial revolution. Educational Studies in Mathematics, 115(1):111–123, January 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10249-6.

Hazzan:2005:RAC

[HZ05] Orit Hazzan and Rina Zazkis. Reducing abstraction: The case of school mathematics. Educational Studies in Mathematics, 58(1):101–119, January 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
3335-x.

Ingram:2019:WSO

[IAP19] Jenni Ingram, Nick Andrews, and Andrea Pitt. When students offer explanations without the teacher explicitly asking them to. Educational Studies in Mathematics, 101(1):51–66, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9873-9;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9873-9.pdf.

Irwin:2005:ANS

[IB05] Kathryn C. Irwin and Murray S. Britt. The algebraic nature of students’ numerical manipulation in the New Zealand Numeracy Project. Educational Studies in Mathematics, 58(2): 169–188, February 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-2755-y.

Izsak:2019:DCA

[IB19] Andrew Izsák and Sybilla Beckmann. Developing a coherent approach to multiplication and measurement. Educational Studies in Mathematics, 101(1):83–103, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09885-8.

Iannone:2020:IHS

[ICR20] Paola Iannone, Christoph Czichowsky, and Johannes Ruf. The impact of high stakes oral performance assessment on students’ approaches to learning: a case study. Educational Studies in Mathematics, 103(3):313–337, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-020-09937-4;   http://link.springer.
com/content/pdf/10.1007/s10649-020-09937-4.pdf.

Innabi:2007:CMT

[IE07] Hanan Innabi and Omar El Sheikh. The change in mathematics teachers’ perceptions of critical thinking after 15 years of educational reform in Jordan. Educational Studies in Mathematics, 64(1):45–68, January 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
9017-x.

Irwin:2005:ADN

[II05] Kathryn C. Irwin and R. John Irwin. Assessing development in numeracy of students from different socio-economic areas: A Rasch analysis of three fundamental tasks. Educational Studies in Mathematics, 58(3):283–298, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-6425-x.

Iannone:2011:DGE

[IIMR^{+}11] Paola Iannone, Matthew Inglis, Juan Pablo Mejía-Ramos, Adrian Simpson, and Keith Weber. Does generating examples aid proof production? Educational Studies in Mathematics, 77(1):1–14, May 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9299-0.

Iannone:2019:GNU

[IM19] Paola Iannone and Dominic Miller. Guided notes for university mathematics and their impact on students’ note-taking behaviour. Educational Studies in Mathematics, 101(3):387– 404, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9872-x.

Inglis:2007:MMA

[IMRS07] Matthew Inglis, Juan Pablo Mejia-Ramos, and Adrian Simpson. Modelling mathematical argumentation: the importance of qualification. Educational Studies in Mathematics, 66(1): 3–21, September 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9059-8.

Iori:2017:OSR

[Ior17] Maura Iori. Objects, signs, and representations in the semio-cognitive analysis of the processes involved in teaching and learning mathematics: A duvalian perspective. Educational Studies in Mathematics, 94(3):275–291, March 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9726-3.

Iori:2018:TAS

[Ior18] Maura Iori. Teachers’ awareness of the semio-cognitive dimension of learning mathematics. Educational Studies in Mathematics, 98(1):95–113, May 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9808-5.

Ikram:2025:BRY

[IPJ25] Fadhil Zil Ikram, M. Azhari Panjaitan, and Ariyanti Jalal. Book review: Yiming Cao (ed.) (2024). Students’ collaborative problem solving in mathematics classrooms: an empirical study. Educational Studies in Mathematics, 120(1):195–204, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10397-x.

Inglis:2008:CIA

[IS08] Matthew Inglis and Adrian Simpson. Conditional inference and advanced mathematical study. Educational Studies in Mathematics, 67(3):187–204, March 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9098-9.

Inglis:2009:CIA

[IS09] Matthew Inglis and Adrian Simpson. Conditional inference and advanced mathematical study: further evidence. Educational Studies in Mathematics, 72(2):185–198, November 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9187-z.

Isaacs:1987:UPS

[Isa87] Ian Isaacs. Using problem solving teaching styles to prepare candidates for the CXC basic proficiency examinations. Educational Studies in Mathematics, 18(2):177–190, May 1987. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00314725.

Ji-Yeong:2024:RAA

[ISJ24] Ji-Yeong I, Ji-Won Son, and Hyunyi Jung. Rethinking anti-Asian racism in mathematics education: the hidden racism made visible through AsianCrit. Educational Studies in Mathematics, 116(3):423–443, July 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10335-3.

Isleyen:2023:LE

[İşl23] Tevfik İşleyen. Letter to the Editor. Educational Studies in Mathematics, 114(1):S9, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https:/

/link.springer.com/article/10.1007/s10649-010-9295-
9.

Iowo-Team:1977:EMG

[IT77] Iowo-Team. Eleven minutes group work — a transcript. Educational Studies in Mathematics, 8(4):377–389, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310942.

Ing:2015:SPE

[IWF^{+}15] Marsha Ing, Noreen M. Webb, Megan L. Franke, Angela C. Turrou, Jacqueline Wong, Nami Shin, and Cecilia H. Fernandez. Student participation in elementary mathematics classrooms: the missing link between teacher practices and student achievement? Educational Studies in Mathematics, 90(3):341– 356, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9625-z.

Jablonski:2023:IAA

[Jab23] Simone Jablonski. Is it all about the setting? — A comparison of mathematical modelling with real objects and their representation. Educational Studies in Mathematics, 113(2):307– 330, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10215-2.

Jahnke:1983:TEE

[Jah83] Hans Niels Jahnke. Technology and education: The example of the computer. Educational Studies in Mathematics, 14(1): 87–100, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00704704.

Jahnke:2001:CCO

[Jah01] Hans Niels Jahnke. Cantor’s cardinal and ordinal infinities: an epistemological and didactic view. Educational Studies in Mathematics, 48(2–3):175–197, November 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016052726876.

Jakuba:1971:EIT

[Jak71] E. G. Jakuba. Experiments on improving the teaching process. Educational Studies in Mathematics, 3(2):161–169, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305444.

Jansson:1974:SLV

[Jan74] Lars C. Jansson. Structural and linguistic variables that contribute to difficulty in the judgment of simple verbal deductive arguments. Educational Studies in Mathematics, 5(4): 493–505, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01420660.

Janvier:1981:USM

[Jan81] Claude Janvier. Use of situations in mathematics education. Educational Studies in Mathematics, 12(1):113–122, February 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386049.

Jankvist:2009:CWH

[Jan09a] Uffe Thomas Jankvist. A categorization of the “whys” and “hows” of using history in mathematics education. Educational Studies in Mathematics, 71(3):235–261, July 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9174-9.

Jansen:2009:PET

[Jan09b] Amanda Jansen. Prospective elementary teachers’ motivation to participate in whole-class discussions during mathematics content courses for teachers. Educational Studies in Mathematics, 71(2):145–160, June 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9168-7.

Jankvist:2010:ESU

[Jan10] Uffe Thomas Jankvist. An empirical study of using history as a ‘goal’. Educational Studies in Mathematics, 74(1):53–74, May 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9227-8.

Jaworski:2002:SCU

[Jaw02] Barbara Jaworski. Sensitivity and challenge in university mathematics tutorial teaching. Educational Studies in Mathematics, 51(1–2):71–94, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022491404298.

Jaworski:2003:RPI

[Jaw03] Barbara Jaworski. Research practice into/influencing mathematics teaching and learning development: Towards a theoretical framework based on co-learning partnerships. Educational Studies in Mathematics, 54(2–3):249–282, December 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000006160.91028.f0.

Jablonka:2021:NDS

[JB21] Eva Jablonka and Christer Bergsten. Numbers don’t speak for themselves: strategies of using numbers in public policy discourse. Educational Studies in Mathematics, 108(3):579–596, November 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10059-8.

Jiang:2022:BRS

[JC22] Chunlian Jiang and Jinfa Cai. Book review: Seeing algebra in arithmetic in cross-cultural contexts. Meixia Ding (2021) Teaching early algebra through example-based problem solving: Insights from Chinese and US elementary classrooms. Educational Studies in Mathematics, 111(1):177–184, September 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10132-2.

Jansen:2023:SMT

[JCM^{+}23] Amanda Jansen, Kelly Curtis, Amanda Mohammad Mirzaei, Catherine E. Cullicott, Ethan P. Smith, and James A. Middleton. Secondary mathematics teachers’ descriptions of student engagement. Educational Studies in Mathematics, 113(3):425– 442, July 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10228-x.

Johnson:2024:PTM

[JDK^{+}24] Heather Lynn Johnson, Courtney Donovan, Robert Knurek, Kristin A. Whitmore, and Livvia Bechtold. Proposing and testing a model relating students’ graph selection and graph reasoning for dynamic situations. Educational Studies in Mathematics, 115(3):387–406, March 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10299-4.

Jurdak:2014:TDS

[JE14] Murad Eid Jurdak and Rabih Raif El Mouhayar. Trends in the development of student level of reasoning in pattern generalization tasks across grade level. Educational Studies in Mathematics, 85(1):75–92, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9494-2.

Jeeves:1971:SES

[Jee71] M. A. Jeeves. Some experiments on structured learning. Educational Studies in Mathematics, 3(3–4):454–476, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302307.

Jerman:1970:CMS

[Jer70] Max Jerman. A counting model for simple addition. Educational Studies in Mathematics, 2(4):438–445, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417626.

Jerman:1973:PLSa

[Jer73a] Max Jerman. Problem length as a structural variable in verbal arithmetic problems. Educational Studies in Mathematics, 5 (1):109–123, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00684692.

Jerman:1973:PLSb

[Jer73b] Max Jerman. Problem length as a structural variable in verbal arithmetic problems. Educational Studies in Mathematics, 5 (2):109–123, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01421020.

Jaremus:2020:GSB

[JGF20] Felicia Jaremus, Jennifer Gore, and Leanne Fray. Girls are still being ‘counted out’: teacher expectations of high-level mathematics students. Educational Studies in Mathematics, 105 (2):219–236, October 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-09986-9.

Jones:2000:GE

[JGM00] Keith Jones, Ángel Gutiérrez, and Maria Alessandra Mariotti. Guest editorial. Educational Studies in Mathematics, 44(1– 3):1–3, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1012706005027.

Jorgensen:2014:SET

[JGR14] Robyn Jorgensen, Peter Gates, and Vanessa Roper. Structural exclusion through school mathematics: using Bourdieu to understand mathematics as a social practice. Educational Studies in Mathematics, 87(2):221–239, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9468-4.

Jiang:2014:CSS

[JHC14] Chunlian Jiang, Stephen Hwang, and Jinfa Cai. Chinese and Singaporean sixth-grade students’ strategies for solving problems about speed. Educational Studies in Mathematics, 87(1): 27–50, September 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9559-x.

Jones:2015:PAP

[JI15] Ian Jones and Matthew Inglis. The problem of assessing problem solving: can comparative judgement help? Educational Studies in Mathematics, 89(3):337–355, July 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9607-1.

Jing:2007:PCH

[Jin07] Tang Jing. The popularization of China’s higher education and its influence on university mathematics education. Educational Studies in Mathematics, 66(1):77–82, September 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9081-5.

Jeannotte:2017:CMM

[JK17] Doris Jeannotte and Carolyn Kieran. A conceptual model of mathematical reasoning for school mathematics. Educational Studies in Mathematics, 96(1):1–16, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9761-8.

Jones:1997:FAN

[JLTM97] Graham A. Jones, Cynthia W. Langrall, Carol A. Thornton, and A. Timothy Mogill. A framework for assessing and nurturing young children’s thinking in probability. Educational Studies in Mathematics, 32(2):101–125, February 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002981520728.

Jerman:1974:LCV

[JM74] Max E. Jerman and Sanford Mirman. Linguistic and computational variables in problem solving in elementary mathematics. Educational Studies in Mathematics, 5(3):317–362, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424553.

Johnson:2018:LBS

[JM18] Heather Lynn Johnson and Evan McClintock. A link between students’ discernment of variation in unidirectional change and their use of quantitative variational reasoning. Educational Studies in Mathematics, 97(3):299–316, March 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-017-9799-7;   https://link.springer.
com/content/pdf/10.1007/s10649-017-9799-7.pdf.

Jankvist:2019:WHW

[JMA19] Uffe Thomas Jankvist, Morten Misfeldt, and Mario Sánchez Aguilar. What happens when CAS procedures are objectified? — the case of “solve” and “desolve”. Educational Studies in Mathematics, 101(1):67–81, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09888-5.

Jones:2008:BGH

[JMD08] Keith Jones and Joanna Mamona-Downs. Brian Griffiths (1927–2008) — his pioneering contributions to mathematics and education. Educational Studies in Mathematics, 69(3): 283–286, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9153-1.

Jenssen:2022:PSP

[JMRW22] Lars Jenßen, Regina Möller, and Bettina Roesken-Winter. Pre-service primary teachers’ shame experiences during their schooling time: characteristics and effects on their subjectchoices at university. Educational Studies in Mathematics, 110(3):435–455, July 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10141-1.

Jankvist:2015:FDR

[JN15] Uffe Thomas Jankvist and Mogens Niss. A framework for designing a research-based “maths counsellor” teacher programme. Educational Studies in Mathematics, 90(3):259–284, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9629-8.

Junker:2025:PSG

[JNF25] Astrid Junker, Guri A. Nortvedt, and Danyal Farsani. Patterning strategies in grade 1 students with low and high number sense proficiency. Educational Studies in Mathematics, 118 (1):29–51, January 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10341-5.

Johnson:2015:TYS

[Joh15] Heather L. Johnson. Together yet separate: Students’ associating amounts of change in quantities involved in rate of change. Educational Studies in Mathematics, 89(1):89–110, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9590-y.

Jones:1982:LMS

[Jon82] Peter L. Jones. Learning mathematics in a second language: A problem with more and less. Educational Studies in Mathematics, 13(3):269–287, August 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311245.

Jones:2000:PFD

[Jon00] Keith Jones. Providing a foundation for deductive reasoning: Students’ interpretations when using dynamic geometry software and their evolving mathematical explanations. Educational Studies in Mathematics, 44(1–3):55–85, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012789201736.

Jones:2016:BRA

[Jon16] Ian Jones. Book review: Algebra teaching around the world. Frederick K. S. Leung, Kyungmee Park, Derek Holton, & David Clarke (Eds.) (2014). Educational Studies in Mathematics, 91(2):289–294, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9650-y.

Jones:2018:PIM

[Jon18] Steven R. Jones. Prototype images in mathematics education: the case of the graphical representation of the definite integral. Educational Studies in Mathematics, 97(3):215–234, March 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-017-9794-z.

Joubert:2013:UDT

[Jou13] Marie Joubert. Using digital technologies in mathematics teaching: developing an understanding of the landscape using three “grand challenge” themes. Educational Studies in Mathematics, 82(3):341–359, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9430-x.

Jouannet:2025:BIM

[Jou25] Paula Jouannet. Beyond intuitive models: exploring the role of gestures in learning multiplication and division. Educational Studies in Mathematics, 120(3):477–496, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10432-x.

Jaworski:2009:BMM

[JP09] Barbara Jaworski and Despina Potari. Bridging the macroand micro-divide: using an activity theory model to capture sociocultural complexity in mathematics teaching and its development. Educational Studies in Mathematics, 72(2):219– 236, November 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9190-4.

Jerman:1972:PRD

[JR72] Max Jerman and Raymond Rees. Predicting the relative difficulty of verbal arithmetic problems. Educational Studies in Mathematics, 4(3):306–323, April 1972. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302580.

Jahnke:1987:BR

[JS87] H. N. Jahnke and Ole Skovsmose. Book review. Educational Studies in Mathematics, 18(1):97–105, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367917.

Jurdak:1999:ESC

[JS99] Murad Jurdak and Iman Shahin. An ethnographic study of the computational strategies of a group of young street vendors in Beirut. Educational Studies in Mathematics, 40(2): 155–172, October 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003894908704.

Jurdak:2001:PSA

[JS01] Murad Jurdak and Iman Shahin. Problem solving activity in the workplace and the school: The case of constructing solids. Educational Studies in Mathematics, 47(3):297–315, September 2001. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1015106804646.

Jahnke:1975:ZRP

[JSV75] H. N. Jahnke, H. Steinbring, and D. Vogel. Zahlbegriff und Rechenfertigkeit — zur Problematik der Entwicklung wissenschaftlicher Begriffe. (German) [Numerical concepts and rendering skills — on the problem of the development of scientific terms]. Educational Studies in Mathematics, 6(2):213– 252, July 1975. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302545.

Jones:1994:MNA

[JTP94] Graham A. Jones, Carol A. Thornton, and Ian J. Putt. A model for nurturing and assessing multidigit number sense among first grade children. Educational Studies in Mathematics, 27(2):117–143, September 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01278918.

Jungwirth:1991:IGF

[Jun91] Helga Jungwirth. Interaction and gender — findings of a microethnographical approach to classroom discourse. Educational Studies in Mathematics, 22(3):263–284, June 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368341.

Jurdak:1977:SLV

[Jur77] Murad E. Jurdak. Structural and linguistic variables in selected inference patterns for bilinguals in grades six to ten. Educational Studies in Mathematics, 8(2):225–238, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241027.

Jurdak:2006:CPP

[Jur06] Murad Eid Jurdak. Contrasting perspectives and performance of high school students on problem solving in real world, situated, and school contexts. Educational Studies in Mathematics, 63(3):283–301, November 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9008-y.

Jones:2017:BRA

[JV17] Keith Jones and Chronoula Voutsina. Book review: Approaches to qualitative research in mathematics education: Examples of methodology and methods, edited by Angelika Bikner–Ahsbahs, Christine Knipping, & Norma Presmeg. (2015). Educational Studies in Mathematics, 96(3):381–390, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-017-9784-1.

Johnson:2025:CMU

[JWFC^{+}25] Estrella Johnson, Keith Weber, Timothy Patrick Fukawa-Connelly, Hamidreza Mahmoudian, and Lisa Carbone. Collaborating with mathematicians to use active learning in university mathematics courses: the importance of attending to mathematicians’ obligations. Educational Studies in Mathematics, 119(1):145–161, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-024-10381-
x.

Kaasila:2007:MBK

[Kaa07] Raimo Kaasila. Mathematical biography and key rhetoric. Educational Studies in Mathematics, 66(3):373–384, November 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9085-1.

Kaeley:1988:SDL

[Kae88] G. S. Kaeley. Sex differences in the learning of postsecondary mathematics in a neo-literate society. Educational Studies in Mathematics, 19(4):435–457, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578693.

Kaeley:1990:ISE

[Kae90] Gurcharn S. Kaeley. The influence of socio-economic status, entry style and instructional variables on the learning of mathematics in a neo-literate society. Educational Studies in Mathematics, 21(4):319–350, August 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304262.

Kaeley:1993:EMA

[Kae93] Gurcharn S. Kaeley. Explaining mathematics achievement of mature internal and external students at the university of Papua New Guinea. Educational Studies in Mathematics, 25(3):251–260, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273864.

See erratum [Kae94].

Kaeley:1994:EEM

[Kae94] Gurcharn S. Kaeley. Erratum: Explaining mathematics achievement of mature internal and external students at the University of Papua New Guinea [short communication], Educational Studies in Mathematics 25, pp. 251–260. Educational Studies in Mathematics, 27(1):111, July 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.

1007/BF01284531.pdf. The author’s name was omitted from the article [Kae93].

Kaimanovich:2013:BRB

[Kai13] Vadim Kaimanovich. Book review: A book about Russian mathematics education. Educational Studies in Mathematics, 84(1):169–175, September 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9479-1.

Klein:2010:MTK

[KAJHB10] Pnina S. Klein, Esther Adi-Japha, and Simcha Hakak-Benizri. Mathematical thinking of kindergarten boys and girls: similar achievement, different contributing processes. Educational Studies in Mathematics, 73(3):233–246, April 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9216-y.

Kalomitsines:1983:TMS

[Kal83] Spyros P. Kalomitsines. Two methods for solving problems. Educational Studies in Mathematics, 14(3):251–274, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410541.

Kapur:1970:CAS

[Kap70] J. N. Kapur. Combinatorial analysis and school mathematics. Educational Studies in Mathematics, 3(1):111–127, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381598.

Kapadia:1974:CEP

[Kap74] Ramesh Kapadia. A critical examination of Piaget–Inhelder’s view on topology. Educational Studies in Mathematics, 5(4): 419–424, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01420655.

Kapadia:1976:STL

[Kap76] Ramesh Kapadia. Set-theory and logic in school. Educational Studies in Mathematics, 6(4):409–413, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00411089.

Kapadia:1977:CDP

[Kap77] Ramesh Kapadia. Curriculum development: Practice and theory. Educational Studies in Mathematics, 8(2):199–210, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00241025.

Kapur:1978:CME

[Kap78] J. N. Kapur. Change in mathematical education since the late 1950’s — ideas and realisation India. Educational Studies in Mathematics, 9(2):245–253, May 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221160.

Kapadia:1980:DSE

[Kap80] Ramesh Kapadia. Developments in statistical education. Educational Studies in Mathematics, 11(4):443–461, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00231216.

Kapa:2001:MSD

[Kap01] Esther Kapa. A metacognitive support during the process of problem solving in a computerized environment. Educational Studies in Mathematics, 47(3):317–336, September 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015124013119.

Kapadia:2009:BRM

[Kap09a] Ramesh Kapadia. Book review: mathematics classrooms in twelve countries. Educational Studies in Mathematics, 71(1): 85–93, May 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-009-9182-4.

Kaput:2009:BII

[Kap09b] James Kaput. Building intellectual infrastructure to expose and understand ever-increasing complexity. Educational Studies in Mathematics, 70(2):211–215, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9169-6.

Karsenty:2002:WDA

[Kar02] Ronnie Karsenty. What do adults remember from their high school mathematics? The case of linear functions. Educational Studies in Mathematics, 51(1–2):117–144, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022429504802.

Karsenty:2010:NMT

[Kar10] Ronnie Karsenty. Nonprofessional mathematics tutoring for low-achieving students in secondary schools: A case study. Educational Studies in Mathematics, 74(1):1–21, May 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9223-z.

Kar:2016:ECA

[Kar16] Tuğrul Kar. An error correction about problem posing on a linear graph. Educational Studies in Mathematics, 91(1):7–8, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9640-0.

Kaufman:1971:B

[Kau71] Burt Kaufman. Background. Educational Studies in Mathematics, 3(3–4):277–280, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302296.

Kazima:2007:MSM

[Kaz07] Mercy Kazima. Malawian students’ meanings for probability vocabulary. Educational Studies in Mathematics, 64(2): 169–189, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9032-6.

Katz:2007:SHA

[KB07] Victor J. Katz and Bill Barton. Stages in the history of algebra with implications for teaching. Educational Studies in Mathematics, 66(2):185–201, October 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9023-7.

Kolikant:2011:EUV

[KB11] Yifat Ben-David Kolikant and Orit Broza. The effect of using a video clip presenting a contextual story on low-achieving students’ mathematical discourse. Educational Studies in Mathematics, 76(1):23–47, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9262-5.

Kjeldsen:2012:BMH

[KB12] Tinne Hoff Kjeldsen and Morten Blomhøj. Beyond motivation: history as a method for learning meta-discursive rules in mathematics. Educational Studies in Mathematics, 80(3): 327–349, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9352-z.

Kazunga:2020:AAS

[KB20] Cathrine Kazunga and Sarah Bansilal. An APOS analysis of solving systems of equations using the inverse matrix method. Educational Studies in Mathematics, 103(3):339–358, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09935-6.

Kullberg:2020:ELA

[KBBK20] Angelika Kullberg, Camilla Björklund, Irma Brkovic, and Ulla Runesson Kempe. Effects of learning addition and subtraction in preschool by making the first ten numbers and their relations visible with finger patterns. Educational Studies in Mathematics, 103(2):157–172, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-019-09927-1;   http://link.springer.
com/content/pdf/10.1007/s10649-019-09927-1.pdf.

Kaiser:2017:EPC

[KBK^{+}17a] Gabriele Kaiser, Sigrid Blömeke, Johannes König, Andreas Busse, Martina Döhrmann, and Jessica Hoth. Erratum to: Professional competencies of (prospective) mathematics teachers — cognitive versus situated approaches. Educational Studies in Mathematics, 94(2):183–184, February 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-016-9724-5.pdf. See [KBK^{+}17b].

Kaiser:2017:PCP

[KBK^{+}17b] Gabriele Kaiser, Sigrid Blömeke, Johannes König, Andreas Busse, Martina Döhrmann, and Jessica Hoth. Professional competencies of (prospective) mathematics teachers — cognitive versus situated approaches. Educational Studies in Mathematics, 94(2):161–182, February 2017. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-016-9713-8. See er-

ratum [KBK^{+}17a].

Konig:2021:LBP

[KBK21] Johannes König, Sigrid Blömeke, and Gabriele Kaiser. The links between pedagogical competence, instructional quality, and mathematics achievement in the lower secondary classroom. Educational Studies in Mathematics, 107(1):189–212, May 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-10021-0.

Kullberg:2024:SNR

[KBK24a] Angelika Kullberg, Camilla Björklund, and Ulla Runesson Kempe. Seeing number relations when solving a three-digit subtraction task. Educational Studies in Mathematics, 115 (2):271–287, February 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10287-0.

Kullberg:2024:ILA

[KBK^{+}24b] Angelika Kullberg, Camilla Björklund, Ulla Runesson Kempe, Irma Brkovic, Maria Nord, and Tuula Maunula. Improvements in learning addition and subtraction when using a structural approach in first grade. Educational Studies in Mathematics, 117(3):399–417, November 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10339-z.

Krygovska:1973:QRC

[KC73] A. Z. Krygovska and Maria Ciosek. Quelques remarques sur le comportement des élèves concernant les problèmes mathématiques. (French) [Some remarks on the behavior of pupils concerning mathematical problems]. Educational Studies in Mathematics, 5(1):39–48, April 1973. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01425410.

Kotsopoulos:2009:IIC

[KC09] Donna Kotsopoulos and Michelle Cordy. Investigating imagination as a cognitive space for learning mathematics. Educational Studies in Mathematics, 70(3):259–274, April 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9154-0.

Kolar:2012:AFI

[KČ12] Vida Manfreda Kolar and Tatjana Hodnik Čadež. Analysis of factors influencing the understanding of the concept of infinity. Educational Studies in Mathematics, 80(3):389–412, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9357-7.

Kus:2022:MIS

[KC22] Mehtap Kus and Erdinc Cakiroglu. Mathematics in the informal setting of an art studio: students’ visuospatial thinking processes in a studio thinking-based environment. Educational Studies in Mathematics, 110(3):545–571, July 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10142-8.

Krawitz:2022:RRC

[KCS22] Janina Krawitz, Yu-Ping Chang, and Stanislaw Schukajlow. The role of reading comprehension in mathematical modelling: improving the construction of a real-world model and interest in Germany and Taiwan. Educational Studies in Mathematics, 109(2):337–359, February 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10058-9.

Kidron:2010:JEC

[KD10] Ivy Kidron and Tommy Dreyfus. Justification enlightenment and combining constructions of knowledge. Educational Studies in Mathematics, 74(1):75–93, May 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9228-7.

Kidron:2014:PI

[KD14] Ivy Kidron and Tommy Dreyfus. Proof image. Educational Studies in Mathematics, 87(3):297–321, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9566-y.

Krause:2021:TTC

[KDM21] Christina M. Krause, Pietro Di Martino, and Judit N. Moschkovich. Tales from three countries: reflections during COVID-19 for mathematical education in the future. Educational Studies in Mathematics, 108(1–2):87–104, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10066-9.

Keitel:1982:CVT

[Kei82] Christine Keitel. Curriculum variables, theory and goals: A comment on Begle’s critical variables in mathematics education. Educational Studies in Mathematics, 13(3):257–267, August 1982. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311244.

Keitel:1993:HFR

[Kei93] Christine Keitel. Hans Freudenthal: Revisiting mathematics education China lecturers. Educational Studies in Mathematics, 25(1–2):161–164, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274107.

Kelly:1971:TTH

[Kel71] Paul J. Kelly. Topology and transformations in high school geometry. Educational Studies in Mathematics, 3(3–4):477– 481, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302308.

Kemme:1981:RSA

[Kem81] S. L. Kemme. References of speech acts as characteristics of mathematical classroom conversation. Educational Studies in Mathematics, 12(1):43–58, February 1981. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386045.

Kieran:2001:GEL

[KFS01] Carolyn Kieran, Ellice Forman, and Anna Sfard. Guest editorial: Learning discourse: Sociocultural approaches to research in mathematics education. Educational Studies in Mathematics, 46(1–3):1–12, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014276102421.

Kent:1980:GT

[KH80] David Kent and Keith Hedger. Growing tall. Educational Studies in Mathematics, 11(2):137–179, May 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304354.

Koichu:2007:TIC

[KH07] Boris Koichu and Guershon Harel. Triadic interaction in clinical task-based interviews with mathematics teachers. Educational Studies in Mathematics, 65(3):349–365, July 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9054-0.

Kaasila:2008:SEO

[KHLP08] Raimo Kaasila, Markku S. Hannula, Anu Laine, and Erkki Pehkonen. Socio-emotional orientations and teacher change. Educational Studies in Mathematics, 67(2):111–123, February 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9094-0.

Konold:2015:DST

[KHRK15] Clifford Konold, Traci Higgins, Susan Jo Russell, and Khalimahtul Khalil. Data seen through different lenses. Educational Studies in Mathematics, 88(3):305–325, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9529-8.

Kwon:2021:GCN

[KHY21] Oh Nam Kwon, Chaereen Han, and Gangwon Yoon. Graphs in the COVID-19 news: a mathematics audit of newspapers in Korea. Educational Studies in Mathematics, 108(1– 2):183–200, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10029-0.

Kidron:2008:ACL

[Kid08] Ivy Kidron. Abstraction and consolidation of the limit procept by means of instrumented schemes: the complementary role of three different frameworks. Educational Studies in Mathematics, 69(3):197–216, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9132-6.

Kidron:2011:TMT

[Kid11] Ivy Kidron. Tacit models, treasured intuitions and the discrete — continuous interplay. Educational Studies in Mathematics, 78(1):109–126, September 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9313-6.

Kidron:2016:ENT

[Kid16] Ivy Kidron. Epistemology and networking theories. Educational Studies in Mathematics, 91(2):149–163, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9666-3.

Kinnear:2025:SAG

[KID25] George Kinnear, Paola Iannone, and Ben Davies. Student approaches to generating mathematical examples: comparing e-assessment and paper-based tasks. Educational Studies in Mathematics, 119(2):179–201, June 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10361-1.

Kieran:1981:CAE

[Kie81] Carolyn Kieran. Concepts associated with the equality symbol. Educational Studies in Mathematics, 12(3):317–326, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311062.

Kieran:2001:MDY

[Kie01] Carolyn Kieran. The mathematical discourse of 13-year-old partnered problem solving and its relation to the mathematics that emerges. Educational Studies in Mathematics, 46(1–3): 187–228, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1014040725558.

Kilpatrick:1985:RR

[Kil85] Jeremy Kilpatrick. Reflection and recursion. Educational Studies in Mathematics, 16(1):1–26, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00354880.

Kilpatrick:2001:UML

[Kil01] Jeremy Kilpatrick. Understanding mathematical literacy: The contribution of research. Educational Studies in Mathematics, 47(1):101–116, May 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017973827514.

Kim:2020:TCH

[Kim20] Yeon Kim. Teacher community for high school mathematics instruction: strengths and challenges. Educational Studies in Mathematics, 104(1):105–125, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09943-6.

King:1974:PA

[Kin74] Amy C. King. Plan ahead. Educational Studies in Mathematics, 5(3):363–370, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01424554.

Kirsch:1969:ACA

[Kir69] Arnold Kirsch. An analysis of commercial arithmetic. Educational Studies in Mathematics, 1(3):300–311, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00558315.

Kissane:1986:SMT

[Kis86] Barry V. Kissane. Selection of mathematically talented students. Educational Studies in Mathematics, 17(3):221–241, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305071.

Komatsu:2022:GMK

[KJ22] Kotaro Komatsu and Keith Jones. Generating mathematical knowledge in the classroom through proof, refutation, and abductive reasoning. Educational Studies in Mathematics, 109 (3):567–591, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10086-5.

Kop:2020:RBG

[KJvD20] Peter M. G. M. Kop, Fred J. J. M. Janssen, and Jan H. van Driel. The relation between graphing formulas by hand and students’ symbol sense. Educational Studies in Mathematics, 105(2):137–161, October 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-020-09970-
3.

Katz:2010:ZIP

[KK10] Karin Usadi Katz and Mikhail G. Katz. Zooming in on infinitesimal 1 − .9 … in a post-triumvirate era. Educational Studies in Mathematics, 74(3):259–273, July 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9239-4.

Kasmer:2012:NSP

[KK12] Lisa Anne Kasmer and Ok-Kyeong Kim. The nature of student predictions and learning opportunities in middle school algebra. Educational Studies in Mathematics, 79(2):175–191, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9336-z.

Koichu:2013:DSS

[KK13] Boris Koichu and Igor Kontorovich. Dissecting success stories on mathematical problem posing: a case of the billiard task. Educational Studies in Mathematics, 83(1):71–86, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9431-9.

Kerr:1982:NLP

[KL82] Donald R. Kerr, Jr. and Frank K. Lester, Jr. A new look at the professional training of secondary school mathematics teachers. Educational Studies in Mathematics, 13(4):431– 441, November 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00366621.

Kynigos:2014:CAT

[KL14] Chronis Kynigos and Jean-Baptiste Lagrange. Cross-analysis as a tool to forge connections amongst theoretical frames in using digital technologies in mathematical learning. Educational Studies in Mathematics, 85(3):321–327, March 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9521-3.

Klein:2020:OMP

[KL20] Sigal Klein and Roza Leikin. Opening mathematical problems for posing open mathematical tasks: what do teachers do and feel? Educational Studies in Mathematics, 105(3):349–365, November 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09983-y.

Klamkin:1968:TMU

[Kla68] Murray S. Klamkin. On the teaching of mathematics so as to be useful. Educational Studies in Mathematics, 1(1–2): 126–160, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00426240.

Klamkin:1971:IRI

[Kla71] Murray S. Klamkin. On the ideal role of an industrial mathematician and its educational implications. Educational Studies in Mathematics, 3(2):244–269, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305452.

Krummheuer:2013:EMC

[KLBMK^{+}13] Götz Krummheuer, Marianne Leuzinger-Bohleber, Marion Müller-Kirchof, Melanie Münz, and Rose Vogel. Explaining the mathematical creativity of a young boy: an interdisciplinary venture between mathematics education and psychoanalysis. Educational Studies in Mathematics, 84(2):183–199, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9505-3.

Klee:1971:URP

[Kle71] Victor Klee. The use of research problems in high school geometry. Educational Studies in Mathematics, 3(3–4):482– 489, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302309.

Kleiner:2001:HIS

[Kle01] Israel Kleiner. History of the infinitely small and the infinitely large in calculus. Educational Studies in Mathematics, 48(2–3): 137–174, November 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1016090528065.

Klein:2002:TMN

[Kle02] M. Klein. Teaching mathematics in/for new times: A poststructuralist analysis of the productive quality of the pedagogic process. Educational Studies in Mathematics, 50(1):63– 78, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1020566020275.

Klein:2012:HIA

[Kle12] Mary Klein. How inconvenient assumptions affect preservice teachers’ uptake of new interactional patterns in mathematics: analysis and aspiration through a bifocal lens. Educational Studies in Mathematics, 80(1–2):25–40, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9390-1.

Knapp:2017:WAF

[KLLJ17] Andrea Knapp, Racheal Landers, Senfeng Liang, and Vetrece Jefferson. We all as a family are graduating tonight: a case for mathematical knowledge for parental involvement. Educational Studies in Mathematics, 95(1):79–95, May 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Kontorovich:2024:TPW

[KLwK24] Igor’ Kontorovich, Nicole Qiusong Liu, and Sun woong Kang. Transitioning to proof via writing scripts on the rules of a new discourse. Educational Studies in Mathematics, 117(1): 143–162, September 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10324-6.

Kaiser-Messmer:1993:RES

[KM93] Gabriele Kaiser-Messmer. Results of an empirical study into gender differences in attitudes towards mathematics. Educational Studies in Mathematics, 25(3):209–233, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273862.

Kelton:2018:RMS

[KM18] Molly L. Kelton and Jasmine Y. Ma. Reconfiguring mathematical settings and activity through multi-party, whole-body collaboration. Educational Studies in Mathematics, 98(2):177– 196, June 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9805-8.

Kollosche:2021:CME

[KM21] David Kollosche and Wolfram Meyerhöfer. COVID-19, mathematics education, and the evaluation of expert knowledge. Educational Studies in Mathematics, 108(1–2):401–417, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10097-2.

Khalloufi-Mouha:2024:ETP

[KM24a] Faten Khalloufi-Mouha. Evolution of teachers’ professional learning when using a technological resource: a case study of a Tunisian primary school teacher. Educational Studies in Mathematics, 116(2):281–306, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-024-10320-
w.

Kuzniak:2024:SDC

[KM24b] Alain Kuzniak and Blandine Masselin. Strongly didactic contracts and mathematical work. Educational Studies in Mathematics, 115(2):289–312, February 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10286-1.

Klingbeil:2025:OLO

[KM25] Katrin Klingbeil and Filip Moons. The omnipresence of the Letter-as-Object misconception: a multilevel Latent Transition Analysis on understanding variables. Educational Studies in Mathematics, 120(3):455–476, November 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10428-7.

Kramarski:2002:EMI

[KMA02] Bracha Kramarski, Zemira R. Mevarech, and Marsel Arami. The effects of metacognitive instruction on solving mathematical authentic tasks. Educational Studies in Mathematics, 49(2):225–250, February 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1016282811724.

Karki:2022:IRN

[KML22] Tomi Kärki, Jake McMullen, and Erno Lehtinen. Improving rational number knowledge using the NanoRoboMath digital game. Educational Studies in Mathematics, 110(1):101– 123, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10120-6.

Kinder:2024:ASM

[KMN24] Cassandra Kinder, Charles Munter, and Phi Nguyen. Aligning school mathematics with a manufactured crisis: re-rendering neoliberal and neoconservative discourses as commonsensical in a rural place. Educational Studies in Mathematics, 117 (2):263–284, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10340-6.

Kanes:2014:PMR

[KMT14] Clive Kanes, Candia Morgan, and Anna Tsatsaroni. The PISA mathematics regime: knowledge structures and practices of the self. Educational Studies in Mathematics, 87(2): 145–165, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9542-6.

Kortenkamp:2016:JMB

[KMT16] Ulrich Kortenkamp, John Monaghan, and Luc Trouche. Jonathan M. Borwein (1951–2016): exploring, experiencing and experimenting in mathematics — an inspiring journey in mathematics. Educational Studies in Mathematics, 93(2): 131–136, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9729-0.

Kuzniak:2021:FGW

[KN21] Alain Kuzniak and Assia Nechache. On forms of geometric work: a study with pre-service teachers based on the theory of Mathematical Working Spaces. Educational Studies in Mathematics, 106(2):271–289, February 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10011-2.

Knijnik:2012:DPL

[Kni12] Gelsa Knijnik. Differentially positioned language games: ethnomathematics from a philosophical perspective. Educational Studies in Mathematics, 80(1–2):87–100, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9396-8.

Knoche:1977:IVC

[Kno77] Norbert Knoche. On the indicative value of the classical goodness criteria of reliability and validity. Educational Studies in Mathematics, 8(4):479–508, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310949.

Kohen:2021:MMT

[KO21] Zehavit Kohen and Doron Orenstein. Mathematical modeling of tech-related real-world problems for secondary school-level mathematics. Educational Studies in Mathematics, 107(1):71– 91, May 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-10020-1.

Kontorovich:2023:HNA

[KO23] Igor’ Kontorovich and Tikva Ovadiya. How narratives about the secondary–tertiary transition shape undergraduate tutors’ sense-making of their teaching. Educational Studies in Mathematics, 113(1):125–146, May 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10211-6.

Kollosche:2016:CFT

[Kol16] David Kollosche. Criticising with Foucault: towards a guiding framework for socio-political studies in mathematics education. Educational Studies in Mathematics, 91(1):73–86, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9648-5.

Kollosche:2018:SFM

[Kol18] David Kollosche. Social functions of mathematics education: a framework for socio-political studies. Educational Studies in Mathematics, 98(3):287–303, July 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9818-3.

Kollosche:2021:SRM

[Kol21] David Kollosche. Styles of reasoning for mathematics education. Educational Studies in Mathematics, 107(3):471–486, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10046-z.

Komatsu:2016:FPR

[Kom16] Kotaro Komatsu. A framework for proofs and refutations in school mathematics: Increasing content by deductive guessing. Educational Studies in Mathematics, 92(2):147–162, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9677-0.

Komatsu:2017:FEE

[Kom17] Kotaro Komatsu. Fostering empirical examination after proof construction in secondary school geometry. Educational Studies in Mathematics, 96(2):129–144, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-016-9731-6.

Kooloos:2022:MSS

[KOMH22] Chris Kooloos, Helma Oolbekkink-Marchand, and Gert Heckman. Making sense of student mathematical thinking: the role of teacher mathematical thinking. Educational Studies in Mathematics, 110(3):503–524, July 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10124-2.

Konior:1972:EDC

[Kon72] Jan Konior. Essai de détermination des causes d’échecs des élèves dans la recherche de démonstrations. (French) [Test to determine the causes of failure of students in the search for demonstrations]. Educational Studies in Mathematics, 4(4): 466–483, December 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00567001.

Konior:1993:RCM

[Kon93] Jan Konior. Research into the construction of mathematical texts. Educational Studies in Mathematics, 24(3):251–256, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01275425.

Kontorovich:2018:WJS

[Kon18] Igor’ Kontorovich. Why Johnny struggles when familiar concepts are taken to a new mathematical domain: towards a polysemous approach. Educational Studies in Mathematics, 97 (1):5–20, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9778-z.

Kontorovich:2020:PPT

[Kon20] Igor’ Kontorovich. Problem-posing triggers or where do mathematics competition problems come from? Educational Studies in Mathematics, 105(3):389–406, November 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09964-1.

Kontorovich:2021:MMD

[Kon21] Igor’ Kontorovich. Minding mathematicians’ discourses in investigations of their feedback on students’ proofs: a case study. Educational Studies in Mathematics, 107(2):213–234, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10035-2. See correction [Kon22].

Kontorovich:2022:CMM

[Kon22] Igor’ Kontorovich. Correction to: Minding mathematicians’ discourses in investigations of their feedback on students’ proofs: a case study. Educational Studies in Mathematics, 109 (1):203–204, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10129-x. See

[Kon21].

Kordaki:2003:ETC

[Kor03] M. Kordaki. The effect of tools of a computer microworld on students’ strategies regarding the concept of conservation of area. Educational Studies in Mathematics, 52(2):177–209, March 2003. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1024065107302.

Kaisari:2010:WDC

[KP10] Maria Kaisari and Tasos Patronis. So we decided to call “straight line” (…): Mathematics students’ interaction and negotiation of meaning in constructing a model of elliptic geometry. Educational Studies in Mathematics, 75(3):253–269, December 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9255-4.

Kaasila:2010:FPS

[KPH10] Raimo Kaasila, Erkki Pehkonen, and Anu Hellinen. Finnish pre-service teachers’ and upper secondary students’ understanding of division and reasoning strategies used. Educational Studies in Mathematics, 73(3):247–261, April 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9213-1.

Karplus:1983:EAP

[KPS83] Robert Karplus, Steven Pulos, and Elizabeth K. Stage. Early adolescents’ proportional reasoning on ‘rate’ problems. Educational Studies in Mathematics, 14(3):219–233, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410539.

Kaspersen:2017:MSS

[KPS17] Eivind Kaspersen, Birgit Pepin, and Svein Arne Sikko. Measuring STEM students’ mathematical identities. Educational Studies in Mathematics, 95(2):163–179, June 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Kaufman:1973:RP

[KR73] Burt Kaufman and Lennart Råde. Relations and probability. Educational Studies in Mathematics, 5(1):49–64, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01425411.

Kramarski:2009:CSR

[KR09] Bracha Kramarski and Tali Revach. The challenge of selfregulated learning in mathematics teachers’ professional training. Educational Studies in Mathematics, 72(3):379–399, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-009-9204-2.

Kuzniak:2011:HDT

[KR11] Alain Kuzniak and Jean-Claude Rauscher. How do teachers’ approaches to geometric work relate to geometry students’ learning difficulties? Educational Studies in Mathematics, 77 (1):129–147, May 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9304-7.

Kallberg:2025:MSP

[KR25] Petra Svensson Källberg and Helena Roos. Meaning(s) of a student perspective in mathematics education research. Educational Studies in Mathematics, 119(2):367–392, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10374-w.

Krainer:1993:PTC

[Kra93] Konrad Krainer. Powerful tasks: A contribution to a high level of acting and reflecting in mathematics instruction. Educational Studies in Mathematics, 24(1):65–93, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273295.

Kraemer:2002:EPM

[Kra02] Jean-Marie Kraemer. Évaluer pour mieux comprendre les enfants et améliorer sa pratique. (French) [Evaluate to better understand children and improve their practice]. Educational Studies in Mathematics, 51(1–2):95–116, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022428829082.

Kapon:2015:PPD

[KRHD15] Shulamit Kapon, Gila Ron, Rina Hershkowitz, and Tommy Dreyfus. Perceiving permutations as distinct outcomes: the accommodation of a complex knowledge system. Educational Studies in Mathematics, 88(1):43–64, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9570-2.

Krummheuer:2013:RBD

[Kru13a] Götz Krummheuer. The relationship between diagrammatic argumentation and narrative argumentation in the context of the development of mathematical thinking in the early years.

Educational Studies in Mathematics, 84(2):249–265, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9471-9.

Krummheuer:2013:RML

[Kru13b] Götz Krummheuer. Research on mathematics learning at the “Center of Individual Development and Adaptive Education” (IDeA) — an introduction. Educational Studies in Mathematics, 84(2):177–181, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9502-6.

Krygovska:1968:PME

[Kry68] A. Z. Krygovska. Processus de la mathématisation dans l’enseignement. (French) [The process of mathematicizing education]. Educational Studies in Mathematics, 1(1–2):9–16, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00426225.

Krygovska:1969:PME

[Kry69a] A. Z. Krygovska. Préparation des maîtres à l’enseignement intégré des sciences. (French) [Preparing teachers for integrated science education]. Educational Studies in Mathematics, 1(3):352–364, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00558322.

Krygowska:1969:TME

[Kry69b] Zofia Krygowska. Le texte mathématique dans l’enseignement. (French) [Mathematical text in teaching]. Educational Studies in Mathematics, 2(2–3):360–370, December 1969. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303469.

Krygowska:1988:MTV

[Kry88a] Anna Zofia Krygowska. À la mémoire de Tamàs Varga. (French) [In memory of Tamàs Varga]. Educational Studies in Mathematics, 19(3):289–290, August 1988. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312448.

Krygowska:1988:CAM

[Kry88b] Zofia Krygowska. Composants de l’activité mathématique qui devraient jouer le rôle essentiel dans la mathématique pour tous. (French) [Components of mathematical activity that should play the essential role in mathematics for all]. Educational Studies in Mathematics, 19(4):423–433, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578692.

Kaufman:1969:CAC

[KS69] Burt A. Kaufman and Hans-G. Steiner. The CSMP approach to a content-oriented, highly individualized mathematics education. Educational Studies in Mathematics, 1(3):312–326, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00558316.

Krawitz:2025:DPP

[KSH25] Janina Krawitz, Stanislaw Schukajlow, and Luisa Hartmann. Does problem posing affect self-efficacy, task value, and performance in mathematical modelling? Educational Studies in Mathematics, 119(3):445–466, July 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10385-1.

Kospentaris:2011:ESS

[KSL11] George Kospentaris, Panagiotis Spyrou, and Dionyssios Lappas. Exploring students’ strategies in area conservation geometrical tasks. Educational Studies in Mathematics, 77(1): 105–127, May 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9303-8.

Leung:2013:TIM

[kSL13] Shuk kwan S. Leung. Teachers implementing mathematical problem posing in the classroom: challenges and strategies. Educational Studies in Mathematics, 83(1):103–116, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9436-4.

Komatsu:2025:DLT

[KSS25] Kotaro Komatsu, Yusuke Shinno, and Gabriel J. Stylianides. Development of local theories in design-based research in mathematics education. Educational Studies in Mathematics, 120(3):407–428, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-025-10398-
w.

Kadluba:2025:HMC

[KSSO25] Alina Kadluba, Anselm Strohmaier, Christian Schons, and Andreas Obersteiner. How much C is in TPACK? A systematic review on the assessment of TPACK in mathematics. Educational Studies in Mathematics, 118(2):169–199, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10357-x.

Keijzer:2001:AAF

[KT01] Ronald Keijzer and Jan Terwel. Audrey’s acquisition of fractions: A case study into the learning of formal mathematics. Educational Studies in Mathematics, 47(1):53–73, May 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017971912662.

Kidron:2015:RVS

[KT15] Ivy Kidron and David Tall. The roles of visualization and symbolism in the potential and actual infinity of the limit process. Educational Studies in Mathematics, 88(2):183–199, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9567-x.

Kuchemann:1981:CDS

[Küc81] Dietmar Küchemann. Cognitive demand of secondary school mathematics items. Educational Studies in Mathematics, 12 (3):301–316, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311061.

Kuntze:2012:PCB

[Kun12] Sebastian Kuntze. Pedagogical content beliefs: global, content domain-related and situation-specific components. Educational Studies in Mathematics, 79(2):273–292, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9347-9.

Kamps:1975:CCC

[KV75] H. J. L. Kamps and J. H. Van Lint. A comparison of a classical calculus test with a similar multiple choice test. Educational Studies in Mathematics, 6(3):259–271, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01793611.

Kemp:2023:SUD

[KV23] Aubrey Kemp and Draga Vidakovic. Students’ understanding and development of the definition of circle in Taxicab and Euclidean geometries: an APOS perspective with schema interaction. Educational Studies in Mathematics, 112(3):567–588, March 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10180-2.

Kazak:2015:IDP

[KWF15] Sibel Kazak, Rupert Wegerif, and Taro Fujita. The importance of dialogic processes to conceptual development in mathematics. Educational Studies in Mathematics, 90(2):105–120, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9618-y.

Kynigos:1993:CIT

[Kyn93] Chronis Kynigos. Children’s inductive thinking during intrinsic and Euclidean geometrical activities in a computer programming environment. Educational Studies in Mathematics, 24(2):177–197, February 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273691.

Khoury:1994:FNS

[KZ94] Helen A. Khoury and Rina Zazkis. On fractions and nonstandard representations: Pre-service teachers’ concepts. Educational Studies in Mathematics, 27(2):191–204, September 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01278921.

Kontorovich:2016:TVS

[KZ16] Igor’ Kontorovich and Rina Zazkis. Turn vs. shape: teachers cope with incompatible perspectives on angle. Educational Studies in Mathematics, 93(2):223–243, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9699-2.

Krainer:2019:ISN

[KZH^{+}19] Konrad Krainer, Stefan Zehetmeier, Barbara Hanfstingl, Franz Rauch, and Tanja Tscheinig. Insights into scaling up a nationwide learning and teaching initiative on various levels. Educational Studies in Mathematics, 102 (3):395–415, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9826-
3;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9826-3.pdf.

Laborde:1978:RAA

[Lab78] C. Laborde. Relations arithmétiques — aspect statique — aspect dynamique. (French) [Arithmetical relations — static aspect — dynamic aspect]. Educational Studies in Mathematics, 9(1):41–50, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00352191.

Laborde:1992:GE

[Lab92] Colette Laborde. Guest editorial. Educational Studies in Mathematics, 23(1):1–4, February 1992. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302311.

Laborde:2000:DGE

[Lab00] Colette Laborde. Dynamic geometry environments as a source of rich learning contexts for the complex activity of proving. Educational Studies in Mathematics, 44(1–3):151–161, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1012793121648.

LaCroix:2014:LSP

[LaC14] Lionel LaCroix. Learning to see pipes mathematically: preapprentices’ mathematical activity in pipe trades training. Educational Studies in Mathematics, 86(2):157–176, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9534-6.

Lambert:2015:CRD

[Lam15] Rachel Lambert. Constructing and resisting disability in mathematics classrooms: a case study exploring the impact of different pedagogies. Educational Studies in Mathematics, 89 (1):1–18, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9587-6.

Lancy:1981:IMP

[Lan81] David F. Lancy. The indigenous mathematics project: An overview. Educational Studies in Mathematics, 12(4):445–453, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00308142.

Langford:1986:AWP

[Lan86a] Peter E. Langford. Arithmetical word problems: Thinking in the head versus thinking on the table. Educational Studies in Mathematics, 17(2):193–199, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311520. See comments [Ril86].

Langford:1986:RR

[Lan86b] Peter E. Langford. Reply to Riley. Educational Studies in Mathematics, 17(2):203–204, May 1986. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311522. See [Lan86a, Ril86].

Landers:2013:TTM

[Lan13] Mara G. Landers. Towards a theory of mathematics homework as a social practice. Educational Studies in Mathematics, 84(3):371–391, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9487-1.

Larsson:1986:GPE

[Lar86] Yvonne Larsson. Governmental policies on the education of gifted and talented children: A world view. Educational Studies in Mathematics, 17(3):213–219, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305070.

Laugwitz:1974:MLA

[Lau74] Detlef Laugwitz. Motivations and linear algebra. Educational Studies in Mathematics, 5(3):243–254, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424547.

Lawson:2000:BRL

[Law00] Alex Lawson. Book review: Liping Ma, Knowing and teaching elementary mathematics: Teachers’ understanding of fundamental mathematics in China and the United States. Educational Studies in Mathematics, 42(1):101–106, April 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004020607375.

Lawrence:2016:BRP

[Law16] Snezana Lawrence. Book review: The power of a great introduction. Karp, Alexander, Schubring, Gert (Eds.) (2014). Handbook on the history of mathematics education. Educational Studies in Mathematics, 92(2):279–285, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9688-5.

Lawrence:2024:BRN

[Law24] Snezana Lawrence. Book review: The new mathematics education of new Europe. Alexander Karp (Ed.) (2020) Eastern European mathematics education in the decades of change. Educational Studies in Mathematics, 116(2):307–312, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10269-2.

Leung:2013:DID

[LBFM13] Allen Leung, Anna Baccaglini-Frank, and Maria Alessandra Mariotti. Discernment of invariants in dynamic geometry environments. Educational Studies in Mathematics, 84(3):439– 460, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9492-4.

Leavy:2023:PPT

[LBH23] Aisling Leavy, Annette Hessen Bjerke, and Mairéad Hourigan. Prospective primary teachers’ efficacy to teach mathematics: measuring efficacy beliefs and identifying the factors that influence them. Educational Studies in Mathematics, 112(3):437– 460, March 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10181-1.

Levenson:2024:CEG

[LBML24] Esther S. Levenson, Ruthi Barkai, Anas Mahamid, and Sigal Levy. Creating equal groups from n items: mathematical creativity in kindergarten. Educational Studies in Mathematics, 116(2):145–163, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10322-8.

Lannin:2007:HSV

[LBT07] John K. Lannin, David D. Barker, and Brian E. Townsend. How students view the general nature of their errors. Educational Studies in Mathematics, 66(1):43–59, September 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-006-9067-8.pdf.

Levenson:2022:EAA

[LBT22] Esther S. Levenson, Ruthi Barkai, and Pessia Tsamir. Exploring adults’ awareness of and suggestions for early childhood numerical activities. Educational Studies in Mathematics, 109 (1):5–21, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10063-y.

Lean:1981:SAV

[LC81] Glen Lean and M. A. (Ken) Clements. Spatial ability, visual imagery, and mathematical performance. Educational Studies in Mathematics, 12(3):267–299, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311060.

Lim:2013:DSF

[LC13] Siew Yee Lim and Elaine Chapman. Development of a short form of the attitudes toward mathematics inventory. Educational Studies in Mathematics, 82(1):145–164, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9414-x.

Lim:2015:EUH

[LC15] Siew Yee Lim and Elaine Chapman. Effects of using history as a tool to teach mathematics on students’ attitudes, anxiety, motivation and achievement in grade 11 classrooms. Educational Studies in Mathematics, 90(2):189–212, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9620-4.

Lean:1990:LPF

[LCD90] Glendon A. Lean, McKenzie A. Clements, and Gina Del Campo. Linguistic and pedagogical factors affecting children’s understanding of arithmetic word problems: A comparative study. Educational Studies in Mathematics, 21(2):165–191, April 1990. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00304900.

Li:2024:MTB

[LCK24] Rangmei Li, Mustafa Cevikbas, and Gabriele Kaiser. Mathematics teachers’ beliefs about their roles in teaching mathematics: orchestrating scaffolding in cooperative learning. Educational Studies in Mathematics, 117(3):357–377, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10359-9.

Lockwood:2020:EPC

[LCW20] Elise Lockwood, John S. Caughman, and Keith Weber. An essay on proof, conviction, and explanation: multiple representation systems in combinatorics. Educational Studies in Mathematics, 103(2):173–189, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09933-8.

Lecoutre:1988:JPM

[LD88] Marie-Paule Lecoutre and Jean-Luc Durand. Jugements probabilistes et modèles cognitifs: Étude d’une situation aléatoire. (French) [Probabilistic judgments and cognitive models: Study of a random situation]. Educational Studies in Mathematics, 19(3):357–368, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312452.

Dorier:1995:MLT

[lD95] Jean luc Dorier. Meta level in the teaching of unifying and generalizing concepts in mathematics. Educational Studies in Mathematics, 29(2):175–197, September 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274212.

lmaz:2021:WAT

[lDlzN21] Zuhal Yılmaz, Hande GülbağcıDede, and Selin Yıldı z Nielsen. Are we all in this together?: mathematics teachers’ perspectives on equity in remote instruction during pandemic. Educational Studies in Mathematics, 108(1–2):307–331, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10060-1.

Lemerise:1992:EEL

[LDM92] Tamara Lemerise, Claude Dessurealt, and Lucille Marcotte. Élaboration d’un environnement logo favorable au développement d’habiletés en manipulation procédurale. (French) [Elaboration of an enabling Logo environment for the development of procedural manipulation skills]. Educational Studies in Mathematics, 23(4):351–382, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302440.

Lagrange:2009:TEG

[LE09] Jean-Baptiste Lagrange and Emel Ozdemir Erdogan. Teachers’ emergent goals in spreadsheet-based lessons: analyzing the complexity of technology integration. Educational Studies in Mathematics, 71(1):65–84, May 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9160-2.

Lecoutre:1992:CMP

[Lec92] Marie-Paule Lecoutre. Cognitive models and problem spaces in “purely random” situations. Educational Studies in Mathematics, 23(6):557–568, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00540060.

Leder:1980:BGM

[Led80] Gilah Leder. Bright girls, mathematics and fear of success. Educational Studies in Mathematics, 11(4):411–422, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00231214.

Leder:1985:SRD

[Led85] Gilah Leder. Sex-related differences in mathematics: An overview. Educational Studies in Mathematics, 16(3):304–309, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00776737.

Leder:1987:RCN

[Led87a] Gilah C. Leder. Response to ‘a cautionary note’. Educational Studies in Mathematics, 18(3):327–328, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386203.

Leder:1987:TSI

[Led87b] Gilah C. Leder. Teacher student interaction: A case study. Educational Studies in Mathematics, 18(3):255–271, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386198.

Leder:1992:MBF

[Led92] Gilah C. Leder. Mathematics before formal schooling. Educational Studies in Mathematics, 23(4):383–396, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302441.

Leder:1995:GE

[Led95] Gilah Leder. Guest editorial. Educational Studies in Mathematics, 28(3):195–198, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274172.

Lee:2017:PST

[Lee17] Mi Yeon Lee. Pre-service teachers’ flexibility with referent units in solving a fraction division problem. Educational Studies in Mathematics, 96(3):327–348, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9771-6.

Lehmann:2023:LMA

[Leh23] Timothy H. Lehmann. Learning to measure the area of composite shapes. Educational Studies in Mathematics, 112 (3):531–565, March 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10191-z.

Lehmann:2024:MMV

[Leh24] Timothy H. Lehmann. Mathematical modelling as a vehicle for eliciting algorithmic thinking. Educational Studies in Mathematics, 115(2):151–176, February 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10275-4.

Lenart:1993:AMD

[Lén93] István Lénárt. Alternative models on the drawing ball. Educational Studies in Mathematics, 24(3):277–312, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01275428.

Lepik:1990:AWP

[Lep90] Madis Lepik. Algebraic word problems: Role of linguistic and structural variables. Educational Studies in Mathematics, 21(1):83–90, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311017.

Leron:1985:DAI

[Ler85] Uri Leron. A direct approach to indirect proofs. Educational Studies in Mathematics, 16(3):321–325, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00776741.

Lerman:1989:CMM

[Ler89] Stephen Lerman. Constructivism, mathematics and mathematics education. Educational Studies in Mathematics, 20 (2):211–223, May 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00579463.

Lerman:1996:GE

[Ler96] Stephen Lerman. Guest editorial. Educational Studies in Mathematics, 31(1–2):1–9, September 1996. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00143924.

Lerman:2001:CDP

[Ler01] Stephen Lerman. Cultural, discursive psychology: A sociocultural approach to studying the teaching and learning of mathematics. Educational Studies in Mathematics, 46(1–3): 87–113, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1014031004832.

Lerman:2014:MEP

[Ler14] Stephen Lerman. Mapping the effects of policy on mathematics teacher education. Educational Studies in Mathematics, 87(2):187–201, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9423-9.

Lesh:1981:AMP

[Les81] Richard Lesh. Applied mathematical problem solving. Educational Studies in Mathematics, 12(2):235–264, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305624.

Leung:1987:SSM

[Leu87] F. K. S. Leung. The secondary school mathematics curriculum in China. Educational Studies in Mathematics, 18(1): 35–57, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00367913.

Leung:1995:MCB

[Leu95] F. K. S. Leung. The mathematics classroom in Beijing, Hong Kong and London. Educational Studies in Mathematics, 29(4): 297–325, December 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273909.

Leung:2001:SEA

[Leu01] Frederick K. S. Leung. In search of an East Asian identity in mathematics education. Educational Studies in Mathematics, 47(1):35–51, May 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017936429620.

Leung:2005:SCE

[Leu05] Frederick Koon Shing Leung. Some characteristics of East Asian mathematics classrooms based on data from the TIMSS 1999 Video Study. Educational Studies in Mathematics, 60(2): 199–215, October 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-3835-8.

Leung:2021:JCC

[Leu21] Frederick K. S. Leung. Jewish culture, Chinese culture, and mathematics education. Educational Studies in Mathematics, 107(2):405–423, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10034-3.

Levi:1971:GAH

[Lev71] Howard Levi. Geometric algebra for the high school program. Educational Studies in Mathematics, 3(3–4):490–500, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302310.

Levin:1981:ETA

[Lev81] James A. Levin. Estimation techniques for arithmetic: Everyday math and mathematics instruction. Educational Studies in Mathematics, 12(4):421–434, November 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00308140.

Levain:1992:RPM

[Lev92] Jean-Pierre Levain. La résolution de problèmes multiplicatifs à la fin du cycle primaire. (French) [Resolution of multiplicative problems at the end of the primary cycle]. Educational Studies in Mathematics, 23(2):139–161, April 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00588053.

Levenson:2010:FGS

[Lev10] Esther Levenson. Fifth-grade students’ use and preferences for mathematically and practically based explanations. Educational Studies in Mathematics, 73(2):121–142, March 2010.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9208-y.

Levenson:2013:EOS

[Lev13] Esther Levenson. Exploring one student’s explanations at different ages: the case of Sharon. Educational Studies in Mathematics, 83(2):181–203, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9447-1.

Levenson:2022:ERB

[Lev22] Esther S. Levenson. Exploring the relationship between teachers’ values and their choice of tasks: the case of occasioning mathematical creativity. Educational Studies in Mathematics, 109(3):469–489, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10101-9.

Laborde:1987:BR

[LG87] Colette Laborde and Lucia Grugnetti. Book reviews. Educational Studies in Mathematics, 18(3):329–337, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386204.

Leikin:2013:TMG

[LG13] Roza Leikin and Dorith Grossman. Teachers modify geometry problems: from proof to investigation. Educational Studies in Mathematics, 82(3):515–531, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9460-4.

Lockwood:2016:CTO

[LG16] Elise Lockwood and Bryan R. Gibson. Combinatorial tasks and outcome listing: Examining productive listing among undergraduate students. Educational Studies in Mathematics, 91(2):247–270, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9664-5.

Lamon:1971:LVS

[LH71] William E. Lamon and Leslie E. Huber. The learning of the vector space structure by sixth grade students. Educational Studies in Mathematics, 4(2):166–181, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303382.

Linchevski:1996:CCG

[LH96] Liora Linchevski and Nicolas Herscovics. Crossing the cognitive gap between arithmetic and algebra: Operating on the unknown in the context of equations. Educational Studies in Mathematics, 30(1):39–65, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00163752.

Leron:1997:WAJ

[LH97] Uri Leron and Orit Hazzan. The world according to Johnny; a coping perspective in mathematics education. Educational Studies in Mathematics, 32(3):265–292, March 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002908608251.

Leron:2006:RDA

[LH06] Uri Leron and Orit Hazzan. The rationality debate: Application of cognitive psychology to mathematics education. Educational Studies in Mathematics, 62(2):105–126, June 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-4833-1.

Lesh:2008:IBV

[LH08] Richard Lesh and Stephen Hegedus. Introduction: Building on the vision of Jim Kaput (1942–2005). Educational Studies in Mathematics, 68(2):81–83, June 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9122-8.

Leron:2009:IVA

[LH09] Uri Leron and Orit Hazzan. Intuitive vs analytical thinking: four perspectives. Educational Studies in Mathematics, 71(3): 263–278, July 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9175-8.

Leavy:2018:RPS

[LH18] Aisling Leavy and Mairead Hourigan. The role of perceptual similarity, context, and situation when selecting attributes: considerations made by 5–6-year-olds in data modeling environments. Educational Studies in Mathematics, 97(2):163–183, February 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-017-9791-2.

Lambert:2022:IAD

[LH22a] Rachel Lambert and Edmund Harriss. Insider accounts of dyslexia from research mathematicians. Educational Studies in Mathematics, 111(1):89–107, September 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10140-2.

Leavy:2022:FPE

[LH22b] Aisling Leavy and Mairéad Hourigan. The Framework for Posing Elementary Mathematics Problems (F-PosE): Supporting teachers to evaluate and select problems for use in elementary mathematics. Educational Studies in Mathematics, 111 (1):147–176, September 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10155-3.

Lepola:2019:SFN

[LHS19] Janne Lepola and Minna Hannula-Sormunen. Spontaneous focusing on numerosity and motivational orientations as predictors of arithmetical skills from kindergarten to grade 2. Educational Studies in Mathematics, 100(3):251–269, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9851-2.

Leron:1995:LGI

[LHZ95] Uri Leron, Orit Hazzan, and Rina Zazkis. Learning group isomorphism: A crossroads of many concepts. Educational Studies in Mathematics, 29(2):153–174, September 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274211.

Libois:1971:MDS

[Lib71] Lucie Libois. Mathematics in the Decroly School. Educational Studies in Mathematics, 4(2):182–186, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303383.

Libeskind:1977:PSA

[Lib77] Shlomo Libeskind. A problem solving approach to teaching mathematics. Educational Studies in Mathematics, 8(2): 167–179, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00241023.

Liebeck:1990:SFI

[Lie90] Pamela Liebeck. Scores and foreits — an intuitive model for integer arithmetic. Educational Studies in Mathematics, 21 (3):221–239, June 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305091.

Lin:1988:SDT

[Lin88] Fou-Lai Lin. Societal differences and their influences on children’s mathematics understanding. Educational Studies in Mathematics, 19(4):471–497, November 1988. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578695.

Lin:1990:ESI

[Lin90] Ng Li Fung Lin. The effect of superfluous information on children’s solution of story arithmetic problems. Educational Studies in Mathematics, 21(6):509–520, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00315942.

Lithner:2000:MRS

[Lit00] Johan Lithner. Mathematical reasoning in school tasks. Educational Studies in Mathematics, 41(2):165–190, February 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003956417456.

Lithner:2003:SMR

[Lit03] J. Lithner. Students’ mathematical reasoning in university textbook exercises. Educational Studies in Mathematics, 52 (1):29–55, January 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1023683716659.

Lithner:2008:RFC

[Lit08] Johan Lithner. A research framework for creative and imitative reasoning. Educational Studies in Mathematics, 67(3): 255–276, March 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9104-2.

Lee:2013:LML

[LJW13] Clare Lee and Sue Johnston-Wilder. Learning mathematics — letting the pupils have their say. Educational Studies in Mathematics, 83(2):163–180, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9445-3.

Lagrange:2014:DTT

[LK14a] Jean-Baptiste Lagrange and Chronis Kynigos. Digital technologies to teach and learn mathematics: Context and recontextualization. Educational Studies in Mathematics, 85(3): 381–403, March 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9525-z.

Lutovac:2014:PST

[LK14b] Sonja Lutovac and Raimo Kaasila. Pre-service teachers’ future-oriented mathematical identity work. Educational Studies in Mathematics, 85(1):129–142, January 2014.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9500-8.

Lee:2016:PST

[LK16] Ji-Eun Lee and Kyoung-Tae Kim. Pre-service teachers’ conceptions of effective teacher talk: their critical reflections on a sample teacher-student dialogue. Educational Studies in Mathematics, 93(3):363–381, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9710-y.

Lutovac:2018:ETN

[LK18] Sonja Lutovac and Raimo Kaasila. An elementary teacher’s narrative identity work at two points in time two decades apart. Educational Studies in Mathematics, 98(3):253–267, July 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9816-5.

Lu:2022:CSM

[LK22] Xiaoli Lu and Gabriele Kaiser. Creativity in students’ modelling competencies: conceptualisation and measurement. Educational Studies in Mathematics, 109(2):287–311, February 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10055-y.

Linchevski:1999:SSR

[LL99] Liora Linchevski and Drora Livneh. Structure sense: The relationship between algebraic and numerical contexts. Educational Studies in Mathematics, 40(2):173–196, October 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003606308064.

Leung:2013:SGP

[LL13] Allen Leung and Arthur Man Sang Lee. Students’ geometrical perception on a task-based dynamic geometry platform. Educational Studies in Mathematics, 82(3):361–377, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9433-7.

Lagrange:2023:CAM

[LL23] Jean-Baptiste Lagrange and Dominique Laval. Connecting algorithmics to mathematics learning: a design study of the intermediate value theorem and the bisection algorithm. Educational Studies in Mathematics, 112(2):225–245, February 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10192-y.

Logan:2014:CTG

[LLD14] Tracy Logan, Tom Lowrie, and Carmel M. Diezmann. Cothought gestures: Supporting students to successfully navigate map tasks. Educational Studies in Mathematics, 87(1): 87–102, September 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9546-2.

Llewellyn:2012:UUR

[Lle12] Anna Llewellyn. Unpacking understanding: the (re)search for the Holy Grail of mathematics education. Educational Studies in Mathematics, 81(3):385–399, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9409-7.

Llewellyn:2016:PPP

[Lle16] Anna Llewellyn. Problematising the pursuit of progress in mathematics education. Educational Studies in Mathematics, 92(3):299–314, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9645-8.

Lonnemann:2013:GDB

[LLHL13] Jan Lonnemann, Janosch Linkersdörfer, Marcus Hasselhorn, and Sven Lindberg. Gender differences in both tails of the distribution of numerical competencies in preschool children. Educational Studies in Mathematics, 84(2):201–208, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9488-0.

Lu:2021:TAI

[LLL21] Xiaoli Lu, Frederick Koon Shing Leung, and Na Li. Teacher agency for integrating history into teaching mathematics in a performance-driven context: a case study of a beginning teacher in China. Educational Studies in Mathematics, 106 (1):25–44, January 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-10006-z.

Lloyd:2006:PTS

[Llo06] Gwendolyn M. Lloyd. Preservice teachers’ stories of mathematics classrooms: Explorations of practice through fictional accounts. Educational Studies in Mathematics, 63(1):57–87, September 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-9004-2.

Leikin:2007:EMT

[LLW07] Roza Leikin and Anat Levav-Waynberg. Exploring mathematics teacher knowledge to explain the gap between theory-based recommendations and school practice in the use of connecting tasks. Educational Studies in Mathematics, 66(3):349– 371, November 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9071-z.

Liu:2020:RBD

[LLZ20] Qimeng Liu, Jian Liu, and Zhikun Zhang. The relationship between domain- and task-specific self-efficacy and mathematical problem posing: a large-scale study of eighth-grade students in China. Educational Studies in Mathematics, 105(3):407–431, November 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09977-w.

Lange:2011:ASS

[LM11] Troels Lange and Tamsin Meaney. I actually started to scream: emotional and mathematical trauma from doing school mathematics homework. Educational Studies in Mathematics, 77 (1):35–51, May 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9298-1.

Lesh:2008:SND

[LMCG08] Richard Lesh, James A. Middleton, Elizabeth Caylor, and Shweta Gupta. A science need: Designing tasks to engage students in modeling complex data. Educational Studies in Mathematics, 68(2):113–130, June 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9118-4.

Long:1970:RME

[LMH70] Richard S. Long, Nancy S. Meltzer, and Peter J. Hilton. Research in mathematics education. Educational Studies in Mathematics, 2(4):446–468, February 1970. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417627.

Luis:1991:CEA

[LMW91] E. Luis, A. Moreno, and Guillermina Waldegg. The conceptual evolution of actual mathematical infinity. Educational Studies in Mathematics, 22(3):211–231, June 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368339.

Liang:2023:SCB

[LNC23] Biyao Liang, Oi-Lam Ng, and Yip-Cheung Chan. Seeing the continuity behind “double discontinuity”: investigating Hong Kong prospective mathematics teachers’ secondary–tertiary transition. Educational Studies in Mathematics, 113(1):107– 124, May 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10197-7.

Lochhead:1992:KBB

[Loc92] Jack Lochhead. Knocking down the building blocks of learning: Constructivism and the ventures program. Educational Studies in Mathematics, 23(5):543–552, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00571472.

Lockwood:2011:SCA

[Loc11] Elise Lockwood. Student connections among counting problems: an exploration using actor-oriented transfer. Educational Studies in Mathematics, 78(3):307–322, December 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9320-7.

Langer-Osuna:2020:WWW

[LOMG^{+}20] Jennifer Langer-Osuna, Jen Munson, Emma Gargroetzi, Immanuel Williams, and Rosa Chavez. “So what are we working on?”: how student authority relations shift during collaborative mathematics activity. Educational Studies in Mathematics, 104(3):333–349, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-020-09962-3.

Lorenz:1982:SPA

[Lor82] Jenz Holger Lorenz. On some psychological aspects of mathematics achievement assessment and classroom interaction. Educational Studies in Mathematics, 13(1):1–19, February 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305495.

Louie:2018:CIM

[Lou18] Nicole L. Louie. Culture and ideology in mathematics teacher noticing. Educational Studies in Mathematics, 97(1):55–69, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-017-9775-2.

Lenz:2024:CWB

[LOW24a] Katja Lenz, Andreas Obersteiner, and Gerald Wittmann. Correction to: Who benefits most from language-responsive learning materials in mathematics? Investigating differential effects in heterogeneous classrooms. Educational Studies in Mathematics, 116(2):213, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-024-10333-

5. See [LOW24b].

Lenz:2024:WBM

[LOW24b] Katja Lenz, Andreas Obersteiner, and Gerald Wittmann. Who benefits most from language-responsive learning materials in mathematics? Investigating differential effects in heterogeneous classrooms. Educational Studies in Mathematics, 116 (2):185–211, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10321-9. See

correction [LOW24a].

Lorenz:1991:BR

[LP91] Jens Holger Lorenz and John P. Pace. Book review. Educational Studies in Mathematics, 22(1):87–103, February 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302718.

Laborde:2005:ITS

[LPG05] Colette Laborde and Marie-Jeanne Perrin-Glorian. Introduction teaching situations as object of research: Empirical studies within theoretical perspectives. Educational Studies in Mathematics, 59(1–3):1–12, July 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5761-1.

Lopez:2004:RME

[LQV04] Jorge M. López, Ana Helvia Quintero, and Wanda Velázquez. A reaction to “Mathematics education: A vision of its evolution” by Ricardo Cantoral and Rosa M. Farfán. Educational Studies in Mathematics, 57(2):291–294, September 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000049270.83790.bc.  See

[CF03].

Lang:1981:GES

[LR81] Bryan Lang and Peter Ruane. Geometry in English secondary schools. Educational Studies in Mathematics, 12(1):123–132, February 1981. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00386050.

Lew:2020:LCM

[LR20] Kristen Lew and Juan Pablo Mejía Ramos. Linguistic conventions of mathematical proof writing across pedagogical contexts. Educational Studies in Mathematics, 103(1):43–62, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-019-09915-5.

leRoux:2016:CDA

[lRA16] Kate le Roux and Jill Adler. A critical discourse analysis of practical problems in a foundation mathematics course at a South African university. Educational Studies in Mathematics, 91(2):227–246, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9656-5.

Lamon:1970:ISE

[LS70] William E. Lamon and Lloyd F. Scott. An investigation of structure in elementary school mathematics: Isomorphism. Educational Studies in Mathematics, 3(1):95–110, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381597.

Lowenthal:1979:LJA

[LS79] F. Lowenthal and R. Severs. Langage, jeu et activité mathématique. (French) [Language, game and mathematical activity]. Educational Studies in Mathematics, 10(2):245–262, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00230989.

Lee:2011:CRC

[LS11] Kyeong-Hwa Lee and Bharath Sriraman. Conjecturing via reconceived classical analogy. Educational Studies in Mathematics, 76(2):123–140, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-010-9274-1.
pdf.

Liebendorfer:2020:QMH

[LS20] Michael Liebendörfer and Stanislaw Schukajlow. Quality matters: how reflecting on the utility value of mathematics affects future teachers’ interest. Educational Studies in Mathematics, 105(2):199–218, October 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09982-z.

Li:2020:LTM

[LSC20] Xinlian Li, Naiqing Song, and Jinfa Cai. Learning to teach mathematics through problem posing: teachers’ beliefs and performance on problem posing. Educational Studies in Mathematics, 105(3):325–347, November 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09981-0.

Lau:2009:CMI

[LSH09] Paul Ngee-Kiong Lau, Parmjit Singh, and Tee-Yong Hwa. Constructing mathematics in an interactive classroom context. Educational Studies in Mathematics, 72(3):307–324, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9196-y.

Lavie:2019:RWL

[LSS19] Irit Lavie, Aya Steiner, and Anna Sfard. Routines we live by: from ritual to exploration. Educational Studies in Mathematics, 101(2):153–176, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9817-
4;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9817-4.pdf.

Lemoyne:1986:AMP

[LT86] G. Lemoyne and C. Tremblay. Addition and multiplication: Problem-solving and interpretation of relevant data. Educational Studies in Mathematics, 17(2):97–123, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311516.

Lambert:2020:BRD

[LT20] Rachel Lambert and Paulo Tan. Book review: David Kollosche, Renato de Souza, Michel Knigge, Miriam Godoy Penteado, and Ole Skovsmose (Eds.) (2019) Inclusive mathematics education: state-of-the-art research from Brazil and Germany. Educational Studies in Mathematics, 103(2):231–239, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-019-09930-6.

Liu:2023:EMP

[LT23] Shimeng Liu and Miwa Aoki Takeuchi. Embodied mathematical pedagogy to liberate racialized and multilingual bodies. Educational Studies in Mathematics, 112(2):267–287, February 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10185-x.

Lunkenbein:1977:RTI

[Lun77] Dieter Lunkenbein. Rationalising teaching interventions — a working model of a process of research in mathematics teaching. Educational Studies in Mathematics, 8(3):271–293, October 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00385926.

Lunkenbein:1982:BR

[Lun82] Dieter Lunkenbein. Book review. Educational Studies in Mathematics, 13(3):347–348, August 1982. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311248.

Lundin:2012:HSL

[Lun12] Sverker Lundin. Hating school, loving mathematics: On the ideological function of critique and reform in mathematics education. Educational Studies in Mathematics, 80(1–2):73– 85, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9366-6.

Lesh:1983:BR

[LW83] Richard Lesh and G. Walther. Book reviews. Educational Studies in Mathematics, 14(3):319–322, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410545.

Lee:1989:AC

[LW89] Lesley Lee and David Wheeler. The arithmetic connection. Educational Studies in Mathematics, 20(1):41–54, February 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00356040.

Lo:1994:LON

[LW94] Jane-Jane Lo and Grayson H. Wheatley. Learning opportunities and negotiating social norms in mathematics class discussion. Educational Studies in Mathematics, 27(2):145–164, September 1994. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01278919.

Linchevski:1999:UIE

[LW99] Liora Linchevski and Julian Williams. Using intuition from everyday life in ‘filling’ the gap in children’s extension of their number concept to include the negative numbers. Educational Studies in Mathematics, 39(1–3):131–147, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003726317920.

Lai:2014:FMP

[LW14] Yvonne Lai and Keith Weber. Factors mathematicians profess to consider when presenting pedagogical proofs. Educational Studies in Mathematics, 85(1):93–108, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9497-z.

Li:2016:BRS

[LX16] Yeping Li and Jianxing Xu. Book review: Student voice in mathematics classrooms around the world. Berinderjeet Kaur, Glenda Anthony, Minoru Ohtani, and David Clarke (Eds.) (2013). Educational Studies in Mathematics, 91(1):141–148, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9647-6.

Lerman:2002:DTM

[LXT02] Stephen Lerman, Guorong Xu, and Anna Tsatsaroni. Developing theories of mathematics education research: the ESM story. Educational Studies in Mathematics, 51(1–2):23–40, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1022412318413.

Lim:2023:DSN

[LYBK23] Woong Lim, Hyunkyoung Yoon, Younggon Bae, and Oh Nam Kwon. The development of sociomathematical norms in the transition to tertiary exam-oriented individualistic mathematics education in an East Asian context. Educational Studies in Mathematics, 113(1):57–78, May 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10203-y.

Lin:2018:MTE

[LYHC18] Fou-Lai Lin, Kai-Lin Yang, Hui-Yu Hsu, and Jian-Cheng Chen. Mathematics teacher educator-researchers’ perspectives on the use of theory in facilitating teacher growth. Educational Studies in Mathematics, 98(2):197–214, June 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9804-9.

Lyness:1968:AME

[Lyn68] R. C. Lyness. Applied mathematics in English schools. Educational Studies in Mathematics, 1(1–2):98–104, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426236.

Larsen:2008:PRU

[LZ08] Sean Larsen and Michelle Zandieh. Proofs and refutations in the undergraduate mathematics classroom. Educational Studies in Mathematics, 67(3):205–216, March 2008.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9106-0.

Ma:1996:ECH

[Ma96] Xin Ma. The effects of cooperative homework on mathematics achievement of Chinese high school students. Educational Studies in Mathematics, 31(4):379–387, December 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00369155.

Mulat:2009:SMW

[MA09] Tiruwork Mulat and Abraham Arcavi. Success in mathematics within a challenged minority: the case of students of Ethiopian origin in Israel (SEO). Educational Studies in Mathematics, 72(1):77–92, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-009-
9186-0.

Milan:2025:BGF

[MA25] Valeria Cruz Milán and Mario Sánchez Aguilar. Between global forces and local values: a Bakhtinian perspective on Mexican mathematics education research. Educational Studies in Mathematics, 119(2):311–333, June 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10391-3.

MacDonald:1978:IIM

[Mac78] I. D. MacDonald. Insight and intuition in mathematics. Educational Studies in Mathematics, 9(4):411–420, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410586.

Macnab:2000:RSM

[Mac00] Donald Macnab. Raising standards in mathematics education: values, vision, and TIMSS. Educational Studies in Mathematics, 42(1):61–80, April 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004190310335.

Maciejewski:2021:TMR

[Mac21] Wes Maciejewski. Teaching math in real time. Educational Studies in Mathematics, 108(1–2):143–159, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10090-9.

Mahdi:2021:BRD

[Mah21a] Fardowsa M. Mahdi. Book review: A direct link between the language of mathematics and teacher preparation programs. Patrick M. Jenlink (Ed.) (2020) The language of mathematics: how the teacher’s knowledge of mathematics affects instruction, Rowman & Littlefield Publishers. 230 pages. Hardcover: 978-1-4758-5479-4 $70.00 US. Paperback: 978-1-4758-5480-0 $35.00 US. eBook: 978-1-4758-5481-7 $33.00 US. Educational Studies in Mathematics, 107(3):607–611, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-10024-x. See correction [Mah21a].

Mahdi:2021:CBR

[Mah21b] Fardowsa M. Mahdi. Correction to: Book Review: A direct link between the language of mathematics and teacher preparation programs. Patrick M. Jenlink (Ed.) (2020) The language of mathematics: how the teacher’s knowledge of mathematics affects instruction. Educational Studies in Mathematics, 107(3):613, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10085-6. See

[Mah21a].

Moreno-Armella:2008:SDM

[MAHK08] Luis Moreno-Armella, Stephen J. Hegedus, and James J. Kaput. From static to dynamic mathematics: historical and representational perspectives. Educational Studies in Mathematics, 68(2):99–111, June 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9116-6.

Maines:1985:PNT

[Mai85] David R. Maines. Preliminary notes on a theory of informal barriers for women in mathematics. Educational Studies in Mathematics, 16(3):314–320, August 1985. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00776740.

Makar:2014:YCE

[Mak14] Katie Makar. Young children’s explorations of average through informal inferential reasoning. Educational Studies in Mathematics, 86(1):61–78, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9526-y.

Mallam:1993:IST

[Mal93] Winifred A. Mallam. Impact of school-type and sex of the teacher on female students’ attitudes toward mathematics in Nigerian secondary schools. Educational Studies in Mathematics, 24(2):223–229, February 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273693.

Malmivuori:2006:ASR

[Mal06] Marja-Liisa Malmivuori. Affect and self-regulation. Educational Studies in Mathematics, 63(2):149–164, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9022-8.

Mala:2023:BRS

[Mal23a] Firdous Ahmad Mala. Book review: Stacy A. Costa, Marcel Danesi, Dragana Martinovic (Eds.) (2020) Mathematics (education) in the information age. Educational Studies in Mathematics, 112(1):199–205, January 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10148-2.

Mala:2023:BRV

[Mal23b] Firdous Ahmad Mala. Book review: Viviane Durand– Guerrier, Reinhard Hochmuth, Elena Nardi, & Carl Winsløw (Eds.) (2021) Research and development in university mathematics education. Educational Studies in Mathematics, 112 (3):589–595, March 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10186-w.

Montecino:2024:UDA

[MAM24] Alex Montecino and Melissa Andrade-Molina. Unpacking discourses about the transition from school to university mathematics: an intensive reading. Educational Studies in Mathematics, 117(3):341–356, November 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10288-z.

Moreno-Arotzena:2021:USU

[MAPHB21] Oihana Moreno-Arotzena, Ion Pombar-Hospitaler, and José Ignacio Barragués. University student understanding of the gradient of a function of two variables: an approach from the perspective of the theory of semiotic representation registers. Educational Studies in Mathematics, 106(1):65–89, January 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-09994-9.

Markouchevitch:1969:CPE

[Mar69] A. Markouchevitch. Certains problèmes de l’enseignement des mathématiques à l’école. (French) [Some problems of teaching mathematics at school]. Educational Studies in Mathematics, 2(2–3):279–289, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00303464.

Marcus:1972:CVF

[Mar72] Philip S. Marcus. Comments on variable-free calculus. Educational Studies in Mathematics, 4(3):324–330, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302581.

Markovits:1986:CEU

[Mar86] Henry Markovits. The curious effect of using drawings in conditional reasoning problems. Educational Studies in Mathematics, 17(1):81–87, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302380.

Mariotti:2000:IPM

[Mar00] Maria Alessandra Mariotti. Introduction to proof: The mediation of a dynamic software environment. Educational Studies in Mathematics, 44(1–3):25–53, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012733122556.

Martin:2013:DBE

[Mar13] Jason Martin. Differences between experts’ and students’ conceptual images of the mathematical structure of Taylor series convergence. Educational Studies in Mathematics, 82(2): 267–283, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9425-7.

Maslova:1969:DIC

[Mas69] G. G. Maslova. Le développement des idées et des concepts mathématiques fondamentaux dans l’enseignement des enfants de 7 à 15 ans. (French) [The development of fundamental mathematical ideas and concepts in the teaching of children aged 7 to 15]. Educational Studies in Mathematics, 2(2–3): 201–211, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303458.

Matthews:1981:IS

[Mat81] Julia Matthews. An investigation into subtraction. Educational Studies in Mathematics, 12(3):327–338, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311063.

Matthews:1983:SES

[Mat83] Julia Matthews. A subtraction experiment with six and seven year old children. Educational Studies in Mathematics, 14 (2):139–154, May 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303683.

Matthews:2003:BOC

[Mat03] Lou Edward Matthews. Babies overboard! The complexities of incorporation culturally relevant teaching into mathematics instruction. Educational Studies in Mathematics, 53 (1):61–82, May 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1024601504028.

Maury:1985:IQE

[Mau85] Sylvette Maury. Influence de la question dans une épreuve relative à la notion d’indépendance. (French) [Influence of the question in a test relating to the notion of independence]. Educational Studies in Mathematics, 16(3):283–301, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00776734.

Maschietto:2009:WAG

[MB09] Michela Maschietto and Maria G. Bartolini Bussi. Working with artefacts: gestures, drawings and speech in the construction of the mathematical meaning of the visual pyramid. Educational Studies in Mathematics, 70(2):143–157, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9162-0.

Magargee:2016:DEN

[MB16] Suzanne D. Magargee and Judith E. Beauford. Do explicit number names accelerate pre-kindergarteners’ numeracy and place value acquisition? Educational Studies in Mathematics, 92(2):179–192, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9682-y.

Mithalal:2019:IDL

[MB19] Joris Mithalal and Nicolas Balacheff. The instrumental deconstruction as a link between drawing and geometrical figure. Educational Studies in Mathematics, 100(2):161–176, February 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9862-z.

Macchioni:2025:WWG

[MBF25] Elena Macchioni and Anna Baccaglini-Frank. When the whole is greater than the sum of its parts: towards a characterization of students’ mathematics discourse participation profiles.

Educational Studies in Mathematics, 120(1):1–31, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10410-3.

Mesa:2021:WRP

[MBV21] Vilma Mesa, Arthur Bakker, and Wim Van Dooren. Writing reviews: perspectives from the editors of Educational Studies in Mathematics. Educational Studies in Mathematics, 108(3): 419–428, November 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10114-4.

McClain:2001:SSA

[MC01] Kay McClain and Paul Cobb. Supporting students’ ability to reason about data. Educational Studies in Mathematics, 45(1– 3):103–129, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1013874514650.

Margolinas:2005:WCT

[MCB05] Claire Margolinas, Lalina Coulange, and Annie Bessot. What can the teacher learn in the classroom? Educational Studies in Mathematics, 59(1–3):205–234, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-3135-3.

McCloskey:2014:PRL

[McC14] Andrea McCloskey. The promise of ritual: a lens for understanding persistent practices in mathematics classrooms. Educational Studies in Mathematics, 86(1):19–38, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9520-4.

McDonald:1989:ASC

[McD89] Janet L. McDonald. Accuracy and stability of cognitive structures and retention of geometric content. Educational Studies in Mathematics, 20(4):425–448, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00315609.

Maass:2019:DWI

[MCKP19] Katja Maass, Paul Cobb, Konrad Krainer, and Despina Potari. Different ways to implement innovative teaching approaches at scale. Educational Studies in Mathematics, 102 (3):303–318, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09920-
8;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09920-8.pdf.

Moler:2025:SLP

[MCL25] José Antonio Moler, Valentina Celi, and Inmaculada Lizasoain. Shedding light on preservice and inservice school teachers’ beliefs and attitudes regarding mathematics. Educational Studies in Mathematics, 120(2):341–369, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10427-8.

McNair:2000:WMF

[McN00] Rodney E. McNair. Working in the mathematics frame: Maximizing the potential to learn from students’ mathematics classroom discussions. Educational Studies in Mathematics, 42(2): 197–209, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004143014409.

Michael–Chrysanthou:2024:ESS

[MCPGE24] Paraskevi Michael-Chrysanthou, Areti Panaoura, Athanasios Gagatsis, and Iliada Elia. Exploring secondary school students’ geometrical figure apprehension: cognitive structure and levels of geometrical ability. Educational Studies in Mathematics, 117(1):23–42, September 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10317-5.

McQualter:1986:BMT

[McQ86] J. W. McQualter. Becoming a mathematics teacher: The use of personal construct theory. Educational Studies in Mathematics, 17(1):1–14, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302375.

McVicar:2024:LSB

[McV24] Elzena L. McVicar. The liberatory stances of Black women mathematics teachers. Educational Studies in Mathematics, 116(3):519–538, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10280-7.

Maracci:2013:SBT

[MCVM13] Mirko Maracci, Claire Cazes, Fabrice Vandebrouck, and Maria Alessandra Mariotti. Synergies between theoretical approaches to mathematics education with technology: A case study through a cross-analysis methodology. Educational Studies in Mathematics, 84(3):461–485, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9495-1.

Makramalla:2025:MES

[MCW25] Mariam Makramalla, Alf Coles, and David Wagner. Mathematics education for sustainable futures: a strengths-based survey of the field to invite further research action. Educational Studies in Mathematics, 119(3):535–556, July 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10389-x.

Mamona-Downs:2001:LIB

[MD01] Joanna Mamona-Downs. Letting the intuitive bear on the formal; a didactical approach for the understanding of the limit of a sequence. Educational Studies in Mathematics, 48(2–3): 259–288, November 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1016004822476.

Mamona-Downs:2004:RTP

[MDD04] Joanna Mamona-Downs and Martin Downs. Realization of techniques in problem solving: The construction of bijections for enumeration tasks. Educational Studies in Mathematics, 56 (2–3):235–253, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040414.
92348.08.

Masingila:1996:MLP

[MDPW96] Joanna O. Masingila, Susana Davidenko, and Ewa Prus-Wisniowska. Mathematics learning and practice in and out of school: A framework for connecting these experiences. Educational Studies in Mathematics, 31(1–2):175–200, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00143931.

Makur:2025:BRE

[MDU^{+}25] Alberta Parinters Makur, Darhim, Patrisius Afrisno Udil, Magdalena Wangge, and Ofirenty Elyada Nubatonis. Book review: Exploring Piaget’s genetic epistemology as a theoretical foundation in mathematics education. Paul Christian Dawkins, Amy J. Hackenberg, & Anderson Norton (Eds.) (2024) Piaget’s genetic epistemology for mathematics education research. Educational Studies in Mathematics, 119 (1):163–175, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10353-1.

Meaney:2013:WRM

[ME13] Tamsin Meaney and Deb Evans. What is the responsibility of mathematics education to the indigenous students that it serves? Educational Studies in Mathematics, 82(3):481–496, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9439-1.

Melhuish:2020:GTS

[MEH20] Kathleen Melhuish, Brittney Ellis, and Michael D. Hicks. Group theory students’ perceptions of binary operation. Educational Studies in Mathematics, 103(1):63–81, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09925-3.

Menger:1971:GRM

[Men71] Karl Menger. The geometry relevant to modern education. Educational Studies in Mathematics, 4(1):1–17, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305793.

Mendes:2007:NLB

[Men07] Jackeline Rodrigues Mendes. Numeracy and literacy in a bilingual context: Indigenous teachers education in Brazil. Educational Studies in Mathematics, 64(2):217–230, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9009-x.

Menghini:2020:BRD

[Men20] Marta Menghini. Book review: Digging into the memory of mathematics education. Fulvia Furinghetti, Alexander Karp (Eds.) (2018). Researching the history of mathematics education. An international overview. Educational Studies in Mathematics, 103(3):383–389, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09940-
9.

Menghini:2023:BRL

[Men23] Marta Menghini. Book review: Learning from the history of mathematics education. Alexander Karp (Ed.) (2022). Advances in the history of mathematics education. Educational Studies in Mathematics, 114(2):359–365, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10251-y.

Mesquita:1989:SED

[Mes89] A. L. Mesquita. Sur une situation d’éveil à la déduction en géométrie. (French) [On a situation of awakening in deduction in geometry]. Educational Studies in Mathematics, 20(1): 55–77, February 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00356041.

Mesa:2004:CPA

[Mes04] Vilma Mesa. Characterizing practices associated with functions in middle school textbooks: An empirical approach. Educational Studies in Mathematics, 56(2–3):255–286, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000040409.63571.56.

Mevarech:1983:DSM

[Mev83] Zemira R. Mevarech. A deep structure model of students’ statistical misconceptions. Educational Studies in Mathematics, 14(4):415–429, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368237.

Meyer:2010:ALV

[Mey10] Michael Meyer. Abduction — a logical view for investigating and initiating processes of discovering mathematical coherences. Educational Studies in Mathematics, 74(2):185–205, June 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9233-x.

Mariotti:1997:DCA

[MF97] Maria Alessandra Mariotti and Efraim Fischbein. Defining in classroom activities. Educational Studies in Mathematics, 34(3):219–248, December 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1002985109323.

Malaspina:2010:RIS

[MF10] Uldarico Malaspina and Vicenç Font. The role of intuition in the solving of optimization problems. Educational Studies in Mathematics, 75(1):107–130, September 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9243-8.

Markovits:2017:TML

[MF17] Zvia Markovits and Helen Forgasz. “Mathematics is like a lion”: Elementary students’ beliefs about mathematics. Educational Studies in Mathematics, 96(1):49–64, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9759-2.

Meaney:2022:UDA

[MFG22] Tamsin Meaney, A. B. Fyhn, and S. R. W. Graham. Unfettering discussions about social justice: the role of conversational prompts in discussions about mathematics education for indigenous students. Educational Studies in Mathematics, 109 (3):549–565, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10089-2.

Miyazaki:2017:SUS

[MFJ17] Mikio Miyazaki, Taro Fujita, and Keith Jones. Students’ understanding of the structure of deductive proof. Educational Studies in Mathematics, 94(2):223–239, February 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-016-9720-9.pdf.

Manouchehri:2000:IMR

[MG00a] Azita Manouchehri and Terry Goodman. Implementing mathematics reform: The challenge within. Educational Studies in Mathematics, 42(1):1–34, April 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004011522216.

Marrades:2000:PPS

[MG00b] Ramón Marrades and Ángel Gutiérrez. Proofs produced by secondary school students learning geometry in a dynamic computer environment. Educational Studies in Mathematics, 44(1–3):87–125, December 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012785106627.

Mesa:2012:TMT

[MG12] Vilma Mesa and Brett Griffiths. Textbook mediation of teaching: an example from tertiary mathematics instructors. Educational Studies in Mathematics, 79(1):85–107, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9339-9.

Melhuish:2023:RPC

[MGD^{+}23] Kathleen Melhuish, Lino Guajardo, Paul C. Dawkins, Holly Zolt, and Kristen Lew. The role of the partitioning and coset algorithm quotient group partial meanings in comprehending the First Isomorphism Theorem and its proof. Educational Studies in Mathematics, 113(3):499–517, July 2023.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10207-2.

Martino:2023:TSUb

[MGI23] Pietro Di Martino, Francesca Gregorio, and Paola Iannone. The transition from school to university in mathematics education research: new trends and ideas from a systematic literature review. Educational Studies in Mathematics, 113(1): 7–34, May 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10194-w.

Mackenzie:1975:LGL

[MGP75] D. E. Mackenzie, J. D. Gray, and S. J. Prokhovnik. Large-group lecturing in mathematics. Educational Studies in Mathematics, 6(3):293–309, November 1975. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01793613.

MacKay:2025:NMU

[MGV^{+}25] Kelsey J. MacKay, Filip Germeys, Wim Van Dooren, Lieven Verschaffel, and Koen Luwel. Numerical magnitude understanding of natural and rational numbers in secondaryschool students: a number line estimation study. Educational Studies in Mathematics, 118(3):429–448, March 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10291-4.

Movshovitz-Hadar:1989:BR

[MH89] Nitsa Movshovitz-Hadar. Book review. Educational Studies in Mathematics, 20(1):111–119, February 1989. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00356045.

Markovits:1997:RAT

[MH97] Zvia Markovits and Rina Hershkowitz. Relative and absolute thinking in visual estimation processes. Educational Studies in Mathematics, 32(1):29–47, January 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002911812669.

Mullen:2024:SLR

[MHC24] Claire Mullen, Emma Howard, and Anthony Cronin. A scoping literature review of the impact and evaluation of mathematics and statistics support in higher education. Educational Studies in Mathematics, 117(1):1–22, September 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10332-6.

Morrone:2004:PID

[MHDC04] Anastasia Steffen Morrone, Shelly S. Harkness, Beatriz D’Ambrosio, and Richard Caulfield. Patterns of instructional discourse that promote the perception of mastery goals in a social constructivist mathematics course. Educational Studies in Mathematics, 56(1):19–38, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000028401.51537.a5.

Movshovitz-Hadar:1990:PEM

[MHH90] Nitsa Movshovitz-Hadar and Rina Hadass. Preservice education of math teachers using paradoxes. Educational Studies in Mathematics, 21(3):265–287, June 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305093.

Molyneux-Hodgson:1999:MMI

[MHRSU99] Susan Molyneux-Hodgson, Teresa Rojano, Rosamund Sutherland, and Sonia Ursini. Mathematical modelling: the interaction of culture and practice. Educational Studies in Mathematics, 39(1–3):167–183, June 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003758310286.

Miezys:2023:CPT

[Mie23] Vytautas Miežys. Cheap plot tricks and plot holes in mathematical stories. Educational Studies in Mathematics, 113 (2):271–285, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10216-1.

Minato:1983:SMA

[Min83] Saburoh Minato. Some mathematical attitudinal data on eighth grade students in Japan measured by a semantic differential. Educational Studies in Mathematics, 14(1):19–38, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00704700.

Mitchelmor:1980:TDG

[Mit80] Michael C. Mitchelmor. Three-dimensional geometrical drawing in three cultures. Educational Studies in Mathematics, 11 (2):205–216, May 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00304356.

Miyazaki:2000:LPL

[Miy00] Mikio Miyazaki. Levels of proof in lower secondary school mathematics. Educational Studies in Mathematics, 41(1):47– 68, January 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003956532587.

Miyakawa:2017:CAN

[Miy17] Takeshi Miyakawa. Comparative analysis on the nature of proof to be taught in geometry: the cases of French and Japanese lower secondary schools. Educational Studies in Mathematics, 94(1):37–54, January 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9711-x.

Misfeldt:2015:RMP

[MJ15] Morten Misfeldt and Mikkel Willum Johansen. Research mathematicians’ practices in selecting mathematical problems. Educational Studies in Mathematics, 89(3):357–373, July 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9605-3.

Matthews:2021:LUP

[MJS21] Lou E. Matthews, Naomi A. Jessup, and Ruthmae Sears. Looking for “us”: power reimagined in mathematics learning for Black communities in the pandemic. Educational Studies in Mathematics, 108(1–2):333–350, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10106-4.

Maqsud:1991:RSS

[MK91] Muhammad Maqsud and Chaudhry M. Khalique. Relationships of some socio-personal factors to mathematics achievement of secondary school and university students in Bophuthatswana. Educational Studies in Mathematics, 22(4): 377–390, August 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00369296.

Mevarech:1997:VDG

[MK97] Zemira R. Mevarech and Bracha Kramarsky. From verbal descriptions to graphic representations: Stability and change in students’ alternative conceptions. Educational Studies in Mathematics, 32(3):229–263, March 1997. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002965907987.

Morgan:2014:DAR

[MK14] Candia Morgan and Chronis Kynigos. Digital artefacts as representations: forging connections between a constructionist and a social semiotic perspective. Educational Studies in Mathematics, 85(3):357–379, March 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9523-1.

Moon:2020:DAM

[ML20] Sung-Jae Moon and Kyeong-Hwa Lee. Deleuzian actualizations of the multiplicative concept: a study of perceptual flow and the transformation of learning assemblages. Educational Studies in Mathematics, 104(2):221–237, June 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09953-4.

Moore:2016:EPU

[MLK16a] Kevin C. Moore, Kevin R. LaForest, and Hee Jung Kim. Erratum to: Putting the unit in pre-service secondary teachers’ unit circle. Educational Studies in Mathematics, 92 (2):243, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-016-9680-0.pdf. See

[MLK16b].

Moore:2016:PUP

[MLK16b] Kevin C. Moore, Kevin R. LaForest, and Hee Jung Kim. Putting the unit in pre-service secondary teachers’ unit circle. Educational Studies in Mathematics, 92(2):221–241, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9671-6. See er-

ratum [MLK16a].

McCloskey:2019:TMI

[MLL19] Andrea McCloskey, Gwendolyn Lloyd, and Courtney Lynch. Theorizing mathematics instruction using ritual: tensions in teaching fractions in a fifth grade classroom. Educational Studies in Mathematics, 101(2):195–213, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9779-y.

Molad:2020:IGM

[MLL20] Osnat Molad, Esther S. Levenson, and Sigal Levy. Individual and group mathematical creativity among post-high school students. Educational Studies in Mathematics, 104(2):201– 220, June 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09952-5.

Moskal:2000:MSW

[MM00] Barbara M. Moskal and Maria E. Magone. Making sense of what students know: Examining the referents, relationships and modes students displayed in response to a decimal task. Educational Studies in Mathematics, 43(3):313– 335, November 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1011983602860.

Magajna:2003:AMT

[MM03] Z. Magajna and J. Monaghan. Advanced mathematical thinking in a technological workplace. Educational Studies in Mathematics, 52(2):101–122, March 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024089520064.

Manizade:2011:UDM

[MM11] Agida Gabil Manizade and Marguerite M. Mason. Using Delphi methodology to design assessments of teachers’ pedagogical content knowledge. Educational Studies in Mathematics, 76 (2):183–207, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-010-9276-z.

Maffia:2020:ASG

[MM20] Andrea Maffia and Maria Alessandra Mariotti. From action to symbols: giving meaning to the symbolic representation of the distributive law in primary school. Educational Studies in Mathematics, 104(1):25–40, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09944-5.

Mathe:2025:FRS

[MM25] Anne-Cécile Mathé and Joris Mithalal. Figure reproduction as a step towards theoretical geometry: analysis of the didactical and a-didactical processes in a classroom setting. Educational Studies in Mathematics, 119(2):203–223, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10370-0.

Martin:2005:ITS

[MMBD05] Tami S. Martin, Sharon M. Soucy McCrone, Michelle L. Wallace Bower, and Jaguthsing Dindyal. The interplay of teacher and student actions in the teaching and learning of geometric proof. Educational Studies in Mathematics, 60(1):95–124, September 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-6698-0.

Maher:2022:ASS

[MMC22] Nicole Maher, Tracey Muir, and Helen Chick. Analysing senior secondary mathematics teaching using the Knowledge Quartet. Educational Studies in Mathematics, 110(2):233– 249, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10125-1.

Maffia:2025:TMA

[MMM25] Andrea Maffia, Carola Manolino, and Elisa Miragliotta. There is more to algebra than meets the eye: the case of blindness. Educational Studies in Mathematics, 119(1):63–77, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10394-0.

Meletiou-Mavrotheris:2015:DSR

[MMP15] Maria Meletiou-Mavrotheris and Efi Paparistodemou. Developing students’ reasoning about samples and sampling in the context of informal inferences. Educational Studies in Mathematics, 88(3):385–404, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9551-5.

Mejias:2021:LSG

[MMR21] Paulina Perez Mejias, Dora Elias McAllister, and Javiera Ravest. A longitudinal study of the gender gap in mathematics achievement: evidence from Chile. Educational Studies in Mathematics, 107(3):583–605, July 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10052-1.

Moreira:1995:UTA

[MN95] Cândida Moreira and Richard Noss. Understanding teachers’ attitudes to change in a logomathematics environment. Educational Studies in Mathematics, 28(2):155–176, March 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01295791.

Marfuah:2023:BRS

[MND23] Marfuah Marfuah, Rita Novita, and Sri Wulandari Danoebroto. Book review: Supporting mathematics educators through and towards online learning. Karen Hollebrands, Robin Anderson, and Kevin Oliver (Eds). (2021) Online learning in mathematics education. Educational Studies in Mathematics, 112(2):383–395, February 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10175-z.

McIntosh:1995:MCP

[MNRR95] Alistair McIntosh, Nobuhiko Nohda, Barbara J. Reys, and Robert E. Reys. Mental computation performance in Australia, Japan and the United States. Educational Studies in Mathematics, 29(3):237–258, October 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274093.

Mellin-Olsen:1981:IEC

[MO81] Stieg Mellin-Olsen. Instrumentalism as an educational concept. Educational Studies in Mathematics, 12(3):351–367, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311065.

Morony:1994:SIA

[MO94] Will Morony and Kevin Olssen. Support for informal assessment in mathematics in the context of standards referenced reporting. Educational Studies in Mathematics, 27(4):387– 399, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273379.

Monaghan:2006:AC

[MO06] John Monaghan and Mehmet Fatih Ozmantar. Abstraction and consolidation. Educational Studies in Mathematics, 62 (3):233–258, July 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-8753-x.

Mogari:2017:UCR

[Mog17] David Mogari. Using culturally relevant teaching in a coeducational mathematics class of a patriarchal community. Educational Studies in Mathematics, 94(3):293–307, March 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9730-7.

Monna:1972:ST

[Mon72] A. F. Monna. Set-theory in 1888. Educational Studies in Mathematics, 4(3):393–397, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302586.

Monaghan:2000:WDD

[Mon00] Frank Monaghan. What difference does it make? Children’s views of the differences between some quadrilaterals. Educational Studies in Mathematics, 42(2):179–196, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004175020394.

Monaghan:2001:YPI

[Mon01] John Monaghan. Young peoples’ ideas of infinity. Educational Studies in Mathematics, 48(2–3):239–257, November 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016090925967.

Moore:1994:MTF

[Moo94] Robert C. Moore. Making the transition to formal proof. Educational Studies in Mathematics, 27(3):249–266, October 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273731.

Moore:2013:MSM

[Moo13] Kevin C. Moore. Making sense by measuring arcs: a teaching experiment in angle measure. Educational Studies in Mathematics, 83(2):225–245, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9450-6.

Moon:2020:NAT

[Moo20] Kyunghee Moon. New approaches for two-variable inequality graphs utilizing the Cartesian Connection and the APOS theory. Educational Studies in Mathematics, 104(3):351–367, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09956-1.

Morgan:2000:BRL

[Mor00] Candia Morgan. Book review: Language in Use in Mathematics Classrooms: Developing Approaches to a Research Domain, H. Steinbring, M. G. Bartolini Bussi and A. Sierpińska. Educational Studies in Mathematics, 41(1):93–99, January 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003891809328.

Morgan:2006:WDS

[Mor06] Candia Morgan. What does social semiotics have to offer mathematics education research? Educational Studies in Mathematics, 61(1–2):219–245, February 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-5477-x.

Morgan:2007:WMN

[Mor07]  Candia Morgan. Who is not multilingual now?  Ed-

ucational Studies in Mathematics, 64(2):239–242, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9064-y.

Morgan:2014:STM

[Mor14a] Candia Morgan. Social theory in mathematics education: Guest editorial. Educational Studies in Mathematics, 87(2): 123–128, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-014-9572-0.pdf.

Morgan:2014:UPM

[Mor14b] Candia Morgan. Understanding practices in mathematics education: structure and text. Educational Studies in Mathematics, 87(2):129–143, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9482-6.

Morris:2021:LPK

[Mor21] Noah Morris. Learning probability in the Kingdom of Tonga: the influence of language and culture. Educational Studies in Mathematics, 107(1):111–134, May 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10022-z.

Moschkovich:1998:SUI

[Mos98] Judit N. Moschkovich. Students’ use of the x-intercept as an instance of a transitional conception. Educational Studies in Mathematics, 37(2):169–197, November 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003539828299.

Moschkovich:2004:AMP

[Mos04] Judit N. Moschkovich. Appropriating mathematical practices: A case study of learning to use and explore functions through interaction with a tutor. Educational Studies in Mathematics, 55(1–3):49–80, March 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017691.13428.b9.

Moseley:2005:SEM

[Mos05] Bryan Moseley. Students’ early mathematical representation knowledge: The effects of emphasizing single or multiple perspectives of the rational number domain in problem solving. Educational Studies in Mathematics, 60(1):37–69, September 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5031-2.

Moschkovich:2007:UTL

[Mos07] Judit Moschkovich. Using two languages when learning mathematics. Educational Studies in Mathematics, 64(2):121–144, February 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-9005-1.

Motte:1971:MAD

[Mot71] Magdeleine Motte. En marge de l’apprentissage de la déduction. (French) [On the margin of learning deduction]. Educational Studies in Mathematics, 4(2):232–242, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303388.

Moyer:2001:WHF

[Moy01] Patricia S. Moyer. Are we having fun yet? How teachers use manipulatives to teach mathematics. Educational Studies in Mathematics, 47(2):175–197, July 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014596316942.

Mason:1984:GES

[MP84] John Mason and David Pimm. Generic examples: Seeing the general in the particular. Educational Studies in Mathematics, 15(3):277–289, August 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312078.

McLeay:1996:MMD

[MP96] Heather McLeay and David Piggins. The mental manipulation of 2-D representations of knots as deformable structures. Educational Studies in Mathematics, 30(4):399–414, June 1996. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00570831.

Martinez:2014:RBI

[MP14] Mara V. Martinez and Bettina Pedemonte. Relationship between inductive arithmetic argumentation and deductive algebraic proof. Educational Studies in Mathematics, 86(1): 125–149, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9530-2.

Marco:2024:TPD

[MP24] Nadav Marco and Alik Palatnik. Teachers pose and design context-based mathematics tasks: what can be learned from product evolution? Educational Studies in Mathematics, 115 (2):223–246, February 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10271-8.

Mata-Pereira:2017:ESM

[MPdP17] Joana Mata-Pereira and João-Pedro da Ponte. Enhancing students’ mathematical reasoning in the classroom: teacher actions facilitating generalization and justification. Educational Studies in Mathematics, 96(2):169–186, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9773-4.

Martinez-Planell:2012:SUG

[MPG12] Rafael Martínez-Planell and María Trigueros Gaisman. Students’ understanding of the general notion of a function of two variables. Educational Studies in Mathematics, 81(3):365– 384, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9408-8.

Martinez-Planell:2012:SCI

[MPGDA12] Rafael Martínez-Planell, Ana Carmen Gonzalez, Gladys Di-Cristina, and Vanessa Acevedo. Students’ conception of infinite series. Educational Studies in Mathematics, 81(2):235– 249, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9401-2.

Maheux:2012:BRR

[MPS12] Jean-François Maheux, Jérôme Proulx, and Elaine Simmt. Book review: A review of the book by SungWon Hwang and Wolff–Michael Roth (2011). Scientific & mathematical bodies: the interface of culture and mind. Educational Studies in Mathematics, 81(2):279–285, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9379-1.

Marco:2022:WLM

[MPS22] Nadav Marco, Alik Palatnik, and Baruch B. Schwarz. When is less more? Investigating gap-filling in proofs without words activities. Educational Studies in Mathematics, 111(2):271–297, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10159-z.

Metaxas:2016:ATP

[MPZ16] N. Metaxas, D. Potari, and T. Zachariades. Analysis of a teacher’s pedagogical arguments using Toulmin’s model and argumentation schemes. Educational Studies in Mathematics, 93(3):383–397, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9701-z.

Meaney:2017:BRH

[MR17] Tamsin Meaney and Toril Eskeland Rangnes. Book review: How research fields change — the documentation of a process. Anjum Halai and Philip Clarkson (Eds.) (2016) Teaching and learning mathematics in multilingual classrooms. Educational Studies in Mathematics, 95(2):219–227, June 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Moutsios-Rentzos:2020:BRA

[MR20] Andreas Moutsios-Rentzos. Book review: Andreas J. Stylianides (2016), Proving in the elementary mathematics classroom. Educational Studies in Mathematics, 103(3):375–381, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09936-5.

Moyer:2018:AHS

[MRC18] John C. Moyer, Victoria Robison, and Jinfa Cai. Attitudes of high-school students taught using traditional and reform mathematics curricula in middle school: a retrospective analysis. Educational Studies in Mathematics, 98(2):115–134, June 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9809-4.

Mellone:2021:ECT

[MRC21] Maria Mellone, Alessandro Ramploud, and Gemma Carotenuto. An experience of cultural transposition of the El’konin– Davydov curriculum. Educational Studies in Mathematics, 106(3):379–396, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-09942-7.

Moore-Russo:2011:CSN

[MRCR11] Deborah Moore-Russo, AnnaMarie Conner, and Kristina I. Rugg. Can slope be negative in 3-space? Studying concept image of slope through collective definition construction. Educational Studies in Mathematics, 76(1):3–21, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9277-y.

Mejia-Ramos:2012:AMP

[MRFW^{+}12] Juan Pablo Mejia-Ramos, Evan Fuller, Keith Weber, Kathryn Rhoads, and Aron Samkoff. An assessment model for proof comprehension in undergraduate mathematics. Educational Studies in Mathematics, 79(1):3–18, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9349-7.

Morton:1994:CST

[MRRF94] Margaret Morton, Barbara Reilly, Elizabeth Robinson, and Sharleen Forbes. A comparative study of two nationwide examinations: Maths with calculus and maths with statistics. Educational Studies in Mathematics, 26(4):367–387, June 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01279521.

Moutsios-Rentzos:2014:ORA

[MRSP14] Andreas Moutsios-Rentzos, Panagiotis Spyrou, and Alexandra Peteinara. The objectification of the right-angled triangle in the teaching of the Pythagorean theorem: an empirical investigation. Educational Studies in Mathematics, 85(1): 29–51, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9498-y.

Mejia-Ramos:2014:WHM

[MRW14] Juan Pablo Mejia-Ramos and Keith Weber. Why and how mathematicians read proofs: further evidence from a survey study. Educational Studies in Mathematics, 85(2):161–173, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9514-2.

MacGregor:1997:SUA

[MS97] Mollie MacGregor and Kaye Stacey. Students’ understanding of algebraic notation: 11–15. Educational Studies in Mathematics, 33(1):1–19, June 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002970913563.

Mason:1999:BMK

[MS99] John Mason and Mary Spence. Beyond mere knowledge of mathematics: The importance of knowing-to act in the moment. Educational Studies in Mathematics, 38(1–3):135–161, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003622804002.

Malisani:2009:ATA

[MS09] Elsa Malisani and Filippo Spagnolo. From arithmetical thought to algebraic thought: The role of the “variable”. Educational Studies in Mathematics, 71(1):19–41, May 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9157-x.

McDonough:2014:SIY

[MS14] Andrea McDonough and Peter Sullivan. Seeking insights into young children’s beliefs about mathematics and learning. Educational Studies in Mathematics, 87(3):279–296, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9565-z.

Maciejewski:2019:JCM

[MS19] Wes Maciejewski and Jon R. Star. Justifications for choices made in procedures. Educational Studies in Mathematics, 101 (3):325–340, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-019-09886-7.

Meyer:2020:WCG

[MS20] Michael Meyer and Susanne Schnell. What counts as a “good” argument in school?— how teachers grade students’ mathematical arguments. Educational Studies in Mathematics, 105(1):35–51, September 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09974-
z;   http://link.springer.com/content/pdf/10.1007/
s10649-020-09974-z.pdf.

Martinez-Sierra:2016:UMS

[MSdSGG16] Gustavo Martínez-Sierra and María del Socorro García-González. Undergraduate mathematics students’ emotional experiences in linear algebra courses. Educational Studies in Mathematics, 91(1):87–106, January 2016. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9634-y.

Martin:2015:GMU

[MT15] Lyndon C. Martin and Jo Towers. Growing mathematical understanding through collective image making, collective image having, and collective property noticing. Educational Studies in Mathematics, 88(1):3–18, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9552-4.

Moll:2016:MOT

[MTBR16] Vicenç Font Moll, María Trigueros, Edelmira Badillo, and Norma Rubio. Mathematical objects through the lens of two different theoretical perspectives: APOS and OSA. Educational Studies in Mathematics, 91(1):107–122, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9639-6.

Mura:1993:IMH

[Mur93] Roberta Mura. Images of mathematics held by university teachers of mathematical sciences. Educational Studies in Mathematics, 25(4):375–385, December 1993. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273907.

Mura:1995:IMH

[Mur95] Roberta Mura. Images of mathematics held by university teachers of mathematics education. Educational Studies in Mathematics, 28(4):385–399, June 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274080.

Murphy:2004:HDC

[Mur04] Carol Murphy. How do children come to use a taught mental calculation strategy? Educational Studies in Mathematics, 56(1):3–18, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000028403.
25781.79.

Murray:2024:BRA

[Mur24] Eileen Murray. Book review: Alan Schoenfeld, Heather Fink, Sandra Zuñiga–Ruiz, Siqi Huang, Xinyu Wei, and Brantina Chirinda (2023) Helping students become powerful thinkers: case studies of teaching for robust understanding. Educational Studies in Mathematics, 117(3):565–575, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10312-w.

Martin:2024:RRR

[MVCV24] Danny Bernard Martin, Luz Valoyes-Chávez, and Paola Valero. Race, racism, and racialization in mathematics education: global perspectives. Educational Studies in Mathematics, 116(3):313–331, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10346-0.

Magiera:2013:ESP

[MvdKM13] Marta T. Magiera, Leigh A. van den Kieboom, and John C. Moyer. An exploratory study of pre-service middle school teachers’ knowledge of algebraic thinking. Educational Studies in Mathematics, 84(1):93–113, September 2013. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9472-8.

Mitchelmore:2000:DAC

[MW00] Michael C. Mitchelmore and Paul White. Development of angle concepts by progressive abstraction and generalisation. Educational Studies in Mathematics, 41(3):209–238, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003927811079.

Miyakawa:2009:DDS

[MW09] Takeshi Miyakawa and Carl Winsløw. Didactical designs for students’ proportional reasoning: an “open approach” lesson and a “fundamental situation”. Educational Studies in Mathematics, 72(2):199–218, November 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9188-y.

Mesa:2019:BDC

[MW19] Vilma Mesa and David Wagner. Behind the door: a critical look at the process of publication in Educational Studies in Mathematics. Educational Studies in Mathematics, 101(3):301–324, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09900-
y;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09900-y.pdf.

Mwakapenda:2002:SCC

[Mwa02] Willy Mwakapenda. The status and context of change in mathematics education in Malawi. Educational Studies in Mathematics, 49(2):251–281, February 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016283520850.

Mao:2017:APU

[MWSS17] Yi Mao, Tyreke White, Philip M. Sadler, and Gerhard Sonnert. The association of precollege use of calculators with student performance in college calculus. Educational Studies in Mathematics, 94(1):69–83, January 2017.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9714-7.

Montiel:2009:USA

[MWVE09] Mariana Montiel, Miguel R. Wilhelmi, Draga Vidakovic, and Iwan Elstak. Using the onto-semiotic approach to identify and analyze mathematical meaning when transiting between different coordinate systems in a multivariate context. Educational Studies in Mathematics, 72(2):139–160, November 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9184-2.

Minato:1984:RBS

[MY84] Saburoh Minato and Shyoichi Yanase. On the relationship between students attitudes towards school mathematics and their levels of intelligence. Educational Studies in Mathematics, 15(3):313–320, August 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00312080.

Mueller:2012:FAC

[MYM12] Mary Mueller, Dina Yankelewitz, and Carolyn Maher. A framework for analyzing the collaborative construction of arguments and its interplay with agency. Educational Studies in Mathematics, 80(3):369–387, July 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9354-x.

Modabbernia:2023:WAE

[MYZ23] Niusha Modabbernia, Xiaoheng Yan, and Rina Zazkis. When algebra is not enough: a dialogue on the composition of even and odd functions. Educational Studies in Mathematics, 112 (3):397–414, March 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10189-7.

Mamolo:2012:SCS

[MZ12] Ami Mamolo and Rina Zazkis. Stuck on convention: a story of derivative relationships. Educational Studies in Mathematics, 81(2):161–177, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9391-0.

Marmur:2018:SFA

[MZ18] Ofer Marmur and Rina Zazkis. Space of fuzziness: avoidance of deterministic decisions in the case of the inverse function. Educational Studies in Mathematics, 99(3):261–275, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9843-2.

Marmur:2021:IGS

[MZ21] Ofer Marmur and Rina Zazkis. Irrational gap: sensemaking trajectories of irrational exponents. Educational Studies in Mathematics, 107(1):25–48, May 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10027-2.

Nardi:2000:MUR

[Nar00]  Elena Nardi.  Mathematics undergraduates’ responses

to semantic abbreviations, ‘geometric’ images and multilevel abstractions in group theory. Educational Studies in Mathematics, 43(2):169–189, September 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012223826388.

Nardi:2016:WFS

[Nar16] Elena Nardi. Where form and substance meet: using the narrative approach of re-storying to generate research findings and community rapprochement in (university) mathematics education. Educational Studies in Mathematics, 92(3):361–377, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9643-x.

Nieminen:2023:DRH

[NBA23] Juuso Henrik Nieminen, Anette Bagger, and Julie Allan. Discourses of risk and hope in research on mathematical learning difficulties. Educational Studies in Mathematics, 112(2):

337–357, February 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10204-x.

Nemirovsky:2004:PSI

[NBD^{+}04] Ricardo Nemirovsky, Marcelo Borba, Cara Dimattia, Ferdinando Arzarello, Ornella Robutti, Marty Schnepp, Daniel Chazan, Chris Rasmussen, Jennifer Olszewski, Kevin Dost, James L. Johnson, Marcelo C. Borba, and Nilce Fatima Scheffer. PME special issue: Bodily activity and imagination in mathematics learning. Educational Studies in Mathematics, 57(3):303–321, November 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-004-
5933-4.

Nunes:2012:TCH

[NBE^{+}12] Terezinha Nunes, Peter Bryant, Deborah Evans, Daniel Bell, and Rossana Barros. Teaching children how to include the inversion principle in their reasoning about quantitative relations. Educational Studies in Mathematics, 79(3):371–388, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9314-5.

Noss:2007:SGW

[NBHK07] Richard Noss, Arthur Bakker, Celia Hoyles, and Phillip Kent. Situating graphs as workplace knowledge. Educational Studies in Mathematics, 65(3):367–384, July 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9058-9.

Nardi:2012:WRI

[NBZ12] Elena Nardi, Irene Biza, and Theodossios Zachariades. ‘Warrant’ revisited: Integrating mathematics teachers’ pedagogical and epistemological considerations into Toulmin’s model for argumentation. Educational Studies in Mathematics, 79(2): 157–173, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9345-y.

Nicol:2006:LTM

[NC06] C. C. Nicol and S. M. Crespo. Learning to teach with mathematics textbooks: How preservice teachers interpret and use curriculum materials. Educational Studies in Mathematics, 62 (3):331–355, July 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-5423-y.

Nieminen:2022:WAD

[NCC22] Juuso Henrik Nieminen, Man Ching Esther Chan, and David Clarke. What affordances do open-ended real-life tasks offer for sharing student agency in collaborative problem-solving? Educational Studies in Mathematics, 109(1):115–136, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10074-9.

Ngware:2015:QTM

[NCMO15] Moses W. Ngware, James Ciera, Peter K. Musyoka, and Moses Oketch. Quality of teaching mathematics and learning achievement gains: evidence from primary schools in Kenya. Educational Studies in Mathematics, 89(1):111–131, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9594-2.

Noss:1988:BR

[ND88] Richard Noss and Régine Douady. Book review. Educational Studies in Mathematics, 19(3):403–417, August 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312456.

Nunez:1999:ECG

[NEM99] Rafael E. Núñez, Laurie D. Edwards, and João Filipe Matos. Embodied cognition as grounding for situatedness and context in mathematics education. Educational Studies in Mathematics, 39(1–3):45–65, June 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003759711966.

Nesher:1976:TDD

[Nes76] Pearla Nesher. Three determinants of difficulty in verbal arithmetic problems. Educational Studies in Mathematics, 7(4): 369–388, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00452220.

Neuman:1999:ELA

[Neu99] Dagmar Neuman. Early learning and awareness of division: A phenomenographic approach. Educational Studies in Mathematics, 40(2):101–128, October 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003852815160.

Newstead:1998:ACM

[New98] Karen Newstead. Aspects of children’s mathematics anxiety. Educational Studies in Mathematics, 36(1):53–71, June 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003177809664.

Nemirovsky:2009:MIE

[NF09] Ricardo Nemirovsky and Francesca Ferrara. Mathematical imagination and embodied cognition. Educational Studies in Mathematics, 70(2):159–174, March 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9150-4.

Nemirovsky:2020:BME

[NFFAP20] Ricardo Nemirovsky, Francesca Ferrara, Giulia Ferrari, and Natividad Adamuz-Povedano. Body motion, early algebra, and the colours of abstraction. Educational Studies in Mathematics, 104(2):261–283, June 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-020-09955-
2;   http://link.springer.com/content/pdf/10.1007/
s10649-020-09955-2.pdf.

Ng:2016:IBL

[Ng16] Oi-Lam Ng. The interplay between language, gestures, dragging and diagrams in bilingual learners’ mathematical communications. Educational Studies in Mathematics, 91(3):307–326, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9652-9.

Ng:2021:SSS

[Ng21a] Clarence Ng. Self-schematic students in mathematics and their contrasting coping responses to mathematical challenges: a cluster analytic investigation. Educational Studies in Mathematics, 107(2):315–338, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10031-
6.

Ng:2021:SCP

[Ng21b] Clarence Ng. Subject choice and perezhivanie in mathematics: a longitudinal case study. Educational Studies in Mathematics, 107(3):547–563, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10050-3.

Nesher:1982:DSC

[NGR82] P. Nesher, J. G. Greeno, and M. S. Riley. The development of semantic categories for addition and subtraction. Educational Studies in Mathematics, 13(4):373–394, November 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00366618.

Nguyen:2024:SMC

[Ngu24] Phi Nguyen. School mathematics as context: examining discourses about the subject in district policymaking. Educational Studies in Mathematics, 117(3):485–509, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10348-y.

Nesher:1994:RST

[NH94] Pearla Nesher and Sara Hershkovitz. The role of schemes in two-step problems: Analysis and research findings. Educational Studies in Mathematics, 26(1):1–23, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273298.

Noll:2015:PPM

[NH15] Jennifer Noll and Stacey Hancock. Proper and paradigmatic metonymy as a lens for characterizing student conceptions of distributions and sampling. Educational Studies in Mathematics, 88(3):361–383, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9547-1.

Niss:2019:MCR

[NH19] Mogens Niss and Tomas Højgaard. Mathematical competencies revisited. Educational Studies in Mathematics, 102(1): 9–28, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-019-09903-9.

Noss:1997:CMM

[NHH97] Richard Noss, Lulu Healy, and Celia Hoyles. The construction of mathematical meanings: Connecting the visual with the symbolic. Educational Studies in Mathematics, 33(2):203– 233, July 1997. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1002943821419.

Nesher:2003:SMT

[NHN03] P. Nesher, S. Hershkovitz, and J. Novotna. Situation model, text base and what else? Factors affecting problem solving. Educational Studies in Mathematics, 52(2):151–176, March 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024028430965.

Nagar:2022:HSM

[NHO22] Gili Gal Nagar, Stephen Hegedus, and Chandra Hawley Orrill. High school mathematics teachers’ discernment of invariant properties in a dynamic geometry environment. Educational Studies in Mathematics, 111(1):127–145, September 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10144-6.

Nicol:1998:LTM

[Nic98] Cynthia Nicol. Learning to teach mathematics: Questioning, listening, and responding. Educational Studies in Mathematics, 37(1):45–66, October 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003451423219.

Nicol:2002:WMP

[Nic02] Cynthia Nicol. Where’s the math? Prospective teachers visit the workplace. Educational Studies in Mathematics, 50(3): 289–309, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1021211207232.

Nilsson:2007:DWW

[Nil07] Per Nilsson. Different ways in which students handle chance encounters in the explorative setting of a dice game. Educational Studies in Mathematics, 66(3):293–315, November 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9062-0.

Niman:1975:GTE

[Nim75] John Niman. Graph theory in the elementary school. Educational Studies in Mathematics, 6(3):351–373, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01793617.

Nimier:1977:MAF

[Nim77] Jacques Nimier. Mathématique et affectivité. (French) [Mathematics and affectivity]. Educational Studies in Mathematics, 8(3):241–250, October 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00385924.

Niss:1999:ANS

[Nis99] Mogens Niss. Aspects of the nature and state of research in mathematics education. Educational Studies in Mathematics, 40(1):1–24, September 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003715913784.

Norqvist:2025:SSA

[NJL25] Mathias Norqvist, Bert Jonsson, and Johan Lithner. Shifts in student attention on algorithmic and creative practice tasks. Educational Studies in Mathematics, 118(3):409–428, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10250-z.

Nesher:1977:SAA

[NK77] Pearla Nesher and Tamar Katriel. A semantic analysis of addition and subtraction word problems in arithmetic. Educational Studies in Mathematics, 8(3):251–269, October 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00385925.

Nardi:2017:CCC

[NK17] Elena Nardi and Eric Knuth. Changing classroom culture, curricula, and instruction for proof and proving: how amenable to scaling up, practicable for curricular integration, and capable of producing long-lasting effects are current interventions? Educational Studies in Mathematics, 96(2):267–274, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9785-0.

Nkhoma:2002:WSB

[Nkh02] Pentecost M. Nkhoma. What successful black South African students consider as factors of their success. Educational Studies in Mathematics, 50(1):103–113, May 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020509528995.

Nunes:1998:SAB

[NM98] Terezinha Nunes and Constanza Moreno. The signed algorithm and its bugs. Educational Studies in Mathematics, 35 (1):85–92, January 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003061009907.

Nemirovsky:1997:MVP

[NN97] Ricardo Nemirovsky and Tracy Noble. On mathematical visualization and the place where we live. Educational Studies in Mathematics, 33(2):99–131, July 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002983213048.

Newton:2007:CEM

[NN07] Douglas P. Newton and Lynn D. Newton. Could elementary mathematics textbooks help give attention to reasons in the classroom? Educational Studies in Mathematics, 64(1):69–84, January 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-9015-z.

Noelting:1980:DPRa

[Noe80a] Gerald Noelting. The development of proportional reasoning and the ratio concept. Part I — differentiation of stages. Educational Studies in Mathematics, 11(2):217–253, May 1980. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00304357.

Noelting:1980:DPRb

[Noe80b] Gerald Noelting. The development of proportional reasoning and the ratio concept. Part II — problem-structure at successive stages; problem-solving strategies and the mechanism of adaptive restructuring. Educational Studies in Mathematics, 11(3):331–363, August 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00697744.

Nolan:2011:BRP

[Nol11] Kathleen Nolan. Book review: Pausing to relate and participate. Educational Studies in Mathematics, 76(2):231–235, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9265-2.

Nolan:2012:DFV

[Nol12] Kathleen Nolan. Dispositions in the field: viewing mathematics teacher education through the lens of Bourdieu’s social field theory. Educational Studies in Mathematics, 80(1– 2):201–215, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9355-9.

Nolan:2016:SNM

[Nol16] Kathleen T. Nolan. Schooling novice mathematics teachers on structures and strategies: a Bourdieuian perspective on the role of “others” in classroom practices. Educational Studies in Mathematics, 92(3):315–329, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9668-1.

Noren:2015:APM

[Nor15] Eva Norén. Agency and positioning in a multilingual mathematics classroom. Educational Studies in Mathematics, 89 (2):167–184, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9603-5.

Noss:1986:CCF

[Nos86] Richard Noss. Constructing a conceptual framework for elementary algebra through logo programming. Educational Studies in Mathematics, 17(4):335–357, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311324.

Noss:1988:CCI

[Nos88] Richard Noss. The computer as a cultural influence in mathematical learning. Educational Studies in Mathematics, 19 (2):251–268, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00751235.

Noss:2001:LMD

[Nos01] Richard Noss. For a learnable mathematics in the digital cultures. Educational Studies in Mathematics, 48(1):21–46, October 2001. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1015528224870.

Nowecki:1992:AZK

[Now92] Bogdan J. Nowecki. Anna Zofia Krygowska. Educational Studies in Mathematics, 23(2):123–137, April 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00588052.

Nesher:1986:SR

[NP86] Pearla Nesher and Irit Peled. Shifts in reasoning. Educational Studies in Mathematics, 17(1):67–79, February 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302379.

Noss:1999:TEM

[NPH99] Richard Noss, Stefano Pozzi, and Celia Hoyles. Touching epistemologies: meanings of average and variation in nursing practice. Educational Studies in Mathematics, 40(1):25– 51, September 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003763812875.

Nilsson:2010:FEC

[NR10] Per Nilsson and Andreas Ryve. Focal event, contextualization, and effective communication in the mathematics classroom. Educational Studies in Mathematics, 74(3):241–258, July 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9236-7.

Nieminen:2024:MBI

[NRV24] Juuso Henrik Nieminen, Daniel L. Reinholz, and Paola Valero. “Mathematics is a battle, but I’ve learned to survive”: becoming a disabled student in university mathematics. Educational Studies in Mathematics, 116(1):5–25, May 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10311-x.

Noster:2025:TEE

[NS25] Norbert Noster and Hans-Stefan Siller. Transforming equations into equivalent equations — an empirical study of the equation transformation capability as a multidimensional construct. Educational Studies in Mathematics, 120(2):249–267, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10419-8.

Nesher:1975:VCI

[NT75] Perla Nesher and Eva Teubal. Verbal cues as an interfering factor in verbal problem solving. Educational Studies in Mathematics, 6(1):41–51, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00590023.

Nemirovsky:2001:CCW

[NT01] Ricardo Nemirovsky and Cornelia Tierney. Children creating ways to represent changing situations: On the development of homogeneous spaces. Educational Studies in Mathematics, 45(1–3):67–102, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1013806228763.

Nachlieli:2019:REO

[NT19] Talli Nachlieli and Michal Tabach. Ritual-enabling opportunitiesto-learn in mathematics classrooms. Educational Studies in Mathematics, 101(2):253–271, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9848-
x.

Ntenza:2006:IFC

[Nte06] S. Philemon Ntenza. Investigating forms of children’s writing in grade 7 mathematics classrooms. Educational Studies in Mathematics, 61(3):321–345, March 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-5891-0.

Nucci:2024:ROL

[Nuc24] Drew Nucci. The role of an online learning environment in teacher care for secondary mathematics students. Educational Studies in Mathematics, 117(3):441–455, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10350-4.

Nunes:1995:BR

[Nun95] T. Nunes. Book review. Educational Studies in Mathematics, 29(3):293–294, October 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274095.

Nunokawa:1996:ALT

[Nun96] Kazuhiko Nunokawa. Applying Lakatos’ theory to the theory of mathematical problem solving. Educational Studies in Mathematics, 31(3):269–293, October 1996. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00376323.

Nolan:2023:BRW

[NX23]  Kathleen Nolan and Constantinos Xenofontos.  Book

review: With whom am I speaking? A perspective on audiences. Tony Brown (2020). A contemporary theory of mathematics education research. Springer Nature. https://www.springer.com/gp/book/9783030550998 https://doi.org/10.1007/978-3-030-55100-1 xxi, 151 pages. Hardcover: ISBN 978-3-030-55099-8, 109.99 EUR. Softcover ISBN 978-3-030-55102-5, 109.99 EUR. eBook, ISBN 978-3-030-55100-1, 39,99 EUR. Educational Studies in Mathematics, 113(2):353–363, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10214-3. See response [Bro23].

Olsher:2025:RBS

[OAS25] Shai Olsher, Rotem Abdu, and Mohamed Shalata. The relationships between student content-specific grouping and teachers-students’ interactions during online collaborative mathematical task solving. Educational Studies in Mathematics, 119(2):249–268, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-024-10382-
w.

OBrien:1970:SNM

[O’B70] Thomas C. O’Brien. Some notes on multiplication of whole numbers. Educational Studies in Mathematics, 3(1):63–67, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00381594.

OBrien:1972:LTA

[O’B72] Thomas C. O’Brien. Logical thinking in adolescents. Educational Studies in Mathematics, 4(4):401–428, December 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00566998.

OBrien:1973:LTC

[O’B73] Thomas C. O’Brien. Logical thinking in college students. Educational Studies in Mathematics, 5(1):71–79, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01425413.

OBrien:1975:DFC

[O’B75] Thomas C. O’Brien. Deformation and the four card problem. Educational Studies in Mathematics, 6(1):23–39, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00590022.

OBrien:1976:TIE

[O’B76] Thomas C. O’Brien. Three informal essays. Educational Studies in Mathematics, 7(1–2):89–108, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144362.

OBrien:1977:IPR

[O’B77] Thomas C. O’Brien. Implications of Piagetian research for teachers of teachers and teachers of children. Educational Studies in Mathematics, 8(4):405–412, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310945.

Ouvrier-Buffet:2006:EMD

[OB06] Cecile Ouvrier-Buffet. Exploring mathematical definition construction processes. Educational Studies in Mathematics, 63 (3):259–282, November 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
9011-3.

Ouvrier-Buffet:2011:MEI

[OB11] Cécile Ouvrier-Buffet. A mathematical experience involving defining processes: in-action definitions and zero-definitions. Educational Studies in Mathematics, 76(2):165–182, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9272-3.

Obersteiner:2019:MPB

[Obe19] Andreas Obersteiner. Multiple pathways between affect and mathematical competence in young children— commentary on the studies in the special issue. Educational Studies in Mathematics, 100(3):317–323, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9853-
0.

OConnor:2001:CFT

[O’C01] M. C. O’Connor. “Can any fraction be turned into a decimal?” a case study of a mathematical group discussion. Educational Studies in Mathematics, 46(1–3):143–185, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014041308444.

Otte:2013:PPD

[OCA13] Michael Friedrich Otte, Tânia M. M. Campos, and Alexandre S. Abido. Plato, Pascal, and the dynamics of personal knowledge. Educational Studies in Mathematics, 82(3):397– 415, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9435-5.

Ozmantar:2025:MTE

[OCB25] Mehmet Fatih Ozmantar, Medine Coşkun, and Ali Bozkurt. Mathematics teachers’ ethical decision-making approaches in instructional practices. Educational Studies in Mathematics, 119(1):117–144, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-025-10396-y.

OptEynde:2006:AEC

[ODV06] Peter Op ’t Eynde, Erik De Corte, and Lieven Verschaffel. “Accepting Emotional Complexity”: A socio-constructivist perspective on the role of emotions in the mathematics classroom. Educational Studies in Mathematics, 63(2):193–207, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9034-4.

Ohuche:1978:CME

[Ohu78] R. Ogbonna Ohuche. Change in mathematics education since the late 1950’s — ideas and realisation Nigeria. Educational Studies in Mathematics, 9(3):271–281, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241030.

Okazaki:2005:CGL

[OK05] Masakazu Okazaki and Masataka Koyama. Characteristics of 5th graders’ logical development through learning division with decimals. Educational Studies in Mathematics, 60(2): 217–251, October 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-8123-0.

Olson:2013:WPM

[OK13] Jo Clay Olson and Libby Knott. When a problem is more than a teacher’s question. Educational Studies in Mathematics, 83 (1):27–36, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9444-4.

Olson:1987:LLL

[OKL87] A. T. Olson, T. E. Kieren, and S. Ludwig. Linking logo, levels and language in mathematics. Educational Studies in Mathematics, 18(4):359–370, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00240985.

Olande:2014:GAT

[Ola14] Oduor Olande. Graphical artefacts: Taxonomy of students’ response to test items. Educational Studies in Mathematics, 85 (1):53–74, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9493-3.

Olsen:2020:MLD

[OLW20] Joe Olsen, Kristen Lew, and Keith Weber. Metaphors for learning and doing mathematics in advanced mathematics lectures. Educational Studies in Mathematics, 105(1):1–17, September 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09968-x.

Oller-Marcen:2024:BRI

[OM24] Antonio M. Oller-Marcén. Book review: International and diverse. failed but seminal. Dirk de Bock (Ed.). (2023) Modern Mathematics. An International Movement? Educational Studies in Mathematics, 115(2):313–320, February 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10257-6.

Oslington:2020:TGP

[OMB20] Gabrielle Oslington, Joanne Mulligan, and Penny Van Bergen. Third-graders’ predictive reasoning strategies. Educational Studies in Mathematics, 104(1):5–24, May 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-020-09949-0.

Ongstad:2006:MME

[Ong06] Sigmund Ongstad. Mathematics and mathematics education as triadic communication? A semiotic framework exemplified. Educational Studies in Mathematics, 61(1–2):247–277, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-8302-7.

Owens:1998:RAP

[OPC^{+}98] Kay Owens, Bob Perry, John Conroy, Noel Geoghegan, and Peter Howe. Responsiveness and affective processes in the interactive construction of understanding in mathematics. Educational Studies in Mathematics, 35(2):105–127, February 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003042716302.

Orlov:1971:EVU

[Orl71] K. Orlov. Experimental verification of the use of the mathematical balance in secondary teaching. Educational Studies in Mathematics, 3(2):192–205, April 1971. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305448.

Orton:1983:SUD

[Ort83a] A. Orton. Students’ understanding of differentiation. Educational Studies in Mathematics, 14(3):235–250, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410540.

Orton:1983:SUI

[Ort83b] A. Orton. Students’ understanding of integration. Educational Studies in Mathematics, 14(1):1–18, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00704699.

Ortiz:2024:LPE

[Ort24] Nickolaus Alexander Ortiz. Lessons in paradise: envisioning a Black liberatory mathematics education. Educational Studies in Mathematics, 116(3):539–550, July 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10263-8.

Osibodu:2024:RTC

[Osi24] Oyemolade Osibodu. Racialization through coloniality in mathematics curricula: problematizing the Cambridge Assessment International Examination in Sub-Saharan Africa. Educational Studies in Mathematics, 116(3):445–461, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10313-9.

Oslund:2012:MTW

[Osl12] Joy A. Oslund. Mathematics-for-teaching: what can be learned from the ethnopoetics of teachers’ stories? Educational Studies in Mathematics, 79(2):293–309, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9348-8.

OBrien:1971:LTL

[OSR71] Thomas C. O’Brien, Bernard J. Shapiro, and Norma C. Reali. Logical thinking-language and context. Educational Studies in Mathematics, 4(2):201–219, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303385.

Osterholm:2006:CRC

[Öst06] Magnus Österholm. Characterizing reading comprehension of mathematical texts. Educational Studies in Mathematics, 63(3):325–346, November 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
9016-y.

Otte:1976:BRD

[Ott76] M. Otte. Book review: Die didaktischen systeme von V. V. Davidov/D. B. Elkonin einerseits und L. V. Zankov andererseits. Educational Studies in Mathematics, 6(4):475–497, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00411093.

Otte:1994:MKP

[Ott94] Michael Otte. Mathematical knowledge and the problem of proof. Educational Studies in Mathematics, 26(4):299–321, June 1994. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01279518.

Otte:2003:CSN

[Ott03] M. Otte. Complementarity, sets and numbers. Educational Studies in Mathematics, 53(3):203–228, September 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026001332585.

Otte:2006:MEP

[Ott06] Michael Otte. Mathematical epistemology from a Peircean semiotic point of view. Educational Studies in Mathematics, 61(1–2):11–38, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
0082-6.

Otte:2007:MHP

[Ott07] Michael Otte. Mathematical history, philosophy and education. Educational Studies in Mathematics, 66(2):243–255, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-007-9079-z.

Otte:2011:ELS

[Ott11] Michael Friedrich Otte. Evolution, learning, and semiotics from a Peircean point of view. Educational Studies in Mathematics, 77(2–3):313–329, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9302-9.

Ott:2020:LGG

[Ott20] Barbara Ott. Learner-generated graphic representations for word problems: an intervention and evaluation study in grade 3. Educational Studies in Mathematics, 105(1):91–113, September 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09978-9;   http://link.
springer.com/content/pdf/10.1007/s10649-020-09978-
9.pdf.

Okumus:2025:PSM

[OVKR25] Samet Okumus, Naďa Vondrová, Tuğrul Kar, and Jarmila Robová. Pre-service mathematics teachers’ handling of a pupil’s incorrect reasoning with a correct answer. Educational Studies in Mathematics, 118(2):249–269, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10369-7.

Otte:1985:C

[OZ85] M. Otte and W. Zawadowski. Creativity. Educational Studies in Mathematics, 16(1):95–97, February 1985. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00354885.

Parker:2014:STS

[PA14] Diane Parker and Jill Adler. Sociological tools in the study of knowledge and practice in mathematics teacher education. Educational Studies in Mathematics, 87(2):203–219, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9421-y.

Palatnik:2018:RMT

[PA18] Alik Palatnik and Dor Abrahamson. Rhythmic movement as a tacit enactment goal mobilizes the emergence of mathematical structures. Educational Studies in Mathematics, 99(3):293– 309, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9845-0.

Pais:2011:CCE

[Pai11] Alexandre Pais. Criticisms and contradictions of ethnomathematics. Educational Studies in Mathematics, 76(2):209–230, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9289-7.

Pais:2013:ICU

[Pai13] Alexandre Pais. An ideology critique of the use-value of mathematics. Educational Studies in Mathematics, 84(1):15–34, September 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9484-4.

Pais:2015:SRM

[Pai15] Alexandre Pais. Symbolising the real of mathematics education. Educational Studies in Mathematics, 89(3):375–391, July 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9602-6.

Pais:2016:IBS

[Pai16] Alexandre Pais. At the intersection between the subject and the political: a contribution to an ongoing discussion. Educational Studies in Mathematics, 92(3):347–359, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9687-6.

Pais:2018:BRL

[Pai18] Alexandre Pais. Book review: The limits of pluralism. Bharath Sriraman (Ed.) (2017) Humanizing mathematics and its philosophy: Essays celebrating the 90th birthday of Reuben Hersh. Educational Studies in Mathematics, 99(2):235–240, October 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9829-0.

Pais:2019:MCB

[Pai19]  Alexandre Pais.  Mathematics, capitalism and bioso-

cial research. Educational Studies in Mathematics, 101 (3):373–386, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9859-
7;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9859-7.pdf.

Palm:2008:IAS

[Pal08] Torulf Palm. Impact of authenticity on sense making in word problem solving. Educational Studies in Mathematics, 67(1): 37–58, January 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9083-3.

Palatnik:2020:PPO

[Pal20] Alik Palatnik. A pedagogical potential of one mathematical inaccuracy. Educational Studies in Mathematics, 103(1): 121–124, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-019-09922-6.

Papy:1969:M

[Pap69] Frédérique Papy. Minicomputer. Educational Studies in Mathematics, 2(2–3):333–345, December 1969. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303467.

Papy:1971:FIN

[Pap71] G. Papy. A first introduction to the notion of topological space. Educational Studies in Mathematics, 4(1):18–30, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305794.

Parzysz:1988:KVS

[Par88] Bernard Parzysz. “Knowing” vs “seeing”. Problems of the plane representation of space geometry figures. Educational Studies in Mathematics, 19(1):79–92, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00428386.

Parzysz:1991:RSS

[Par91] Bernard Parzysz. Representation of space and students’ conceptions at high school level. Educational Studies in Mathematics, 22(6):575–593, December 1991. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312716.

Park:2015:EDF

[Par15a] Jungeun Park. Erratum to: Is the derivative a function? If so, how do we teach it? Educational Studies in Mathematics, 90 (2):231, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-015-9630-2.pdf. See

[Par15b].

Park:2015:DFI

[Par15b] Jungeun Park. Is the derivative a function? If so, how do we teach it? Educational Studies in Mathematics, 89(2):233– 250, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9601-7. See er-

ratum [Par15a].

Park:2016:CAS

[Par16] Jungeun Park. Communicational approach to study textbook discourse on the derivative. Educational Studies in Mathematics, 91(3):395–421, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9655-6.

Pomeroy:2024:SES

[PAR^{+}24] David Pomeroy, Mahdis Azarmandi, Matiu Tai Ratima, Sara Tolbert, Kay-Lee Jones, Nathan Riki, and Te Hurinui Karaka-Clarke. Shame, entitlement, and the systemic racism of mathematics “ability” grouping in Aotearoa New Zealand. Educational Studies in Mathematics, 116(3):463–478, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10266-5.

Park:2025:EPC

[Par25] JinHyeong Park. Exploring the probabilities of compound events and the equiprobability bias by navigating between enumeration and probability operations. Educational Studies in Mathematics, 120(2):371–395, October 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10430-z.

Pastre:1978:ODM

[Pas78] Dominique Pastre. Observation du mathématicien: Aide à l’enseignement et à la démonstration automatique de théorèmes. (French) [Observation of the mathematician: Aid for teaching and the automatic demonstration of theorems]. Educational Studies in Mathematics, 9(4):461–502, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410590.

Patronis:1985:CSR

[Pat85] T. Patronis. Convexity and shortest road. Educational Studies in Mathematics, 16(4):425–430, November 1985.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417196.

Pfannkuch:2015:WSQ

[PAW15] Maxine Pfannkuch, Pip Arnold, and Chris J. Wild. What i see is not quite the way it really is: students’ emergent reasoning about sampling variability. Educational Studies in Mathematics, 88(3):343–360, March 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9539-1.

Pehkonen:1984:BR

[PB84] Erkki Pehkonen and C. A. Brown. Book reviews. Educational Studies in Mathematics, 15(3):321–323, August 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312081.

Picker:2000:IPI

[PB00] Susan H. Picker and John S. Berry. Investigating pupils’ images of mathematicians. Educational Studies in Mathematics, 43(1):65–94, July 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017523230758.

Pierce:2009:PMA

[PB09] Robyn Pierce and Lynda Ball. Perceptions that may affect teachers’ intention to use technology in secondary mathematics classes. Educational Studies in Mathematics, 71(3):299– 317, July 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9177-6.

Pedersen:2021:TCF

[PB21] Pernille Ladegaard Pedersen and Mette Bjerre. Two conceptions of fraction equivalence. Educational Studies in Mathematics, 107(1):135–157, May 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10030-

Polo-Blanco:2025:RGE

[PBCPP25] Irene Polo-Blanco, Maria Chimoni, and Demetra Pitta-Pantazi. Representations and generalization in early algebra: a comparative study of autistic students and their non-autistic peers. Educational Studies in Mathematics, 120(1):33–55, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10416-x.

Peressini:2004:CFL

[PBR^{+}04] Dominic Peressini, Hilda Borko, Lew Romagnano, Eric Knuth, and Christine Willis. A conceptual framework for learning to teach secondary mathematics: A situative perspective. Educational Studies in Mathematics, 56(1):67–96, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000028398.80108.87.

Pape:2003:DMT

[PBY03] S. J. Pape, C. V. Bell, and İE. Yetkin. Developing mathematical thinking and self-regulated learning: A teaching experiment in a seventh-grade mathematics classroom. Educational Studies in Mathematics, 53(3):179–202, September 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026062121857.

Philippou:1998:EPM

[PC98] George N. Philippou and Constantinos Christou. The effects of a preparatory mathematics program in changing prospective teachers’ attitudes towards mathematics. Educational Studies in Mathematics, 35(2):189–206, February 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003030211453.

Pittalis:2010:TRG

[PC10] Marios Pittalis and Constantinos Christou. Types of reasoning in 3D geometry thinking and their relation with spatial ability. Educational Studies in Mathematics, 75(2):191–212, November 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9251-8.

Pinto:2023:CRF

[PC23] Alon Pinto and Jason Cooper. “This cannot be” — refutation feedback and its potential affordances for proof comprehension. Educational Studies in Mathematics, 113(2):287– 306, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10190-0.

Passolunghi:2019:RBC

[PCP19] Maria Chiara Passolunghi, Elisa Cargnelutti, and Sandra Pellizzoni. The relation between cognitive and emotional factors and arithmetic problem-solving. Educational Studies in Mathematics, 100(3):271–290, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9863-
y.

Pea:1991:BR

[PD91] Roy D. Pea and A. J. Sandy Dawson. Book review. Educational Studies in Mathematics, 22(5):481–501, October 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367910.

Pittalis:2023:EID

[PD23] Marios Pittalis and Paul Drijvers. Embodied instrumentation in a dynamic geometry environment: eleven-year-old students’ dragging schemes. Educational Studies in Mathematics, 113(2):181–205, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10222-3.

Peters:2012:CUS

[PDT^{+}12] Greet Peters, Bert De Smedt, Joke Torbeyns, Pol Ghesquière, and Lieven Verschaffel. Children’s use of subtraction by addition on large single-digit subtractions. Educational Studies in Mathematics, 79(3):335–349, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9308-3.

Pijls:2007:RCM

[PDV07] Monique Pijls, Rijkje Dekker, and Bernadette Van Hout-Wolters. Reconstruction of a collaborative mathematical learning process. Educational Studies in Mathematics, 65(3):309– 329, July 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9051-3.

Peaker:1976:CID

[Pea76] G. F. Peaker. A commentary from IEA on Dr. Freudenthal’s article in: Educational Studies in Mathematics, vol. 6, no. 2. Educational Studies in Mathematics, 7(4):523–527, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00452228. See [Fre75].

Pedemonte:2007:HCR

[Ped07] Bettina Pedemonte. How can the relationship between argumentation and proof be analysed? Educational Studies in Mathematics, 66(1):23–41, September 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9057-x.

Pettersson:1991:PMP

[Pet91] Astrid Pettersson. Pupils’ mathematical performance in grades 3 and 6. A longitudinal study. Educational Studies in Mathematics, 22(5):439–450, October 1991. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00367907.

Pinxten:2011:PIC

[PF11] Rik Pinxten and Karen François. Politics in an Indian canyon? Some thoughts on the implications of ethnomathematics. Educational Studies in Mathematics, 78(2):261–273, November 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9328-z.

Prediger:2019:CMC

[PFSS19] Susanne Prediger, Claudia Fischer, Christoph Selter, and Christian Schöber. Combining material- and communitybased implementation strategies for scaling up: the case of supporting low-achieving middle school students. Educational Studies in Mathematics, 102(3):361–378, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9835-2.

Perrenet:1993:HAH

[PG93] Jacob Perrenet and Wim Groen. A hint is not always a help. Educational Studies in Mathematics, 25(4):307–329, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273904.

Papadopoulos:2020:EWR

[PG20] Ioannis Papadopoulos and Robert Gunnarsson. Exploring the way rational expressions trigger the use of “mental” brackets by primary school students. Educational Studies in Mathematics, 103(2):191–207, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-019-09929-z.

Pantziara:2009:UDT

[PGE09] Marilena Pantziara, Athanasios Gagatsis, and Iliada Elia. Using diagrams as tools for the solution of non-routine mathematical problems. Educational Studies in Mathematics, 72(1): 39–60, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9181-5.

Park:2013:TPT

[PGM13] Jungeun Park, Beste Güçler, and Raven McCrory. Teaching prospective teachers about fractions: historical and pedagogical perspectives. Educational Studies in Mathematics, 82(3): 455–479, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9440-8.

Paoletti:2022:CSC

[PGV22] Teo Paoletti, Allison L. Gantt, and Madhavi Vishnubhotla. Constructing a system of covariational relationships: two contrasting cases. Educational Studies in Mathematics, 110 (3):413–433, July 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10134-0.

Pitkethly:1996:RRR

[PH96] Anne Pitkethly and Robert Hunting. A review of recent research in the area of initial fraction concepts. Educational Studies in Mathematics, 30(1):5–38, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00163751.

Phelps:2010:FPS

[Phe10] Christine M. Phelps. Factors that pre-service elementary teachers perceive as affecting their motivational profiles in mathematics. Educational Studies in Mathematics, 75(3):293– 309, December 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-010-9257-2.

Prusak:2012:VRL

[PHS12] Naomi Prusak, Rina Hershkowitz, and Baruch B. Schwarz. From visual reasoning to logical necessity through argumentative design. Educational Studies in Mathematics, 79(1):19–40, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9335-0.

Phythian:1971:MKT

[Phy71] J. E. Phythian. Mathematical kujitegemea in Tanzania. Educational Studies in Mathematics, 4(2):187–200, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303384.

Pickert:1971:IMU

[Pic71] Günter Pickert. The introduction of metric by the use of conics. Educational Studies in Mathematics, 4(1):31–47, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305795.

Pimm:1990:PRT

[Pim90] David Pimm. Problems of representation in the teaching and learning of mathematics. Educational Studies in Mathematics, 21(1):91–99, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311018.

Pimm:1991:BR

[Pim91] David Pimm. Book review. Educational Studies in Mathematics, 22(4):391–405, August 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00369297.

Pimm:1993:BR

[Pim93] David Pimm. Book review. Educational Studies in Mathematics, 25(3):261–277, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273865.

Pimm:1995:AP

[Pim95] David Pimm. The advance party. Educational Studies in Mathematics, 29(2):97–122, September 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274209.

Piroi:2025:EAS

[Pir25] Margherita Piroi. The emergence of the analytic-structural way of thinking in linear algebra as a blended space. Educational Studies in Mathematics, 118(1):101–117, January 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10354-0.

Pisot:1968:EMP

[Pis68] Charles Pisot. Enseignement des mathématiques pour physiciens. (French) [Teaching mathematics for physicists]. Educational Studies in Mathematics, 1(1–2):89–92, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426234.

Potari:2023:TUM

[PJP23] Despina Potari, Barbara Jaworski, and Georgia Petropoulou. Theorizing university mathematics teaching: the Teaching Triad within an Activity Theory perspective. Educational Studies in Mathematics, 114(1):1–16, September 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10244-x.

Pirie:1992:CCE

[PK92] Susan Pirie and Thomas Kieren. Creating constructivist environments and constructing creative mathematics. Educational Studies in Mathematics, 23(5):505–528, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00571470.

Pirie:1994:GMU

[PK94] Susan Pirie and Thomas Kieren. Growth in mathematical understanding: How can we characterise it and how can we represent it? Educational Studies in Mathematics, 26(2–3): 165–190, March 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273662.

Palatnik:2017:SMC

[PK17] Alik Palatnik and Boris Koichu. Sense making in the context of algebraic activities. Educational Studies in Mathematics, 95 (3):245–262, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Park:2023:FFO

[PK23a] JinHyeong Park and Dong-Won Kim. Facilitating factors and obstacles in the coordination of theoretical probability and relative frequency estimates of probability. Educational Studies in Mathematics, 114(3):439–460, November 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10253-w.

Park:2023:IQT

[PK23b] Mimi Park and Yeon Kim. Instructional quality of two beginning mathematics teachers for three years: what professional competency makes a difference? Educational Studies in Mathematics, 112(3):505–529, March 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10195-9.

Pinto:2023:DPE

[PK23c] Alon Pinto and Boris Koichu. Diverse perspectives and experiences of university mathematics teachers on improving the secondary–tertiary transition. Educational Studies in Mathematics, 113(1):147–164, May 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-022-10196-
8.

Paoletti:2018:TQI

[PKP^{+}18] Teo Paoletti, Victoria Krupnik, Dimitrios Papadopoulos, Joseph Olsen, Tim Fukawa-Connelly, and Keith Weber. Teacher questioning and invitations to participate in advanced mathematics lectures. Educational Studies in Mathematics, 98(1):1–17, May 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9807-6.

Planas:2014:OST

[Pla14] Núria Planas. One speaker, two languages: Learning opportunities in the mathematics classroom. Educational Studies in Mathematics, 87(1):51–66, September 2014. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9553-3.

Planas:2018:LRK

[Pla18] Núria Planas. Language as resource: a key notion for understanding the complexity of mathematics learning. Educational Studies in Mathematics, 98(3):215–229, July 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9810-y;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9810-y.pdf.

Pluvinage:1979:LQP

[Plu79] F. Pluvinage. Loto-questionnaires (pour l’évaluation et l’autocontrôle en mathématiques). (French) [Loto-questionnaires (for evaluation and self-checking in mathematics)]. Educational Studies in Mathematics, 10(4):443–485, November 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417090.

Pollatsek:1981:CCS

[PLW81] A. Pollatsek, S. Lima, and A. D. Well. Concept or computation: Students’ understanding of the mean. Educational Studies in Mathematics, 12(2):191–204, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305621.

Pirie:1997:EWE

[PM97] Susan E. B. Pirie and Lyndon Martin. The equation, the whole equation and nothing but the equation! One approach to the teaching of linear equations. Educational Studies in Mathematics, 34(2):159–181, November 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003051829991.

Porter:2000:EEW

[PM00] Mary K. Porter and Joanna O. Masingila. Examining the effects of writing on conceptual and procedural knowledge in calculus. Educational Studies in Mathematics, 42(2):165–177, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1004166811047.

Passelaigue:2015:STD

[PM15] Dominique Passelaigue and Valérie Munier. Schoolteacher trainees’ difficulties about the concepts of attribute and measurement. Educational Studies in Mathematics, 89(3):307– 336, July 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9610-6.

Peck:2016:RFD

[PM16] Frederick Peck and Michael Matassa. Reinventing fractions and division as they are used in algebra: the power of preformal productions. Educational Studies in Mathematics, 92 (2):245–278, June 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9690-y.

Proulx:2019:ESE

[PM19] Jérôme Proulx and Jean-François Maheux. Effect sizes, epistemological issues, and identity of mathematics education research: a commentary on editorial 102(1). Educational Studies in Mathematics, 102(2):299–302, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09913-7. See [BCE^{+}19].

Ponte:1994:TSV

[PMG^{+}94] João P. Ponte, João F. Matos, Henrique M. Guimarães, Leonor C. Leal, and Ana P. Canavarro. Teachers’ and students’ views and attitudes towards a new mathematics curriculum: A case study. Educational Studies in Mathematics, 26 (4):347–365, June 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01279520.

Pozzi:1998:TPM

[PNH98] Stefano Pozzi, Richard Noss, and Celia Hoyles. Tools in practice, mathematics in use. Educational Studies in Mathematics, 36(2):105–122, July 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003216218471.

Pollak:1968:SPT

[Pol68] H. O. Pollak. On some of the problems of teaching applications of mathematics. Educational Studies in Mathematics, 1(1– 2):24–30, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00426228.

Pollak:1969:HCW

[Pol69] H. O. Pollak. How can we teach applications of mathematics? Educational Studies in Mathematics, 2(2–3):393–404, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00303471.

Pollak:1976:WIW

[Pol76] H. O. Pollak. What industry wants a mathematician to know and how we want them to know it. Educational Studies in Mathematics, 7(1–2):109–112, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144363.

Popkewitz:1988:IIS

[Pop88] Thomas S. Popkewitz. Institutional issues in the study of school mathematics: Curriculum research. Educational Studies in Mathematics, 19(2):221–249, May 1988. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00751234.

Porteous:1990:WDC

[Por90] Keith Porteous. What do children really believe? Educational Studies in Mathematics, 21(6):589–598, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00315946.

Pantziara:2012:LSC

[PP12] Marilena Pantziara and George Philippou. Levels of students’ “conception” of fractions. Educational Studies in Mathematics, 79(1):61–83, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9338-x.

Pitta-Pantazi:2009:CSD

[PPC09] Demetra Pitta-Pantazi and Constantinos Christou. Cognitive styles, dynamic geometry and measurement performance. Educational Studies in Mathematics, 70(1):5–26, January 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9139-z.

Pitta-Pantazi:2025:SSS

[PPDS^{+}25] Demetra Pitta-Pantazi, Eleni Demosthenous, Maike Schindler, Achim J. Lilienthal, and Constantinos Christou. Structure sense in students’ quantity comparison and repeating pattern extension tasks: an eye-tracking study with first graders. Educational Studies in Mathematics, 118(3):339–357, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10290-5.

Phillips:1995:ESU

[PPG95] Richard J. Phillips, Daniel Pead, and John Gillespie. Evolving strategies for using interactive video resources in mathematics classrooms. Educational Studies in Mathematics, 28 (2):133–154, March 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01295790.

Paparistodemou:2014:PTA

[PPPP14] Efi Paparistodemou, Despina Potari, and Demetra Pitta-Pantazi. Prospective teachers’ attention on geometrical tasks. Educational Studies in Mathematics, 86(1):1–18, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9518-y.

Potari:2019:CDR

[PPSZ19] Despina Potari, Giorgos Psycharis, Charalampos Sakonidis, and Theodossios Zachariades. Collaborative design of a reform-oriented mathematics curriculum: contradictions and boundaries across teaching, research, and policy. Educational Studies in Mathematics, 102(3):417–434, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9834-3.

Presmeg:2008:SSR

[PR08] Norma Presmeg and Luis Radford. On semiotics and subjectivity: a response to Tony Brown’s “Signifying ‘students’, ‘teachers’, and ‘mathematics’: a reading of a special issue”. Educational Studies in Mathematics, 69(3):265–276, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9146-0.

Pedemonte:2011:RAP

[PR11] Bettina Pedemonte and David Reid. The role of abduction in proving processes. Educational Studies in Mathematics, 76 (3):281–303, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-010-9275-0.

Presmeg:1986:VMG

[Pre86] Norma C. Presmeg. Visualisation and mathematical giftedness. Educational Studies in Mathematics, 17(3):297–311, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305075.

Presmeg:1988:SMC

[Pre88] Norma C. Presmeg. School mathematics in culture-conflict situations. Educational Studies in Mathematics, 19(2):163– 177, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00751230.

Presmeg:1992:BMB

[Pre92a] Norma C. Presmeg. The body in the mind: The bodily basis of meaning, imagination and reason. Educational Studies in Mathematics, 23(3):307–314, June 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02309535.

Presmeg:1992:PMM

[Pre92b] Norma C. Presmeg. Prototypes, metaphors, metonymies and imaginative rationality in high school mathematics. Educational Studies in Mathematics, 23(6):595–610, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00540062.

Presmeg:2003:CMD

[Pre03] Norma Presmeg. Creativity, mathematizing, and didactizing: Leen Streefland’s work continues. Educational Studies in Mathematics, 54(1):127–137, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000005255.04769.89.

Presmeg:2006:SCS

[Pre06] Norma Presmeg. Semiotics and the “Connections” Standard: Significance of semiotics for teachers of mathematics. Educational Studies in Mathematics, 61(1–2):163–182, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-3365-z.

Presmeg:2007:BRP

[Pre07] Norma Presmeg. Book review: Pattern and beauty in mathematics and music. Educational Studies in Mathematics, 65 (1):121–124, May 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9068-7.

Presmeg:2009:E

[Pre09] Norma Presmeg. Editorial. Educational Studies in Mathematics, 70(1):1–3, January 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-008-9171-z.
pdf.

Presmeg:2010:E

[Pre10] Norma Presmeg. Editorial. Educational Studies in Mathematics, 73(1):1–2, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-009-9224-y.
pdf.

Presmeg:2011:E

[Pre11] Norma Presmeg. Editorial. Educational Studies in Mathematics, 76(1):1–2, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-010-9287-9.
pdf.

Presmeg:2012:E

[Pre12] Norma Presmeg. Editorial. Educational Studies in Mathematics, 79(1):1–2, January 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-011-9369-3.
pdf.

Presmeg:2013:Ea

[Pre13a] Norma Presmeg. Editorial. Educational Studies in Mathematics, 82(1):1–3, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-012-9451-5.
pdf.

Presmeg:2013:Eb

[Pre13b] Norma Presmeg. Editorial. Educational Studies in Mathematics, 84(3):279–280, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-013-9511-5.
pdf.

Presmeg:2016:CLR

[Pre16] Norma Presmeg. Commognition as a lens for research. Educational Studies in Mathematics, 91(3):423–430, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9676-1.

Presmeg:2023:RLE

[Pre23] Norma Presmeg. Reply to the letter to the Editor. Educational Studies in Mathematics, 114(1):S10, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-011-9312-7.

Proia:1969:ETG

[Pro69] Lina Mancini Proia. An experiment in teaching geometry. Educational Studies in Mathematics, 1(4):378–385, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02396310.

Proia:1970:ELC

[Pro70] Lina Mancini Proia. Une expérience dans un lycée classique. (French) [An experience in a classical high school]. Educational Studies in Mathematics, 3(1):43–62, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381593.

Proulx:2013:MME

[Pro13] Jérôme Proulx. Mental mathematics, emergence of strategies, and the enactivist theory of cognition. Educational Studies in Mathematics, 84(3):309–328, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9480-8.

Paries:2008:ASC

[PRR08] M. Pariès, A. Robert, and J. Rogalski. Analyses de séances en classe et stabilité des pratiques d’enseignants de mathématiques expérimentés du second degré. (French) [Analyzes of classroom sessions and stability of the practices of experimental-mathematics teachers of the second level]. Educational Studies in Mathematics, 68(1):55–80, May 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9108-y.

Przenioslo:2004:ILF

[Prz04] Małgorzata Przenioslo. Images of the limit of function formed in the course of mathematical studies at the university. Educational Studies in Mathematics, 55(1–3):103–132, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017667.70982.05.

Przenioslo:2005:ICC

[Prz05] Małgorzata Przenioslo. Introducing the concept of convergence of a sequence in secondary school. Educational Studies in Mathematics, 60(1):71–93, September 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5325-4.

Pirie:1988:MDM

[PS88] S. E. B. Pirie and R. L. E. Schwarzenberger. Mathematical discussion and mathematical understanding. Educational Studies in Mathematics, 19(4):459–470, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578694.

Polotskaia:2021:SMS

[PS21] Elena Polotskaia and Annie Savard. Some multiplicative structures in elementary education: a view from relational paradigm. Educational Studies in Mathematics, 106(3):447– 469, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09979-8.

Pollitt:1987:FAG

[PSE87] A. B. Pollitt, J. W. Searl, and A. S. El Sheikh. Factors affecting girls’ learning of mathematics in Sudan. Educational Studies in Mathematics, 18(4):439–444, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00240989.

Paoletti:2018:IFP

[PSH^{+}18] Teo Paoletti, Irma E. Stevens, Natalie L. F. Hobson, Kevin C. Moore, and Kevin R. LaForest. Inverse function: Pre-service teachers’ techniques and meanings. Educational Studies in Mathematics, 97(1):93–109, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9787-y.

Petronzi:2019:FDC

[PSS^{+}19] Dominic Petronzi, Paul Staples, David Sheffield, Thomas E. Hunt, and Sandra Fitton-Wilde. Further development of the Children’s Mathematics Anxiety Scale UK (CMAS-UK) for ages 4–7 years. Educational Studies in Mathematics, 100(3):231–249, March 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9860-
1;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9860-1.pdf.

Perrenet:2009:MES

[PT09] Jacob Perrenet and Ruurd Taconis. Mathematical enculturation from the students’ perspective: shifts in problem-solving beliefs and behaviour during the bachelor programme. Educational Studies in Mathematics, 71(2):181–198, June 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-008-9166-9.pdf.

Pugalee:2004:CVW

[Pug04] David K. Pugalee. A comparison of verbal and written descriptions of students’ problem solving processes. Educational Studies in Mathematics, 55(1–3):27–47, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017666.11367.c7.

Purakam:1978:CME

[Pur78] Orasri Purakam. Change in mathematics education since the late 1950’s — ideas and realisation Thailand. Educational Studies in Mathematics, 9(3):331–337, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241035.

Pais:2012:RRM

[PV12] Alexandre Pais and Paola Valero. Researching research: mathematics education in the political. Educational Studies in Mathematics, 80(1–2):9–24, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9399-5.

Pais:2014:WST

[PV14] Alexandre Pais and Paola Valero. Whither social theory? Educational Studies in Mathematics, 87(2):241–248, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9573-z.

Presmeg:2003:LSL

[PvdHP03] Norma Presmeg and Marja van den Heuvel-Panhuizen. Leen Streefland’s legacy. Educational Studies in Mathematics, 54 (1):1–4, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000005235.
20840.14.

Peltenburg:2012:SES

[PvdHPR12] Marjolijn Peltenburg, Marja van den Heuvel-Panhuizen, and Alexander Robitzsch. Special education students’ use of indirect addition in solving subtraction problems up to 100— a proof of the didactical potential of an ignored procedure. Educational Studies in Mathematics, 79(3):351–369, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-011-9351-0.pdf.

Pimm:2003:IME

[PW03] David Pimm and David Wagner. Investigation, mathematics education and genre: An essay review of Candia Morgan’s Writing Mathematically: the Discourse of Investigation. Educational Studies in Mathematics, 53(2):159–178, July 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1025531028661.

Prediger:2022:ESB

[PW22] Susanne Prediger and David Wagner. Editorial: Singleblind or double-blind review processes? Educational Studies in Mathematics, 111(1):1–3, September 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10171-3.

Prediger:2024:E

[PW24] Susanne Prediger and David Wagner. Editorial. Educational Studies in Mathematics, 115(1):1–3, January 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10304-w.

Pepin:2017:DDU

[PXTW17] Birgit Pepin, Binyan Xu, Luc Trouche, and Chongyang Wang. Developing a deeper understanding of mathematics teaching expertise: an examination of three Chinese mathematics teachers’ resource systems as windows into their work and expertise. Educational Studies in Mathematics, 94(3):257–274, March 2017. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-016-9727-2.pdf.

Popper:2022:OEB

[PY22] Porat Popper and Michal Yerushalmy. Online examplebased assessment as a resource for teaching about quadrilaterals. Educational Studies in Mathematics, 110(1):83–100, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10109-1.

Quesada:1994:EUG

[QM94] Antonio R. Quesada and Mary E. Maxwell. The effects of using graphing calculators to enhance college students’ performance in precalculus. Educational Studies in Mathematics, 27(2):205–215, September 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01278922.

Quaye:2022:SCI

[QP22] Jeffery Quaye and David Pomeroy. Social class inequalities in attitudes towards mathematics and achievement in mathematics cross generations: a quantitative Bourdieusian analysis. Educational Studies in Mathematics, 109(1):155–175, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10078-5.

Quan:1992:CAM

[Qua92] Wang Lin Quan. Chinese advancements in mathematics education. Educational Studies in Mathematics, 23(3):287–298, June 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02309533.

r:2020:CES

[rA20] Aslı Çakır and Hatice Akkoç. Correction to: Examining sociomathematical norms related to problem posing: a case of a gifted and talented mathematics classroom. Educational Studies in Mathematics, 105(3):433, November 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10001-4.

Rotem:2021:EIH

[RA21] Sigal-Hava Rotem and Michal Ayalon. Exploring Israeli high school graduates’ explanations for the spread of the coronavirus. Educational Studies in Mathematics, 108(1–2): 161–181, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10042-3.

Radford:2000:SMS

[Rad00] Luis Radford. Signs and meanings in students’ emergent algebraic thinking: A semiotic analysis. Educational Studies in Mathematics, 42(3):237–268, May 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017530828058.

Radford:2003:ELL

[Rad03] Luis Radford. On the epistemological limits of language: Mathematical knowledge and social practice during the Renaissance. Educational Studies in Mathematics, 52(2):123–150, March 2003. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1024029808871.

Radford:2006:AM

[Rad06] Luis Radford. The anthropology of meaning. Educational Studies in Mathematics, 61(1–2):39–65, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-7136-7.

Radford:2009:WDG

[Rad09] Luis Radford. Why do gestures matter? Sensuous cognition and the palpability of mathematical meanings. Educational Studies in Mathematics, 70(2):111–126, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9127-3.

Radford:2011:BRC

[Rad11] Luis Radford. Book review: Classroom Interaction: Why is it Good, Really? Baruch Schwarz, Tommy Dreyfus and Rina Hershkowitz (Eds.) (2009) Transformation of knowledge through classroom interaction. Educational Studies in Mathematics, 76(1):101–115, January 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9271-4.

Radford:2012:EIE

[Rad12] Luis Radford. Education and the illusions of emancipation. Educational Studies in Mathematics, 80(1–2):101–118, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9380-8.

Radford:2014:RRA

[Rad14] Luis Radford. On the role of representations and artefacts in knowing and learning. Educational Studies in Mathematics, 85 (3):405–422, March 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9527-x.

Radford:2021:DCC

[Rad21] Luis Radford. Davydov’s concept of the concept and its dialectical materialist background. Educational Studies in Mathematics, 106(3):327–342, March 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09959-y.

Raman:2003:KIW

[Ram03] M. Raman. Key ideas: What are they and how can they help us understand how people view proof? Educational Studies in Mathematics, 52(3):319–325, April 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024360204239.

Rossner:2025:SPT

[RBK25] Michael Rößner, Karin Binder, and Stefan Krauss. Students’ performance and typical errors in filling empty probabilistic visualizations with probabilities or frequencies. Educational Studies in Mathematics, 120(1):137–167, September 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10372-y.

Radovic:2018:TCC

[RBWS18] Darinka Radovic, Laura Black, Julian Williams, and Christian E. Salas. Towards conceptual coherence in the research on mathematics learner identity: a systematic review of the literature. Educational Studies in Mathematics, 99(1):21–42, September 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-018-9819-2.

Ruthven:1994:SAS

[RC94] Kenneth Ruthven and Robert Coe. A structural analysis of students’ epistemic views. Educational Studies in Mathematics, 27(1):101–109, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01284530.

Rowlands:2002:WWF

[RC02] Stuart Rowlands and Robert Carson. Where would formal, academic mathematics stand in a curriculum informed by ethnomathematics? A critical review of ethnomathematics. Educational Studies in Mathematics, 50(1):79–102, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020532926983. See [AAB03] and re-

sponse [RC04].

Rowlands:2004:ORA

[RC04] Stuart Rowlands and Robert Carson. Our response to Adam, Alangui and Barton’s “A Comment on Rowlands & Carson ‘Where would Formal, Academic Mathematics stand in a Curriculum informed by Ethnomathematics? A Critical Review’ ”. Educational Studies in Mathematics, 56(2–3): 329–342, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040370.

10717.82. See [AAB03, RC02].

Ryan:2020:JPR

[RC20] Ulrika Ryan and Anna Chronaki. A joke on precision? Revisiting “precision” in the school mathematics discourse.

Educational Studies in Mathematics, 104(3):369–384, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-020-09963-2;   http://link.springer.
com/content/pdf/10.1007/s10649-020-09963-2.pdf.

Ran:2025:AUR

[RCH25] Hua Ran, Jinfa Cai, and Stephen Hwang. An analysis of unexpected responses in middle school students’ mathematical problem posing from the perspective of problem-posing processes. Educational Studies in Mathematics, 119(3), July 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10399-9.

Robitaille:1992:TIM

[RD92] David F. Robitaille and J. Stuart Donn. The Third International Mathematics and Science Study (TIMSS): A brief introduction. Educational Studies in Mathematics, 23(2):203– 210, April 1992. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00588056.

Ruthven:2009:UGS

[RDH09] Kenneth Ruthven, Rosemary Deaney, and Sara Hennessy. Using graphing software to teach about algebraic forms: a study of technology-supported practice in secondary-school mathematics. Educational Studies in Mathematics, 71(3):279–297, July 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9176-7.

Ron:2010:PCC

[RDH10] Gila Ron, Tommy Dreyfus, and Rina Hershkowitz. Partially correct constructs illuminate students’ inconsistent answers. Educational Studies in Mathematics, 75(1):65–87, September 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9241-x.

Radford:2009:IBW

[REA09] L. Radford, L. Edwards, and F. Arzarello. Introduction: beyond words. Educational Studies in Mathematics, 70(2):91–95, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-008-9172-y.pdf.

Relich:1996:GSC

[Rel96] Joe Relich. Gender, self-concept and teachers of mathematics: Effects on attitudes to teaching and learning. Educational Studies in Mathematics, 30(2):179–195, March 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302629.

Ron-Ezra:2025:EHS

[REL25] Maya Ron-Ezra and Esther S. Levenson. Exploring how students with learning disorders engage with open arithmetic tasks using a mathematical creativity lens. Educational Studies in Mathematics, 120(3):587–608, November 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10426-9.

Reuter:2023:EMA

[Reu23] Friederike Reuter. Explorative mathematical argumentation: a theoretical framework for identifying and analysing argumentation processes in early mathematics learning. Educational Studies in Mathematics, 112(3):415–435, March 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10199-5.

Revuz:1968:PEM

[Rev68] André Revuz. Les pièges de l’enseignement des mathématiques. (French) [The pitfalls of teaching mathematics]. Educational Studies in Mathematics, 1(1–2):31–36, May 1968. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426229.

Revuz:1969:PAF

[Rev69] André Revuz. Les premiers pas en analyse. (French) [The first steps in analysis]. Educational Studies in Mathematics, 2(2–3): 270–278, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303463.

Revuz:1971:PGM

[Rev71] André Revuz. The position of geometry in mathematical education. Educational Studies in Mathematics, 4(1):48–52, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305796.

Revuz:1972:NCE

[Rev72] A. Revuz. La notion de continuité dans l’enseignement du second degré. (French) [The notion of continuity in secondary education]. Educational Studies in Mathematics, 4(3):281–298, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302578.

Revuz:1976:SPA

[Rev76] A. Revuz. Stratégies pour une approche de Z. (French) [Strategies for an approach to Z]. Educational Studies in Mathematics, 7(1–2):113–120, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00144364.

Revuz:1978:CME

[Rev78] A. Revuz. Change in mathematics education since the late 1950’s — ideas and realisation France. Educational Studies in Mathematics, 9(2):171–181, May 1978. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221157.

Radford:2007:ITM

[RFK07] Luis Radford, Fulvia Furinghetti, and Victor Katz. Introduction the topos of meaning or the encounter between past and present. Educational Studies in Mathematics, 66(2):107–110, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9076-7.

Rojano:2014:ISP

[RFP14] Teresa Rojano, Eugenio Filloy, and Luis Puig. Intertextuality and sense production in the learning of algebraic methods. Educational Studies in Mathematics, 87(3):389–407, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9561-3.

Recio:2001:IPM

[RG01] Angel M. Recio and Juan D. Godino. Institutional and personal meanings of mathematical proof. Educational Studies in Mathematics, 48(1):83–99, October 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015553100103.

Robertson:2019:EMT

[RG19] Sally-Ann Robertson and Mellony Graven. Exploratory mathematics talk in a second language: a sociolinguistic perspective. Educational Studies in Mathematics, 101(2):215– 232, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9840-5. See correction

[RG20].

Robertson:2020:CEM

[RG20] Sally-Ann Robertson and Mellony Graven. Correction to: Exploratory mathematics talk in a second language: a sociolinguistic perspective. Educational Studies in Mathematics, 103(1):135, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09914-
6;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09914-6.pdf. See [RG19].

Ruthven:2002:PMU

[RH02] Kenneth Ruthven and Sara Hennessy. A practitioner model of the use of computer-based tools and resources to support mathematics teaching and learning. Educational Studies in Mathematics, 49(1):47–88, January 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016052130572.

Ryve:2019:EPI

[RH19] Andreas Ryve and Kirsti Hemmi. Educational policy to improve mathematics instruction at scale: conceptualizing contextual factors. Educational Studies in Mathematics, 102 (3):379–394, November 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09887-
6;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09887-6.pdf.

Roy:2017:MRD

[RIA17] Somali Roy, Matthew Inglis, and Lara Alcock. Multimedia resources designed to support learning from written proofs: an eye-movement study. Educational Studies in Mathematics, 96(2):249–266, October 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-017-9754-7.pdf;   http://link.springer.
com/article/10.1007/s10649-017-9754-7.

Richard:2004:LFR

[Ric04] Philippe R. Richard. L’inférence figurale: Un pas de raisonnement discursivo-graphique. (French) [The figurative inference: A step of discursive-graphic reasoning]. Educational Studies in Mathematics, 57(2):229–263, September 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000049272.75852.c4.

Rico:2010:BRF

[Ric10] Luis Rico. Book review: The first century of the International Commission on Mathematical Instruction. Educational Studies in Mathematics, 73(2):205–210, March 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9211-3.

Rich:2020:BRD

[Ric20] Kathryn M. Rich. Book review: Digital technologies in designing mathematics education tasks: Potential and pitfalls. Allen Leung and Anna Baccaglini–Frank (Eds.): (2017). Educational Studies in Mathematics, 103(1):103–108, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09918-2.

Riley:1986:CLA

[Ril86] Mary S. Riley. Comments on Langford’s article. Educational Studies in Mathematics, 17(2):200–202, May 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311521. See [Lan86a, Lan86b].

Rivera:2007:ASS

[Riv07] Ferdinand D. Rivera. Accounting for students’ schemes in the development of a graphical process for solving polynomial inequalities in instrumented activity. Educational Studies in Mathematics, 65(3):281–307, July 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9052-2.

Rivera:2010:VTP

[Riv10] F. D. Rivera. Visual templates in pattern generalization activity. Educational Studies in Mathematics, 73(3):297–328, April 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9222-0.

Ruef:2020:WIL

[RJWB20] Jennifer L. Ruef, Michelle M. Jacob, G. Keith Walker, and Virginia R. Beavert. Why indigenous languages matter for mathematics education: a case study of Ichishkíin. Educational Studies in Mathematics, 104(3):313–332, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09957-0.

Remillard:2017:KCE

[RK17] Janine Remillard and Ok-Kyeong Kim. Knowledge of curriculum embedded mathematics: exploring a critical domain of teaching. Educational Studies in Mathematics, 96(1):65– 81, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9757-4.

Raymond:2000:CAR

[RL00] Anne M. Raymond and Marylin Leinenbach. Collaborative action research on the learning and teaching of algebra: a story of one mathematics teacher’s development. Educational Studies in Mathematics, 41(3):283–307, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004091631981.

Roth:2004:IUG

[RL04] Wolff-Michael Roth and Yew Jin Lee. Interpreting unfamiliar graphs: A generative, activity theoretic model. Educational Studies in Mathematics, 57(2):265–290, September 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000049276.37088.e4.

Rosa:2011:ROM

[RL11] Maurício Rosa and Stephen Lerman. Researching online mathematics education: opening a space for virtual learner identities. Educational Studies in Mathematics, 78(1):69–90, September 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9310-9.

Robinson:2012:IRB

[RL12] Katherine M. Robinson and Jo-Anne LeFevre. The inverse relation between multiplication and division: Concepts, procedures, and a cognitive framework. Educational Studies in Mathematics, 79(3):409–428, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9330-5.

Rigo-Lemini:2013:ESE

[RL13] Mirela Rigo-Lemini. Epistemic schemes and epistemic states. A study of mathematics convincement in elementary school classes. Educational Studies in Mathematics, 84(1):71–91, September 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9466-6.

Roth:2015:VIM

[RM15] Wolff-Michael Roth and Jean-François Maheux. The visible and the invisible: mathematics as revelation. Educational Studies in Mathematics, 88(2):221–238, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9582-y.

Rubel:2021:CMW

[RM21] Laurie H. Rubel and Andrea V. McCloskey. Contextualization of mathematics: which and whose world? Educational Studies in Mathematics, 107(2):383–404, June 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10041-4.

Ruffell:1998:SAM

[RMA98] Moyra Ruffell, John Mason, and Barbara Allen. Studying attitude to mathematics. Educational Studies in Mathematics, 35 (1):1–18, January 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003019020131.

Rodriguez-Muniz:2022:UCM

[RMAGMR22] Luis J. Rodríguez-Muñiz, Álvaro Aguilar-González, and Laura Muñiz-Rodríguez. Undergraduates’ conceptions of mathematics teaching and learning: an empirical study. Educational Studies in Mathematics, 109(3):523–547, March 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10105-5.

Rubel:2021:CMP

[RNC21] Laurie H. Rubel, Cynthia Nicol, and Anna Chronaki. A critical mathematics perspective on reading data visualizations: reimagining through reformatting, reframing, and renarrating. Educational Studies in Mathematics, 108(1–2):249–268, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10087-4.

Ryve:2012:EMT

[RNM12] Andreas Ryve, Per Nilsson, and John Mason. Establishing mathematics for teaching within classroom interactions in teacher education. Educational Studies in Mathematics, 81(1): 1–14, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-011-9371-9.

Ryve:2013:AEC

[RNP13] Andreas Ryve, Per Nilsson, and Kerstin Pettersson. Analyzing effective communication in mathematics group work: The role of visual mediators and technical terms. Educational Studies in Mathematics, 82(3):497–514, March 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9442-6.

Rodriguez-Nieto:2023:CUE

[RNRVM23] Camilo Andrés Rodríguez-Nieto, Flor Monserrat Rodríguez-Vásquez, and Vicenç Font Moll. Combined use of the extended theory of connections and the onto-semiotic approach to analyze mathematical connections by relating the graphs of f and f′. Educational Studies in Mathematics, 114(1):63–88, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10246-9.

Riordain:2009:RBP

[RO09] Máire Ní Ríordáin and John O’Donoghue. The relationship between performance on mathematical word problems and language proficiency for students learning through the medium of Irish. Educational Studies in Mathematics, 71(1):43–64, May 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9158-9.

Roberge:1975:DCL

[Rob75] James J. Roberge. Development of comprehension of logical connectives in symbolic or verbal form. Educational Studies in Mathematics, 6(2):207–212, July 1975. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302544.

Robinet:1986:RQM

[Rob86] J. Robinet. Les réels: Quels modèles en ont les élèves?. (French) [Reals: What models do students have?]. Educational Studies in Mathematics, 17(4):359–386, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311325.

Robitaille:1990:ACB

[Rob90] David F. Robitaille. Achievement comparisons between the first and second IEA studies of mathematics. Educational Studies in Mathematics, 21(5):395–414, October 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00398860.

Robert:1995:ADN

[Rob95]  Aline Robert.  Analyse des discours non strictement

mathématiques accompagnant des cours de mathématiques. (French) [Analysis of non-strictly mathematical discourses accompanying mathematics courses]. Educational Studies in Mathematics, 28(1):73–86, January 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273857.

Robotti:2012:NLT

[Rob12] Elisabetta Robotti. Natural language as a tool for analyzing the proving process: the case of plane geometry proof. Educational Studies in Mathematics, 80(3):433–450, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9383-0.

Rodd:2006:CME

[Rod06] Melissa Rodd. Commentary: Mathematics, emotion and special needs. Educational Studies in Mathematics, 63(2):227– 234, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-9014-0.

Roganovskij:1971:AAT

[Rog71] N. M. Roganovskij. Axiomatic approach to the teaching of solid geometry in grade IX. Educational Studies in Mathematics, 3(2):170–179, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305445.

Rogerson:1985:MMS

[Rog85] A. Rogerson. MISP — the Mathematics in Society Project. Educational Studies in Mathematics, 16(2):220–221, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02400940.

Roh:2008:SIT

[Roh08] Kyeong Hah Roh. Students’ images and their understanding of definitions of the limit of a sequence. Educational Studies in Mathematics, 69(3):217–233, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9128-2.

Roh:2010:ESS

[Roh10] Kyeong Hah Roh. An empirical study of students’ understanding of a logical structure in the definition of limit via the ε-strip activity. Educational Studies in Mathematics, 73 (3):263–279, April 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9210-4.

Roman:1974:OMR

[Rom74] T. Roman. Les olympiades mathématiques en Roumanie. (French) [The mathematical olympiads in Romania]. Educational Studies in Mathematics, 5(4):425–440, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01420656.

Ronda:2015:GPL

[Ron15] Erlina Ronda. Growth points in linking representations of function: a research-based framework. Educational Studies in Mathematics, 90(3):303–319, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9631-1.

Room:1971:GAR

[Roo71] Thomas G. Room. The geometry and algebra of reflections (and of 2 × 2 matrices). Educational Studies in Mathematics, 4(1):53–75, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305797.

Roos:2019:IME

[Roo19] Helena Roos. Inclusion in mathematics education: an ideology, a way of teaching, or both? Educational Studies in Mathematics, 100(1):25–41, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-018-9854-
z;   http://link.springer.com/content/pdf/10.1007/
s10649-018-9854-z.pdf.

Roos:2023:SVI

[Roo23] Helena Roos. Students’ voices of inclusion in mathematics education. Educational Studies in Mathematics, 113(2):229– 249, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10213-4.

Rosenbloom:1969:VS

[Ros69] Paul C. Rosenbloom. Vectors and symmetry. Educational Studies in Mathematics, 2(2–3):405–414, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303472.

Rostami:1978:CME

[Ros78] Badiollah Rostami. Change in mathematics education since the late 1950’s — ideas and realisation Iran. Educational Studies in Mathematics, 9(2):255–260, May 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221161.

Roth:2012:RWS

[Rot12] Wolff-Michael Roth. Re/writing the subject: a contribution to post-structuralist theory in mathematics education. Educational Studies in Mathematics, 80(3):451–473, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9375-5.

Roth:2014:RBB

[Rot14] Wolff-Michael Roth. Rules of bending, bending the rules: the geometry of electrical conduit bending in college and workplace. Educational Studies in Mathematics, 86(2):177–192, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9376-4.

Roth:2016:GMM

[Rot16] Wolff-Michael Roth. Growing-making mathematics: a dynamic perspective on people, materials, and movement in classrooms. Educational Studies in Mathematics, 93(1):87– 103, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9695-6.

Roth:2017:APC

[Rot17] Wolff-Michael Roth. Astonishment: a post-constructivist investigation into mathematics as passion. Educational Studies in Mathematics, 95(1):97–111, May 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Rott:2021:IDJ

[Rot21] Benjamin Rott. Inductive and deductive justification of knowledge: epistemological beliefs and critical thinking at the beginning of studying mathematics. Educational Studies in Mathematics, 106(1):117–132, January 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10004-1.

Roumanet:1968:EEM

[Rou68]  André Roumanet.  Une expérience d’enseignement de

mathématique avec des enfants de 11 à 13 ans. (French) [A mathematics-teaching experiment with 11-to-13-year-old children]. Educational Studies in Mathematics, 1(1–2):222– 236, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00426245;   http:/
/link.springer.com/article/10.1007/BF00426245.

Roumanet:1969:CMM

[Rou69] A. Roumanet. Une classe de mathématique: Motivations et méthodes. (French) [A class of mathematics: Motivations and methods]. Educational Studies in Mathematics, 2(2–3):212– 231, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303459.

Rowland:1995:HMT

[Row95] Tim Rowland. Hedges in mathematics talk: Linguistic pointers to uncertainty. Educational Studies in Mathematics, 29(4): 327–353, December 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273910.

Rowland:2008:PDU

[Row08] Tim Rowland. The purpose, design and use of examples in the teaching of elementary mathematics. Educational Studies in Mathematics, 69(2):149–163, October 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9148-y.

Royoux:1981:ERT

[Roy81] F. Royoux. Sur l’enseignement de la relation transitive à l’école élémentaire. (French) [On the teaching of the transitive relation in the elementary school]. Educational Studies in Mathematics, 12(1):89–111, February 1981. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386048.

Radford:2007:SMS

[RP07] Luis Radford and Luis Puig. Syntax and meaning as sensuous, visual, historical forms of algebraic thinking. Educational Studies in Mathematics, 66(2):145–164, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9024-6.

Reinholz:2021:SSP

[RP21] Daniel L. Reinholz and Mary E. Pilgrim. Student sensemaking of proofs at various distances: the role of epistemic, rhetorical, and ontological distance in the peer review process. Educational Studies in Mathematics, 106(2):211–229, February 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-10010-3.

Rotem:2024:UCI

[RPP24] Sigal-Hava Rotem, Despina Potari, and Giorgos Psycharis. Using critical incidents as a tool for promoting prospective teachers’ noticing during reflective discussions in a fieldwork-based university course. Educational Studies in Mathematics, 117 (1):67–95, September 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10336-2.

Resek:1980:CMC

[RR80] Diane Resek and William H. Rupley. Combatting ‘mathophobia’ with a conceptual approach toward mathematics. Educational Studies in Mathematics, 11(4):423–441, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00231215.

Robert:2005:CAM

[RR05] Aline Robert and Janine Rogalski. A cross-analysis of the mathematics teacher’s activity. An example in a French 10thgrade class. Educational Studies in Mathematics, 59(1–3):269– 298, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-5890-6.

Radford:2011:IEC

[RR11] Luis Radford and Wolff-Michael Roth. Intercorporeality and ethical commitment: an activity perspective on classroom interaction. Educational Studies in Mathematics, 77(2–3):227– 245, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9282-1.

Radford:2017:AME

[RR17] Luis Radford and Wolff-Michael Roth. Alienation in mathematics education: a problem considered from neo-Vygotskian approaches. Educational Studies in Mathematics, 96(3):367– 380, November 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9769-0.

Rupnow:2023:NMD

[RR23] Rachel Rupnow and Brooke Randazzo. Norms of mathematical definitions: imposing constraints, permitting choice, or both? Educational Studies in Mathematics, 114(2):297–314, October 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10227-y.

Reys:1991:EPS

[RRP91] Barbara J. Reys, Robert E. Reys, and Alfinio Flores Peñafiel. Estimation performance and strategy use of Mexican 5th and 8th grade student sample. Educational Studies in Mathematics, 22(4):353–375, August 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00369295.

Rupnow:2022:SAV

[RS22] Rachel Rupnow and Peter Sassman. Sameness in algebra: views of isomorphism and homomorphism. Educational Studies in Mathematics, 111(1):109–126, September 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10162-4.

Rum:2023:BRE

[RSD23] Ahbi Mahdianing Rum, Ahmad Hasan Saifurrisal, and Mia Roosmalisa Dewi. Book review: Exploring how to teach multiplication through the Japanese and Ibero–American theories. Masami Isoda & Raimundo Olfos (Eds.). (2021) Teaching multiplication with lesson study: Japanese and Ibero– American theories for international mathematics education. Educational Studies in Mathematics, 114(3):525–536, November 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10238-9.

Rellensmann:2017:MDE

[RSL17] Johanna Rellensmann, Stanislaw Schukajlow, and Claudia Leopold. Make a drawing. Effects of strategic knowledge, drawing accuracy, and type of drawing on students’ mathematical modelling performance. Educational Studies in Mathematics, 95(1):53–78, May 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Ramirez-Sanchez:2023:SMQ

[RSRVRVVR23] Camilo Andrés Ramírez-Sánchez, Avenilde Romo-Vázquez, Rebeca Romo-Vázquez, and Diana Velásquez-Rojas. Study of modeling questions in a first-year university mathematics online course. Educational Studies in Mathematics, 114(3): 503–524, November 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10261-w.

Radford:2011:SMM

[RSS11] Luis Radford, Gert Schubring, and Falk Seeger. Signifying and meaning-making in mathematical thinking, teaching, and learning. Educational Studies in Mathematics, 77(2–3):149– 156, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9322-5.

Rabijewska:1985:MCB

[RT85] Barbara Rabijewska and Mieczysław Tra̧d. A mathematical camp for bright pupils. Educational Studies in Mathematics, 16(1):41–57, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00354882.

Roth:2009:BEM

[RT09] Wolff-Michael Roth and Jennifer S. Thom. Bodily experience and mathematical conceptions: from classical views to a phenomenological reconceptualization. Educational Studies in Mathematics, 70(2):175–189, March 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9138-0.

Roth:2014:UVD

[RT14] Wolff-Michael Roth and Shelby Temple. On understanding variability in data: a study of graph interpretation in an advanced experimental biology laboratory. Educational Studies in Mathematics, 86(3):359–376, July 2014.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9535-5.

Roschelle:2008:RSR

[RTSK08] J. Roschelle, D. Tatar, N. Shechtman, and J. Knudsen. The role of scaling up research in designing for and evaluating robustness. Educational Studies in Mathematics, 68(2):149– 170, June 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9119-3.

Ruede:2013:HSL

[Rüe13] Christian Rüede. How secondary level teachers and students impose personal structure on fractional expressions and equations — an expert-novice study. Educational Studies in Mathematics, 83(3):387–408, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9462-2.

Ruthven:1987:ASM

[Rut87] Kenneth Ruthven. Ability stereotyping in mathematics. Educational Studies in Mathematics, 18(3):243–253, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386197.

Ruthven:1989:EAA

[Rut89a] Kenneth Ruthven. An exploratory approach to advanced mathematics. Educational Studies in Mathematics, 20(4):449– 467, November 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00315610.

Ruthven:1989:GE

[Rut89b] Kenneth Ruthven. Guest editorial. Educational Studies in Mathematics, 20(3):233–235, August 1989. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310872.

Ruthven:1990:IGC

[Rut90] Kenneth Ruthven. The influence of graphic calculator use on translation from graphic to symbolic forms. Educational Studies in Mathematics, 21(5):431–450, October 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00398862.

Ruthven:1994:BJR

[Rut94] Kenneth Ruthven. Better judgement: Rethinking assessment in mathematics education. Educational Studies in Mathematics, 27(4):433–450, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273382.

Ruthven:1996:E

[Rut96] Kenneth Ruthven. Editorial. Educational Studies in Mathematics, 30(1):1–2, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00163749.

Ruthven:1998:APAa

[Rut98a] Kenneth Ruthven. Advice to prospective authors. Educational Studies in Mathematics, 36(2):III–IV, July 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003417109009.

Ruthven:1998:APAb

[Rut98b] Kenneth Ruthven. Advice to prospective authors. Educational Studies in Mathematics, 37(2):i–iv, November 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003172004833.

Ruthven:1999:APAa

[Rut99a] Kenneth Ruthven. Advice to prospective authors. Educational Studies in Mathematics, 40(1):v–vi, September 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017185819455.

Ruthven:1999:APAb

[Rut99b] Kenneth Ruthven. Advice to prospective authors. Educational Studies in Mathematics, 40(2):iii–iv, October 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017127104434.

Ruthven:1999:APAc

[Rut99c] Kenneth Ruthven. Advice to prospective authors. Educational Studies in Mathematics, 40(3):iii–iv, November 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017183206251.

Ruthven:2000:APA

[Rut00a] Kenneth Ruthven. Advise to prospective authors. Educational Studies in Mathematics, 42(1):v–vi, April 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017271606196.

Ruthven:2000:Ea

[Rut00b] Kenneth Ruthven. Editorial. Educational Studies in Mathematics, 41(1):vii, January 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017218718754.

Ruthven:2000:Eb

[Rut00c] Kenneth Ruthven. Editorial. Educational Studies in Mathematics, 42(1):vii, April 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1004161223096.

Ruthven:2002:BRB

[Rut02] Kenneth Ruthven. Book review: Barbin, E., Duval, R., Giorgiutti, I., Houdebine, J. and Laborde, C. (Eds.), Produire et lire des textes de démonstration, Paris: Ellipses, 2001. Educational Studies in Mathematics, 51(1–2):145–147, July 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022408100104.

Ruthven:2005:AID

[Rut05] Kenneth Ruthven. Alternative interpretation of the dataset on K–12 calculator usage and college grades as analysed by Wilson and Naiman (2003, 2004). Educational Studies in Mathematics, 60(3):383–387, November 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-0445-4.

Ruthven:2010:BRS

[Rut10] Kenneth Ruthven. Book review: Summing up mathematics teacher education. Educational Studies in Mathematics, 73 (1):87–97, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-009-9209-x.

Radu:2011:ERM

[RW11a] Iuliana Radu and Keith Weber. Erratum to: Refinements in mathematics undergraduate students’ reasoning on completed infinite iterative processes. Educational Studies in Mathematics, 78(2):181–182, November 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-011-9332-3.pdf. See [RW11b].

Radu:2011:RMU

[RW11b] Iuliana Radu and Keith Weber. Refinements in mathematics undergraduate students’ reasoning on completed infinite iterative processes. Educational Studies in Mathematics, 78 (2):165–180, November 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-

9318-1. See erratum [RW11a].

Roth:2019:AEM

[RW19] Wolff-Michael Roth and Margaret Walshaw. Affect and emotions in mathematics education: toward a holistic psychology of mathematics education. Educational Studies in Mathematics, 102(1):111–125, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09899-

Rasmussen:2015:EIC

[RWZ15] Chris Rasmussen, Megan Wawro, and Michelle Zandieh. Examining individual and collective level mathematical progress. Educational Studies in Mathematics, 88(2):259–281, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9583-x.

Ryve:2004:CCC

[Ryv04] Andreas Ryve. Can collaborative concept mapping create mathematically productive discourses? Educational Studies in Mathematics, 56(2–3):157–177, July 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000040395.17555.c2.

Ryve:2006:MEA

[Ryv06] Andreas Ryve. Making explicit the analysis of students’ mathematical discourses — revisiting a newly developed methodological framework. Educational Studies in Mathematics, 62 (2):191–209, June 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-4834-0.

Setati:2000:BLD

[SA00] Mamokgethi Setati and Jill Adler. Between languages and discourses: Language practices in primary multilingual mathematics classrooms in South Africa. Educational Studies in Mathematics, 43(3):243–269, November 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1011996002062.

Shayer:2007:FCD

[SA07] Michael Shayer and Mundher Adhami. Fostering cognitive development through the context of mathematics: Results of the CAME Project. Educational Studies in Mathematics, 64 (3):265–291, March 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-9037-1.

Shvarts:2021:EIL

[SAD21] Anna Shvarts, Rosa Alberto, and Paul Drijvers. Embodied instrumentation in learning mathematics as the genesis of a body-artifact functional system. Educational Studies in Mathematics, 107(3):447–469, July 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10053-0.

Saenz:2009:RCC

[Sáe09] César Sáenz. The role of contextual, conceptual and procedural knowledge in activating mathematical competencies (PISA). Educational Studies in Mathematics, 71(2):123–143, June 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9167-8.

Sajka:2003:SSS

[Saj03] M. Sajka. A secondary school student’s understanding of the concept of function — a case study. Educational Studies in Mathematics, 53(3):229–254, September 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026033415747.

Samurcay:1985:SFD

[Sam85] Renan Samurçay. Signification et fonctionnement du concept de variable informatique chez des élèves debutants. (French) [Meaning and functioning of the computer variable concept in early students]. Educational Studies in Mathematics, 16 (2):143–161, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02400935.

Santi:2011:OSF

[San11] George Santi. Objectification and semiotic function. Educational Studies in Mathematics, 77(2–3):285–311, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9296-8.

Sarrazy:2002:HEM

[Sar02] Bernard Sarrazy. Les hétérogénéités dans l’enseignement des mathématiques. (French) [Heterogeneities in the teaching of mathematics]. Educational Studies in Mathematics, 49(1): 89–117, January 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1016006805418.

Schukajlow:2019:DCM

[SAR19] Stanislaw Schukajlow, Kay Achmetli, and Katrin Rakoczy. Does constructing multiple solutions for real-world problems affect self-efficacy? Educational Studies in Mathematics, 100 (1):43–60, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9847-y.

Sayers:1994:GDM

[Say94] Roy Sayers. Gender differences in mathematics education in Zambia. Educational Studies in Mathematics, 26(4):389–403, June 1994. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01279522.

Schmidt:1997:RAA

[SB97] Sylvine Schmidt and Nadine Bednarz. Raisonnements arithmétiques et algébriques dans un contexte de résolution de problèmes: difficultés rencontrées par les futurs enseignants. (French) [Arithmetic and algebraic reactions in a context of problem solving: difficulties encountered by future teachers]. Educational Studies in Mathematics, 32(2):127–155, February 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002946500851.

Stinson:2012:CPT

[SB12] David W. Stinson and Erika C. Bullock. Critical postmodern theory in mathematics education research: a praxis of uncertainty. Educational Studies in Mathematics, 80(1–2):41– 55, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9386-x.

Schindler:2020:AFD

[SB20] Maike Schindler and Arthur Bakker. Affective field during collaborative problem posing and problem solving: a case study. Educational Studies in Mathematics, 105(3):303–324, November 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09973-0.

Schou:2022:UHV

[SBA22] Marit Hvalsøe Schou and Angelika Bikner-Ahsbahs. Unpacking hidden views: seven ways to treat your formula. Educational Studies in Mathematics, 109(3):639–659, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10092-7.

Shvarts:2024:RAA

[SBDD24] Anna Shvarts, Rogier Bos, Michiel Doorman, and Paul Drijvers. Reifying actions into artifacts: process–object duality from an embodied perspective on mathematics learning. Educational Studies in Mathematics, 117(2):193–214, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10310-y.

Scandura:1969:URB

[SBDM69] Joseph M. Scandura, Joan Barksdale, John H. Durnin, and Robert Mcgee. An unexpected relationship between failure and subsequent mathematics learning. Educational Studies in Mathematics, 1(3):247–251, January 1969. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00558310.

Satyam:2022:AGL

[SBL22] V. Rani Satyam, Younggon Bae, and Mariana Levin. Affect graphing: leveraging graphical representations in the study of students’ affect in mathematics. Educational Studies in Mathematics, 110(3):481–501, July 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10131-3.

Sierpinska:2011:TAV

[SBP11] Anna Sierpinska, Georgeana Bobos, and Andreea Pruncut. Teaching absolute value inequalities to mature students. Educational Studies in Mathematics, 78(3):275–305, December 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9325-2.

Spalt:1986:BR

[SBS86] Detlef D. Spalt, Leone Burton, and Erhard Scholz. Book review. Educational Studies in Mathematics, 17(4):443–452, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311329.

Swetz:1983:MEE

[SC83] Frank J. Swetz and Albert Y. Chi. Mathematics entrance examinations in Chinese institutions of higher education. Educational Studies in Mathematics, 14(1):39–54, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00704701.

Simmons:1990:FKA

[SC90] Malcolm Simmons and Peter Cope. Fragile knowledge of angle in turtle geometry. Educational Studies in Mathematics, 21(4): 375–382, August 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00304265.

Simmons:1993:ARE

[SC93] Malcolm Simmons and Peter Cope. Angle and rotation: Effects of different types of feedback on the quality of response. Educational Studies in Mathematics, 24(2):163–176, February 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273690.

Sarouphim:2017:MEL

[SC17] Ketty M. Sarouphim and Madona Chartouny. Mathematics education in Lebanon: gender differences in attitudes and achievement. Educational Studies in Mathematics, 94(1):55– 68, January 2017. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-016-9712-9.

Scandura:1971:RBT

[Sca71] Joseph M. Scandura. A research basis for teacher education. Educational Studies in Mathematics, 3(2):229–243, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305451.

Scandura:1975:HDM

[Sca75] Joseph M. Scandura. How does mathematics learning take place? Educational Studies in Mathematics, 6(3):375–385, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01793618.

Schuster:1971:TGP

[Sch71] Seymour Schuster. On the teaching of geometry a potpourri. Educational Studies in Mathematics, 4(1):76–86, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305798.

Schipper:1983:TPT

[Sch83] Wilhelm Schipper. The topological primacy thesis: Genetic and didactic aspects. Educational Studies in Mathematics, 14 (3):285–296, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00410543.

Schneider:1992:PAD

[Sch92] Maggy Schneider. À propos de l’apprentissage du taux de variation instantané. (French) [On learning about the instantaneous rate of change]. Educational Studies in Mathematics, 23(4):317–350, August 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00302439.

Schubring:2011:CRE

[Sch11a] Gert Schubring. Conceptions for relating the evolution of mathematical concepts to mathematics learning — epistemology, history, and semiotics interacting. Educational Studies in Mathematics, 77(1):79–104, May 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9301-x.

Schubring:2011:HTB

[Sch11b] Gert Schubring. On how n and i turned out to become indices in mathematical sequences and formulae. Educational Studies in Mathematics, 77(2–3):331–350, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9286-x.

Schneider:2012:CKI

[Sch12] Michael Schneider. Commentary 2: knowledge integration in mathematics learning: the case of inversion. Educational Studies in Mathematics, 79(3):447–453, March 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-011-9373-7.pdf.

Schreiber:2013:SPC

[Sch13] Christof Schreiber. Semiotic processes in chat-based problemsolving situations. Educational Studies in Mathematics, 82 (1):51–73, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9417-7.

Scheiner:2016:NLO

[Sch16] Thorsten Scheiner. New light on old horizon: Constructing mathematical concepts, underlying abstraction processes, and sense making strategies. Educational Studies in Mathematics, 91(2):165–183, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9665-4.

Scheiner:2020:DOT

[Sch20] Thorsten Scheiner. Dealing with opposing theoretical perspectives: knowledge in structures or knowledge in pieces? Educational Studies in Mathematics, 104(1):127–145, May 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09950-7.

Scheiner:2023:SWP

[Sch23] Thorsten Scheiner. Shifting the ways prospective teachers frame and notice student mathematical thinking: from deficits to strengths. Educational Studies in Mathematics, 114(1): 35–61, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10235-y. See

correction [Sch25].

Schoenfeld:2024:BRC

[Sch24] Alan H. Schoenfeld. Book review: Charting a career and its influence. Jinfa Cai, Gabriel J. Stylianides, and Patricia Ann Kenney (Eds.) (2023) Research studies on learning and teaching of mathematics: dedicated to Edward A. Silver. Educational Studies in Mathematics, 117(2):329–338, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10316-6.

Scheiner:2025:ACS

[Sch25] Thorsten Scheiner. Author correction: Shifting the ways prospective teachers frame and notice student mathematical thinking: from deficits to strengths. Educational Studies in Mathematics, 119(2):419–420, June 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10402-3. See [Sch23].

Shama:1994:VAM

[SD94] Gilli Shama and Tommy Dreyfus. Visual, algebraic and mixed strategies in visually presented linear programming problems. Educational Studies in Mathematics, 26(1):45–70, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273300.

Schwarz:1995:NAU

[SD95] Baruch Schwarz and Tommy Dreyfus. New actions upon old objects: A new ontological perspective on functions. Educational Studies in Mathematics, 29(3):259–291, October 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274094.

Shield:2013:APM

[SD13] Malcolm Shield and Shelley Dole. Assessing the potential of mathematics textbooks to promote deep learning. Educational Studies in Mathematics, 82(2):183–199, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9415-9.

Splinter:2025:SCP

[SDVT25] Suzanne Elise Splinter, Fien Depaepe, Lieven Verschaffel, and Joke Torbeyns. Supporting cardinality principle development: the contribution of instruction type and its interplay with child characteristics. Educational Studies in Mathematics, 119 (2):335–365, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-025-10392-2.

Shillor:1993:KAC

[SE93] Irith Shillor and Bridget Egan. The King Alfred’s College maths game: Problem solving and mathematical activity. Educational Studies in Mathematics, 24(3):313–317, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01275429.

Singer:2013:PPR

[SEC13] Florence Mihaela Singer, Nerida Ellerton, and Jinfa Cai. Problem-posing research in mathematics education: new questions and directions. Educational Studies in Mathematics, 83(1):1–7, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9478-2.

Schwarts:2023:STD

[SECP^{+}23] Gil Schwarts, Avital Elbaum-Cohen, Birte Pöhler, Susanne Prediger, Abraham Arcavi, and Ronnie Karsenty. The servants of two discourses: how novice facilitators draw on their mathematics teaching experience. Educational Studies in Mathematics, 112(2):247–266, February 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10182-0.

Sedivy:1969:TEL

[Šed69] Jaroslav Šedivý. Teaching of elements of logic by means of finite graphs. Educational Studies in Mathematics, 1(4):422– 439, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02396316.

Sedivy:1976:NRP

[Šed76] Jaroslav Šedivý. A note on the role of parameters in mathematics teaching. Educational Studies in Mathematics, 7(1– 2):121–126, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00144365.

Seeger:2001:RDM

[See01] Falk Seeger. Research on discourse in the mathematics classroom: A commentary. Educational Studies in Mathematics, 46(1–3):287–297, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014092200410.

Seeger:2011:MMM

[See11] Falk Seeger. On meaning making in mathematics education: social, emotional, semiotic. Educational Studies in Mathematics, 77(2–3):207–226, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9279-9.

Sekiguchi:2021:ASA

[Sek21] Yasuhiro Sekiguchi. Activity systems analysis of classroom teaching and learning of mathematics: a case study of Japanese secondary schools. Educational Studies in Mathematics, 108(3):493–511, November 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10069-6.

Selter:1998:BCM

[Sel98] Christoph Selter. Building on children’s mathematics — a teaching experiment in grade three. Educational Studies in Mathematics, 36(1):1–27, June 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003111425076.

Selter:2000:BRV

[Sel00] Christoph Selter. Book review: Verschaffel, L., Greer, B., and de Corte, E., Making Sense of Word Problems. Educational Studies in Mathematics, 42(2):211–213, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004190927303.

Selter:2001:AST

[Sel01] Christoph Selter. Addition and subtraction of three-digit numbers: German elementary children’s success, methods and strategies. Educational Studies in Mathematics, 47(2):145– 173, July 2001. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1014521221809.

Semusin:1970:SGF

[Sem70] A. D. Semušin. Study of geometry in the fourth grade. Educational Studies in Mathematics, 3(1):68–83, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381595.

Semadeni:1984:PCP

[Sem84] Zbigniew Semadeni. A principle of concretization permanence for the formation of arithmetical concepts. Educational Studies in Mathematics, 15(4):379–395, November 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311113.

Semadeni:2008:DIL

[Sem08] Zbigniew Semadeni. Deep intuition as a level in the development of the concept image. Educational Studies in Mathematics, 68(1):1–17, May 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-007-
9105-1.

Sensevy:1996:FPF

[Sen96] Gerard Sensevy. Fabrication de problèmes de fraction par des éléves à la fin de l’enseignement élémentaire. (French) [Manufacturing of fraction problems by pupils at the end of elementary education]. Educational Studies in Mathematics, 30 (3):261–288, April 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00304568.

Senger:1998:RRM

[Sen98] Elizabeth Smith Senger. Reflective reform in mathematics: The recursive nature of teacher change. Educational Studies in Mathematics, 37(3):199–221, December 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003539410030.

Servais:1968:CEM

[Ser68] W. Servais. Comment enseigner la mathématique pour qu’elle soit utile?. (French) [How can we teach mathematics to be useful?]. Educational Studies in Mathematics, 1(1–2):37–54, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426230.

Servais:1969:LEM

[Ser69] W. Servais. Logique et enseignement mathématique. (French) [Logic and mathematical education]. Educational Studies in Mathematics, 2(2–3):160–179, December 1969. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303455.

Servais:1976:PAM

[Ser76] W. Servais. Problématique dans l’apprentissage de la mathématique. (French) [Problems in the learning of mathematics]. Educational Studies in Mathematics, 7(1–2):127– 137, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00144366.

Sevinc:2022:TRM

[Sev22] Serife Sevinc. Toward a reconceptualization of model development from models-and-modeling perspective in mathematics education. Educational Studies in Mathematics, 109(3):611– 638, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10096-3.

Schmidt:1984:BR

[SF84] Siegbert Schmidt and T. J. Fletcher. Book reviews. Educational Studies in Mathematics, 15(1):95–102, February 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00380441.

Segura:2023:PST

[SF23] Carlos Segura and Irene Ferrando. Pre-service teachers’ flexibility and performance in solving Fermi problems. Educational Studies in Mathematics, 113(2):207–227, June 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10220-5.

Sfard:1991:DNM

[Sfa91] Anna Sfard. On the dual nature of mathematical conceptions: Reflections on processes and objects as different sides of the same coin. Educational Studies in Mathematics, 22(1): 1–36, February 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00302715.

Sfard:2001:TMD

[Sfa01] Anna Sfard. There is more to discourse than meets the ears: Looking at thinking as communicating to learn more about mathematical learning. Educational Studies in Mathematics, 46(1–3):13–57, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014097416157.

Sfard:2005:WCM

[Sfa05] Anna Sfard. What could be more practical than good research? Educational Studies in Mathematics, 58(3):393–413, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-4818-5.

Sfard:2009:WAF

[Sfa09] Anna Sfard. What’s all the fuss about gestures? A commentary. Educational Studies in Mathematics, 70(2):191–200, March 2009. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-008-9161-1.pdf.

Sfard:2013:AYA

[Sfa13] Anna Sfard. Almost 20 years after: Developments in research on language and mathematics. Review of J. N. Moschkovich (Ed.) (2010) Language and mathematics education: Multiple perspectives and directions for research. Educational Studies in Mathematics, 82(2):331–339, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9446-2.

Sfard:2015:BRW

[Sfa15] Anna Sfard. Book review: When Oksapmin number system met Western arithmetic — answering the question of how cognitive and cultural change induce each other. [Saxe, G. B. (2012) Cultural Development of Mathematical Ideas. Papua New Guinea studies Cambridge, UK: Cambridge University Press. 393 pages. Hardback: ISBN: 978-0-521-76166-6. Paperback: ISBN: 978-1-107-68569-7. Online ISBN: 978-1-139-

04536-0. Book DOI: https://doi.org/10.1017/CBO9781139045360.

£64]. Educational Studies in Mathematics, 88(2):283–290, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9586-7.

Shield:1998:ASE

[SG98a] Mal Shield and Peter Galbraith. The analysis of student expository writing in mathematics. Educational Studies in Mathematics, 36(1):29–52, June 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003109819256.

Stillman:1998:AMR

[SG98b]  Gloria A. Stillman and Peter L. Galbraith.  Apply-

ing mathematics with real world connections: metacognitive characteristics of secondary students. Educational Studies in Mathematics, 36(2):157–194, July 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003246329257.

Schorr:2008:SEA

[SG08] Roberta Y. Schorr and Gerald A. Goldin. Students’ expression of affect in an inner-city simcalc classroom. Educational Studies in Mathematics, 68(2):131–148, June 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9117-5.

Sanchez:2014:SMN

[SG14] Victoria Sánchez and Mercedes García. Sociomathematical and mathematical norms related to definition in pre-service primary teachers’ discourse. Educational Studies in Mathematics, 85(2):305–320, February 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-013-
9516-0.

Sherman:2025:CAW

[SG25] Julie Sherman and José F. Gutiérrez. Classifying arithmetic word problems from the Estelle Reel Papers collection. Educational Studies in Mathematics, 120(3):609–634, November 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10421-0.

Shaughnessy:2021:FAP

[SGB21] Meghan Shaughnessy, Nicole M. Garcia, and Deborah Loewenberg Ball. Formatively assessing prospective teachers’ skills in leading mathematics discussions. Educational Studies in Mathematics, 108(3):451–472, November 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10070-z.

Scheiner:2022:MTA

[SGC22] Thorsten Scheiner, Juan D. Godino, and Nuria Climent. On metaphors in thinking about preparing mathematics for teaching. Educational Studies in Mathematics, 111(2):253–270, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10154-4.

Soneira:2018:ASR

[SGCA18] Carlos Soneira, José Antonio González-Calero, and David Arnau. An assessment of the sources of the reversal error through classic and new variables. Educational Studies in Mathematics, 99(1):43–56, September 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9828-1.

Soneira:2023:EAL

[SGCA23] Carlos Soneira, José Antonio González-Calero, and David Arnau. Effect of algebraic language and problem text wording on problem model accuracy when solving age word problems. Educational Studies in Mathematics, 114(1):109–127, September 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10236-x.

Sanchez:2024:DHS

[SGRS24] Ernesto Sánchez, Victor Nozair García-Ríos, and Francisco Sepúlveda. Development of high school students’ conceptions of sampling distribution in the context of learning significance tests with technology. Educational Studies in Mathematics, 117(2):215–238, October 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10330-8.

Shayshon:2014:TMT

[SGTK14] Bruria Shayshon, Hagar Gal, Bertha Tesler, and Eun-Sung Ko. Teaching mathematically talented students: a cross-cultural study about their teachers’ views. Educational Studies in Mathematics, 87(3):409–438, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9568-9.

Shaughnessy:1977:MPE

[Sha77] J. Michael Shaughnessy. Misconceptions of probability: An experiment with a small-group, activity-based, model building approach to introductory probability at the college level. Educational Studies in Mathematics, 8(3):295–316, October 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00385927.

Sharfunddin:1978:CME

[Sha78] S. M. Sharfunddin. Change in mathematics education since the late 1950’s — ideas and realisation Bangladesh. Educational Studies in Mathematics, 9(2):159–170, May 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00221156.

Shama:1998:UPP

[Sha98] Gilli Shama. Understanding periodicity as a process with a gestalt structure. Educational Studies in Mathematics, 35(3): 255–281, March 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1003173805389.

Stacey:2001:EEF

[SHAC01] Kaye Stacey, Sue Helme, Shona Archer, and Caroline Condon. The effect of epistemic fidelity and accessibility on teaching with physical materials: A comparison of two models for teaching decimal numeration. Educational Studies in Mathematics, 47(2):199–221, July 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1014590319667.

Sherman:1987:CN

[She87] B. F. Sherman. A cautionary note. Educational Studies in Mathematics, 18(3):325–326, August 1987. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386202.

Solomon:2021:RAG

[SHG21] Yvette Solomon, Susan Hough, and Stephen Gough. The role of appropriation in guided reinvention: establishing and preserving devolved authority with low-attaining students. Educational Studies in Mathematics, 106(2):171–188, February 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-09998-5.

Shongwe:2021:LBA

[Sho21] Benjamin Shongwe. Learners’ beliefs about the functions of proof: building an argument for validity. Educational Studies in Mathematics, 107(3):503–523, July 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10047-y.

Shriki:2010:WLR

[Shr10] Atara Shriki. Working like real mathematicians: developing prospective teachers’ awareness of mathematical creativity through generating new concepts. Educational Studies in Mathematics, 73(2):159–179, March 2010. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9212-2.

Shvarts:2022:BRT

[Shv22] Anna Shvarts. Book review: Transforming mathematics education: from embodied experiences to an ethical commitment. Luis Radford (2021) The theory of objectification: a Vygotskian perspective on knowing and becoming in mathematics teaching and learning, Brill. xvi, 259 pages. Hardcover: ISBN: 978-90-04-45965-6, EUR 130.00. Paperback: ISBN: 978-90-04-45964-9, EUR 60.00. eBook (PDF): ISBN: 978-90-04-45966-3, EUR 130.00. Educational Studies in Mathematics, 109(3):677– 686, March 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10095-4.

Sianturi:2024:TBI

[Sia24] Iwan A. J. Sianturi. Teachers’ beliefs on integrating children’s literature in mathematics teaching and learning in Indonesia. Educational Studies in Mathematics, 117(1):97–119, September 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10334-4.

Sidoli:2005:RNS

[Sid05] Nathan Sidoli. Reviel Netz, the shaping of deduction in Greek mathematics: A study in cognitive history. Educational Studies in Mathematics, 58(2):277–282, February 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-4033-4.

Sierpinska:1987:HSE

[Sie87] Anna Sierpińska. Humanities students and epistemological obstacles related to limits. Educational Studies in Mathematics, 18(4):371–397, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00240986.

Sierpinska:2001:E

[Sie01] Anna Sierpinska. Editorial. Educational Studies in Mathematics, 47(1):1–5, May 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017959528957.

Sierpinska:2002:E

[Sie02] Anna Sierpinska. Editorial. Educational Studies in Mathematics, 50(1):i–ii, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1020975325670.

Sierpinska:2003:E

[Sie03] Anna Sierpinska. Editorial. Educational Studies in Mathematics, 53(1):3–4, May 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1024614913886.

Sierpinska:2004:BSR

[Sie04] Anna Sierpinska. Breaking the silence and revealing the backstage. Educational Studies in Mathematics, 57(3):301, November 2004. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-004-5932-5.

Silver:2013:PPR

[Sil13] Edward A. Silver. Problem-posing research in mathematics education: looking back, looking around, and looking ahead. Educational Studies in Mathematics, 83(1):157–162, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9477-3.

Silver:2025:BRM

[Sil25] Edward A. Silver. Book review: Meeting the challenge of mathematical challenge in mathematics education. Roza Leikin (Ed.) (2023) Mathematical Challenges for All. Educational Studies in Mathematics, 120(2):??, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10415-y.

Simon:1994:LML

[Sim94] Martin A. Simon. Learning mathematics and learning to teach: Learning cycles in mathematics teacher education. Educational Studies in Mathematics, 26(1):71–94, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273301.

Simon:1996:BID

[Sim96] Martin A. Simon. Beyond inductive and deductive reasoning: The search for a sense of knowing. Educational Studies in Mathematics, 30(2):197–210, March 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302630.

Simpson:2015:AMP

[Sim15] Adrian Simpson. The anatomy of a mathematical proof: Implications for analyses with Toulmin’s scheme. Educational Studies in Mathematics, 90(1):1–17, September 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9616-0.

Simon:2017:EMC

[Sim17] Martin A. Simon. Explicating mathematical concept and mathematical conception as theoretical constructs for mathematics education research. Educational Studies in Mathematics, 94(2):117–137, February 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9728-1.

Simpson:2020:MES

[Sim20] Adrian Simpson. On the misinterpretation of effect size. Educational Studies in Mathematics, 103(1):125–133, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09924-4.

Singh:2000:UCP

[Sin00] Parmjit Singh. Understanding the concepts of proportion and ratio constructed by two grade six students. Educational Studies in Mathematics, 43(3):271–292, November 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1011976904850.

Sinclair:2003:SIR

[Sin03] M. P. Sinclair. Some implications of the results of a case study for the design of pre-constructed, dynamic geometry sketches and accompanying materials. Educational Studies in Mathematics, 52(3):289–317, April 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1024305603330.

Siu:1993:PPA

[Siu93] Man-Keung Siu. Proof and pedagogy in ancient China: Examples from Liu Hui’s commentary on Jiu Zhang Suan Shu. Educational Studies in Mathematics, 24(4):345–357, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273370.

Siwek:1973:LFR

[Siw73] Héléna Siwek. Logique formelle et raisonnement naturel des élèves dans l’enseignement de la mathématique. (French) [Formal logic and natural reasoning of pupils in the teaching of mathematics]. Educational Studies in Mathematics, 5(1):23– 37, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01425409.

Steele:2004:STV

[SJ04] Diana F. Steele and Debra I. Johanning. A schematic– theoretic view of problem solving and development of algebraic thinking. Educational Studies in Mathematics, 57 (1):65–90, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000047054.
90668.f9.

Sangwin:2017:ASA

[SJ17] Christopher J. Sangwin and Ian Jones. Asymmetry in student achievement on multiple-choice and constructed-response items in reversible mathematics processes. Educational Studies in Mathematics, 94(2):205–222, February 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-016-9725-4.pdf.

Shappeck:2012:BRU

[SJS12] Marco Shappeck, Gwendolyn Johnson, and Ali Shalaih. Book review: Understanding culture and social identity as the key to success in mathematics. Na’ilah Suad Nasir and Paul Cobb (Eds.) (2007). Improving access to mathematics: diversity and equity in the classroom. Educational Studies in Mathematics, 81(1):127–132, September 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9378-2.

Steiner:1969:CGO

[SK69] Hans-G. Steiner and Burt A. Kaufman. Checker games in operational systems as media for an inductive approach to teaching algebra. Educational Studies in Mathematics, 1(4): 445–483, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02396318.

Shaffer:1998:MVC

[SK98] David Williamson Shaffer and James J. Kaput. Mathematics and virtual culture: an evolutionary perspective on technology and mathematics education. Educational Studies in Mathematics, 37(2):97–119, November 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003590914788.

Stathopoulou:2007:LCM

[SK07] Charoula Stathopoulou and Fragiskos Kalabasis. Language and culture in mathematics education: Reflections on observing a Romany class in a Greek school. Educational Studies in Mathematics, 64(2):231–238, February 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-4496-y.

Solares:2013:ASN

[SK13] Armando Solares and Carolyn Kieran. Articulating syntactic and numeric perspectives on equivalence: the case of rational expressions. Educational Studies in Mathematics, 84(1):115– 148, September 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-013-9473-7.

Steflitsch:2025:SPC

[SK25] Daniela Steflitsch and David Kollosche. Students’ perceptions of Critical Mathematics Education: an exploration. Educational Studies in Mathematics, 120(2):225–247, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10423-y.

Solstad:2025:EMC

[SKE25] Trygve Solstad, Eivind Kaspersen, and Magnus Eggen. Eye movements in conceptual and non-conceptual thinking. Educational Studies in Mathematics, 118(3):323–338, March 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10358-w.

Schukajlow:2023:OMP

[SKK^{+}23] Stanislaw Schukajlow, Janina Krawitz, Jonas Kanefke, Werner Blum, and Katrin Rakoczy. Open modelling problems: cognitive barriers and instructional prompts. Educational Studies in Mathematics, 114(3):417–438, November 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10265-6.

Stovner:2021:ISW

[SKN21] Roar B. Stovner, Kirsti Klette, and Guri A. Nortvedt. The instructional situations in which mathematics teachers provide substantive feedback. Educational Studies in Mathematics, 108 (3):533–551, November 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10065-w.

Skovsmose:1985:MEV

[Sko85] Ole Skovsmose. Mathematical education versus critical education. Educational Studies in Mathematics, 16(4):337–354, November 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00417191.

Skovsmose:1988:MPT

[Sko88] Ole Skovsmose. Mathematics as part of technology. Educational Studies in Mathematics, 19(1):23–41, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00428383.

Skovsmose:1990:MED

[Sko90] Ole Skovsmose. Mathematical education and democracy. Educational Studies in Mathematics, 21(2):109–128, April 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304897.

Skovsmose:1994:TCM

[Sko94] Ole Skovsmose. Towards a critical mathematics education. Educational Studies in Mathematics, 27(1):35–57, July 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01284527.

Skott:2004:FAM

[Sko04] Jeppe Skott. The forced autonomy of mathematics teachers. Educational Studies in Mathematics, 55(1–3):227–257, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017670.35680.88.

Skovsmose:2012:SPR

[Sko12] Ole Skovsmose. Symbolic power, robotting, and surveilling. Educational Studies in Mathematics, 80(1–2):119–132, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9388-8.

Skovsmose:2016:IIM

[Sko16] Ole Skovsmose. An intentionality interpretation of meaning in mathematics education. Educational Studies in Mathematics, 92(3):411–424, July 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9644-9.

Skovsmose:2021:MC

[Sko21] Ole Skovsmose. Mathematics and crises. Educational Studies in Mathematics, 108(1–2):369–383, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10037-0.

Simon:2016:CPR

[SKPS16] Martin A. Simon, Melike Kara, Nicora Placa, and Hakan Sandir. Categorizing and promoting reversibility of mathematical concepts. Educational Studies in Mathematics, 93(2): 137–153, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9697-4.

Schukajlow:2015:EPM

[SKR15] Stanisław Schukajlow, André Krug, and Katrin Rakoczy. Effects of prompting multiple solutions for modelling problems on students’ performance. Educational Studies in Mathematics, 89(3):393–417, July 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9608-0.

Shumway:1974:NIA

[SL74] Richard J. Shumway and Frank K. Lester. Negative instances and the acquisition of the mathematical concepts of commutativity and associativity. Educational Studies in Mathematics, 5 (3):301–315, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01424552.

Sfard:1994:GPR

[SL94] Anna Sfard and Liora Linchevski. The gains and the pitfalls of reification — the case of algebra. Educational Studies in Mathematics, 26(2–3):191–228, March 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273663.

Saenz-Ludlow:1995:AFS

[SL95] Adalira Sáenz-Ludlow. Ann’s fraction schemes. Educational Studies in Mathematics, 28(2):101–132, March 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01295789.

Saenz-Ludlow:2006:CIG

[SL06] Adalira Sáenz-Ludlow. Classroom interpreting games with an illustration. Educational Studies in Mathematics, 61(1–2): 183–218, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-5760-x.

Schindler:2019:DSI

[SL19] Maike Schindler and Achim J. Lilienthal. Domain-specific interpretation of eye tracking data: towards a refined use of the eye-mind hypothesis for the field of geometry. Educational Studies in Mathematics, 101(1):123–139, May 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-9878-z;   http://link.springer.com/
content/pdf/10.1007/s10649-019-9878-z.pdf.

Slavit:1997:ARR

[Sla97] David Slavit. An alternate route to the reification of function. Educational Studies in Mathematics, 33(3):259–281, September 1997. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1002937032215.

Slavit:1998:ROS

[Sla98] David Slavit. The role of operation sense in transitions from arithmetic to algebraic thought. Educational Studies in Mathematics, 37(3):251–274, December 1998. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003602322232.

Soldano:2019:TBI

[SLAY19] Carlotta Soldano, Yael Luz, Ferdinando Arzarello, and Michal Yerushalmy. Technology-based inquiry in geometry: semantic games through the lens of variation. Educational Studies in Mathematics, 100(1):7–23, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9841-4.

Silver:1995:GMS

[SLC95] Edward A. Silver, Shukkwan S. Leung, and Jinfa Cai. Generating multiple solutions for a problem: A comparison of the responses of U.S. and Japanese students. Educational Studies in Mathematics, 28(1):35–54, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273855.

Slesnick:1982:ASV

[Sle82] Twila Slesnick. Algorithmic skill vs. conceptual understanding. Educational Studies in Mathematics, 13(2):143–154, May 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00460708.

Suppes:1969:PSC

[SLJ69] Patrick Suppes, Elizabeth F. Loftus, and Max Jerman. Problem-solving on a computer-based teletype. Educational Studies in Mathematics, 2(1):1–15, July 1969. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368982.

Sueur:1979:LID

[SLM79] M. Sueur, J.-P. Lamarche, and P. Marthe. Locutions inductrices et distractrices: ‘de plus que’, ‘de moins que’. (French) [Inductive and distracting dispositions: ‘more than’,‘less than’]. Educational Studies in Mathematics, 10(4):421–434, November 1979. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00417088.

Saenz–Ludlow:2006:GES

[SLP06] Adalira Sáenz-Ludlow and Norma Presmeg. Guest editorial semiotic perspectives on learning mathematics and communicating mathematically. Educational Studies in Mathematics, 61(1–2):1–10, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
9001-5.

Schukajlow:2012:TMM

[SLP^{+}12] Stanisław Schukajlow, Dominik Leiss, Reinhard Pekrun, Werner Blum, Marcel Müller, and Rudolf Messner. Teaching methods for modelling problems and students’ task-specific enjoyment, value, interest and self-efficacy expectations. Educational Studies in Mathematics, 79(2):215–237, February 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9341-2.

Safstrom:2024:DDF

[SLP^{+}24] Anna Ida Säfström, Johan Lithner, Torulf Palm, Björn Palmberg, Johan Sidenvall, Catarina Andersson, Erika Boström, and Carina Granberg. Developing a diagnostic framework for primary and secondary students’ reasoning difficulties during mathematical problem solving. Educational Studies in Mathematics, 115(2):125–149, February 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10278-1.

Stylianou:2024:SMG

[SLR^{+}24] Despina A. Stylianou, Boram Lee, Ingrid Ristroph, Eric Knuth, Maria Blanton, Ana Stephens, and Angela Gardiner. Semiotic mediation of gestures in the teaching of early algebra: the case of the equal sign. Educational Studies in Mathematics, 116(2):257–279, June 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10319-3.

Searle:1974:SVA

[SLS74] Barbara W. Searle, Paul Lorton, Jr., and Patrick Suppes. Structural variables affecting CaI performance on arithmetic word problems of disadvantaged and deaf students. Educational Studies in Mathematics, 5(3):371–384, March 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424555.

Svenson:1983:DUS

[SLW83] I. F. Svenson, J. A. Lawrence, and S. G. Willis. Distance university students’ processing of mathematics exercises. Educational Studies in Mathematics, 14(1):73–85, February 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00704703.

Saenz-Ludlow:1998:TGI

[SLW98] Adalira Sáenz-Ludlow and Catherine Walgamuth. Third graders’ interpretations of equality and the equal symbol. Educational Studies in Mathematics, 35(2):153–187, February 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003086304201.

Scandura:1972:EIB

[SM72] Joseph M. Scandura and Robert McGee. An exploratory investigation of basic mathematical abilities of kindergarten children. Educational Studies in Mathematics, 4(3):331–345, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302582.

Schrage:1983:BR

[SM83] G. Schrage and Curtis McKnight. Book reviews. Educational Studies in Mathematics, 14(4):431–433, November 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368238.

Schuler-Meyer:2017:SDS

[SM17] Alexander Schüler-Meyer. Students’ development of structure sense for the distributive law. Educational Studies in Mathematics, 96(1):17–32, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:/

/link.springer.com/article/10.1007/s10649-017-9765-
4.

Sjoblom:2021:PGL

[SM21] Marie Sjöblom and Tamsin Meaney. “I am part of the group, the others listen to me”: theorising productive listening in mathematical group work. Educational Studies in Mathematics, 107(3):565–581, July 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10051-
2.

Schuler-Meyer:2022:HTS

[SM22a] Alexander Schüler-Meyer. How transition students relearn school mathematics to construct multiply quantified statements. Educational Studies in Mathematics, 110(2):291–311, June 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10127-z.

Sharpe:2022:SRF

[SM22b] Sheree T. Sharpe and Dalton D. Marsh. A systematic review of factors associated with high schoolers’ algebra achievement according to HSLS:09 results. Educational Studies in Mathematics, 110(3):457–480, July 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10130-4.

Sanchez-Matamoros:2019:RAP

[SMFL19] Gloria Sánchez-Matamoros, Ceneida Fernández, and Salvador Llinares. Relationships among prospective secondary mathematics teachers’ skills of attending, interpreting and responding to students’ understanding. Educational Studies in Mathematics, 100(1):83–99, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9855-y.

Smith:1970:CCE

[Smi70] David A. Smith. A calculus-with-computer experiment. Educational Studies in Mathematics, 3(1):1–11, September 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00381591.

Strohmaier:2020:ETM

[SMOR20] Anselm R. Strohmaier, Kelsey J. MacKay, Andreas Obersteiner, and Kristina M. Reiss. Eye-tracking methodology in mathematics education research: A systematic literature review. Educational Studies in Mathematics, 104(2):147–200, June 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09948-1;   http://link.
springer.com/content/pdf/10.1007/s10649-020-09948-
1.pdf.

Street:2022:CSS

[SMS22] Karin E. S. Street, Lars-Erik Malmberg, and Gabriel J. Stylianides. Changes in students’ self-efficacy when learning a new topic in mathematics: a micro-longitudinal study. Educational Studies in Mathematics, 111(3):515–541, November 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-022-10165-1.

Sandefur:2013:EGU

[SMSW13a] J. Sandefur, J. Mason, G. J. Stylianides, and A. Watson. Erratum to: Generating and using examples in the proving process. Educational Studies in Mathematics, 84(3):487, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10649-013-9510-6.pdf. See [SMSW13b].

Sandefur:2013:GUE

[SMSW13b] J. Sandefur, J. Mason, G. J. Stylianides, and A. Watson. Generating and using examples in the proving process. Educational Studies in Mathematics, 83(3):323–340, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9459-x. See [SMSW13a].

Sinclair:2011:DSV

[SMW11] Margaret Sinclair, Ami Mamolo, and Walter J. Whiteley. Designing spatial visual tasks for research: the case of the filling task. Educational Studies in Mathematics, 78(2):135–163, November 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9315-4.

Siregar:2025:BRC

[SMWM25] Hesty Marwani Siregar, Talisadika Serrisanti Maifa, Ambarsari Kusuma Wardani, and Sitti Busyrah Muchsin. Book review: Catherine P. Vistro–Yu & Tin Lam Toh (eds.) (2019) School mathematics curricula – Asian perspectives and glimpses of reform. Educational Studies in Mathematics, 118 (2):295–305, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10338-0.

Shapiro:1973:QCL

[SO73] Bernard J. Shapiro and Thomas C. O’Brien. Quasi-child logics. Educational Studies in Mathematics, 5(2):181–184, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01421024.

Span:1986:IPI

[SOC86] Pieter Span and Ruth Overtoom-Corsmit. Information processing by intellectually gifted pupils solving mathematical problems. Educational Studies in Mathematics, 17(3):273–295, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305074.

Solar:1995:IPM

[Sol95] Claudie Solar. An inclusive pedagogy in mathematics education. Educational Studies in Mathematics, 28(3):311–333, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01274179.

Solomon:2006:DDR

[Sol06] Yvette Solomon. Deficit or difference? The role of students’ epistemologies of mathematics in their interactions with proof.

Educational Studies in Mathematics, 61(3):373–393, March 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-6927-1.

Solomon:2012:FVN

[Sol12] Yvette Solomon. Finding a voice? Narrating the female self in mathematics. Educational Studies in Mathematics, 80(1– 2):171–183, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9384-z.

Son:2013:HPT

[Son13] Ji-Won Son. How preservice teachers interpret and respond to student errors: ratio and proportion in similar rectangles. Educational Studies in Mathematics, 84(1):49–70, September 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9475-5.

Son:2016:MBT

[Son16] Ji-Won Son. Moving beyond a traditional algorithm in whole number subtraction: Preservice teachers’ responses to a student’s invented strategy. Educational Studies in Mathematics, 93(1):105–129, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9693-8.

Souviney:1983:MAL

[Sou83] Randall J. Souviney. Mathematics achievement, language and cognitive development: Classroom practices in Papua New Guinea. Educational Studies in Mathematics, 14(2):183–212, May 1983. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00303685.

Sandoval:2016:ADR

[SP16] Ivonne Sandoval and Edgar Possani. An analysis of different representations for vectors and planes in R^{3}. Educational Studies in Mathematics, 92(1):109–127, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9675-2.

Scheiner:2019:EPM

[SP19] Thorsten Scheiner and Márcia M. F. Pinto. Emerging perspectives in mathematical cognition: contextualizing, complementizing, and complexifying. Educational Studies in Mathematics, 101(3):357–372, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-019-9879-y.

Samper:2016:DUE

[SPC^{+}16] Carmen Samper, Patricia Perry, Leonor Camargo, Adalira Sáenz-Ludlow, and Óscar Molina. A dilemma that underlies an existence proof in geometry. Educational Studies in Mathematics, 93(1):35–50, September 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-016-
9683-x.

Speer:2005:IMT

[Spe05] Natasha M. Speer. Issues of methods and theory in the study of mathematics teachers’ professed and attributed beliefs. Educational Studies in Mathematics, 58(3):361–391, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-2745-0.

Straehler-Pohl:2014:SMR

[SPFGF14] Hauke Straehler-Pohl, Saínza Fernández, Uwe Gellert, and Lourdes Figueiras. School mathematics registers in a context of low academic expectations. Educational Studies in Mathematics, 85(2):175–199, February 2014. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9503-5.

Selter:2012:TDD

[SPNH12] Christoph Selter, Susanne Prediger, Marcus Nührenbörger, and Stephan Hußmann. Taking away and determining the difference — a longitudinal perspective on two models of subtraction and the inverse relation to addition. Educational Studies in Mathematics, 79(3):389–408, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9305-6.

Stouraitis:2017:CDO

[SPS17] Konstantinos Stouraitis, Despina Potari, and Jeppe Skott. Contradictions, dialectical oppositions and shifts in teaching mathematics. Educational Studies in Mathematics, 95(2):203– 217, June 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Singer:1992:RPR

[SR92] Janice Ann Singer and Lauren B. Resnick. Representations of proportional relationships: Are children part–part or part– whole reasoners? Educational Studies in Mathematics, 23(3): 231–246, June 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02309531.

Solomon:2016:CAV

[SRB16] Yvette Solomon, Darinka Radovic, and Laura Black. “I can actually be very feminine here”: contradiction and hybridity in becoming a female mathematician. Educational Studies in Mathematics, 91(1):55–71, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9649-4.

Sriraman:2018:BRA

[Sri18] Bharath Sriraman. Book review: Anthropomorphizing mathematics education. Review of Hauke Staehler–Pohl, Nina Bohlmann, and Alexandre Pais (Eds). (2017). The disorder of mathematics education: challenging the sociopolitical dimensions of research. 329 pp. Educational Studies in Mathematics, 97(2):209–213, February 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-017-9786-z.

Stephan:2021:PUM

[SRP21] Michelle Stephan, Jordan Register, and David Pugalee. People use math as a weapon: critical mathematics consciousness in the time of COVID-19. Educational Studies in Mathematics, 108(3):513–532, November 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10062-z.

Strohmaier:2022:DCW

[SRR22] Anselm R. Strohmaier, Frank Reinhold, and Kristina Reiss. Different complex word problems require different combinations of cognitive skills. Educational Studies in Mathematics, 109(1):89–114, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10079-4.

Swadener:1988:VME

[SS88] Marc Swadener and R. Soedjadi. Values, mathematics education, and the task of developing pupils’ personalities: an Indonesian perspective. Educational Studies in Mathematics, 19 (2):193–208, May 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00751232.

Simon:1991:TCP

[SS91] Martin A. Simon and Deborah Schifter. Towards a constructivist perspective: An intervention study of mathematics teacher development. Educational Studies in Mathematics, 22(4):309–331, August 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00369293.

Simon:1993:TCP

[SS93a] Martin A. Simon and Deborah Schifter. Toward a constructivist perspective: The impact of a mathematics teacher inservice program on students. Educational Studies in Mathematics, 25(4):331–340, December 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273905.

Sinclair:1993:UWN

[SS93b] Anne Sinclair and Nora Scheuer. Understanding the written number system: 6 year-olds in Argentina and Switzerland. Educational Studies in Mathematics, 24(2):199–221, February 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273692.

Selden:1995:ULM

[SS95] John Selden and Annie Selden. Unpacking the logic of mathematical statements. Educational Studies in Mathematics, 29(2):123–151, September 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274210.

Sinclair:2003:CRW

[SS03] N. Sinclair and M. Schiralli. A constructive response to ‘where mathematics comes from’. Educational Studies in Mathematics, 52(1):79–91, January 2003. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023673520853.

Sadovsky:2005:AIP

[SS05a] Patricia Sadovsky and Carmen Sessa. The adidactic interaction with the procedures of peers in the transition from arithmetic to algebra: A milieu for the emergence of new questions. Educational Studies in Mathematics, 59(1–3):85– 112, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-5886-2.

Sierpinska:2005:TIA

[SS05b] Anna Sierpinska and Marie Sheldon. Technical instructions for authors. Educational Studies in Mathematics, 58(1):137–143, January 2005. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-005-7273-4.

Simpson:2006:AMS

[SS06] Adrian Simpson and Nada Stehlíková. Apprehending mathematical structure: A case study of coming to understand a commutative ring. Educational Studies in Mathematics, 61 (3):347–371, March 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-006-1300-y.

Stylianides:2009:PCE

[SS09] Andreas J. Stylianides and Gabriel J. Stylianides. Proof constructions and evaluations. Educational Studies in Mathematics, 72(2):237–253, November 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-009-
9191-3.

Son:2010:HRC

[SS10] Ji-Won Son and Sharon L. Senk. How reform curricula in the USA and Korea present multiplication and division of fractions. Educational Studies in Mathematics, 74(2):117–142, June 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9229-6.

Santos:2015:DMW

[SS15] Leonor Santos and Sílvia Semana. Developing mathematics written communication through expository writing supported by assessment strategies. Educational Studies in Mathematics, 88(1):65–87, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-014-
9557-z.

Stylianides:2017:RBI

[SS17] Gabriel J. Stylianides and Andreas J. Stylianides. Researchbased interventions in the area of proof: the past, the present, and the future. Educational Studies in Mathematics, 96(2):119–127, October 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10649-017-
9782-3.pdf;   http://link.springer.com/article/10.
1007/s10649-017-9782-3.

Swidan:2020:DMR

[SSA20] Osama Swidan, Cristina Sabena, and Ferdinando Arzarello. Disclosure of mathematical relationships with a digital tool: a three layer-model of meaning. Educational Studies in Mathematics, 103(1):83–101, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09926-
2.

Schindler:2025:IET

[SSL25] Maike Schindler, Anna Shvarts, and Achim J. Lilienthal. Introduction to eye tracking in mathematics education: interpretation, potential, and challenges. Educational Studies in Mathematics, 118(3):309–321, March 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10393-1.

Sensevy:2005:AMT

[SSLM^{+}05] Gérard Sensevy, Maria-Luisa Schubauer-Leoni, Alain Mercier, Florence Ligozat, and Gérard Perrot. An attempt to model the teacher’s action in the mathematics class. Educational Studies in Mathematics, 59(1–3):153–181, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-5887-1.

Stylianides:2004:USU

[SSP04] Andreas J. Stylianides, Gabriel J. Stylianides, and George N. Philippou. Undergraduate students’ understanding of the contraposition equivalence rule in symbolic and verbal contexts. Educational Studies in Mathematics, 55(1–3):133–162, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000017671.47700.0b.

Scheibling-Seve:2020:ACK

[SSPS20] Calliste Scheibling-Sève, Elena Pasquinelli, and Emmanuel Sander. Assessing conceptual knowledge through solving arithmetic word problems. Educational Studies in Mathematics, 103 (3):293–311, March 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-020-09938-3.

Saldanha:2002:CST

[ST02] L. Saldanha and P. Thompson. Conceptions of sample and their relationship to statistical inference. Educational Studies in Mathematics, 51(3):257–270, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023692604014.

Sdrolias:2008:TSS

[ST08] Konstantinos A. Sdrolias and Triandafillos A. Triandafillidis. The transition to secondary school geometry: can there be a “chain of school mathematics”? Educational Studies in Mathematics, 67(2):159–169, February 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9093-1.

Sinclair:2010:DSM

[ST10] Nathalie Sinclair and Shiva Gol Tabaghi. Drawing space: mathematicians’ kinetic conceptions of eigenvectors. Educational Studies in Mathematics, 74(3):223–240, July 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9235-8.

Shahbari:2020:FMP

[ST20] Juhaina Awawdeh Shahbari and Michal Tabach. Features of modeling processes that elicit mathematical models represented at different semiotic registers. Educational Studies in Mathematics, 105(2):115–135, October 2020. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09971-2.

Stacey:1989:FUP

[Sta89] Kaye Stacey. Finding and using patterns in linear generalising problems. Educational Studies in Mathematics, 20(2): 147–164, May 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00579460.

Stacey:2011:BRJ

[Sta11] Kaye Stacey. Book review: A journey through the world of technology in mathematics education. Celia Hoyles and Jean-Baptiste Lagrange (Eds.) (2010). Mathematics education and technology — Rethinking the terrain. Educational Studies in Mathematics, 78(1):127–134, September 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9316-3.

Steiner:1968:EEM

[Ste68] Hans-Georg Steiner. Examples of exercises in mathematization on the secondary school level. Educational Studies in Mathematics, 1(1–2):181–201, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00426243.

Steiner:1969:EEM

[Ste69a] Hans-Georg Steiner. Examples of exercises in mathematization. Educational Studies in Mathematics, 1(3):289–299, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00558314.

Steiner:1969:MRN

[Ste69b] Hans-Georg Steiner. Magnitudes and rational numbers — a didactical analysis. Educational Studies in Mathematics, 2(2–3): 371–392, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303470.

Steiner:1971:FEG

[Ste71] Hans-Georg Steiner. A foundation of Euclidean geometry by means of congruence mappings. Educational Studies in Mathematics, 4(1):87–90, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00305799.

Steiner:1976:FGN

[Ste76] Hans Georg Steiner. Finite geometries and non-measurable voting bodies. Educational Studies in Mathematics, 7(1–2): 139–146, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00144367.

Steffe:1983:CAS

[Ste83] Leslie P. Steffe. Children’s algorithms as schemes. Educational Studies in Mathematics, 14(2):109–125, May 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00303681.

Steinbring:1991:CCE

[Ste91] Heinz Steinbring. The concept of chance in everyday teaching: Aspects of a social epistemology of mathematical knowledge. Educational Studies in Mathematics, 22(6):503–522, December 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312713.

Steiner:1994:EW

[Ste94] Hans-Georg Steiner. Elfriede Wenzelburger 1947–1993. Educational Studies in Mathematics, 26(1):103, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273303.

Steinbring:1997:EIC

[Ste97] Heinz Steinbring. Epistemological investigation of classroom interaction in elementary mathematics teaching. Educational Studies in Mathematics, 32(1):49–92, January 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002919830949.

Stevenson:1998:RCE

[Ste98] Ian Stevenson. Radical constructivism. Ernst von Glasersfeld. Educational Studies in Mathematics, 35(1):93–104, January 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003018912036.

Steinbring:2005:AMT

[Ste05] Heinz Steinbring. Analyzing mathematical teaching–learning situations — the interplay of communicational and epistemological constraints. Educational Studies in Mathematics, 59(1– 3):313–324, July 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-4819-4.

Steinbring:2006:WMS

[Ste06] Heinz Steinbring. What makes a sign a mathematical sign? — an epistemological perspective on mathematical interaction. Educational Studies in Mathematics, 61(1–2):133–162, February 2006. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-5892-z.

Stein:2013:RRH

[Ste13] Robert George Stein. Review of Rewriting the history of school mathematics in North America 1607–1861 by Nerida Ellerton and M. A. (Ken) Clements (2012). The central role of cyphering books. Educational Studies in Mathematics, 82(1): 165–167, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9418-6.

Sullivan:2006:PDS

[STM06] Peter Sullivan, Steve Tobias, and Andrea McDonough. Perhaps the decision of some students not to engage in learning mathematics in school is deliberate. Educational Studies in Mathematics, 62(1):81–99, May 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-1348-8.

Stone:1971:LTA

[Sto71] Marshall Stone. Learning and teaching axiomatic geometry. Educational Studies in Mathematics, 4(1):91–103, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305800.

Stohlmann:2022:BRG

[Sto22] Micah Stohlmann. Book review: Growing the research base for mathematical modeling with young learners. Jennifer Suh, Megan Wickstrom, & Lyn English (Eds.) (2021) Exploring mathematical modeling with young learners. Educational Studies in Mathematics, 109(1):195–201, January 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10091-8.

Streefland:1978:SOR

[Str78] Leen Streefland. Some observational results concerning the mental constitution of the concept of fraction. Educational Studies in Mathematics, 9(1):51–73, February 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00352192.

Streefland:1982:SFD

[Str82] Leen Streefland. Subtracting fractions with different denominators. Educational Studies in Mathematics, 13(3):233–255, August 1982. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00311243.

Streefland:1984:SRR

[Str84] Leen Streefland. Search for the roots of ratio: Some thoughts on the long term learning process (towards … a theory). Educational Studies in Mathematics, 15(4):327–348, November 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311111.

Streefland:1985:SRR

[Str85] Leen Streefland. Search for the roots of ratio: Some thoughts on the long term learning process (towards … a theory). Educational Studies in Mathematics, 16(1):75–94, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00354884.

Straker:1987:VTM

[Str87] Neil Straker. Videodisc technology and mathematics education. Educational Studies in Mathematics, 18(4):445–450, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00240990.

Streefland:1993:DMC

[Str93a] L. Streefland. The design of a mathematics course. A theoretical reflection. Educational Studies in Mathematics, 25(1–2): 109–135, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274105.

Streefland:1993:ELF

[Str93b] Leen Streefland. Editorial: The legacy of Freudenthal. Educational Studies in Mathematics, 25(1–2):viii–7, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274098.

Streefland:2003:LHT

[Str03] L. Streefland. Learning from history for teaching in the future. Educational Studies in Mathematics, 54(1):37–62, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000005236.36499.fc.

Strasser:2008:RPF

[Str08] Rudolf Sträßer. Review of the proceedings of the 2001, 2003 and 2005 French summer schools in didactics of mathematics. Educational Studies in Mathematics, 69(3):277–281, November 2008. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-008-9152-2.

Stylianides:2007:NPC

[Sty07] Andreas J. Stylianides. The notion of proof in the context of elementary school mathematics. Educational Studies in Mathematics, 65(1):1–20, May 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-006-
9038-0.

Stylianou:2011:EMS

[Sty11] Despina A. Stylianou. An examination of middle school students’ representation practices in mathematical problem solving through the lens of expert work: towards an organizing scheme. Educational Studies in Mathematics, 76(3):265–280, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9273-2.

Sun:2011:VPT

[Sun11] Xuhua Sun. “Variation problems” and their roles in the topic of fraction division in Chinese mathematics textbook examples. Educational Studies in Mathematics, 76(1):65–85, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9263-4.

Surany:1977:LTC

[Sur77] János Surány. “I have lost the thread completely … ”. Educational Studies in Mathematics, 8(1):35–39, April 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302503.

Sutherland:1989:PCB

[Sut89] Rosamund Sutherland. Providing a computer based framework for algebraic thinking. Educational Studies in Mathematics, 20(3):317–344, August 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00310876.

Sutherland:1993:CTP

[Sut93] Rosamund Sutherland. Connecting theory and practice: Results from the teaching of logo. Educational Studies in Mathematics, 24(1):95–113, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273296.

Stacey:2009:MRE

[SV09] Kaye Stacey and Jill Vincent. Modes of reasoning in explanations in Australian eighth-grade mathematics textbooks. Educational Studies in Mathematics, 72(3):271–288, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9193-1.

Singer:2013:PSC

[SV13] Florence Mihaela Singer and Cristian Voica. A problemsolving conceptual framework and its implications in designing problem-posing tasks. Educational Studies in Mathematics, 83(1):9–26, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9422-x.

Schreiter:2025:SLV

[SV25] Saskia Schreiter and Markus Vogel. Students’ local vs. global views of data distributions: a cross-grade-level analysis using eye-tracking. Educational Studies in Mathematics, 119(1):1– 22, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10352-2.

Supply:2020:AYC

[SVLO20] Anne-Sophie Supply, Wim Van Dooren, Stephanie Lem, and Patrick Onghena. Assessing young children’s ability to compare probabilities. Educational Studies in Mathematics, 103 (1):27–42, January 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-019-09917-3.

Supply:2023:PCC

[SVVO23] Anne-Sophie Supply, Elien Vanluydt, Wim Van Dooren, and Patrick Onghena. Out of proportion or out of context? Comparing 8- to 9-year-olds’ proportional reasoning abilities across fair-sharing, mixtures, and probability contexts. Educational Studies in Mathematics, 113(3):371–388, July 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10212-5.

Steffe:1992:RPM

[SW92] Leslie P. Steffe and Heide G. Wiegel. On reforming practice in mathematics education. Educational Studies in Mathematics, 23(5):445–465, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00571467.

Steffe:1994:CPM

[SW94] Leslie P. Steffe and Heide G. Wiegel. Cognitive play and mathematical learning in computer microworlds. Educational Studies in Mathematics, 26(2–3):111–134, March 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273660.

Smidt:1995:SSO

[SW95] Siegbert Smidt and Werner Weiser. Semantic structures of one-step word problems involving multiplication or division. Educational Studies in Mathematics, 28(1):55–72, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273856.

Swanson:2014:MAM

[SW14] David Swanson and Julian Williams. Making abstract mathematics concrete in and out of school. Educational Studies in Mathematics, 86(2):193–209, June 2014. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-014-9536-4.pdf.

Swetz:1977:ODP

[Swe77] Frank Swetz. The ‘open-door’ policy and the reform of mathematics education in the People’s Republic of China. Educational Studies in Mathematics, 8(4):461–478, December 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00310948.

Swetz:1995:KTM

[Swe95] Frank Swetz. To know and to teach: Mathematical pedagogy from a historical context. Educational Studies in Mathematics, 29(1):73–88, July 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273901.

Speiser:2005:GMS

[SWG05] Bob Speiser, Chuck Walter, and Tiffini Glaze. Getting at the mathematics: Sara’s journal. Educational Studies in Mathematics, 58(2):189–207, February 2005. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-1525-1.

Supply:2023:IPP

[SWVO23] Anne-Sophie Supply, Nore Wijns, Wim Van Dooren, and Patrick Onghena. It is probably a pattern: does spontaneous focusing on regularities in preschool predict reasoning about randomness four years later? Educational Studies in Mathematics, 112(1):3–24, January 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10187-9.

Swetz:1979:MOP

[SY79] Frank Swetz and Ying-King Yu. Mathematical olympiads in the People’s Republic of China. Educational Studies in Mathematics, 10(4):435–442, November 1979. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417089.

Sirotic:2007:ING

[SZ07] Natasa Sirotic and Andrina Zazkis. Irrational numbers: The gap between formal and intuitive knowledge. Educational Studies in Mathematics, 65(1):49–76, May 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9041-5.

Shreyar:2010:TAT

[SZP10] Sam Shreyar, Betina Zolkower, and Silvia Pérez. Thinking aloud together: A teacher’s semiotic mediation of a whole-class conversation about percents. Educational Studies in Mathematics, 73(1):21–53, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-009-
9203-3.

Tarim:2008:ECL

[TA08] Kamuran Tarim and Fikri Akdeniz. The effects of cooperative learning on Turkish elementary students’ mathematics achievement and attitude towards mathematics using TAI and STAD methods. Educational Studies in Mathematics, 67(1): 77–91, January 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9088-y.

Tabach:2013:CWG

[Tab13] Michal Tabach. Commentary: The whole is greater than the sum of its parts. Educational Studies in Mathematics, 84(2): 273–277, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10649-013-9508-0.pdf.

Tabach:2022:BRF

[Tab22] Michal Tabach. Book review: A fruitful resource emerging from collaboration within an ERME topic conference. Alison Clark-Wilson, Ana Donevska-Todorova, Eleonora Faggiano, Jana Trgalová and Hans-Georg Weigand (Eds.) (2021) Mathematics education in the digital age: learning, practice and theory. Educational Studies in Mathematics, 110(3):573– 580, July 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10126-0.

Tabach:2008:TAD

[TAH08] Michal Tabach, Abraham Arcavi, and Rina Hershkowitz. Transitions among different symbolic generalizations by algebra beginners in a computer intensive environment. Educational Studies in Mathematics, 69(1):53–71, September 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9125-5.

Takeuchi:2018:PII

[Tak18] Miwa Aoki Takeuchi. Power and identity in immigrant parents’ involvement in early years mathematics learning. Educational Studies in Mathematics, 97(1):39–53, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9781-4.

Tall:1980:NIM

[Tal80] David Tall. The notion of infinite measuring number and its relevance in the intuition of infinity. Educational Studies in Mathematics, 11(3):271–284, August 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00697740.

Tall:2001:NFI

[Tal01] David Tall. Natural and formal infinities. Educational Studies in Mathematics, 48(2–3):199–238, November 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016000710038.

Tall:2008:JJK

[Tal08] David Tall. James J. Kaput (1942–2005) imagineer and futurologist of mathematics education. Educational Studies in Mathematics, 68(2):185–193, June 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9121-9.

Tao:1986:PIG

[Tao86] Billy Tao. Parental involvement in gifted education. Educational Studies in Mathematics, 17(3):313–321, August 1986.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305076.

Tarim:2009:ECL

[Tar09] Kamuran Tarim. The effects of cooperative learning on preschoolers’ mathematics problem-solving ability. Educational Studies in Mathematics, 72(3):325–340, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9197-x.

Tavernier:1968:TPA

[Tav68] Jean Tavernier. The thinking of a physicist about mathematics. Educational Studies in Mathematics, 1(1–2):55–60, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426231.

Taylor:1991:IIA

[Tay91] Nick Taylor. Independence and interdependence: Analytical vectors for defining the mathematics curriculum of schools in a democratic society. Educational Studies in Mathematics, 22 (2):107–123, April 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00555719.

Taylor:1996:MMM

[Tay96] Peter Charles Taylor. Mythmaking and mythbreaking in the mathematics classroom. Educational Studies in Mathematics, 31(1–2):151–173, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00143930.

Traore:2009:MVQ

[TB09] Kalifa Traoré and Nadine Bednarz. Mathématiques de la vie quotidienne au Burkina Faso: une analyse de la pratique sociale de comptage et de vente de mangues. (French) [Mathematics of everyday life in Burkina Faso: an analysis of the social practice of counting and selling mangoes]. Educational Studies in Mathematics, 72(3):359–378, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9200-6.

Thurm:2022:TMT

[TB22] Daniel Thurm and Bärbel Barzel. Teaching mathematics with technology: a multidimensional analysis of teacher beliefs. Educational Studies in Mathematics, 109(1):41–63, January 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-021-10072-x.

Tchoshanov:2011:RBT

[Tch11] Mourat A. Tchoshanov. Relationship between teacher knowledge of concepts and connections, teaching practice, and student achievement in middle grades mathematics. Educational Studies in Mathematics, 76(2):141–164, March 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9269-y.

Towers:2002:SO

[TD02] Jo Towers and Brent Davis. Structuring occasions. Educational Studies in Mathematics, 49(3):313–340, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020245320040.

Temple:2012:DFM

[TD12] Codruta Temple and Helen M. Doerr. Developing fluency in the mathematical register through conversation in a tenthgrade classroom. Educational Studies in Mathematics, 81(3): 287–306, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9398-6.

Turner:2013:LBT

[TDEM13] Erin Elizabeth Turner, Higinio Dominguez, Susan Empson, and Luz Angelica Maldonado. Latino/a bilinguals and their teachers developing a shared communicative space. Educational Studies in Mathematics, 84(3):349–370, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9486-2.

Torbeyns:2009:AUS

[TDGV09] Joke Torbeyns, Bert De Smedt, Pol Ghesquière, and Lieven Verschaffel. Acquisition and use of shortcut strategies by traditionally schooled children. Educational Studies in Mathematics, 71(1):1–17, May 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-008-
9155-z.

Tirosh:1997:DDC

[TE97] Dina Tirosh and Ruhama Even. To define or not to define: The case of (−8)^{1/3}. Educational Studies in Mathematics, 33(3):321–330, September 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002916606955.

Tsatsaroni:2014:ANT

[TE14] Anna Tsatsaroni and Jeff Evans. Adult numeracy and the totally pedagogised society: PIAAC and other international surveys in the context of global educational policy on lifelong learning. Educational Studies in Mathematics, 87(2):167–186, October 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9470-x.

Teppo:2020:BRI

[Tep20] Anne R. Teppo. Book review: International and national reflections on The Netherlands didactics of mathematics education. Educational Studies in Mathematics, 104(2):285–298, June 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-020-09960-5.

Tirosh:1998:SAE

[TER98] Dina Tirosh, Ruhama Even, and Naomi Robinson. Simplifying algebraic expressions: Teacher awareness and teaching approaches. Educational Studies in Mathematics, 35(1):51–64, January 1998. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1003011913153.

Tartre:1995:MAG

[TF95] Lindsay Anne Tartre and Elizabeth Fennema. Mathematics achievement and gender: A longitudinal study of selected cognitive and affective variables [grades 6–12]. Educational Studies in Mathematics, 28(3):199–217, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274173.

Tirosh:1989:PET

[TG89] Dina Tirosh and Anna O. Graeber. Preservice elementary teachers’ explicit beliefs about multiplication and division. Educational Studies in Mathematics, 20(1):79–96, February 1989. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00356042.

Tirpakova:2023:DCT

[TGWN23] Anna Tirpáková, Dalibor Gonda, Adriana Wiegerová, and Hana Navrátilová. Developing the concept of task substitution and transformation by defining own equivalences. Educational Studies in Mathematics, 114(3):483–502, November 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10264-7.

Ticha:2013:DTS

[TH13] Marie Tichá and Alena Hošpesová. Developing teachers’ subject didactic competence through problem posing. Educational Studies in Mathematics, 83(1):133–143, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9455-1.

Thanheiser:2010:IFP

[Tha10] Eva Thanheiser. Investigating further preservice teachers’ conceptions of multidigit whole numbers: refining a framework. Educational Studies in Mathematics, 75(3):241–251, December 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9252-7.

Thompson:1984:RTC

[Tho84] Alba Gonzalez Thompson. The relationship of teachers’ conceptions of mathematics and mathematics teaching to instructional practice. Educational Studies in Mathematics, 15(2): 105–127, May 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305892.

Thornton:1985:BBN

[Tho85] E. B. C. Thornton. Between behaviour and neurology. Educational Studies in Mathematics, 16(1):98–101, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00354886.

Thornton:1990:SSS

[Tho90] Carol A. Thornton. Solution strategies: Subtraction number facts. Educational Studies in Mathematics, 21(3):241– 263, June 1990. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00305092.

Thompson:1993:QRC

[Tho93] Patrick W. Thompson. Quantitative reasoning, complexity, and additive structures. Educational Studies in Mathematics, 25(3):165–208, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273861.

Thompson:1994:YCI

[Tho94a] Ian Thompson. Young children’s idiosyncratic written algorithms for addition. Educational Studies in Mathematics, 26 (4):323–345, June 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01279519.

Thompson:1994:IRO

[Tho94b] Patrick W. Thompson. Images of rate and operational understanding of the fundamental theorem of calculus. Educational Studies in Mathematics, 26(2–3):229–274, March 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273664.

Threlfall:2002:FMC

[Thr02] John Threlfall. Flexible mental calculation. Educational Studies in Mathematics, 50(1):29–47, May 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020572803437.

Tabach:2006:CCA

[THS06] Michal Tabach, Rina Hershkowitz, and Baruch Schwarz. Constructing and consolidating of algebraic knowledge within dyadic processes: A case study. Educational Studies in Mathematics, 63(3):235–258, November 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9012-2.

Thurston:1972:WE

[Thu72] Hugh Thurston. What exactly is dy/dx? Educational Studies in Mathematics, 4(3):358–367, April 1972. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302584.

Thwaites:1969:RCS

[Thw69] Bryan Thwaites. The role of the computer in school mathematics. Educational Studies in Mathematics, 2(2–3):346–359, December 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00303468.

Ticknor:2012:SLA

[Tic12] Cindy S. Ticknor. Situated learning in an abstract algebra classroom. Educational Studies in Mathematics, 81(3):307– 323, November 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9405-y.

Tiedemann:2000:GRB

[Tie00]  Joachim Tiedemann.  Gender-related beliefs of teach-

ers in elementary school mathematics. Educational Studies in Mathematics, 41(2):191–207, February 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003953801526.

Tiedemann:2002:TGS

[Tie02] Joachim Tiedemann. Teachers’ gender stereotypes as determinants of teacher perceptions in elementary school mathematics. Educational Studies in Mathematics, 50(1):49–62, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020518104346.

Tirosh:1999:FMK

[Tir99] Dina Tirosh. Forms of mathematical knowledge: Learning and teaching with understanding. Educational Studies in Mathematics, 38(1–3):1–9, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017237429271.

Tanner:1994:UPS

[TJ94] Howard Tanner and Sonia Jones. Using peer and selfassessment to develop modelling skills with students aged 11 to 16: A socio-constructive view. Educational Studies in Mathematics, 27(4):413–431, December 1994. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273381.

Tsay:2011:CSC

[TJHD11] Jenq-Jong Tsay, April B. Judd, Shandy Hauk, and Mark K. Davis. Case study of a college mathematics instructor: patterns of classroom discourse. Educational Studies in Mathematics, 78(2):205–229, November 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9323-4.

Tall:2014:CAC

[TK14] David Tall and Mikhail Katz. A cognitive analysis of Cauchy’s conceptions of function, continuity, limit and infinitesimal, with implications for teaching the calculus. Educational Studies in Mathematics, 86(1):97–124, May 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9531-9.

Tainio:2015:EWA

[TL15] Liisa Tainio and Anu Laine. Emotion work and affective stance in the mathematics classroom: the case of IRE sequences in Finnish classroom interaction. Educational Studies in Mathematics, 89(1):67–87, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9591-5.

Thurm:2024:PDT

[TLB^{+}24] Daniel Thurm, Shuhui Li, Bärbel Barzel, Lianghuo Fan, and Na Li. Professional development for teaching mathematics with technology: a comparative study of facilitators’ beliefs and practices in China and Germany. Educational Studies in Mathematics, 115(2):247–269, February 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10284-3.

Traore:2007:QEP

[TLM07] Kalifa Traoré, Caroline Lajoie, and Roberta Mura. Quelques erreurs pouvant être liées à une difficulté à concevoir un ensemble comme un objet distinct de ses éléments chez des étudiants et des étudiantes universitaires. (French) [Some errors that can be related to a difficulty in conceiving a set as an object distinct from its elements in university students]. Educational Studies in Mathematics, 64(3):247–264, March 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9018-9.

Thanheiser:2021:NFM

[TMH21] Eva Thanheiser, Kathleen Melhuish, and Ruth Heaton. Networking frameworks: a method for analyzing the complexities of classroom cultures focusing on justifying. Educational Studies in Mathematics, 107(2):285–314, June 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10026-3.

Trigueros:2010:GRL

[TMP10] Maria Trigueros and Rafael Martínez-Planell. Geometrical representations in the learning of two-variable functions. Educational Studies in Mathematics, 73(1):3–19, January 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9201-5.

Tabach:2015:CET

[TN15] Michal Tabach and Talli Nachlieli. Classroom engagement towards using definitions for developing mathematical objects: the case of function. Educational Studies in Mathematics, 90(2):163–187, October 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9624-0.

Tabach:2016:CPL

[TN16] Michal Tabach and Talli Nachlieli. Communicational perspectives on learning and teaching mathematics: prologue. Educational Studies in Mathematics, 91(3):299–306, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9638-7.

Tsamir:2013:USG

[TO13] Pessia Tsamir and Regina Ovodenko. University students’ grasp of inflection points. Educational Studies in Mathematics, 83(3):409–427, July 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9463-1.

Tsamir:2024:EPC

[TOT24] Pessia Tsamir, Regina Ovodenko, and Dina Tirosh. Extrema points: concept images, mis-in and mis-out examples. Educational Studies in Mathematics, 115(1):13–33, January 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10273-6.

Touyarot:1971:IMR

[Tou71] M. A. Touyarot. Inquiry into the meanings of ‘relation’. Educational Studies in Mathematics, 3(2):270–276, April 1971.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305453.

Tourniaire:1986:PES

[Tou86] Françoise Tourniaire. Proportions in elementary school. Educational Studies in Mathematics, 17(4):401–412, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311327.

Toumasis:1990:EEG

[Tou90] Charalampos Toumasis. The epos of Euclidean geometry in Greek secondary education (1836–1985): Pressure for change and resistance. Educational Studies in Mathematics, 21(6): 491–508, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00315941.

Tourniaire:1985:PRR

[TP85] Françoise Tourniaire and Steven Pulos. Proportional reasoning: A review of the literature. Educational Studies in Mathematics, 16(2):181–204, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02400937.

Thom:2002:PPM

[TP02] Jennifer S. Thom and Susan E. B. Pirie. Problems, perseverance, and mathematical residue. Educational Studies in Mathematics, 50(1):1–28, May 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020507300013.

Triantafillou:2010:MPT

[TP10] Chrissavgi Triantafillou and Despina Potari. Mathematical practices in a technological workplace: the role of tools. Educational Studies in Mathematics, 74(3):275–294, July 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9237-6.

Triantafillou:2014:RPV

[TP14] Chrissavgi Triantafillou and Despina Potari. Revisiting the place value concept in the workplace context: the issue of transfer development. Educational Studies in Mathematics, 86 (3):337–358, July 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-014-9543-5.

Tunc-Pekkan:2015:AES

[TP15] Zelha Tunç-Pekkan. An analysis of elementary school children’s fractional knowledge depicted with circle, rectangle, and number line representations. Educational Studies in Mathematics, 89(3):419–441, July 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9606-2.

Tondorf:2022:CCE

[TP22] Alexandra Tondorf and Susanne Prediger. Connecting characterizations of equivalence of expressions: design research in grade 5 by bridging graphical and symbolic representations. Educational Studies in Mathematics, 111(3):399–422, November 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10158-0.

Threlfall:2007:IAP

[TPHS07] John Threlfall, Peter Pool, Matthew Homer, and Bronwen Swinnerton. Implicit aspects of paper and pencil mathematics assessment that come to light through the use of the computer. Educational Studies in Mathematics, 66(3):335–348, November 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9078-5.

Thom:2011:RES

[TR11] Jennifer S. Thom and Wolff-Michael Roth. Radical embodiment and semiotics: toward a theory of mathematics in the flesh. Educational Studies in Mathematics, 77(2–3):267–284, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9293-y.

Tuanaya:2024:BRE

[TR24] Rugaya Tuanaya and Heri Retnawati. Book review: Empowering primary mathematics educators: strategies for addressing anxiety in the classroom; Karen Wicks (2021) Tackling anxiety in primary mathematics teachers. Educational Studies in Mathematics, 117(1):163–166, September 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10300-0.

Tabach:2020:TAG

[TRDA20] Michal Tabach, Chris Rasmussen, Tommy Dreyfus, and Naneh Apkarian. Towards an argumentative grammar for networking: a case of coordinating two approaches. Educational Studies in Mathematics, 103(2):139–155, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-020-09934-7.

Trelinski:1983:SMS

[Tre83] G. Trelinski. Spontaneous mathematization of situations outside mathematics. Educational Studies in Mathematics, 14(3): 275–284, August 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00410542.

Treffers:1987:ICA

[Tre87] A. Treffers. Integrated column arithmetic according to progressive schematisation. Educational Studies in Mathematics, 18(2):125–145, May 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00314723.

Treffers:1991:MIP

[Tre91] Adrian Treffers. Meeting innumeracy at primary school. Educational Studies in Mathematics, 22(4):333–352, August 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00369294.

Treffers:1993:WFR

[Tre93] A. Treffers. Wiskobas and Freudenthal realistic mathematics education. Educational Studies in Mathematics, 25(1–2):

89–108, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274104.

Triandafillidis:1995:CVL

[Tri95] T. A. Triandafillidis. Circumventing visual limitations in teaching the geometry of shapes. Educational Studies in Mathematics, 29(3):225–235, October 1995. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274092.

Trouche:2000:PDG

[Tro00] Luc Trouche. La parabole du gaucher et de la casserole à bec verseur: ètude des processus d’apprentissage dans un environnement de calculatrices symboliques. (French) [The parabola of the left-handed and spout-casserole: study of learning processes in an environment of symbolic calculators]. Educational Studies in Mathematics, 41(3):239–264, March 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003939314034.

Trouche:2004:EIM

[Tro04] Luc Trouche. Environnements informatisés et mathématiques: quels usages pour quels apprentissages?. (French) [Computerized and mathematical environments: what uses for what learning?]. Educational Studies in Mathematics, 55(1–3): 181–197, March 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000017674.
82796.62.

Trumpower:2015:AFY

[Tru15] David L. Trumpower. Aspects of first year statistics students’ reasoning when performing intuitive analysis of variance: effects of within- and between-group variability. Educational Studies in Mathematics, 88(1):115–136, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9574-y.

Trzcieniecka-Schneider:1993:SRC

[TS93] Irena Trzcieniecka-Schneider. Some remarks on creating mathematical concepts. Educational Studies in Mathematics, 24 (3):257–264, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01275426.

Tirosh:1999:IRW

[TS99] Dina Tirosh and Ruth Stavy. Intuitive rules: A way to explain and predict students’ reasoning. Educational Studies in Mathematics, 38(1–3):51–66, March 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003436313032.

Tabach:2018:PDM

[TS18] Michal Tabach and Baruch B. Schwarz. Professional development of mathematics teachers toward the facilitation of smallgroup collaboration. Educational Studies in Mathematics, 97 (3):273–298, March 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9796-x.

Tsamir:1999:TCF

[Tsa99] Pessia Tsamir. The transition from comparison of finite to the comparison of infinite sets: Teaching prospective teachers. Educational Studies in Mathematics, 38(1–3):209–234, March 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003514208428.

Tsamir:2001:WSP

[Tsa01] Pessia Tsamir. When ‘the same’ is not perceived as such: The case of infinite sets. Educational Studies in Mathematics, 48(2–3):289–307, November 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016034917992.

Tsamir:2002:BRL

[Tsa02] Pessia Tsamir. Book review: Learning and Teaching Number Theory: Research in Cognition and Instruction by S. R. Campbell and R. Zazkis (eds.). Educational Studies in Mathematics, 51(3):271–286, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023682216523.

Tsamir:2007:WIB

[Tsa07] Pessia Tsamir. When intuition beats logic: Prospective teachers’ awareness of their same sides–same angles solutions. Educational Studies in Mathematics, 65(3):255–279, July 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9053-1.

Teule-Sensacq:1982:RPD

[TSV82] P. Teule-Sensacq and G. Vinrich. Résolution de problèmes de division au cycle élémentaire dans deux types de situations didactiques. (French) [Solving division problems at the elementary cycle in two types of teaching situations]. Educational Studies in Mathematics, 13(2):177–203, May 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00460710.

Tall:1991:EVT

[TT91] David Tall and Michael Thomas. Encouraging versatile thinking in algebra using the computer. Educational Studies in Mathematics, 22(2):125–147, April 1991. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00555720.

Tall:2001:INE

[TT01] David Tall and Dina Tirosh. Infinity — the never-ending struggle. Educational Studies in Mathematics, 48(2–3):129– 136, November 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1016019128773.

Thomaidis:2007:NHP

[TT07] Yannis Thomaidis and Constantinos Tzanakis. The notion of historical “parallelism” revisited: historical evolution and students’ conception of the order relation on the number line.

Educational Studies in Mathematics, 66(2):165–183, October 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9077-6.

Tsamir:2023:EME

[TT23] Pessia Tsamir and Dina Tirosh. Mis-in and mis-out concept images: the case of even numbers. Educational Studies in Mathematics, 112(2):207–224, February 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10183-z.

Tsamir:2008:INC

[TTL08] Pessia Tsamir, Dina Tirosh, and Esther Levenson. Intuitive nonexamples: the case of triangles. Educational Studies in Mathematics, 69(2):81–95, October 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9133-5.

Tirosh:2013:EYC

[TTL^{+}13] Dina Tirosh, Pessia Tsamir, Esther Levenson, Michal Tabach, and Ruthi Barkai. Exploring young children’s self-efficacy beliefs related to mathematical and nonmathematical tasks performed in kindergarten: abused and neglected children and their peers. Educational Studies in Mathematics, 83(2):309– 322, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9458-y.

Tsamir:2017:RPK

[TTL^{+}17] Pessia Tsamir, Dina Tirosh, Esther S. Levenson, Ruthi Barkai, and Michal Tabach. Repeating patterns in kindergarten: findings from children’s enactments of two activities. Educational Studies in Mathematics, 96(1):83–99, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9762-7.

Tan:2021:HHH

[TTS21] Da Yang Tan, Eng Guan Tay, and Paul M. E. Shutler. Hands, Head and Heart (3H) framework for curriculum review: emergence and nesting phenomena. Educational Studies in Mathematics, 106(2):189–210, February 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10003-2.

Tsamir:2010:MSM

[TTTL10] Pessia Tsamir, Dina Tirosh, Michal Tabach, and Esther Levenson. Multiple solution methods and multiple outcomes — is it a task for kindergarten children? Educational Studies in Mathematics, 73(3):217–231, April 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9215-z.

Tunis:2004:LE

[Tun04] Harry B. Tunis. Letter to the Editor. Educational Studies in Mathematics, 57(1):145–146, July 2004. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000047092.85145.50.

Turnau:1976:CRI

[Tur76] Stefan Turnau. On a class-room incident. Educational Studies in Mathematics, 7(1–2):147–155, July 1976. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144368.

Turnau:1980:MTY

[Tur80] Stefan Turnau. The mathematical textbook for young students. Educational Studies in Mathematics, 11(4):393–410, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00231213.

Turnau:1988:ASK

[Tur88] Stefan Turnau. Anna Sophia Krygowska. Educational Studies in Mathematics, 19(4):420–422, November 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00578691.

Turnau:1993:MER

[Tur93] Stefan Turnau. Mathematics education research in Poland. Educational Studies in Mathematics, 24(3):235–238, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01275423.

Tall:1981:CIC

[TV81] David Tall and Shlomo Vinner. Concept image and concept definition in mathematics with particular reference to limits and continuity. Educational Studies in Mathematics, 12(2): 151–169, May 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305619.

Tunks:2009:CPC

[TW09] Jeanne Tunks and Kirk Weller. Changing practice, changing minds, from arithmetical to algebraic thinking: an application of the concerns-based adoption model (CBAM). Educational Studies in Mathematics, 72(2):161–183, November 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9189-x.

Talmon:2004:UDB

[TY04] Varda Talmon and Michal Yerushalmy. Understanding dynamic behavior: Parent–child relations in dynamic geometry environments. Educational Studies in Mathematics, 57 (1):91–119, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000047052.
57084.d8.

Taole:1995:GIM

[TZLT95] James K. Taole, Marlies Zonneveld, and Lebohang Letsie-Taole. Gender interaction in mathematics classrooms: Reflection and transformation. Educational Studies in Mathematics, 28(3):263–274, April 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01274176.

Tzur:2007:FGA

[Tzu07] Ron Tzur. Fine grain assessment of students’ mathematical understanding: participatory and anticipatory stagesin learning a new mathematical conception. Educational Studies in Mathematics, 66(3):273–291, November 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-007-9082-4.

Urhan:2023:HCR

[UB23] Selin Urhan and Ali Bülbül. Habermas’ construct of rationality in the analysis of the mathematical problem-solving process. Educational Studies in Mathematics, 112(1):175–197, January 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10188-8.

Uhlig:2002:RPC

[Uhl02] Frank Uhlig. The role of proof in comprehending and teaching elementary linear algebra. Educational Studies in Mathematics, 50(3):335–346, May 2002. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1021245213997. See comments [DRR02,

Bur02a] and response [Uhl03].

Uhlig:2003:ARC

[Uhl03] Frank Uhlig. Author’s response to the comments on ‘The Role of Proof in … Teaching … Elementary Linear Algebra’, by Frank Uhlig, this Journal, vol. 50.3 (2002), p. 335–346. Educational Studies in Mathematics, 53(3):271–274, September 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026094903762. See [Uhl02].

Uscanga:2024:STA

[UMC24] Rosaura Uscanga, Kathleen Melhuish, and John Paul Cook. Students’ techniques for approaching defining properties of functions. Educational Studies in Mathematics, 117(3):457– 484, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10344-2.

ElRMS:1973:APF

[Une73] Une Équipe de l’I.R.E.M. de Strasbourg. Sur l’assimilation des programmes de 6ème–5ème. (French) [On the assimilation of programs of 6th–5th]. Educational Studies in Mathematics, 5(2):207–242, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01421027.

ElS:1974:ASN

[Une74] Une Équipe de l’Irem de Strasbourg. Sur l’acquisition des structures numériques en fin de 3^{ème}. (French) [On the acquisition of the numerical structures at the end of the third [grade]]. Educational Studies in Mathematics, 5(4):441–459, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01420657.

Uegatani:2025:DMM

[UOF25] Yusuke Uegatani, Hiroki Otani, and Taro Fujita. Decentralising mathematics: Mutual development of spontaneous and mathematical concepts via informal reasoning. Educational Studies in Mathematics, 118(2):229–248, February 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10366-w.

Urbanska:1993:NCS

[Urb93] Aleksandra Urbańska. On the numerical competence of sixyears-old children. Educational Studies in Mathematics, 24 (3):265–275, September 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01275427.

Usiskin:1974:SCP

[Usi74] Zalman Usiskin. Some corresponding properties of real numbers and implications for teaching. Educational Studies in Mathematics, 5(3):279–290, March 1974. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01424550.

Usiskin:1985:DPM

[Usi85] Zalman Usiskin. Didactical phenomenology of mathematical structures. Educational Studies in Mathematics, 16(2):223– 228, May 1985. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02400941.

Venkat:2018:MPM

[VA18] Hamsa Venkat and Mike Askew. Mediating primary mathematics: theory, concepts, and a framework for studying practice. Educational Studies in Mathematics, 97(1):71–92, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-017-9776-1.

Venkat:2021:SSD

[VAM21] Hamsa Venkat, Mike Askew, and Samantha Morrison. Shapeshifting Davydov’s ideas for early number learning in South Africa. Educational Studies in Mathematics, 106(3):397–412, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09993-w.

VanDormolen:1969:UVD

[Van69a] J. Van Dormolen. The uselessness of Venn diagrams. Educational Studies in Mathematics, 1(4):402–407, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02396313.

VanHiele:1969:QAD

[Van69b] P. M. Van Hiele. Quelques aspects didactiques du développement de la pensée des enfants dans les mathématiques et la physique. (French) [Some didactic aspects of the development of children’s thinking in mathematics and physics]. Educational Studies in Mathematics, 1(3):343–346, January 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00558320.

VanHiele:1974:SST

[Van74] P. M. Van Hiele. System separation and transfer. Educational Studies in Mathematics, 5(4):413–417, December 1974. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01420654.

VanHiele:1976:WKM

[Van76] P. M. Van Hiele. Wie kann man im Mathematikunterricht den Denkstufen Rechnung tragen?. (German) [How can one in mathematics teaching teach thought levels?]. Educational Studies in Mathematics, 7(1–2):157–169, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144369.

VanDormolen:1977:LUW

[Van77] J. Van Dormolen. Learning to understand what giving a proof really means. Educational Studies in Mathematics, 8(1):27– 34, April 1977. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00302502.

VanDenBrink:1984:NCF

[Van84] F. J. Van Den Brink. Numbers in contextual frameworks. Educational Studies in Mathematics, 15(3):239–257, August 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00312076.

VanDenBrink:1993:DAD

[Van93a] Jan Van Den Brink. Different aspects in designing mathematics education: Three examples from the Freudenthal Institute. Educational Studies in Mathematics, 24(1):35–64, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01273294.

VanEst:1993:HFS

[Van93b] W. T. Van Est. Hans Freudenthal (17 September 1905–13 October 1990). Educational Studies in Mathematics, 25(1– 2):59–69, March 1993. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274102.

VanDenHeuvel-Panhuizen:1994:IDA

[Van94] Marja Van Den Heuvel-Panhuizen. Improvement of (didactical) assessment by improvement of problems: An attempt with respect to percentage. Educational Studies in Mathematics, 27(4):341–372, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273377.

VanAmerom:2003:FIS

[Van03a] Barbara A. Van Amerom. Focusing on informal strategies when linking arithmetic to early algebra. Educational Studies in Mathematics, 54(1):63–75, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000005237.72281.bf.

VanDenHeuvel-Panhuizen:2003:DUM

[Van03b] Marja Van Den Heuvel-Panhuizen. The didactical use of models in realistic mathematics education: An example from a longitudinal trajectory on percentage. Educational Studies in Mathematics, 54(1):9–35, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000005212.03219.dc.

Varga:1972:LPL

[Var72] T. Varga. Logic and probability in the lower grades. Educational Studies in Mathematics, 4(3):346–357, April 1972. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302583.

Varga:1976:PST

[Var76] Tamás Varga. On primary school teachers’ mathematics. Educational Studies in Mathematics, 7(1–2):171–177, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144370.

Vaughan:1971:DEG

[Vau71] Herbert E. Vaughan. The development of Euclidean geometry in terms of translations. Educational Studies in Mathematics, 4(1):104–110, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305801.

Vermeulen:2021:TBV

[VBE21] Jorine A. Vermeulen, Anton Béguin, and Theo J. H. M. Eggen. Task beliefs and the voluntary use of the empty number line in third-grade subtraction and addition. Educational Studies in Mathematics, 106(2):231–249, February 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-10016-x.

vanBommel:2019:BRB

[vBP19] Jorryt van Bommel and Hanna Palmér. Book review: Building the foundation: Whole numbers in the primary grades. Maria G. Bartolini Bussi and Xu Hua Sun (Eds.) (2018) The 23rd ICMI study — new ICMI study series. Educational Studies in Mathematics, 100(2):193–199, February 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9866-8.

Verschaffel:2012:I

[VBT12] Lieven Verschaffel, Peter Bryant, and Joke Torbeyns. Introduction. Educational Studies in Mathematics, 79(3):327–334, March 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9381-2.

Valoyes-Chavez:2019:MNM

[VC19] Luz Valoyes-Chávez. On the making of a new mathematics teacher: professional development, subjectivation, and resistance to change. Educational Studies in Mathematics, 100(2): 177–191, February 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9869-5.

Valoyes-Chavez:2024:IEP

[VCD24] Luz Valoyes-Chávez and Lisa Darragh. Interrogating the equity promise for Black immigrant students in reformed mathematics classrooms. Educational Studies in Mathematics, 116 (3):479–499, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10314-8.

Vithal:1995:PWU

[VCS95] Renuka Vithal, Iben Christiansen, and Ole Skovsmose. Project work in university mathematics education. Educational Studies in Mathematics, 29(2):199–223, September 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01274213.

Vandewiele:1978:QDM

[VD78] M. Vandewiele and W. D’Hondt. Quelques difficultés en mathématique et leur perception par les élèves en classe de troisième dans l’enseignement secondaire à Dakar. (French) [Some difficulties in mathematics and their perception by pupils in third grade in secondary education in Dakar]. Educational Studies in Mathematics, 9(4):439–460, November 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410589.

Vandewiele:1979:PQD

[VD79] M. Vandewiele and W. D’Hondt. La perception de quelques difficultés en mathématiques par les professeurs en classe de troisième dans l’enseignement secondaire au Sénégal. (French) [The perception of some difficulties in mathematics by thirdgrade teachers in secondary education in Senegal]. Educational Studies in Mathematics, 10(2):199–225, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00230987.

VanDooren:2003:ILE

[VDD^{+}03] Wim Van Dooren, Dirk De Bock, Fien Depaepe, Dirk Janssens, and Lieven Verschaffel. The illusion of linearity: Expanding the evidence towards probabilistic reasoning. Educational Studies in Mathematics, 53(2):113–138, July 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1025516816886.

vanDijke-Droogers:2021:SMP

[vDDDB21] Marianne van Dijke-Droogers, Paul Drijvers, and Arthur Bakker. Statistical modeling processes through the lens of instrumental genesis. Educational Studies in Mathematics, 107 (2):235–260, June 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-10023-y.

vandenHeuvel-Panhuizen:2013:PSS

[vdHPKR13] Marja van den Heuvel-Panhuizen, Angeliki Kolovou, and Alexander Robitzsch. Primary school students’ strategies in early algebra problem solving supported by an online game. Educational Studies in Mathematics, 84(3):281–307, November 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-013-9483-5.

vandenKieboom:2022:PST

[vdKG22] Leigh A. van den Kieboom and Sally V. Groleau. Pre-service teacher planning for differentiation of instruction in mathematics classrooms. Educational Studies in Mathematics, 111 (2):225–252, October 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10149-1.

vandeSande:2011:DCS

[vdS11] Carla van de Sande. A description and characterization of student activity in an open, online, mathematics help forum. Educational Studies in Mathematics, 77(1):53–78, May 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9300-y.

VanDerAuwera:2024:CSA

[VDTV24] Stijn Van Der Auwera, Bert De Smedt, Joke Torbeyns, and Lieven Verschaffel. Children’s subtraction by addition strategy use and their subtraction-related conceptual knowledge. Educational Studies in Mathematics, 115(1):35–50, January 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-023-10276-3.

VanDooren:2004:PPI

[VDWV04] Wim Van Dooren, Dirk De Bock, Dave Weyers, and Lieven Verschaffel. The predictive power of intuitive rules: A critical analysis of the impact of ‘more a–more b’ and ‘same a–same b’. Educational Studies in Mathematics, 56(2–3):179–207, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000040379.26033.0d.

Vroom:2024:SSS

[VE24] Kristen Vroom and Brittney Ellis. Sociomathematical scaffolding as students engage in revising draft definitions, conjectures, and proofs. Educational Studies in Mathematics, 116(1): 27–47, May 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-023-10294-1.

Vroom:2024:TPS

[VEBS24] Kristen Vroom, Tenchita Alzaga Elizondo, José Saúl Barbosa, and Stephen Strand II. Teaching practices that support revising definition drafts to adhere to mathematical norms. Educational Studies in Mathematics, 117(2):285–302, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-024-10331-7.

Veksler:1971:DCP

[Vek71] S. I. Veksler. Developing the capabilities of the pupil. Educational Studies in Mathematics, 3(2):180–183, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305446.

Vergnaud:1979:AAC

[Ver79] Gérard Vergnaud. The acquisition of arithmetical concepts. Educational Studies in Mathematics, 10(2):263–274, May 1979. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00230990.

Vergnaud:2012:C

[Ver12] Gerard Vergnaud. Commentary 1. Educational Studies in Mathematics, 79(3):439–445, March 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9382-1.

Vest:1969:CMO

[Ves69] Floyd R. Vest. A catalog of models for the operations of addition and subtraction of whole numbers. Educational Studies in Mathematics, 2(1):59–68, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00368985.

Vest:1971:CMM

[Ves71] Floyd R. Vest. A catalog of models for multiplication and division of whole numbers. Educational Studies in Mathematics, 3 (2):220–228, April 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305450.

Vest:1973:UMOa

[Ves73a] Floyd Vest. Using models of operations and equations. Educational Studies in Mathematics, 5(1):147–155, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00684694.

Vest:1973:UMOb

[Ves73b] Floyd Vest. Using models of operations and equations. Educational Studies in Mathematics, 5(2):147–155, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01421022.

Vest:1976:TPS

[Ves76] Floyd Vest. Teaching problem solving as viewed through a theory of models. Educational Studies in Mathematics, 6(4): 395–408, March 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00411088.

vonGlasersfeld:1992:GE

[vG92] Ernst von Glasersfeld. Guest editorial. Educational Studies in Mathematics, 23(5):443–444, October 1992. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00571466.

Voutsina:2019:MAY

[VGJ19] Chronoula Voutsina, Lois George, and Keith Jones. Microgenetic analysis of young children’s shifts of attention in arithmetic tasks: underlying dynamics of change in phases of seemingly stable task performance. Educational Studies in Mathematics, 102(1):47–74, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/article/10.1007/s10649-019-09883-
w;   http://link.springer.com/content/pdf/10.1007/
s10649-019-09883-w.pdf.

Venenciano:2016:PTM

[VH16] Linda Venenciano and Ronald Heck. Proposing and testing a model to explain traits of algebra preparedness. Educational Studies in Mathematics, 92(1):21–35, May 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9672-5.

Villa:1971:LAT

[Vil71] Mario Villa. A logical approach to the teaching of geometry at the secondary level. Educational Studies in Mathematics, 4 (1):111–134, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00305802.

Vilela:2010:DPB

[Vil10] Denise Silva Vilela. Discussing a philosophical background for the ethnomathematical program. Educational Studies in Mathematics, 75(3):345–358, December 2010. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9261-6.

Vinner:1975:NPA

[Vin75] Shlomo Vinner. The naive platonic approach as a teaching strategy in arithmetics. Educational Studies in Mathematics, 6(3):339–350, November 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01793616.

Vinner:1976:NCD

[Vin76] Shlomo Vinner. The naive concept of definition in mathematics. Educational Studies in Mathematics, 7(4):413–429, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00452223.

Vinner:1977:CEU

[Vin77] Shlomo Vinner. The concept of exponentiation at the undergraduate level and the definitional approach. Educational Studies in Mathematics, 8(1):17–26, April 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00302501.

Vinner:1997:PCP

[Vin97] Shlomo Vinner. The pseudo-conceptual and the pseudoanalytical thought processes in mathematics learning. Educational Studies in Mathematics, 34(2):97–129, November 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002998529016.

Vale:2004:SVC

[VL04] Colleen M. Vale and Gilah C. Leder. Student views of computer-based mathematics in the middle years: Does gender make a difference? Educational Studies in Mathematics, 56(2– 3):287–312, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040411.
94890.56.

Vlassis:2002:BMH

[Vla02] Joëlle Vlassis. The balance model: Hindrance or support for the solving of linear equations with one unknown. Educational Studies in Mathematics, 49(3):341–359, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020229023965.

vanLieshout:2018:PRS

[vLXD18] Ernest C. D. M. van Lieshout and Iro Xenidou-Dervou. Pictorial representations of simple arithmetic problems are not always helpful: a cognitive load perspective. Educational Studies in Mathematics, 98(1):39–55, May 2018. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-017-9802-3;   http://link.springer.com/
content/pdf/10.1007/s10649-017-9802-3.pdf.

Vysotskaya:2021:CPD

[VLY21] Elena Vysotskaya, Anastasia Lobanova, and Maria Yanishevskaya. The challenge of proportion: does it require rethinking of the measurement paradigm? Educational Studies in Mathematics, 106(3):429–446, March 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-09987-8.

Vidal-Madjar:1978:TMS

[VM78] A. Vidal-Madjar. Teaching mathematics and statistics to adults who are keen on psychology. Educational Studies in Mathematics, 9(4):381–390, November 1978. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00410584.

Verzosa:2013:LSA

[VM13] Debbie Bautista Verzosa and Joanne Mulligan. Learning to solve addition and subtraction word problems in English as an imported language. Educational Studies in Mathematics, 82(2):223–244, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9420-z.

Villarroel:2011:OCS

[VMN11] José Domingo Villarroel, Margarita Miñón, and Teresa Nuño. The origin of counting: A study of the early meaning of ‘one’, ‘two’ and ‘three’ among Basque- and Spanish-speaking children. Educational Studies in Mathematics, 76(3):345–361, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-010-9291-0.

Viirman:2019:NDD

[VN19] Olov Viirman and Elena Nardi. Negotiating different disciplinary discourses: biology students’ ritualized and exploratory participation in mathematical modeling activities. Educational Studies in Mathematics, 101(2):233–252, June 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9861-0;   http://link.springer.com/
content/pdf/10.1007/s10649-018-9861-0.pdf.

vanOers:2001:EFI

[vO01] Bert van Oers. Educational forms of initiation in mathematical culture. Educational Studies in Mathematics, 46(1–3):59–85, March 2001. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1014031507535.

vanOers:2010:EMT

[vO10] Bert van Oers. Emergent mathematical thinking in the context of play. Educational Studies in Mathematics, 74(1):23– 37, May 2010. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10649-009-9225-x.pdf.

vanOers:2013:CIM

[vO13] Bert van Oers. Challenges in the innovation of mathematics education for young children. Educational Studies in Mathematics, 84(2):267–272, October 2013. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-013-9509-z.pdf.

Villarroel:2017:SRS

[VO17] José Domingo Villarroel and Olga Sanz Ortega. A study regarding the spontaneous use of geometric shapes in young children’s drawings. Educational Studies in Mathematics, 94(1): 85–95, January 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9718-3.

Vogeli:1970:SAC

[Vog70] Bruce R. Vogeli. Sweep away all cows, ghosts, dragons and devils. Educational Studies in Mathematics, 2(4):496–500, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00417631.

Vogel:2013:MSP

[Vog13] Rose Vogel. Mathematical situations of play and exploration. Educational Studies in Mathematics, 84(2):209–225, October 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9504-4.

Voigt:1994:NMM

[Voi94] Jörg Voigt. Negotiation of mathematical meaning and learning mathematics. Educational Studies in Mathematics, 26(2–3): 275–298, March 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273665.

Vollrath:1969:SAA

[Vol69] Hans-Joachim Vollrath. Some algebraic aspects in analysis teaching. Educational Studies in Mathematics, 1(4):440–444, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02396317.

Vollrath:1976:PGM

[Vol76] H. J. Vollrath. The place of geometry in mathematics teaching: An analysis of recent developments. Educational Studies in Mathematics, 7(4):431–442, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00452224.

Vollrath:1977:USS

[Vol77] Hans-Joachim Vollrath. The understanding of similarity and shape in classifying tasks. Educational Studies in Mathematics, 8(2):211–224, August 1977. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00241026.

Vollrath:1986:SSI

[Vol86] Hans-Joachim Vollrath. Search strategies as indicators of functional thinking. Educational Studies in Mathematics, 17(4):

387–400, November 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00311326.

Voutsina:2012:PCC

[Vou12] Chronoula Voutsina. Procedural and conceptual changes in young children’s problem solving. Educational Studies in Mathematics, 79(2):193–214, February 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9334-1.

Voutsina:2016:OCS

[Vou16] Chronoula Voutsina. Oral counting sequences: a theoretical discussion and analysis through the lens of representational redescription. Educational Studies in Mathematics, 93(2):175– 193, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9705-8.

VanHarpen:2013:IRB

[VP13] Xianwei Y. Van Harpen and Norma C. Presmeg. An investigation of relationships between students’ mathematical problem-posing abilities and their mathematical content knowledge. Educational Studies in Mathematics, 83(1):117– 132, May 2013. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-012-9456-0.

Venenciano:2021:IMP

[VPR21] Linda Venenciano, Elena Polotskaia, and Luis Radford. An introduction to multiple perspectives on Davydov’s approach in the XXI century. Educational Studies in Mathematics, 106 (3):323–326, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10040-5.

Vlaardingerbroek:1990:ETR

[VR90] Barend Vlaardingerbroek and Lawrence Ros. Educational transition rates and upper secondary students’ arithmetical ability in five developing countries. Educational Studies in Mathematics, 21(5):451–460, October 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00398863.

Vredenduin:1969:CFV

[Vre69] P. G. J. Vredenduin. The conclusive force of Venn diagrams, and the implication. Educational Studies in Mathematics, 1 (4):394–401, March 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02396312.

Valenta:2024:FRS

[VRK24] Anita Valenta, Kirsti Rø, and Sigrid Iversen Klock. A framework for reasoning in school mathematics: analyzing the development of mathematical claims. Educational Studies in Mathematics, 116(1):91–111, May 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10309-5.

Vahey:2012:CDA

[VRP^{+}12] Philip Vahey, Ken Rafanan, Charles Patton, Karen Swan, Mark van ’t Hooft, Annette Kratcoski, and Tina Stanford. A cross-disciplinary approach to teaching data literacy and proportionality. Educational Studies in Mathematics, 81(2): 179–205, October 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9392-z.

VanDenBrink:1979:YCR

[VS79] Jan Van Den Brink and Leen Streefland. Young children (6– 8)-ratio and proportion. Educational Studies in Mathematics, 10(4):403–420, November 1979. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00417087.

Vithal:1997:EIC

[VS97] Renuka Vithal and Ole Skovsmose. The end of innocence: A critique of ‘ethnomathematics’. Educational Studies in Mathematics, 34(2):131–157, November 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002971922833.

VanHarpen:2013:CMP

[VS13] Xianwei Y. Van Harpen and Bharath Sriraman. Creativity and mathematical problem posing: an analysis of high school students’ mathematical problem posing in China and the USA. Educational Studies in Mathematics, 82(2):201–221, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9419-5.

Voutsina:2023:PCC

[VS23] Chronoula Voutsina and Debbie Stott. Preschool children’s conceptions of the meanings and use of written numerals in everyday life: a phenomenographic study of the nature and structure of qualitative variation. Educational Studies in Mathematics, 114(2):249–275, October 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10232-1.

Voutsina:2025:YCN

[VS25] Chronoula Voutsina and Debbie Stott. Young children’s notations for representing ordinal position and quantity: a longitudinal study of cross-task and within-task qualitative variations. Educational Studies in Mathematics, 120(2):317–340, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10422-z.

vanSchalkwijk:2000:LPI

[vSBvR00] Louis van Schalkwijk, Theo Bergen, and Arnoud van Rooij. Learning to prove by investigations: a promising approach in Dutch secondary education. Educational Studies in Mathematics, 43(3):293–311, November 2000. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1011963320625.

Voica:2020:HMS

[VSS20] Cristian Voica, Florence Mihaela Singer, and Emil Stan. How are motivation and self-efficacy interacting in problem-solving and problem-posing? Educational Studies in Mathematics, 105 (3):487–517, November 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-10005-0.

Vizek:2024:AQC

[VSS24] Lukáš Vízek, Libuše Samková, and Jon R. Star. Assessing the quality of conceptual knowledge through dynamic constructions. Educational Studies in Mathematics, 117(2):167–191, October 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10349-x.

Vallejo-Vargas:2025:REP

[VV25] Estela A. Vallejo-Vargas. The role of examples in proving and disproving: extending an existing framework. Educational Studies in Mathematics, 119(1):101–116, May 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10378-6.

Vicente:2022:AWP

[VVM22] Santiago Vicente, Lieven Verschaffel, and David Múñez. Arithmetic word problem solving. Analysis of Singaporean and Spanish textbooks. Educational Studies in Mathematics, 111 (3):375–397, November 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-022-10169-x.

Vamvakoussi:2013:BRE

[VVV13] Xenia Vamvakoussi, Wim Van Dooren, and Lieven Verschaffel. Brief report. Educated adults are still affected by intuitions about the effect of arithmetical operations: evidence from a reaction-time study. Educational Studies in Mathematics, 82(2):323–330, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-012-
9432-8.

VanHoof:2015:IAN

[VVV15] Jo Van Hoof, Lieven Verschaffel, and Wim Van Dooren. Inappropriately applying natural number properties in rational number tasks: characterizing the development of the natural number bias through primary and secondary education. Educational Studies in Mathematics, 90(1):39–56, September 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-015-9613-3.

Vanluydt:2022:RRP

[VVV22] Elien Vanluydt, Lieven Verschaffel, and Wim Van Dooren. The role of relational preference in word-problem solving in 6- to 7-year-olds. Educational Studies in Mathematics, 110(3):393– 411, July 2022. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10139-9.

Vanluydt:2021:ECM

[VWV21] Elien Vanluydt, Nore Wijns, and Wim Van Dooren. Early childhood mathematical development: the association between patterning and proportional reasoning. Educational Studies in Mathematics, 107(1):93–110, May 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-020-10017-w.

Vysin:1975:ECE

[Vyš75] J. Vyšín. An experience of the Czechoslovakian experimental center. Educational Studies in Mathematics, 6(1):119–123, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00590030.

Venenciano:2021:DRT

[VYZ21] Linda C. H. Venenciano, Seanyelle L. Yagi, and Fay K. Zenigami. The development of relational thinking: a study of Measure Up first-grade students’ thinking and their symbolic understandings. Educational Studies in Mathematics, 106(3):413–428, March 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-020-10014-z.

Weber:2004:SSP

[WA04] Keith Weber and Lara Alcock. Semantic and syntactic proof productions. Educational Studies in Mathematics, 56(2–3): 209–234, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000040410.
57253.a1.

Wagner:2012:OMT

[Wag12] David Wagner. Opening mathematics texts: resisting the seduction. Educational Studies in Mathematics, 80(1–2):153– 169, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9372-8.

Waits:1970:RET

[Wai70] Bert K. Waits. Relative effectiveness of two different television techniques and one large lecture technique for teaching large enrollment college mathematics courses. Educational Studies in Mathematics, 2(4):476–477, February 1970. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00417629.

Wake:2014:MSM

[Wak14] Geoff Wake. Making sense of and with mathematics: the interface between academic mathematics and mathematics in practice. Educational Studies in Mathematics, 86(2):271–290, June 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9540-8.

Walker:1968:SUC

[Wal68] R. J. Walker. Student use of a computer. Educational Studies in Mathematics, 1(1–2):111–117, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426238.

Wallrabenstein:1973:DSG

[Wal73a] Hartmut Wallrabenstein. Development and signification of a geometry test. Educational Studies in Mathematics, 5(1):81– 89, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01425414.

Wallrabenstein:1973:ETI

[Wal73b] Hartmut Wallrabenstein. Experiments in teaching intuitive topology in the 5th and 6th grades. Educational Studies in Mathematics, 5(1):91–108, April 1973. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01425415.

Walther:1976:BGT

[Wal76] G. Walther. Balancierte Graphen — ein Thema für die didaktische Diskussion zum Geometrieunterricht. (German) [Balanced graphs — a topic for didactic discussion on teaching of geometry]. Educational Studies in Mathematics, 7(4):465– 495, December 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00452226.

Walshaw:2010:BRT

[Wal10] Margaret Anne Walshaw. Book review: Teacher-centredness and student-centredness under interrogation. Educational Studies in Mathematics, 73(1):99–103, January 2010. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9221-1.

Walshaw:2012:BRI

[Wal12] Margaret Walshaw. Book review: Interpreting how students come to understand mathematics in the digital environment. Nigel Calder (2011). Processing mathematics through digital technologies: The primary years. Educational Studies in Mathematics, 81(3):401–405, November 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9403-0.

Walshaw:2017:BRL

[Wal17] Margaret Walshaw. Book review: Leveraging the potential of research on classroom discourse for enhancing mathematics teaching and learning, Jill Adler and Anna Sfard (Eds.) (2017) Research for Educational Change: Transforming Researchers ’ Insights Into Improvement in Mathematics Teaching and Learning. Educational Studies in Mathematics, 95 (3):353–357, July 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Watson:2018:CSU

[WAL18] Anne Watson, Michal Ayalon, and Stephen Lerman. Comparison of students’ understanding of functions in classes following English and Israeli national curricula. Educational Studies in Mathematics, 97(3):255–272, March 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9798-8.

Wang:2025:BRA

[Wan25] Haoyi Wang. Book review: Applying values in mathematics education. Yüksel Dede, Gosia Marschall, and Philip Clarkson (Eds.) (2024) Values and valuing in mathematics education: moving forward into practice. Educational Studies in Mathematics, 118(1):139–151, January 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10329-1.

Warren:2006:CML

[War06] Elizabeth Warren. Comparative mathematical language in the elementary school: A longitudinal study. Educational Studies in Mathematics, 62(2):169–189, June 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-4627-5.

Wasserman:2015:UTM

[Was15] Nicholas H. Wasserman. Unpacking teachers’ moves in the classroom: navigating micro- and macro-levels of mathematical complexity. Educational Studies in Mathematics, 90(1): 75–93, September 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9615-1.

Wasserman:2022:UFM

[Was22] Nicholas H. Wasserman. Unpacking foreshadowing in mathematics teachers’ planned practices. Educational Studies in Mathematics, 111(3):423–443, November 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10152-6.

Watson:1980:IEB

[Wat80] Ivan Watson. Investigating errors of beginning mathematicians. Educational Studies in Mathematics, 11(3):319–329, August 1980. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00697743.

Watson:1987:LAN

[Wat87] Helen Watson. Learning to apply numbers to nature. Educational Studies in Mathematics, 18(4):339–357, November 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00240984.

Watson:2000:MTA

[Wat00] Anne Watson. Mathematics teachers acting as informal assessors: Practices, problems and recommendations. Educational Studies in Mathematics, 41(1):69–91, January 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003933431489.

Watson:2001:LDI

[Wat01] Jane M. Watson. Longitudinal development of inferential reasoning by school students. Educational Studies in Mathematics, 47(3):337–372, September 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015158813656.

Watson:2002:IRI

[Wat02] J. M. Watson. Inferential reasoning and the influence of cognitive conflict. Educational Studies in Mathematics, 51(3):225– 256, November 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1023622017006.

Watt:2005:AUA

[Wat05] Helen M. G. Watt. Attitudes to the use of alternative assessment methods in mathematics: A study with secondary mathematics teachers in Sydney, Australia. Educational Studies in Mathematics, 58(1):21–44, January 2005.

CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-3228-z.

Watson:2007:RCC

[Wat07] Jane M. Watson. The role of cognitive conflict in developing students’ understanding of average. Educational Studies in Mathematics, 65(1):21–47, May 2007. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9043-3.

Waywood:1994:IWL

[Way94] Andrew Waywood. Informal writing-to-learn as a dimension of a student profile. Educational Studies in Mathematics, 27(4): 321–340, December 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01273376.

Winter:1984:BR

[WB84] Heinrich Winter and Alan J. Bishop. Book reviews. Educational Studies in Mathematics, 15(4):443–446, November 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311117.

Walter:2011:SBM

[WB11] Janet G. Walter and Tara Barros. Students build mathematical theory: semantic warrants in argumentation. Educational Studies in Mathematics, 78(3):323–342, December 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9326-1.

Walshaw:2012:APM

[WB12] Margaret Walshaw and Tony Brown. Affective productions of mathematical experience. Educational Studies in Mathematics, 80(1–2):185–199, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-011-
9370-x.

Walkowiak:2014:IOM

[WBM^{+}14] Temple A. Walkowiak, Robert Q. Berry, J. Patrick Meyer, Sara E. Rimm-Kaufman, and Erin R. Ottmar. Introducing an observational measure of standards-based mathematics teaching practices: Evidence of validity and score reliability. Educational Studies in Mathematics, 85(1):109–128, January 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-013-9499-x.

Wagner:2020:WCW

[WBM^{+}20] David Wagner, Arthur Bakker, Tamsin Meaney, Vilma Mesa, Susanne Prediger, and Wim Van Dooren. What can we do against racism in mathematics education research? Educational Studies in Mathematics, 104(3):299–311, July 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/
10.1007/s10649-020-09969-w;   http://link.springer.
com/content/pdf/10.1007/s10649-020-09969-w.pdf.

Warren:2008:GPR

[WC08] Elizabeth Warren and Tom Cooper. Generalising the pattern rule for visual growth patterns: Actions that support 8-year olds’ thinking. Educational Studies in Mathematics, 67(2): 171–185, February 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9092-2.

Watson:2005:PTD

[WD05] Anne Watson and Els De Geest. Principled teaching for deep progress: Improving mathematical learning beyond methods and materials. Educational Studies in Mathematics, 58(2): 209–234, February 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-005-2756-x.

Wagner:2010:FNG

[WD10] David Wagner and Brent Davis. Feeling number: grounding number sense in a sense of quantity. Educational Studies in Mathematics, 74(1):39–51, May 2010. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9226-9.

Watson:2014:DIC

[WD14] Anne Watson and Els De Geest. Department-initiated change. Educational Studies in Mathematics, 87(3):351–368, November 2014. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9549-z.

Worthington:2019:DMA

[WDvO19] Maulfry Worthington, Marjolein Dobber, and Bert van Oers. The development of mathematical abstraction in the nursery. Educational Studies in Mathematics, 102(1):91–110, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-019-09898-3;   http://link.
springer.com/content/pdf/10.1007/s10649-019-09898-
3.pdf.

Warren:1995:FPS

[WE95] Elizabeth Warren and Lyn English. Facility with plane shapes: A multifaceted skill. Educational Studies in Mathematics, 28 (4):365–383, June 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01274079.

Wearne:1990:AMD

[Wea90] Diana Wearne. Acquiring meaning for decimal fraction symbols: A one year follow-up. Educational Studies in Mathematics, 21(6):545–564, December 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00315944.

Weber:2001:SDC

[Web01] Keith Weber. Student difficulty in constructing proofs: The need for strategic knowledge. Educational Studies in Mathematics, 48(1):101–119, October 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015535614355.

Weber:2018:BRB

[Web18a] Keith Weber. Book review: Baruch B. Schwarz and Michael J. Baker (Eds) (Foreword by L. B. Resnick with F. Schantz)

(2017) Dialogue, Argumentation and Education: History, Theory and Practice. Educational Studies in Mathematics, 97(1): 111–118, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-017-9780-5. See correction

[Web18b].

Weber:2018:CBR

[Web18b] Keith Weber. Correction to: Book review: Baruch B. Schwarz and Michael J. Baker (Foreword by L. B. Resnick with F. Schantz) (2017) Dialogue, Argumentation and Education: History, Theory and Practice. Educational Studies in Mathematics, 97(1):119, January 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10649-017-
9792-1.pdf;   http://link.springer.com/article/10.
1007/s10649-017-9792-1. See [Web18a].

Wedege:1999:KKM

[Wed99] Tine Wedege. To know or not to know — mathematics, that is a question of context. Educational Studies in Mathematics, 39(1–3):205–227, June 1999. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003871930181.

Weigand:1991:IST

[Wei91] Hans-Georg Weigand. Iteration sequences and their representations. Educational Studies in Mathematics, 22(5):411–437, October 1991. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00367906.

Weiland:2017:PSL

[Wei17] Travis Weiland. Problematizing statistical literacy: An intersection of critical and statistical literacies. Educational Studies in Mathematics, 96(1):33–47, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9764-5.

Wittmann:2013:AMM

[WFB13] Michael C. Wittmann, Virginia J. Flood, and Katrina E. Black. Algebraic manipulation as motion within a landscape. Educational Studies in Mathematics, 82(2):169–181, February 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9428-4.

Weber:2023:WML

[WFC23] Keith Weber and Timothy Fukawa-Connelly. What mathematicians learn from attending other mathematicians’ lectures. Educational Studies in Mathematics, 112(1):123–139, January 2023. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-022-10177-x.

Weinberg:2015:CIG

[WFCW15] Aaron Weinberg, Tim Fukawa-Connelly, and Emilie Wiesner. Characterizing instructor gestures in a lecture in a proof-based mathematics class. Educational Studies in Mathematics, 90(3): 233–258, November 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9623-1.

Walter:2007:TPA

[WG07] Janet G. Walter and Hope Gerson. Teachers’ personal agency: Making sense of slope through additive structures. Educational Studies in Mathematics, 65(2):203–233, June 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9048-y.

Weiss:2015:ERT

[WH15a] Michael Weiss and Patricio Herbst. Erratum to: The role of theory building in the teaching of secondary geometry. Educational Studies in Mathematics, 89(2):231, June 2015. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-015-9604-4.pdf. See [WH15b].

Weiss:2015:RTB

[WH15b] Michael Weiss and Patricio Herbst. The role of theory building in the teaching of secondary geometry. Educational Studies in Mathematics, 89(2):205–229, June 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-015-9599-x. See er-

ratum [WH15a].

Wilkie:2024:GAS

[WH24] Karina J. Wilkie and Sarah Hopkins. Generalizing actions with the subtraction-compensation property: primary students’ algebraic thinking with tasks involving vertical towers of blocks. Educational Studies in Mathematics, 115(3):433–458, March 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10303-x.

Whang:1996:IEK

[Wha96] Woo-Hyung Whang. The influence of English–Korean bilingualism in solving mathematics word problems. Educational Studies in Mathematics, 30(3):289–312, April 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00304569.

Weiss:2009:TPA

[WHC09] Michael Weiss, Patricio Herbst, and Chialing Chen. Teachers’ perspectives on “authentic mathematics” and the twocolumn proof form. Educational Studies in Mathematics, 70 (3):275–293, April 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9144-2.

Wheatley:1992:RRM

[Whe92] Grayson H. Wheatley. The role of reflection in mathematics learning. Educational Studies in Mathematics, 23(5):529–541, October 1992. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00571471.

Wheeler:1996:BR

[Whe96] David Wheeler. Book review. Educational Studies in Mathematics, 31(3):339–348, October 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF00376326.

Wagner:2008:JDS

[WHE08] David Wagner and Beth Herbel-Eisenmann. “Just don’t”: The suppression and invitation of dialogue in the mathematics classroom. Educational Studies in Mathematics, 67(2):143– 157, February 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9097-x.

Wagner:2009:RMM

[WHE09] David Wagner and Beth Herbel-Eisenmann. Re-mythologizing mathematics through attention to classroom positioning. Educational Studies in Mathematics, 72(1):1–15, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9178-5.

Whitney:1987:CAS

[Whi87] Hassler Whitney. Coming alive in school math and beyond. Educational Studies in Mathematics, 18(3):229–242, August 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386196.

White:2009:EOD

[Whi09] Tobin White. Encrypted objects and decryption processes: problem-solving with functions in a learning environment based on cryptography. Educational Studies in Mathematics, 72(1):17–37, September 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-008-9180-y.
pdf.

Williams:2001:AAM

[WI01] Steven R. Williams and Kathy M. C. Ivey. Affective assessment and mathematics classroom engagement: A case study. Educational Studies in Mathematics, 47(1):75–100, May 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017987500929.

Williams:1971:PPA

[Wil71] John Williams. Problems and possibilities in the assessment and investigation of mathematics learning. Educational Studies in Mathematics, 4(1):135–149, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305803.

Williams:1978:CME

[Wil78a] A. Williams. Change in mathematics education since the late 1950’s — ideas and realisation Sierra Leone. Educational Studies in Mathematics, 9(3):297–302, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241032.

Wilson:1978:CME

[Wil78b] B. J. Wilson. Change in mathematics education since the late 1950’s — ideas and realisation West Indies. Educational Studies in Mathematics, 9(3):355–379, August 1978. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00241037.

Wilensky:1997:WNA

[Wil97] Uri Wilensky. What is normal anyway? Therapy for epistemological anxiety. Educational Studies in Mathematics, 33 (2):171–202, July 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1002935313957.

Wilson:2005:SRT

[Wil05] W. Stephen Wilson. Short response to Tunis’s letter to the Editor on technology in college. Educational Studies in Mathematics, 58(3):415–420, March 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
6424-5.

Williams:2009:EMM

[Wil09] Julian Williams. Embodied multi-modal communication from the perspective of activity theory. Educational Studies in Mathematics, 70(2):201–210, March 2009. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9164-y.

Williams:2012:UEV

[Wil12] Julian Williams. Use and exchange value in mathematics education: contemporary CHAT meets Bourdieu’s sociology. Educational Studies in Mathematics, 80(1–2):57–72, May 2012. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9362-x.

Wilkie:2016:SUV

[Wil16a] Karina J. Wilkie. Students’ use of variables and multiple representations in generalizing functional relationships prior to secondary school. Educational Studies in Mathematics, 93(3): 333–361, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-016-9703-x.

Williams:2016:AME

[Wil16b] Julian Williams. Alienation in mathematics education: critique and development of neo-Vygotskian perspectives. Educational Studies in Mathematics, 92(1):59–73, May 2016. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-015-9659-2.pdf.

Wilkie:2021:SQN

[Wil21] Karina J. Wilkie. Seeing quadratics in a new light: secondary mathematics pre-service teachers’ creation of figural growing patterns. Educational Studies in Mathematics, 106(1):91–116, January 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-020-09997-6.

Winslow:2003:SDV

[Win03] C. Winsløw. Semiotic and discursive variables in cas-based didactical engineering. Educational Studies in Mathematics, 52 (3):271–288, April 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1024201714126.

Wing:2011:BRP

[Win11] Tony Wing. Book review: Purifying the dialect of the tribe. Anna Sfard (2008). Thinking as communicating: Human development, the growth of discourses, and mathematizing. Educational Studies in Mathematics, 76(3):363–369, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9283-0.

Wittmann:1969:DSR

[Wit69] Erich Wittmann. The development of self-reliant thinking in mathematics teaching. Educational Studies in Mathematics, 2 (1):123–133, July 1969. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00368989.

Wittmann:1971:CAP

[Wit71] Erich Wittmann. Complementary attitudes in problem solving. Educational Studies in Mathematics, 4(2):243–251, December 1971. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00303389.

Wittmann:1973:CGJa

[Wit73a] Erich Wittmann. The concept of grouping in Jean Piaget’s psychology — formalization and applications. Educational Studies in Mathematics, 5(1):125–146, April 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00684693.

Wittmann:1973:CGJb

[Wit73b] Erich Wittmann. The concept of grouping in Jean Piaget’s psychology — formalization and applications. Educational Studies in Mathematics, 5(2):125–146, June 1973. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01421021.

Wittmann:1975:NNG

[Wit75] Erich Wittmann. Natural numbers and groupings. Educational Studies in Mathematics, 6(1):53–75, March 1975. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00590024.

Wittmann:1976:MAC

[Wit76] Erich Wittmann. Mathematizing around convexity. Educational Studies in Mathematics, 7(1–2):179–184, July 1976. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00144371.

Wittmann:1980:DAA

[Wit80] Erich Wittmann. ‘Discrete analysis’ — an approach to fundamental ideas of analysis. Educational Studies in Mathematics, 11(4):383–392, November 1980. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00231212.

Wittmann:1981:CRI

[Wit81] Erich Wittmann. The complementary roles of intuitive and reflective thinking in mathematics teaching. Educational Studies in Mathematics, 12(3):389–397, August 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00311068.

Wittmann:1982:BR

[Wit82] Erich Wittmann. Book review. Educational Studies in Mathematics, 13(1):113–116, February 1982. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305500.

Wittmann:1984:TUI

[Wit84] Erich Wittmann. Teaching units as the integrating core of mathematics education. Educational Studies in Mathematics, 15(1):25–36, February 1984. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00380437.

Wittmann:1995:MED

[Wit95] Erich Ch. Wittmann. Mathematics education as a ‘design science’. Educational Studies in Mathematics, 29(4):355–374, December 1995. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01273911.

Wittmann:2001:DME

[Wit01] Erich Ch. Wittmann. Developing mathematics education in a systemic process. Educational Studies in Mathematics, 48 (1):1–20, October 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1015538317850.

Witten:2005:DMC

[Wit05] Gareth Q. Witten. Designing a mathematics course for chemistry and geology students. Educational Studies in Mathematics, 58(1):1–19, January 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
2719-2.

Wilkerson-Jerde:2011:HDM

[WJW11] Michelle H. Wilkerson-Jerde and Uri J. Wilensky. How do mathematicians learn math?: resources and acts for constructing and understanding mathematics. Educational Studies in Mathematics, 78(1):21–43, September 2011. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9306-5.

Watson:2004:SVC

[WK04] Jane M. Watson and Ben A. Kelly. Statistical variation in a chance setting: A two-year study. Educational Studies in Mathematics, 57(1):121–144, July 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000047053.96987.5f.

Wubbels:1997:PTR

[WKB97] Theo Wubbels, Fred Korthagen, and Harrie Broekman. Preparing teachers for realistic mathematics education. Educational Studies in Mathematics, 32(1):1–28, January 1997. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1002900522457.

Whitman:1990:SDT

[WL90] Nancy C. Whitman and Morris K. Lai. Similarities and differences in teachers’ beliefs about effective teaching of mathematics: Japan and Hawai’i. Educational Studies in Mathematics, 21(1):71–81, February 1990. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00311016.

Watson:1998:BSI

[WM98] Jane M. Watson and Jonathan B. Moritz. The beginning of statistical inference: Comparing two data sets. Educational Studies in Mathematics, 37(2):145–168, November 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003594832397.

Watson:2001:DRA

[WM01] Jane M. Watson and Jonathan B. Moritz. Development of reasoning associated with pictographs: Representing, interpreting, and predicting. Educational Studies in Mathematics, 48 (1):47–81, October 2001. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1015594414565.

Wheeler:2009:CCS

[WM09] Denna L. Wheeler and Diane Montgomery. Community college students’ views on learning mathematics in terms of their epistemological beliefs: a Q method study. Educational Studies in Mathematics, 72(3):289–306, December 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9192-2.

Wagner:2025:E

[WM25] David Wagner and Vilma Mesa. Editorial. Educational Studies in Mathematics, 118(1):1–4, January 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10384-2.

Winslow:2013:SRC

[WMM13] Carl Winsløw, Yves Matheron, and Alain Mercier. Study and research courses as an epistemological model for didactics. Educational Studies in Mathematics, 83(2):267–284, June 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-012-9453-3.

Wilkins:2022:CSU

[WMN22] Jesse L. M. Wilkins, Beth L. MacDonald, and Anderson Norton. Construction of subitized units is related to the construction of arithmetic units. Educational Studies in Mathematics, 109(1):137–154, January 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10076-7. See correction [WMN25].

Wilkins:2025:CCS

[WMN25] Jesse L. M. Wilkins, Beth L. MacDonald, and Anderson Norton. Correction to: Construction of subitized units is related to the construction of arithmetic units. Educational Studies in Mathematics, 118(2):307, February 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10375-9. See [WMN22].

Weber:2008:LOG

[WMPL08] Keith Weber, Carolyn Maher, Arthur Powell, and Hollylynne Stohl Lee. Learning opportunities from group discussions: warrants become the objects of debate. Educational Studies in Mathematics, 68(3):247–261, July 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9114-8.

Weber:2011:WHM

[WMR11] Keith Weber and Juan Pablo Mejia-Ramos. Why and how mathematicians read proofs: an exploratory study. Educational Studies in Mathematics, 76(3):329–344, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9292-z.

Wilson:2004:KCU

[WN04] W. Stephen Wilson and Daniel Q. Naiman. K–12 calculator usage and college grades. Educational Studies in Mathematics, 56(1):119–122, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000028405.
03127.55.

Wolters:1983:PWS

[Wol83] Miriam A. D. Wolters. The part–whole schema and artithmetical problems. Educational Studies in Mathematics, 14 (2):127–138, May 1983. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00303682.

Wood:1996:ELM

[Woo96] Terry Wood. Events in learning mathematics: Insights from research in classrooms. Educational Studies in Mathematics, 30(1):85–105, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00163754.

Wood:2012:PCC

[Woo12] Leigh N. Wood. Practice and conceptions: communicating mathematics in the workplace. Educational Studies in Mathematics, 79(1):109–125, January 2012. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10649-011-9340-3.pdf.

Wood:2016:RRA

[Woo16] Marcy B. Wood. Rituals and right answers: barriers and supports to autonomous activity. Educational Studies in Mathematics, 91(3):327–348, March 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9653-8.

Wagner:2022:E

[WP22] David Wagner and Susanne Prediger. Editorial. Educational Studies in Mathematics, 109(1):1–4, January 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10138-w.

Wagner:2023:E

[WP23] David Wagner and Susanne Prediger. Editorial. Educational Studies in Mathematics, 112(1):1–2, January 2023. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10208-1.

Wagner:2024:EPC

[WP24] David Wagner and Susanne Prediger. Editorial: paper collections, research retrieval, serendipity. Educational Studies in Mathematics, 116(1):1–3, May 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10328-2.

Wagner:2023:FME

[WPA^{+}23] David Wagner, Susanne Prediger, Michèle Artigue, Angelika Bikner-Ahsbahs, Gail Fitzsimons, Tamsin Meaney, Vilma Mesa, Demetra Pitta-Pantazi, Luis Radford, and Michal Tabach. The field of mathematics education research and its boundaries. Educational Studies in Mathematics, 114(3): 367–369, November 2023. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10270-9.

Winarti:2024:UPS

[WPL24] Destina Wahyu Winarti, Sitti Maesuri Patahuddin, and Tom Lowrie. Unleashing the potential: spatializing middle school mathematics for enhanced learning. Educational Studies in Mathematics, 117(3):419–439, November 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-024-10343-3.

Wittmann:1985:BR

[WQTV85] E. Ch. Wittmann, Douglas Quadling, G. Törner, and J. H. Van Lint. Book reviews. Educational Studies in Mathematics, 16(3):327–334, August 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00776742.

Wang:2018:FIH

[WqWLR18] Ke Wang, Xiao qin Wang, Yeping Li, and Michael S. Rugh. A framework for integrating the history of mathematics into teaching in Shanghai. Educational Studies in Mathematics, 98 (2):135–155, June 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9811-x.

Wenger:1977:POM

[WR77] Robert B. Wenger and Charles R. Rhyner. A problemoriented mathematical optimization course. Educational Studies in Mathematics, 8(3):359–368, October 1977. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00385930.

Wheatley:1996:CAU

[WR96] Grayson H. Wheatley and Anne Reynolds. The construction of abstract units in geometric and numeric settings. Educational Studies in Mathematics, 30(1):67–83, January 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00163753.

Wright:1994:SND

[Wri94] Robert J. Wright. A study of the numerical development of 5-year-olds and 6-year-olds. Educational Studies in Mathematics, 26(1):25–44, January 1994. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273299.

Watson:2008:ULG

[WS08] Anne Watson and Steve Shipman. Using learner generated examples to introduce new concepts. Educational Studies in Mathematics, 69(2):97–109, October 2008. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9142-4.

Wilkie:2018:EIE

[WS18] Karina J. Wilkie and Peter Sullivan. Exploring intrinsic and extrinsic motivational aspects of middle school students’ aspirations for their mathematics learning. Educational Studies in Mathematics, 97(3):235–254, March 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9795-y.

Wawro:2025:WMI

[WS25] Megan Wawro and Kaitlyn Stephens Serbin. “What makes it eigen-esque-ish?”: a form-function analysis of the development of eigentheory concepts in a quantum mechanics course. Educational Studies in Mathematics, 119(2):287–310, June 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10390-4.

Wawro:2011:SLA

[WSR11] Megan Wawro, George F. Sweeney, and Jeffrey M. Rabin. Subspace in linear algebra: investigating students’ concept images and interactions with the formal definition. Educational Studies in Mathematics, 78(1):1–19, September 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9307-4.

Weixler:2019:ELL

[WSU19] Simon Weixler, Daniel Sommerhoff, and Stefan Ufer. The empirical law of large numbers and the hospital problem: systematic investigation of the impact of multiple task and person characteristics. Educational Studies in Mathematics, 100 (1):61–82, January 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/article/10.1007/s10649-018-9856-x.

Weber:2014:SIT

[WT14] Eric Weber and Patrick W. Thompson. Students’ images of two-variable functions and their graphs. Educational Studies in Mathematics, 87(1):67–85, September 2014. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9548-0.

Weber:2022:IRM

[WT22] Keith Weber and Fenner Stanley Tanswell. Instructions and recipes in mathematical proofs. Educational Studies in Mathematics, 111(1):73–87, September 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10156-2.

Wubbels:1992:BR

[Wub92] Theo Wubbels. Book review. Educational Studies in Mathematics, 23(6):619–624, December 1992. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00540064.

Wijaya:2015:OLC

[WvdHPD15] Ariyadi Wijaya, Marja van den Heuvel-Panhuizen, and Michiel Doorman. Opportunity-to-learn context-based tasks provided by mathematics textbooks. Educational Studies in Mathematics, 89(1):41–65, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10649-015-9595-1.
pdf.

Walen:2002:MTE

[WW02] Sharon B. Walen and Steven R. Williams. A matter of time: Emotional responses to timed mathematics tests. Educational Studies in Mathematics, 49(3):361–378, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020258815748.

Williams:2007:BBW

[WW07a] Julian Williams and Geoff Wake. Black boxes in workplace mathematics. Educational Studies in Mathematics, 64(3):317– 343, March 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9039-z.

Williams:2007:MMT

[WW07b] Julian Williams and Geoff Wake. Metaphors and models in translation between college and workplace mathematics. Educational Studies in Mathematics, 64(3):345–371, March 2007. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9040-6.

Weinberg:2011:UMT

[WW11] Aaron Weinberg and Emilie Wiesner. Understanding mathematics textbooks through reader-oriented theory. Educational Studies in Mathematics, 76(1):49–63, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9264-3.

Weinberg:2016:MLN

[WWFC16] Aaron Weinberg, Emilie Wiesner, and Tim Fukawa-Connelly. Mathematics lectures as narratives: insights from network graph methodology. Educational Studies in Mathematics, 91(2):203–226, February 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-015-
9663-6.

Wagner:1986:IFM

[WZ86] Harald Wagner and Bernd Zimmermann. Identification and fostering of mathematically gifted students. Educational Studies in Mathematics, 17(3):243–260, August 1986. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305072.

Wang:2021:HPE

[WZC21] Guangming Wang, Sheng Zhang, and Jinfa Cai. How are parental expectations related to students’ beliefs and their perceived achievement? Educational Studies in Mathematics, 108 (3):429–450, November 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-021-10073-w.

Xenofontos:2020:DCM

[XA20] Constantinos Xenofontos and Paul Andrews. The discursive construction of mathematics teacher self-efficacy. Educational Studies in Mathematics, 105(2):261–283, October 2020. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-09990-z.

Xu:2019:SSC

[XC19] Lihua Xu and David Clarke. Speaking or not speaking as a cultural practice: analysis of mathematics classroom discourse in Shanghai, Seoul, and Melbourne. Educational Studies in Mathematics, 102(1):127–146, September 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09901-x.

Xu:1988:CSS

[XH88] Shudao Xu and Biao Huang. Current status of secondary school mathematics curriculum development in China. Educational Studies in Mathematics, 19(1):93–104, February 1988. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00428387.

Xie:2017:EIB

[XM17] Jinxia Xie and Joanna O. Masingila. Examining interactions between problem posing and problem solving with prospective primary teachers: A case of using fractions. Educational Studies in Mathematics, 96(1):101–118, September 2017. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-017-9760-9.

Xie:2021:MCP

[XXL21] Zhiyong Xie, Leifeng Xiao, and Jian Liu. Micro classes as a primary school-level mathematics education response to COVID-19 pandemic in China: students’ degree of approval and perception of digital equity. Educational Studies in Mathematics, 108(1–2):65–85, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10111-7.

Yahya:2025:DME

[Yah25] Aehsan Haj Yahya. Discussing mathematical events to foster teachers’ noticing of geometric definitions. Educational Studies in Mathematics, 120(3):497–523, November 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-025-10436-7.

Yamskulna:2009:BRB

[Yam09] Gaywalee Yamskulna. Book review: The beautiful Monster. Educational Studies in Mathematics, 70(3):295–298, April 2009. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9179-4.

Yanik:2011:PMS

[Yan11] H. Bahadir Yanik. Prospective middle school mathematics teachers’ preconceptions of geometric translations. Educational Studies in Mathematics, 78(2):231–260, November 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9324-3.

Yang:2012:SCM

[Yan12] Kai-Lin Yang. Structures of cognitive and metacognitive reading strategy use for reading comprehension of geometry proof. Educational Studies in Mathematics, 80(3):307–326, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9350-1.

Yang:2019:DTT

[Yan19] Der-Ching Yang. Development of a three-tier number sense test for fifth-grade students. Educational Studies in Mathematics, 101(3):405–424, July 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/article/10.1007/s10649-018-9874-8.

Yerushalmy:1990:OVO

[YC90] Michal Yerushalmy and Daniel Chazan. Overcoming visual obstacles with the aid of the supposer. Educational Studies in Mathematics, 21(3):199–219, June 1990. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305090.

Yang:1995:CCI

[YC95] Ma Tzu-Lin Yang and Paul Cobb. A cross-cultural investigation into the development of place-value concepts of children in Taiwan and the United States. Educational Studies in Mathematics, 28(1):1–33, January 1995. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01273854.

Yopp:2016:WDA

[YE16] David A. Yopp and Rob Ely. When does an argument use a generic example? Educational Studies in Mathematics, 91 (1):37–53, January 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-015-9633-z.

Yerushalmy:2000:PSS

[Yer00] Michal Yerushalmy. Problem solving strategies and mathematical resources: A longitudinal view on problem solving in a function based approach to algebra. Educational Studies in Mathematics, 43(2):125–147, September 2000. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017566031373.

Yerushalmy:1999:SCR

[YG99] Michal Yerushalmy and Shoshana Gilead. Structures of constant rate word problems: A functional approach analysis. Educational Studies in Mathematics, 39(1–3):185–203, June 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003793926216.

Yang:2015:EEP

[YHL^{+}15] Kai-Lin Yang, Hui-Yu Hsu, Fou-Lai Lin, Jian-Cheng Chen, and Ying-Hao Cheng. Exploring the educative power of an experienced mathematics teacher educator-researcher. Educational Studies in Mathematics, 89(1):19–39, May 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9589-4.

Yim:2010:CSD

[Yim10] Jaehoon Yim. Children’s strategies for division by fractions in the context of the area of a rectangle. Educational Studies in Mathematics, 73(2):105–120, March 2010. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-009-9206-0.

Yang:2020:BRP

[YK20] Xinrong Yang and Gabriele Kaiser. Book review: Professional development of mathematics teachers- an Asian perspective. Berinderjeet Kaur, Oh Nam Kwon, Yew Hoong Leong (Eds.) (2017). Educational Studies in Mathematics, 103(2):241–249, February 2020. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
article/10.1007/s10649-019-09923-5.

Yolcu:2024:SJM

[YK24] Ayşe Yolcu and Kathryn L. Kirchgasler. Social (justice) mathematics: racializing effects of ordering pedagogies and their inherited regimes of truth. Educational Studies in Mathematics, 116(3):351–370, July 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10289-y.

Yang:2008:MRC

[YL08] Kai-Lin Yang and Fou-Lai Lin. A model of reading comprehension of geometry proof. Educational Studies in Mathematics, 67(1):59–76, January 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-007-
9080-6.

Yi:2024:CMT

[YLH24] Gyuhee Yi, Jihyun Lee, and Jihyun Hwang. Cultivating mindfulness through conditional tasks in mathematics classrooms. Educational Studies in Mathematics, 117(3):379–398, November 2024. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-024-10323-7.

Yau:2016:ISP

[YM16] King Woon Yau and Ida Ah Chee Mok. Insights from students’ private work in their notebooks: how do students learn from the teacher’s examples? Educational Studies in Mathematics, 93(3):275–292, November 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9702-y.

Yoon:2025:TOF

[Yoo25] Caroline Yoon. Treasures from our field: sharing mathematics education research through story and essay. Educational Studies in Mathematics, 120(2):205–223, October 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-025-10418-9.

Young:1968:CC

[You68] Gail Young. The computer and the calculus. Educational Studies in Mathematics, 1(1–2):105–110, May 1968. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00426237.

Yerushalmy:2012:SAG

[YS12] Michal Yerushalmy and Osama Swidan. Signifying the accumulation graph in a dynamic and multi-representation environment. Educational Studies in Mathematics, 80(3):287– 306, July 2012. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-011-9356-8.

Yang:2022:PSM

[YSL22] Xinrong Yang, Björn Schwarz, and Issic K. C. Leung. Preservice mathematics teachers’ professional modeling competencies: a comparative study between Germany, Mainland China, and Hong Kong. Educational Studies in Mathematics, 109(2):409–429, February 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://

link.springer.com/article/10.1007/s10649-021-10064-
x.

Yackel:2003:DCW

[YSRU03] Erna Yackel, Michelle Stephan, Chris Rasmussen, and Diana Underwood. Didactising: Continuing the work of Leen Streefland. Educational Studies in Mathematics, 54(1):101– 126, November 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AEDUC.0000005213.
85018.34.

Yang:2022:TPS

[YST22] Der Ching Yang, Iwan Andi Jonri Sianturi, and Natthapoj Vincent Trakulphadetkrai. Taiwanese primary school teachers’ perceived enablers for and barriers to the integration of children’s literature in mathematics teaching and learning. Educational Studies in Mathematics, 110(1):125–148, May 2022. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10115-3.

Yoon:2011:GBM

[YTD11] Caroline Yoon, Michael O. J. Thomas, and Tommy Dreyfus. Grounded blends and mathematical gesture spaces: developing mathematical understandings via gestures. Educational Studies in Mathematics, 78(3):371–393, December 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-011-9329-y.

Zahner:2015:RRC

[Zah15] William Zahner. The rise and run of a computational understanding of slope in a conceptually focused bilingual algebra class. Educational Studies in Mathematics, 88(1):19–41, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-014-9575-x.

Zassenhaus:1971:GDC

[Zas71] Hans Zassenhaus. Genetic development of the congruence axioms. Educational Studies in Mathematics, 4(1):150–152, June 1971. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00305804.

Zaslavsky:2005:SOC

[Zas05] Orit Zaslavsky. Seizing the opportunity to create uncertainty in learning mathematics. Educational Studies in Mathematics, 60(3):297–321, November 2005. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-005-
0606-5.

Zavaleta:2021:ECB

[Zav21] Luis Miguel Maraví Zavaleta. Emergent curriculum in basic education for the new normality in Peru: orientations proposed from mathematics education. Educational Studies in Mathematics, 108(1–2):291–305, October 2021. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-021-10100-w.

Zazkis:1998:OEO

[Zaz98] Rina Zazkis. Odds and ends of odds and evens: An inquiry into students’ understanding of even and odd numbers. Educational Studies in Mathematics, 36(1):73–89, June 1998. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003149901409.

Zazkis:1999:IRN

[Zaz99] Rina Zazkis. Intuitive rules in number theory: Example of ‘the more of a, the more of b’ rule implementation. Educational Studies in Mathematics, 40(2):197–209, October 1999. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1003711828805.

Zhao:2019:BRW

[ZB19] Dongchen Zhao and Christian Bokhove. Book review: What is the most important source of teachers’ knowledge? Lianghuo Fan (2014) Investigating the pedagogy of mathematics: how do teachers develop their knowledge? Educational Studies in Mathematics, 102(2):289–298, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09912-8.

Zehavi:1985:QEM

[ZBBZ85] Nurit Zehavi, Maxim Bruckheimer, and Ruth Ben-Zvi. Qualitative evaluation of mathematical activity and its relation to effective guidance. Educational Studies in Mathematics, 16(1): 27–40, February 1985. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00354881.

Zan:2006:AME

[ZBEH06] Rosetta Zan, Laurinda Brown, Jeff Evans, and Markku S. Hannula. Affect in mathematics education: An introduction. Educational Studies in Mathematics, 63(2):113–121, October 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-006-9028-2.

Zbiek:2006:BME

[ZC06] Rose Mary Zbiek and AnnaMarie Conner. Beyond motivation: Exploring mathematical modeling as a context for deepening students’ understandings of curricular mathematics. Educational Studies in Mathematics, 63(1):89–112, September 2006. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9002-4.

Zazkis:2008:WMC

[ZC08] Rina Zazkis and Egan J. Chernoff. What makes a counterexample exemplary? Educational Studies in Mathematics, 68 (3):195–208, July 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-007-9110-4.

Zhu:2019:EHN

[ZC19] Jinxin Zhu and Ming Ming Chiu. Early home numeracy activities and later mathematics achievement: early numeracy, interest, and self-efficacy as mediators. Educational Studies in Mathematics, 102(2):173–191, October 2019. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-019-09906-6.

Zhuang:2022:TUR

[ZC22] Yuling Zhuang and AnnaMarie Conner. Teachers’ use of rational questioning strategies to promote student participation in collective argumentation. Educational Studies in Mathematics, 111(2):345–365, October 2022. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-022-10160-6.

Zachariades:2013:RSA

[ZCPP13] Theodossios Zachariades, Constantinos Christou, and Demetra Pitta-Pantazi. Reflective, systemic and analytic thinking in real numbers. Educational Studies in Mathematics, 82(1): 5–22, January 2013. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-012-9413-y.

Zengin:2018:ECD

[Zen18] Yılmaz Zengin. Examination of the constructed dynamic bridge between the concepts of differential and derivative with the integration of GeoGebra and the ACODESA method. Educational Studies in Mathematics, 99(3):311–333, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9832-5.

Zepp:1981:RBM

[Zep81] Raymond A. Zepp. Relationships between mathematics achievement and various English language proficiencies. Educational Studies in Mathematics, 12(1):59–70, February 1981. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00386046.

Zepp:1982:BUL

[Zep82] Raymond A. Zepp. Bilinguals’ understanding of logical connectives in English and Sesotho. Educational Studies in Mathematics, 13(2):205–221, May 1982. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00460711.

Zandieh:2017:CUN

[ZER17] Michelle Zandieh, Jessica Ellis, and Chris Rasmussen. A characterization of a unified notion of mathematical function: the case of high school function and linear transformation. Educational Studies in Mathematics, 95(1):21–38, May 2017. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic).

Zevenbergen:1996:CLB

[Zev96] Robyn Zevenbergen. Constructivism as a liberal bourgeois discourse. Educational Studies in Mathematics, 31(1–2):95– 113, September 1996. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00143928.

Zevenbergen:2004:TNI

[Zev04] Robyn Zevenbergen. Technologizing numeracy: Intergenerational differences in working mathematically in new times.

Educational Studies in Mathematics, 56(1):97–117, May 2004. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AEDUC.0000028399.76056.91.

Zevenbergen:2006:SPC

[Zev06] Robyn Zevenbergen. Search of a pedagogy of conflict and dialogue for mathematics education. Educational Studies in Mathematics, 62(1):101–104, May 2006. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-005-9010-4.

Zevenbergen:2011:YWT

[Zev11] Robyn Jorgensen Zevenbergen. Young workers and their dispositions towards mathematics: tensions of a mathematical habitus in the retail industry. Educational Studies in Mathematics, 76(1):87–100, January 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s10649-010-
9267-0.

Zack:2001:MMM

[ZG01] Vicki Zack and Barbara Graves. Making mathematical meaning through dialogue: “Once you think of it, the Z minus three seems pretty weird”. Educational Studies in Mathematics, 46(1–3):229–271, March 2001. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1014045408753.

Zhu:2025:DDP

[ZHT25] Jiawen Zhu, Xingfeng Huang, and Luc Trouche. The development of didactic praxeologies in lesson design through social resource interactions: evidence from two novice mathematics teachers. Educational Studies in Mathematics, 119(3):487– 513, July 2025. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-025-10403-2.

Zazkis:2015:FDI

[ZK15] Rina Zazkis and Boris Koichu. A fictional dialogue on infinitude of primes: introducing virtual duoethnography. Educational Studies in Mathematics, 88(2):163–181, February 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9580-0.

Ziols:2021:HPB

[ZK21] Ryan Ziols and Kathryn L. Kirchgasler. Health and pathology: a brief history of the biopolitics of US mathematics education. Educational Studies in Mathematics, 108(1–2):123–142, October 2021. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL https://link.springer.com/
article/10.1007/s10649-021-10110-8.

Ziols:2024:SER

[ZK24] Ryan Ziols and Kathryn L. Kirchgasler. Slow emergencies of racism in mathematics education. Educational Studies in Mathematics, 116(3):371–389, July 2024. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10293-2.

Zhu:2018:GEM

[ZKC18] Yan Zhu, Gabriele Kaiser, and Jinfa Cai. Gender equity in mathematical achievement: the case of China. Educational Studies in Mathematics, 99(3):245–260, November 2018. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/article/10.
1007/s10649-018-9846-z.

Zazkis:1991:CCT

[ZL91] Rina Zazkis and Uri Leron. Capturing congruence with a turtle. Educational Studies in Mathematics, 22(3):285–295, June 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00368342.

Zazkis:2002:GPT

[ZL02] Rina Zazkis and Peter Liljedahk. Generalization of patterns: the tension between algebraic thinking and algebraic notation. Educational Studies in Mathematics, 49(3):379–402, March 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020291317178.

Zazkis:2008:EDC

[ZL08] Rina Zazkis and Roza Leikin. Exemplifying definitions: a case of a square. Educational Studies in Mathematics, 69(2): 131–148, October 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s10649-008-9131-7.

Ziqiang:1991:MEC

[ZM91] Zhu Ziqiang and Eula Ewing Monroe. Mathematics education in China today: Four problem areas. Educational Studies in Mathematics, 22(2):205–208, April 1991. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00555723.

Zwanch:2025:AEM

[ZM25] Karen Zwanch and Brooke Mullins. The (activity-)effect of manipulatives on algebraic generalizations: a constructivist teaching experiment. Educational Studies in Mathematics, 119(1):41–61, May 2025. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-024-10371-z.

Zepp:1987:CLE

[ZML87] R. Zepp, J. Monin, and C. L. Lei. Common logical errors in English and Chinese. Educational Studies in Mathematics, 18 (1):1–17, February 1987. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00367911.

Zaki:1991:DRS

[ZP91] Moncef Zaki and François Pluvinage. Démarches de résolution et de simulation face au problème de la ruine d’un joueur. (French) [Solving and simulating the problem of the ruin of a player]. Educational Studies in Mathematics, 22(2):149–181, April 1991. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00555721.

Zazkis:2018:OOC

[ZR18] Rina Zazkis and Annette Rouleau. Order of operations: On convention and met-before acronyms. Educational Studies in Mathematics, 97(2):143–162, February 2018. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10649-017-9789-9.

Zolkower:2007:TMT

[ZS07] Betina Zolkower and Sam Shreyar. A teacher’s mediation of a thinking-aloud discussion in a 6th grade mathematics classroom. Educational Studies in Mathematics, 65(2):177– 202, June 2007. CODEN EDSMAN. ISSN 0013-1954 (print),

1573-0816 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s10649-006-9046-0.

Zagorianakos:2015:RIP

[ZS15] Andonis Zagorianakos and Anna Shvarts. The role of intuition in the process of objectification of mathematical phenomena from a Husserlian perspective: a case study. Educational Studies in Mathematics, 88(1):137–157, January 2015. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-014-9576-9.

Zaslavsky:2002:BSA

[ZSL02] Orit Zaslavsky, Hagit Sela, and Uri Leron. Being sloppy about slope: The effect of changing the scale. Educational Studies in Mathematics, 49(1):119–140, January 2002. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1016093305002.

Zhang:2025:UCP

[ZSWC25] Ling Zhang, Naiqing Song, Guowei Wu, and Jinfa Cai. Understanding the cognitive processes of mathematical problem posing: evidence from eye movements. Educational Studies in Mathematics, 118(3):449–478, March 2025. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10649-023-10262-9.

Zurina:2011:GO

[ZW11] Harun Zurina and Julian Williams. Gesturing for oneself. Educational Studies in Mathematics, 77(2–3):175–188, July 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9294-x.

Zwanch:2021:RCS

[ZW21] Karen Zwanch and Jesse L. M. Wilkins. Releasing the conceptual spring to construct multiplicative reasoning. Educational Studies in Mathematics, 106(1):151–170, January 2021. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL https://link.springer.com/article/
10.1007/s10649-020-09999-4.

Zhang:2024:DEP

[ZWC^{+}24] Yuhuan Zhang, Xiaohui Wu, Shuang Chen, Chengcheng Cui, Yahan He, and Lidong Wang. Differential effects of private tutoring on groups of students’ mathematics achievements: a longitudinal study. Educational Studies in Mathematics, 115 (2):197–222, February 2024. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL https://link.

springer.com/article/10.1007/s10649-023-10282-5.

Zazkis:2016:BGB

[ZWMR16] Dov Zazkis, Keith Weber, and Juan Pablo Mejía-Ramos. Bridging the gap between graphical arguments and verbalsymbolic proofs in a real analysis context. Educational Studies in Mathematics, 93(2):155–173, October 2016. CO-DEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9698-3.

Zodik:2008:CTC

[ZZ08] Iris Zodik and Orit Zaslavsky. Characteristics of teachers’ choice of examples in and for the mathematics classroom. Educational Studies in Mathematics, 69(2):165–182, October 2008. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-008-9140-6.

Zazkis:2011:SMK

[ZZ11] Rina Zazkis and Dov Zazkis. The significance of mathematical knowledge in teaching elementary methods courses: perspectives of mathematics teacher educators. Educational Studies in Mathematics, 76(3):247–263, April 2011. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 REFERENCES 622

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-010-9268-z.